A Bibliography of Papers in Lecture Notes in Computer Science (1994)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

11 March 2017 Version 1.27

Title word cross-reference

(1+, λ) [3101]. (s, t) [2700]. 0 [2716]. 1 [2716]. 16 [2245]. 2 [3492, 874, 633, 1887, 1891, 1116, 1872, 872]. 2k − 1 [1202]. 3 [1074, 3282, 1048, 2075, 2480, 1883, 1097, 3284, 1887, 1891, 2760, 1116, 1059, 1885, 1870, 1088, 1203, 1872, 1119, 1082, 3377, 1037, 1091, 1108, 1121, 876]. 8 [539]. ^{k} [339]. 2 [2217]. D [1202, 2162]. f [2291, 2037]. h [2070]. k [1202, 1733, 354, 2593, 1974, 2053, 1848, 306, 2312, 2039, 2077, 2162, 1865]. k/n [573]. L [3560]. L_{p} [374]. λ_{Δ} [754]. m [106]. NC^{k}(NP) = AC^{k−1}(NP) [341]. μ [3317]. N [3482, 2207, 821, 1720, 898, 2593, 2586, 62]. o [167]. O(log log N) [1201, 1861]. O(n) [3543]. p [2640]. π [2136, 2294, 2138, 1568, 2137, 1585]. pi/4 [554]. q [346]. r [2631]. t [1507]. v [2070]. w [1175]. Z [1227].

- [1202, 2070, 2716, 2162]. -adic [2640]. -ary [2593, 2207]. -Best [2077]. -Body [3482, 898]. -Calculus [2294, 3317, 1568, 2136, 2138, 2137, 754, 1585]. -categorical [633]. -Clustering [354]. -Coloring [2037]. -Cube [2586]. -cubes [2593]. -D [3492, 1048, 2480, 1887, 1116, 1059, 1091, 62]. -Dimensional [539, 1203, 3377, 1097]. -DQPSK [554]. -Edge-Connect [2039]. -Grammars [346]. -Interpretation [2168]. -invariants [1507]. -Label [306]. -Languages [1175]. -Matching [2291].

-Metric [374]. -MKRP [1476]. -Nodes [167]. -Overlap [2312]. -Pairwise [2053]. -Paths [2700]. -Reductions [3560]. -Round [2245]. -Sequences [106]. -sets [1974]. -Strategies [3101]. -Structures [1733]. -Style [1227]. -Trees [1848, 1865].

/ [185].

0/1 [3225]. 011 [1773]. 011-AC [1773]. 011-DC [1773].

1 [182, 879, 982, 3504, 2822]. 1-k [335]. 1-Reader [2822]. 1-Writer [2822]. 1.2 [2031]. 120 [231]. 15 [1009].

2 [950, 2656, 1356, 2406, 1360, 3027, 1294, 2355, 3253, 3517, 1606, 2398, 2713, 2617]. 2-Adic [1294]. 2-class [3446]. 2-d [2739]. 2-Optimal [2713]. 2-Structures [1356, 1360]. 200 [3518, 893]. 2DT [650].

3 [485].

5 [3489, 118, 148, 2602, 965, 3518, 893, 910]. 5/4 [2694]. 5/4-Approximation [2694]. 5/ CM [3484, 3521]. 5E [3484, 3521].

6 [205]. 64 [1274]. 64-bit [1297]. 65 [572]. 65-MHz [572].

802.12 [2397].

90 [998, 1004]. 90D [152]. 90D/HPF [152]. 93 [1485].

= [1379].

A-Codes [92]. Abduction [2190, 3299, 1928]. Abductive [1304, 1782, 2498, 1925]. abelian [3449]. Ability [1592]. AbleProfessionals [2933]. Abnormality [2959]. Abstract [1182, 3046, 655, 385, 387, 1780, 729, 3043, 2297, 3053, 1741, 1742, 3360, 654, 1162, 2016, 3065, 988, 1407, 503, 743, 3337, 3430, 758, 1233, 739, 2010, 627, 1453, 1458, 722, 638, 724, 2148]. Abstracted [1156]. Abstracting [780, 2353, 1831, 2351, 1745, 2329]. Abstraction [728, 3169, 1164, 1330, 1304, 1653, 2104, 486, 483, 3551, 3326, 503, 2890, 2320]. Abstractions [2795, 1633, 516, 1342, 1623, 2143]. abstractor [701]. AC [2374, 1455, 2372, 499, 1773]. AC-superposition [1455]. AC-Unification [2374, 2372]. AC-unifiers [1455]. Academic [32, 23]. Accelerated [2652]. Accelerating [838]. Acceleration [164]. Acceptance [359, 360, 3430]. Accepted [2284]. Access [568, 2926, 2525, 564, 3418, 1678, 2652, 490, 3415, 532, 2940, 549, 2905, 2649, 542, 1903]. Accesses [149]. Accessibility [2917]. Accessible [2924, 2917]. Accessing [2936]. Accommodation [2933]. Accounting [2842]. Accumulators [116]. Accuracy [846, 3528]. Accurate [1307, 1879, 3139]. Achieved [1000]. Achievement [1791]. Achieving [2091, 2572]. Acquisition [3163, 597, 2986, 3159, 3158, 3165]. Across [670, 2348]. Act [1333, 1331]. Action [3330, 1959, 3087, 1910, 747, 178, 2956, 1994, 3357, 644, 643, 2359, 2010, 3257, 2140, 3263, 3242]. Action-Based [643]. Actions [2938, 1402, 2886, 3233]. Activation [3018, 1580]. Activation-Oriented [3018]. Active [1680, 2768, 1035, 2298, 2852, 2769, 2851, 2770, 2771, 1577, 1054, 1037, 1057, 1144, 1108]. Activities [1663]. Activity [202]. Actor [1597, 394]. Actor-Based [1597]. ACTRESS [644]. Acyclic [2221, 651]. AD [3273]. Ada [9]. Adaptable [2165]. Adaptation [3188, 2166, 3080, 84, 3187, 3114, 3115, 2165, 3193]. Adaptations [2511]. Adapted [3429, 1116]. Adapting [978, 2969, 2897, 815, 548].

Adaptive [340, 1573, 148, 2594, 1653, 2611, 2593, 2006, 2554, 2178, 2392, 2609, 2592, 1563, 3085, 3142, 532, 3176, 546, 3118, 1127]. Adaptively [3141, 254, 253]. Add [1006]. Add-On [1006]. Added [2531]. Adder [2571]. Adding [2202, 3366, 1767, 1552, 525, 660, 3379]. Additional [3294, 348, 3296]. Additive [2269]. Address [1, 2600]. Adequacy [1829, 1151, 628]. adic [2640, 1294]. adjunctions [635]. Adjusting [2064]. Adjustment [74]. Admissible [1409]. Admission [2398]. Admits [2017]. Admitting [2639]. ADMS [460]. Advance [3179]. Advanced [779, 2515, 993, 2526, 2513, 564, 3102, 2811, 28, 2423, 57, 1374, 872, 803]. Advances [416, 2901, 1043]. Advantages [512, 1057]. Advection [3481]. Advection-Chemistry [3481]. Adventure [17]. Adverbials [1922]. Advice [2766, 1479]. ADVISE [2629]. AE [1424]. AEA [907]. Aeronautical [860]. Aerospace [937, 935, 936]. AF2 [1996]. Affine [1126, 1715, 1867, 1881, 1872, 1049, 1117, 1080, 1116, 1067, 1084, 1069, 1047, 1037]. Affinity [157, 1587]. AFORIZM [3168]. Aft [6]. After [1968]. Against [1869, 253, 805, 254]. Agent [2791, 3159, 1956, 3439, 1957, 2100, 1593, 1271, 1268, 1965, 3232, 3240]. Agent-Based [3159]. Agents [2522, 1964, 191, 3145, 1960, 2976, 2954, 1927]. AGG [403]. Aggregate [1662]. Aggregating [283]. Aggregation [1330, 3423, 3387]. Aggregations [1958]. Agley [6]. Agreement [2832, 257]. Agricultural [2629]. AI [2415, 2896, 1315]. aid [3229]. Aided [57, 66, 58, 87, 90, 3234]. Aids [2902]. Air [3481, 3491, 2225, 454, 1660]. Aircraft [855, 2804, 936]. Airline [2716]. AKL [1012]. Al [2937, 703]. ALF [1226]. Algebra [1181, 1982, 1944, 1873, 3369, 845, 479, 498, 1528, 3551, 2736, 3521, 1192, 698, 2146, 2142, 1458, 624]. Algebra-Based [845]. Algebraic [3059, 1941, 1898, 3486, 2490, 1725, 519, 632, 315, 1526, 1891, 686, 2002, 2275, 2861, 1191, 1590, 3549, 1872, 72, 58, 403, 2761, 2762, 1199, 727, 1442, 3038, 1170, 3531, 626, 3241, 622, 634, 3455, 639, 801, 640, 641, 3469]. Algebras [2313, 793, 1378, 2130, 633, 638, 2132]. ALGOL [2320]. ALGOL-like [2320]. Algorithm [229, 951, 3306, 2385, 2092, 3186, 1837, 1234, 1839, 1947, 3308, 3079, 2463, 504, 2990, 2714, 2703, 2702, 3150, 676, 960, 1842, 3129, 2717, 2087, 2063, 2644, 2654, 2973, 1728, 3075, 3501, 2689, 2439, 2031, 3147, 298, 1609, 3317, 606, 919, 587, 1649, 3151, 2042, 1644, 920, 355, 1274, 687, 2050, 893, 790, 1538, 1582, 3543, 1281, 3109, 2713, 1861, 2854, 309, 2701, 2633, 2831, 3328, 217, 3107, 2835, 2377, 1208, 839, 2716, 596, 1289, 3124, 2062, 3099, 458, 2699, 1865, 3448, 3451, 3461, 3248, 1139, 2117, 824, 3462, 1070, 1110, 3456, 3464, 802, 2214]. Algorithmic [1710, 78, 2885, 1374, 2687, 2040]. Algorithms [2601, 1835, 3558, 3104, 3136, 1237, 3548, 581, 3106, 1573, 2698, 3092, 3137, 2064, 2875, 3080, 354, 2593, 3103, 1845, 2662, 1591, 2626, 1847, 3500, 3542, 2690, 2688, 2689, 1198, 2712, 1848, 3078, 1704, 3508, 3133, 1852, 3083, 1257, 2634, 2027, 3082, 161, 2045, 892, 2033, 1200, 1615, 353, 1242, 3146, 446, 2026, 1855, 2312, 2603, 2643, 2676, 3091, 250, 3450, 3521, 1858, 2888, 3072, 3557, 3081, 331, 2995, 417, 3564, 2978, 2727, 1592, 2683, 3144, 3073, 2721, 1377, 1969, 2067, 3524, 882, 2169, 1219, 567, 3093, 2999, 2086, 2084, 3289, 1078, 1093, 825, 3225, 1509]. algorithms [3477, 1077]. Aliasing [2569, 2506]. Align [159]. Alignment [1237, 1888, 1234, 1235, 154, 1241, 1240, 1135]. Alignment-Distribution [154]. Alignments [1236]. all-digital [803]. All/ Nothing [113]. Allcache [1575]. Allegories [1387]. Allocation [513, 1759, 2828, 529, 2544, 986, 1607, 1963, 1588]. Allows [1243]. Almost [98]. ALOHA [822, 566]. ALOHA-based [822]. Alpha [497, 1230]. Alpha-Conversion [497, 1230]. Alphabet [1244]. Alphabet-Independent [1244]. Alphabetic [1708]. Alphabets [1250]. Altera [2437, 2468]. Altera-PLDs [2468]. Alternate [133, 134]. Alternating [136, 1712]. Alternation [2309, 1404, 323]. Alternative [116, 1240, 41, 482, 2561, 21]. Alternatives [447]. Always [327]. Amalgamated [403]. ambiguities [1072]. Ambiguity [371, 2854, 3223]. Ambiguous [3321]. America [2921]. among [3546, 2709]. Amorphous [894]. Amount [2091]. AMPHION [3244]. Amplifiers [555, 546]. Amplitude [544]. analog [1070]. Analogical [2417, 603, 2151, 1312, 2152]. Analogies [3300]. Analogue [95]. Analogues [3513, 3076]. Analogy [2167, 3300]. Analyser [2098]. Analyses [2869]. Analysing [2064, 1613]. Analysis [380, 2280, 2494, 2574, 2696, 1326, 3059, 2283, 843, 421, 2759, 2847, 1761, 840, 655, 3335, 912, 1790, 3069, 1985, 514, 296, 149, 562, 666, 3102, 2501, 1211, 3054, 174, 173, 171, 3066, 3044, 36, 3393, 3067, 1732, 1596, 1721, 2178, 3064, 3063, 2674, 2582, 1695, 84, 586, 1334, 3055, 1242, 833, 88, 2675, 2568, 2850, 2764, 687, 2616, 3383, 1760, 2844, 501, 3048, 2592, 166, 3403, 2275, 1672, 690, 540, 143, 2697, 3395, 172, 2636, 827, 668, 940, 1702, 67, 2354, 3464, 3068, 694, 795, 1773, 656, 2671, 739, 90, 3289, 2029, 1491, 1078, 1495]. analysis [1498, 1499, 3461, 1430, 1489, 2119, 711, 713, 715, 821, 1504, 718, 1090, 1147, 1110, 3252, 1088, 1510, 1512, 1082, 1516, 1485, 820]. Analytic [1561, 375, 588, 1467]. Analytical [794, 1132]. Analyzing [3106, 850, 3376]. Ancestors [1860]. And-Parallel [1569]. And-parallelism [3061]. Angular [2690]. animation [640]. Animations [2872].

Annealing

[1608, 3137, 2682, 2428, 3135, 3134, 898]. Annotated [2340]. Annotation [1933]. Annotation-Based [1933]. Anomalies [1762, 2825]. Anomalous [2959]. Anonymous [113]. Answer [305, 1026]. Answering [1805, 3230, 1438]. Ant [1956]. Antenna [534]. Anti [1788]. Anti-Links [1788]. Anticipatory [3087]. Antiprenexing [1775]. APIs [2525]. apparent [1039]. Appendix [2264]. Applicability [2512, 581, 1430]. Application [2993, 280, 368, 1669, 2378, 2653, 957, 2177, 2537, 2988, 1671, 3205, 1673, 2529, 1956, 2001, 1867, 2352, 269, 1687, 1807, 1393, 3095, 2804, 2054, 1663, 2735, 3512, 3187, 975, 3146, 1620, 1795, 1269, 1995, 1768, 990, 2983, 2716, 933, 2194, 3001, 2028, 1428, 1489, 3216, 1040]. Applications [3181, 3396, 3586, 3195, 570, 2668, 2759, 1911, 994, 1658, 3196, 1670, 993, 2669, 958, 889, 2395, 921, 3559, 2630, 2235, 2555, 1689, 1935, 2588, 3272, 2797, 3334, 2516, 3585, 2018, 965, 832, 576, 86, 394, 3398, 572, 265, 1745, 1322, 2394, 1880, 987, 3380, 2254, 2267, 3179, 969, 3590, 1347, 2743, 1009, 1119, 2375, 532, 973, 902, 2787, 757, 1917, 2009, 3281, 2014, 3172, 1659, 816, 1046, 3475]. Applied [2425, 2340, 3504, 892, 3519, 1407, 987, 936, 62, 3229, 1077]. Applying [206, 3145, 1596, 3348, 208, 1090, 3173, 3088, 1672, 2794]. Approach [1381, 445, 2997, 2574, 3168, 2530, 1386, 2163, 3487, 1628, 2208, 1944, 2847, 2353, 152, 3138, 3170, 3017, 2550, 1165, 46, 3121, 1832, 1272, 2829, 2863, 3349, 385, 3424, 564, 1275, 2289, 2957, 1932, 2754, 473, 600, 2555, 1929, 3498, 1831, 2421, 2772, 1807, 901, 845, 1821, 12, 2043, 174, 3020, 730, 2853, 2725, 2178, 1922, 1333, 905, 89, 2150, 1668, 2760, 1963, 2884, 1265, 609, 373, 2243, 3350, 2864, 59, 1345, 267, 2312, 177, 1822, 2396, 66, 2861, 1768, 448, 591, 188, 665, 3120, 990, 248, 2535, 3304, 2727, 1613, 2628, 1550, 517, 972, 1340]. Approach [1776, 1560, 21, 1743, 87, 2637, 3228, 3230, 1066, 3264, 2129, 1506, 1426, 1895, 3256, 3224]. Approaches [3339, 2944, 3158, 867, 2850, 2937, 2154, 913, 2393, 2260]. Approaching [2694]. Appropriate [2751]. Approximability [358]. Approximable [1988]. Approximate [1254, 1256, 1261, 3324, 2041, 3219, 3224]. Approximately [1239, 3323]. Approximating [2023]. Approximation [2060, 419, 3558, 1237, 2702, 2490, 3559, 1841, 1853, 3560, 2725, 1398, 1703, 2061, 2049, 2312, 3557, 418, 2694]. Approximations [2270, 3056, 1850, 306, 2216, 2322, 2715, 3463, 3262]. AQDT [3253]. AQDT-2 [3253]. Arbitrary [3101, 2027, 3546, 142]. Arbitration [2591, 246, 2228]. Arc [2078, 2077, 1498, 1506]. Arches [6]. Architectural [520, 2531, 772, 1008, 2393, 2535]. Architecture [575, 2171, 182, 1959, 563, 2791, 2523, 4, 3416, 2401, 1561, 2669, 1016, 960, 1575, 2473, 2529, 844, 3435, 1684, 180, 462, 1656, 2445, 455, 1604, 2509, 2627, 1603, 506, 3575, 2657, 1602, 2444, 464, 613, 2551, 2427, 2481, 2950, 1006, 517, 3524, 3525, 3570, 3422, 2572, 3280, 3233, 3235]. Architectures [1748, 991, 2537, 989, 3499, 984, 2682, 2661, 953, 3510, 932, 2520, 3146, 1614, 1537, 3089, 3142, 969, 278, 882, 1601, 924, 2893, 2586, 1496]. Archiving [925, 3209, 1676, 3570]. Arctic [2612]. Area [2724, 573, 2391, 1129]. Area-Efficient [573]. ARGS [3432]. Argument [1995]. Arguments [288, 786]. Aristotelian [2093]. Arithmetic [2431, 2450, 3568, 2640, 2346, 2239, 795, 804, 1437]. Arity [1404, 1989]. Arjuna [295]. Arm [3085, 1130]. Arnoldi [957]. Array [664, 2673, 2579, 2632, 2480, 3054, 663, 1603, 146, 169, 3395, 172, 1006, 2633, 1560]. Arrays [2236, 149, 2425, 1363, 2441, 2052, 1207, 2581, 2459, 2652, 534, 2478]. Arrival [831]. Arrowhead [2681]. Art [2929, 803, 743]. Artefact [3426]. articulated [1094, 1040]. Artifacts [2774]. Artificial [2003, 2899, 1964, 3131, 2486, 1961, 2942]. ARTMAP [3132]. ary [2207, 2593, 540]. ASIC [2460]. ASICs [2453]. Asides [2151]. Asking [1313]. Aspect [1922, 1122]. Aspects [1941, 830, 1185, 2533, 2896, 1991, 1369, 1003, 2846, 1029, 1872]. Assembly [975]. assertion [1468]. Assertions [1744, 797]. Assessing [842]. Assessment [2920, 3173, 41, 18, 907]. Assigning [2084]. Assignment [1562, 168, 2399, 2408, 737]. Assignments [2003, 744]. Assimilation [902]. Assistance [2172]. Assistant [3429, 3037]. Assisted [1985, 269, 3004]. Assistive [2904, 2923, 2913]. Association [1107, 2921]. Associative [1454, 3230]. Associative-Commutative [1454]. Assumption [1786, 3268, 3470, 2134]. assumption-based [3268]. assumption-commitment [2134]. Assumptions [224, 1228, 223]. Assurance [2541, 3422]. astronomical [1439]. Asymmetric [1174]. Asymptotically [1851, 2676]. Asynchronous [2624, 1698, 2834, 212, 2026, 1699, 549, 698, 1450, 812]. Asynchronously [2285]. Asynchrony [3035]. Atlanta [2933]. ATLAS [1526]. ATM [575, 834, 3586, 3189, 3188, 2829, 3184, 3194, 836, 3177, 2406, 3202, 3193, 820]. Atmospheric [901]. ATMSWSIM [836]. Atomic [2496, 1747, 2822]. Atomicity [1685]. Attack [2240, 1943, 2245, 234, 2246, 805]. Attacks [127, 2218, 255, 1292, 254, 253]. Attaining [254, 253]. Attempt [2917]. Attention [3307, 1035]. Attitudes [1315, 2101]. Attribute [2419, 645, 646, 1813, 3228]. attribute-oriented [3228]. Attributes [586]. Auctions [1372]. Audio [3013, 2914, 562]. Auditory [2916].

Augmentation [2039]. Authentication [237, 2226, 247, 3407, 3406, 246, 2227, 2228, 3410, 234, 137, 93]. Author [138, 259, 1769, 2251]. Authoring [3441]. Auto [823]. Auto-regressive [823]. Autoepistemic [1303, 2186, 1306, 1309, 2209, 2201, 3268]. Autogenerating [3432]. Automata [1381, 1617, 2092, 420, 1628, 1583, 878, 1691, 422, 328, 3129, 2380, 1358, 1694, 310, 73, 3555, 1698, 2057, 2308, 685, 1619, 2310, 325, 1625, 1699, 3341, 3556, 1377, 757, 3118, 3039, 2126, 2149]. Automata-Driven [3556]. Automata-Theoretic [1628]. Automated [2198, 597, 917, 1264, 2197, 1641, 1329, 2968, 1938, 1479, 1416, 2365, 1428, 3265, 1480, 1485]. Automath [2870]. Automatic [1826, 493, 1636, 144, 1622, 151, 1488, 1002, 959, 3496, 74, 2626, 896, 2658, 2018, 298, 2985, 866, 833, 169, 1654, 3365, 924, 3244]. Automatically [847, 1661, 1421]. Automaticity [366]. Automating [68, 2890]. Automation [2871, 2863, 71, 1290, 3254]. Automaton [72]. Autonomous [3145, 2976, 2940, 1064]. AUTONOMY [2904]. Availability [153]. Available [280, 305]. Avalanche [100]. Average [2836, 2032, 1986, 2031, 1973, 1736, 1702, 3289]. Average-case [1702]. Averaging [590]. Avoid [1943, 120]. Avoidance [1013, 1027, 3583]. Awareness [21]. AWGN [806]. Axiomatic [1265, 1408]. Axiomatisation [1173]. Axiomatisations [749]. Axiomatising [1158]. Axiomatization [361]. Axiomatizing [1914]. Axioms [1304, 1023].

B [1491, 1837, 2341, 3152, 3351, 1121]. B-spline [1121]. B-Trees [1837]. B-W [1491]. Babel [2361]. Back [1921, 648]. Background [262]. Backmarking [3373]. Backtrack [3558, 426, 1986]. Backtrack-Free [426]. Backtracking [1181, 3391]. backward [1491]. Baden [928]. Baden-Wurttemberg [928]. Balanced [1589, 218]. Balancing [3490, 971, 2686, 982, 1963, 2685, 1612, 1556, 1613, 458]. Baldwin [3096]. Ballistic [2726]. BAN [112, 133, 134]. BAN-logic [134, 133]. Bandwidth [2399, 2410, 511, 2413, 1588]. Banking [2547]. Banyan [2584]. Barcelona [828]. Bargaining [1372]. Base [2955, 534, 2599, 928, 2191, 1948]. Base-Station [534]. Based [2170, 3047, 1942, 2997, 2574, 2171, 2153, 2385, 322, 2455, 2163, 421, 2653, 2242, 1302, 2219, 96, 241, 2569, 2164, 2177, 3170, 598, 147, 2698, 2166, 655, 2842, 2470, 966, 3129, 562, 2222, 95, 1597, 3159, 1932, 2456, 425, 3158, 1578, 2618, 1831, 830, 1339, 2784, 845, 3298, 1285, 2939, 2006, 2907, 2474, 2521, 1820, 2469, 997, 2682, 2428, 36, 3393, 3284, 3067, 3147, 249, 3397, 2183, 2853, 2357, 2178, 2472, 3300, 274, 3406, 1333, 1661, 2883, 2804, 2179, 2621, 2462, 161, 2151, 2485, 684, 3301, 2150, 86, 2227, 1599, 2782, 3344, 609, 3444, 1242, 2172, 396]. Based [833, 1610, 1889, 1933, 2764, 2432, 59, 3383, 2160, 2937, 2185, 2452, 2444, 1563, 3441, 2168, 643, 66, 3002, 3114, 63, 448, 658, 2174, 248, 3395, 523, 3304, 2687, 1804, 3427, 2978, 443, 401, 2158, 214, 1335, 2180, 58, 694, 2181, 215, 1810, 1566, 203, 1340, 1676, 3570, 3422, 2400, 316, 2561, 2169, 2802, 3165, 2011, 2617, 91, 2182, 2155, 2162, 2454, 2420, 76, 62, 2010, 85, 3390, 314, 1194, 1428, 1464, 1477, 626, 1494, 3227, 3228, 2143, 3229, 763, 2118, 3230, 1097, 1503, 1066, 3122, 3066, 1094, 3266, 3250, 3268, 1504, 3027, 1668, 1070, 1059, 1089, 172]. based [1484, 822, 817, 159, 1884, 2449, 978, 989, 914, 2434, 1825, 882, 2650]. Bases [2807, 1336, 3161, 442, 462, 2860, 413, 3463, 3251, 1438]. Basic [2001, 998, 751, 670, 2253, 2102]. Basis [957, 173, 3468, 637]. Batch [1591]. Bayesian [3248, 1891, 2989, 3240]. BBN [1007]. BCH [810]. BCI [2918]. BDDs [3265, 1451]. Be [1000, 1918, 1742, 1741, 2946, 111, 2950]. Beacon [3406]. Beam [912]. Bearings [3519]. Beats [2328]. Beginning [34]. Behavior [1959, 1677, 1575, 2878, 3097, 3048, 3160, 1550]. Behavioral [2849, 1802, 1342, 449, 1796]. Behaviors [420, 1173, 700, 2149]. Behaviour [2639, 3087, 2821, 3130, 198, 471, 3341, 1161, 3346, 1340, 2129, 1511, 1894, 642]. Behavioural [2492, 501, 701, 634]. Behaviours [2370, 2161, 2145]. behind [1096]. Belief [2192, 2190, 382, 2418, 607, 1926]. Beliefs [1306, 2209, 3270]. belongs [3470]. below [2287]. Benchmark [951, 950]. Benchmarking [1899]. Benchmarks [2458]. Bendix [330]. Benefits [2910, 3043]. Benes [986]. Benes-Type [986]. Bent [99, 1295]. BER [557]. BERKOM [3197]. Berlekamp [802]. Best [1234, 2077]. BETA [781]. Better [2691]. Between [1179, 1772, 2654, 281, 1221, 23, 2307, 1189, 3089, 2161, 1325, 1693, 436, 3241, 3268, 92, 2781, 1425, 3564, 1050]. Beyond [3378, 2852, 883, 1634, 2527, 3160, 407, 1740, 3264, 1095]. Bi [452, 3154, 2985, 636]. Bi-Gram [2985]. Bi-Partitioning [3154]. bi-rewriting [636]. Bias [3150, 2998]. Biases [620, 619]. Biconnected [423, 2831]. Bidiagonal [3529]. Bidirectional [2973]. Bidomains [1716]. Big [2289]. Bijective [124]. bilattices [3270]. Binary [2707, 1543, 2017, 2286, 370, 1838, 318, 2881, 427, 2081, 2305, 2311, 2070, 3564, 2080, 1864, 2368, 1435, 3464, 817]. BIND [2800]. Binding [522, 3575, 2806, 3231, 713]. binding-time [713]. binocular [1112, 1077]. BinWAM [2330]. Biochemical [3509]. Biological [3076, 890, 2182]. Biology [889]. Biomedical [2800]. Biomolecular [1678]. Bipartite [2076]. Birational [255, 214]. Birkhoff [1014]. Birthday [1351]. Bisection [2307]. Bisimilarity [363, 2117]. Bisimulation [3022, 1156, 1370, 2141, 1153, 362, 2138, 2137]. BIST [2437, 2568, 2454]. BIST-PST [2437]. Bistructure [2279]. Bistructures [1716]. Bit [129, 238, 1967, 2232, 558, 1602, 571, 464, 548, 494, 239, 1297]. Bit-Sequential [1602]. Bit-Serial [571]. Bit-Sliced [464]. Bit-Vector [1967]. Bits [2047]. BLAS [860]. Blind [2914, 106, 2896, 2927, 2929, 2898]. Blinde [2930]. Block [2244, 1275, 1298, 249, 1276, 1292, 536, 1274, 125, 248, 2427, 1297, 876, 817, 812]. Block-Ciphering [1274]. Blocking [1551, 2619]. Blocks [2665, 3141, 670, 873]. Blowfish [1297]. Board [2485, 3027, 2660, 1006]. Body [3482, 3517, 898]. Book [1966]. Bookreading [2934]. Books [2935]. Boolean [3059, 421, 2341, 100, 2435, 2051, 1510, 301, 218, 1712, 2586]. Boolean-valued [301]. bordered [1516]. BOSS [3434]. BOSS-System [3434]. Both [2317, 3384]. Bottlenecks [2404]. Bottom [1437, 272, 715, 585, 1785, 440, 2335, 1473]. Bottom-Up [585, 2335, 1437, 272, 715, 1785, 440, 1473]. Bound [2724, 1733, 1844, 2292, 2644, 678, 1945, 2228, 1245, 547, 1738]. Boundary [2218, 1856, 819, 1101]. Bounded [2056, 2693, 2283, 2710, 1709, 388, 2222, 1846, 321, 1989, 1850, 350, 1507]. Bounded-arity [1989]. Bounded-Degree [1850]. Bounding [123]. Bounds [1381, 1692, 2284, 2236, 2698, 1842, 370, 3535, 1847, 1735, 2690, 2302, 1259, 2722, 1688, 2261, 432, 3549, 849, 2028]. box [2131, 1505]. Boxes [1160, 2247]. Boxings [3051]. BPA [2116]. BPA-systems [2116]. Brain [1677, 2918]. Brain-Computer [2918]. Branch [2644]. Branching [1628, 1900, 1923, 1918, 3567]. Branching-Time [1628, 1918]. Breaking [117, 2238, 2705]. Breakpoints [2838]. Bridge [2562]. bridges [3241]. Bridging [3588, 2781, 441, 25]. brightness [1116]. Bringing [1211]. British [3333]. Broadband [3592, 2550, 3205, 2399, 3572]. Broadcast [129, 258, 1852, 1214]. Broadcasting [562, 2681, 1213, 2680, 725]. Broader [476]. Broker [1684]. Browsing [1978]. BSP [1605]. BST [1702]. Buchberger [2385]. Buchi [2057]. Buechi [1699]. Buffer [2608, 3191]. Bugs [2619]. Build [852, 1432]. Building [3141, 3241, 2521, 2183, 33, 944, 1322, 592, 295, 2787, 1421]. Built [337]. Built-Ins [337]. Bulk [1289]. Bundle [58]. Burst [1946, 558]. Bus [2680, 1578, 1212, 1552, 2477, 70, 3361]. Bus-Based [1578]. Business [1326, 210]. Butterfly [1007]. Buy [1920, 2689]. Bypass [1552]. Byte [1274]. Byte-Oriented [1274].

C [2536, 543, 2529, 2000, 2742, 663, 2855, 2477]. C-TORI [2742]. C/TDMA [543]. CA [1773]. Cable [3181]. Cache [1564, 1561, 1633, 2651, 669, 3048, 1563, 29, 2650]. Cached [273]. Caches [1562]. Cachesim [29]. Caching [460, 932]. CAD [2378, 864, 2469, 2442, 481, 66, 2950, 2753]. CAD-Plans [2950]. CAFE [3412]. Calculable [785]. Calculate [849]. Calculating [3471]. Calculi [1781, 1725, 679, 411, 1221, 1393, 1023, 1994, 1191, 1997, 748, 3541]. Calculus [2281, 3293, 3050, 753, 322, 3012, 1386, 1939, 1583, 2622, 1783, 1772, 1395, 2294, 1393, 272, 2381, 2841, 1397, 3317, 2965, 744, 1262, 3409, 2371, 1995, 1407, 2195, 2888, 1585, 789, 3037, 1645, 2979, 1568, 741, 1925, 1651, 3040, 677, 1980, 702, 2136, 729, 1483, 1460, 2131, 719, 2138, 696, 2135, 2137, 754]. calibration [1128, 1133, 1130, 1144]. Call [3050, 688, 1492]. Call-by-Name [688]. Call-by-Value [688, 3050]. Camera [917, 1128, 1144, 1881, 1133, 1054]. Cameraman [2018]. cameras [1145]. Can [1432, 224, 1000, 1918, 279, 111, 2314, 732, 852, 223, 1136]. Canadian [2895]. Cancellation [541, 540]. Cancer [1680, 610]. Candidate [2256]. Canonic [1148]. Canonical [389, 3124]. Canons [2105]. cap [1773]. Capabilities [288, 2733, 1812, 1414, 2572]. Capability [55]. Capacity [538, 531, 574]. CAPTools [966]. Capture [567, 566]. Capturing [2272, 1331, 1749, 1328]. car [1076]. Card [1686]. Cardinality [2941, 322, 1498]. Care [1254, 2548, 2177, 3196]. Carlo [908, 2470, 879, 2654]. Carnap [2186]. Carrying [497]. Cartan [1867]. Cascades [3141]. Case [2170, 2153, 3010, 3127, 2234, 2163, 3426, 2164, 2177, 2159, 2166, 2594, 3229, 2546, 38, 3362, 3298, 2945, 2939, 2183, 2178, 1736, 274, 2179, 2151, 2150, 1225, 2549, 1528, 2160, 2184, 2185, 2168, 2174, 503, 2262, 3304, 3337, 1590, 2950, 985, 295, 2180, 2165, 2181, 2802, 1415, 1558, 2182, 2155, 2162, 2010, 3289, 1869, 3230, 3242, 1702, 10, 1334, 1320, 2540, 87]. Case-Based [2170, 2153, 2163, 2164, 2177, 2166, 3298, 2183, 2178, 274, 2179, 2151, 2150, 2160, 2185, 2168, 2174, 3304, 2180, 2181, 2182, 2155, 2162, 2010, 3229, 3230]. Case-Deliverer [2184]. Case-retrieval [2950]. Cases [2156, 2373, 2184, 2173, 3327, 2165]. Cash [243, 424]. CAST [71, 67]. CAST-Tool [67]. Catalog [943]. Catalyst [2732]. Catch [788]. Categorical [1177, 759, 633]. categorically [2125]. Categories [1163, 635]. Category [1176, 1348]. Cauchy [177]. Causal [1205, 3214, 607, 588, 3171, 1143, 3238]. Causality [2316, 1518]. cause [1416]. CBR [576, 2154]. CC [1218, 2496, 3384, 3383]. CCB [27]. CCB-Directed [27]. CCS [492]. CDMA [538, 533, 562, 541, 534, 531, 540, 574, 561]. CDMA-Based [562]. Cell [575, 3098, 2473, 550, 3192, 574]. Celloids [3131]. Cellular [543, 533, 422, 3129, 2425, 1363, 2308, 528, 1699, 534, 531, 560, 561]. Center [2927, 2928, 945, 1862]. Centered [182, 3170, 2735, 1321, 212]. Centers [941, 3213]. Central [2703]. Centralized [2479]. Centres [2920]. CERFACS [3495]. CERN [911]. Certain [1710, 569, 3471, 3466]. certification [630]. Certified [1229]. CFB [235]. CFD [1011]. CG [954]. Chain [2475, 2062, 2084]. Chain-Like [2084]. Chaining [2244, 1447]. Challange [934]. Challenge [3505, 2576, 909]. Challenges [199, 261, 3444]. Chance [178]. Chang [132]. Chang-Wu-Chen [132]. Change [199, 3405, 2085, 2956, 82, 2191, 3263]. Changes [2876, 2685, 1098]. Changing [330]. Channel [529, 555, 551, 569, 2595, 536, 541, 807, 552, 3409, 537, 113, 2481, 565, 809, 810, 820]. Channels [1713, 556, 538, 535, 557, 547, 1951, 1497, 806, 813, 811]. Chaos [2242, 823]. Chaotic [3095, 3063]. Characteristic [582, 1257, 2050, 3453]. Characteristics [943, 2470, 558, 124, 251, 218]. Characterizable [294]. Characterization [1972, 2983]. Characterizations [292, 339, 2148]. Characterizing [701, 581, 2138]. Characters [1254]. Charging [2546]. Charts [2970]. Cheap [2689]. Cheaper [2839]. Cheating [101]. Chebyshev [1574]. Check [217]. Checked [736, 1432]. Checking [3045, 1382, 2837, 1643, 1386, 1628, 1947, 1594, 1650, 1653, 679, 1641, 3385, 1223, 1630, 3360, 480, 2792, 2745, 2575, 1634, 1620, 1632, 1647, 1645, 1219, 1917, 1712, 3472, 2121, 706, 2122]. Checkpointing [2573, 2583, 276]. Checks [1635]. Chemical [884, 1182, 1162]. Chemistry [3481, 905, 883, 870]. Chen [132]. Chernoff [2040]. Chervonenkis [304]. Chinese [2922]. Chip [2612, 2224, 2482, 3151, 572, 571, 804].

Chip-Matched-Filter [572]. Chips

[486, 2277]. Choice [1696, 3077, 2961, 2146]. Choices [775]. Chordal [2703, 2643]. Chords [1843]. CHORUS [289, 2395]. Chosen [2245, 254, 253]. Chromosome [1259]. CIAO [3381]. CIE [1089]. CIM [2810, 2812, 87]. CIO [3198, 3591]. CIPF [591]. Cipher [2244, 1277, 1275, 1298, 1276, 126, 125, 1297]. Ciphering [1274]. Ciphers [1280, 105, 1286, 249, 136, 1299, 248, 1284, 217]. Ciphertext [254, 2246, 253]. Circuit [2878, 835]. Circuits [2569, 1387, 2269, 2880, 3147, 2013, 2467, 1200, 3537, 2568, 2432, 1654, 301, 2127]. Circular [2078, 2077]. Circular-Arc [2077]. Circulation [901]. Circumscribing [1912, 3263]. circumscription [1436]. Circumscriptions [2188]. Circumscriptive [2960]. Citizens [1015]. clarification [3218]. Class [1617, 2867, 2724, 1015, 476, 2985, 1342, 1625, 2603, 790, 518, 2335, 2145, 3446, 1424, 3471, 716, 722, 639, 3470, 834]. Classes [3290, 99, 2269, 3297, 1394, 1848, 360, 2272, 290, 1348, 2598, 517, 352, 2037, 710, 2849]. Classic [2953]. Classical [753, 2199, 3261, 321, 730, 3271]. Classification [581, 3311, 2939, 3300, 1889, 1348, 176, 2155, 3248, 824, 1052, 3223]. Classified [582]. Classifier [3121, 3087, 3123, 615, 3074]. Classifiers [606]. Classifying [943, 1378]. Classroom [27, 8]. Clausal [2367]. Clause [2363, 2212, 318, 2347, 1993, 1435]. Clauses [1016, 1993, 621, 1233, 1427]. Cleaning [1723]. Cleanup [1971]. Client [305, 2843, 114, 291]. Client-Independent [114]. Client-Server [305, 291]. Climbing [3125, 3134]. Cliques [1866, 376]. Clock [1288, 3021, 558, 571, 1857, 2091]. Clocks [2839, 2128, 698]. Closed [1930, 1786]. Closing [509, 617]. Closure [3306, 768, 1722, 1230]. Closures [1777].

Clothing [2753]. clp [2341, 2343, 1017, 2351, 3383, 3379, 1776, 2379]. CLP-Language [1776]. CLP-Scheme [1017]. Cluster [3502, 2053, 875, 861, 897, 3526, 947, 76]. Cluster-Based [76]. Clustered [3106]. Clustering [1826, 2997, 1680, 3293, 2415, 2470, 354, 1827, 2738, 3031, 611, 1105]. Clusters [3480, 3482, 3489, 934, 284, 969, 973, 1595, 314]. CM [3484, 3489, 148, 3449, 2602, 965, 3518, 893, 1606, 3450, 3521, 910]. CM-2 [1606]. CM-200 [3518, 893]. CM-5 [3484, 3489, 148, 2602, 965, 3518, 893, 3521, 910]. CM-5/CM-5E [3484, 3521]. CM-fields [3450]. CML [3156]. CMOS [3147, 571, 2562]. Co [3086, 3297, 59, 3100, 1152, 1426]. Co-Design [59]. Co-evolutionary [3100]. Co-Evolving [3086]. Co-Induction [1152]. Co-Learnability [3297]. Co-Operative [3086]. Coalgebra [1192]. Coalgebras [1398]. Coarse [2663, 2087]. Coarse-Grained [2087]. Coarsest [1649]. Code [1011, 3482, 2665, 144, 651, 662, 3507, 1334, 2228, 665, 1547, 2889, 532, 3351, 549, 863, 2321, 872]. Coded [3183, 1951, 816, 809, 825, 806, 815, 820]. Codes [1942, 245, 1955, 1947, 966, 1950, 1946, 247, 1737, 246, 92, 372, 536, 573, 537, 165, 234, 1949, 1948, 810, 814, 817, 812]. Codesign [2472]. Coding [3094, 344, 533, 798, 807, 803]. CODIT [563]. Coevolutionary [3083, 3120]. Cognitive [3439]. Coherence [3443, 2651, 1563]. Coherency [1600]. Coherent [1564, 1610, 2961]. Coherent-Cache [1564]. Cohesive [8]. cohomology [3460]. Coins [242, 121]. Collaboration [3592, 22]. Collaborative [3440, 37, 1321, 212, 935]. Collapses [2021]. Collecting [3327]. Collection [3420]. Collective [1300, 3150]. Collector [2331, 2479, 1577]. Collectors [2332]. Collins [1443]. Collision [2712, 569, 887, 1068]. Collisions [118, 120, 119]. Colonies [1956, 1365]. Color [3192, 1664]. Colored [311]. Coloring [1562, 1975, 3544, 2067, 2037, 1865]. colors [1089]. Coloured [1497, 1185, 686, 849, 1493, 1506]. Colouring [2314]. Combination [1434, 331, 1443]. Combinations [1888, 1574, 3316, 307, 738]. Combinatorial [1842, 2296, 2015, 2299, 686, 3557]. Combinatorics [2080]. Combinators [1982, 766]. Combinatory [1527, 1530]. Combine [3169, 1097]. Combined [2574, 3498, 1066]. Combining [2350, 1521, 1470, 2341, 3367, 1725, 2949, 2349, 730, 1678, 2850, 1760, 1647, 1907, 1517]. Comes [2963, 1968]. Comfortable [2934]. Comfortable-Digital [2934]. Command [3432]. Commercial [859, 921, 2689, 10, 2852, 3398, 886, 8]. Commitment [238, 2232, 239, 2134]. Common [2060, 1248, 2971, 370, 1249, 1338, 1703, 302, 1311, 2276, 1860, 3465, 1134]. Common-Sense [2971, 1311]. CommonKADS [3156]. Commonsense [1810]. Communicating [1583, 2285, 1157, 2954, 3553]. Communication [1381, 2514, 2673, 775, 3188, 2843, 2526, 2607, 2292, 3212, 925, 2402, 2619, 156, 153, 2661, 219, 2517, 965, 1333, 2533, 2590, 558, 2485, 2576, 2042, 1534, 858, 3199, 2273, 949, 2676, 1535, 1537, 1688, 3198, 3591, 2817, 110, 1588, 3206, 3579, 848, 2903, 565, 2617, 3584, 3400, 956, 1497]. Communication-Automata [1381]. Communication/Computation [2673]. Communications [529, 2519, 2550, 1355, 946, 1555, 526, 1212, 1541, 2895, 3580, 2527, 1033, 2671, 542]. Community [43, 929]. Commutative [3310, 3071, 343, 1454, 3537]. Compact [3161, 1392, 2444]. Compactness [2187]. Companies [2813]. Company [2813]. Comparative [3332, 3133, 1078].

Comparator [2290]. Comparing [3339, 2644, 1397, 1362, 1238, 1413, 1530]. Comparison [2653, 232, 3491, 2698, 211, 2407, 2292, 2944, 2654, 3158, 3078, 2308, 1812, 2994, 588, 2154, 3089, 2995, 968, 554, 2552, 882, 3011, 1139, 1050]. Comparison-Based [2698]. Comparisons [1109, 799, 1043]. Compatible [2249]. compensating [1098]. Competing [3115]. Competitive [2280, 2064, 219]. Competitiveness [941]. Compilation [152, 995, 708, 3020, 1527, 1538, 1537]. Compilative [675]. Compile [3402, 156, 669, 165, 2671, 3057]. Compile-Time [3402, 156, 165, 3057]. Compile-Time/Run-Time [165]. Compiler [661, 155, 664, 2268, 269, 648, 2358, 2334, 663, 644, 643, 659]. Compiler-Assisted [269]. Compilers [152, 1000]. Compiling [1536, 2354, 652]. Complement [2066]. Complementation [1699]. Complete [2056, 1543, 2286, 2270, 3385, 347, 3569, 361, 2195, 1173, 936, 761, 595, 2145, 2058]. Completeness [1924, 2491, 2313, 3054, 1396, 3563, 349, 1916, 1640, 1486, 635, 1710]. Completion [2385, 2964, 372, 330, 1031, 2498, 1464]. completion-based [1464]. Complex [2465, 2759, 2431, 1794, 605, 410, 573, 88, 2955, 2771, 1793, 61, 67, 461, 404, 1495, 1119]. Complexity [1381, 1979, 2836, 3290, 3094, 421, 2284, 2083, 379, 1235, 2569, 2730, 364, 3536, 1388, 2269, 3534, 2291, 2292, 2644, 1390, 408, 2457, 3554, 2296, 1694, 104, 321, 1697, 1198, 2726, 2031, 3123, 2299, 360, 1362, 1215, 2068, 1363, 2704, 2069, 1736, 2272, 2243, 2311, 430, 324, 409, 1314, 66, 1688, 301, 2318, 3549, 2830, 366, 616, 304, 303, 1199, 3479, 1456, 806, 3454, 3465, 811]. Component [3571, 3332, 521, 2878, 1609, 2520]. Components [3162, 423, 9, 3574, 1373, 3427, 295, 2831, 2048]. Composable [3570].

Composed [1638]. Composing

[2665, 1639]. Composite [1667, 251, 437]. Composite-Object [437]. Composition [780, 2344, 147, 1747, 392, 2392, 2149, 2121, 2132, 1439]. Compositional [2494, 1643, 1160, 1652, 2614, 1922, 3033, 2275, 2888, 1611, 1516, 2114]. Compositionality [490]. Compositions [1799, 2122]. Comprehensions [3054, 1896]. Comprehensive [19, 1504]. Compressed [1705, 1243]. Compression [2696, 1955, 2569, 118, 2440, 2066, 3192, 1242, 1243, 2330, 119]. Computation [2639, 3483, 1636, 2664, 2035, 3129, 83, 240, 1735, 2205, 2358, 170, 3067, 2641, 2634, 1678, 2248, 1603, 3537, 2058, 3077, 77, 3108, 3076, 1246, 1786, 1917, 3062, 1898, 1470, 1490]. Computation/Communication [2664]. Computational [3480, 2560, 966, 2546, 224, 3058, 1206, 2189, 2644, 390, 78, 871, 895, 1362, 1399, 3574, 3552, 1151, 2267, 1690, 3276, 2073, 223, 1087]. Computationally [913]. Computations [880, 2559, 2838, 144, 2723, 3497, 174, 1724, 392, 2322, 1554, 511, 1544, 1753, 1590, 2727, 1863, 2648, 1193, 3527, 3528, 3472, 3446, 3458, 2133, 2139, 3465]. Compute [2708, 1871, 1861]. Computed [678]. Computer [834, 3413, 1873, 828, 2919, 2914, 2654, 2811, 38, 1539, 984, 479, 2918, 2897, 2909, 3027, 1174, 736, 3438, 2423, 1004, 2660, 2921, 1612, 39, 2484, 6, 57, 66, 29, 34, 945, 887, 5, 58, 1752, 890, 2766, 2929, 3004, 2084, 87, 90, 1496, 1111, 3234, 3477]. Computer-Aided [57, 66, 87, 90, 3234]. Computer-Checked [736]. Computer-Supported [3438, 1752]. Computerization [57]. Computers [2574, 943, 878, 948, 3491, 2268, 2607, 1578, 3503, 506, 275, 3519, 969, 1006, 904, 3524, 1726, 877, 956]. Computing [1572, 1833, 1834, 979, 2446, 1301, 3485, 829, 282, 2910, 1873, 3536, 1236, 3494, 2470, 3459, 167, 3474, 3449, 946, 856, 900, 3502, 2021, 1020, 927, 942, 1257, 1678, 944, 2576, 2685, 883, 974, 2238, 3515, 2036, 2050, 1729, 945, 3269, 928, 2578, 1862, 317, 616, 304, 972, 3155, 2831, 3526, 776, 3469, 510, 933, 745, 2818, 3530, 1065]. CONCAT [2764]. Concatenated [533, 1950, 814]. Concatenating [1616]. Concatenation [245]. Concatenations [3319]. Conceiving [913]. Concentration [1551]. Concept [1300, 2208, 1336, 1305, 582, 2421, 605, 3324, 3341, 523, 614, 3255, 2915, 314, 815]. Concepts [2941, 2536, 190, 2104, 197, 2773, 2852, 1310, 1829, 82, 3325, 2551, 2253, 2255, 594, 139, 2321, 777, 631]. Conceptual [2112, 1330, 2110, 2107, 1814, 2109, 2760, 2100, 2103, 2097, 1341, 3156, 2094, 2102, 413, 2095, 1800, 1904]. Concerning [1363, 863, 3239]. Conclusion [2325, 2488]. Concrete [2654, 3430]. Concurrency [445, 1383, 775, 1984, 1201, 1719, 2860, 1399, 685, 525, 3357, 1189, 778, 1192, 3034, 1586, 2141, 748, 2134]. Concurrent [2281, 2624, 884, 765, 2938, 147, 2656, 2496, 515, 1183, 1700, 729, 3352, 1935, 31, 3381, 766, 516, 3348, 1187, 2660, 1745, 770, 1930, 397, 3384, 3383, 2370, 2324, 2487, 2369, 2580, 1746, 2497, 745, 2329, 468, 1032, 1194, 1488, 2119, 2133, 2123]. Condition [2882]. Conditional [1784, 2384, 2957, 2500, 3032]. Conditionals [1316, 3257, 3266]. Conditions [2481, 1718]. Conference [3590]. Conferencing [3195, 3587, 3204]. Configurable [3175]. Configuration [298, 2581, 2900, 2562]. Configurations [1768]. Confirmer [2220]. Conflict [2822, 3373, 3438, 1514]. Conflict-free [2822]. Conflicts [2938, 2976, 2892]. Confluence [2500, 689, 1407, 1415]. Confluent [2371]. Conforming [2599]. Congestion [2602, 3583, 3584]. Congestion-Free [2602]. Congruence [666, 361, 1173, 362, 2138, 2144]. Congruences [3051]. Congruity [2019]. Conjectures [1774]. Conjugate [955].

Connect [2039]. Connected

[2682, 1609, 815]. Connecting [3442, 1856]. Connection [2827, 2195, 1425, 3484, 3485, 3489, 965, 3518, 893, 1606, 3521]. Connectionist [2997, 2422, 139]. Connectionist-Symbolic [2997]. Connections [1552, 1207]. Connectivity [1840, 1728, 2069, 652]. Connotation [2764]. Conquer [926]. Conscious [513]. Consensus [1235, 2821, 2834, 2817]. Consensusless [2435]. Consequences [2513, 135]. Conservative [3497, 1436]. Conservativeness [789]. Conservativity [1221]. Consideration [3419, 87]. Considerations [64, 1278]. Considering [1327]. Consistency [1388, 2379, 2189, 3374, 1120, 2745, 1425]. Consistent [3313, 2049, 2583, 1030, 303, 1325]. Constant [667, 1210, 2695, 2678, 332, 3108, 2705, 2722, 1460]. Constant-Space [2705]. Constant-Time [1210]. constants [3478]. Constrained [2203, 2667, 3555, 2196, 2715, 1094, 1471]. Constraint [3047, 2385, 3388, 2378, 319, 1840, 2759, 675, 889, 2341, 1783, 2342, 426, 2340, 2358, 3385, 3381, 3375, 3083, 654, 3373, 733, 3316, 3392, 3398, 2792, 2376, 397, 2452, 2324, 2487, 690, 3394, 2856, 2806, 3100, 3395, 3380, 3372, 3382, 2375, 2369, 3389, 2377, 2329, 3390, 699, 706, 707, 819, 723, 2123, 2382]. Constraint-Based [3047, 2385, 2452, 3395, 3390]. Constraint-Generating [3388]. Constraint-Programming [2856]. Constraints [3386, 1979, 3396, 2350, 1797, 1723, 2387, 328, 672, 2380, 2389, 201, 1937, 3374, 3393, 3371, 2390, 684, 1991, 2386, 2085, 1756, 2675, 1581, 1870, 2371, 2370, 3379, 490, 3387, 2429, 3376, 592, 2365, 3377, 1437, 1097, 3233, 1455, 1044, 1454]. Constructing [3494, 3473, 2789, 2311, 3326]. Construction [2693, 1837, 957, 3046, 2645, 2076, 2822, 168, 3006, 246, 2305, 1793, 2257, 3004, 1479, 3256]. Constructions [1995, 107, 629]. Constructive [800, 675, 2274, 3302, 591, 2947, 1458]. Constructor [689]. Constructor-Sharing [689]. Constructs [2614, 1696, 734]. Consumer [2009]. Containment [2998, 1648]. Content [431, 444]. Contention [568, 2820, 847, 848]. Contents [455, 1664]. Context [3195, 579, 1626, 2990, 2005, 1357, 2998, 1021, 682, 1721, 1238, 1366, 2764, 2728, 18, 1258, 3329, 2999, 2120, 2117, 2122]. Context-Free [1626, 2990, 1357, 2998, 2728, 1258, 3329, 1366, 2120, 2117, 2122]. Contexts [2106, 2104, 2105, 2772, 3258]. Contextual [2109, 797, 397, 2959, 1061]. Continuation [1179, 1587, 723]. Continuations [2337, 722]. Continuing [2897]. Continuous [2395, 3374, 1529, 2413, 3142, 2464]. contour [1039, 1142, 1041]. contours [1113]. Contraction [165]. Contractions [2975]. Contradicting [3372]. Contradiction [1013, 1266, 1027]. contradictory [3236]. contrapositives [1422]. Contribution [3169]. Control [445, 559, 580, 3013, 598, 2902, 1622, 1594, 3092, 917, 2345, 1154, 1772, 2001, 855, 2911, 3362, 3086, 2860, 914, 649, 3509, 71, 2462, 3191, 2913, 1600, 3097, 2337, 2767, 1758, 193, 778, 3139, 2398, 560, 2354, 3415, 2400, 2903, 2009, 3584, 739, 62, 90, 561, 1042, 1081, 1489, 1035, 1503, 1451, 1057, 803, 3224]. Control-Flow [739]. Control-Parallel [1594]. Controllable [2218]. Controlled [380, 1288, 1364, 3423, 3036, 3231]. Controller [2006, 2469, 2012]. Controllers [2556]. Controlling [2769, 591, 3119, 3116, 312]. Controls [3424]. Conventional [3372]. Convergence [3104, 772, 3111, 3107]. Convergency [2281, 2029]. convergent [1093]. converging [1087]. Conversational [3440]. Conversion [497, 1230]. Convex [1839, 3548, 2710, 2087, 1392, 3056, 3516, 3549]. Convexity [1210]. Convolution [1528]. Convolutional [1950, 536, 573, 812]. Cook [349, 1349]. Cooperate [2769]. Cooperating [1319, 1354, 2946, 1373]. Cooperation [3574, 3203]. Cooperative [979, 3205, 1656, 1323, 191, 2742, 3438, 937, 3120, 1782, 2743, 1513, 1752]. Coordinate [2484]. Coordinated [198, 1963, 1965]. Coordination [775]. Coping [2820]. Coplanar [1890]. Copying [383, 2331]. Coq [1780, 1225, 1229]. CORDIC [2484]. Core [1695]. Core-ML [1695]. COREA [1583]. Coreferent [2105]. Cores [2083]. corners [1053]. Corpus [3002]. Correct [3016, 662, 1946, 2334, 3019, 3020, 3323]. Correcting [1942, 92, 545]. Correction [1774, 1120]. Correctness [1385, 1594, 453, 3006, 657, 1712, 628]. Correlation [3102, 107]. Correspondence [1177, 1890, 1870, 1061, 1065]. Cost [1692, 2032, 3058, 2456, 666, 856, 1020, 608, 263, 2042, 3108, 1537, 2687, 1006, 956]. Cost-Sensitive [608]. Costs [3043, 1737]. Cosy [661, 661]. Cosynthesis [2455]. Cottage [2931]. Could [2689]. Count [2332]. Countable [2124]. Counterexamples [2274]. Counters [1610]. Counting [2719, 2089, 2287, 1204, 3453, 1456]. Coupled [2803, 2605, 1721, 646, 3144]. Coupled-Context-Free [1721]. Coupling [2852, 905, 646]. Course [14, 2483, 12, 33, 34]. Courses [15, 11, 14, 13]. Cover [1972]. Coverage [2557, 2563, 574]. covered [1507]. Coverings [1952]. Covers [2435, 2027, 3259]. CPS [677]. CPS-Translation [677]. CPU [3185]. CQL [1686]. Craig [1411]. CRANIUM [2609]. Crashworthiness [858]. Cray [3508, 1004, 949, 3520, 2610]. Creating [3168]. Creation [2521, 2508, 2534]. Creativity [2150]. credible [3217]. Crew [2716]. crisis [1416]. Criteria [2175, 2419, 3083, 1619]. Criterion [1385, 100, 312, 813]. Critic [3142, 1417]. Critical [556, 3489, 2166, 3353, 2442, 3029, 2315, 2566, 1465]. Cross [1868, 1871, 375, 1889, 706]. cross-fertilization [706]. Cross-Ratios [1868]. Crossover [3116, 3118]. Cryptanalysis [233, 2242, 132, 1286, 2226, 1288, 2216, 2215, 3410, 2213, 126, 2214, 235]. Cryptanalytic [127]. Crypto [2224]. Crypto-Chip [2224]. Cryptographic [2219, 241, 2240, 2225, 2233, 1296, 1371, 1372]. Cryptography [2234, 3272, 257, 97]. Cryptologist [103]. Cryptosystem [2092, 96, 2245]. Cryptosystems [2242, 1290]. Crystal [896, 897]. Cs [361, 950]. CS-2 [950]. Cs-Congruence [361]. CSCW [2741]. CSP [1598]. CTA [3200]. CTL [1643]. Cube [2596, 2586]. Cubes [1413, 2593]. CUE [3165]. cues [1116, 1114]. Cultivating [3443]. Culture [47]. Cumulative [1014]. cure [1416]. Curing [2959]. Current [3489, 2897, 3235]. Curriculum [22]. Curvature [2710, 1110]. Curve [95, 1123, 1075, 3469]. Curved [1849, 1882]. Curves [2707, 1876, 2218, 1867, 1891, 3533, 1880, 2048, 1126, 3448, 1124, 3473, 3453, 1046, 801, 3452, 1053]. Custom [2446, 2470]. Customising [1759]. Cut [379, 1028, 2701]. Cuts [1727, 2036, 2077]. Cutting [1709, 2753]. Cycle [2027, 2045]. Cycles [2719, 3368, 1731, 3451]. Cyclic [1204, 1565].

d [2739, 1074, 3282, 3492, 1048, 2075, 2480, 1883, 874, 997, 3284, 1887, 3153, 1891, 2760, 1116, 1059, 1885, 1870, 1088, 1872, 1119, 1082, 1037, 1091, 1108, 62, 1121, 872, 876]. D-optimal [3153]. D0 [2198]. D2014 [3201]. DAG [2718]. Damped [2611]. Dangerous [1369]. Data [2776, 1683, 65, 368, 1384, 2449, 3169, 568, 1797, 1180, 144, 667, 556, 305, 2622, 151, 2669, 2451, 1830, 533, 3339, 1001, 439, 1002, 2859, 1653, 666, 1005, 1662, 3497, 2731, 982, 1687, 550, 153, 2602, 1223, 3273, 2411, 3095, 454, 1661, 3274, 3285, 158, 903, 1252, 607, 3319, 3277, 1571, 260, 3321, 3278, 2108, 2042, 3055, 1534, 1242, 2706, 528, 850, 501, 3324, 77, 988, 1679, 165, 2748, 1509, 1822, 1538, 2214, 1606, 1567, 308, 1003, 145, 172, 1996, 503, 1685, 659, 3275, 3582, 961, 1337, 440, 694, 2330, 2567, 902, 567, 924, 3276, 1289, 3280]. Data [3281, 1805, 3040, 1565, 3400, 1904, 2142, 3256, 2213, 1606]. data-driven [3256]. Data-Flow [1180, 667, 666, 694]. Data-Oriented [3280]. Data-Parallel [2622, 1001, 982, 145]. Data-Procedural [2449]. Database [2776, 1346, 3386, 2795, 381, 1680, 1667, 2419, 1669, 2788, 1939, 1675, 1924, 2808, 2796, 3421, 3282, 1658, 2843, 1832, 2842, 3420, 438, 439, 921, 3370, 1673, 2733, 2768, 2803, 453, 1656, 2797, 3385, 1820, 2786, 2841, 2734, 2742, 2800, 2804, 733, 2769, 1678, 922, 410, 3277, 1668, 2760, 2922, 1600, 1808, 1679, 1606, 2801, 1686, 2793, 1682, 1825, 2951, 2765, 1660, 2752, 2750, 2794, 3279, 404, 1665, 1800, 458, 1659, 3267, 1906, 1899, 1903, 1901, 1905]. Database-Coupled [2803]. Databases [2190, 2810, 1754, 888, 436, 2112, 2744, 1670, 1819, 452, 1828, 1684, 3283, 213, 2780, 2298, 2785, 434, 759, 2804, 3285, 1829, 2755, 3392, 2805, 609, 2792, 2756, 449, 3423, 1348, 2771, 464, 2746, 1657, 407, 2757, 3422, 1925, 1801, 314, 1897, 1898, 1900, 1896, 3228, 3217, 1902, 1895]. Dataflow [3017, 1599]. Dataflow-Based [1599]. Datalog [408, 1437]. Datapath [2482]. Datapaths [2445]. Datarol [1604]. Datarol-II [1604]. Datatypes [2873]. David [478]. Davis [2368]. Day [2920]. daylight [1089]. DBJ [458]. DBMS [1677, 1575, 1672, 437, 2758, 1676]. DBMS-Based [1676]. DBPL [1681].

DBROUGH [3250]. DBS3 [1575]. DC [1773]. DDAS [1569]. Deadline [3042]. Dealing [1913, 3239]. Debugging [991, 175, 665, 3376]. Decade [3554]. Decades [3305]. Decentralized [116]. Decidability [679, 2389, 363, 1156, 1624, 1927, 3553, 2115]. Decidable [763, 684, 2373, 1625, 1461, 2317, 3039, 2116]. Deciding [382, 1691, 1357, 2117]. Decision [3164, 2288, 2881, 498, 2081, 608, 313, 585, 2311, 2541, 590, 3564, 495, 3549, 1965, 2080, 2368, 1926, 3229, 3253, 3224]. Decision-Making [1965]. decisions [3246]. Declarations [2965]. Declarative [1755, 3236, 1789]. Decoders [573]. Decoding [1952, 1947, 215, 1951, 818, 813, 817]. Decomposability [382]. Decomposing [3105, 777]. Decomposition [1326, 3484, 3421, 676, 960, 3547, 423, 1360, 965, 2306, 903, 3455, 2062, 314]. DECT [530]. Dedicated [2521, 1283, 3201]. Deduction [2336, 1933, 1938, 1231, 2365, 1428, 3265, 1480, 1454]. Deductive [2776, 1754, 3287, 1781, 439, 2733, 2298, 410, 2805, 2792, 1341, 1439, 404, 1898, 1481, 3271]. Deep [2889]. Deepness [1211]. Default [1301, 2202, 1306, 1308, 2958, 1316, 2959, 3269]. Defaults [1303, 405, 2952]. Definable [2301]. Define [2917]. Defined [394, 475, 787]. Defining [755, 2509, 626]. Definite [2363, 2347, 3243]. Definition [2083, 1557, 1345, 469, 703, 3260, 1502]. Definitional [1023, 1031]. Definitions [755, 3538, 1305, 1018, 3345, 476, 2103, 1151, 1410, 704, 1483, 3249, 1426]. Deforestation [3070, 694, 726]. deformable [1134]. Degree [2044, 2693, 2218, 49, 2291, 1850, 3543, 3459]. Degrees [472, 2300]. Delay [2563, 3513, 2429, 2403, 549, 2091]. Delayed [3087, 653, 3088]. Delays [3186]. Deletion [1364]. Deletions [2064]. Deliberative [1927]. Deliverer [2184]. Delivery [2802, 21]. Delta [293, 3148, 1176, 1473, 3201]. Delta-4 [293]. delta-Functors [1176]. delta011 [1773]. delta011-CA [1773]. Demand [2397, 2408, 2409, 668]. Demonstrating [3206]. Demonstration [3436]. Demonstrational [3431]. Denotational [740, 659, 707]. Dense [958, 2066, 3521, 1107]. Densities [2178]. Density [2465, 3446, 2033, 2464]. Density/ Performance [2464]. Dependability [2574, 2576]. Dependable [2578]. Dependence [513, 667, 2352, 2674, 663, 2675, 3395]. Dependence-Conscious [513]. Dependences [172]. Dependencies [3388, 3060, 3414, 1801, 1565]. Dependency [3059, 2806]. Dependency-Directed [2806]. Dependent [1643, 512, 831, 1790, 869, 1238, 2232, 3326, 1569, 1498]. Deployment [2442]. Deposition [2726]. Depth [1833, 1551, 332, 3537, 301, 1116, 1134]. Derivation [435, 2625, 3032, 1483]. Derivational [2167]. Derivations [1001, 385, 386, 389, 1408]. Derivative [3117]. Derived [3291, 2881, 1798, 3255]. Deriving [2145, 3060, 3370, 2626, 1584, 1114, 1341, 2332]. Describing [1662, 1793]. Description [472, 3577, 387, 2207, 2542, 365, 2900, 1760, 2952, 1886, 1297, 70, 1438]. Descriptional [1362, 366]. Descriptions [582, 64, 3160]. Descriptor [77]. Design [1012, 2559, 3010, 1669, 2436, 3426, 2668, 35, 2471, 2164, 2606, 3282, 3170, 1761, 37, 881, 2218, 2231, 3367, 2876, 1275, 224, 2611, 3207, 2, 519, 1286, 539, 2457, 1391, 1807, 2483, 31, 338, 2474, 2434, 3381, 3147, 1327, 1596, 10, 2183, 2442, 189, 2815, 300, 3438, 208, 2884, 3437, 275, 2570, 2466, 782, 1808, 59, 3354, 1269, 2783, 115, 3394, 2861, 448, 990, 2888, 1749, 1283, 3152, 2478, 1344, 2978, 67, 2628, 58, 2765, 203, 2752, 3425, 992, 2572, 1798, 2650, 3041, 90, 1194, 705, 223, 634].

Designated [2220]. Designing [1346, 3141, 3404, 3016, 2843, 3414, 2, 2555, 391, 2906, 2746, 2247, 2230]. Designs [1574, 3153, 2885]. Desk [2391]. Desnakification [2721]. Destruction [1793]. Destructive [402]. Detailed [870]. Detecting [2707, 1431, 2619, 587, 2950, 1592]. Detection [356, 3402, 543, 2539, 2571, 2570, 2026, 82, 1375, 544, 2580, 839, 3057, 808, 1071, 1070, 1052, 1101, 1137, 822]. Detector [532]. Detectors [2834, 542, 1055]. DETERM [1811]. Determinant [2022]. Determinants [1868]. Determination [2985, 2096, 1099, 1138]. Determine [3088]. Determining [2019, 288]. Determinism [3384, 3401, 2124]. Deterministic [1784, 2285, 1395, 1811, 1696, 2833, 2052, 2057, 325, 2722, 692, 335, 788]. Determinizable [1617]. Determinizing [1698]. Develop [1670]. Developing [2755, 1229, 658, 973, 1655, 87]. Development [3092, 2862, 3349, 2630, 2457, 2618, 3366, 946, 507, 1549, 2469, 2472, 1329, 2735, 3512, 3344, 2848, 3030, 975, 2864, 2540, 450, 2801, 2858, 3008, 2846, 3209, 187, 1324, 474, 1325, 3356, 1476, 1484]. Developmental [2629]. Developments [954, 1474]. DEVI [3377]. Devices [456, 865, 2408, 2461]. DFR [3209]. DFS [2718]. DFS-Forests [2718]. Diagnoser [1789]. Diagnosis [2574, 2163, 2172, 2584, 85]. Diagnostic [3163, 2174, 2677, 1491]. Diagonal [1207, 2697]. Diagonals [334]. Diagram [1839, 2708, 2080]. Diagrams [2289, 2881, 2081, 2016, 374, 764, 2311, 1190, 3564, 2368]. Dialog [2946]. Dialogue [51]. Dialogues [1782]. Diameter [1706, 1862]. diamond [1125]. Dictionaries [1250]. Dictionary [1250, 1661]. Dictionary-Matching [1250]. Difference [3513, 3340, 2337]. Different [3133, 867, 3564, 352]. Differential [233, 3486, 1646, 83, 1286, 2215, 3519, 234, 1624, 544, 808, 1117, 1075]. Differential-Linear [2215]. Differentially [97]. Differentiation [959, 1956]. Differing [503]. Difficult [200]. Difficulties [1876]. difficulty [3445]. Diffie [2238]. Diffraction [912]. Diffuse [3132]. Diffusion [78, 2777, 1099]. Digest [2439]. Digital [1942, 116, 2221, 3412, 2463, 1767, 2483, 2935, 2392, 2404, 2884, 572, 528, 2484, 571, 2934, 131, 2443, 2561, 1540, 3093, 810, 3215, 803, 130]. Digraphs [311]. Dilated [2591]. Dilation [1202, 2678]. Dilemma [576, 1960]. Dimension [717, 304, 1969, 301, 1090]. Dimensional [1705, 1868, 539, 1005, 1946, 1244, 865, 2597, 1203, 1953, 3377, 1916, 1601, 877, 1097, 1038, 1118]. Dimensioning [987]. Dimensions [2019, 374, 2200, 310, 616]. Dining [2835]. Diploid [3099]. DIPOL92 [615]. Direct [1179, 879, 1112, 2443, 549, 542, 1102]. Direct-Sequence [549, 542]. Directed [2221, 3080, 651, 2718, 1727, 2806, 3148, 27]. Direction [2749, 1098]. Directional [3045]. Directions [3396, 37]. Directly [1243]. Disabilities [2895]. Disability [2922]. Disabled [2926, 2919]. Discarding [2405]. Discipline [2, 2398]. Disclosure [3423, 2775]. discontinuities [1099]. discontinuous [1062]. Discourse [2096]. discover [1465]. Discovering [599, 756, 588, 314]. Discovers [3129]. Discovery [2544, 2769, 3321, 3228, 3250, 3254, 3235]. Discreet [244]. Discrete [229, 2219, 2238, 1620, 612, 3448, 3477]. Discriminant [296]. Discriminate [2974]. Discussion [24, 128]. Disequality [2350, 328]. Disjoint [2017, 2038, 1852, 2307, 1729, 2700, 1434]. Disjunctive [408, 2797, 1785]. Disk [2225, 273, 161]. Disk-Based [161]. Disorder [2283]. Disparity [1073]. Disparity-space [1073]. Dispatcher [2479]. Dispatcher-Collector [2479]. Dispatching [1765]. Dispelling [3338]. dispersion [3478]. Display [3273]. Dissemination [1215, 2413]. Dissipation [2429]. Distance [3200, 51, 123, 2034, 1866, 2516, 1259, 2643, 1949, 1050]. Distance-Hereditary [1866]. Distances [2963]. Distinct [1829]. Distinguishing [2925]. distortion [1110]. distortions [1116]. Distribute [159]. Distribute-based [159]. Distributed [1572, 906, 2826, 2836, 2814, 1205, 384, 2838, 991, 2560, 1487, 3490, 781, 3492, 2740, 2241, 3416, 3213, 2537, 938, 151, 2875, 522, 1666, 955, 3414, 149, 959, 3205, 989, 1204, 784, 2662, 838, 3499, 772, 783, 896, 1553, 874, 1633, 3272, 1589, 174, 1757, 434, 649, 3083, 3023, 2027, 1570, 3180, 9, 932, 866, 1571, 1963, 2538, 2685, 3572, 3204, 2652, 867, 2706, 3031, 782, 974, 2812, 1961, 3199, 1581, 1765, 2783, 1538, 983, 987, 293, 2888, 1577, 945, 1556, 1611, 1592, 3144, 444, 295, 999, 3206, 972, 2831, 776, 2589, 933, 3240, 2818, 774]. Distributed [964, 876, 3594, 1906]. Distributed-Memory [896]. Distribution [256, 237, 199, 1734, 132, 151, 154, 1700, 1002, 3207, 773, 158, 3325, 124, 251, 137, 1071, 2133]. Distribution-Free [3325]. Distributions [831, 997, 1265]. Distributive [2302]. Disunification [2373]. divergence [1417]. Divergent [3021]. diverging [1087]. Diversification [3077]. Diversity [2570, 2572, 3099]. Divide [926]. Divide-and-Conquer [926]. Divided [1143]. Division [2727, 532, 549, 3449, 3447]. divisionless [802]. divisor [3465]. Dlogtime [340]. DNA [2277, 309]. DNF [3309]. Do [150, 2270, 2734, 2590, 988, 1585]. Do-Loop-Surface [988]. Document [2788, 1674, 1512]. Documentation [2179]. Documented [3432]. Documents [2993, 2924, 2926, 2740, 2925, 3570]. Does [2769, 2845, 852]. DOF [1040]. Doing [13]. Domain [488, 2163, 2523, 1715, 1776, 1478]. Domains [209, 1015, 3046, 1164, 1356, 681, 1168, 1169, 3378, 1663, 605, 685, 1153, 1170, 3216]. Dominating [1866, 376]. Don’t [1254, 1635]. Door [645]. Double [1873, 385, 1292]. Down [43, 2737, 1473]. DP1 [611]. DPL [778]. DPSK [809]. DQPSK [554]. Drawing [416, 1120]. Drawings [2691, 2690, 1970, 2070]. Dream [576]. Driven [2170, 2550, 2451, 84, 3322, 3556, 2429, 2761, 3042, 1123, 3043, 3256]. Driver [2937]. Drives [2225]. Driving [3177, 2336, 796]. Drug [881]. DS [538, 572, 540, 812]. DS-CDMA [538, 540]. DS-Spread [572]. DS/CDMA [812]. DSA [110]. DSP [2472, 2443]. Duality [1980, 281]. Duals [2104]. Duper [496]. Durability [1685]. Duration [3014, 3037, 3009, 2135]. Durations [3028]. During [1342, 450, 3428, 330]. Duumvirate [2337]. Dwell [548]. Dynamic [2807, 2795, 1667, 559, 529, 3246, 2414, 884, 225, 1763, 971, 1356, 2686, 986, 1591, 2352, 2718, 830, 1847, 3126, 550, 3391, 2773, 68, 2769, 1612, 3519, 1582, 2004, 143, 1793, 2996, 1860, 1588, 1332, 2400, 2454, 1743, 2790, 3390, 458, 2129, 1504, 3232, 1054]. Dynamical [1693, 3523]. Dynamically [1724, 2411, 3570, 1764]. Dynamically-Typed [1724]. DynamicPVM [971]. Dynamics [3480, 2190, 3489, 3490, 878, 3492, 599, 871, 3509, 1957, 3512, 3097, 893, 897, 886, 894, 3119, 3073, 3527, 139, 3518]. Dynamization [426].

E- [1464, 1462]. E-Complete [2270]. E-Unification [1778, 2195, 1463]. Early [141, 610, 3171]. Ears [2306]. earth [3213, 3213]. Easy [2904, 2300, 110]. Easy-to-Use [2904]. EASYPVM [970]. ECA [1894]. echocardiographic [1079]. ECMWF [899]. Eco [1365]. ECOLOG [2382]. Economics [506]. Economy [365]. Eden [2726]. Edge [2707, 1709, 2288, 2038, 1852, 684, 3113, 2039, 2700, 2699, 2037, 1865, 1055]. Edge-Coloring [2037, 1865]. Edge-Disjoint [2038, 1852, 2700]. Edge-Ranking [2699]. Edit [1240]. Editing [2731]. Editor [1226, 1483]. Editors [2924]. Education [32, 26, 40, 48, 44, 3, 51, 30, 38, 2516, 23, 41, 3234]. Educational [3442, 2459, 18, 2929]. EECA [2770]. EER [1809]. EFECT [3511]. Effect [555, 2595, 3096]. Effective [1386, 1873, 2579, 3367, 856, 2860, 1586]. Effectively [958]. Effectively-Dense [958]. Effectiveness [263]. Effects [3428, 620, 3140]. Efficiency [543, 2341, 3318, 1420]. Efficient [3472, 3059, 3010, 2693, 2665, 1205, 270, 3286, 1234, 2331, 3308, 3295, 1781, 3411, 1763, 2046, 2035, 2327, 2632, 1209, 3547, 354, 2197, 2235, 2293, 2626, 240, 2878, 836, 1641, 3231, 170, 3385, 1696, 2682, 427, 649, 2071, 454, 516, 3064, 2133, 2066, 2781, 161, 2045, 2306, 1259, 1903, 825, 3533, 1571, 2756, 573, 2042, 1534, 2739, 1664, 325, 1739, 3048, 3325, 2433, 2603, 165, 1535, 113, 3521, 143, 658, 2405, 3556, 945, 659, 3564, 1786, 214, 3579, 1864, 2377, 3528, 1917, 3062, 1712, 1764, 2790, 2084, 2699, 1502, 3192]. Efficiently [167, 1973]. egomotion [1080]. Eiffel [3349]. Eigenproblems [860]. Eigensolution [3524]. Eigenstructure [3483]. Eigenvalues [3485]. Elasticity [290]. Elderly [2904]. Election [1851, 687, 113]. Electromagnetic [864]. Electron [912]. Electronic [880, 2924, 3440, 2902, 3411, 424, 553, 2936]. Electrosmog [864]. Element [2653, 859, 869, 1211, 3503, 3504, 2178, 965, 892, 1672, 928, 863]. Elementary [3015, 498, 1374, 1161]. elements [1127, 818]. Elf [1481]. Eliminating [1564, 495]. Elimination [514, 1224, 3071, 1422]. Elliptic [2218, 95, 3473, 3453, 3452]. Embedded [1017, 856, 2442, 299, 302]. Embedding [661, 472, 1202, 2286, 1848, 2889, 1203]. Emerging [297]. Emotion [1959]. Empirical [3078, 3444, 588, 2154, 3427, 1325]. Employers [2933]. Empty [2309, 3546]. EMS [1669]. Emulation [2471, 2453]. Emulations [2722]. Emulator [2454]. Enablers [289]. Enabling [2522, 553, 1494]. Enactment [202, 199, 192, 193]. Encapsulated [737, 3382]. Encapsulation [2339, 522]. Encipherment [2214]. Encoding [2600, 1967, 1227]. Encrypting [1282]. Encryption [258, 240, 2213, 2220, 1281, 1289, 810]. encyclopedia [1432]. End [570, 2532, 2524, 3576, 648, 2804, 2393, 3279]. End-systems [2393]. End-to-End [2532, 2524, 3576]. End-User [3279]. Endeavour [2929]. Endstations [3178]. Endsystem [3588]. Energy [912]. Enforcement [2804, 1901]. Engine [855, 1226]. Engineering [1683, 42, 26, 40, 48, 3, 4, 51, 1819, 2167, 43, 53, 38, 2646, 12, 210, 28, 1820, 23, 3433, 2781, 1334, 2884, 2510, 6, 13, 41, 56, 34, 3427, 5, 518, 87, 623, 3254, 49]. English [3002]. Enhanced [2519, 2401, 2914, 1544, 970, 528]. Enhanced-TDMA [528]. enhancement [3248]. Enhancements [3507]. Enhancing [3076, 1685, 1800, 1565]. Enough [318]. Enriching [2744, 2948]. ENSOLV [1011]. Enterprise [282, 1331]. Entities [1300, 1798]. Entity [237, 1819, 1811, 2772, 140, 1822, 1806, 1804, 1810, 1813]. Entity-Relationship [1819, 1811, 140, 1822, 1806, 1810, 1813]. Entrenchment [2961]. entropy [825]. entropy-coded [825]. Enumerable [3290, 3297]. Enumeration [2998]. Enumerations [1394]. enumerative [1122]. enumerator [1478]. Environment [906, 280, 556, 2471, 3442, 938, 2451, 2528, 114, 844, 784, 3005, 1714, 3211, 1821, 2508, 2382, 1549, 3433, 2769, 944, 937, 2629, 3572, 3204, 1321, 1615, 867, 975, 88, 1739, 2616, 1795, 2534, 3166, 1547, 29, 1825, 980, 3582, 554, 1595, 3183, 1324, 2775, 3526, 2762, 916, 2182, 2753, 2566, 485, 566, 3248, 1476]. Environmental [1766, 1668, 3281, 928]. Environments [3480, 2740, 3494, 2863, 1794, 2521, 10, 414, 790, 3441, 66, 560, 968, 3206, 546]. ENZO [3138]. ENZO-M [3138]. EOL [2318]. EP [951]. Ephemeral [3573]. Epipolar [1036, 1869, 1146, 1047]. Epistemic [2953, 1305, 3540, 2961, 1965, 3257, 3259]. EPOCA [844]. Epoch [1482]. EPOS [1319]. equal [1514]. Equality [429, 1408, 1996, 2196]. Equality-Testing [429]. Equalization [546, 803, 811]. Equalizers [545]. Equals [1013, 1027]. Equation [2495, 1712, 1380]. Equational [2494, 2387, 2966, 1724, 3312, 1434, 1456, 1427]. Equationally [394]. Equations [2867, 3486, 2271, 83, 2635, 2293, 2068, 2304, 1367, 3519, 873, 1720, 1379, 2014]. Equations-Solving [2014]. Equilibratory [1963]. Equipments [2903]. Equiprobable [1737]. Equivalence [1643, 1981, 420, 3309, 1990, 794, 1908, 1366, 2238, 441, 688, 2728, 2318]. Equivalences [1717, 751, 732, 2115]. Equivalent [2336, 1394, 2220]. ER-Based [1825]. ER-Model [1661]. Erasure [536]. Eratosthenes [1225]. EREW [2717]. Ergonomics [3439]. Ergosemiotics [3425]. Erlangen [60]. Errata [2000]. Error [1942, 809, 285, 2565, 547, 803, 1592, 548, 2580, 303]. Error-control [803]. Error-Correcting [1942]. Error-Detecting [1592]. Errors [1946, 92, 558]. ERT [2758]. ERT-SQL [2758]. ESM [392]. ESPRIT [3412, 1010, 857]. Essence [796, 524]. Essential [3369, 2103, 1045]. Essentials [3242, 1749]. Establishing [2827]. Establishment [3409]. Esterel [1652]. Estimates [1020]. Estimating [3105, 3058, 847, 586]. Estimation [3012, 2990, 2987, 547, 837, 457, 1103, 1062, 1048, 1112, 1066, 1094, 1080, 1116, 1043, 1045, 811]. Estimators [549]. Estuary [904]. Ethics [38]. Euclidean [1867, 1878]. Euler [963, 874, 875, 873]. Euler/Navier [875]. EUODHILOS [1482]. EUODHILOS-II [1482]. Eurescom [3182]. Europe [3200, 929, 852]. European [857, 935, 3202, 941, 2545]. EUROPORT [857]. Evaluate [2509]. Evaluating [2673, 1574, 3157]. Evaluation [2475, 950, 1521, 2455, 994, 460, 2737, 1819, 2558, 2593, 3248, 838, 3345, 1639, 953, 530, 557, 796, 2410, 922, 952, 1644, 867, 3429, 3359, 3148, 2161, 3119, 67, 440, 826, 3262, 1437, 722, 726]. Evaluations [962]. Evaluators [646]. Even [106, 2983, 2328, 1731]. Event [147, 1666, 1811, 2206, 2565, 2621, 3436, 612, 1375, 400, 2761, 1925, 1895]. Event-Based [147, 2621]. Event-Driven [2761]. event-oriented [1895]. Event-Tuned [1811]. Events [1921, 1332, 3239]. Eventual [2496]. Every [2017]. evidence [3222]. Evidential [2747]. Evolution [3101, 3138, 3079, 2842, 3098, 1867, 2773, 3111, 1342, 3117, 449, 527, 1348, 63, 2774, 407, 890, 3143, 3096, 1659, 1123, 720]. Evolutionary [3104, 3136, 2530, 3090, 3122, 3078, 3151, 1034, 3077, 3091, 3076, 3152, 3148, 3109, 3155, 3074, 3100]. Evolvable [3145]. Evolving [183, 2791, 1958, 3086, 414, 2745, 2775, 3140]. Exact [2730, 2640, 307, 172, 339, 2727]. Examination [3346, 2677]. Example [272, 2895, 2913, 88, 205]. Examples [2865, 270, 3286, 2208, 3308, 3295, 1352, 3442, 83, 582, 3312, 3313, 604, 3320]. Exaptive [3130]. ExBase [3278]. Exceptions [2768, 653, 3057]. Exchange [3404, 108, 109]. Exclusion [504, 1858]. Exclusions [1015]. Executed [575]. Execution [1543, 2537, 1750, 653, 269, 1580, 2620, 1593, 1751, 143, 1548, 2416, 1569, 916, 839, 3185, 3231]. Exemplar [3516]. Existence [2083, 238, 595, 239, 1418, 624]. Existentially [2235]. Existing [3285, 1685, 3590]. Exotic [2821]. Expandable [836]. Expander [425]. Expanders [2604]. Expansion [1317, 2485]. Expansive [3094]. Expectations [1260]. Expected [370, 1550, 3544]. Experience [1011, 15, 4, 2539, 2656, 148, 3495, 33, 2466, 13, 2932]. Experiences [32, 3486, 2463, 141, 828, 840, 19, 3499, 1334, 977, 3395, 27, 3337, 204, 2929, 863, 3518]. Experiment [206, 2583, 3139]. Experimental [3355, 1016, 3205, 982, 2213, 827]. Experimenting [1964]. Experiments [3163, 1421, 995, 1957]. Expert [597, 2791, 2777, 2767, 81, 2802, 2753]. Expertise [3168]. experts [3217]. Explained [2156]. Explanation [2170, 268, 2163, 3301, 133, 2176, 3561, 1428, 134]. Explanation-Based [2163, 3301, 1428]. Explanation-Driven [2170]. Explanatory [3442]. Explicit [2294, 874, 3345, 731, 1228, 1266, 1997, 2099, 3466]. Explode [1635]. Exploiting [2111, 3005, 1714, 153, 2809, 3077, 1863]. Exploration [2795, 1646, 2859, 3278, 3281]. Exploratory [2705]. Exploring [1458, 3431]. Exponential [2282, 1171]. Exponentiation [2223]. Exporting [1453]. Express [778, 968]. Expressed [1809]. Expressing [1894]. Expression [3577, 1400]. Expressions [270, 2673, 3308, 3295, 1644, 1505, 1140]. Expressive [408, 1396, 1695, 1228, 3417]. Expressiveness [1696, 409]. Expressivity [1897, 2622]. Extended [1819, 2733, 3342, 140, 1452, 2786, 2069, 1661, 3274, 1822, 1786, 2080, 1497, 1509, 1894, 3192]. Extendibility [784]. Extending [383, 3186, 702, 2212, 2395, 2840, 160, 1807, 479, 1827, 3436, 3565]. Extends [574]. Extensible [740, 1795].

Extension

[2663, 1472, 2218, 3418, 2916, 743, 1421]. Extensional [1725, 1397, 1460]. Extensionality [1224]. Extensions [1301, 1330, 242, 1802, 36, 2674, 586, 732, 1532, 2959, 1506]. Extent [3055]. Exterior [3504]. External [3566]. Extra [1784, 2469]. Extra-Fast [2469]. extracted [1107]. Extracting [3061, 1069, 2099, 3246]. Extraction [1779, 753, 1092, 785, 1115, 1140]. Extrapolation [3486, 2636, 1556]. Extreme [2826, 1392]. Eye [2942, 1057, 1144]. eye-head [1057].

F [3227]. F-logic [3227]. Face [1098]. facial [1140]. Facilitate [3168, 1342]. Facility [1346, 2588, 1810]. Facing [261]. Factor [2421]. Factoring [94, 2816, 3456, 3477, 1442]. Factorization [231, 3467, 3454, 1051]. Factorizations [96]. Factors [1596, 3515]. Factory [2751]. Fading [538, 535, 541, 537, 540, 560, 565, 809, 806]. Fail [238, 239]. Fail-Stop [238, 239]. Failure [43, 1308, 2834, 2835]. Failures [3550, 302, 133, 134]. Fair [1717, 3532]. Fairness [1648, 3552, 1598, 2135]. fall [1143]. Fallacies [2274]. Falling [43]. False [156, 1774]. Families [245, 2981]. Family [2465, 1007, 746, 1001]. Far [549]. Fast [2692, 2591, 3487, 1277, 1954, 2819, 2046, 1239, 522, 3069, 3090, 2111, 1728, 1974, 1707, 2435, 2712, 2469, 1721, 1276, 2052, 1310, 2685, 1243, 2239, 2824, 1295, 1674, 2259, 2448, 3271, 2580, 1978, 2086, 2117, 3462, 2214]. Faster [2696, 3566, 2590, 334, 327, 1731]. Fat [2724, 2709]. Fat-Trees [2724]. Fault [2559, 3010, 297, 2557, 2569, 2606, 305, 2579, 2558, 261, 1213, 284, 2605, 2588, 271, 2053, 2424, 2554, 2853, 299, 3022, 273, 3027, 2581, 302, 3007, 287, 2587, 263, 289, 3030, 265, 2570, 2564, 2026, 2584, 293, 295, 2556, 278, 992, 2572, 2586]. Fault-Based [2853].

Fault-Tolerance [992, 2586]. Fault-Tolerant [2559, 3010, 2606, 2579, 1213, 2605, 2554, 3022, 3007, 3030, 2556]. Faults [1692, 2250, 2723, 2832, 2571, 2563, 287, 2454, 2566]. Faulty [2585, 2676, 2481, 1690]. Favorite [3554]. FD [2379, 2343]. FDM [529, 1903]. FDM/ TDMA [529]. FE [859]. Feasible [358, 2581]. Feature [2796, 2539, 2984, 600, 1885, 2747, 1115, 1140, 3256]. Features [2388, 1912, 2840, 1338, 1004, 1890, 3263]. Federating [2780, 1816]. Federation [1682]. Feedback [559, 569, 3431, 3428]. Feelable [2909]. FeelMouse [2909]. Feistel [1299]. fertilization [706]. FFT [2631, 1291]. FFT-Hashing [1291]. FIAT [171]. Fiber [1553]. Fibonacci [2065]. Fibrational [759]. fibred [635]. Field [230, 2470, 2480, 2682, 2428, 2482, 2459, 937, 531, 1880, 2478, 804, 1111, 1046]. Field-Programmable [2482]. Fields [229, 3055, 77, 2048, 872, 3448, 1062, 800, 3449, 3467, 1036, 3471, 1079, 1107, 3473, 3453, 801, 3466, 3450]. Fife [1350]. File [2814, 3416, 2410, 932, 1243, 2409, 2790]. Files [464]. Filter [2463, 572, 2413, 545, 571]. Filtered [2243, 3119]. Filtering [1954, 454, 3193, 814]. Filters [780, 1762, 1540, 1067]. FIN [288, 3297]. FIN-Identifiability [3297]. FIN-Type [288]. Final [1149, 1192]. Financial [1670, 1671]. Find [120]. FINDER [1478]. Finding [2826, 2719, 358, 368, 1234, 2703, 1843, 2290, 3313, 2027, 3063, 378, 2049, 1731, 3445, 1113]. Fine [2663, 680, 1604, 1580, 2486, 2466, 2346, 446, 165]. Fine-Grain [1604, 1580, 2346]. Fine-Grained [2486, 2466]. Fine-Granularity [446]. Finest [2321]. Fingerprint [924]. Finitary [1018]. Finite [229, 2092, 1386, 2653, 96, 1015, 859, 3017, 1183, 422, 869, 1017, 2296, 1211, 3503, 3504, 360, 2437, 2178, 310, 3378, 3063, 965, 892, 2308, 372, 536, 1403, 343, 290, 1672, 1173, 2317, 482, 1478, 1290, 3545, 1776, 2048, 863, 3448, 800, 1488, 804, 3467, 1127, 3473, 3453, 801, 1416, 1469]. Finite-Domain [1776]. Finite-State [1386, 1183, 482, 3545, 1488]. Finitely [2035, 3475]. Finiteness [1718]. Firms [2752]. First [742, 911, 1302, 2492, 1015, 3079, 2871, 1725, 3332, 411, 321, 895, 716, 1919, 2213, 2185, 790, 688, 34, 2995, 164, 1060, 1325, 3247, 1429, 722, 1450, 1427]. First-Class [1015, 790, 716, 722]. First-Order [742, 2492, 2871, 321, 1919, 688, 1429, 1450, 1427]. Fish [1277]. Fit [1970, 2713, 1071]. Fitch [1479]. Fitch-style [1479]. Fitness [3101]. Fitt [3426]. Five [42, 1972, 261]. Fix [1917]. Fix-Points [1917]. fixating [1112]. Fixed [1251, 2271, 1989, 3067, 1999, 697, 825]. Fixed-point [1989]. Fixed-Queries [1251]. fixed-rate [825]. fixedpoint [1426]. Fixpoint [3069, 1644, 3062]. Fixpoints [3063]. Flang [2364]. Flatness [1916]. Flexibility [2465]. Flexible [2807, 3200, 522, 2904, 2627, 2223, 782, 1320, 446, 1348, 613, 2551, 1100]. Flight [2804, 1082]. Floating [2571]. Floorplanning [584]. Flora [2736]. Flow [1011, 3402, 1180, 667, 3424, 1154, 666, 2974, 892, 1793, 936, 2354, 3523, 694, 1660, 2567, 656, 739, 3193, 872, 876, 1058, 1062, 1064, 1107, 1508, 1043, 1099, 1044, 1060, 861]. Flow-Fields [872]. Flows [869, 879, 904, 870]. Flowshop [3092]. fluent [3263]. Fluents [1912, 3263]. Fluid [3480, 871, 3523]. Fly [167, 1647, 3048]. FM8501 [3593]. FMC [3532]. FNT [2580]. Focus [3307, 2174]. Foil [3410]. FOL [792]. Folded [2676, 1535]. Following [1053]. force [3229]. forces [1108]. Forecast [899]. Forests [1864, 2718]. Forgery [1942]. Form [168, 690, 1590, 696, 3255, 802]. Formal [3355, 2436, 1631, 2557, 3338, 3577, 3170, 2874, 2876, 3370, 3331, 53, 3366, 3332, 2772, 2625, 3334, 2785, 36, 2442, 1329, 3348, 2542, 2884, 2885, 2553, 2968, 3350, 1760, 69, 658, 1793, 2952, 3072, 503, 3157, 2991, 1550, 3365, 469, 70, 776, 3041, 3042]. Formalisation [2539, 2886, 1337]. Formalised [2889]. Formalising [3369, 3333]. Formalism [1830, 2011, 3242]. Formality [472]. Formalization [1302, 587, 3340, 483, 490, 3260]. Formalizing [207, 473]. Formally [487]. Format [750, 362]. Formation [1198, 605, 909]. formedness [1514]. Forming [2993, 853, 1310]. Forms [1330, 703]. Formula [1643, 3032]. Formula-Dependent [1643]. Formulae [1028]. Formulas [3309, 161, 1424, 3220, 3466]. Formulated [958, 792]. Formulation [759, 3073, 1058, 715]. Forth [1921]. Fortran [664, 152, 994, 993, 998, 997, 1004, 996]. Forward [276, 561]. Forwarding [2412]. FOSSIL [583]. Foundation [3540, 2510, 1025, 2369]. Foundational [699]. Foundations [3566, 1391, 3392, 3387, 3425]. Founded [1309, 1402, 1785, 1898]. Four [1840, 301, 799]. Four-Connectivity [1840]. Fourier [2243]. FPGA [2475, 2455, 2471, 2456, 2430, 2457, 2460, 2439, 2434, 2445, 2442, 2441, 2467, 2486, 2484, 2444, 2427, 2429, 2448, 2454]. FPGA-Based [2434]. FPGA-to-ASIC [2460]. FPGAs [2431, 2479, 2483, 2469, 2428, 2424, 2458, 2466, 2432, 2433, 2481, 2443, 2426]. FPL [2449, 2477]. FPL-Based [2449]. FPMCM [2482]. Fractal [2440, 1689, 3177]. Fractional [537]. Fractional-Rate [537]. Fragment [1383, 2870]. Fragmentation [2555]. Fragmentation-Redundancy-Scattering [2555]. Frame [2971, 1867, 2405]. Frame-Induced [2405]. Frames [2803, 1103]. Framework [1826, 206, 1779, 3174, 1766, 2984, 2531, 14, 2402, 2667, 153, 171, 3066, 336, 2211, 3348, 410, 2755, 212, 670, 177, 1793, 3190, 1782, 3203, 1033, 3155, 2080, 2750, 1925, 2412, 3172, 85, 700, 1115, 634, 3268, 720, 1134]. Frameworks [211]. Free [1920, 1350, 1710, 1626, 2990, 1357, 2460, 426, 2998, 2602, 682, 1721, 923, 343, 3325, 2024, 2314, 1857, 2728, 1258, 440, 618, 3329, 2999, 1949, 2120, 2870, 2822, 2117, 2122, 1366, 2825]. Free-Text [923]. Freeness [3061]. Frequency [538, 535, 551, 557, 807]. frequency-hopped [807]. Frequency-Phase [535]. Frequency-Selective [538]. Front [570, 2804]. Frontier [434]. Fronts [1211, 1034]. FSM [2437, 75]. FTM [2583]. FTMPS [992]. FTMPS-Project [992]. Full [728, 1164, 463, 2063, 2472, 920, 396, 3517, 2320, 3254]. Full-Speed [2472]. Full-Text [463]. Fully [225, 2917, 729, 2297, 724, 2148]. Fully-Accessible [2917]. Function [421, 118, 1677, 117, 3063, 2503, 2965, 1745, 307, 2314, 120, 3120, 3119, 548, 618, 2246, 352, 3096, 2420, 119, 3457, 1058, 703, 823, 1460]. Function-Based [2420]. Function-Free [618]. Functional [2665, 2362, 35, 3539, 3353, 2207, 473, 2358, 2797, 2658, 3044, 647, 730, 1571, 2348, 2736, 737, 2890, 985, 2354, 2757, 1801, 787, 709, 715, 722, 723]. Functional-Logic [2348]. Functional-Style [2736]. Functionals [77, 1405]. Functions [3290, 3059, 3195, 2363, 3101, 2236, 245, 2221, 232, 1164, 1165, 99, 1646, 100, 678, 2059, 3297, 390, 2435, 3111, 249, 2081, 798, 1292, 2051, 1620, 2103, 1283, 248, 2687, 1284, 301, 1155, 107, 218, 1877, 1884, 2420, 1494, 1441, 3455]. Functors [1176, 721]. Fundamental [2707, 1876, 1147]. Fundamentals [140, 2013, 933]. Furniture [2811]. Further [2901, 2260, 2073]. Fusion [158]. Future [3396, 3502, 2517, 1919, 2576, 2534, 69, 936, 278, 2094, 803]. Fuzzy [2415, 2419, 2414, 2417, 3121, 2001, 2007, 2006, 2469, 2013, 3300, 2422, 1668, 3218, 2002, 3132, 2004, 2416, 2009, 2420, 2012, 3226, 3220].

GA [3154, 3097, 3119, 3099]. GA-Individuals [3097]. Gaining [3328]. GAITS [3234]. Galaxies [909]. Galois [3460]. Game [2388, 1983, 1960]. Games [244, 760, 1401, 1370, 1688, 3545, 828]. Gang [6]. Gap [23, 509, 2781]. Gaps [25]. Garbage [2331, 3420, 1577, 2332]. GARF [783]. Gas [878, 3493, 2974, 861]. Gas-Liquid [2974]. Gastric [610]. Gate [2425, 2480, 2459, 2562, 2478]. Gated [168]. Gateway [1681]. GATOSTAR [2588]. Gauge [908]. Gaussian [3461, 3462, 1067]. GC [2642]. GCD [2441, 2634, 3464]. GCel [968]. GCLA [1019]. GEANT [1010]. Gemmell [2226]. General [2847, 3442, 230, 655, 60, 2473, 1018, 2626, 480, 2943, 883, 2564, 1367, 88, 63, 2448, 1592, 1925, 2152, 85, 1898, 3237, 1506, 1484, 1463]. General-Purpose [2473, 480]. generalisation [1428]. Generalised [1999, 1465, 1066]. Generalization [1330, 1770, 272, 603, 587]. Generalizations [368, 1845, 680, 1945, 569]. Generalized [419, 1551, 1502, 1533, 3051, 2741, 1576, 1786, 616, 1801, 262, 964, 3259, 3470]. Generalized-Concentration [1551]. Generalizing [1349, 1887, 1765, 1718]. Generally [2560]. Generate [1748, 1661]. generated [821, 3475]. Generating [3388, 888, 2063, 3070, 3299, 3568, 161, 2885, 659, 2145]. Generation [3292, 1941, 129, 3287, 144, 1670, 1355, 662, 2402, 1549, 2424, 2302, 866, 1334, 2239, 1672, 3434, 2561, 2915, 3229, 1419, 1469]. Generations [2582]. Generator [2338, 216, 2882, 1278, 644, 643, 2889]. Generators [2338, 1285, 3468]. Generic [2208, 2456, 390, 3381, 1789, 1345, 1995, 1940]. Genericity [2282, 791]. GENESIS [950].

Genetic [3105, 3127, 3106, 3079, 3092, 3137, 3150, 3080, 3129, 3000, 599, 3103, 3084, 3075, 3126, 3147, 3095, 3083, 3125, 3128, 3082, 919, 3146, 3225, 3134, 3112, 3081, 3139, 2683, 3144, 3073, 3107, 1077, 3149, 3093, 3124, 2999, 3099, 3248]. Genetics [889]. Genome [264, 2746]. Genotype [3127, 3098]. Genotype-Phenotype [3098]. Genotype-Phenotype-Mapping [3127]. Gentzen [1262]. genus [3448]. Geographic [2776, 1667, 1766]. Geographical [2777]. Geometric [529, 245, 354, 600, 2712, 2725, 2651, 1855, 2259, 2255, 317, 1872, 1129]. geometries [1148]. Geometry [2034, 1206, 1867, 1845, 1689, 1123, 2113, 2760, 2073, 1869, 1146, 1443, 1097, 1446, 1122, 1047, 1442]. geometry-based [1097]. Geometry-driven [1123]. George [42]. German [2977]. Gestalts [2950]. gestures [1106]. Get [3186, 1920, 1585]. GHz [557]. GIG [3000]. Girard [1383]. GIS [2778]. Given [2719, 3473]. GKS [3337]. Glamis [832]. glasshouse [1892]. Glider [629]. Global [513, 2837, 129, 512, 3571, 3489, 348, 2654, 651, 647, 2428, 3374, 1908, 2923, 690, 3223]. Global/Local [3489]. Globally [1573]. GLS [3235]. GMSK [554]. GNU [1482]. Goal [1326, 202, 1790, 2966, 2100, 3326, 3235]. Goal-Dependent [3326]. Goal-seeking [2100]. Goals [13, 1431, 3252]. Goedel [3296]. GOL [3259]. GOLD [2621]. Golden [6]. gone [1892]. Good [3295, 2980, 3320]. Gossip [1852, 1216]. Gossiping [2307]. GPC [726]. Graded [2203]. Gradient [955]. Gradually [2069]. Graduate [42, 22]. Grain [2663, 1604, 1580, 2346, 165]. Grained [2087, 2486, 2466]. Gram [2985, 2632, 1380]. Grammar [384, 1354, 398, 1373, 694, 403, 752, 3001, 2999, 3004, 715, 1365]. Grammar-Based [694]. Grammar/Bigram [3001]. Grammars [2993, 380, 2992, 2990, 2984, 2288, 346, 1357, 673, 2998, 645, 1366, 3002, 646, 1258, 400, 2989, 656, 1435]. Grammatical [2993, 2997, 3000, 919, 2996]. Granularity [1913, 772, 2658, 446]. GRAPE [909]. Graph [380, 383, 384, 1564, 2691, 2656, 154, 2703, 422, 385, 386, 3534, 3418, 1932, 673, 1728, 1591, 2076, 416, 390, 2107, 3154, 2069, 1162, 1814, 2066, 378, 394, 2023, 395, 3031, 396, 2348, 2103, 398, 467, 2315, 399, 2978, 400, 403, 2039]. Graph-Based [1932]. Graph-Grammar [384]. Graph-theoretical [2978]. Graphic [989]. Graphical [2514, 2866, 2897, 1938, 3431, 29, 980, 2898, 696]. Graphs [1840, 667, 2221, 2112, 1972, 2078, 2874, 386, 388, 423, 377, 1866, 376, 2111, 651, 1845, 1846, 2718, 1727, 2038, 391, 1850, 1732, 1648, 2071, 1241, 2725, 2503, 2066, 2054, 684, 585, 1200, 1975, 2100, 1581, 163, 1729, 2643, 1858, 2079, 176, 402, 58, 3544, 2067, 2080, 2077, 2561, 2094, 2700, 2095, 2037, 1493, 2126, 1122]. Grasping [1039]. Gravity [3152]. Graz [2918]. greater [3453]. greatest [3465]. Greedy [3558, 2720, 2075]. Grey [1562]. GRH [3467]. Grid [2054, 873, 1856]. Grids [1543, 334, 1203, 963]. Griesmer [1945]. Ground [328, 1637]. groundness [718]. Groundwater [905]. Group [3188, 3199, 1722, 433, 3473]. grouping [1141]. Groups [1710, 96, 2569, 136, 782, 1885, 3446, 1092, 3475]. Groupware [1324]. Growing [43, 3430, 286]. Growth [678, 896, 2726, 897, 3449]. GSAT [3391]. GSM [556, 527]. GSOS [2145]. GSPN [840, 1509]. GTS [1579]. Guaranteeing [2605, 402]. Guarantees [3307]. GUI [2906, 2905]. Guidance [2164]. Guide [2514, 1757]. Guided [914, 45, 2996, 1429]. Guideline [868]. Guidelines [3167].

H [2351]. H0 [3517]. Hadamard [1948]. Halfplane [1974]. Hall [1722]. Hamiltonian [2692, 2016, 1200]. Hammock [2306]. Hammock-on-Ears [2306]. Hand [575, 2900, 2184, 1106, 1040]. Hand-off [575]. Handicapped [2904, 2897, 2929, 2932]. Handling [2776, 568, 2005, 2342, 2211, 3264]. Handoff [574]. Handshake [2827]. Handshaking [2314]. Handwriting [2338]. Hankel [3483]. Hard [241, 2020, 662, 2725, 1858]. Harder [2817]. Hardest [2967]. Hardness [1853, 2023]. Hardware [496, 472, 1631, 2455, 2471, 305, 2451, 1637, 473, 2571, 471, 3020, 3145, 509, 208, 2570, 495, 1638, 2461, 1640, 1518]. Hardware-Software [2455]. Hardware/ Software [2471]. HardWire [2460]. Harmonic [421]. Harmonization [3200]. Harmonizing [1675]. Harnessing [2525]. Harvesting [3443]. Hash [245, 117, 2233, 249, 1292, 120, 1283, 248, 1284, 458]. Hash-Values [2233]. Hashing [2234, 1646, 224, 1707, 2227, 115, 2697, 2217, 223, 1291]. Haskell [710, 3064]. HASTE [990]. Having [372]. HDM2 [1807]. Head [1016, 708, 1057, 1130, 1144]. head-eye [1144]. Health [2548, 2177, 3196]. Healthcare [1673]. Heap [2645]. Heaps [2065, 2047]. Heegner [3458]. Heidelberg [3185]. HEIPOET [3185]. held [3457]. Hellman [2238]. Helmholtz [3504]. Help [884]. Helping [2030]. Helps [382, 2752]. Hereditarily [1405]. Hereditary [1866, 1973]. Heritability [3105]. Hermes [784]. Hermes/ST [784]. Heterogeneity [3588]. Heterogeneous [3482, 2859, 1597, 1684, 3498, 1815, 944, 441, 59, 3584, 1906]. Heuristic [2691, 2682, 2428, 3142, 2260, 964]. Heuristics [2257, 2259, 3252, 3254]. Hidden [3012, 298]. Hiding [2525, 2231, 238, 239]. Hierarchic [79, 1337]. Hierarchical [2997, 2574, 1495, 2604, 2127, 1831, 1854, 1580, 606, 1127, 3436, 3437, 2575, 2568, 1620, 2452, 3134, 738, 1886, 2589]. Hierarchies [453, 1989, 1375, 3376, 3255]. Hierarchy [960, 2021, 868, 1931, 320, 2825, 692, 351, 2981, 2377, 323]. High [951, 2465, 529, 829, 664, 2569, 3412, 3581, 2525, 2268, 3494, 912, 2345, 522, 2607, 3207, 855, 389, 52, 1539, 946, 856, 2646, 3505, 854, 927, 942, 2482, 516, 2410, 922, 932, 2855, 2438, 3515, 2396, 2942, 1548, 936, 928, 3422, 929, 3528, 2572, 933, 626, 1495, 632, 806, 1040, 994, 993, 996]. High-Accuracy [3528]. High-Assurance [3422]. High-Bandwidth [2410]. High-Level [2525, 2345, 522, 389, 516, 2855, 1548, 626]. High-Performance [951, 3581, 3494, 2607, 946, 856, 927, 942, 2482, 922, 928, 933, 806, 994, 993, 996]. High-Speed [2438]. Higher [2019, 2384, 3294, 384, 1522, 2871, 504, 2874, 2875, 100, 1154, 1780, 1778, 1711, 2503, 2883, 683, 374, 2373, 1528, 1994, 1529, 2371, 467, 2372, 1029, 164, 482, 495, 2354, 2369, 1523, 1530, 1415, 1531, 2894, 721, 1462, 1461]. Higher-Order [2384, 384, 1522, 2874, 2875, 1780, 1778, 1711, 2503, 2373, 1528, 1529, 2371, 2372, 1029, 482, 495, 2369, 1523, 1530, 1415, 1531, 1994, 721, 1462, 1461]. Higher-Order-Logic [2871]. Highlighting [2870]. Highly [280, 305]. Highly-Available [305]. Hilbert [3471]. Hill [3125, 3134]. Hill-Climbing [3125]. HIRLAM [3522]. Historical [2757]. Histories [1923, 1584, 3436]. Hoeffding [2040]. HOL [488, 2865, 466, 472, 2867, 2868, 2869, 500, 491, 2877, 2879, 476, 2881, 479, 498, 480, 2886, 481, 2887, 492, 475, 469, 2892, 487, 474, 470, 2894, 2893, 494, 468]. HOL-ML [470]. HOL-UNITY [466]. HOL-Voss [480]. HOL90 [499]. Holomorphic [1171]. homogeneous [2655]. Homomorphic [1358]. Homomorphisms [1406, 2249]. Homophonic [2321]. hopped [807]. Horn [2494, 1383, 2212, 318, 760, 1993, 618]. Hot [2720, 2052]. Hot-Potato [2720, 2052]. Hough [1109, 1050]. House [550]. HPC [941]. HPCN [851, 939]. HPF [995]. HS [2790]. HSP [1014]. HSP-Theorem [1014]. Hull [2087]. Hulls [3549]. Human [2919, 1596, 2746, 1140, 1040]. Human-Computer [2919]. HW [2472]. HW/SW [2472]. Hy0 [1666]. Hybrid [2997, 381, 3012, 1626, 3138, 3079, 3137, 3084, 3005, 1553, 3056, 2357, 3025, 2172, 537, 1625, 3033, 1623, 3142, 3008, 1624, 1638, 596, 3040, 3041, 485, 954, 2126, 819]. hyper [1429]. hyper-linking [1429]. Hyperbolic [2034]. Hypercube [1552, 1217, 1245, 2636, 432]. Hypercubes [1546, 2585, 1202, 1533, 2053, 1534, 1203]. Hypercubic [1738]. hyperelliptic [3448]. HyperGrids [1576]. Hypermedia [1807, 2784, 2516, 2785, 2786, 3441, 3190, 3589, 1752, 3183, 2787]. Hypermedia-Based [3441]. Hyperplane [3106]. Hyperrational [1316]. Hyperrectangle [596]. Hyperspectral [3281]. Hypertext [3443, 1339, 3444]. Hypertext-Based [1339, 3444]. Hypertransition [342]. Hypertree [376]. Hyperuniverses [2295]. hypotheses [1419]. Hypothesis [2203, 279, 3470].

I-CASE [1334]. I/O [2149]. IBC [2512, 2538, 2535]. IBCN [2543]. IBM [2601, 3499, 3504]. Icon [3430]. ICSP [3552]. Id [163]. Idea [3592, 236, 125]. IDEAL [3433]. Ideas [266]. Idempotence [2605]. Identifiability [3297]. Identification [1239, 2233, 2051, 72, 2230, 215, 2101, 823]. Identify [2751]. Identifying [2763, 3310, 601, 3296, 260]. Identities [1359]. Identity [2222, 1756, 624, 1801]. Identity-Based [2222]. IEEE [2397]. if [349, 571]. IF-Filter [571]. II [3479, 2512, 564, 2407, 2918, 1604, 1482, 1027, 3234]. IKUMI [1324]. ILIAS [1554]. illuminant [1089]. illumination [1087, 1098]. Illustrations [576]. ILP [3245]. IM [87]. IM-Systems [87]. Image [380, 2475, 2446, 1868, 2440, 854, 1664, 1881, 1870, 3206, 3281, 1100, 1048, 1097, 1066, 1084, 1101, 1086, 1082, 1054, 1083, 1104, 1121]. image-based [1097]. Images [2707, 1871, 1128, 1136, 1079, 1073, 1102, 1059, 1088, 1089, 1131, 1114, 1060, 1118]. Immediate [3088]. Immune [107]. Impact [1575, 1314, 969]. Impacts [3212]. Impaired [2927, 2928]. Imperative [1349]. Imperfectly [2156]. Implement [2008]. Implementation [2475, 1579, 2523, 1494, 3493, 230, 2656, 3079, 917, 938, 993, 460, 2327, 1209, 346, 3207, 519, 3576, 1929, 963, 2482, 654, 2781, 1571, 1599, 1649, 2587, 920, 2486, 3031, 163, 1295, 3521, 448, 2333, 437, 2636, 205, 2448, 1006, 985, 2765, 3208, 2335, 992, 1798, 1491, 1518]. Implementational [2892]. Implementations [3481, 2344, 2662, 2626, 3500, 874, 649, 2533, 2462]. Implemented [960, 1070]. Implementing [207, 650, 775, 3016, 2360, 2463, 2843, 1016, 3414, 2450, 2402, 2786, 2441, 2852, 636, 2851, 487, 1776, 699, 1426]. Implication [578, 22]. Implications [20, 2523, 2490]. Implicit [869, 786, 1156, 2099]. Implies [1993, 1415]. Imply [2590]. Importance [302]. Important [1369]. Imprecise [2559, 1805]. Imprecise/Incomplete [1805]. Improve [3111, 2162, 3251]. Improved [122, 1533, 1850, 1644, 75, 250, 3476, 2067, 567]. Improvement [1646, 1551, 2968, 3149, 17]. Improving [2343, 1937, 1648, 158, 2792, 1059, 2346, 2258, 3148, 1569]. IMPS [1440]. Impulse [552]. IMS1 [3182]. In-and [3384]. in-Approximations [306]. In-House [550]. In-Line [940]. Inaccuracies [3314]. inaccurate [3217]. Incarnations [3060]. Include [1801]. Including [1298, 546]. Inclusion [3553]. Inclusions [1624]. Incomplete [3309, 3321, 1534, 1962]. Incompressible [869]. Inconsistencies [2971]. Inconsistency [2005, 1987, 3264]. Incorporate [1883]. Incorporating [3195, 2960, 6, 2168, 1999, 1676].

INCOSOFT [2813]. increase [822]. Increased [538]. Increases [574]. Increasing [2926, 2069, 511]. Increment [2819]. Incremental [2234, 580, 2731, 3542, 3267, 267, 1645, 2831, 3476]. Indefinite [3392, 2119]. Independence [2904]. Independent [2569, 1790, 114, 3061, 1850, 92, 2304, 1244, 3322, 3135, 1068, 2637, 821, 637, 1034]. Index [138, 259, 1769, 2251, 2266, 923, 2748]. Index-Server [923]. Indexed [1164]. Indexing [447, 2750, 1452]. INDIA [1687]. Indices [2950]. Indirect [149]. Individual [2515]. Individuals [3097]. Indoor [551, 557, 552, 546]. Induced [2111, 2405]. Inducing [2989]. Induction [2491, 1423, 3288, 1780, 3568, 585, 3302, 591, 2947, 475, 1152, 316, 3124, 2999, 314, 1418, 1419]. Inductive [3290, 3291, 368, 270, 3286, 2417, 3538, 2287, 582, 1018, 3084, 1263, 3324, 279, 1151, 3304, 2991, 277, 3116, 618, 593, 3329, 620, 619, 2802, 621, 1486, 1424, 1483, 1426, 1427]. Industrial [3480, 3338, 3367, 856, 2508, 576, 2767, 987, 2578, 2732, 90]. Industry [26, 48, 2931, 860, 49, 3331, 851, 521, 942, 935, 941, 2753]. Inequalities [2376, 2040]. inequality [636]. Inexact [1265]. Inference [3047, 2993, 2997, 3290, 3291, 368, 2987, 3000, 2982, 2107, 2998, 2187, 1310, 294, 919, 3324, 279, 589, 790, 1782, 1030, 1409, 761, 3329, 1640, 2145, 711, 1485]. Inferences [2984]. Inferring [3377]. Infinite [1382, 1954, 1220, 1360, 1362, 3360, 3314, 3545, 1193, 323, 2139]. Infinite-State [1382]. Infinity [1996]. Influence [26, 1004, 2427]. InfoMall [946]. Informal [3332, 36, 3354, 1232]. Informal-Based [36]. Informality [194]. Informatics [264]. Informatik [2930]. Information [2807, 3294, 3402, 2926, 2857, 1336, 463, 2862, 211, 3424, 2859, 851, 3296, 3311, 3061, 170, 2797, 1168, 3284, 455, 1327, 1215, 414, 1333, 375, 2542, 2777, 2896, 1668, 1334, 2895, 2789, 609, 373, 3574, 930, 2779, 3423, 431, 2097, 409, 1331, 450, 3114, 2858, 1674, 448, 1825, 412, 1341, 928, 444, 2180, 402, 3183, 3328, 2775, 1318, 852, 1655, 3038, 2790, 1962, 3237, 3260, 1500, 1451, 824, 3254, 3239, 1085]. information-preserving [1085]. Infrared [550, 551]. Infrastructure [3200, 942, 930, 1655]. Ingres [1681]. Inherent [3276]. Inheritance [1762, 1755, 639, 625]. Initial [3079, 3084, 1192, 1532]. Initiative [39]. Initiatives [2509]. Initio [883]. Injecting [2566]. Injector [2564]. Injury [1709]. Inlining [3051]. Inner [1950]. Innovation [2732]. Input [270, 3286, 3032]. Input/ Output [270, 3286]. Inputs [1988]. Ins [337]. INSAT [3027]. INSAT-2 [3027]. Inserting [2091]. Insertion [1364]. Inside-Outside [2990]. Insight [1005]. Inspection [2018]. Instance [609]. Instance-Based [609]. Instantiated [440]. institution [642]. Institutions [627]. Instruction [3049, 670, 669, 1006, 164, 2893]. Instruction-Set [2893]. Instructions [184, 653]. Instructor [21]. Instructor-Led [21]. Instruments [21]. Insufficiency [607]. Integer [1937, 2634, 1914, 3109, 2040, 3452, 3454]. Integers [94, 2675, 1199, 1444]. Integral [1691]. Integrated [2171, 3182, 2550, 3424, 2507, 844, 2811, 2183, 3166, 2426, 2813, 1082, 2903, 3001, 3578]. Integrated-Company-Software [2813]. Integrating [2362, 2163, 150, 2669, 460, 2803, 2948, 1821, 3273, 2179, 1891, 528, 2856, 2746, 774, 1035]. Integration [2807, 3174, 570, 1336, 2863, 1042, 3500, 584, 2511, 759, 414, 1814, 1815, 2423, 1885, 1808, 81, 613, 3590, 3275, 1817, 2951, 1340, 3254]. Integrity [855, 3385, 2804, 2792, 217, 1901]. Intel [829, 1009]. Intellectual [415]. Intelligence [2003, 1961]. Intelligent [1012, 1669, 2859, 885, 1687, 71, 2781, 2896, 2751, 2534, 2541, 2783, 2753, 1489, 3215, 3234]. Intensional [602]. Intensity [912, 1104]. Intensive [1687, 913]. Intentions [1782, 1033]. Inter [2523, 2843, 522, 2809, 1905]. Inter-Client [2843]. Inter-Domain [2523]. Inter-Modular [522]. inter-operability [1905]. Inter-Operation [2809]. Interact [2356, 193]. Interact/Intermediate [193]. Interacting [3131]. Interaction [746, 1519, 256, 2539, 1964, 2674, 1190, 2976, 1793, 1566, 771, 2113]. Interaction-Based [1566]. Interactions [780, 3098, 3278]. Interactive [3181, 493, 2030, 422, 3534, 224, 223, 424, 247, 507, 3585, 584, 1824, 219, 2629, 115, 940, 1483]. Interconnect [2464]. Interconnecting [506]. Interconnection [1543, 3184, 2680, 1215, 2273, 1614, 1535, 2584, 1726, 1601, 2596]. Interconnections [1553]. Interface [2514, 3282, 2902, 2524, 2608, 1767, 486, 3284, 2742, 2918, 2855, 300, 2477, 2609, 166, 3425, 1659, 705, 1902, 1472]. Interfaces [664, 2513, 3432, 3435, 2897, 2542, 2916, 3437, 3434, 70, 2898]. Interfacing [2112, 2919]. Interference [3424, 541, 547, 540, 3401, 1951, 813]. Interleavers [536]. interleaving [2115]. Interlocking [3364]. Internet [2406]. Internetworking [3189]. Internetworks [2407]. Interoperability [1828]. Interoperable [2800]. Interpolation [2166, 1411, 3456]. Interpret [612]. Interpretation [2665, 1177, 1524, 3043, 2879, 3053, 654, 84, 3065, 2764, 2168, 1411, 739, 1139]. Interpretation-Driven [84, 3043]. interpretation-tree [1139]. Interpreter [3539, 655]. Interpreters [657, 2901, 658]. Interpreting [1922]. Interprocedural [667, 171, 668]. Interprocessor [949]. interrupted [821]. Interrupts [767]. Intersecting [1856]. Intersection [763, 762, 1849, 2086]. Intersections [2089].

Intersymbol [547, 1951, 813]. Intertwine [1730]. Interval [1302, 1968, 1573, 1976, 2679, 2878, 1919, 1200, 2739, 3514, 1938, 2079, 2979, 849, 1493]. Interval-Based [1302]. Interval-Semantic [2878]. Intervals [1968, 1977]. Intra [2562]. Intra-Gate [2562]. Intrinsic [1124]. Introducing [52, 47, 1598]. Introduction [3587, 3538, 1185, 2252]. intrusive [174]. Intuitionistic [1715, 1987, 734, 1438]. Intuitive [2274]. Invalidation [1564]. Invariance [1888, 3014, 1882]. Invariant [1353, 375, 1877, 1884, 1129, 1890]. Invariants [1636, 2142, 1873, 3562, 1871, 1887, 1891, 1131, 1886, 1875, 3009, 1499, 1507, 1446, 1142, 1037]. Invent [294]. Invention [593]. Inverse [2207]. Inversion [1259]. Inverting [578, 2633]. Investigating [199, 976]. Investigation [3095, 1155]. Invited [2123]. Invocation [779]. Involving [2712, 1849]. IP [3583]. irreducible [800]. Irredundance [2079]. Irregular [2760]. ISA [453]. ISABEL [3205]. Isabelle [505, 1940]. ISDN [3184]. Iserver [968]. Iserver-Occam [968]. Isn’t [305]. isogeny [3451]. Isolation [3112]. Isomorphic [603]. ISPW [205]. ISPW-6 [205]. Issues [16, 202, 520, 2001, 2840, 939, 3506, 3381, 3439, 66, 145, 2892, 2999, 699]. ISTL [1908]. Items [799]. Iterated [249, 1960, 1199]. Iteration [1167, 3069, 1361, 3063, 697]. Iterative [2419, 37, 2635, 3067, 1603, 1061, 814]. ITU [3587]. ITU-T [3587].

Jacobian [3469]. Jacobians [3448]. Jeep [1351]. Jitter [2403]. JNNIE [953]. Job [2343, 2479]. Job-Shop [2343]. Jobs [2278]. Join [459, 457, 458]. Joint [543, 2189, 240, 535]. Jordan [1590]. Juggling [1196]. Junction [1052]. Jungle [387]. Jupiter [1905]. Juricas [2766]. Just [14]. Justification [2958].

K-64 [1274]. k-TTI [2995]. k0t0h [1834]. KAML [576]. Karlsruhe [2927]. Karlsruhe/Germany [2927]. KARO [3166]. Karp [349]. Karp-Levin [349]. KBS [1893]. Keeping [2457]. KEIM [1480]. Kernel [2395, 2686, 1321, 291, 3178, 250, 2659, 1861]. Kernighan [964]. Kernighan-Lin [964]. Key [1941, 2092, 2926, 256, 237, 96, 3404, 128, 132, 3405, 257, 2220, 137, 1297, 1290, 1113]. key-points [1113]. Keynote [1]. Keys [127, 2241, 236, 2230, 2101]. Keystream [1285]. Khufu [2245]. KIDS [3344]. KIDS/ VDM [3344]. Killer [38]. Kind [1791]. Kinds [292, 91]. kinetic [1134]. Kiosk [3444, 2936]. kit [626]. KITP [1485]. KITP-93 [1485]. KL1 [2327]. Kleene [1165, 2313, 1719, 1441, 2943]. Kleisli [1163]. Knapsack [120, 3225]. Know [1033]. Know-How [1033]. knowing [3259]. Knowledge [3174, 2153, 382, 3163, 2163, 597, 1911, 580, 931, 3170, 598, 3442, 3228, 3161, 2231, 2167, 2110, 224, 3159, 2803, 2111, 3158, 1831, 1270, 2205, 1893, 2093, 1195, 2781, 2769, 2777, 79, 86, 2782, 2172, 2955, 2770, 2008, 577, 3441, 115, 3175, 3427, 221, 2978, 3160, 1335, 2165, 2099, 3165, 91, 413, 2753, 262, 3241, 223, 220, 3250, 3255]. Knowledge-Based [3170, 598, 1831, 86, 2782, 2172, 3441, 3427, 2978, 1335, 91]. Knowledge-Level [2955]. Knowledge-Representation [2803]. Known [222]. Knuth [330]. KoMeT [1475]. Kripke [1717, 2943]. Kripke-Kleene [2943]. Kruskal [793]. Krylov [957, 956]. KSR1 [1575, 3496, 961]. Kuperee [3407].

L [3128]. L-System [3128]. L1 [1862]. L1-Diameter [1862]. L2 [2873]. Label [306]. Labeled [1384, 1356, 1360, 2111, 2114]. Labeling [2828, 1844, 1609]. Labelled [1718]. Laboratory [2483]. Lack [1071].

Lack-of-fit [1071]. Lamarckian [3096]. Lambda [3050, 753, 2870, 1395, 1725, 1393, 1397, 744, 1407, 1997, 789, 2281, 702, 703, 1772, 729, 272, 1460, 2965, 677, 483, 467, 1024]. Lambda-Calculus [1395, 1397, 2281, 1772, 272, 2965, 702, 729]. lambda-definition [703]. Lambda-free [2870]. lambda-mu-Calculus [677]. Lambda-Prolog [1024]. lambdasigma [789]. lambdasigma-Calculus [789]. Lambert [1095]. Lamda [1221]. LAN [570, 3184, 3146, 553, 882]. LAN-Based [882]. LAN-Message [3146]. Lander [3139]. Landin [1177]. Landmark [2011]. Landmark-Based [2011]. Language [384, 322, 2622, 775, 1761, 1164, 3539, 1178, 740, 1016, 2733, 3003, 519, 1017, 2754, 1935, 2614, 662, 2948, 2797, 2000, 2658, 2785, 1526, 3273, 734, 1648, 2232, 2985, 2109, 1695, 3316, 292, 2621, 1571, 1599, 3320, 2900, 2922, 1554, 2885, 1808, 1809, 290, 1406, 2736, 2333, 658, 523, 1232, 659, 1803, 3157, 737, 3156, 2320, 2757, 517, 2940, 3155, 2099, 518, 2981, 1566, 2761, 2758, 3001, 3553, 3289, 468, 1904, 629, 723, 3254, 1481, 1776]. Language-Inclusion [3553]. Languages [3386, 1300, 381, 472, 368, 2449, 1180, 2284, 1924, 359, 2339, 1175, 993, 2268, 1936, 520, 1305, 1719, 1689, 3310, 601, 3298, 507, 682, 2301, 1721, 730, 2302, 2357, 3315, 274, 733, 3318, 294, 3319, 3322, 2348, 2856, 1722, 2728, 145, 693, 3327, 2991, 985, 2983, 2354, 3035, 1378, 3329, 776, 1764, 2145, 2123, 641]. Large [1011, 979, 207, 2814, 860, 2668, 888, 869, 1339, 2688, 862, 3285, 903, 410, 300, 2685, 459, 2097, 873, 464, 1674, 2259, 2845, 972, 3524, 924, 413, 3448, 1893, 1073, 3473]. Large-Scale [860, 862, 972, 2814, 2097]. Largest [2060, 3546]. Late [2294]. Latency [2051]. LaTeR [3237]. Lattice [908, 878, 3493, 3447, 861, 2312, 1978, 3463, 3461]. Lattice-Gas [861]. law [1095, 3426]. laws [3246]. Layer [3188, 3187, 3584]. Layered [3026, 2504, 267, 2008, 2572]. Layering [3023]. Layers [2513, 2010]. Laying [2678]. Layout [2410, 2427, 2426]. Lazy [2350, 387, 711, 2349, 712, 1407, 1419]. LCF [2865]. LCS [700]. LDL [657]. Lead [2453]. Lead-Time [2453]. Lean [1477]. leanTAP [1477]. Learn [281]. Learnability [3307, 3297, 3298, 589]. Learnable [2981]. Learned [207]. Learner [3150, 583]. Learners [281, 2981]. Learning [3303, 2997, 3174, 2171, 3306, 3294, 1100, 3440, 2414, 420, 3200, 241, 580, 3247, 3138, 581, 3308, 3295, 598, 1352, 2992, 3087, 3309, 288, 582, 3084, 2974, 3086, 315, 3126, 3122, 1307, 602, 3123, 2299, 3133, 584, 3312, 2980, 283, 3315, 274, 604, 3316, 292, 3318, 606, 607, 3319, 313, 3320, 3322, 609, 610, 2172, 3323, 1369, 3252, 3253, 264, 612, 13, 267, 3132, 3325, 307, 2154, 577, 3441, 3002, 3085, 3089, 3142, 613, 3175, 3327, 2991, 3116, 2983, 309, 3305, 617, 594, 312, 3328, 303, 2169, 2802, 286, 621, 2420, 262, 311, 3289, 2029, 3248, 3249]. Learnt [293]. Least [1720, 1999, 1110]. Lecture [2123]. Led [21]. Left [1368, 3464]. left-shift [3464]. Legal [2765, 2766]. Lemma [791]. Length [2719, 1250, 370, 2233, 798, 1292, 1977, 3134, 1297, 2833, 1129]. Lessons [207, 37, 293, 2577]. Letter [1737]. Letters [1248]. Level [2574, 2696, 2153, 2827, 581, 2525, 2345, 522, 2608, 389, 2878, 516, 84, 2855, 3191, 2955, 2396, 1548, 3175, 164, 3160, 2890, 626, 1495, 632, 1115, 1468, 2564, 1509, 1758, 3224]. Leveling [981]. Levels [2917, 52, 503]. Levin [349]. Lexicon [2986]. LFSR [2227]. LFSR-Based [2227]. Liability [3413]. Liberation [1547]. Libraries [2664, 3494, 3505, 3510, 3215, 3244, 1439]. Library [494, 1433, 970]. Life [2539, 2222, 1632]. Life-Span [2222]. Lifschitz [785]. Lifting [1163]. light [630, 631]. Like [83, 3097, 2084, 2841, 597, 2320]. Likelihood [547, 1066, 1038]. limit [2139]. Limitation [112]. Limited [3558, 1611, 652]. Limited-Backtrack [3558]. Limiting [3515]. Limits [2462, 1668, 3565, 2124]. Lin [964]. Linda [2662]. Lindenmayerian [1361]. Lindstroem [2272]. Line [1835, 1836, 2017, 243, 2708, 2714, 3432, 2450, 1844, 121, 856, 2994, 2706, 1856, 940, 2278, 839, 3306, 991, 1120, 3454]. line-drawing [1120]. Linear [3386, 358, 2867, 1383, 268, 1385, 555, 1944, 1171, 1955, 1626, 1574, 1236, 2703, 2064, 676, 60, 3547, 1843, 1389, 2635, 1213, 1184, 1391, 2625, 104, 1918, 3056, 998, 1145, 1169, 2823, 2216, 1399, 760, 3568, 2016, 2215, 2640, 3514, 2243, 126, 307, 2050, 1406, 2214, 1623, 3520, 3521, 1716, 615, 2983, 2633, 159, 546, 2497, 1172, 2700, 2028, 3009, 3531, 1032, 1062, 1494, 1099, 817]. Linear-Time [2703, 2050, 2823]. linearity [718]. Lines [2089]. LinguControl [2908]. Linguistic [2864, 2097]. Linguistics [1390]. Link [2711, 2406, 1223, 553, 2748, 531, 2903, 574, 561]. Linking [481, 1429]. Links [1788]. Liquid [2974, 894]. Lisp [999]. List [2280, 2064, 1844]. Lists [313, 711]. Literate [3356]. lived [2824]. Liveness [1701, 1507, 490, 2135]. LL [660]. Load [3490, 971, 2686, 2588, 982, 1589, 2685, 1612, 1556, 2687, 1613, 981]. Load-Balanced [1589]. Load-Balancing [3490]. Load-Prediction [2687]. Local [419, 697, 2828, 2837, 2971, 1256, 3161, 2122, 3125, 326, 2366, 3376, 1712, 1112, 1116, 1101, 3114, 1065]. Locales [1168]. Localities [3550]. Locality [157, 1700, 158, 2615, 748, 961, 1863, 2835]. Localization [1330]. Locally [2496, 3375, 3319, 1127]. Located [1392]. Location [3546, 2643, 1758, 2709]. Locking [447, 446]. Locomorphism [794]. Loef [1224]. Log [2620, 1060]. log-polar [1060]. Logarithmic [1197, 1738]. Logarithms [229, 2219, 2238, 3448, 3477]. Logic [3045, 2192, 1014, 1383, 268, 1384, 1385, 1784, 1909, 2378, 2362, 319, 1302, 2492, 1171, 1911, 2417, 112, 1715, 2202, 675, 2339, 2871, 504, 2874, 2875, 489, 1830, 1016, 889, 3060, 1790, 2470, 1637, 1783, 2199, 2425, 3058, 1389, 2957, 1932, 1390, 1792, 2005, 2418, 1184, 1018, 2342, 1929, 1910, 2340, 1391, 345, 1771, 1987, 1915, 2186, 1306, 321, 54, 2358, 2206, 602, 178, 1937, 2469, 3044, 647, 785, 1222, 1396, 3381, 734, 1021, 3397, 2211, 654, 2883, 2109, 1399, 760, 2462, 2458, 1022, 3563, 2386, 1402, 1919, 2943, 2886, 3028, 3540, 3219, 2962, 1789, 867]. Logic [1404, 2348, 2502, 2506, 1933, 1938, 2376, 2325, 2488, 1930, 1025, 3032, 1593, 2002, 1269, 1370, 1406, 1753, 133, 2275, 1934, 2806, 1785, 1267, 467, 1028, 1716, 1548, 2638, 2209, 1782, 3380, 1271, 1029, 2562, 3304, 2201, 2427, 2359, 1409, 2416, 482, 401, 2375, 440, 618, 593, 620, 619, 2498, 2464, 2461, 1412, 2377, 1414, 1916, 2497, 2894, 1172, 91, 792, 1743, 2191, 1317, 2194, 1926, 262, 2012, 2959, 468, 1032, 1897, 3236, 699, 3227, 3259, 706, 1430, 707, 3263, 3266, 3267, 3268, 1441, 3243, 3218, 637, 723, 1450, 134]. Logic-Based [3397]. logic-independent [637]. Logical [1779, 2175, 3293, 3015, 1272, 2984, 1777, 2013, 1991, 2200, 1166, 1193, 3276, 3260, 707, 3249, 720, 2123]. Logics [382, 2207, 2197, 1221, 1694, 1989, 1261, 320, 2952, 1315, 1411, 3039, 3220, 1467, 1484, 3271, 2382]. LOGSPACE [2287, 1728]. LOLITA [2864]. Long [2987, 3110, 2634, 33, 2824]. Long-lived [2824]. Long-Range [2987]. Longest [370, 1703, 1400, 2276]. Look [1763, 1650, 3353, 3154, 1764]. Look-Up [1763, 1764]. Lookahead [2280]. Lookup [2458]. Loop [1607, 158, 988, 162, 159]. Loops [150, 160, 2670, 652]. Loosely [3144]. Losses [552]. Lossy [1242]. LOTOS [3336]. Low [2693, 2456, 219, 84, 2033, 806, 431, 1006, 2830, 1115]. Low-complexity [806]. Low-Cost [1006]. Low-Density [2033]. Low-Level [84]. Lower [1381, 2724, 1733, 2698, 2218, 1844, 2292, 3535, 678, 1847, 1735, 2302, 1738, 2261, 432, 3549, 2028]. Lowest [1860]. Lowness [2082]. LSE [2493]. LT [1863]. LT-RAM [1863]. LTL [1650]. Lucifer [233]. Lukasiewicz [3218]. LUT [2432]. LUT-Based [2432].

M [3138, 1293, 3474, 540]. M-ary [540]. M-Sequences [1293]. MAAL [3187]. MAAMAW [1966]. MACAO [3169]. Machine [1279, 3484, 512, 3489, 598, 2988, 2451, 1985, 387, 2470, 2641, 1162, 965, 2883, 1605, 923, 3321, 2657, 610, 2655, 612, 577, 1606, 464, 3521, 1231, 3337, 999, 2802, 2637, 876, 1081, 1901, 722, 3485, 722]. Machine-Dependent [512]. Machine-Independent [2637]. Machines [155, 2449, 1386, 2673, 1182, 1589, 3505, 2437, 283, 922, 2615, 1536, 2666, 327, 279, 3518, 893, 2393, 1376, 482, 2323, 1208, 1559, 323, 1899]. Macintosh [2485]. Macro [1630]. Made [2479, 228]. MAGE [1593]. Magnetohydrodynamics [957, 910]. Mahalanobis [1050]. Mail [3569, 1676, 3570]. Mail-Based [3570]. Main [2808]. Maintaining [2250, 2832, 1706]. Maintenance [30, 2718, 1342, 2766]. Make [2924, 266, 2675, 3215]. Makes [2980, 275]. Making [2909, 2184, 1965]. Malicious [3306]. Manage [774]. Management [181, 2512, 2814, 2548, 2530, 779, 2532, 2523, 1675, 2796, 1658, 2522, 2843, 3335, 438, 2546, 2531, 19, 2529, 2524, 2608, 3576, 830, 3194, 210, 1323, 2786, 1021, 530, 2392, 2542, 1580, 3277, 2042, 2543, 2739, 2706, 2549, 446, 2540, 1822, 2547, 308, 2757, 1660, 2181, 433, 2403, 2545, 2552, 2752, 947, 924, 3356, 1508]. Manager [830, 2891, 3237]. Managers [1685]. Managing [1892, 2740, 2772, 409, 3573]. Manifestation [2565]. Manifesto [3029]. manifold [1045, 1041]. manifold-solid [1041]. manipulation [1510]. MANNA [999]. Manpower [2085]. manufactured [1125]. Manufacturing [2811, 71, 2752]. map [814, 2246]. Mapped [2670]. Mapping [3127, 1608, 1005, 2430, 2682, 1823, 2433, 2683, 882, 964]. Mappings [2103, 97, 1295, 124, 3326]. Maps [1667, 1169, 2011]. MARC [2485]. Mark [3420]. Mark-and-Sweep [3420]. Market [2923]. Markets [297, 1964]. Marking [2642, 1498]. marking-dependent [1498]. Markings [1883]. Markov [1079, 136, 298, 3478, 1111]. Martin [1224]. Martin-Loef [1224]. MASCOT [907]. Mason [42]. MasPar [982]. Mass [2003, 456]. Massey [802]. Massive [934, 1578]. Massively [2574, 2573, 943, 2673, 3491, 840, 1539, 871, 984, 2661, 2613, 1604, 2627, 923, 1599, 920, 1593, 2160, 2647, 3528, 992, 924]. Massively-Parallel [871]. Master [49]. Matched [572, 542]. matches [1047]. Matching [1254, 1705, 1251, 1250, 2730, 1256, 460, 333, 1239, 2291, 3535, 2729, 1704, 1253, 1887, 1245, 1242, 331, 417, 1702, 1377, 2041, 1255, 1874, 1061, 1139, 1456, 1063]. Mate [3077]. Materials [895]. Mathematical [3355, 3356]. Mathematics [1232, 1432, 1416]. Mating [3088]. Matrices [1834, 3483, 3485, 2626, 2643, 1590, 3524]. Matrix [3487, 958, 438, 3418, 1603, 1554, 2839, 2633, 1255, 1379, 1559, 1147]. Matrix-Vector [1559]. Matroid [2086]. Matroids [102]. Matter [1752]. Mature [50]. Maturity [55]. MAUDE [631]. max [1056, 976]. max-min [1056]. Maxcut [2702]. Maximal [1249, 2049]. Maximizing [158, 1592]. Maximum [419, 358, 379, 2291, 2051, 547, 1856, 1949]. May [2281]. Maze [3089]. Mazes [3142]. Mbit [3182, 553]. Mbit/s [3182, 553]. MCCP [2470]. MD [118, 119]. MD-5 [118, 119]. MD5 [2439]. MDL [591]. Mean [3104, 2682, 2428, 2563]. Meaning [2764, 317]. Meaningful [2458]. Means [2992, 2654, 3056, 410, 2193, 1749, 697, 3463, 1503]. Measure [2282, 579, 2158, 366, 3223]. Measurement [912, 557, 949, 3185, 1125]. Measurements [551]. Measures [421, 2157]. Measuring [558, 1067]. Mechanical [880, 3488, 2875, 3071, 3517, 502, 3073, 1059]. Mechanically [1443, 2888, 1432]. Mechanics [966]. Mechanisation [497]. Mechanism [2345, 788, 2599, 3573, 2556]. Mechanisms [192, 2686, 2782, 3076, 3578, 1962]. Mechanistic [281]. Mechanizable [1424]. mechanization [1441]. Mechanized [508]. Mechanizing [1166, 468]. MEDEA [850]. Media [3577, 2395, 3180, 2413, 2936]. Median [542]. Mediation [1684]. Medical [2799, 465, 2801, 2098, 1059]. Medium [3415, 2813, 1752]. Medium-sized [2813]. MEDTOOL [81]. Meeting [48, 49]. Megaprogramming [52]. Meiko [950]. Meiosis [3099]. Membership [2695, 364, 3309, 2589, 2420]. Memetic [3072]. Memoirs [103]. Memories [2704, 29]. Memory [906, 1562, 2171, 266, 2653, 2868, 2808, 2159, 1561, 151, 2723, 955, 960, 959, 2654, 3499, 1974, 3500, 896, 1553, 1633, 156, 846, 1589, 976, 2469, 2641, 2357, 1854, 1570, 1571, 2620, 1612, 2652, 1610, 1581, 2655, 511, 175, 1538, 977, 1556, 2361, 999, 2649, 2650, 2073, 1949, 1948, 876, 2834]. Memory-Based [2171]. Mental [1315]. Mercury [1587]. mereology [3221]. Merge [150, 3541]. Merging [2992, 2088, 432, 2989]. Mesh [2720, 966, 2682, 2178, 965, 2721]. Mesh-Based [966]. Mesh-Connected [2682]. Meshes [1202, 1210, 333, 1545, 1544, 335, 2319, 1097]. Mesoscale [901]. Message [222, 899, 2653, 2654, 240, 3500, 2439, 967, 428, 2660, 883, 2609, 234, 2599, 969, 882, 1764, 3146]. Message-Digest [2439]. Message-Efficient [240]. Message-Passing-Systems [969].

Messaging [3569]. Meta [2668, 581, 185, 1824, 1758, 1749, 1816, 3224, 1481]. meta-language [1481]. Meta-Level [581, 1758, 3224]. Meta-Parallelizer [2668]. Meta-Process [185]. Metacomputer [972, 3532]. Metadata [3273]. Metal [853]. Metals [894]. metamathematics [1453]. Metaphors [2513]. METATEM [1935]. metavariables [1459]. Meteorology [902]. Mether [977]. Mether-NFS [977]. Method [2574, 957, 1709, 2992, 955, 879, 3000, 1867, 1211, 3504, 2434, 3405, 1238, 609, 254, 1791, 126, 1627, 2008, 172, 417, 1592, 3351, 1764, 915, 926, 2790, 1464, 1421, 253, 1443, 1063, 1059, 1137, 1051, 1517]. Methodical [3010]. Methodological [3169, 3381]. Methodologies [1019, 36, 1343, 2845, 2552]. Methodology [575, 1631, 185, 2425, 156, 832, 189, 161, 3173, 987, 487, 656, 1558, 3216]. Methods [1683, 3486, 3163, 3338, 1763, 3137, 3331, 2635, 2292, 53, 2911, 31, 3334, 3078, 3021, 3397, 2969, 1891, 294, 1744, 2745, 2553, 2968, 3514, 588, 69, 2540, 2275, 613, 1556, 1344, 1258, 70, 311, 956, 954, 3261, 2118, 1132]. Metric [206, 2156, 3028, 374, 1153]. Metrics [2160]. MHEG [3569, 3183]. MHz [572]. MICE [3195]. Micro [2395, 291]. Micro-Kernel [2395, 291]. Microcoded [484]. Microkernel [289]. Micronet [2393]. Microprocessor [1622, 2474, 3027, 485, 3593]. microSR [468]. Midimew [2678]. Migration [2460, 89, 3135]. Milner [736]. MIMD [2665, 960, 959, 1536, 3522, 876]. MIMD/ SIMD [1536]. Min [3373, 2312, 2701, 1056, 2650]. Min- [2312]. MIN-Based [2650]. Min-Conflict [3373]. MIND [1609]. MINE [2621]. Minimal [368, 3296, 1249, 2503, 2054, 378, 2049, 1977, 2603, 2592, 2209, 1462]. Minimisation [1988]. Minimization [2430, 2310, 1247, 177, 1255]. Minimizing [2620, 2042]. Minimum [2695, 2711, 2290, 1197, 1214, 72, 2091, 2039, 2062]. Minimum-Link [2711]. Mining [3285, 1560]. Miss [1563]. mission [3213]. MITRE [3334]. MIX [1385]. Mixed [2664, 1151, 1825, 1464]. MKRP [1476]. ML [2663, 700, 765, 731, 716, 1695, 576, 2887, 3357, 2370, 790, 743, 469, 470, 3057]. ML-ART [743]. MLDD [2572]. Mobile [2514, 575, 529, 543, 538, 564, 562, 2832, 558, 571, 137, 748, 554, 2940, 567, 2012, 566, 810, 811]. Mobile-Executed [575]. Mobility [1651]. Modal [1980, 1384, 1388, 2197, 2186, 1396, 1934, 2359, 1409, 1411, 1645, 1743, 1467]. Modal-calculus [1980]. Modalities [2210]. Mode [3044, 2307, 302, 235]. Model [661, 1382, 2153, 207, 3169, 1643, 184, 899, 1888, 1422, 1386, 1628, 667, 2353, 1715, 2159, 1561, 2744, 3098, 1650, 1984, 3418, 1653, 19, 2987, 3003, 1662, 408, 679, 3158, 2118, 1263, 830, 2784, 1641, 390, 901, 140, 1802, 2006, 3364, 2741, 1094, 3273, 2554, 1630, 1824, 3360, 1827, 557, 2985, 480, 1854, 3274, 2906, 2848, 1634, 1620, 1809, 59, 291, 1632, 2770, 589, 55, 1647, 1806, 1796, 1793, 2174, 2793, 451, 339, 2562, 3175, 1702, 3417, 1638, 1337, 837, 1645, 461, 2989, 1810, 139, 1917, 1586, 1884, 85, 2028, 314, 1898, 1900, 2121, 706, 1500]. model [2122, 1906, 1087, 1490, 1509, 467, 1661]. Model-Based [3158, 830, 2006, 59, 2174, 85, 2118, 1884]. Model-Checking [1386, 480, 1647]. Model-Theoretic [2353, 451]. Modeling [1766, 2558, 1750, 1956, 1811, 486, 835, 2424, 68, 1661, 2651, 1663, 1829, 1812, 2805, 1632, 1751, 1258, 1613, 1550, 1816, 1489, 1034]. Modelling [1572, 183, 202, 3174, 2518, 3170, 2537, 3164, 1330, 1496, 211, 2546, 190, 830, 3167, 845, 201, 1323, 191, 868, 2399, 1338, 3348, 905, 2760, 833, 88, 1508, 3033, 398, 1331, 81, 3341, 188, 187, 1511, 3157, 186, 412, 3156, 200, 2761, 1332, 848, 196, 413, 494, 76, 1518].

Modellversuch [2930]. Models

[1826, 3481, 1797, 1171, 2569, 3491, 2268, 1819, 195, 1830, 1154, 2957, 2644, 3369, 838, 2878, 78, 1198, 846, 80, 1711, 1397, 759, 298, 832, 1399, 903, 89, 1635, 1812, 1545, 1529, 3033, 1823, 1189, 2636, 1804, 3346, 2743, 1341, 1532, 1816, 3107, 1813, 1558, 771, 2279, 3001, 1100, 1421, 2127, 1504, 1102, 1111, 2141, 1086, 1108, 1469]. Modern [1279, 880]. Modernization [1318]. Modes [1287, 1852, 2852, 3419, 1088]. Modification [1346]. Modified [798, 3113]. Modifying [1218]. Modular [3330, 232, 522, 676, 2641, 2392, 1930, 2583, 2372, 3201, 3236, 705, 3459, 634, 1462, 3456]. Modularity [2500, 689, 2383]. Modularization [2955]. modulated [808]. Modulation [555, 539, 535, 544, 1951, 816, 806, 815, 803]. Modulations [554]. Module [2791, 522, 2456, 2482, 2887, 523]. Modules [3481, 1629, 2469, 3151, 1024]. Modulo [3562, 1435]. Modulo-Place-Invariants [3562]. Molecular [3490, 878, 888, 3492, 889, 3518, 893, 897, 886, 894, 887, 3527]. Molecular-Dynamics [893, 897]. Mollusc [1484]. moments [1069]. Monadic [345]. Monitoring [989, 856, 2619, 174, 2621, 2403]. Monogenic [3329]. Monoids [1719, 3310, 343, 2317]. Monolithic [570]. Monopole [3489]. Monotone [1834, 307]. Monotonic [3186, 1844, 3315, 1267, 1316, 2194]. Monotonicity [3318]. Monte [908, 2470, 879, 2654]. Monte-Carlo [2470, 2654]. MooZ [3349]. Morphisms [393, 2728]. Morphology [2980]. MOSS [3210]. Most [1543]. Motion [1101, 1880, 1046, 1047, 1045, 1103, 1049, 1074, 1078, 1048, 1072, 1066, 1107, 1080, 1070, 1084, 1088, 1051, 1132, 1068, 1134, 1105]. Mount [917]. mounted [1130]. movable [819]. Movement [2942]. movements [1107]. Movie [2404]. Movies [2408]. Moving [1867, 2712, 1103]. MP [982, 3520]. MP-1 [982]. MPEG [3585]. MPEG2 [3586]. MPI [967]. MPP [853]. MPP-Systems [853]. MRF [1070]. MSO [2301]. MST [2717]. MT [2970]. Mu [1386, 1645, 679, 677]. mu-Calculi [679]. Mu-Calculus [1386, 1645]. Much [2734]. Multi [555, 2791, 2229, 51, 1763, 2600, 2702, 1272, 1016, 2528, 1956, 1005, 3103, 462, 3211, 1694, 3407, 3439, 3083, 2823, 2482, 1957, 3151, 2008, 1271, 3376, 1965, 1601, 196, 2572, 876, 3232, 1905, 3240, 820, 1787, 2328]. Multi-address [2600]. Multi-Agent [2791, 1956, 3439, 1957, 1271, 1965, 3232, 3240]. Multi-Block [876]. Multi-Channel [555, 820]. Multi-Chip [2482, 3151]. Multi-Constraint [3083]. Multi-Criteria [3083]. multi-database [1905]. Multi-Dimensional [1005, 1601]. Multi-Head [1016]. Multi-Layered [2008, 2572]. Multi-Methods [1763]. Multi-Operator [3211]. Multi-parent [3103]. Multi-Pebble [1694]. Multi-Provider [2528]. Multi-reader [2823]. Multi-secret [2229]. Multi-Service [3407]. Multi-SLD [1787, 2328]. Multi-Sources [1272]. Multi-Threaded [462]. Multi-User [51]. Multi-View [196]. Multi-Way [2702, 3376]. Multi-writer [2823]. Multiagent [1033, 3594]. Multiblock [875, 872]. Multicast [2600, 3207, 3580, 2412]. Multicomputer [991, 835, 2598, 973, 981]. Multicomputers [3483, 978, 983, 1588]. Multiconnectivity [1729]. Multidatabase [1671, 443]. Multidestination [2599]. Multidimensional [3535]. Multidisciplinary [2669]. Multifunctions [768]. Multigrid [873, 961]. Multilevel [3421, 3416, 3420, 1806, 451, 1817, 3422, 806]. Multilingual [2098]. Multimedia [3181, 3197, 3195, 3182, 3592, 3571, 3196, 2401, 3581, 3587, 3588, 3212, 3194, 2397, 2411, 2392, 3210, 3180, 3569, 3191, 3575, 3572, 3204, 3187, 3574, 3199, 3178, 2394, 2396, 2858, 2393, 3190, 3589, 3202, 3209, 3179, 3198, 3591, 1676, 3570, 3185, 3584, 1665, 3578, 1500]. Multimedia-Object [3210]. Multimodal [2915]. Multiparadigm [517]. Multiparty [3579]. Multipath [557, 565]. Multiple [2807, 568, 1237, 2532, 2645, 3060, 1021, 1253, 1338, 2216, 2581, 1960, 1885, 2325, 2488, 613, 2598, 2176, 532, 549, 544, 1044, 542, 1927, 698, 1103, 1133, 1076, 1148, 1052]. Multiple-Access [542]. Multiple-Conclusion [2325, 2488]. Multiple-Symbol [544]. Multiplexing [2604]. Multiplication [3487, 1199, 1559]. Multiplicative [1385, 1389]. Multiplicity [420, 1366]. Multiplier [2224]. multipliers [818]. Multiply [1217]. Multiply-Twisted [1217]. Multipoint [3188, 3581]. Multiprocessing [1553]. Multiprocessor [1562, 908, 650, 151, 1553, 2617, 458]. Multiprocessors [266, 148, 2087, 1587, 1552, 2575, 1581, 1556, 1613, 848, 2649, 2650]. Multiresolution [380]. Multiround [2226]. Multisearching [1546]. Multisecret [226, 227]. Multiset [2826]. Multispace [2043]. Multistage [2584]. Multistrategy [584]. Multistreamed [837]. Multitemporal [1828]. Multithreaded [1580, 1614, 163]. Multitone [565]. Multitransputer [3497]. Multiuser [532, 822]. Multivalued [2822]. Multivariate [2641, 3456]. Multiversion [447, 2860]. Multiway [1727, 2088]. MUSE [2660]. Must [2281]. Mutation [3104, 3148, 3143]. Mutual [504, 475, 3328]. My [3554]. Myths [3338].

N [2351, 2717, 310, 939]. NA48 [911]. Name [497, 688, 1191]. Name-Carrying [497]. Name-Passing [1191]. Naming [2024]. Naor [2226]. Narrowing [2350, 2384, 387, 2349, 2383, 2493, 1425]. Narrowings [2348]. NAS [951]. National [1655, 2895, 930]. Natural [1286, 1808, 1809, 2333, 2940, 2099, 1081]. Nature [183, 317]. Naval [2804]. NavChair [2913]. Navier [861, 877, 876]. Navigating [2912]. Navigation [1049, 195, 1739, 2011, 1064]. NC [2269, 339]. NC-Real [2269]. NC02 [1200]. nCUBE [2617, 2656]. nCube/2 [2656]. nCUBE2 [3496]. Near [549]. Near-Far [549]. Nearest [1834, 1544, 596]. Nearest-Hyperrectangle [596]. Nearest-Neighbor [1544, 596]. Nearly [1839, 3101, 3296]. Necessary [1983]. Need [852]. Needed [756, 1455]. Neededness [3063]. Needs [48, 49, 213, 2511]. Negation [675, 1987, 1308, 1266, 1786, 91]. Negative [3000, 2389, 3313, 2045, 2144]. Negotiation [3580, 1962]. Neighbor [1544, 596]. Neighborhood [1805]. Neighborhood/ Conceptual [1805]. Neighbors [1834]. Neighbourhood [3081]. Nested [160, 652, 461, 1917, 1316]. Net [1558, 1504, 1510]. Nets [843, 3015, 1389, 83, 1184, 679, 1929, 2007, 2297, 1618, 1990, 363, 1584, 1185, 760, 397, 849, 3140, 1493, 626, 1494, 1495, 1497, 1498, 1499, 1489, 1500, 1502, 632, 1503, 1506, 1507, 1490, 696, 1511, 1513, 2130]. NetSim [835]. Network [3181, 1636, 878, 948, 3098, 995, 3414, 2608, 847, 1831, 2948, 862, 2411, 3131, 2533, 607, 2088, 2921, 3572, 1282, 2486, 2391, 2534, 1614, 2541, 2597, 1729, 949, 2584, 2942, 3202, 2025, 3179, 969, 2830, 2181, 1601, 915, 2596, 3584, 2084, 2028, 877, 1070]. Network-Architectures [969]. Networked [3492, 3501, 3183]. Networking [939, 927, 935, 2405, 3582, 928, 929, 933]. Networks [575, 834, 3586, 568, 529, 1543, 3399, 2606, 3141, 305, 931, 2286, 3138, 3017, 3188, 2522, 1958, 3443, 2678, 2829, 1551, 3205, 2290, 2507, 2524, 2605, 425, 986, 2680, 3194, 1687, 3177, 2688, 835, 918, 3133, 1212, 1215, 2399, 2517, 2027, 332, 932, 2789, 3575, 3409, 2609, 2026, 2273, 2036, 534, 2008, 2676, 1535, 137, 2413, 1859, 1196, 1738, 936, 2996, 3583, 2951, 1616, 1690, 2580, 1726, 3143, 567, 2009, 2182, 3584, 3193, 3578, 3553, 2029, 1141, 1489, 1893, 3240, 820]. NeuDB [2951]. NeuDB-System [2951]. Neumann [1605]. Neural [2997, 3141, 3138, 1958, 296, 83, 2007, 918, 3133, 2789, 2486, 2008, 2942, 2951, 1690, 3143, 916, 915, 2009, 2182, 3140, 2028, 2029, 1141, 3254]. Neuro [914, 3139]. Neuro-Control [914]. Neuro-Genetic [3139]. Neuron [2448]. Neuronal [78]. Neutral [3127]. NewLink [2931]. Newspaper [2934]. Newspapers [2936]. NewSpeak [3358]. NewsReader [2934]. Newton [177]. Newton-Cauchy [177]. Next [1670, 2402, 1549, 1672]. Next-Generation [1549]. NFS [977]. NIAM [1812]. Niches [3088]. NIIT [930]. NIKS [2240]. Nine [3413]. NL [2864]. No [372, 1455]. Node [1727, 2657, 918]. NODEN [2879]. Nodes [167, 1856, 819]. Nogood [3373]. Noise [3306, 3307]. Noise-Tolerance [3307]. Noisy [3308, 3111, 3095, 1118]. Non [1234, 35, 1061, 1551, 1395, 3562, 424, 2635, 2625, 3061, 174, 1397, 1022, 3537, 2655, 3384, 788, 3114, 1856, 2705, 1267, 3401, 546, 2194, 2753, 1065, 1062, 1431, 2124, 1434, 1109, 2115, 1099, 3271]. Non-Blocking [1551]. non-classical [3271]. Non-commutative [3537]. Non-Determinism [3384, 2124]. Non-Deterministic [1395, 788]. non-disjoint [1434]. Non-Exploratory [2705]. Non-Extensional [1397]. Non-Functional [35]. Non-homogeneous [2655]. Non-Interactive [424]. Non-Interference [3401]. non-interleaving [2115]. Non-Intersecting [1856]. Non-intrusive [174]. Non-iterative [1061]. Non-Linear [2625, 546, 1062, 1099]. Non-local [3114]. Non-Monotonic [1267, 2194]. Non-Overlapping [1234]. Non-parametric [1065]. non-probabilistic [1109]. Non-Procedural [1022]. non-provable [1431]. Non-Reachability [3562]. Non-Rectangular [2753]. Non-Simultaneous [2635]. Non-strict [3061]. Nonbinary [3075]. Nonblocking [986]. noncoherent [808]. Noncooperative [2946]. Nondeterminism [369]. Nondeterministic [2313, 3024, 685, 636]. Nonexistence [595]. Nonlineaer [1951]. Nonlinear [98, 580, 862, 3503, 3509, 177, 58]. Nonlinearity [107]. Nonlinearly [2243, 218]. Nonmasking [3010]. Nonmonotonic [2208, 2418, 2187, 1314, 1026, 3261]. Nonmonotonicity [2963]. Nonnegative [1379]. Nonnormal [860]. Nonperfect [102]. Nonrecursive [1375]. nonregular [2120]. nonrigid [1088]. Nonsense [934]. Nontraditional [757]. Nonuniform [2082]. Norm [1255]. Normal [1789, 1590, 752, 621, 703, 3468, 1467]. Normalisation [1995, 1997]. Normalization [1218, 1876, 1395, 329, 1520, 1126, 1436]. Normalizing [741]. Normative [1303]. normed [2117]. North [904, 2921]. Norwegian [103]. Notation [1662, 2883, 3337, 1896]. Notational [2900]. Note [3309, 386, 2962]. Notes [769, 2033]. Notion [1401, 3259]. Notions [349, 352, 1427]. Novel [2969, 553, 3524, 2330, 1450]. NP [2020, 357, 415, 2725, 349, 2058, 3470]. NP- [2725]. NP-Complete [2058]. NPCo [3470]. NPCo-NP [3470]. NTH [15]. NTS [1357]. NUBUS [2485]. NUBUS-Expansion [2485]. nuclear [1508, 3254]. Null [1412]. NUMA [1537]. Number [230, 1353, 3314, 873, 1856, 235, 3472, 3479, 2119, 3471, 3453, 3476, 3470, 3466]. Numbers [2821, 3296, 785, 2640, 1296, 1590]. Numerical [3494, 3499, 3500, 3508, 927, 2351, 2423]. NXLib [1595].

o [2717, 2149]. O.S.S. [111]. O2 [2843]. OASIS [643]. Oberon [524, 525, 523]. OBJ3 [2334]. Object [2776, 1346, 780, 1755, 181, 1759, 1868, 2810, 1754, 2208, 779, 2857, 2759, 2847, 3421, 781, 1761, 463, 1178, 1766, 436, 2339, 881, 2843, 1818, 2862, 2744, 1330, 1896, 2842, 3420, 2984, 2863, 1750, 1673, 2733, 3435, 1684, 2754, 1662, 2840, 2555, 784, 769, 453, 1587, 1767, 2780, 901, 462, 1802, 2860, 1747, 1757, 36, 2841, 2734, 2853, 1329, 2109, 3210, 3180, 1891, 2735, 2855, 1815, 3512, 2756, 2848, 405, 1745, 1756, 782, 1885, 2850, 3350, 2864, 2851, 1760, 2844, 1795, 449, 1679, 1765, 1822, 2736, 1348, 1758, 1753, 1672, 2771, 2858, 2861, 1751, 1768]. Object [1796, 437, 1793, 2846, 1749, 2793, 523, 778, 1803, 1886, 2845, 412, 2746, 407, 517, 461, 433, 3208, 203, 2752, 3422, 2762, 947, 1328, 1801, 2420, 745, 1897, 1904, 3227, 1035, 622, 1090, 1902, 642, 1037]. Object- [1761]. Object-Affinity [1587]. Object-Based [2842, 3422]. Object-Centered [2735]. Object-Groups [782]. Object-Identity [1801]. Object-Orientation [3350]. Object-Oriented [2776, 1346, 1755, 2810, 2857, 2759, 2847, 3421, 781, 1178, 1766, 436, 881, 2843, 1818, 2862, 1330, 3420, 2984, 2863, 1750, 1673, 1684, 2754, 1662, 2840, 2555, 784, 769, 453, 1767, 2780, 901, 1757, 36, 2841, 2734, 2853, 1329, 2855, 3512, 2756, 2848, 405, 1745, 782, 2850, 2864, 1760, 2844, 1795, 449, 1679, 1822, 1348, 1753, 1672, 2858, 2861, 1751, 2846, 1749, 2793, 523, 1803, 2845, 2746, 3208, 2752, 2762, 947, 745, 2109, 1897, 1904, 1896, 1902]. Objectives [3425]. Objects [742, 1522, 3421, 2035, 456, 1220, 1794, 1910, 1883, 772, 2712, 1849, 336, 773, 2651, 2760, 2652, 2709, 2856, 1793, 1577, 743, 3389, 1740, 1752, 1746, 1882, 695, 1103, 630, 1094, 1120, 1903, 1063, 1902, 1119, 1134, 727, 1038]. Oblivious [3408, 585]. observational [2128]. Observations [3439]. Observed [2161]. Observers [1626, 243, 122]. Observing [770]. Obstacles [2711, 3546]. Obtain [495]. Obtained [1881, 1120]. Occam [968, 3140]. occluding [1041]. Occlusion [1072, 1073, 1076]. occlusions [1096, 1143]. Occurrence [1493, 397]. Ocean [903]. OCR [2994]. Odd [2833]. Odd-length [2833]. ODEs [1375]. Off [243, 121, 2994, 2464, 839, 575, 2133, 2413]. Off-Line [243, 121, 2994, 839]. Office [2863, 2743]. Offs [428]. OffScreen [2906]. Offset [2596]. Oikos [181]. Oil [906]. Olden [141]. Oligopolistic [1964]. Olympic [828]. omega [2057, 2310, 1616]. omega-Automata [2310]. OMNI [2552]. OMNI-Point [2552]. OMNIS [2788]. OMNIS/Myriad [2788]. OMT [2849]. OMT-Classes [2849]. On-Board [3027, 2660]. On-Chip [571]. On-Line [1835, 1836, 2714, 1844, 856, 2706, 2278, 3306, 991]. On-The-Fly [167, 1647, 3048]. Once [585]. One [222, 116, 1920, 2221, 1355, 318, 3535, 2293, 2059, 1284, 1953, 303, 1290, 2246, 1432, 1136]. One-Dimensional [1953]. One-Key [1290]. One-Sided [303]. One-Way [116, 2221, 1355, 2059, 1284, 2246]. Ongoing [2935]. Online [2863, 842, 1582]. Only [2246, 1561, 3259]. onto [1130, 1536, 1581, 2433]. Ontologies [3166]. Ontology [3165]. Ontology-Based [3165]. OO [1821, 1823, 2540]. OODB [2798]. OODBMS [2852]. OOPS [2853]. OPAL [519]. Open [3479, 3399, 931, 2786, 1363, 942, 3204, 3409, 1722, 827, 3206, 947, 404, 418, 1892]. Opening [2241]. OpenMol [885]. operability [1905]. Operating [2686, 2626, 2435, 952, 2391, 2583, 2409]. Operation [435, 1287, 2809, 671, 3371, 343, 3428, 3419, 2080]. Operation-Triggering [671]. Operational [337, 1178, 730, 2131, 1228, 732, 774, 3260, 2144]. Operations [2344, 2046, 2187, 2655, 142]. Operative [3086]. Operator [3211, 1999, 3113, 3541, 3220, 3252]. Operators [768, 2431, 3102, 3084, 998, 3117, 2337, 3552, 3143, 595]. Opportunities [851]. optic [1060]. Optical [2522, 1553, 2018, 552, 553, 1043, 1058, 1064, 1107, 1099, 1044]. Optics [551]. Optimal [2065, 1705, 2672, 2559, 1543, 1837, 1839, 2714, 2737, 2711, 333, 1197, 2270, 2973, 1846, 2679, 3194, 1211, 2729, 2667, 1988, 1704, 1851, 1852, 1244, 2305, 2042, 353, 355, 2311, 267, 3108, 2676, 1544, 1708, 1203, 93, 2055, 2835, 1208, 3468, 3153, 1056, 545, 2713]. Optimally [1307]. Optimisation [3083, 3077]. Optimisations [960, 669]. Optimised [1281]. Optimising [3091]. Optimistic [2860]. Optimization [3136, 512, 522, 3003, 2342, 3043, 2688, 862, 2043, 173, 2809, 647, 3147, 3397, 2357, 3373, 2458, 2798, 459, 3112, 3120, 3152, 3557, 461, 2677, 3096, 1559, 1457]. Optimizations [785, 2615]. Optimize [153]. Optimized [879, 2437, 2610, 2465]. Optimizer [460]. Optimizing [3138, 456, 648, 984, 663, 643, 2671]. Optimizing/Parallelizing [663]. OPTIMUL [1553]. Optimum [1202, 2822]. Optoelectronic [2480, 2596]. OPUS [3350]. OR-Parallel [1570]. OR-SML [2797]. Oracles [348, 1992]. Order [742, 2384, 3294, 384, 1522, 1302, 2492, 2871, 1976, 504, 2874, 2875, 100, 1154, 1780, 1725, 2111, 411, 345, 321, 1778, 1711, 1724, 2503, 2883, 683, 1919, 1931, 2373, 1404, 1528, 1529, 2371, 688, 1647, 467, 2372, 1029, 164, 482, 495, 1532, 2354, 2320, 2369, 142, 1523, 1530, 1415, 1531, 2894, 1194, 695, 702, 3247, 1429, 2509, 1505, 3473, 721, 1994, 1462, 1461, 641, 1450, 1060, 1427]. Order-Sorted [1724]. Ordered [1447, 1967, 2081, 2311, 1529, 2080, 2368, 1460]. Ordering [672, 2199, 1641, 2998, 793, 326, 2390, 330, 1466, 3252]. orderings [1447, 1449, 1423, 1418]. Orders [1833, 2219, 1968, 1976, 2032, 1974, 758, 3468]. Ordinary [3519]. ORES [2758]. Organic [2185]. Organisation [946]. Organization [649, 2751, 2743, 3232]. Organize [2666]. Organizing [2684]. Orientation [3350, 1971, 2833, 1110, 1052, 1114, 1054]. Oriented [2776, 1346, 1755, 11, 202, 2810, 2857, 2759, 2847, 3421, 781, 1761, 1178, 1766, 1781, 436, 2339, 881, 2843, 1818, 2862, 2744, 1330, 3420, 2984, 2863, 1750, 1673, 2733, 2966, 3435, 1684, 2754, 1662, 14, 2840, 2555, 784, 2618, 769, 453, 1767, 3018, 2780, 901, 12, 1757, 2445, 36, 2841, 2734, 2853, 1329, 1259, 2855, 3512, 2756, 2848, 405, 1745, 782, 1274, 2850, 2864, 1760, 2844, 1795, 449, 1679, 1822, 1348, 1753, 1672, 2858, 2861, 1751, 2846, 1749, 2793, 523, 1803, 2845, 2746, 1268, 3208, 203, 2752, 2762, 947, 3280, 745, 1897, 1904, 1896]. oriented [3228, 2109, 3253, 1902, 1895]. Origin [1523]. Originate [515]. Orthogonal [2691, 2236, 1400, 329, 540, 2072]. Orthogonalisation [2632]. Orthogonality [1415]. orthography [1078]. OSA [2550]. OSI [2542]. OSIMIS [2525]. OSTIA [2988]. Other [377, 2306, 2921, 2727, 2154]. OTTER [1421, 1471]. Outer [1950]. Output [2907]. Outputs [2569]. Overall [2042]. Overflow [819]. Overhead [2651]. Overlap [1350, 2673, 2312]. Overlap-Free [1350]. Overlapping [1234]. Overloading [741, 625, 1460]. Overview [1007, 155, 645, 1934]. Oz [3382].

p [1379, 1501, 1618, 163, 350]. P-RISC [163]. P-Selective [350]. P/T [1501, 1618]. P02 [1890]. P02-Invariant [1890]. P7519 [1010]. PAC [267, 589, 3328]. PAC-Learnability [589]. PAC-Learning [3328]. Pace [2457]. Package [3189, 859, 912, 552, 2477, 475]. Packaging [2604]. Packet [2609, 2405, 812, 820]. Packing [2713]. PACLIB [2659]. PAFEC [859]. PAFEC-FE [859]. Page [1835]. Paged [1974]. PAGEIN [939]. Paired [2986]. Pairs [1881, 2315, 1465]. Pairwise [1394, 2053]. Palace [1010]. Palindromes [356]. PAM [2450]. Pan [2545]. Pan-European [2545]. Panel [24, 3398]. pans [1465]. PAPS [841]. Paraconsistency [3264]. Paradigm [3426, 958, 3082, 203, 709, 622, 1102]. Paradigms [2639, 197, 3272, 517, 771, 1897]. Paradox [2962]. Paragon [1009, 981]. ParaGraph [978]. parallax [1080]. Parallel [1572, 2639, 2574, 155, 3482, 356, 2044, 3483, 3136, 1543, 3486, 860, 1564, 1542, 1839, 2573, 2808, 1583, 943, 957, 931, 1608, 859, 2673, 2622, 888, 3491, 3493, 3492, 1573, 2121, 150, 2664, 1594, 840, 917, 2046, 2645, 3137, 889, 1001, 921, 1355, 1209, 1841, 1575, 3547, 2607, 869, 2717, 354, 1206, 423, 2291, 2630, 879, 971, 1197, 2654, 855, 3497, 2686, 847, 1578, 838, 1539, 2626, 2588, 982, 2614, 3500, 925, 900, 2729, 1555, 963, 1707, 874, 1986, 2646, 862, 1198, 2726, 3503, 170, 1848, 871, 1549, 2658, 3506, 984, 1704, 2682, 854, 3381, 2661]. Parallel [1596, 2641, 2224, 953, 2071, 868, 751, 2613, 2068, 3594, 875, 2069, 1736, 1609, 832, 1854, 3082, 1244, 2305, 2045, 2306, 1604, 922, 905, 892, 3511, 2627, 962, 932, 944, 923, 1571, 1599, 2615, 3512, 3135, 952, 1649, 2640, 1612, 883, 1600, 858, 1615, 1554, 920, 3514, 355, 327, 833, 975, 3097, 2070, 2655, 2616, 2160, 165, 2452, 897, 2643, 3519, 886, 1606, 873, 464, 1582, 3521, 1003, 143, 910, 990, 145, 1547, 1548, 945, 827, 339, 969, 659, 894, 1590, 417, 2022, 176, 980, 1860, 1006, 1291, 1862, 1613, 985, 3549, 2278, 1550, 999, 904]. Parallel [2647, 1595, 2633, 972, 2648, 3155, 3524, 890, 870, 179, 3527, 1726, 1874, 3528, 992, 916, 915, 1199, 1558, 841, 924, 3529, 926, 1865, 2637, 3531, 877, 872, 956, 1491, 1900, 2122, 1436, 1899, 3233, 1901, 722, 1570, 1569]. Parallel-Algorithms [417]. Parallelisation [2638]. Parallelism [2663, 934, 148, 2669, 1578, 2625, 2809, 158, 395, 175, 1567, 164, 1565, 3061, 2133, 1450]. Parallelization [911, 1010, 966, 2635, 901, 896, 173, 903, 353, 961, 3522, 902, 863, 907, 876]. Parallelized [861, 665]. Parallelizer [2668]. Parallelizing [1011, 1000]. Parameter [84, 1103]. Parameterfree [1085]. Parameterized [1235, 2491, 357, 1390, 1253, 2891, 1486]. Parameters [1354, 3327, 1379, 1054]. Parametric [3047, 1846, 1241, 724, 1065]. parametrization [3459]. Paramodulation [2964]. paraperspective [1051]. PARASPAR [3531]. parcPlan [3233]. PARCS [925]. parent [3103]. Parix [968]. parLisp [999]. Parnas [478]. PARSE [2646]. Parsytec [968]. Part [1013, 43, 1027]. Partial [1521, 2032, 2231, 3262, 2005, 1018, 2336, 3051, 1848, 796, 393, 1505, 1931, 1368, 2337, 1647, 1407, 3524, 2642, 1865, 1194, 1949, 1948, 2146, 1483, 633, 1441, 726]. Partially [3310, 1967, 603, 2927, 343, 2249, 440]. Partially-Commutative [3310]. Particle [3129]. Particle-Based [3129]. Partition [1649, 2817, 2321]. Partitioned [3507]. Partitioning [2702, 666, 3154, 1538, 873]. Partitions [1838, 1641, 77, 2312]. Parts [1791]. PASS [3132]. Passing [899, 2653, 2654, 1587, 3500, 967, 2660, 883, 3146, 2609, 492, 2599, 1191, 969, 882, 1764, 2143]. Passive [65, 2768, 2298, 2771]. Password [3405]. Past [3554, 320, 278, 2094]. Patching [3161]. Path [380, 2711, 2379, 2460, 3542, 3110, 1852, 1619, 540, 2012, 1452]. Path-Controlled [380]. path-indexing [1452]. Paths [2038, 2306, 2307, 1200, 355, 1729, 2700]. Pattern [368, 333, 1355, 1198, 274, 1244, 866, 2925, 3132, 1602, 1702, 2041, 1864, 2617, 3289]. Pattern-Matching [1702]. Pattern-Types [2925]. Patterns [1748, 1005, 1253, 3131, 3313, 2504, 260, 369, 1749, 2372, 2176]. Pay [271]. Payment [3412, 3411]. PBL [2169]. PC [2477, 2616]. PCB [2262]. PCF [728, 1164, 2279]. PCKS [722]. PCKS-machine [722]. PDEs [3513]. PDS [423]. Pe [813]. Pe-criterion [813]. Pebble [1694]. Pebbling [1692]. Peers [1487]. Pegase [940]. PEI [2623]. [185]. 1 [3225]. 2 [2905]. 4-Approximation [2694]. Analogy [2152]. Bigram [3001]. CDMA [812]. CM-5E [3484, 3521]. Communication [2664]. Computation [2673]. Conceptual [1805]. Fold [2505, 2360, 2347]. Germany [2927]. Hardware [795]. HPF [152]. Incomplete [1805]. Intermediate [193]. Local [3489]. Mellor [3369]. Myriad [2788]. Navier [875]. NJ [648]. Non-stochastic [3012]. Nothing [113]. Output [270, 3286]. Parallelizing [663]. Performance [2464]. Relational [1675]. Resource [2544]. RT [3369]. Run-Time [165]. s [3182, 553]. S-15 [1009]. SIMD [1536]. Software [2471]. ST [784]. SW [2472]. T [1501, 1618]. TDMA [543, 529]. transition [1494, 1503, 1490]. Tyxt [750]. VDM [3344]. WSI [2579]. People [2511, 2896, 2898, 2932]. PEPA [845]. Peplom0d [2733]. perceptive [1141]. Perceptual [312, 1057]. Perfect [98, 2076, 246, 221, 220]. Perform [3124]. Performability [2558, 830]. Performance [951, 2475, 559, 829, 664, 948, 3581, 828, 2268, 460, 1819, 2594, 3494, 1646, 2470, 2607, 562, 539, 2593, 810, 838, 2618, 1139, 946, 856, 845, 835, 1549, 3505, 927, 1596, 530, 942, 3154, 2482, 2404, 922, 1004, 2657, 1576, 1615, 3515, 2346, 3429, 2583, 534, 988, 1614, 2592, 893, 450, 2396, 2636, 2427, 1284, 1009, 560, 928, 837, 554, 3525, 826, 546, 849, 915, 567, 2617, 1558, 841, 933, 926, 3400, 806, 3251, 994, 993, 996]. Performances [3415]. Period [74, 2091]. Periodic [1621, 1175, 1732, 332, 1949, 3239]. Periodicity [680, 1953]. Permanent [2571]. Permit [246]. Permutation [255, 377, 2038, 2224, 818, 2319]. Permutations [98, 2594, 1298, 251, 214, 2055]. Permuted [250]. Perpetual [1400]. Perpetuality [329]. Persistent [463]. Personal [2519, 3212, 3211, 526, 2533, 2527, 1686, 1006]. Persons [2904, 2927, 2895]. Perspective [42, 40, 2518, 265, 1808, 1881, 437, 1872, 1874, 1078]. Perspectives [202, 2515, 935, 826]. Perturbation [174]. Perturbations [317]. Perturbed [3514]. Petersen [2676, 1535]. Petri [843, 1160, 626, 1495, 1497, 1498, 1499, 1489, 1500, 1502, 632, 1184, 679, 1929, 2297, 1990, 363, 1185, 760, 1504, 2131, 1506, 1507, 1490, 1510, 1511, 849, 1558, 2130]. PGM [2224]. Phase [661, 535, 545, 808, 824]. Phase-Correcting [545]. Phased [3565]. Phases [2974, 839]. Phenotype [3127, 3098]. Philosophers [2835]. Philosophical [2896]. Phoneme [2986]. Photochemical [1572]. photometric [1091]. Physical [3033]. Physically [2919]. Physician [1673]. Pi [1483, 696, 1651, 3140, 719, 3361]. PI-Bus [3361]. Pi-Calculus [1651, 719, 696]. Pi-nets [696]. PICL [978, 850]. Picture [1352, 925, 693]. Pictures [345]. Piecewise [615, 1120]. pilot [1511]. Pin [2564]. Pin-level [2564]. Pipeline [2892]. Pipelined [1622]. Pipelining [160, 2667]. Piranha [148]. Pitfalls [2247, 3249]. Pittsburgh [3121]. PLA [2476]. Place [3562, 2429, 1494, 1490, 1516]. place-bordered [1516]. place/transition [1494, 1490]. Placement [2452]. Plaintext [2245]. Plan [3292, 3287, 3164, 3252]. Planar [1876, 2495, 1840, 2034, 2690, 2700, 1126, 1124, 1120, 1142]. Plane [1842, 1838, 2061, 3533, 1874, 3469]. Planes [2495, 1867]. planet [3213]. planner [3242].

Planning [1913, 2163, 2177, 1781, 1319, 550, 2969, 1056, 2185, 2010, 3233]. Plans [178, 1774, 2416, 2011, 2950]. Plant [3428]. Platform [2446, 2656, 2525, 3588, 3198, 3591]. Platforms [1572, 297, 3522]. Platonic [2093]. Playout [3191]. PLD [2465]. PLDs [2468]. plus [3483]. PLUSS [628]. PMIPS [184]. PML [197]. Point [2060, 2019, 2271, 2571, 3067, 2709, 1999, 697, 3458, 1989, 1063, 1047, 2552]. Pointers [1915]. Pointing [735]. Points [1868, 2075, 1392, 1917, 1113, 3453, 1131, 3452]. Poisson [821, 2697]. polar [1060]. Policies [2595]. Policy [1607, 3088, 2398, 3573, 2403]. Pollution [1572, 3481, 3491]. Polygon [2089, 1839, 1843, 1861, 1862, 2074]. Polylogtime [340]. Polymer [892, 3512]. Polymere [891]. Polymers [880]. Polymorphic [3060, 713, 3273, 2370, 702]. Polymorphism [3047, 731, 1568, 714, 716, 724]. Polynomial [3386, 2271, 2270, 2021, 1732, 2641, 1257, 2675, 1722, 351, 3477, 3544, 1379, 2014, 62]. Polynomial-Time [1257]. Polynomials [1574, 800, 3467, 3456]. Popperian [288]. Popular [1543]. Population [3097, 3325, 3108, 3099]. Populations [3086]. Portabel [1686]. Portable [2327, 836, 648, 3505, 2616]. Portation [3496]. Portfolio [2518]. Porting [857, 877]. Pose [1138, 1108, 1094]. Positive [368, 3000, 2389, 3313, 3319, 2051, 3324, 3428, 3368]. Possibilistic [2005, 73, 1262]. Possibilities [1313, 803]. Possibility [2418, 3266]. possibility-based [3266]. Possible [2957, 523]. Post [22]. Post-graduate [22]. Postgres [2786]. Postulates [2194]. pot [1465]. Potato [2720, 2052]. Potential [3425]. Power [559, 555, 348, 359, 1354, 2289, 408, 2430, 1306, 1168, 1361, 454, 326, 1695, 1364, 2581, 343, 1228, 3552, 3379, 2317, 2429, 3417, 560, 1690, 546, 510, 561, 2128]. Powerful [2188, 3039]. PPL [2211].

PPOST [2808]. PRACMA [2946]. Practical [1011, 1631, 2708, 46, 3090, 108, 109, 2935, 1278, 665, 2263, 111]. Practically [1299]. Practice [261, 47, 28, 23, 2980, 441, 187, 745, 2818, 1462]. Practitioners [20, 22]. Pragmatic [188, 2095]. Pragmatics [140, 1440]. PRAM [2723, 1545, 2722]. PRAMs [2815]. Pre [2475, 1685, 758]. Pre-Existing [1685]. Pre-Orders [758]. Pre-Processing [2475]. Precedence [170]. Precedents [2763]. Precise [2675, 3346]. Precision [1539, 2942]. Precomputation [2223]. Precondition [2622]. Preconditioned [955]. pred [660]. pred-LL [660]. Predicate [768, 1402, 3117, 3326, 593, 1503]. predicate/transition [1503]. Predicates [489, 2005, 660, 2144]. Predicative [1403, 1159, 1568]. Predict [2977, 2802]. Predictability [2465]. Predicting [3015]. Prediction [900, 927, 2154, 2687, 2161, 915, 2617, 841, 1068]. Predictive [1955]. Predictor [3185]. Preemptive [2714]. PREENS [916]. Preference [3117]. Preferences [2342]. Prefetching [462]. Prefix [2730, 2729, 1737, 1704, 1736]. Prefix-Matching [2730]. Pregnant [2802]. Prejudices [3338]. Preliminary [3486, 148, 3538, 1242]. premises [2144]. Preorders [674, 2148]. PREPARE [664]. Preprocessing [2828, 131, 1864, 130]. preprocessor [1473]. Prerequisite [2542]. Preselection [2939]. Presence [2250, 2711, 2832, 3314, 3321]. Present [1351, 278, 2094]. Presentation [2914, 54, 3091]. Presentations [3347, 1724]. Preserve [732]. Preserving [3559, 1409, 3099, 2029, 1129, 1085]. Preterm [2802]. Primal [1435]. Primary [2817]. Prime [1170, 3476]. Primer [2893]. Primes [2239]. Primitive [742, 371, 695, 818]. Primitives [241, 1252, 3357]. Primitivity [1733]. Principal [2381, 2312]. Principle [1177, 1056]. Principles [3413, 2212, 895, 3439, 2543, 1008, 1283]. Print [2926]. Printed [2925]. Prior [1265, 1923]. Prior-Thomason [1923]. priori [2178]. Priorities [2202]. Prioritized [2972, 2201, 2959, 3269]. Priority [2046, 1201, 2397, 1589, 1584, 3551, 725]. PRISM [2552]. Prisoner [1960]. Privacy [3411, 122]. Private [2524, 2248]. Privatization [169]. Privilege [3418]. Probabilistic [1381, 1492, 3003, 2644, 1533, 80, 681, 1168, 3024, 2408, 325, 588, 687, 2160, 2958, 2147, 1690, 2989, 316, 1380, 907, 3038, 1109, 2149, 3240, 2148]. Probability [101, 547, 1702, 548, 812]. Probably [3323]. Probe [3152]. Probe-B [3152]. Problem [2280, 1835, 183, 1383, 1546, 3163, 2284, 379, 580, 1608, 1351, 2730, 3162, 364, 2291, 3102, 426, 3542, 3501, 3504, 2090, 1990, 3154, 428, 2018, 1736, 1214, 3555, 2045, 1366, 3135, 2310, 287, 2023, 3146, 3517, 250, 1374, 613, 2174, 1708, 2317, 2983, 72, 1642, 594, 3176, 3113, 2101, 2835, 2152, 2753, 1491, 3225, 3253, 1098, 1433]. Problem-Independent [3135]. problem-oriented [3253]. Problem-Solving [594]. Problems [1713, 1833, 419, 2836, 2056, 2367, 3488, 1710, 3548, 884, 241, 1210, 3162, 357, 3416, 422, 2288, 354, 1390, 1591, 3075, 2967, 78, 1988, 3503, 2712, 1849, 1848, 3506, 3110, 2725, 1363, 363, 2061, 2306, 410, 1649, 2058, 430, 2312, 3394, 3089, 827, 2259, 2254, 668, 417, 3549, 3523, 2715, 826, 2086, 418, 1469, 3479, 800, 1470, 1456, 3216, 1461]. Procedural [2449, 1022]. Procedure [2385, 664, 3490, 272, 498, 3568, 2368, 1926, 2753, 1492]. Procedures [322, 965, 482, 495, 2335, 1503]. Process [206, 181, 183, 16, 1326, 202, 207, 2146, 1909, 184, 182, 1181, 1982, 199, 1160, 500, 3092, 2579, 195, 185, 190, 213, 3369, 845, 210, 191, 3022, 189, 3348, 208, 1321, 212, 1745, 1320, 50, 3077, 2767, 3551, 188, 1343, 205, 18, 187, 3346, 1337, 204, 194, 200, 203, 17, 698, 2142, 2115, 2132, 3271]. Process-Centered [182, 1321, 212]. Process-Programming [189]. Process-to-Function [1745]. Processes [1149, 1158, 515, 211, 2285, 2295, 856, 201, 1711, 853, 1630, 751, 3023, 3024, 3317, 770, 2684, 1582, 2161, 186, 1155, 748, 1157, 887, 1153, 196, 3185, 3038, 3553, 76, 90, 2120, 2121, 1042, 2119, 2117, 2122, 821, 2147, 2130, 2148]. Processing [380, 2475, 2446, 664, 1954, 931, 888, 1597, 855, 3497, 3498, 851, 842, 953, 2538, 2097, 1606, 2413, 443, 999, 3093, 3267]. Processor [831, 2632, 2470, 2473, 2440, 3049, 2447, 2052, 3026, 2815, 1603, 2438, 837, 2633]. Processors [2665, 884, 2579, 2613, 2575, 164, 487]. Prodigy [2152]. Prodigy/Analogy [2152]. Produce [2032]. Produces [3161]. Product [1661, 161, 2504, 1768]. Production [2784, 2521, 2422, 2262, 3351, 2752, 3594]. Productivity [39]. Products [1358, 104, 203, 2009]. Professional [23]. Profiles [839]. Program [466, 1779, 911, 157, 3292, 753, 2338, 2344, 3493, 49, 60, 1750, 885, 838, 3366, 2352, 2626, 896, 1771, 3066, 785, 1596, 2357, 3314, 796, 683, 2627, 883, 2956, 2968, 975, 797, 2616, 3048, 3134, 2326, 2489, 690, 1229, 1751, 1548, 2638, 176, 1550, 5, 1335, 1998, 841, 3124, 907, 1457, 1430, 1508, 1485]. Programmability [2613]. Programmable [2470, 2425, 2480, 2482, 2459, 2478, 2464, 2461]. Programmed [2905]. Programmers [45]. Programming [1755, 3127, 384, 2378, 2414, 2653, 700, 3489, 2622, 781, 1178, 3442, 889, 1783, 2733, 1597, 519, 1792, 599, 2342, 1019, 784, 769, 2614, 662, 2340, 783, 3126, 3272, 2797, 2000, 3053, 3381, 3273, 734, 1021, 3095, 730, 2853, 3125, 3312, 3128, 2211, 189, 2109, 733, 1022, 3511, 3436, 2386, 3398, 1599, 3540, 1225, 1615, 782, 3340, 2376, 3516, 1025, 397, 1593, 2371, 3134, 1758, 1753, 1934, 2324, 2487, 2856, 2806, 1267, 1003, 1028, 1782, 3380, 1029, 743, 2359, 3109, 980, 3382, 1268, 2375, 968, 2369, 517, 2040, 618, 593, 1595, 518, 620, 619]. Programming [1412, 2377, 2497, 2716, 2194, 2329, 774, 468, 2028, 2637, 699, 706, 707, 709, 3244, 2123]. Programming-by-Difference [3340]. Programs [1713, 3045, 2192, 2494, 1014, 1349, 3010, 2624, 1784, 3402, 2363, 319, 147, 675, 2496, 1594, 3060, 1790, 3370, 3058, 3288, 3496, 847, 2618, 1714, 1555, 2347, 2107, 170, 602, 173, 984, 3044, 647, 2882, 654, 760, 3301, 2943, 1187, 2356, 2921, 2620, 1789, 833, 2502, 2506, 2193, 2325, 2488, 3302, 1930, 3384, 175, 2452, 2275, 1785, 1229, 162, 440, 2498, 2567, 179, 1317, 262, 3057, 964, 3236, 1437, 3243, 722, 42]. Progress [2866, 3363]. Progressions [2239]. Project [15, 11, 484, 3195, 3592, 3412, 1010, 564, 19, 14, 12, 28, 2935, 89, 1334, 13, 1325, 3201, 3254, 992]. Project-Oriented [11, 14, 12]. Projection [435, 1792]. Projections [2048, 1120]. Projective [1876, 2495, 1867, 1871, 1879, 691, 1872, 1142, 1126, 1148]. Projects [2512, 3182, 857, 33, 27, 2845]. Prolog [1569, 1032, 1987, 2109, 2331, 3043, 3299, 1570, 3301, 2356, 719, 1024, 2210, 162, 1776]. PROMAL [2359]. Promised [219]. Proof [2030, 222, 322, 735, 2872, 2231, 2285, 1440, 1391, 1223, 1020, 2969, 219, 3344, 1226, 1774, 2952, 482, 1532, 1998, 3037, 2891, 791, 474, 1773, 1380, 2062, 3042, 1457, 1432, 1476, 1425, 1445, 1479, 1484, 641, 3271, 2134]. Proof-Based [3344]. proof-checked [1432]. Proof-Checking [1223]. proof-development [1484]. Proof-Theoretical [1773]. Proofs [1981, 2491, 1770, 1908, 752, 1531, 2497, 2894, 1486, 1424, 1459, 1468, 1418, 727]. Prop [3060]. Propagation [667, 552, 2413, 2996, 3376, 218, 549, 3231]. Properties [3059, 749, 3014, 3015, 2539, 2360, 1691, 1270, 3363, 2500, 1973, 3111, 2187, 3025, 1931, 1188, 686, 290, 691, 490, 2952, 366, 618, 2120, 1498, 1117, 1075, 1055, 727]. Property [1357, 1361, 1619, 343, 2188, 788, 1407, 2317, 1411, 404, 2062]. Proposal [1276, 1024, 111]. Propositional [322, 1388, 2206, 1406, 3258, 3266]. Propositions [1030]. Props [3440]. prosaic [3216]. PROSPECTOR [597]. PROSPECTOR-like [597]. Protecting [3411]. Protection [2512, 1757, 375, 3365]. Protein [1679, 1472]. Protocol [1542, 563, 3013, 224, 2226, 1223, 3187, 2238, 1627, 1632, 1563, 115, 3415, 3208, 2589, 1492, 223]. Protocols [2219, 1734, 3404, 123, 3587, 422, 2231, 2407, 2402, 3410, 1371, 1372]. Prototype [3196, 2804, 657, 1768, 2169, 3227, 1905]. Prototype-Based [2169]. Prototypes [2471]. Prototyping [1755, 2455, 1522, 2456, 1549, 2434, 2472, 2477, 1795, 2453, 2443, 2762]. Provability [2192, 1411]. Provable [657, 2239, 1431]. Provably [108, 662, 2334, 3019, 3020, 109, 1093]. Prove [505, 477, 478]. Prover [479, 480, 1940, 1472]. provers [1473]. Provide [1684, 2517]. Provider [2528]. Providers [2532, 2517]. provides [2464]. Providing [2869, 2855, 2535, 3582]. Proving [2198, 1218, 2866, 493, 2492, 2871, 2490, 2966, 3562, 1451, 272, 3025, 1188, 502, 503, 1477, 1487, 1429, 1470, 1430, 1459, 1443, 1420, 1442]. Proving-by-Example [272]. Provisioning [2518, 2528, 2541]. Proximity [1251, 2712]. Prserving [2505]. Pruning [1012, 423, 608, 702]. Pseudo [3099]. Pseudo-Meiosis [3099]. Pseudoduplication [2571]. Pseudorandom [1296]. psi [3474]. PSK [537]. PSK-TCM [537]. PSPACE [1710, 2725]. PSPACE-completeness [1710]. PSPACE-Hard [2725]. PST [2437]. pSystem [1615]. Public [2092, 96, 128, 2517, 257, 2220]. Public-Key [96, 257, 2220]. Publishing [2740]. pudding [1465]. Pulsed [1141]. Pumping [1723]. punctured [816]. Pure [1410, 3329, 1219]. Purely [737, 1540]. Purpose [2847, 655, 2473, 480, 2564, 2448, 3237]. pushbroom [1145]. Pushdown [2121, 1376]. Pushout [385, 393, 394, 395]. Putnam [2368]. Putting [28, 1752]. PVM [3481, 3493, 875, 970, 968, 3530]. PVS [478, 3037].

Q [3122]. Q-Learning [3122]. qcomp [842]. QOS [2400, 830, 3580, 3193]. QPN [843]. QPN-Tool [843]. QUAD [3379]. QUAD-CLP [3379]. Quadratic [2219, 1293, 219, 3379, 2028, 3446, 1055, 3466]. Quadrature [1573]. quadric [1119]. Quadtrees [2696]. Qualified [731]. Qualitative [843, 83, 2948, 80, 1195, 2008, 409, 2846, 76, 3239]. Quality [3577, 2401, 3207, 3576, 2646, 854, 1327, 905, 2543, 2541, 286, 1800]. Quantifiers [2272, 1992, 495, 1999]. Quantify [2581]. Quantitative [843, 1125, 2846, 3239]. Quantization [920, 825]. Quantum [880, 3488, 884, 883, 2058, 3517, 3477]. Quasi [1167, 2063, 2015, 1882, 1170]. Quasi-Full [2063]. Quasi-Invariance [1882]. Quasi-Prime [1170]. Quasi-Random [2015]. Quasi-Varieties [1167]. Quasigroups [3536]. Quasilinear [324]. Queries [1251, 2953, 420, 348, 436, 3309, 2737, 2711, 1305, 601, 3313, 459, 307, 1313, 461, 2072, 3276, 3226, 3227, 3231, 3269]. Query [1346, 3386, 1924, 460, 1936, 2754, 2809, 2841, 733, 1252, 2798, 1803, 1347, 443, 2758, 1805, 1896, 3230, 1436, 1438]. Querying [2419, 2796, 1666, 1668, 1904]. Questions [363]. Queue [1201, 2403]. Queueing [843, 832]. Queues [831, 2046, 1589, 2595, 2319].

R [3379, 2748, 2749]. R-Link [2748]. R-trees [2749]. r.e [3315]. R2 [2611]. RACE [2512, 564]. RACE-II [564]. Radio [559, 563, 543, 530, 2527, 571, 137, 2443, 560, 554, 2903, 566, 811]. Radiosity [2648]. Radix [2631]. RAID [2911, 273]. Rail [3333]. Railway [556, 3364, 501, 3149]. Railways [3331]. Rainbow [2314]. RAISE [3362]. Rake [548]. RAM [1863]. Ramanujan [1470]. Random [1248, 348, 2020, 2967, 1988, 2015, 2568, 1627, 1729, 3544, 2067, 1079, 1111]. Randomized [2836, 2092, 1728, 1735, 1855, 2024, 1858, 3557]. Randomness [2056, 1734, 1353, 2225, 2248]. Randomness-Rounds [2248]. Range [3490, 3492, 2987, 2709, 877, 1102]. Rank [1401, 692]. Ranking [2284, 377, 1674, 2699]. Rapid [2456]. Rapidly [548]. Rapidly-Adapting [548]. Rarefied [879]. Rate [3186, 2667, 573, 537, 1563, 816, 825]. Rate- [573]. Rate-Monotonic [3186]. Rate-Optimal [2667]. Rates [948, 831, 949, 3449]. Rating [555]. Ratio [1871, 1889, 2562, 1066]. Rational [1175, 2303, 2640, 2100, 3160, 2954, 3448, 3452]. Ratios [1868, 1091]. Rayleigh [535, 541, 806, 537, 540]. Razor [3140]. Re [1683, 210, 2168]. Re- [210]. Re-Engineering [1683]. Reachability [1542, 422, 3562, 326, 430, 1491]. Reaching [2832, 1327]. Reacting [870]. Reactive [1629, 1631, 2810, 3488, 1583, 338, 2434, 3030, 3032, 1161, 67, 3034, 3035, 1042]. Read [585, 2316]. Read-Once [585]. Read-Write [2316]. Readable [2883, 2355]. Reader [2907, 2905, 2915, 2823, 2822]. Reader/2 [2905]. Readers [2914]. Real [1762, 2559, 1909, 1386, 2219, 3015, 2539, 3016, 994, 1658, 3092, 1964, 1158, 1081, 3366, 662, 486, 3018, 856, 31, 1618, 12, 2880, 498, 3006, 299, 3023, 2392, 3026, 3027, 300, 302, 3029, 2467, 1188, 1632, 2901, 2396, 2793, 1347, 1611, 1159, 3526, 2400, 2794, 3038, 3553, 3446, 1104, 2269]. Real-Life [2539]. Real-Quadratic [2219]. Real-Time [1762, 1386, 3015, 3016, 1658, 1158, 662, 3018, 856, 31, 1618, 299, 3023, 2392, 3026, 300, 302, 3029, 1188, 2901, 2396, 2793, 1611, 1159, 2400, 2794, 3038, 3553, 3092]. Real-World [486, 12]. Realise [779]. Realistic [3303, 1581]. Realizability [1218, 1629]. Realization [2815, 75]. Realize [2493]. Realized [2432]. Realizing [2044, 3543]. Reals [479]. Rearrangement [3322]. Rearrangement-Independent [3322]. Reason [479]. Reasoner [1939]. Reasoning [2170, 2867, 3355, 2171, 2153, 3257, 2971, 2163, 2557, 2208, 2417, 2177, 2166, 1272, 3260, 1304, 1264, 2197, 2204, 1714, 1263, 2340, 2948, 411, 1260, 2205, 2949, 1261, 2880, 603, 2178, 3300, 1308, 1309, 2151, 2504, 410, 3028, 1265, 1262, 2956, 1930, 1312, 2160, 1269, 1314, 2168, 1267, 2174, 1928, 3304, 475, 1611, 1638, 2958, 3565, 2181, 1810, 469, 2802, 2152, 91, 2182, 2162, 1316, 3038, 2010, 3214, 1143, 3229, 3261, 3263, 3230, 1076, 3219, 1416, 3240, 3224]. Reasons [1986]. Receiver [570, 548, 567]. Receivers [3584]. Receptive [77]. recipe [1503]. recipe-based [1503]. Recognition [1888, 3164, 1891, 1885, 1140, 1602, 2643, 924, 1884, 2420, 1035, 1092, 1090, 1098, 1137, 1135, 1138, 1142, 1037]. Recognizability [345]. Recognizable [388, 1719]. Recognizing [1106, 2071, 3217]. Recombination [3103, 3113]. Recomposing [777]. Recomputing [1692, 1241]. Reconciliation [128]. Reconfigurable [1210, 2447, 2485, 2444, 2453, 2461]. Reconfiguration [3010, 2585, 2573, 2579]. Reconsidering [1351]. Reconstructing [1468]. Reconstruction [1205, 1878, 1695, 1879, 1881, 1785, 697, 1132, 1134, 1041, 1118, 1121]. Record [2381, 743]. Records [465, 2804]. recover [1102]. recover-and-select [1102].

Recoverable [2748]. Recovering [1110, 1875]. Recovery [445, 269, 285, 1870, 1089, 1882, 1146, 1102, 1034, 1051, 1104]. Recruiting [2933]. Rectangle [3546]. Rectangles [2713]. Rectangular [1624, 2753]. Rectilinear [2694, 1862]. Recurrence [2068, 1403, 3520]. Recurrences [962, 164]. Recursion [1725, 1697]. Recursive [3290, 1062, 2737, 762, 3297, 2625, 476, 367, 3299, 604, 683, 3317, 3322, 587, 1084, 3302, 761, 1540, 3231]. Redesign [1832]. Rediscovering [1376]. Redistribution [1560]. Reduce [1542]. Reduced [235, 1951, 811]. Reduced-complexity [811]. Reduced-State [1951]. Reducer [2658]. Reduces [750]. Reducibilities [2300]. Reducibility [328, 3559, 352]. Reducible [431, 350]. Reducing [3463, 666, 156, 23]. Reduction [2499, 3094, 1564, 232, 2656, 2218, 288, 1162, 1527, 1400, 3537, 2348, 788, 1563, 1682, 1530, 2623, 956, 3461, 809, 708, 3462, 1509]. Reductions [340, 150, 1652, 756, 3560, 1400, 1647]. reducts [3246]. Redundancy [2555, 3071, 2722]. Redundant [2376]. Reed [1947]. Reengineering [1326, 1338]. Reference [1204, 3283, 1915, 846, 1610, 2332, 1119]. References [2265]. Referential [1384]. refined [1463]. Refinement [3169, 3017, 3342, 3362, 1929, 1910, 747, 1222, 3024, 1186, 3354, 3341, 186, 1157, 1159, 3361, 508, 595, 2140, 1459, 3243, 1108]. refinement-tree [1459]. Refinements [2199, 351]. Refitting [1547]. Reflecting [594, 1453]. Reflection [912, 2212, 1796, 1031]. Reflective [2171, 779, 3176]. Reformulation [268]. Refutably [3323]. Refutation [3368]. Refute [279]. Regarding [1200]. Regions [157, 2581, 670, 1126, 1081]. Register [513, 1285, 104, 2885, 75]. registering [1119]. Registers [2819, 1288, 2823, 1294, 2666, 503]. Registration [1075, 1665, 1117, 1063, 1059]. regressive [823]. Regular [2997, 359, 3308, 3295, 2992, 1646, 3000, 2982, 473, 3310, 601, 294, 1025, 2991, 1378, 2981, 3545, 2145]. regularities [3255]. Regularity [2116]. Regularization [3506]. regularized [1121]. Regulation [3098]. Rehabilitation [2921, 2922]. Reification [3339]. Reinforcement [2414, 3150, 3088]. REKAP [3173]. Related [1833, 209, 127, 2877, 363, 2254, 3549, 2321]. Relating [768, 400]. Relation [746, 2362, 438, 1772, 3342, 1641, 1584, 92, 3328, 1325]. Relational [1522, 3017, 2112, 2744, 1691, 1819, 3370, 1828, 2840, 2342, 583, 2780, 2786, 759, 2804, 393, 1649, 3423, 1823, 2371, 2736, 1606, 437, 451, 1804, 3226, 3227, 1903]. Relations [1300, 1693, 2269, 2207, 685, 587, 1745, 2739, 2749, 475, 617, 2940, 1166, 3249, 3267]. Relationship [1205, 1819, 1811, 140, 1827, 2307, 1822, 1806, 1804, 1810, 1813, 3268]. Relationships [1662, 2549, 588, 1189, 3238]. Relative [1797, 2587, 1121]. Relativity [1328]. Relativization [1709]. Release [108, 109]. Relevance [2175, 763, 3262]. Relevant [2184]. Reliability [3534, 2582, 826]. Reliable [266, 282, 3188, 533, 2607, 3207, 2290, 2555, 275, 291, 1729, 3445, 2133]. RELIEF [586]. Remaining [287]. Remark [252]. Remote [2242, 1492]. Removal [1013, 1266, 1027]. Removing [1565]. Renaming [2824]. Rendering [2692, 854, 2647]. Reordering [2080]. Reorganization [427]. Repair [3084, 269, 3373]. Repaired [111]. Repeated [2376, 1870]. Replaceability [1642]. Replacement [389]. Replay [2620, 175]. Replicated [2814, 2042, 308, 3422]. Replication [1835, 2242, 766]. Report [484, 1010, 2935]. Reporting [2089, 355]. Repository [1824, 66]. Repository-Based [66]. Represent [2106]. Representability [3298]. Representable [2035]. Representation [3059, 2175, 1868, 2857, 1911, 1976, 386, 1923, 2803, 2421, 1307, 2093, 1195, 274, 1333, 79, 2900, 1977, 1269, 577, 3134, 63, 1030, 58, 3183, 2330, 2011, 3124, 3241, 1501, 1127, 3216]. Representational [3150, 2999]. Representations [2157, 1353, 2986, 1883, 2007, 1358, 1216, 1620, 1890, 1312, 2996, 1978, 1148, 3460]. Representing [1457, 2938, 178, 2081, 450, 619, 1412, 2894, 2010]. Reprogrammable [2440]. Requests [1836]. Required [2827, 799]. Requirement [1818, 1809]. Requirements [2788, 35, 2523, 3196, 2526, 2167, 2524, 192, 2508, 3585, 1327, 1329, 2509, 2355, 265, 858, 3199, 1331, 450, 67, 3279, 3400]. Research [3439, 2897, 891, 935, 2732, 929, 3425, 916]. Reservation [2407, 3179]. Reservoir [906]. Reset [2837]. Residual [2332]. Residuosity [219]. Resilient [2236, 2723]. Resistant [549]. resistive [1070]. Resolution [568, 2199, 1932, 2690, 734, 3438, 3533, 277, 2196, 2854, 1787, 3271]. Resolving [1143]. Resource [2056, 2828, 1781, 1709, 1389, 3194, 321, 2667, 3585, 530, 2211, 2392, 1963, 2715]. Resource-Bounded [321]. Resource-Constrained [2667]. Resource-Oriented [1781]. Resources [944, 2535, 3179, 1611, 3233]. Respect [853]. Response [1954, 551, 552]. restoration [1085, 1083]. Restricted [3307, 3088, 3255]. Restrictions [2941, 2560, 2207, 3034, 1466]. Restrictiveness [500]. Restructuring [3251]. Result [460, 457]. Resultant [2727]. Resultants [2641]. Results [3396, 911, 3101, 3079, 2708, 1016, 1838, 346, 3084, 3332, 982, 2389, 1694, 1853, 1639, 2299, 682, 557, 2897, 541, 612, 2583, 873, 531, 1325, 1093, 2119, 3235]. Resynchronization [105]. Retail [3429].

Retrieval [2788, 2156, 463, 2159, 2111, 3284, 2742, 923, 2789, 1664, 2097, 2749, 1674, 1825, 2180, 2165, 3183, 3276, 2162, 2790, 2950]. Retrieving [2165, 3570]. Reusable [2456, 9, 1749, 413]. Reuse [44, 46, 1818, 2862, 1770, 1339, 47, 2521, 1338, 3301, 45, 3160, 1337, 1893, 634]. Reusing [2861, 3166]. Reverse [1819, 1820, 531, 2713, 574]. Reverse-Fit [2713]. Revision [2418, 2193, 1313, 3356, 1317, 2194, 3245]. Revisited [1708, 632, 1448]. Revisiting [2078]. Rewrite [1784, 2500, 2657, 1520, 1530, 1531]. Rewriting [2384, 383, 1521, 1350, 749, 387, 1725, 2352, 2303, 623, 393, 329, 1536, 395, 396, 2502, 2851, 689, 2315, 2383, 738, 399, 400, 401, 1523, 1586, 704, 705, 636]. Rician [565]. Riemann [3470]. RIFLE [2564]. rig [1146]. Rights [114]. Rigid [2964, 1117, 1778, 2195, 1464, 1063]. Rigorous [3349, 2844]. Ring [2833, 2026, 2584]. Ring-Banyan [2584]. Ring-orientation [2833]. Rings [2495, 1851, 2833, 2055, 1199]. rippling [1449]. RISC [3026, 163, 2892, 487]. RISC-Pipeline [2892]. Risk [2241, 2460]. Risk-Free [2460]. RML [2333]. RNA [1258]. Road [2937, 1137, 1068]. Robot [2911, 38, 914, 1739, 2912, 3085, 2416, 2940, 2012, 1130]. Robotic [456]. Robust [1146, 583, 3405, 1076, 1080, 2026, 2947, 2416, 1103, 1074, 1137]. Role [3163, 598, 1260, 2779, 77, 57, 443, 2165, 1816, 1813, 3276, 1113]. Role-Attribute [1813]. Role-Based [443]. Roles [2207, 1743]. Roll [276]. Roll-Forward [276]. Rollback [269]. Rolling [3519]. ROMM [2603]. rotating [1133]. Rotation [2034, 2438, 2484]. Rotor [1279]. Rough [3221, 2977, 1271, 316, 3250, 3219]. Round [2245]. Rounds [2248, 235]. Routability [2465, 2481]. Route [1245, 2429]. Router [2606, 2607, 2611, 2602, 2595, 2592].

Routers [2591, 2594, 2590, 2598]. Routes [2074]. Routing [2601, 1836, 2720, 2612, 2829, 2593, 425, 1533, 2679, 2053, 2428, 2602, 2052, 2054, 334, 2467, 3151, 3146, 2609, 2603, 2599, 2610, 1616, 444, 2055, 335, 2319]. RPL2 [3155]. RPM [1008]. RSA [95, 231, 252]. RSA-120 [231]. RSL [3363]. RSVP [2407]. RT [2995]. RTL [2433, 3368]. RTSORAC [2793]. Rule [3084, 3122, 2881, 3299, 2804, 2657, 448, 2158, 3228, 3248, 3251, 3222]. Rule-Based [2804, 448, 2158, 3122, 3228]. Rule-Generating [3299]. rule-sets [3248]. Rules [2415, 2436, 406, 1985, 1305, 2803, 2001, 3086, 750, 2298, 3567, 3333, 2173, 2771, 1409, 316, 3253, 1894, 2134]. Run [2357, 1559, 1071]. run-distribution [1071]. Run-Time [1559]. Runnable [2355]. Runs [3021]. Runtime [2614, 3192]. Runtime-efficient [3192]. RUTH [3245].

s [703, 1952, 2353, 2883, 1914]. S-Semantics [2353]. S-Sets [1952]. S4 [1026]. SA [3369]. SA/RT [3369]. Sacks [2666]. Safari [2074]. Safe [1642, 402]. SAFER [1274]. Safety [556, 3363, 2442, 3025, 501, 2478, 907, 2566, 3260]. Safety-Critical [2442]. Salesman [3102, 3501, 2267, 3113]. Salience [2747]. Salosauna [1351]. Same [2563, 1977]. Sample [589, 303]. Sampled [65, 3040, 1060]. Samples [3000, 2041]. Sampling [74, 267]. SAMSON [2548]. SAT [2967, 2031]. Satellite [562]. Satellite-Mobile [562]. Sather [518]. Satisfaction [426, 3375, 3373, 3394, 3100, 3372]. Satisfiability [419, 2367, 3567]. Satisfying [100]. Saturday [54]. SBF [3172]. Scalability [842, 2613, 1576, 833]. Scalable [155, 266, 829, 958, 2604, 2829, 869, 1206, 946, 854, 953, 3510, 1602, 915]. Scalar [879, 1058]. Scale [979, 860, 1339, 862, 903, 1055, 1674, 2845, 972, 924, 2814, 2097, 1129, 1053].

Scale-space [1055]. scale-spaces [1129]. Scaleable [991]. Scaled [2411, 1078]. Scaling [3180, 3584]. Scanning [2670]. Scattering [3488, 2555]. SCDBR [1939]. Scenario [1326, 2548, 2547, 1325]. Scenarios [30, 38, 936]. Scene [84]. scenes [1138]. Scheduler [736, 2394, 3042]. Scheduling [3186, 2343, 3092, 2714, 1587, 160, 173, 3049, 2595, 2085, 3029, 1615, 3031, 670, 1581, 166, 1582, 983, 1859, 2398, 2683, 2278, 2715, 2716]. Scheldt [904]. Schema [3558, 1981, 2842, 1831, 759, 1814, 3274, 1808, 441, 449, 1804, 1817, 1340, 3227]. Schemas [442, 453, 465, 1815, 1227, 1823]. Schemata [3106]. Scheme [2707, 779, 533, 2240, 3559, 2222, 539, 2235, 425, 3576, 986, 2061, 3192, 252, 1243, 2599, 113, 2697, 3589, 2687, 1613, 1616, 215, 544, 2330, 3124, 926, 819, 111, 1450, 1017, 1607]. Schemes [1942, 256, 1944, 225, 2229, 101, 255, 238, 2233, 1021, 3560, 2725, 226, 2066, 3568, 683, 1216, 102, 3112, 276, 1799, 131, 214, 135, 2230, 3107, 239, 227, 130]. Schmidt [2632]. Schnorr [131, 130]. Schoof [3451]. School [52]. Schur [802]. SCIDDLE [979]. Science [895, 1174, 6, 34, 2929, 3281, 3213]. Scientific [2668, 2664, 1555, 3273, 3510, 3274, 179, 902, 3244]. Scoping [734]. Score [1234]. SCOTT [1471]. Screen [2914, 2907, 2905, 2915]. Screen-Readers [2914]. Screening [610]. Screens [2909]. script [1440]. Sea [904]. Seamless [2848]. Search [419, 2343, 3150, 3000, 2982, 2973, 1846, 1974, 426, 1391, 1986, 2043, 427, 1244, 2305, 920, 2239, 3077, 1025, 459, 3134, 3142, 3081, 3382, 1420]. searches [1515]. Searching [2696, 184, 2689, 1849, 2090, 1243, 2709, 3089, 1799]. Second [184, 345, 1404, 1532, 641, 2320, 695, 702]. Second-Order [345, 1532, 641, 695]. Secondary [2073]. Secret [1944, 225, 128, 108, 220, 228, 102, 257, 221, 135, 2229, 109].

Secret-Key [257]. Secure [3404, 3196, 3416, 3420, 3414, 108, 2555, 3332, 240, 3405, 2232, 1299, 3409, 2801, 3419, 2249, 3415, 2818, 109]. SecureNet [2512]. Security [3413, 2512, 3195, 2548, 2244, 3412, 2869, 3196, 2250, 3335, 3414, 2524, 1285, 1327, 249, 254, 2549, 125, 3403, 2547, 1806, 3415, 3400, 253]. Sediment [904]. SEE [182]. Seeing [1096, 1095]. seeking [2100]. Seeks [2921]. Segment [2595]. Segmentation [1103, 1079, 1104, 1882, 1081, 1068]. Segments [2017, 2708, 1856, 2072]. Seismic [3497]. select [1102]. Selected [3521]. Selecting [3427]. Selection [1834, 3104, 2419, 144, 1573, 2645, 600, 651, 3126, 3078, 2178, 3153, 1207, 3135, 3077, 3114, 3112, 2025, 3107]. Selective [538, 557, 3052, 350]. Self [2631, 2837, 1947, 2064, 3177, 1133, 2437, 2833, 2661, 1130, 2684, 431, 1857, 2705, 18, 3176, 1054]. Self-Adaptive [3176]. Self-Adjusting [2064]. Self-Assessment [18]. Self-calibration [1133]. Self-Checking [1947]. Self-Organizing [2684]. self-orientation [1054]. Self-Reducible [431]. Self-similar [3177]. Self-Sorting [2631]. Self-Stabilization [2705, 2837]. Self-Stabilizing [1857, 2833]. Self-Testable [2437]. Self-Timed [2661]. Semantic [1826, 765, 2744, 3443, 1830, 629, 1828, 2986, 2948, 2878, 747, 2741, 2357, 2179, 1466, 1812, 2798, 660, 637, 1901]. Semantic-Based [2357]. Semantical [2975]. Semantically [2744, 1429, 1471]. Semantics [1218, 381, 384, 1160, 1983, 2353, 1178, 2496, 452, 1985, 1179, 491, 1652, 729, 2944, 408, 3369, 2294, 1831, 2297, 1525, 2000, 2211, 1309, 2885, 396, 1228, 3384, 3383, 3357, 3358, 1313, 3359, 1785, 1266, 767, 2333, 451, 1585, 778, 1192, 1159, 1328, 2095, 1233, 1032, 3236, 700, 701, 2136, 3258, 707, 3262, 622, 2131, 1505, 721, 3216, 723, 2114, 2144]. Semester [4]. Semi [1696, 3544, 3475].

Semi-Deterministic [1696]. semi-groups [3475]. Semi-Random [3544]. Semiautomatic [966]. Semiconductor [865]. Semigroups [2304]. Semiorders [1977]. Semirings [347]. Sense [2971, 1311]. Sensitive [608, 3423]. sensitivity [1110]. Sensor [2937]. Sensor-Based [2937]. Sentences [1305]. Separability [2059]. Separable [814, 691]. Separate [739]. Separating [1842, 349]. Separation [773]. Separators [378]. Sequence [1237, 4, 1235, 2291, 547, 2443, 549, 882, 542, 822]. Sequences [2044, 2665, 1293, 370, 2986, 106, 104, 1238, 587, 2243, 2049, 3543, 3039, 1100, 1084, 1101, 1086, 1082, 1121]. Sequencing [2277]. Sequent [1783, 1262, 1484]. sequent-based [1484]. Sequential [2499, 3402, 1164, 1841, 3547, 1772, 1848, 2045, 905, 260, 1554, 327, 813, 1247, 1602, 1405, 882, 3553, 90, 2132]. Sequential-Parallel [905]. Sequentiality [1165, 1150]. Sequentially [3565]. Serial [1594, 571]. Series [1361, 3095, 2071, 3516, 2154, 3587]. Series-Parallel [2071]. Server [3571, 305, 500, 3188, 2479, 114, 2799, 3407, 3210, 2410, 923, 291, 1686]. Server-Process [500]. Servers [1667, 444]. Service [2536, 2532, 3577, 831, 2518, 2550, 2401, 2537, 2513, 2528, 3288, 2524, 2544, 1684, 3576, 3211, 2521, 2508, 3407, 2517, 2179, 2509, 2553, 2510, 2549, 2534, 3178, 2541, 2409, 2933, 2551]. Service-Driven [2550]. Service/Resource [2544]. Services [2514, 181, 2530, 3182, 305, 2518, 3188, 3581, 2526, 2863, 2531, 2529, 2507, 2786, 2392, 3180, 3572, 3204, 3199, 2535, 3578]. Serving [1836]. SESAME [2512]. Session [209, 32, 26, 20, 24, 16, 11, 40, 48, 44, 7, 3, 185, 190, 180, 197, 3398, 204]. Set [1979, 3396, 2017, 2387, 2877, 1392, 2389, 3393, 730, 2468, 1991, 2386, 1668, 429, 3322, 1993, 399, 316, 3430, 792, 2893, 3250, 1056, 1515, 1517]. Set-Based [3393, 1668]. Set-Driven [3322].

Set-Theoretic [399]. SETHEO [1474]. Sets [578, 2060, 2019, 3291, 1952, 2415, 1621, 2020, 1352, 388, 1017, 2270, 582, 1967, 1850, 2300, 431, 1026, 1722, 2977, 2256, 1786, 350, 404, 3248, 1974, 3219]. Setting [1779]. Seven [3338, 3444]. Several [3535, 679]. Severely [2919]. Sexual [3077]. SGML [1675]. Shading [1875, 1093]. Shadow [560]. Shallow [472]. Shape [2857, 1078, 1093, 1116, 61, 1086, 1875, 1048, 1112, 1125, 714, 1127, 1034, 1051, 1118]. Shape-adapted [1116]. Shapely [714]. Shared [266, 2653, 2654, 3500, 1553, 156, 976, 2641, 2913, 2620, 1612, 2834, 2652, 175, 977, 2361, 3573, 135, 2650]. Shared-Control [2913]. Shared-Memory [2620, 175, 2650, 2834]. Shares [101]. Sharing [1944, 225, 2229, 831, 2588, 156, 3061, 228, 102, 2348, 689, 221, 3160, 220, 718]. Shell [1320, 3431, 1484]. Shestakov [1943]. SHGCs [1104]. Shift [1285, 1288, 104, 1294, 2085, 75, 503, 3464]. Shift-Register [104]. Shop [2343]. Short [1981, 3490, 3492, 3095, 228, 1729, 2230]. Short-Range [3490, 3492]. Shorter [3186, 2319]. Shortest [1248, 2711, 1841, 3542, 1703, 2306, 2091, 2074]. Should [1432, 271, 1741, 1742, 2517, 477]. Shrinking [216, 1278]. Shuffle [1368]. Sided [303]. Sidel’nikov [1943]. Sidel’nikov-Shestakov [1943]. Sieve [230, 1225, 3457]. Sieves [94, 3476]. sieving [3447]. Sighted [2927]. Sigma [3140]. Sigma-Pi [3140]. Sign [2900, 1137]. Signal [3093, 824, 3031]. Signalling [3331, 3333, 501]. signals [808, 820]. Signature [1942, 255, 2222, 2235, 2476, 252, 464, 131, 214, 2790, 111, 130]. Signatures [116, 2221, 3408, 238, 108, 2587, 2220, 239, 109]. Signed [3220]. Signification [2513]. Signing [2234]. SIMD [2665, 3491, 840, 94, 2625, 871, 1536, 1602, 1208, 1601, 1559]. SIMD-Systems [871]. similar [3177].

Similarity [2156, 2163, 579, 2164, 2157, 1256, 2160, 2161, 2158]. Simple [2331, 2622, 3308, 1524, 1707, 505, 2066, 1619, 275, 355, 686, 1448, 3325, 523, 738, 1862, 2556, 909, 2701, 791, 2074, 1864, 3074, 2959, 2028]. Simplicity [2341, 300]. Simplification [2390, 2366, 712]. Simplifications [2349]. Simplified [2231]. Simplifies [115]. Simplify [224, 2752, 223]. Simplifying [2367, 2675, 2889]. Simply [3050]. Simply-Typed [3050]. Simulate [2314]. Simulated [1608, 3137, 3135, 3134, 898]. Simulating [918, 2468, 2562, 890, 2454]. Simulation [834, 2574, 911, 3012, 556, 704, 881, 3130, 19, 879, 1956, 838, 896, 80, 3020, 853, 868, 1186, 865, 937, 1545, 541, 552, 858, 867, 1188, 1296, 325, 612, 3048, 897, 3519, 886, 81, 990, 503, 936, 1512, 3523, 904, 870, 3526, 3171, 2182]. Simulation-Based [2574]. Simulations [908, 489, 869, 3535, 894, 887, 915, 2147]. Simulator [906, 3428, 836]. Simulators [2566]. Simultaneous [2284, 2635, 3375]. Singapore [1655]. Single [242, 121, 3542, 168, 2565, 393, 365, 394, 2923, 395, 540, 1712]. Single-Assignment [168]. Single-Pushout [395]. Single-Source [3542]. Single-Term [242]. Single-Valued [365]. Singular [3484, 960, 801]. Singularities [3533]. Situation [1939, 2539, 1273, 3173, 2979]. Situations [3168, 2010]. Size [101, 1551, 2081, 2304, 1730, 2620, 3108, 175, 1373, 2230, 1877, 3114]. Size-Depth [1551]. sized [2813]. Sizes [2595, 457]. Sizing [2672]. Skeletons [985]. Skew [3537]. SLD [1787, 2328]. sleep [1517]. Sleepers [2345]. Sliced [464]. Slicing [149, 2352]. Slotted [569, 566]. Slow [538]. Small [578, 582, 1706, 2027, 349, 2036, 2813, 2752]. Small- [2813]. Smaller [523]. SMART [863]. SMIsC [3443]. SML [648, 2797]. SML/NJ [648]. SMoLCS [626]. smooth [1117, 1041]. smoothing [1116]. SMU [49].

Snapshot [2823, 2704]. SNMP [2522]. SOCCA [198]. Social [1959, 1958, 1273, 1956, 3097]. Social-Like [3097]. Societies [1956, 412]. Society [2779]. Soft [2559, 2796, 3153, 574]. Soft-Real-Time [2559]. Software [206, 181, 183, 26, 40, 48, 44, 3, 2455, 3426, 35, 1277, 4, 51, 881, 30, 46, 2669, 2451, 2558, 3494, 912, 211, 49, 43, 19, 2, 885, 53, 844, 38, 3362, 3499, 946, 521, 1339, 160, 31, 2646, 12, 2667, 507, 47, 210, 28, 23, 509, 10, 285, 1338, 3510, 2582, 9, 3192, 937, 2629, 287, 2570, 2510, 50, 6, 13, 63, 41, 55, 2846, 29, 34, 945, 969, 1281, 1344, 2628, 5, 204, 194, 2813, 203, 795, 947, 196, 17, 2566, 623, 1439, 42]. Software- [969]. Software-Faults [287]. Software-Optimised [1281]. Software/ Hardware [795]. Solid [2726, 1041]. Solid-on-Solid [2726]. Solitary [244]. Solomon [1947]. Solution [3486, 1671, 856, 2306, 2640, 1367, 2813, 2753, 872, 1502, 2464]. Solutions [3416, 1988, 2021, 2258, 2267, 1720]. Solve [2834]. Solver [1011, 963, 875, 861, 961, 877, 876]. Solvers [874, 3176, 3529, 3531]. Solving [2495, 3163, 860, 884, 2271, 2938, 580, 3162, 1723, 2341, 2293, 2068, 2390, 3555, 410, 3519, 3520, 613, 2174, 2263, 668, 594, 2152, 2014, 2753, 1491]. Some [266, 1167, 1182, 994, 2698, 2387, 1016, 1470, 1390, 3084, 315, 2389, 78, 3381, 682, 1363, 363, 2061, 1278, 3513, 2373, 3216, 430, 477, 1003, 1688, 2577, 2317, 478, 2194, 418, 1693, 2462, 2154, 3475]. Sometime [1910]. Sometimes [1910, 1369]. Sort [799]. Sorted [1724, 2965]. Sorting [2631, 3566, 332, 1207, 334, 1738, 2721]. Sorts [2965, 2210, 1460, 624]. SOS [749, 704]. Sound [3428, 722]. Soundness [3563]. Source [3542]. Sources [1272, 1087]. Southern [904]. SP [879, 3504]. SP-1 [879, 3504]. SP1 [2601]. SP2 [3525]. Space [1381, 222, 1234, 2695, 1832, 1236, 3137, 1646, 1838, 2982, 2421, 1735, 2670, 2822, 2357, 2660, 325, 3088, 446, 2705, 2323, 323, 1073, 1055, 1509, 1053]. Space-Efficient [325]. Space-optimum [2822]. Space-Time [2670]. Spacecraft [3027, 3152]. Spaces [1270, 306, 3063, 1025, 279, 3142, 592, 1081, 1129]. Spacing [3131]. SPADE [182]. SPADE-1 [182]. Span [2222, 106]. Spanner [2693]. Spanning [1197, 1706]. Spark [2658]. Sparse [3485, 144, 2626, 3507, 1576, 3524, 1978, 1559, 3531]. Spatial [2035, 2914, 2949, 1195, 2777, 79, 2738, 2748, 3112, 2976, 1657, 2940]. Spatially [1087]. spatiotemporal [1066]. Special [2511, 2813, 87]. Specialization [1408]. Specific [88]. Specification [1762, 3012, 2536, 1939, 2431, 1761, 3017, 3335, 198, 3018, 338, 2785, 1526, 2883, 2621, 3026, 2627, 1187, 405, 2553, 1528, 3032, 1161, 503, 659, 1344, 2361, 3034, 3434, 3361, 3365, 3036, 1907, 2552, 3390, 3241, 629, 633, 719, 640, 641]. Specifications [337, 2436, 1522, 740, 1818, 3353, 3370, 3288, 2618, 3332, 2625, 3049, 3360, 1329, 2542, 2355, 1156, 3344, 2193, 2325, 2488, 1809, 1529, 398, 658, 3035, 1193, 2762, 1800, 362, 627, 3259, 628, 630, 636, 720, 639]. Specificity [2202, 2960]. Specified [1598, 2039]. Specifying [3352, 3363, 3007, 1344, 2545, 403, 3039, 2893]. Spectrum [570, 543, 1525, 572, 2443, 565]. Speculative [653, 269]. Speech [568, 1954, 1955, 2899, 2914, 2907, 1333, 1331]. Speed [2472, 932, 2438, 2432, 936, 929]. Speed-Up [2432]. Speedup [1557, 1986]. spheres [1128]. SPIKE [1486]. SPILL [2355]. SPILL-2 [2355]. Spiral [12]. Spirit [1966]. Splaying [3561]. spline [1121]. Split [2546, 3547]. SPMD [2618, 987, 964]. SPN [2560]. Spoken [3002]. SPP [3516]. Spread [570, 2443, 565, 572]. Spread-Spectrum [570]. Spreadsheet [520]. SProc [2125]. SQL [1675, 2809, 2841, 2734, 1899, 2758]. SQL-Like [2841]. Squares [1246, 1110]. ST [2407, 2406]. ST-2 [2406]. ST-II [2407].

Stability [1038, 3529, 3074, 1147]. Stabilization [2705, 2837]. stabilizers [1124]. Stabilizing [1857, 2833]. Stable [1715, 2944, 408, 273, 2279, 3262]. Stack [1349, 2474, 567]. Stack-Based [2474]. Stackability [3050]. Stages [3171]. Staggered [1607]. Stalk [362]. Standard [2492, 386, 1306, 967, 3569, 3209, 3427, 295, 3057, 1050, 2213, 469]. Standardization [433]. Standardizing [2112]. Standards [2507]. Standpoint [1374]. Star [2286, 343, 2317, 1665]. State [1382, 3012, 1386, 831, 2992, 1832, 1183, 1646, 1653, 2206, 2437, 3088, 1598, 1173, 482, 3573, 1951, 3545, 2929, 3039, 1488, 1509, 803]. State-Dependent [831]. State-Space [1832]. Statecharts [491, 396, 767, 3011, 2114]. Statements [664]. States [310]. Static [1555, 3049, 2357, 3048, 2454, 3252]. Statically [1178]. Statically-Typed [1178]. Station [534]. Statistical [2990, 438, 821, 2994, 2991, 3073, 2567, 820, 1066]. Statistically [238, 239]. Status [1010]. Steiner [2694, 2063, 2061, 2054]. Steno [2901]. Step [3114, 2162]. Step-size [3114]. Steps [521]. Stepwise [3362, 3030, 1682, 186, 508]. Stereo [1868, 917, 1146, 1130, 1073, 1132, 1114, 1091]. Stereoscopy [2515]. stereovision [1077]. STIMS [81]. STIMS-MEDTOOL [81]. Stit [1927]. Stochastic [3012, 2992, 2990, 3003, 1502, 2985, 88, 1258, 1105, 1101]. Stochastic/Non [3012]. Stochastic/ Non-stochastic [3012]. Stock [2753]. Stokes [875, 861, 877, 876]. Stokes-Flow [861]. Stop [238, 694, 239]. Storage [456, 2756, 2408, 3276, 1512]. Store [3571, 463]. stored [1903]. Stories [465]. story [3238]. Str [1444]. Straight [3454]. Straight-line [3454]. Strategic [1964, 2779]. Strategies [3163, 3101, 1573, 705, 2197, 1391, 2975, 2939, 3111, 2658, 1020, 2661, 2399, 3315, 2410, 313, 1739, 3354, 2638, 3545, 882, 3276, 1420]. Strategy [2499, 529, 3079, 1763, 2071, 1552, 2895, 82, 2405, 3115, 3351, 1901]. Stratification [1314]. Stratified [2792]. Stream [1277, 1280, 105, 1286, 217, 1500]. Streams [2411, 1225]. Strengthening [3410]. Stress [2977]. Strict [100, 1962, 3061]. Strictly [1551]. Strictness [3069, 3067, 3064, 3068, 711]. String [1254, 356, 2730, 1256, 2296, 2090, 1245, 1242, 1246, 417, 1377]. String-Matching [1245]. Strings [1248, 1257, 2050, 3380]. Strip [1560]. Strong [1218, 1395, 329, 2082, 1995, 1520, 708, 1441]. Strongly [1467, 815]. strongly-connected [815]. Structural [2164, 2849, 2998, 2900, 1228, 1152, 2318, 2567, 774, 1499, 1509]. Structure [880, 2924, 1868, 144, 357, 1677, 1273, 376, 2731, 2618, 2718, 315, 471, 2179, 2542, 607, 2108, 2096, 1870, 1679, 2748, 401, 3143, 3276, 1049, 1048, 1116, 1084, 2123]. Structured [2993, 2157, 2740, 604, 1252, 639, 2144]. Structures [1384, 1717, 1733, 2558, 2110, 1005, 473, 2015, 367, 2437, 649, 393, 394, 2467, 2756, 2002, 1870, 2433, 143, 400, 1872, 440, 1874, 404, 2102, 2454, 3253, 1509, 1040, 1356, 1360]. Structuring [2926, 2509, 2782, 2543]. stubborn [1515, 1517]. Students [28, 2927, 2928, 2929]. Studies [295]. Study [3010, 3127, 43, 3362, 3332, 2945, 3133, 2927, 2738, 1225, 3444, 2510, 2549, 1528, 2767, 2185, 1407, 503, 2262, 3337, 1590, 3427, 985, 1325, 1558, 3242]. Studying [3098, 2887]. Stumps [590]. Stuttering [1930]. Style [2514, 1227, 2736, 502, 1479, 2134]. Styles [3430, 2362]. Sub [1164, 545]. Sub-Language [1164]. Sub-optimal [545]. Subclass [309]. Subclasses [1743]. Subcodes [1948]. Subexponential [229, 3448, 3467]. Subgraph [1858]. subgroup [3448]. Subgroups [1710].

Subject [788, 1752]. Subjunctive [1303, 1313]. Sublanguage [146]. Sublattices [614]. Subliminal [110]. Sublinear [2305]. Sublogarithmic [2323, 323]. subnets [1516]. Suboptimal [1236, 1242]. Subpages [1600]. Subpopulations [3115]. Subroutine [970, 3244, 1439]. Subsequence [370]. Subsequences [1249, 1703, 2276]. Subset [64, 3126, 2033, 2320]. Subspace [957]. Substitution [3371, 1997, 2247]. Substitutions [2294, 3555]. Substrate [2607]. Substrings [1239]. Subsumption [1302, 436, 410, 440]. Subsumption-Free [440]. Subsystem [273, 3414]. Subsystems [358, 2456, 2442]. Subterm [3556]. Subtrees [2060]. Subtyped [761]. Subtyping [1755, 762, 741]. Success [1325, 812]. Successful [22, 805]. Successive [2582, 540]. Succinctness [364]. Sufficient [2491, 1048, 1486]. Suffix [709]. Suggested [30]. Suitable [2434, 1295, 1491]. Suite [950, 655, 2630]. Sum [2033]. Super [496, 2083, 1738]. Super-Logarithmic [1738]. Superclasses [1741, 1742]. Supercompilation [3070, 726]. Supercomputers [936, 909]. Supercomputing [861, 891, 935]. Superhighways [852]. Superlinear [1986]. Superposed [1502]. Superposition [693, 821, 1455]. superquadric [1102]. Superscalar [2447, 837]. Supersequences [1249, 1703]. Superstrings [1248, 1841]. Superunitary [2639]. Supervised [3126, 611]. Support [3355, 1336, 51, 2159, 3581, 2537, 3164, 151, 1319, 439, 2395, 2608, 2904, 845, 3433, 2411, 1329, 1814, 2179, 3285, 3572, 782, 974, 1320, 3199, 2391, 3429, 2928, 1345, 2937, 2541, 2540, 490, 29, 2409, 2598, 3582, 1786, 3579, 1324, 2787, 3276, 508, 1504]. Support-Environment [3572]. Support-for-Negation [1786]. Supported [3438, 2888, 1752]. Supporting [906, 2663, 3440, 3577, 463, 2768, 3194, 2392, 663, 446, 194, 2775, 3422, 3279, 2929]. Supportive [2927]. Supports [1680]. Surface [1883, 988, 1112, 1075, 1110, 1089, 1119, 1085, 1121]. Surfaces [2034, 1970, 1891, 1117, 1036, 1053]. Surfactants [886]. Surveillance [2512]. Survey [1935, 624]. Susceptibility [1942]. Suspension [1625]. SVR4 [2394]. SW2000 [2577]. Sweep [3420, 2289]. SWitch [836, 2650]. Switched [835]. Switching [2582]. SWNs [1496]. Symbol [544]. Symbolic [2997, 1621, 2872, 1723, 3069, 706, 2380, 2358, 173, 1639, 2674, 1632, 665, 503, 999, 1515, 2155, 697, 2136, 1470]. Symbolic-Model [1632]. Symlog [1479]. Symmetric [753, 1284, 1496]. Symmetry [1150, 2705, 1515, 1118]. Symmetry-Breaking [2705]. Synchronisation [3577, 3191, 3190, 3589]. Synchronization [106, 1857, 145, 2830, 2649, 3578, 2329]. Synchronized [2899, 3015, 1923]. Synchronous [1180, 1583, 105, 3053, 1541, 3574, 2432, 3203, 532, 2567, 2589, 542, 1083, 1497, 2127, 2817]. Synchrony [3035]. Syndrome [215]. synergistic [718]. Syntactic [2198, 755, 2374, 1772, 747, 2994, 660, 1378]. Syntactical [2501]. Syntax [381, 1780, 497, 1172, 700, 638]. Synthesis [1629, 3292, 2624, 270, 3286, 2899, 2431, 2360, 3106, 3288, 2437, 1418, 3314, 3317, 2484, 2185, 1003, 2426]. Synthesized [2914]. Synthetic [248]. System [2574, 3355, 559, 1913, 1669, 1631, 1521, 563, 2791, 543, 2899, 51, 3170, 1766, 2656, 132, 3416, 3411, 828, 3121, 3087, 3092, 2843, 3335, 912, 3349, 763, 2859, 2285, 562, 989, 3207, 1597, 424, 2803, 3497, 2686, 2840, 3084, 3332, 2904, 925, 1656, 1618, 1307, 338, 1757, 914, 479, 2786, 3381, 584, 219, 3128, 2946, 965, 480, 2582, 2804, 2781, 1957, 2485, 2735, 2777, 3511, 2896, 3027, 89, 2927, 2855, 1668, 866, 744, 525, 2884, 2789, 2356, 3278, 2913, 1612, 405, 883, 2408, 481, 2172, 2751, 782, 974, 2887, 2364, 1885, 2812, 2391, 1938, 3429]. System [2864, 3516, 3032, 1795, 1593, 2583, 2097, 1765, 2541, 2444, 2783, 3441, 2453, 2326, 2489, 2858, 1229, 63, 540, 977, 3152, 1996, 2409, 2934, 3590, 3175, 3427, 2361, 67, 928, 2890, 3434, 1613, 761, 1335, 3365, 402, 58, 1816, 3183, 518, 403, 1566, 1324, 474, 70, 795, 2903, 2787, 3201, 2617, 3356, 3280, 2420, 1665, 787, 1640, 362, 3245, 1486, 1495, 705, 1112, 634, 2800, 3250, 1144, 3253, 1905, 3234, 1512, 1485, 3235, 2951]. Systematic [3046, 300, 2570, 2628, 1506]. Systems [742, 1346, 65, 908, 358, 1382, 2867, 2559, 342, 1941, 2030, 1693, 746, 1519, 222, 1784, 2814, 3330, 1205, 384, 829, 3012, 597, 199, 2573, 282, 1675, 2471, 3412, 1336, 2796, 1626, 3014, 3015, 3016, 538, 2862, 151, 456, 2842, 3420, 533, 1183, 211, 1700, 921, 60, 955, 1354, 3367, 869, 959, 1725, 2635, 971, 389, 3352, 2352, 3086, 2877, 2614, 772, 946, 3018, 315, 2784, 1701, 896, 2107, 345, 3122, 31, 2500, 156, 2358, 3056, 2434, 3123, 2741, 3019, 1168, 3284, 3021, 434, 854, 2841, 2554, 1327, 2734, 3006, 2853, 3509, 1824, 3439, 68, 516]. Systems [299, 392, 654, 414, 1333, 1661, 1570, 1527, 2179, 3025, 2304, 2422, 558, 2404, 2485, 9, 3569, 1400, 329, 2627, 86, 300, 394, 2423, 1004, 302, 3007, 2782, 1265, 952, 2660, 3029, 2685, 287, 2798, 3444, 2848, 2640, 1600, 3030, 2834, 265, 2652, 1188, 3514, 2502, 88, 3199, 528, 2864, 441, 446, 59, 2851, 82, 2767, 3033, 1331, 2540, 450, 1822, 689, 1623, 57, 66, 81, 2801, 3520, 63, 1373, 2004, 1598, 448, 56, 987, 3008, 29, 1624, 1682, 945, 738, 61, 1271, 827, 969, 732, 3419, 2478, 1347, 3275, 412, 93, 1341, 560, 980, 3034, 1611]. Systems [443, 400, 2951, 1532, 1159, 1638, 401, 1410, 295, 1033, 3036, 309, 973, 2545, 549, 2552, 776, 2750, 2561, 2794, 947, 1520, 1219, 1718, 1523, 2802, 1530, 1531, 565, 2589, 2766, 3528, 1712, 898, 992, 916, 1558, 404, 1318, 933, 792, 3074, 62, 3040, 3041, 87, 85, 774, 3390, 458, 3531, 3400, 1194, 695, 2145, 699, 2140, 1892, 704, 1488, 2143, 1501, 2119, 3594, 1906, 1365, 636, 3219, 2393, 1481, 822, 641, 3271, 1514, 2114, 811, 871, 853, 87, 2116]. Systolic [1574, 2661, 1215, 2441, 2634, 1006, 2628].

T [3587, 918]. T-node [918]. T.120 [3587]. T3D [3508, 949, 2610]. Table [2945, 2458, 350]. Table-Lookup [2458]. tableau [1477]. tableau-based [1477]. Tableaux [1937, 2962, 1466, 1467]. Tables [1952, 375, 3246]. Tabu [459]. Tabulated [375]. Tabulation [2363]. TabVer [2945]. Tactic [496, 1770, 1459, 1229]. Tactics [1771]. Tailoring [1697]. Talking [2935]. Tally [2300]. Tamarack [485]. Tamarack-3 [485]. Taming [1764]. Tape [3535]. Tapes [3535]. TAS [2198, 2240]. TAS-D0 [2198]. Task [2669, 986, 2618, 1581, 2184, 1859, 909]. Tasks [2163, 2177, 2988, 917, 2969, 2659, 2084, 3248]. TAU [2616]. Tautologies [1981]. Taxonomy [1804]. TCM [537]. TDGs [3065]. TDMA [559, 564, 528, 811]. Teach [10]. Teaching [37, 46, 19, 54, 36, 2299, 9, 29]. Team [2966, 3305]. Teams [283]. Technical [2179, 86, 2172]. Technique [1542, 3577, 2306, 1580, 541, 2348, 2432, 2851, 2433, 1537, 186, 2335, 1066]. Techniques [419, 880, 568, 2557, 2653, 2537, 3367, 2948, 486, 160, 862, 28, 1596, 285, 2442, 2472, 2769, 2738, 3343, 610, 2570, 2994, 2956, 1188, 833, 3336, 1808, 612, 1760, 3521, 827, 1907, 1816, 826, 992, 702, 1434, 623]. Technological [521]. Technologies [2576, 2913, 553]. Technology [1680, 2920, 2268, 1673, 2811, 213, 2430, 2896, 208, 1334, 506, 2923, 2895, 2779, 2767, 57, 66, 2746, 204, 1660, 433, 2794, 2921]. Tele [3203]. Tele-Cooperation [3203]. Telecommunication [3211, 2520]. Telecommunications [297, 2540]. Teleconferencing [3573]. Telecooperation [3588]. Teleservice [3571, 3209, 1676, 3570]. Teleservices [3197, 3196, 3212]. Teletutoring [3202]. Television [3181, 803]. Telework [2932]. tell [1136]. Telling [465]. Telmat [918]. Temporal [1913, 1302, 1924, 3015, 1936, 489, 452, 1637, 1932, 1792, 2754, 2204, 1929, 1910, 2340, 1802, 1915, 1937, 1186, 1829, 2755, 1931, 2886, 3028, 320, 2756, 764, 1933, 1930, 3032, 2096, 1269, 409, 1406, 1934, 1928, 1030, 1803, 1907, 1895, 2757, 1660, 2758, 1925, 1916, 1801, 1926, 2120, 3237, 1894, 3239]. Ten [3554, 1139]. Tense [1922, 2096, 795]. Tensor [161]. Term [1521, 749, 242, 121, 2352, 329, 2965, 587, 2502, 689, 2315, 2383, 738, 2330, 1523, 1423, 704, 1501]. Termination [1449, 2505, 2501, 2297, 2500, 1446, 738, 1531, 1448, 3271]. Terminological [1301, 2949]. Terminologies [406]. Terminology [2799]. Terms [307]. Ternary [2495]. Terrain [415]. Test [514, 2425, 2424, 2485, 866, 867, 2568, 531, 2561, 1071, 1066]. Testability [2476, 75, 1540]. Testable [2437, 3319]. Testbed [930]. Testbench [2434]. Testing [2143, 2630, 1973, 2853, 3024, 663, 429, 1627, 2453, 2728, 2567, 2566, 2148]. Testing-based [2143]. Tests [3385, 1101]. Tetrahedralization [2075]. Texas [21]. Text [1675, 463, 2106, 2731, 2689, 2301, 923, 1243, 2097, 2901, 464, 2765]. Text/ Relational [1675]. Texts [3318, 2098, 2099]. Texture [82, 1110, 1069]. TFDL [2757]. Their [2284, 993, 2269, 2001, 2772, 1711, 392, 558, 393, 1975, 69, 176, 218, 1378, 403, 620, 1884, 2523, 1694, 2587, 2549, 1532, 752]. Them [266, 852]. Themes [2025]. Theorem [2198, 1350, 2871, 2313, 2966, 1719, 680, 479, 793, 480, 1940, 503, 1152, 1477, 1487, 1429, 1470, 1430, 1459, 1420, 1473, 2144, 1442, 1014]. Theorem-Prover [480]. Theorems [493, 2492, 3554, 505, 798, 1163, 477, 478, 1443]. Theoretic [1628, 2353, 1270, 336, 89, 609, 451, 399, 3479, 3224]. Theoretical [2708, 64, 1185, 1369, 2154, 1033, 1773, 2978]. Theories [3303, 908, 1301, 1167, 2387, 2203, 1388, 2212, 2960, 2204, 1191, 1408, 2194, 1927, 3247, 1434, 3269]. Theory [742, 488, 493, 2003, 1472, 3426, 3101, 2868, 765, 1715, 2874, 422, 1984, 3538, 1220, 60, 1772, 1273, 2418, 3347, 1524, 2877, 411, 756, 367, 23, 1224, 730, 794, 1814, 2520, 1311, 373, 1960, 1227, 441, 2103, 264, 1462, 324, 63, 56, 2952, 1232, 2977, 1965, 70, 316, 1726, 757, 2102, 792, 2191, 62, 745, 2818, 2959, 85, 695, 3245, 3472, 1458, 640]. Therapy [3292]. Thesauri [2764]. Thinned [168]. Thinning [1208]. third [1136]. Thirty [799]. Thomason [1923]. Threaded [462]. Threads [2396, 2400]. Three [1868, 232, 3334, 2582, 1244, 1829, 865, 3517, 2597, 3305, 2498, 1153, 877, 3453, 1038, 1118]. Three-Dimensional [1868, 865, 2597, 877, 1118]. Three-Valued [2498]. Threshold [101, 2081, 226, 252, 301, 227]. Thresholds [2562]. Throughput [2273, 822]. Throw [788]. Thue [309]. Thunkification [3052]. Tickets [3407]. Tidal [904]. Tight [1692, 1733, 1844, 2722, 2715]. Tightly [2605]. Tightly-Coupled [2605]. Tile [2672]. Tiling [345]. Tilings [1953]. Time [1762, 2282, 2559, 3104, 1913, 1909, 3402, 2838, 1386, 1628, 1921, 1210, 3015, 831, 3016, 2695, 1658, 2166, 1158, 2703, 3547, 869, 1843, 1923, 1500, 1197, 2270, 14, 3366, 662, 3018, 856, 2729, 31, 1618, 156, 2670, 1918, 2031, 178, 1732, 3006, 2357, 1922, 2704, 428, 1257, 299, 1214, 3023, 2392, 2305, 3026, 1829, 3027, 2016, 300, 302, 1156, 3029, 1311, 1188, 669, 324, 2050, 2154, 165, 2453, 2901, 2396, 1544, 2793, 1914, 3034, 1611, 1159, 548, 3544, 3526, 2400, 2794, 2700, 2671, 3038, 3553, 1559, 3257, 3092, 1081, 3467, 713, 2823, 3477, 3057]. Time-Critical [2166]. Time-Dependent [869]. Time-Message [428]. Time-Optimal [1544]. Time/Run [165].

Timed [1617, 1626, 3014, 1929, 1701, 2661, 1619, 767, 1654, 849, 3039, 3553, 1493, 2126, 1500, 2129, 2130]. Times [2560, 3185]. Timetable [3091]. Timetables [3149]. Timetabling [3090, 3148]. Timewise [1157]. Timing [2436, 3049]. TINA [2536, 2529]. TINA-C [2536, 2529]. tiny [723]. TIP [3189]. TMN [2530, 2523, 2525, 2522, 2531, 2524, 2542]. Toeplitz [3483]. Toeplitz-plus-Hankel [3483]. Tolerance [3307, 2558, 261, 284, 271, 299, 302, 2587, 263, 289, 293, 992, 2572, 2586]. Tolerant [2559, 3010, 297, 2557, 2606, 305, 2579, 1213, 2605, 2588, 2053, 2554, 3022, 273, 3027, 3007, 287, 3030, 265, 2584, 295, 2556, 278, 1962]. Tomorrow [1910]. Tool [2866, 493, 979, 2872, 1873, 2630, 1956, 1339, 836, 845, 54, 842, 835, 2469, 984, 584, 1596, 414, 2468, 2956, 2540, 1547, 2628, 3171, 1651, 3246, 843, 67, 3185]. Tool-Set [2468]. Toolkit [1767, 1480]. Tools [1683, 7, 2378, 829, 978, 3592, 938, 655, 1183, 2457, 2882, 71, 826, 8, 1659, 2118]. Toolset [841]. TOOPLE [1178]. TOOSQL [1803]. Top [2737, 261, 2786, 682, 2852, 1322, 291, 1776, 1482, 1473]. Top-Context-Free [682]. Top-down [2737, 1473]. TOP02 [2630]. Topological [1154, 1270, 1041]. Topologies [2597]. Topology [347, 1174, 1722, 2029, 3219, 1034]. topology-independent [1034]. TORI [2742]. Toroidal [1732]. Torus [2681]. Total [3290, 2501, 3297, 375, 2543, 1447]. Totality [3068]. Totally [2076]. Totally-Perfect [2076]. Tours [2710]. TPS [493]. TPTP [1433]. Trace [2140, 850, 175, 1173, 2317]. Trace-and-Replay [175]. Traces [344, 1666, 718]. Tracing [2237, 2774]. Tracking [2876, 2773, 2942, 1523, 1074, 1076, 1040]. Tractable [2869]. Trade [7, 428, 2464, 2133]. Trade-Off [2464, 2133]. Trade-Offs [428]. Trademark [1665].

Tradeoff [1551, 2248]. Trading [2413]. Traffic [568, 529, 3194, 3177, 156, 3585, 1282, 528, 1563, 2598, 1660, 1489, 821, 1138]. Train [3365]. Training [578, 2920, 3126, 2516, 2921, 2787, 3201, 2898, 3001]. Traitors [2237]. Trajectory [1639]. Trans [3202]. Trans-European [3202]. Transaction [445, 1911, 2878, 842, 2538, 446, 778, 1341, 1900]. Transaction-Level [2878]. Transactional [784, 777]. Transactions [1319, 1900]. transceiver [807]. Transducers [673, 365, 1247, 692]. Transductions [678, 1359, 684]. Transfer [2970, 3168, 2885, 2767, 3582, 3208, 823]. Transferring [3343]. Transfinitely [787]. Transform [3487, 2243, 2697, 1067, 1050, 820]. Transformation [1349, 157, 1521, 2990, 1179, 1389, 2626, 1820, 796, 2023, 797, 1809, 2326, 2489, 3417, 3564, 3034, 1335, 1998, 72, 1430, 637, 1069]. Transformation-Based [1820]. Transformational [1776]. Transformations [2198, 2360, 2347, 171, 2357, 3022, 394, 644, 166, 162, 403, 1340, 159, 1457]. Transformers [768, 3070]. Transforming [319, 3227]. Transforms [3075, 3530, 1109, 1065]. Transient [2250, 285, 861]. Transis [280]. Transition [1693, 2560, 1641, 1718, 362, 2114]. Transitions [2972, 767]. Transitive [1971]. Translates [1842]. Translating [631, 1306, 1881]. Translation [2988, 677]. translations [724]. Transmission [3586, 948, 533, 550, 820]. Transmitter [570]. Transparent [3184]. Transport [3189, 1351, 2401, 3194, 671, 905]. Transport-Triggering [671]. Transports [904]. Transputer [1016, 978, 914]. Transputer-Based [978, 914]. Transversal [2086]. Trapezoid [1845]. TRAPPER [980]. Trapping [653]. Traps [3249]. Traveling [3102, 3501, 2018, 2267].

Travelling [3113]. Treatment [3488, 302]. Tree [3482, 1542, 2017, 388, 2380, 678, 673, 1846, 750, 1358, 682, 3313, 604, 2061, 1240, 3555, 1252, 355, 687, 2748, 1708, 692, 1459, 1139]. Tree-Structured [1252]. Tree-Walking [692]. Tree-Width [388, 1846]. Trees [2198, 3010, 1251, 1543, 1837, 2694, 2724, 2286, 672, 2063, 1197, 2205, 476, 1848, 427, 1706, 1240, 2305, 2054, 608, 2070, 2413, 1702, 3549, 142, 2162, 2699, 1865, 709, 2749]. Treewidth [2322]. Trellis [1955, 573, 817, 1951, 816, 815]. Trellis-based [817]. trellis-coded [816, 815]. trial [3447]. Triangle [1975]. Triangles [2072]. Triangular [1211]. Triangulating [1840]. Triangulations [2692, 2034, 2062]. Tridiagonal [2633]. Tries [2696]. Triggering [671]. Trigram [2987]. Trilinearity [1135]. TROLL [630, 631]. True [748]. True-concurrency [748]. Truly [770]. Trust [3399, 1961, 135]. Trustworthy [2882]. Truth [2002, 350]. Truth-Table [350]. TSP [2263, 2267]. TSPLIB [2264]. TSPs [2262]. TTI [2995]. Tuned [1811]. Tuning [1596, 2173, 3327]. Tuplespace [2662]. Turbulence [2225]. Turing [1392, 2323]. Turning [3242]. Tutorial [2126, 2118, 54, 55]. Tutoring [2781]. Twisted [1217]. Two [1705, 2512, 322, 199, 4, 2569, 99, 1280, 370, 2292, 3158, 1591, 1946, 1871, 3507, 2302, 285, 33, 2033, 588, 2850, 13, 133, 2200, 3089, 351, 1616, 1916, 91, 2644, 1136, 134]. Two-Dimensional [1705, 1946, 1916]. Two-Semester [4]. Tyft [750]. Tyft/Tyxt [750]. Type [3047, 435, 3045, 674, 493, 2910, 3060, 3538, 1220, 514, 763, 2111, 1524, 986, 756, 476, 710, 3066, 3067, 1224, 336, 1695, 744, 1227, 790, 3403, 523, 1232, 732, 761, 1410, 402, 518, 1219, 697, 711, 1458, 288]. Type-Based [3067, 3066]. Type-labeled [2111]. Type-Theoretic [336]. Typeable [1520]. Typed [3050, 1178, 3418, 1395, 769, 1221, 1393, 272, 1526, 1724, 1397, 3312, 1997, 3541, 1746, 1764, 1897]. Typeflow [514]. Types [127, 1171, 3162, 762, 2105, 1222, 2381, 360, 731, 3316, 2581, 1403, 2925, 1996, 503, 3564, 761, 1337, 787, 714, 717, 624]. Typical [546]. Typing [731, 1566].

U [1914]. UCD [484]. UIS [928].

Ultimately [1175]. Ultrasonic [2912]. UMTS [563, 564]. Unacceptable [2394]. Unambiguity [1689]. Unary [3417]. Unbounded [1250]. Uncalibrated [1878, 1146, 1132, 1131]. Uncaught [3057]. Uncertain [3038]. Uncertainties [1264]. Uncertainty [2763, 2803, 2005, 178, 3540, 2805, 3152]. Unconstrained [177]. uncountable [2124]. Undecidable [1713, 751, 1993]. Undefined [755]. Underdetermined [3345]. Undergraduate [3, 4, 52, 33, 5]. Underlying [2159]. Understand [2734]. Understanding [3238, 2150, 3214]. Undirected [2874, 2036, 2700]. Unequal [1737]. Unexpected [2910, 1688]. Unfold [2505, 2360, 2347]. Unfold/Fold [2505, 2360, 2347]. Unfolding [2346]. Unforgeable [2235]. Unification [2964, 2199, 1778, 1460, 3555, 2965, 3002, 2195, 3556, 499, 262, 1464, 712, 1462, 1461, 1435, 1463, 2374, 2372]. Unification-Based [3002]. Unified [445, 2707, 425, 2725, 1922, 3082, 638, 177, 3190, 3304, 517, 2750, 2011, 2412, 808]. unifiers [1455]. Uniform [1250, 2902, 1802, 2833, 1721, 292, 398, 97, 1796, 2497]. Uniform-Length [1250]. Uniform-Modelling [398]. Unifying [2163, 1830, 336, 2025, 311, 726, 2134]. Unimonotone [1165]. Union [1976, 762]. Unions [368]. Uniprocessor [2328]. Uniquely [2021]. Uniqueness [3054, 402, 1093]. Unison [428]. Unit [2665, 1245, 1949, 1948, 804]. units [3466]. UNITY [2866, 466, 3563]. Universal [1942, 2505, 3211, 2108, 1170, 1464]. Universes [1149]. University [42, 2927]. UNIX [305, 2394, 989]. Unix-Based [989]. Unknown [1739, 76]. Unpredictable [129]. Unrecognizable [601]. Unreliable [1713]. Unstructured [871]. Unsupervised [1081, 611]. Untimed [1701]. Untraceable [243]. Untransferable [114]. Untyped [742, 1393]. Update [2280, 1610, 1313, 1563]. Update-Based [1610, 1563]. Updates [1939, 2972, 402]. Updating [2419, 2717, 2957, 2704, 3222]. upkeep [1081]. Upper [370, 3535, 1245, 547, 1688, 2028]. Upset [2565]. UPT [2546, 2549]. Upward [1970]. urban [1489]. Usability [3433]. USD [2108]. Use [3440, 288, 2531, 38, 2904, 851, 1871, 2897, 3569, 2459, 2210, 2547, 3114, 945, 3427, 2578, 403, 2898, 1884, 8, 1064, 815]. Used [2233, 3341, 2950]. User [1346, 2514, 3282, 51, 3170, 2526, 2513, 3435, 2608, 1767, 3284, 2674, 2897, 2916, 3437, 3429, 2396, 3434, 3430, 3425, 3279, 3276, 2898, 1659]. User-Adapted [3429]. User-Level [2608, 2396]. User-Oriented [1346]. Users [2914, 2897]. Using [1572, 435, 780, 2970, 157, 2828, 2924, 1981, 1251, 950, 3094, 860, 1522, 127, 667, 2471, 1608, 1626, 3015, 2353, 51, 581, 3491, 3106, 2463, 2112, 3121, 3087, 2990, 1670, 3309, 2645, 3150, 514, 2914, 1653, 3207, 2966, 2106, 94, 2911, 3362, 3352, 3369, 838, 2626, 1097, 856, 756, 3311, 2483, 2435, 1618, 2646, 1633, 1937, 914, 2437, 1648, 10, 3560, 2178, 2823, 3313, 875, 557, 1814, 1661, 3555, 2216, 505, 3568, 3285, 2468, 313, 3191, 1334, 260, 3135, 405, 1600, 610, 2666, 2834, 2885, 2549, 2575, 2100, 3146, 1610, 553, 2778, 3088, 2346, 2864, 2239, 2325, 2488]. Using [612, 1632, 2312, 2540, 2643, 571, 450, 1758, 1623, 3002, 3091, 2749, 3076, 2174, 778, 3209, 2977, 936, 3337, 2478, 1347, 1886, 478, 186, 2416, 1258, 2249, 2443, 985, 592, 1337, 2354, 295, 110, 1965, 3351, 3183, 870, 70, 2903, 2246, 2802, 849, 567, 2671, 2162, 3118, 2999, 2753, 926, 2012, 1882, 466, 1049, 2362, 1423, 1429, 3060, 1143, 2555, 1117, 1071, 1443, 1075, 3061, 823, 824, 3063, 1127, 818, 1102, 1059, 1067, 3065, 1056, 163, 1510, 1099, 503, 1511, 1047, 1054, 2114, 1037, 1091, 1108]. Utility [593, 1965]. Utilization [3185]. Utilizing [1342, 932, 2940, 1118].

V3.2 [1474]. Vague [2005]. Validating [2485, 3171]. Validation [1542, 3364, 1808, 1800, 1492, 3213]. validity [1427]. Valuation [3399]. Valuations [1689]. Value [1683, 3484, 960, 2531, 2189, 1775, 2689, 997, 307, 492, 688, 172, 217, 3050, 2143]. Value-Based [997, 172]. Value-Passing [492, 2143]. Valued [365, 2498, 301]. Values [2419, 2233, 2351, 2002, 1412]. Vapnik [304]. Variable [364, 816, 452, 1298, 1356, 2293, 2822, 2506, 3134, 2562, 1297]. Variable-Length [1297]. Variable-rate [816]. Variables [1784, 3014, 338, 2806, 1598, 490, 404, 76]. Variance [3105]. variant [3462]. Variants [2944, 3011]. Variation [3127]. variations [1139]. Varied [1667]. Varieties [1167, 3449]. Various [2902, 3316, 292, 2049, 2037]. Varying [1537, 560, 1087]. Vault [1606]. VBR [3585]. VC [2399, 310, 1090, 301, 616]. VC-Dimension [301, 1090]. VC-Dimensions [310, 616]. VDM [3342, 3343, 3358]. VDM-SL [3358]. ve [1444]. Vector [1967, 760, 1004, 920, 511, 339, 2649, 1559, 825]. Vectorization [2631]. Vectors [494]. Vegas [353]. vehicle [1107]. Vehicles [2937, 1068, 1138]. Verb [2096]. Verbalization [2945]. verbs [1107]. Verifiable [3351]. Verification [1713, 466, 484, 2436, 1631, 2455, 1621, 3013, 1626, 2120, 3539, 1622, 504, 2875, 1183, 3367, 2876, 1637, 1652, 1197, 3005, 2878, 3167, 1503, 471, 1618, 1633, 3056, 3021, 2882, 1186, 736, 1187, 3343, 2956, 764, 3336, 1528, 1632, 1623, 503, 1654, 2890, 495, 1642, 3036, 3390, 485, 1488, 2118]. Verified [2882, 3593]. Verifiers [222]. Verifying [3014, 3352, 3007, 2022, 2892, 487, 3470]. Versatile [2345]. Version [899, 3538, 2364, 592, 1463]. Versioning [1794, 1804]. Versions [2772, 2778, 668]. Versus [364, 1960, 349, 1962, 3400, 3318, 2510]. Vertex [1972, 2703, 377, 2307, 1729]. Vertex-Disjoint [2307]. Vertices [2039]. Very [951, 464, 627]. VHDL [2468, 2361]. Via [3069, 245, 1987, 3407, 1308, 158, 3219, 797, 1795, 1151]. Viable [1578]. Vibration [1088]. Video [3207, 1767, 455, 2410, 3192, 2408, 2409, 2405, 3583]. Video-on-Demand [2408, 2409]. Videocommunications [2515]. View [1684, 982, 393, 1808, 1343, 196, 808, 2139]. Viewing [2203, 2326, 2489]. viewpoint [1035]. Views [1667, 1888, 1684, 2849, 1878, 2093, 1338, 1885, 2103, 437, 1872, 1874, 1133, 1148, 1056, 1119, 1038]. VINO [1298]. Viper [484]. Virtual [2653, 2451, 2524, 442, 156, 976, 1605, 977, 2659]. Virtually [1612, 2817]. Virtually-Shared [1612]. Virtually-synchronous [2817]. vis [2057, 3273]. vis-a-vis [2057]. VIS-AD [3273]. Visibility [3548, 2648]. Visible [1843]. Vision [2550, 1873, 2444, 913, 62, 1112, 1144, 1111]. Vision-Based [62]. VISTA [166]. Visual [2899, 3431, 82, 166, 1040, 1347, 1324, 1042, 1135, 1065]. Visualisation [2629]. Visualization [3442, 938, 439, 2110, 1549, 3285, 3277, 937, 988, 1548, 3275, 179, 3280]. Visualizations [1550, 3279]. Visualize [2356]. Visualizing [3489, 1666, 3283]. Visually [914, 2897, 2927, 2928, 312, 2929]. Visually-Guided [914]. Viterbi [573]. Vivid [91]. VLIW [2666, 2655]. VLIWs [652]. VLSI [2378, 2579, 804, 3147, 2054, 481, 2568]. VLSI/WSI [2579]. VN [2399]. Voice [2910, 2479, 2901]. Voltan [295]. Volume [2647]. Voronoi [1839, 2708, 2289, 2016, 374]. Voss [480]. Voting [1541, 1371, 2249]. VP [2399]. VPN [2547]. Vs [3035, 3386, 202, 1790, 3424, 671, 2133, 1404, 409, 175, 3415].

W [1491]. Wait [2024, 2314, 1857, 2825]. Wait-Free [2024, 2314, 1857, 2825]. Waiting [831]. Walk [1627]. Walking [692]. Walks [686]. WALKY [2912]. Wallets [243, 122]. Wanna [2689]. Ward [3369]. Ward/Mellor [3369]. WASMII [2451]. waste [1508]. Watchdog [2575]. Water [2777]. Wave [3502]. Wavefront [2289]. Wavelet [940, 3530, 820]. Way [116, 2221, 2702, 1355, 1005, 1578, 2059, 1345, 1284, 3376, 2246]. Way-of-Working [1345]. WDM [1553]. Weak [236, 2877, 428, 1309, 2132, 1415, 1898]. weakening [3226]. Weakest [2622]. Weakly [2056, 2569, 1843]. Weakly-Visible [1843]. Weaknesses [96, 105]. Weather [899, 900, 927]. Week [21]. Weight [2693, 2062]. Weighted [1727, 1851, 3154, 2952, 2079, 1494]. Well [793, 1309, 1402, 1785, 1514, 1898]. Well-formedness [1514]. Well-Founded [1309, 1402, 1785, 1898]. Well-Ordering [793]. Western [904]. Where [515, 2317, 1379]. Which [2689, 718, 1507]. While [2670]. Whose [1635]. Width [388, 1846, 2307]. Wildcard [2504]. Wilf [680]. Win [2388]. Windows [2907]. Wireless [575, 568, 570, 550, 551, 553, 2908, 807]. Wisdom [3372]. Within [2924, 2089, 1428, 2914, 2524, 2772, 3206, 2753]. Without [569, 617, 3104, 1422, 129, 224, 223, 1063, 257, 1140]. Witness [2231]. witnesses [3445]. Women [2802]. WooRKS [1322]. Word [2293, 2445, 1367, 2977, 494]. Word-Oriented [2445]. Words [1350, 1175, 1737, 1362, 371]. Work [2717, 3205, 2966, 2729, 1323, 212, 2326, 2489, 2845, 1752, 1208, 1508]. Work-Optimal [1208]. Work-Time [2729]. Workbench [845, 1651, 3281]. Workflow [1323, 1322]. Working [1345, 936]. Workplaces [3172]. Workshop [2577]. Workstation [3480, 934, 2917, 1673, 2911, 861, 969, 973, 1595]. Workstations [3493, 3492, 995, 284, 3501, 875, 897, 877]. World [486, 12, 2535, 1347, 1786, 617, 3214]. Worlds [2773]. Wormhole [2599]. Worst [2594]. Worth [1962]. Write [2316, 3422]. Write-up [3422]. writer [2823, 2822]. Wu [132, 1443]. Wurttemberg [928].

x [3474, 2524, 1379]. X-Interface [2524]. X0T [1379]. XBLOX [2463]. XILINX [2471, 2439, 2462, 2443]. XP [1009]. XP/S [1009]. XP/S-15 [1009]. Xpress [3208]. XTP [3207]. Xunet [2398].

Y-MP [3520]. Year [42, 33]. Year-Long [33]. Years [15].

Z [3460, 3352, 3343, 3359]. ZAREPTA [2453]. Zero [2569, 224, 2453, 115, 221, 223, 220]. Zero-Knowledge [224, 115, 221, 223, 220]. Zoomable [1677]. ZPL [146].

References

Anonymous:1994:KA

[1] Anonymous. Keynote address. Lecture Notes in Computer Science, 750:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denning:1994:DDS

[2] Peter J. Denning. Designing a discipline of software design. Lecture Notes in Computer Science, 750:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1994:SUS

[3] Anonymous. Session 1: Undergraduate software engineering education. Lecture Notes in Computer Science, 750:3–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boardman:1994:TSU

[4] David B. Boardman and Aditya P. Mathur. A two-semester undergraduate sequence in software engineering: Architecture and experience. Lecture Notes in Computer Science, 750:5–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shepard:1994:SEU

[5] Terry Shepard. Software engineering in an undergraduate computer engineering program. Lecture Notes in Computer Science, 750:23–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Modesitt:1994:WGA

[6] Kenneth L. Modesitt. When the golden arches gang aft agley: Incorporating software engineering into computer science. Lecture Notes in Computer Science, 750:35–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1994:STT

[7] Anonymous. Session 2: Tools of the trade. Lecture Notes in Computer Science, 750:63–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zalewski:1994:CUC

[8] Janusz Zalewski. Cohesive use of commercial tools in a classroom. Lecture Notes in Computer Science, 750:65–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kermarrec:1994:ARS

[9] Yvon Kermarrec. Ada reusable software components for teaching distributed systems. Lecture Notes in Computer Science, 750:77–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horton:1994:UCC

[10] Thomas B. Horton. Using commercial CASE environments to teach software design. Lecture Notes in Computer Science, 750:97–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1994:SPO

[11] Anonymous. Session 3: Projectoriented courses. Lecture Notes in Computer Science, 750:117–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gotterbarn:1994:RWS

[12] Donald Gotterbarn and Robert Riser. Real-world software engineering: A spiral approach to a project-oriented course. Lecture Notes in Computer Science, 750:119–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moore:1994:LDG

[13] Melody Moore and Colin Potts. Learning by doing: Goals and experience of two software engineering project courses. Lecture Notes in Computer Science, 750:151–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Epp:1994:JTC

[14] Edward C. Epp. A “just in time” course framework: Project-oriented courses. Lecture Notes in Computer Science, 750:165–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersen:1994:PCN

[15] Rudolf Andersen, Reidar Conradi, John Krogstie, Guttorm Sindre, and Arne Solvberg. Project courses at the NTH: 20 years of experience. Lecture Notes in Computer Science, 750:177– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1994:SPI

[16] Anonymous. Session 4: Process issues. Lecture Notes in Computer Science, 750:189–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Werth:1994:ASP

[17] Laurie Honour Werth. An adventure in software process improvement. Lecture Notes in Computer Science, 750: 191–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robillard:1994:PSA

[18] Pierre N. Robillard, Jean Mayrand, and Jean-Normand Drouin. Process self-assessment in an educational context. Lecture Notes in Computer Science, 750:211–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deininger:1994:TSP

[19] Marcus Deininger and Kurt Schneider. Teaching software project management by simulation– experiences with a comprehensive model. Lecture Notes in Computer Science, 750:227–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1994:SIP

[20] Anonymous. Session 5: Implications of practitioners. Lecture Notes in Computer Science, 750:243–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weinberger:1994:AWT

[21] Barbara J. Weinberger. Awareness week at Texas Instruments: An alternative approach to instructor-led delivery. Lecture Notes in Computer Science, 750:245–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guyard:1994:IPP

[22] Jacques Guyard, Jean-Pierre Jacquot, Bernard Amar, and Yves Benoit. Implication of practitioners in a postgraduate curriculum, a successful collaboration. Lecture Notes in Computer Science, 750:251–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartvigsen:1994:RGB

[23] Gunnar Hartvigsen. Reducing the gap between academic theory and professional practice in software engineering education. Lecture Notes in Computer Science, 750:263–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1994:SPD

[24] Anonymous. Session 6: Panel discussion. Lecture Notes in Computer Science, 750:275–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naveda:1994:BG

[25] J. Fernando Naveda, John Beidler, James E. Cardow, Everald Mills, and Frances Van Scoy. Bridging the gaps. Lecture Notes in Computer Science, 750:277–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1994:SII

[26] Anonymous. Session 7: Industry influence in software engineering education. Lecture Notes in Computer Science, 750:283–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Purtilo:1994:ECD

[27] James M. Purtilo and Stan Siegel. Experiences with CCB-Directed projects in the classroom. Lecture Notes in Computer Science, 750:285–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habra:1994:PPA

[28] Naji Habra and Eric Dubois. Putting into practice advanced software engineering techniques through students project. Lecture Notes in Computer Science, 750:303–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prete:1994:CGS

[29] Cosimo Antonio Prete. Cachesim: A graphical software environment to support the teaching of computer systems with cache memories. Lecture Notes in Computer Science, 750:317–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calliss:1994:SSS

[30] Frank W. Calliss and Debra Trantina Calliss. Suggested scenarios of software maintenance education. Lecture Notes in Computer Science, 750:329– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gomaa:1994:SDM

[31] Hassan Gomaa. Software design methods for concurrent and real-time systems. Lecture Notes in Computer Science, 750:341–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1994:SEA

[32] Anonymous. Session 9: Experiences in academic education. Lecture Notes in Computer Science, 750:343–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johnston:1994:BEU

[33] Lorraine Johnston and Philip Dart. Building on experience: An undergraduate course with two year-long projects. Lecture Notes in Computer Science, 750:345–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prey:1994:SEB

[34] Jane C. Prey, James P. Cohoon, and Greg Fife. Software engineering beginning in the first computer science course. Lecture Notes in Computer Science, 750:359–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berztiss:1994:NFR

[35] Alfs T. Berztiss. Non-functional requirements in the design of software. Lecture Notes in Computer Science, 750:375–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartrum:1994:TFE

[36] Thomas C. Hartrum and Paul D. Bailor. Teaching formal extensions of informal-based object-oriented analysis methodologies. Lecture Notes in Computer Science, 750:389–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruegge:1994:TIC

[37] Bernd Bruegge and Robert F. Coyne. Teaching iterative and collaborative design: Lessons and directions. Lecture Notes in Computer Science, 750:411– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Epstein:1994:UCE

[38] Richard G. Epstein. The use of computer ethics scenarios in software engineering education: The case of the killer robot. Lecture Notes in Computer Science, 750:429–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maly:1994:CPI

[39] Kurt J. Maly, Dennis E. Ray, J. Christian Wild, Irwin B. Levinstein, Stephan Olariu, C. Michael Overstreet, Tijen Ireland, George Kantsios, and Nageswara S. V. Rao. Computer productivity initiative. Lecture Notes in Computer Science, 750:441–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1994:SPS

[40] Anonymous. Session 11: Perspective on software engineering education. Lecture Notes in Computer Science, 750:461–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paul:1994:AAS

[41] Jody Paul. Alternative assessment for software engineering education. Lecture Notes in Computer Science, 750: 463–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ammann:1994:FYP

[42] Paul Ammann, Hassan Gomaa, Jeff Offutt, David Rine, and Bo Sanden. A five year perspective on Software Engineering Graduate Programs at George Mason University. Lecture Notes in Computer Science, 750:473–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dalcher:1994:FPG

[43] Darren Dalcher. Falling down is part of growing up; the study of failure and the software engineering community. Lecture Notes in Computer Science, 750:489–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1994:SSR

[44] Anonymous. Session 12: Software reuse education. Lecture Notes in Computer Science, 750:497–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pei:1994:GRP

[45] Hsia Pei, Richard H. S. Wang, and Dave Chenho Kung. Guided reuse for programmers. Lecture Notes in Computer Science, 750:499–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cardow:1994:PAT

[46] James E. Cardow and William D. Watson. A practical approach to teaching software reuse. Lecture Notes in Computer Science, 750:517–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grechenig:1994:ISR

[47] Thomas Grechenig and Stefan Biffl. Introducing a software reuse culture in practice. Lecture Notes in Computer Science, 750:527–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1994:SSE

[48] Anonymous. Session 13: Software engineering education meeting industry needs. Lecture Notes in Computer Science, 750:541–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coyle:1994:MNI

[49] Frank P. Coyle, Edward Forest, Murat M. Tanik, and Dennis J. Frailey. Meeting the needs of industry: SMU’s Master’s Degree Program in Software Engineering. Lecture Notes in Computer Science, 750:543–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCarty:1994:HMY

[50] William B. McCarty and G. Thomas Plew. How mature is your software process? Lecture Notes in Computer Science, 750:555–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bralick:1994:UMU

[51] William A. Bralick and John S. Robinson. Using a multi-user dialogue system to support software engineering distance education. Lecture Notes in Computer Science, 750:565–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eward:1994:IMH

[52] Mary Eward and Steven Wartik. Introducing megaprogramming at the high school and undergraduate levels. Lecture Notes in Computer Science, 750: 583–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dodani:1994:FMS

[53] Mahesh H. Dodani. Formal methods for software engineering. Lecture Notes in Computer Science, 750:597– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gries:1994:STP

[54] David Gries and Fred Schneider. Saturday tutorial presentation teaching logic as a tool. Lecture Notes in Computer Science, 750:599–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paulk:1994:CMM

[55] Mark C. Paulk. The capability maturity model for software: A tutorial. Lecture Notes in Computer Science, 750:601–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pichler:1994:STE

[56] F. Pichler. Systems theory and engineering. Lecture Notes in Computer Science, 763:2–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oren:1994:CASa

[57] T. I. Oren. Computer-aided systems technology: Its role in advanced computerization. Lecture Notes in Computer Science, 763:11–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sommer:1994:CAN

[58] H. J. Sommer and H. Hahn. Computer aided nonlinear system design based on algebraic system representation and on nonlinear bundle graphs. Lecture Notes in Computer Science, 763:21–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monjau:1994:NMB

[59] D. Monjau, St. Kahlert, K. Buchenrieder, and Ch. Veith. A new modelbased approach to the co-design of heterogeneous systems. Lecture Notes in Computer Science, 763:42–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Creutzburg:1994:TEP

[60] R. Creutzburg, V. G. Labunets, and E. V. Labunets. Towards an “erlangen program” for general linear systems theory. Lecture Notes in Computer Science, 763:52–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rattray:1994:SCS

[61] Ch Rattray. The shape of complex systems. Lecture Notes in Computer Science, 763:72–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ylinen:1994:PST

[62] R. Ylinen. Polynomial systems theory for n-D systems applied to vision-based control. Lecture Notes in Computer Science, 763:83–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parets:1994:RSS

[63] J. Parets, A. Anaya, M. J. Rodriguez, and P. Paderewski. A representation of software systems evolution based on the theory of the general system. Lecture Notes in Computer Science, 763: 96–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fiol:1994:TCA

[64] G. Fiol and J. Miro-Nicolau. Theoretical considerations about subset descriptions. Lecture Notes in Computer Science, 763:111–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albertos:1994:SDP

[65] P. Albertos. Sampled data passive systems. Lecture Notes in Computer Science, 763:118–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oren:1994:CASb

[66] T. I. Oren and D. G. King. Computeraided systems technology for CAD environments: Complexity issues and a repository-based approach. Lecture Notes in Computer Science, 763:131– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaffer:1994:RCT

[67] Ch. Schaffer and H. Prahofer. On requirements for a CAST-Tool for complex, reactive system analysis, design and evaluation. Lecture Notes in Computer Science, 763:137–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hyotyniemi:1994:AMD

[68] H. Hyotyniemi. Automating the modeling of dynamic systems. Lecture Notes in Computer Science, 763:160– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Musgrave:1994:FMT

[69] G. Musgrave, S. Finn, M. Francis, R. Harris, and R. B. Hughes. Formal methods and their future. Lecture Notes in Computer Science, 763: 180–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thurner:1994:FDB

[70] E. M. Thurner. Formal description of bus interfaces using methods of system theory. Lecture Notes in Computer Science, 763:190–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacak:1994:CTI

[71] W. Jacak and J. Rozenblit. CAST tools for intelligent control in manufacturing automation. Lecture Notes in Computer Science, 763:203–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sierocki:1994:ATM

[72] I. Sierocki. An algebraic transformation of the minimum automaton identification problem. Lecture Notes in Computer Science, 763:220–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joslyn:1994:PA

[73] C. Joslyn. On possibilistic automata. Lecture Notes in Computer Science, 763:231–??, 1994. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Dormido:1994:AAS

[74] S. Dormido, I. Lopez, F. Morilla, and M. A. Canto. On automatic adjustment of the sampling period. Lecture Notes in Computer Science, 763:243– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mueller-Wipperfuerth:1994:FSR

[75] Th. Mueller-Wipperfuerth, J. Scharinger, and F. Pichler. FSM shift register realization for improved testability. Lecture Notes in Computer Science, 763:254–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ylinen:1994:CBM

[76] R. Ylinen. Cluster-based modelling of processes with unknown qualitative variables. Lecture Notes in Computer Science, 763:268–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Munoz-Blanco:1994:RPF

[77] J. A. Munoz-Blanco, J. C. Quevedo-Losada, and O. Bolivar-Toledo. The role of partitions and functionals in descriptor computation for data receptive fields. Lecture Notes in Computer Science, 763:282–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giorno:1994:SAC

[78] V. Giorno, A. G. Nobile, and L. M. Ricciardi. On some algorithmic and computational problems for neuronal diffusion models. Lecture Notes in Computer Science, 763:293–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kieronska:1994:HRS

[79] D. H. Kieronska and S. Venkatesh. Hierarchic representation for spatial knowledge. Lecture Notes in Computer Science, 763:314–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grossmann:1994:PMQ

[80] W. Grossmann and H. Werthner. Probabilistic models in qualitative simulation. Lecture Notes in Computer Science, 763:331–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Otero:1994:SMI

[81] R. P. Otero, A. Barreiro, H. Praehofer, F. Pichler, and J. Mira. STIMS-MEDTOOL: Integration of expert systems with systems modelling and simulation. Lecture Notes in Computer Science, 763:347–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Munoz-Blanco:1994:SCV

[82] J. A. Munoz-Blanco, C. Garcia, F. Alayon, and S. Candela. Systems concepts for visual texture change detection strategy. Lecture Notes in Computer Science, 763:357–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delgado:1994:QCN

[83] A. Delgado, R. Moreno-Diaz, and J. Mira. Qualitative computation with neural nets: Differential equations like examples. Lecture Notes in Computer Science, 763:366–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kilger:1994:IDL

[84] M. Kilger and T. Dietl. Interpretationdriven low-level parameter adaptation in scene analysis. Lecture Notes in Computer Science, 763:380–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zdrahal:1994:GST

[85] Z. Zdrahal. General systems theory as a framework for model-based diagnosis. Lecture Notes in Computer Science, 763:388–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kopacek:1994:TAK

[86] P. Kopacek. Technical applications of knowledge-based systems. Lecture Notes in Computer Science, 763:399– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zauner:1994:CCA

[87] M. Zauner. CASE — computer-aided systems engineering, a new approach for developing IM-Systems with special consideration of CIM systems. Lecture Notes in Computer Science, 763:406– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marshall:1994:MAC

[88] G. J. Marshall, A. Behrooz, and F. M. Clayton. Modelling and analysis of complex stochastic systems: from a specific example to a general environment. Lecture Notes in Computer Science, 763:418–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kitzmuller:1994:STA

[89] K. Kitzmuller. System theoretic approach to migration of project models. Lecture Notes in Computer Science, 763:427–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zenger:1994:CAA

[90] K. Zenger. Computer-aided analysis and design of sequential control in industrial processes. Lecture Notes in Computer Science, 763:440–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1994:VLK

[91] Gerd Wagner. Vivid logic. knowledgebased reasoning with two kinds of negation. Lecture Notes in Computer Science, 764:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johansson:1994:RBC

[92] Thomas Johansson, Gregory Kabatianskii, and Bernard Smeets. On the relation between A-codes and codes correcting independent errors. Lecture Notes in Computer Science, 765: 1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650001.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/0765/07650001.
pdf.

Safavi-Naini:1994:OAS

[93] R. Safavi-Naini and L. Tombak. Optimal authentication systems. Lecture Notes in Computer Science, 765:12– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650012.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650012.
pdf.

Dixon:1994:FIU

[94] Brandon Dixon and Arjen K. Lenstra. Factoring integers using SIMD sieves. Lecture Notes in Computer Science, 765:28–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0765/07650028.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0765/07650028.
pdf.

Demytko:1994:NEC

[95] N. Demytko. A new elliptic curve based analogue of RSA. Lecture Notes in Computer Science, 765:40– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650040.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650040.
pdf.

Blackburn:1994:WPK

[96] Simon Blackburn, Sean Murphy, and Jacques Stern. Weaknesses of a publickey cryptosystem based on factorizations of finite groups. Lecture Notes in Computer Science, 765:50– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650050.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650050.
pdf.

Nyberg:1994:DUM

[97] Kaisa Nyberg. Differentially uniform mappings for cryptography. Lecture Notes in Computer Science, 765:55– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650055.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650055.
pdf.

Beth:1994:APN

[98] Thomas Beth and Cunsheng Ding. On almost perfect nonlinear permutations. Lecture Notes in Computer Science, 765:65–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0765/07650065.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0765/07650065.
pdf.

Carlet:1994:TNC

[99] Claude Carlet. Two new classes of bent functions. Lecture Notes in Computer Science, 765:77–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650077.
pdf.

Cusick:1994:BFS

[100] Thomas W. Cusick. Boolean functions satisfying a higher order strict avalanche criterion. Lecture Notes in Computer Science, 765:102–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650102.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650102.
pdf.

Carpentieri:1994:SSP

[101] Marco Carpentieri, Alfredo De Santis, and Ugo Vaccaro. Size of shares and probability of cheating in threshold schemes. Lecture Notes in Computer Science, 765:118–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650118.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650118.
pdf.

Kurosawa:1994:NSS

[102] Kaoru Kurosawa, Koji Okada, Keiichi Sakano, Wakaha Ogata, and Shigeo Tsujii. Nonperfect secret sharing schemes and matroids. Lecture Notes in Computer Science, 765:126– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650126.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650126.
pdf.

Selmer:1994:MNC

[103] Ernst S. Selmer. From the memoirs of a Norwegian cryptologist. Lecture Notes in Computer Science, 765:142– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650142.
pdf.

Gottfert:1994:LCP

[104] Rainer Göttfert and Harald Niederreiter. On the linear complexity of products of shift-register sequences. Lecture Notes in Computer Science, 765:151– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650151.
pdf.

Daemen:1994:RWS

[105] Joan Daemen, René Govaerts, and Joos Vandewalle. Resynchronization weaknesses in synchronous stream ciphers. Lecture Notes in Computer Science, 765:159–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0765/07650159.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0765/07650159.
pdf.

Games:1994:BSS

[106] Richard A. Games and Joseph J. Rushanan. Blind synchronization of m-sequences with even span. Lecture Notes in Computer Science, 765:168– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650168.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650168.
pdf.

Seberry:1994:CNC

[107] Jennifer Seberry, Xian-Mo Zhang, and Yuliang Zheng. On constructions and nonlinearity of correlation immune functions. Lecture Notes in Computer Science, 765:181–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650181.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650181.
pdf.

Damgaard:1994:PPS

[108] Ivan Bjerre Damgård. Practical and provably secure release of a secret and exchange of signatures. Lecture Notes in Computer Science, 765:200– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650200.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650200.
pdf.

Damgard:1994:PPS

[109] I. B. Damgard. Practical and provably secure release of a secret and exchange of signatures. Lecture Notes in Computer Science, 765:200–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simmons:1994:SCE

[110] Gustavus J. Simmons. Subliminal communication is easy using the DSA. Lecture Notes in Computer Science, 765: 218–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650218.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650218.
pdf.

Naccache:1994:CRP

[111] David Naccache. Can O.S.S. be repaired? — proposal for a new practical signature scheme–. Lecture Notes in Computer Science, 765:233– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650233.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650233.
pdf.

Boyd:1994:LBL

[112] Colin Boyd and Wenbo Mao. On a limitation of BAN logic. Lecture Notes in Computer Science, 765:240– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650240.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650240.
pdf.

Park:1994:EAC

[113] Choonsik Park, Kazutomu Itoh, and Kaoru Kurosawa. Efficient anonymous channel and all/nothing election scheme. Lecture Notes in Computer Science, 765:248–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650248.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650248.
pdf.

Domingo-Ferrer:1994:URC

[114] Josep Domingo-Ferrer. Untransferable rights in a client-independent server environment. Lecture Notes in Computer Science, 765:260–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/0765/07650260.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650260.
pdf.

Ostrovsky:1994:IHS

[115] Rafail Ostrovsky, Ramarathnam Venkatesan, and Moti Yung. Interactive hashing simplifies zero-knowledge protocol design. Lecture Notes in Computer Science, 765:267–273, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650267.
pdf.

Benaloh:1994:OWA

[116] Josh Cohen Benaloh and Michael de Mare. One-way accumulators: A decentralized alternative to digital signatures. Lecture Notes in Computer Science, 765:274–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650274.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650274.
pdf.

Damgaard:1994:BAH

[117] Ivan Bjerre Damgård and Lars R. Knudsen. The breaking of the AR hash function. Lecture Notes in Computer Science, 765:286–292, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0765/07650286.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0765/07650286.
                                        pdf.

Boer:1994:CCF

[118] Bert den Boer and Antoon Bosselaers. Collisions for the compression function of MD-5. Lecture Notes in Computer Science, 765:293– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0765/07650293.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0765/07650293.
                                        pdf.

denBoer:1994:CCF

[119] Bert den Boer and Antoon Bosselaers. Collisions for the compression function of MD-5. Lecture Notes in Computer Science, 765:293– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0765/07650293.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0765/07650293.
                                        pdf.

Patarin:1994:HFA

[120] Jacques Patarin. How to find and avoid collisions for the knapsack hash function. Lecture Notes in Computer Science, 765:305–317, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/0765/07650305.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650305.
pdf.

Ferguson:1994:STL

[121] Niels Ferguson. Single term off-line coins. Lecture Notes in Computer Science, 765:318–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0765/07650318.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0765/07650318.pdf.

Cramer:1994:IPW

[122] Ronald J. F. Cramer and Torben P. Pedersen. Improved privacy in wallets with observers. Lecture Notes in Computer Science, 765:329–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650329.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650329.
pdf.

Brands:1994:DBP

[123] Stefan Brands and David Chaum. Distance bounding protocols. Lecture Notes in Computer Science, 765:344– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650344.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/0765/07650344.
                                        pdf.

OConnor:1994:DCB

[124] Luke O’Connor. On the distribution of characteristics in bijective mappings. Lecture Notes in Computer Science, 765:360–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0765/07650360.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0765/07650360.
                                        pdf.

Meier:1994:SIB

[125] Willi Meier. On the security of the IDEA block cipher. Lecture Notes in Computer Science, 765:371– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0765/07650371.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0765/07650371.
                                        pdf.

Matsui:1994:LCM

[126] Mitsuru Matsui. Linear cryptanalysis method for DES cipher. Lecture Notes in Computer Science, 765:386– 397, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0765/07650386.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0765/07650386.
                                        pdf.

Biham:1994:NTC

[127] Eli Biham. New types of cryptanalytic attacks using related keys. Lecture Notes in Computer Science, 765: 398–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650398.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650398.
pdf.

Brassard:1994:SKR

[128] Gilles Brassard and Louis Salvail. Secret key reconciliation by public discussion. Lecture Notes in Computer Science, 765:410–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650410.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650410.
pdf.

Beaver:1994:GUB

[129] Donald Beaver and Nicol So. Global, unpredictable bit generation without broadcast. Lecture Notes in Computer Science, 765:424–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650424.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650424.
pdf.

deRooij:1994:SPD

[130] Peter de Rooij. On Schnorr’s preprocessing for digital signature schemes. Lecture Notes in Computer Science, 765:435–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/0765/07650435.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/0765/07650435.
                                        pdf.

Rooij:1994:SPD

[131] Peter de Rooij. On Schnorr’s preprocessing for digital signature schemes. Lecture Notes in Computer Science, 765:435–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/0765/07650435.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/0765/07650435.
                                        pdf.

Burmester:1994:CCW

[132] Mike V. D. Burmester. Cryptanalysis of the Chang-Wu-Chen key distribution system. Lecture Notes in Computer Science, 765:440–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0765/07650440.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0765/07650440.
                                        pdf.

Oorschot:1994:AET

[133] Paul C. van Oorschot. An alternate explanation of two BANlogic “failures”. Lecture Notes in Computer Science, 765:443–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650443.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650443.
pdf.

vanOorschot:1994:AET

[134] Paul C. van Oorschot. An alternate explanation of two BANlogic “failures”. Lecture Notes in Computer Science, 765:443–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0765/07650443.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0765/07650443.
pdf.

Simmons:1994:CTS

[135] Gustavus J. Simmons. The consequences of trust in shared secret schemes. Lecture Notes in Computer Science, 765:448–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0765/07650448.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0765/07650448.pdf.

Hornauer:1994:MCA

[136] G. Hornauer, W. Stephan, and Ralph Wernsdorf. Markov ciphers and alternating groups. Lecture Notes in Computer Science, 765:453–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0765/07650453.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0765/07650453.
                                        pdf.

Park:1994:KDA

[137] Choonsik Park, Kaoru Kurosawa, Tatsuaki Okamoto, and Shigeo Tsujii. On key distribution and authentication in mobile radio networks. Lecture Notes in Computer Science, 765:461– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0765/07650461.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0765/07650461.
                                        pdf.

Anonymous:1994:AIa

[138] Anonymous. Author index. Lecture Notes in Computer Science, 765:467– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/0765/0765indx.pdf.

VanLoocke:1994:DCC

[139] Philip R. Van Loocke. The dynamics of concepts. A connectionist model. Lecture Notes in Computer Science, 766:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gogolla:1994:EER

[140] Martin Gogolla. An extended entityrelationship model fundamentals and pragmatics. Lecture Notes in Computer Science, 767:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carlisle:1994:EEO

[141] Martin C. Carlisle, Anne Rogers, John H. Reppy, and Laurie J. Hendren. Early experiences with olden. Lecture Notes in Computer Science, 768:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Solworth:1994:AOO

[142] Jon A. Solworth and Bryan B. Reagan. Arbitrary order operations on trees. Lecture Notes in Computer Science, 768:21–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plevyak:1994:ADS

[143] John Plevyak, Andrew A. Chien, and Vijay Karamcheti. Analysis of dynamic structures for efficient parallel execution. Lecture Notes in Computer Science, 768:37–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bik:1994:ADS

[144] Aart J. C. Bik and Harry A. G. Wijshoff. On automatic data structure selection and code generation for sparse computations. Lecture Notes in Computer Science, 768:57–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prakash:1994:SID

[145] Sundeep Prakash, Maneesh Dhagat, and Rajive Bagrodia. Synchronization issues in data-parallel languages. Lecture Notes in Computer Science, 768: 76–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1994:ZAS

[146] Calvin Lin and Lawrence Snyder. ZPL: An array sublanguage. Lecture Notes in Computer Science, 768:96–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Browne:1994:EBC

[147] Raju Browne and James C. Pandey. Event-based composition of concurrent programs. Lecture Notes in Computer Science, 768:115–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carriero:1994:APM

[148] Nicholas Carriero, Eric Freeman, and David Gelernter. Adaptive parallelism on multiprocessors: Preliminary experience with Piranha on the CM-5. Lecture Notes in Computer Science, 768:139–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Das:1994:SAI

[149] Raja Das, Joel Saltz, and Reinhard von Hanxleden. Slicing analysis and indirect accesses to distributed arrays. Lecture Notes in Computer Science, 768:152–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bwolen:1994:DMI

[150] Yang Bwolen, Jon Webb, David R. Stichnoth, James M. O’Hallaron, and Thomas Gross. Do and merge: Integrating parallel loops and reductions. Lecture Notes in Computer Science, 768:169–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chapman:1994:ASD

[151] Barbara M. Chapman, Thomas Fahringer, and Hans P. Zima. Automatic support for data distribution on distributed memory multiprocessor systems. Lecture Notes in Computer Science, 768:184–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bozkus:1994:CAF

[152] Zeki Bozkus, Alok Choudhary, Geoffrey Fox, Tomasz Haupt, and Sanjay Ranka. A compilation approach for Fortran 90D/HPF compilers. Lecture Notes in Computer Science, 768:200– 215, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1994:FED

[153] Manish Gupta and Edith Schonberg. A framework for exploiting data availability to optimize communication. Lecture Notes in Computer Science, 768: 216–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chatterjee:1994:ADG

[154] Siddhartha Chatterjee, Robert Schreiber, and John R. Gilbert. The alignmentdistribution graph. Lecture Notes in Computer Science, 768:234–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amarasinghe:1994:OCS

[155] Saman P. Amarasinghe, Jennifer M. Anderson, Monica S. Lam, and Amy W. Lim. An overview of a compiler for scalable parallel machines. Lecture Notes in Computer Science, 768:253–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Granston:1994:TCT

[156] Elana D. Granston. Toward a compiletime methodology for reducing false sharing and communication traffic in shared virtual memory systems. Lecture Notes in Computer Science, 768: 273–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Appelbe:1994:PTL

[157] Bill Appelbe, Charles Hardnett, and Sri Doddapaneni. Program transformation for locality using affinity regions. Lecture Notes in Computer Science, 768:290–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kennedy:1994:MLP

[158] Ken Kennedy and Kathryn S. McKinley. Maximizing loop parallelism and improving data locality via loop fusion and distribution. Lecture Notes in Computer Science, 768:301–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Torres:1994:ADB

[159] Jordi Torres, Eduard Ayguade, Jesus Labarta, and Mateo Valero. Align and distribute-based linear loop transformations. Lecture Notes in Computer Science, 768:321–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gao:1994:ESP

[160] Guang R. Gao, Ning Qi, and Vincent Van Dongen. Extending software pipelining techniques for scheduling nested loops. Lecture Notes in Computer Science, 768:340–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaushik:1994:MGE

[161] S. D. Kaushik, C.-H. Huang, P. Sadayappan, and R. W. Johnson. A methodology for generating efficient disk-based algorithms from tensor product formulas. Lecture Notes in Computer Science, 768:358–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sehr:1994:LTP

[162] David C. Sehr, Laxmikant V. Kale, and David A. Padua. Loop transformations for Prolog programs. Lecture Notes in Computer Science, 768:374–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nikhil:1994:MII

[163] Rishiyur S. Nikhil. A multithreaded implementation of Id using P-RISC graphs. Lecture Notes in Computer Science, 768:390–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schlansker:1994:AFH

[164] Michael Schlansker and Vinod Kathail. Acceleration of first and higher order recurrences on processors with instruction level parallelism. Lecture Notes in Computer Science, 768:406–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neves:1994:ECT

[165] Richard Neves and Robert B. Schnabel. Efficient compile-time/run-time contraction of fine grain data parallel codes. Lecture Notes in Computer Science, 768:430–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Novack:1994:VVI

[166] Steven Novack and Alexandru Nicolau. VISTA: The visual interface for scheduling transformations and analysis. Lecture Notes in Computer Science, 768:449–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cytron:1994:ECN

[167] Ron K. Cytron and Jeanne Ferrante. Efficiently computing o-nodes on-the-fly. Lecture Notes in Computer Science, 768:461–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Havlak:1994:CTG

[168] Paul Havlak. Construction of thinned gated single-assignment form. Lecture Notes in Computer Science, 768:477– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peng:1994:AAP

[169] Tu Peng and David Padua. Automatic array privatization. Lecture Notes in Computer Science, 768:500–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grunwald:1994:ECP

[170] Dirk Grunwald and Harini Srinivasan. Efficient computation of precedence information in parallel programs. Lecture Notes in Computer Science, 768: 502–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hall:1994:FFI

[171] Mary W. Hall, John M. Mellor-Crummey, Alan Carle, and Rene G. Rodriguez. FIAT: A framework for interprocedural analysis and transformations. Lecture Notes in Computer Science, 768:522–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pugh:1994:EMA

[172] William Pugh and David Wonnacott. An exact method for analysis of valuebased array data dependences. Lecture Notes in Computer Science, 768:546– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haghighat:1994:SAB

[173] Mohammad R. Haghighat and Constantine D. Polychronopoulos. Symbolic analysis: A basis for parallelization, optimization, and scheduling of programs. Lecture Notes in Computer Science, 768:567–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1994:TNI

[174] Rajiv Gupta and Madalene Spezialetti. Towards a non-intrusive approach for monitoring distributed computations through perturbation analysis. Lecture Notes in Computer Science, 768:586– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Netzer:1994:TSV

[175] Robert H. B. Netzer. Trace size vs. parallelism in trace-and-replay debugging of shared-memory programs. Lecture Notes in Computer Science, 768: 617–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sarkar:1994:PPG

[176] Vivek Sarkar and Barbara Simons. Parallel program graphs and their classification. Lecture Notes in Computer Science, 768:633–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nazareth:1994:NCF

[177] J. L. Nazareth. The Newton-Cauchy framework: A unified approach to unconstrained nonlinear minimization. Lecture Notes in Computer Science, 769:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haddawy:1994:RPU

[178] Peter Haddawy. Representing plans under uncertainty A logic of time, chance, and action. Lecture Notes in Computer Science, 770:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tomas:1994:VSP

[179] Gerald Tomas and Christoph W. Ueberhuber. Visualization of scientific parallel programs. Lecture Notes in Computer Science, 771:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Estublier:1994:AS

[180] J. Estublier. Architecture session. Lecture Notes in Computer Science, 772: 1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambriola:1994:OSO

[181] V. Ambriola, G. A. Cignoni, and C. Montangero. The Oikos services for object management in the software process. Lecture Notes in Computer Science, 772:2–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bandinelli:1994:ASP

[182] S. Bandinelli, M. Braga, A. Fuggetta, and L. Lavazza. The architecture of SPADE-1 process-centered SEE. Lecture Notes in Computer Science, 772: 15–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1994:NSP

[183] M. Anderson and P. Griffiths. The nature of the software process modelling problem is evolving. Lecture Notes in Computer Science, 772:31–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldi:1994:SPP

[184] M. Baldi, S. Gai, M. L. Jaccheri, and P. Lago. Searching for PMIPS: Process model instructions per second. Lecture Notes in Computer Science, 772:35–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conradi:1994:MPM

[185] R. Conradi. Meta-process / methodology session. Lecture Notes in Computer Science, 772:38–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sa:1994:MPU

[186] J. Sa and B. C. Warboys. Modelling processes using a stepwise refinement technique. Lecture Notes in Computer Science, 772:40–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rodden:1994:PMD

[187] T. Rodden, V. King, J. Hughes, and I. Sommerville. Process modelling and development practice. Lecture Notes in Computer Science, 772:59–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Phalp:1994:PAP

[188] K. Phalp and M. Shepperd. A pragmatic approach to process modelling. Lecture Notes in Computer Science, 772:65–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Junkermann:1994:DMP

[189] G. Junkermann and W. Schafer. A design methodology for processprogramming. Lecture Notes in Computer Science, 772:69–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Derniame:1994:PMC

[190] J.-C. Derniame. Process modelling concepts session. Lecture Notes in Computer Science, 772:74–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heineman:1994:PMC

[191] G. T. Heineman. Process modelling with cooperative agents. Lecture Notes in Computer Science, 772:75–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dowson:1994:TRE

[192] M. Dowson and C. Fernstrom. Towards requirements for enactment mechanisms. Lecture Notes in Computer Science, 772:90–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perry:1994:ECI

[193] D. E. Perry. Enactment control in interact/intermediate. Lecture Notes in Computer Science, 772:107–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sommerville:1994:SIS

[194] I. Sommerville and S. Monk. Supporting informality in the software process. Lecture Notes in Computer Science, 772:114–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chroust:1994:NPM

[195] G. Chroust. Navigation in process models. Lecture Notes in Computer Science, 772:119–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verlage:1994:MVM

[196] M. Verlage. Multi-view modelling of software processes. Lecture Notes in Computer Science, 772:123–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuggetta:1994:PCP

[197] A. Fuggetta. PML concepts and paradigms session. Lecture Notes in Computer Science, 772:127–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engels:1994:SCB

[198] G. Engels and L. Groenewegen. Specification of coordinated behaviour by SOCCA. Lecture Notes in Computer Science, 772:128–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berrington:1994:DCI

[199] N. Berrington, D. De Roure, R. Greenwood, and P. Henderson. Distribution and change: Investigating two challenges for process enactment systems. Lecture Notes in Computer Science, 772:152–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Starke:1994:WPM

[200] G. Starke. Why is process modelling so difficult? Lecture Notes in Computer Science, 772:163–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greenwood:1994:MPC

[201] R. M. Greenwood. Modelling processes with constraints. Lecture Notes in Computer Science, 772:167–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arbaoui:1994:GOV

[202] S. Arbaoui and F. Oquendo. Goal oriented vs. activity oriented process modelling and enactment: Issues and perspectives. Lecture Notes in Computer Science, 772:171–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tankoano:1994:SPD

[203] J. Tankoano, J.-C. Derniame, and A. B. Kaba. Software process design based on products and the object oriented paradigm. Lecture Notes in Computer Science, 772:177–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sommerville:1994:ESP

[204] I. Sommerville. Experiences with software process technology session. Lecture Notes in Computer Science, 772: 186–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robertson:1994:IIP

[205] I. Robertson. An implementation of the ISPW-6 process example. Lecture Notes in Computer Science, 772:187– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambriola:1994:AMF

[206] V. Ambriola, R. Di Meglio, V. Gervasi, and B. Mercurio. Applying a metric framework to the software process: an experiment. Lecture Notes in Computer Science, 772:207–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aumaitre:1994:LLF

[207] J.-M. Aumaitre, M. Dowson, and D.- R. Harjani. Lessons learned from formalizing and implementing a large process model. Lecture Notes in Computer Science, 772:227–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kramer:1994:APT

[208] B. Kramer and B. Dinler. Applying process technology to hardware design. Lecture Notes in Computer Science, 772:240–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambriola:1994:RDS

[209] V. Ambriola. Related domains session. Lecture Notes in Computer Science, 772:249–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gruhn:1994:SPM

[210] V. Gruhn. Software process management and business process (re- )engineering. Lecture Notes in Computer Science, 772:250–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conradi:1994:CMF

[211] R. Conradi, G. M. Hoydalsvik, and G. Sindre. A comparison of modelling frameworks for software processes and information systems. Lecture Notes in Computer Science, 772:254–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lonchamp:1994:PCF

[212] J. Lonchamp. A process-centered framework for asynchronous collaborative work. Lecture Notes in Computer Science, 772:261–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Estublier:1994:WPT

[213] J. Estublier. What process technology needs from databases. Lecture Notes in Computer Science, 772:270–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shamir:1994:ESS

[214] Adi Shamir. Efficient signature schemes based on birational permutations. Lecture Notes in Computer Science, 773:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0773/07730001.
htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
0773/07730001.pdf.

Stern:1994:NIS

[215] Jacques Stern. A new identification scheme based on syndrome decoding. Lecture Notes in Computer Science, 773:13–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0773/07730013.
pdf.

Coppersmith:1994:SG

[216] Don Coppersmith, Hugo Krawczyk, and Yishay Mansour. The shrinking generator. Lecture Notes in Computer Science, 773:22–39, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0773/07730022.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0773/07730022.pdf.

Taylor:1994:ICV

[217] Richard Taylor. An integrity check value algorithm for stream ciphers. Lecture Notes in Computer Science, 773: 40–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730040.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0773/07730040.
pdf.

Seberry:1994:NBB

[218] Jennifer Seberry, Xian-Mo Zhang, and Yuliang Zheng. Nonlinearly balanced Boolean functions and their propagation characteristics. Lecture Notes in Computer Science, 773:49– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0773/07730049.
pdf.

Itoh:1994:LCC

[219] Toshiya Itoh, Masafumi Hoshi, and Shigeo Tsujii. A low communication competitive interactive proof system for promised quadratic residuosity. Lecture Notes in Computer Science, 773: 61–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0773/07730061.
pdf.

DeSantis:1994:SSP

[220] A. De Santis, G. Di Crescenzo, and G. Persiano. Secret sharing and perfect zero-knowledge. Lecture Notes in Computer Science, 773:73–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santis:1994:SSP

[221] Alfredo De Santis, Giovanni Di Crescenzo, and Giuseppe Persiano. Secret sharing and perfect zeroknowledge. Lecture Notes in Computer Science, 773:73–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0773/07730073.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0773/07730073.pdf.

Aumann:1994:OMP

[222] Yonatan Aumann and Uriel Feige. One message proof systems with known space verifiers. Lecture Notes in Computer Science, 773:85–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730085.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0773/07730085.
pdf.

Damgard:1994:IHC

[223] Ivan B. Damgard. Interactive hashing can simplify zero-knowledge protocol design without computational assumptions. Lecture Notes in Computer Science, 773:100–109, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damgaard:1994:IHC

[224] Ivan Bjerre Damgård. Interactive hashing can simplify zero-knowledge protocol design without computational assumptions. Lecture Notes in Computer Science, 773:100–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730100.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/0773/07730100.
pdf.

Blundo:1994:FDS

[225] Carlo Blundo, Antonella Cresti, Alfredo De Santis, and Ugo Vaccaro. Fully dynamic secret sharing schemes. Lecture Notes in Computer Science, 773:110–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0773/07730110.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0773/07730110.
pdf.

Jackson:1994:MTS

[226] Wen-Ai Jackson, Keith M. Martin, and Christine M. O’Keefe. Multisecret threshold schemes. Lecture Notes in Computer Science, 773:126– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730126.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0773/07730126.
pdf.

Wen-Ai:1994:MTS

[227] Jackson Wen-Ai, Keith M. Martin, and Christine M. O’Keefe. Multisecret threshold schemes. Lecture Notes in Computer Science, 773:126–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krawczyk:1994:SSM

[228] Hugo Krawczyk. Secret sharing made short. Lecture Notes in Computer Science, 773:136–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730136.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0773/07730136.
pdf.

Adleman:1994:SAD

[229] Leonard M. Adleman and Johnathan DeMarrais. A subexponential algorithm for discrete logarithms over all finite fields. Lecture Notes in Computer Science, 773:147–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730147.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0773/07730147.
pdf.

Buchmann:1994:IGN

[230] Johannes A. Buchmann, J. Loho, and J. Zayer. An implementation of the general number field sieve. Lecture Notes in Computer Science, 773:159– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730159.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0773/07730159.
pdf.

Denny:1994:FR

[231] T. Denny, Bruce Dodson, Arjen K. Lenstra, and Mark S. Manasse. On the factorization of RSA-120. Lecture Notes in Computer Science, 773:166– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730166.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0773/07730166.
pdf.

Bosselaers:1994:CTM

[232] Antoon Bosselaers, René Govaerts, and Joos Vandewalle. Comparison of three modular reduction functions. Lecture Notes in Computer Science, 773:175– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730175.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0773/07730175.
pdf.

Ben-Aroya:1994:DCL

[233] Ishai Ben-Aroya and Eli Biham. Differential cryptanalysis of Lucifer. Lecture Notes in Computer Science, 773: 187–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730187.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0773/07730187.
pdf.

Ohta:1994:DAM

[234] Kazuo Ohta and Mitsuru Matsui. Differential attack on message authentication codes. Lecture Notes in Computer Science, 773:200–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0773/07730200.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0773/07730200.
                                        pdf.

Preneel:1994:CCM

[235] Bart Preneel, Marnix Nuttin, Vincent Rijmen, and Johan Buelens. Cryptanalysis of the CFB mode of the DES with a reduced number of rounds. Lecture Notes in Computer Science, 773: 212–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0773/07730212.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0773/07730212.
                                        pdf.

Daemen:1994:WKI

[236] Joan Daemen, René Govaerts, and Joos Vandewalle. Weak keys for IDEA. Lecture Notes in Computer Science, 773: 224–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0773/07730224.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0773/07730224.
                                        pdf.

Bellare:1994:EAK

[237] Mihir Bellare and Phillip Rogaway. Entity authentication and key distribution. Lecture Notes in Computer Science, 773:232–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730232.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0773/07730232.
pdf.

Damgaard:1994:ESH

[238] Ivan Bjerre Damgård, Torben Pryds Pedersen, and Birgit Pfitzmann. On the existence of statistically hiding bit commitment schemes and fail-stop signatures. Lecture Notes in Computer Science, 773:250–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730250.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0773/07730250.
pdf.

Damgard:1994:ESH

[239] Ivan B. Damgard, Torben P. Pedersen, and Birgit Pfitzmann. On the existence of statistically hiding bit commitment schemes and fail-stop signatures. Lecture Notes in Computer Science, 773:250–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franklin:1994:JEM

[240] Matthew Franklin and Stuart Haber. Joint encryption and message-efficient secure computation. Lecture Notes in Computer Science, 773:266–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730266.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0773/07730266.
                                        pdf.

Blum:1994:CPB

[241] Avrim Blum, Merrick Furst, Michael Kearns, and Richard J. Lipton. Cryptographic primitives based on hard learning problems. Lecture Notes in Computer Science, 773:278–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0773/07730278.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0773/07730278.
                                        pdf.

Ferguson:1994:EST

[242] Niels Ferguson. Extensions of single-term coins. Lecture Notes in Computer Science, 773:292–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0773/07730292.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0773/07730292.
                                        pdf.

Brands:1994:ULC

[243] Stefan Brands. Untraceable off-line cash in wallets with observers. Lecture Notes in Computer Science, 773: 302–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0773/07730302.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/0773/07730302.
pdf.

Crepeau:1994:DSG

[244] Claude Crépeau and Joe Kilian. Discreet solitary games. Lecture Notes in Computer Science, 773:319–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730319.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0773/07730319.
pdf.

Bierbrauer:1994:FHF

[245] Jürgen Bierbrauer, Thomas Johansson, Gregory A. Kabatianskii, and Bernard J. M. Smeets. On families of hash functions via geometric codes and concatenation. Lecture Notes in Computer Science, 773:331– 342, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730331.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0773/07730331.
pdf.

Johansson:1994:CPA

[246] Thomas Johansson. On the construction of perfect authentication codes that permit arbitration. Lecture Notes in Computer Science, 773:343– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730343.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/0773/07730343.
                                        pdf.

Gemmell:1994:CIA

[247] Peter Gemmell and Moni Naor. Codes for interactive authentication. Lecture Notes in Computer Science, 773:355– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0773/07730355.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0773/07730355.
                                        pdf.

Preneel:1994:HFB

[248] Bart Preneel, René Govaerts, and Joos Vandewalle. Hash functions based on block ciphers: A synthetic approach. Lecture Notes in Computer Science, 773:368–378, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0773/07730368.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0773/07730368.
                                        pdf.

Hohl:1994:SIH

[249] Walter Hohl, Xuejia Lai, Thomas Meier, and Christian Waldvogel. Security of iterated hash functions based on block ciphers. Lecture Notes in Computer Science, 773:379–390, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0773/07730379.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/0773/07730379.
pdf.

Patarin:1994:IAP

[250] Jacques Patarin and Pascal Chauvaud. Improved algorithms for the permuted kernel problem. Lecture Notes in Computer Science, 773:391– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730391.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0773/07730391.
pdf.

OConnor:1994:DCC

[251] Luke O’Connor. On the distribution of characteristics in composite permutations. Lecture Notes in Computer Science, 773:403–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730403.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0773/07730403.
pdf.

Li:1994:RTR

[252] Chuan-Ming Li, Tzonelih Hwang, and Narn-Yih Lee. Remark on the threshold RSA signature scheme. Lecture Notes in Computer Science, 773:413– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730413.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/0773/07730413.
                                        pdf.

Chae:1994:AMA

[253] Hoon Lim Chae and Joong Lee Pil. Another method for attaining security against adaptively chosen ciphertext attacks. Lecture Notes in Computer Science, 773:420–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lim:1994:AMA

[254] Chae Hoon Lim and Pil Joong Lee. Another method for attaining security against adaptively chosen ciphertext attacks. Lecture Notes in Computer Science, 773:420–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0773/07730420.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0773/07730420.
                                        pdf.

Coppersmith:1994:ABP

[255] Don Coppersmith, Jacques Stern, and Serge Vaudenay. Attacks on the birational permutation signature schemes. Lecture Notes in Computer Science, 773:435–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0773/07730435.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0773/07730435.
                                        pdf.

Beimel:1994:IKD

[256] Amos Beimel and Benny Chor. Interaction in key distribution schemes. Lecture Notes in Computer Science, 773: 444–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730444.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0773/07730444.
pdf.

Leighton:1994:SKA

[257] Tom Leighton and Silvio Micali. Secret-key agreement without publickey cryptography. Lecture Notes in Computer Science, 773:456–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730456.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0773/07730456.
pdf.

Fiat:1994:BE

[258] Amos Fiat and Moni Naor. Broadcast encryption. Lecture Notes in Computer Science, 773:480–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0773/07730480.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0773/07730480.
pdf.

Anonymous:1994:AIb

[259] Anonymous. Author index. Lecture Notes in Computer Science, 773:492– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/0773/0773indx.pdf.

Laird:1994:IUP

[260] P. Laird. Identifying and using patterns in sequential data. Lecture Notes in Computer Science, 774:1–18, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chillarege:1994:TFC

[261] Ram Chillarege. Top five challenges facing the practice of fault tolerance. Lecture Notes in Computer Science, 774:3–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamamoto:1994:GUB

[262] A. Yamamoto. Generalized unification as background knowledge in learning logic programs. Lecture Notes in Computer Science, 774:11–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levendel:1994:FTC

[263] Y. Levendel. Fault tolerance cost effectiveness. Lecture Notes in Computer Science, 774:13–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miyano:1994:LTT

[264] S. Miyano. Learning theory toward genome informatics. Lecture Notes in Computer Science, 774:19–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Locke:1994:FTA

[265] C. Douglass Locke. Fault tolerant applications systems; A requirements perspective. Lecture Notes in Computer Science, 774:21–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Banatre:1994:SSM

[266] Michel Banatre, Alain Gefflaut, and Christine Morin. Scalable shared memory multiprocessors: Some ideas to make them reliable. Lecture Notes in Computer Science, 774:29–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muggleton:1994:OLL

[267] S. Muggleton. Optimal layered learning: A PAC approach to incremental sampling. Lecture Notes in Computer Science, 774:37–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arima:1994:REL

[268] J. Arima and H. Sawamura. Reformulation of explanation by linear logic — toward logic for explanation. Lecture Notes in Computer Science, 774:45–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuchs:1994:ACA

[269] W. Kent Fuchs, Wen mei Hwu, and Neal J. Alewine. Application of compiler-assisted rollback recovery to speculative execution repair. Lecture Notes in Computer Science, 774:45–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barzdins:1994:TEIa

[270] J. Barzdins, G. Barzdins, K. Apsitis, and U. Sarkans. Towards efficient inductive synthesis of expressions from input/output examples. Lecture Notes in Computer Science, 774:59–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gleeson:1994:FTW

[271] Barry J. Gleeson. Fault tolerance: Why should I pay for it? Lecture Notes in Computer Science, 774:66–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hagiya:1994:TLC

[272] M. Hagiya. A typed lambda-calculus for proving-by-example and bottomup generalization procedure. Lecture Notes in Computer Science, 774:73–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1994:SDF

[273] Jeremy Jones and Brian Coghlan. Stable disk — A fault tolerant cached RAID subsystem. Lecture Notes in Computer Science, 774:78–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jantke:1994:CBR

[274] K. P. Jantke and S. Lange. Case-based representation and learning of pattern languages. Lecture Notes in Computer Science, 774:87–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liddell:1994:SDM

[275] David Liddell. Simple design makes reliable computers. Lecture Notes in Computer Science, 774:91–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pradhan:1994:RFC

[276] Dhiraj K. Pradhan, Debendra Das Sharma, and Nitin H. Vaidya. Rollforward checkpointing schemes. Lecture Notes in Computer Science, 774:95–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sato:1994:IR

[277] T. Sato and S. Akiba. Inductive resolution. Lecture Notes in Computer Science, 774:101–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stiffler:1994:FTA

[278] Jack J. Stiffler. Fault tolerant architectures — past, present and (?) future. Lecture Notes in Computer Science, 774:117–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mukouchi:1994:IIM

[279] Y. Mukouchi and S. Arikawa. Inductive inference machines that can refute hypothesis spaces. Lecture Notes in Computer Science, 774:123–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amir:1994:HAA

[280] Ofir Amir, Yair Amir, and Danny Dolev. A highly available application in the Transis environment. Lecture Notes in Computer Science, 774:125– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freivalds:1994:DBM

[281] R. Freivalds and C. H. Smith. On the duality between mechanistic learners and what it is they learn. Lecture Notes in Computer Science, 774:137– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Birman:1994:REC

[282] Kenneth P. Birman. Reliable enterprise computing systems. Lecture Notes in Computer Science, 774:140–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jain:1994:ATL

[283] S. Jain and A. Sharma. On aggregating teams of learning machines. Lecture Notes in Computer Science, 774: 150–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elnozahy:1994:FTC

[284] Elmootazbellah N. Elnozahy. Fault tolerance for clusters of workstations. Lecture Notes in Computer Science, 774: 151–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1994:TTT

[285] Yennun Huang, Chandra Kintala, and Pankaj Jalote. Two techniques for transient software error recovery. Lecture Notes in Computer Science, 774:159– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Viksna:1994:LGQ

[286] J. Viksna. Learning with growing quality. Lecture Notes in Computer Science, 774:164–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1994:SFR

[287] Peter A. Lee. Software-faults: The remaining problem in fault tolerant systems? Lecture Notes in Computer Science, 774:171–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daley:1994:URA

[288] R. Daley and B. Kalyanasundaram. Use of reduction arguments in determining popperian FIN-Type learning capabilities. Lecture Notes in Computer Science, 774:173–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lipkis:1994:FTE

[289] Jim Lipkis and Marc Rozier. Fault tolerance enablers in the CHORUS microkernel. Lecture Notes in Computer Science, 774:182–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moriyama:1994:PLC

[290] T. Moriyama and M. Sato. Properties of language classes with finite elasticity. Lecture Notes in Computer Science, 774:187–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muller:1994:RCS

[291] Gilles Muller. A reliable client-server model on top of a micro-kernel. Lecture Notes in Computer Science, 774: 191–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapur:1994:UCV

[292] S. Kapur. Uniform characterizations of various kinds of language learning. Lecture Notes in Computer Science, 774:197–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Powell:1994:DFT

[293] David Powell. Distributed fault tolerance — lessons learnt from delta-4. Lecture Notes in Computer Science, 774:199–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knuutila:1994:HIC

[294] T. Knuutila. How to invent characterizable inference methods for regular languages. Lecture Notes in Computer Science, 774:209–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shrivastava:1994:AVC

[295] Santosh K. Shrivastava. Arjuna and voltan: Case studies in building fault tolerant distributed systems using standard components. Lecture Notes in Computer Science, 774:218–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cuellar:1994:NDA

[296] J. R. Cuellar and H. U. Simon. Neural discriminant analysis. Lecture Notes in Computer Science, 774:223–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bark:1994:FTP

[297] Rod Bark. Fault tolerant platforms for emerging telecommunications markets. Lecture Notes in Computer Science, 774:229–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iwayama:1994:NAA

[298] M. Iwayama, N. Indurkhya, and H. Motoda. A new algorithm for automatic configuration of hidden Markov models. Lecture Notes in Computer Science, 774:237–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jahanian:1994:FTE

[299] Farman Jahanian. Fault tolerance in embedded real-time systems. Lecture Notes in Computer Science, 774:237– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kopetz:1994:SDL

[300] Hermann Kopetz. The systematic design of large real-time systems or interface simplicity. Lecture Notes in Computer Science, 774:250–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakurai:1994:VDD

[301] A. Sakurai. On the VC-Dimension of depth four threshold circuits and the complexity of Boolean-valued functions. Lecture Notes in Computer Science, 774:251–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lala:1994:FTE

[302] Jaynarayan H. Lala and Richard E. Harper. Fault tolerance in embedded real-time systems: Importance and treatment of common mode failures. Lecture Notes in Computer Science, 774:263–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takimoto:1994:SCC

[303] E. Takimoto and A. Maruoka. On the sample complexity of consistent learning with one-sided error. Lecture Notes in Computer Science, 774:265– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shinohara:1994:CCV

[304] A. Shinohara. Complexity of computing Vapnik–Chervonenkis dimension. Lecture Notes in Computer Science, 774:279–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borr:1994:HAD

[305] Andrea Borr and Carol Wilhelmy. Highly-available data services for UNIX client-server networks: Why fault tolerant hardware isn’t the answer. Lecture Notes in Computer Science, 774:285–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hasegawa:1994:ALS

[306] S. Hasegawa, H. Imai, and M. Ishiguro. in-approximations of k-label spaces. Lecture Notes in Computer Science, 774:288–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakamura:1994:ELL

[307] A. Nakamura and N. Abe. Exact learning of linear combinations of monotone terms from function value queries. Lecture Notes in Computer Science, 774:300–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paris:1994:MRD

[308] Jean-François Paris. The management of replicated data. Lecture Notes in Computer Science, 774:305–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Siromoney:1994:TSD

[309] R. Siromoney, D. G. Thomas, K. G. Subramanian, and V. R. Dare. Thue systems and DNA — A learning algorithm for a subclass. Lecture Notes in Computer Science, 774:314–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishigami:1994:VDF

[310] Y. Ishigami and S. Tani. The VC-Dimensions of finite automata with n states. Lecture Notes in Computer Science, 774:328–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yoshida:1994:ULM

[311] K. Yoshida, H. Motoda, and N. Indurkhya. Unifying learning methods by colored digraphs. Lecture Notes in Computer Science, 774:342–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suwa:1994:PCV

[312] M. Suwa and H. Motoda. A perceptual criterion for visually controlling learning. Lecture Notes in Computer Science, 774:356–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kobayashi:1994:LSU

[313] S. Kobayashi. Learning strategies using decision lists. Lecture Notes in Computer Science, 774:370–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhong:1994:DBI

[314] N. Zhong and S. Ohsuga. A decomposition based induction model for discovering concept clusters from databases. Lecture Notes in Computer Science, 774:384–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ganascia:1994:ASS

[315] J.-G. Ganascia. Algebraic structure of some learning systems. Lecture Notes in Computer Science, 774:398– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsumoto:1994:IPR

[316] S. Tsumoto and H. Tanaka. Induction of probabilistic rules based on rough set theory. Lecture Notes in Computer Science, 774:410–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seidel:1994:NMP

[317] Raimund Seidel. The nature and meaning of perturbations in geometric computing. Lecture Notes in Computer Science, 775:3–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Devienne:1994:OBH

[318] Ph. Devienne, P. Lebegue, J-Ch. Routier, and J. Wurtz. One binary Horn clause is enough. Lecture Notes in Computer Science, 775:21–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bensaou:1994:TCL

[319] N. Bensaou and I. Guessarian. Transforming constraint logic programs. Lecture Notes in Computer Science, 775: 33–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laroussinie:1994:HTL

[320] F. Laroussinie and Ph. Schnoebelen. A hierarchy of temporal logics with past. Lecture Notes in Computer Science, 775:47–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goubault:1994:CRB

[321] J. Goubault. The complexity of resource-bounded first-order classical logic. Lecture Notes in Computer Science, 775:59–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benhamou:1994:TPP

[322] B. Benhamou, L. Sais, and P. Siegel. Two proof procedures for a cardinality based language in propositional calculus. Lecture Notes in Computer Science, 775:71–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vonBraunmuhl:1994:AHM

[323] B. von Braunmuhl, R. Gengler, and R. Rettinger. The alternation hierarchy for machines with sublogarithmic space is infinite. Lecture Notes in Computer Science, 775:85–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naik:1994:QTC

[324] A. V. Naik, K. W. Regan, and D. Sivakumar. Quasilinear time complexity theory. Lecture Notes in Computer Science, 775:97–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Macarie:1994:SED

[325] I. I. Macarie. Space-efficient deterministic simulation of probabilistic automata. Lecture Notes in Computer Science, 775:109–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Immerman:1994:RPL

[326] K. Immerman and N. Etessami. Reachability and the power of local ordering. Lecture Notes in Computer Science, 775:123–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mak:1994:PMA

[327] L. Mak. Are parallel machines always faster than sequential machines? Lecture Notes in Computer Science, 775: 137–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Comon:1994:GRA

[328] H. Comon and F. Jacquemard. Ground reducibility and automata with disequality constraints. Lecture Notes in Computer Science, 775:151–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khasidashvili:1994:PSN

[329] Z. Khasidashvili. Perpetuality and strong normalization in orthogonal term rewriting systems. Lecture Notes in Computer Science, 775:163– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sattler-Klein:1994:ACO

[330] A. Sattler-Klein. About changing the ordering during Knuth–Bendix completion. Lecture Notes in Computer Science, 775:175–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ringeissen:1994:CMA

[331] C. Ringeissen. Combination of matching algorithms. Lecture Notes in Computer Science, 775:187–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kik:1994:PCD

[332] M. Kik, M. Kutylowski, and G. Stachowiak. Periodic constant depth sorting networks. Lecture Notes in Computer Science, 775:201–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chlebus:1994:OPM

[333] B. S. Chlebus and L. Gasieniec. Optimal pattern matching on meshes. Lecture Notes in Computer Science, 775: 213–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kunde:1994:FSR

[334] M. Kunde, R. Niedermeier, and P. Rossmanith. Faster sorting and routing on grids with diagonals. Lecture Notes in Computer Science, 775:225– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sibeyn:1994:DKR

[335] J. F. Sibeyn and M. Kaufmann. Deterministic 1-k routing on meshes. Lecture Notes in Computer Science, 775: 237–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofmann:1994:UTT

[336] M. Hofmann and B. Pierce. A unifying type-theoretic framework for objects. Lecture Notes in Computer Science, 775:251–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Avenhaus:1994:OSB

[337] J. Avenhaus and K. Becker. Operational specifications with built-ins. Lecture Notes in Computer Science, 775: 263–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gribomont:1994:RVS

[338] E. P. Gribomont and J. Hagelstein. Reactive variables for system specification and design. Lecture Notes in Computer Science, 775:275–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Regan:1994:NPV

[339] K. W. Regan. A new parallel vector model, with exact characterizations of NC^{k}. Lecture Notes in Computer Science, 775:289–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alvarez:1994:ADP

[340] C. Alvarez and B. Jenner. On adaptive dlogtime and polylogtime reductions. Lecture Notes in Computer Science, 775:301–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ogiwara:1994:NKA

[341] M. Ogiwara. NC^{k}(NP) = AC^{k−1}(NP). Lecture Notes in Computer Science, 775:313–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arnold:1994:HS

[342] Andre Arnold. Hypertransition systems. Lecture Notes in Computer Science, 775:327–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Metivier:1994:SOF

[343] Y. Metivier and G. Richomme. On the star operation and the finite power property in free partially commutative monoids. Lecture Notes in Computer Science, 775:341–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruyere:1994:CT

[344] V. Bruyere, C. De Felice, and G. Guaiana. Coding with traces. Lecture Notes in Computer Science, 775:353– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giammarresi:1994:MSO

[345] D. Giammarresi, A. Restivo, S. Seibert, and W. Thomas. Monadic second-order logic over pictures and recognizability by tiling systems. Lecture Notes in Computer Science, 775:365–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delest:1994:GRI

[346] M. Delest and J-Ph. Dubernard. qgrammars: Results, implementation. Lecture Notes in Computer Science, 775:377–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karner:1994:TCS

[347] G. Karner. A topology for complete semirings. Lecture Notes in Computer Science, 775:389–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Book:1994:GPA

[348] R. V. Book, J. H. Lutz, and D. M. Martin. The global power of additional queries to random oracles. Lecture Notes in Computer Science, 775: 403–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lutz:1994:CVK

[349] J. H. Lutz and E. Mayordomo. Cook versus Karp-Levin: Separating completeness notions if NP is not small. Lecture Notes in Computer Science, 775:415–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thierauf:1994:SBT

[350] Th. Thierauf, S. Toda, and O. Watanabe. On sets bounded truth-table reducible to P-selective sets. Lecture Notes in Computer Science, 775:427– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Selivanov:1994:TRP

[351] V. L. Selivanov. Two refinements of the polynomial hierarchy. Lecture Notes in Computer Science, 775:439–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vollmer:1994:DRN

[352] H. Vollmer. On different reducibility notions for function classes. Lecture Notes in Computer Science, 775:449– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luby:1994:OPV

[353] M. Luby and W. Ertel. Optimal parallelization of Las Vegas algorithms. Lecture Notes in Computer Science, 775: 463–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Datta:1994:EPA

[354] A. Datta. Efficient parallel algorithms for geometric k-clustering problems. Lecture Notes in Computer Science, 775:475–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maheshwari:1994:SOP

[355] A. Maheshwari and A. Lingas. A simple optimal parallel algorithm for reporting paths in a tree. Lecture Notes in Computer Science, 775:487–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Apostolico:1994:PDA

[356] A. Apostolico, D. Breslauer, and Z. Galil. Parallel detection of all palindromes in a string. Lecture Notes in Computer Science, 775:497–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cai:1994:SPP

[357] L. Cai, J. Chen, R. Downey, and M. Fellows. On the structure of parameterized problems in NP. Lecture Notes in Computer Science, 775:509– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amaldi:1994:AFM

[358] E. Amaldi and V. Kann. On the approximability of finding maximum feasible subsystems of linear systems. Lecture Notes in Computer Science, 775: 521–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borchert:1994:APR

[359] B. Borchert. On the acceptance power of regular languages. Lecture Notes in Computer Science, 775:533–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hertrampf:1994:CCF

[360] U. Hertrampf. Complexity classes with finite acceptance types. Lecture Notes in Computer Science, 775:543– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parrow:1994:CAC

[361] J. Parrow and P. Sjodin. The complete axiomatization of cs-congruence. Lecture Notes in Computer Science, 775: 557–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanOostrom:1994:TSS

[362] V. van Oostrom and E. P. de Vink. Transition system specifications in stalk format with bisimulation as a congruence. Lecture Notes in Computer Science, 775:569–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jancar:1994:DQB

[363] P. Jancar. Decidability questions for bisimilarity of Petri nets and some related problems. Lecture Notes in Computer Science, 775:581–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buntrock:1994:VMP

[364] G. Buntrock and K. Lorys. The variable membership problem: Succinctness versus complexity. Lecture Notes in Computer Science, 775:595– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klemm:1994:EDS

[365] A. Klemm and R. Weber. Economy of description for single-valued transducers. Lecture Notes in Computer Science, 775:607–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shallit:1994:APM

[366] J. Shallit and Y. Breitbart. Automaticity: Properties of a measure of descriptional complexity. Lecture Notes in Computer Science, 775:619–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harel:1994:TTR

[367] David Harel. Towards a theory of recursive structures. Lecture Notes in Computer Science, 775:633–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arimura:1994:FMG

[368] H. Arimura, T. Shinohara, and S. Otsuki. Finding minimal generalizations for unions of pattern languages and its application to inductive inference from positive data. Lecture Notes in Computer Science, 775:649–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mateescu:1994:NP

[369] A. Mateescu and A. Salomaa. Nondeterminism in patterns. Lecture Notes in Computer Science, 775:661–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dancik:1994:UBE

[370] V. Paterson and M. Dancik. Upper bounds for the expected length of a longest common subsequence of two binary sequences. Lecture Notes in Computer Science, 775:669–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petersen:1994:APW

[371] H. Petersen. The ambiguity of primitive words. Lecture Notes in Computer Science, 775:679–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1994:CHN

[372] N. H. Lam. On codes having no finite completion. Lecture Notes in Computer Science, 775:691–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lipton:1994:NAI

[373] R. J. Lipton. A new approach to information theory. Lecture Notes in Computer Science, 775:699–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Le:1994:VDM

[374] N.-M. Le. On Voronoi diagrams in the L_{p}-metric in higher dimensions. Lecture Notes in Computer Science, 775: 711–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kao:1994:TPA

[375] M.-Y. Kao. Total protection of analytic invariant information in cross tabulated tables. Lecture Notes in Computer Science, 775:723–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dragan:1994:DCG

[376] F. F. Dragan and A. Brandstadt. Dominating cliques in graphs with hypertree structure. Lecture Notes in Computer Science, 775:735–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deogun:1994:VRP

[377] J. S. Deogun, T. Kloks, D. Kratsch, and H. Muller. On vertex ranking for permutation and other graphs. Lecture Notes in Computer Science, 775: 747–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kloks:1994:FAM

[378] T. Kloks and D. Kratsch. Finding all minimal separators of a graph. Lecture Notes in Computer Science, 775: 759–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodlaender:1994:CMC

[379] H. L. Bodlaender and K. Jansen. On the complexity of the maximum cut problem. Lecture Notes in Computer Science, 775:769–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aizawa:1994:PCG

[380] K. Aizawa and A. Nakamura. Pathcontrolled graph grammars for multiresolution image processing and analysis. Lecture Notes in Computer Science, 776:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andries:1994:SSH

[381] M. Andries and G. Engels. Syntax and semantics of hybrid database languages. Lecture Notes in Computer Science, 776:19–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arnborg:1994:DHD

[382] S. Arnborg. Decomposability helps for deciding logics of knowledge and belief. Lecture Notes in Computer Science, 776:37–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barendsen:1994:EGR

[383] E. Barendsen and S. Smetsers. Extending graph rewriting with copying. Lecture Notes in Computer Science, 776: 51–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barthelmann:1994:GGS

[384] K. Barthelmann and G. Schied. Graphgrammar semantics of a higher-order programming language for distributed systems. Lecture Notes in Computer Science, 776:71–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corradini:1994:AGD

[385] A. Corradini, H. Ehrig, M. Lowe, U. Montanari, and F. Rossi. Abstract graph derivations in the double pushout approach. Lecture Notes in Computer Science, 776:86–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corradini:1994:NSR

[386] A. Corradini, H. Ehrig, M. Lowe, U. Montanari, and F. Rossi. Note on standard representation of graphs and graph derivations. Lecture Notes in Computer Science, 776:104–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corradini:1994:JRA

[387] A. Corradini and D. Wolz. Jungle rewriting: an abstract description of a lazy narrowing machine. Lecture Notes in Computer Science, 776:119– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courcelle:1994:RSG

[388] B. Courcelle and J. Lagergren. Recognizable sets of graphs of bounded tree-width. Lecture Notes in Computer Science, 776:138–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrig:1994:CDH

[389] H. Ehrig, H.-J. Kreowski, and G. Taentzer. Canonical derivations for high-level replacement systems. Lecture Notes in Computer Science, 776: 153–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gemis:1994:CMG

[390] M. Gemis, J. Paredaens, P. Peelman, and J. Van den Bussche. A computational model for generic graph functions. Lecture Notes in Computer Science, 776:170–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grabska:1994:GD

[391] E. Grabska. Graphs and designing. Lecture Notes in Computer Science, 776:188–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janssens:1994:ESC

[392] D. Janssens. ESM systems and the composition of their computations. Lecture Notes in Computer Science, 776:203–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kawahara:1994:RST

[393] Y. Kawahara and Y. Mizoguchi. Relational structures and their partial morphisms in view of single pushout rewriting. Lecture Notes in Computer Science, 776:218–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Korff:1994:SPT

[394] M. Korff. Single pushout transformations of equationally defined graph structures with applications to actor systems. Lecture Notes in Computer Science, 776:234–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lowe:1994:PSP

[395] M. Lowe and J. Dingel. Parallelism in single-pushout graph rewriting. Lecture Notes in Computer Science, 776: 248–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maggiolo-Schettini:1994:SFS

[396] A. Maggiolo-Schettini and A. Peron. Semantics of full statecharts based on graph rewriting. Lecture Notes in Computer Science, 776:265–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montanari:1994:CON

[397] U. Montanari and F. Rossi. Contextual occurrence nets and concurrent constraint programming. Lecture Notes in Computer Science, 776:280–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nagl:1994:UMG

[398] M. Nagl. Uniform-modelling in graph grammar specifications. Lecture Notes in Computer Science, 776:296–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raoult:1994:STG

[399] J.-C. Raoult and F. Voisin. Settheoretic graph rewriting. Lecture Notes in Computer Science, 776:312– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schied:1994:RRS

[400] G. Schied. On relating rewriting systems and graph grammars to event structures. Lecture Notes in Computer Science, 776:326–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schurr:1994:LBS

[401] A. Schurr. Logic based structure rewriting systems. Lecture Notes in Computer Science, 776:341–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smetsers:1994:GSD

[402] S. Smetsers, E. Barendsen, M. v. Eekelen, and R. Plasmeijer. Guaranteeing safe destructive updates through a type system with uniqueness information for graphs. Lecture Notes in Computer Science, 776:358–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taentzer:1994:AGT

[403] G. Taentzer and M. Beyer. Amalgamated graph transformations and their use for specifying AGG — an algebraic graph grammar system. Lecture Notes in Computer Science, 776:380– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walter:1994:CSD

[404] B. Walter and S. Benzschawel. Complex structures in (deductive) database systems: Variables and open property sets. Lecture Notes in Computer Science, 777:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lipeck:1994:OOS

[405] U. W. Lipeck and S. Brass. Objectoriented system specification using defaults. Lecture Notes in Computer Science, 777:22–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burckert:1994:TR

[406] H.-J Burckert. Terminologies and rules. Lecture Notes in Computer Science, 777:44–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scholl:1994:ETB

[407] M. H. Scholl and M. Tresch. Evolution towards, in, and beyond object databases. Lecture Notes in Computer Science, 777:64–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eiter:1994:EPC

[408] T. Eiter, G. Gottlob, and H. Mannila. Expressive power and complexity of disjunctive datalog under the stable model semantics. Lecture Notes in Computer Science, 777:83–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nebel:1994:MQT

[409] B. Nebel and H.-J. Burckert. Managing qualitative temporal information: Expressiveness vs. complexity. Lecture Notes in Computer Science, 777:104– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kiebling:1994:DRD

[410] W. Kiebling and U. Guntzer. Database reasoning — A deductive framework for solving large and complex problems by means of subsumption. Lecture Notes in Computer Science, 777: 118–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Furbach:1994:TRF

[411] U. Furbach. Theory reasoning in first order calculi. Lecture Notes in Computer Science, 777:139–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saake:1994:MIS

[412] G. Saake and T. Hartmann. Modelling information systems as object societies. Lecture Notes in Computer Science, 777:157–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wielinga:1994:CML

[413] B. J. Wielinga and A. T. Schreiber. Conceptual modelling of large reusable knowledge bases. Lecture Notes in Computer Science, 777:181–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jarke:1994:TIE

[414] M. Jarke, H. W. Nissen, and K. Pohl. Tool integration in evolving information systems environments. Lecture Notes in Computer Science, 777:201– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartmanis:1994:ITA

[415] J. Hartmanis and S. Chari. On the intellectual terrain around NP. Lecture Notes in Computer Science, 778:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garg:1994:AGD

[416] A. Garg and R. Tamassia. Advances in graph drawing. Lecture Notes in Computer Science, 778:12–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sahinalp:1994:PAM

[417] S. C. Sahinalp and U. Vishkin. On a parallel-algorithms method for string matching problems. Lecture Notes in Computer Science, 778:22–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yannakakis:1994:SOP

[418] M. Yannakakis. Some open problems in approximation. Lecture Notes in Computer Science, 778:33–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alimonti:1994:NLS

[419] P. Alimonti. New local search approximation techniques for maximum generalized satisfiability problems. Lecture Notes in Computer Science, 778:40–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergadano:1994:LBA

[420] F. Bergadano and S. Varricchio. Learning behaviors of automata from multiplicity and equivalence queries. Lecture Notes in Computer Science, 778:54–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernasconi:1994:MBF

[421] A. Bernasconi and B. Codenotti. Measures of Boolean function complexity based on harmonic analysis. Lecture Notes in Computer Science, 778:63–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clementi:1994:GTI

[422] A. Clementi and R. Impagliazzo. Graph theory and interactive protocols for reachability problems on finite cellular automata. Lecture Notes in Computer Science, 778:73–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dekel:1994:PPD

[423] E. Dekel and J. Hu. Parallel pruning decomposition (PDS) and biconnected components of graphs. Lecture Notes in Computer Science, 778:91–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiCrescenzo:1994:NIE

[424] G. Di Crescenzo. A non-interactive electronic cash system. Lecture Notes in Computer Science, 778:109–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Even:1994:USR

[425] S. Even and A. Litman. A unified scheme for routing in expander based networks. Lecture Notes in Computer Science, 778:125–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frigioni:1994:DBF

[426] D. Frigioni, A. Marchetti-Spaccamela, and U. Nanni. Dynamization of backtrack-free search for the constraint satisfaction problem. Lecture Notes in Computer Science, 778:136–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofri:1994:ERB

[427] M. Hofri and H. Shachnai. Efficient reorganization of binary search trees. Lecture Notes in Computer Science, 778:152–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Israeli:1994:TMT

[428] A. Israeli, E. Kranakis, D. Krizanc, and N. Santoro. Time-message trade-offs for the weak unison problem. Lecture Notes in Computer Science, 778:167– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1994:SET

[429] T. W. Lam and K. H. Lee. On set equality-testing. Lecture Notes in Computer Science, 778:179–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monti:1994:CSR

[430] A. Monti and A. Roncato. On the complexity of some reachability problems. Lecture Notes in Computer Science, 778:192–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mundhenk:1994:SRS

[431] M. Mundhenk. On self-reducible sets of low information content. Lecture Notes in Computer Science, 778:203– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rub:1994:LBM

[432] C. Rub. Lower bounds for merging on the hypercube. Lecture Notes in Computer Science, 778:213–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stone:1994:OMG

[433] C. M. Stone. The object management group: Standardization of object technology. Lecture Notes in Computer Science, 779:1, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herbert:1994:DDS

[434] A. Herbert. Databases in distributed systems: The new frontier. Lecture Notes in Computer Science, 779:2–6, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agrawal:1994:TDU

[435] R. Agrawal and L. G. Demichiel. Type derivation using the projection operation. Lecture Notes in Computer Science, 779:7–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buchheit:1994:SBQ

[436] M. Buchheit, M. A. Jeusfeld, W. Nutt, and M. Staudt. Subsumption between queries to object-oriented databases. Lecture Notes in Computer Science, 779:15–22, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pirahesh:1994:COV

[437] H. Pirahesh, B. Mitschang, N. Südkamp, and B. Lindsay. Composite-object views in relational DBMS: An implementation perspective. Lecture Notes in Computer Science, 779:23–30, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cicchetti:1994:MRS

[438] R. Cicchetti and L. Lakhal. Matrix relation for statistical database management. Lecture Notes in Computer Science, 779:31–44, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Consens:1994:DDS

[439] M. P. Consens, A. O. Mendelzon, and D. Vista. Deductive database support for data visualization. Lecture Notes in Computer Science, 779:45–58, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Somogyi:1994:SFB

[440] Z. Somogyi, D. Kemp, J. Harland, and K. Ramamohanarao. Subsumptionfree bottom-up evaluation of logic programs with partially instantiated data structures. Lecture Notes in Computer Science, 779:59–72, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miller:1994:SEH

[441] R. J. Miller, Y. E. Ioannidis, and R. Ramakrishnan. Schema equivalence in heterogeneous systems: Bridging theory and practice. Lecture Notes in Computer Science, 779:73–80, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DosSantos:1994:VSB

[442] C. S. Dos Santos, S. Abiteboul, and C. Delobel. Virtual schemas and bases. Lecture Notes in Computer Science, 779:81–94, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scheuermann:1994:RBQ

[443] P. Scheuermann and E. I. Chong. Role-based query processing in multidatabase systems. Lecture Notes in Computer Science, 779:95–108, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sheldon:1994:CRD

[444] M. A. Sheldon, A. Duda, R. Weiss, J. W. O’Toole, and D. K. Gifford. Content routing for distributed information servers. Lecture Notes in Computer Science, 779:109–122, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alonso:1994:UAC

[445] G. Alonso, R. Vingralek, D. Agrawal, Y. Breitbart, A. El Abbadi, H. Schek, and G. Weikum. A unified approach to concurrency control and transaction recovery. Lecture Notes in Computer Science, 779:123–130, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohan:1994:AFS

[446] C. Mohan and D. Haderle. Algorithms for flexible space management in transaction systems supporting finegranularity locking. Lecture Notes in Computer Science, 779:131–144, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bober:1994:IAM

[447] P. M. Bober and M. J. Carey. Indexing alternatives for multiversion locking. Lecture Notes in Computer Science, 779:145–158, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petrounias:1994:RBA

[448] I. Petrounias and P. Loucopoulos. A rule-based approach for the design and implementation of information systems. Lecture Notes in Computer Science, 779:159–172, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morsi:1994:BSE

[449] M. M. A. Morsi, S. B. Navathe, and J. Shilling. On behavioral schema evolution in object-oriented databases. Lecture Notes in Computer Science, 779:173–186, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nixon:1994:RUP

[450] B. A. Nixon. Representing and using performance requirements during the development of information systems. Lecture Notes in Computer Science, 779:187–200, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Qian:1994:MTS

[451] X. Qian. A model-theoretic semantics of the multilevel relational model. Lecture Notes in Computer Science, 779:201–214, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clifford:1994:SBT

[452] J. Clifford and T. Isakowitz. On the semantics of (bi)temporal variable databases. Lecture Notes in Computer Science, 779:215–230, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Formica:1994:CIH

[453] A. Formica and M. Missikoff. Correctness of ISA hierarchies in objectoriented database schemas. Lecture Notes in Computer Science, 779:231– 244, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Imielinski:1994:PEF

[454] T. Imielinski, S. Viswanathan, and B. R. Badrinath. Power efficient filtering of data on air. Lecture Notes in Computer Science, 779:245–258, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hjelsvold:1994:VIC

[455] R. Hjelsvold. Video information contents and architecture. Lecture Notes in Computer Science, 779:259–272, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1994:OSO

[456] L. T. Chen and D. Rotem. Optimizing storage of objects on mass storage systems with robotic devices. Lecture Notes in Computer Science, 779:273– 286, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Swami:1994:EJR

[457] A. Swami and K. B. Schiefer. On the estimation of join result sizes. Lecture Notes in Computer Science, 779:287– 300, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhao:1994:DDB

[458] X. Zhao, R. G. Johnson, and N. J. Martin. DBJ — A dynamic balancing hash join algorithm in multiprocessor database systems. Lecture Notes in Computer Science, 779:301–308, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morzy:1994:TSO

[459] T. Morzy, M. Matysiak, and S. Salza. Tabu search optimization of large join queries. Lecture Notes in Computer Science, 779:309–322, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1994:IPE

[460] C. M. Chen and N. Roussopoulos. The implementation and performance evaluation of the ADMS query optimizer: Integrating query result caching and matching. Lecture Notes in Computer Science, 779:323–336, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steenhagen:1994:ONQ

[461] H. J. Steenhagen, P. M. G. Apers, and H. M. Blanken. Optimization of nested queries in a complex object model. Lecture Notes in Computer Science, 779:337–350, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerlhof:1994:MTA

[462] C. A. Gerlhof and A. Kemper. A multithreaded architecture for prefetching in object bases. Lecture Notes in Computer Science, 779:351–364, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1994:SFT

[463] E. W. Brown, J. P. Callan, W. B. Croft, and J. E. B. Moss. Supporting fulltext information retrieval with a persistent object store. Lecture Notes in Computer Science, 779:365–378, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Panagopoulos:1994:BSS

[464] G. Panagopoulos and C. Faloutsos. Bitsliced signature files for very large text databases on a parallel machine architecture. Lecture Notes in Computer Science, 779:379–392, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goble:1994:STS

[465] C. Goble and P. Crowther. Schemas for telling stories in medical records. Lecture Notes in Computer Science, 779:393–406, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersen:1994:PVU

[466] F. Andersen, K. D. Petersen, and J. S. Pettersson. Program verification using HOL-UNITY. Lecture Notes in Computer Science, 780:1–15, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petersen:1994:GML

[467] K. D. Petersen. Graph model of LAMBDA in higher order logic. Lecture Notes in Computer Science, 780:16–28, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1994:MPL

[468] C. Zhang, R. Shaw, R. Olsson, and K. Levitt. Mechanizing a programming logic for the concurrent programming language microSR in HOL. Lecture Notes in Computer Science, 780:29–42, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Syme:1994:RFD

[469] D. Syme. Reasoning with the formal definition of Standard ML in HOL. Lecture Notes in Computer Science, 780:43–58, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanInwegen:1994:HM

[470] M. VanInwegen and E. Gunter. HOL-ML. Lecture Notes in Computer Science, 780:59–72, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goossens:1994:SBH

[471] K. Goossens. Structure and behaviour in hardware verification. Lecture Notes in Computer Science, 780:73–86, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Angelo:1994:DFS

[472] C. M. Angelo, L. Claesen, and H. De Man. Degrees of formality in shallow embedding hardware description languages in HOL. Lecture Notes in Computer Science, 780:87–98, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eisenbiegler:1994:FAF

[473] D. Eisenbiegler, K. Schneider, and R. Kumar. A functional approach for formalizing regular hardware structures. Lecture Notes in Computer Science, 780:99–112, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thery:1994:PDS

[474] L. Thery. A proof development system for HOL. Lecture Notes in Computer Science, 780:113–126, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roxas:1994:HPR

[475] R. Roxas. A HOL package for reasoning about relations defined by mutual induction. Lecture Notes in Computer Science, 780:127–139, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gunter:1994:BCT

[476] E. Gunter. A broader class of trees for recursive type definitions for HOL. Lecture Notes in Computer Science, 780:140–153, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parnas:1994:STW

[477] D. Parnas. Some theorems we should prove. Lecture Notes in Computer Science, 780:154–161, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rushby:1994:UPP

[478] J. Rushby and M. Srivas. Using PVS to prove some theorems of David Parnas. Lecture Notes in Computer Science, 780:162–173, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harrison:1994:EHT

[479] J. Harrison and L. Thery. Extending the HOL theorem prover with a computer algebra system to reason about the reals. Lecture Notes in Computer Science, 780:174–184, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joyce:1994:HVS

[480] J. Joyce and C. Seger. The HOL-Voss system: Model-checking inside a general-purpose theorem-prover. Lecture Notes in Computer Science, 780: 185–198, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1994:LHV

[481] J. Lu and S. Chin. Linking HOL to a VLSI CAD system. Lecture Notes in Computer Science, 780:199–212, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneider:1994:APP

[482] K. Schneider, R. Kumar, and T. Kropf. Alternative proof procedures for finitestate machines in higher-order logic. Lecture Notes in Computer Science, 780:213–226, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McIsaac:1994:FAL

[483] A. McIsaac. A formalization of abstraction in LAMBDA. Lecture Notes in Computer Science, 780:227–238, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arora:1994:RUM

[484] T. Arora, T. Leung, K. Levitt, and T. Schubert. Report on the UCD microcoded viper verification project. Lecture Notes in Computer Science, 780:239–252, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhu:1994:VTM

[485] Z. Zhu, J. Joyce, and C. Seger. Verification of the Tamarack-3 microprocessor in a hybrid verification environment. Lecture Notes in Computer Science, 780:253–266, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fura:1994:ATM

[486] D. A. Fura, P. J. Windley, and A. K. Somani. Abstraction techniques for modeling real-world interface chips. Lecture Notes in Computer Science, 780:267–280, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tahar:1994:IMF

[487] S. Tahar and R. Kumar. Implementing a methodology for formally verifying RISC processors in HOL. Lecture Notes in Computer Science, 780:281– 294, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agerholm:1994:DTH

[488] S. Agerholm. Domain theory in HOL. Lecture Notes in Computer Science, 780:295–309, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chou:1994:PTL

[489] C.-T. Chou. Predicates, temporal logic, and simulations. Lecture Notes in Computer Science, 780:310–323, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prasetya:1994:FVA

[490] W. Prasetya. Formalization of variables access constraints to support compositionality of liveness properties. Lecture Notes in Computer Science, 780:324– 337, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Day:1994:SSH

[491] N. Day and J. Joyce. The semantics of statecharts in HOL. Lecture Notes in Computer Science, 780:338–351, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nesi:1994:VPC

[492] M. Nesi. Value-passing CCS in HOL. Lecture Notes in Computer Science, 780:352–365, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andrews:1994:TIA

[493] P. Andrews, M. Bishop, S. Issar, and D. Nesmith. TPS: An interactive and automatic tool for proving theorems of type theory. Lecture Notes in Computer Science, 780:366–371, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wong:1994:MBV

[494] W. Wong. Modelling bit vectors in HOL: the word library. Lecture Notes in Computer Science, 780:372–385, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneider:1994:EHO

[495] K. Schneider, R. Kumar, and T. Kropf. Eliminating higher-order quantifiers to obtain decision procedures for hardware verification. Lecture Notes in Computer Science, 780:386–399, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aagaard:1994:TSD

[496] M. Aagaard, M. Leeser, and P. Windley. Toward a super duper hardware tactic. Lecture Notes in Computer Science, 780:400–413, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gordon:1994:MNC

[497] A. Gordon. A mechanisation of name-carrying syntax up to alphaconversion. Lecture Notes in Computer Science, 780:414–426, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harrison:1994:HDP

[498] J. Harrison. A HOL decision procedure for elementary real algebra. Lecture Notes in Computer Science, 780:427– 437, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Slind:1994:AUH

[499] K. Slind. AC unification in HOL90. Lecture Notes in Computer Science, 780:438–450, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brackin:1994:SPR

[500] S. Brackin and S.-K. Chin. Serverprocess restrictiveness in HOL. Lecture Notes in Computer Science, 780:451– 464, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morley:1994:SRS

[501] M. J. Morley. Safety in railway signalling data: A behavioural analysis. Lecture Notes in Computer Science, 780:465–475, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prasetya:1994:SMP

[502] W. Prasetya. On the style of mechanical proving. Lecture Notes in Computer Science, 780:476–489, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rajan:1994:ADT

[503] S. Rajan, J. Joyce, and C. Seger. From abstract data types to shift registers: A case study in formal specification and verification at differing levels of abstraction using theorem proving and symbolic simulation. Lecture Notes in Computer Science, 780:490–501, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carreno:1994:VHO

[504] V. Carreno. Verification in higher order logic of mutual exclusion algorithm. Lecture Notes in Computer Science, 780:502–515, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalvala:1994:UIP

[505] S. Kalvala. Using Isabelle to prove simple theorems. Lecture Notes in Computer Science, 780:516–517, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lampson:1994:ICA

[506] Butler W. Lampson. Interconnecting computers: Architecture, technology, and economics. Lecture Notes in Computer Science, 782:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graham:1994:LIS

[507] Susan Graham. Languages and interactive software development. Lecture Notes in Computer Science, 782:21–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vandeSnepscheut:1994:MSS

[508] Jan L. A. van de Snepscheut. Mechanized support for stepwise refinement. Lecture Notes in Computer Science, 782:35–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoare:1994:HSC

[509] C. A. R. Hoare and Innes Page. Hardware and software: The closing gap. Lecture Notes in Computer Science, 782:49–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vuillemin:1994:CP

[510] Jean E. Vuillemin. On computing power. Lecture Notes in Computer Science, 782:69–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moyer:1994:IMB

[511] Sally A. Moyer, Steven A. McKee, William A. Wulf, and Charles Hitchcock. Increasing memory bandwidth for vector computations. Lecture Notes in Computer Science, 782:87–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benitez:1994:AMD

[512] Manuel E. Benitez and Jack W. Davidson. The advantages of machinedependent global optimization. Lecture Notes in Computer Science, 782:105– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambrosch:1994:DCG

[513] Wolfgang Ambrosch, M. Anton Ertl, Felix Beer, and Andreas Krall. Dependence-conscious global register allocation. Lecture Notes in Computer Science, 782:129–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corney:1994:TTE

[514] Diane Corney and John Gough. Type test elimination using typeflow analysis. Lecture Notes in Computer Science, 782:137–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chrobot:1994:WCP

[515] Stanislaw Chrobot. Where concurrent processes originate. Lecture Notes in Computer Science, 782:151–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jagannathan:1994:HLA

[516] Suresh Jagannathan and James Philbin. High-level abstractions for efficient concurrent systems. Lecture Notes in Computer Science, 782:171– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spinellis:1994:LAP

[517] Diomidis Spinellis, Sophia Drossopoulou, and Susan Eisenbach. Language and architecture paradigms as object classes: A unified approach towards multiparadigm programming. Lecture Notes in Computer Science, 782:191– 207, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Szyperski:1994:EPL

[518] Clemens A. Szyperski, Stephen Omohundro, and Stephan Murer. Engineering a programming language: The type and class system of sather. Lecture Notes in Computer Science, 782:208– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Didrich:1994:ODI

[519] Klaus Didrich, Andreas Fett, Carola Gerke, Wolfgang Grieskamp, and Peter Pepper. OPAL: Design and implementation of an algebraic programming language. Lecture Notes in Computer Science, 782:228–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohn:1994:AIS

[520] Alan G. Cohn and David L. Yoder. Architectural issues in spreadsheet languages. Lecture Notes in Computer Science, 782:245–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franz:1994:TST

[521] Michael Franz. Technological steps toward a software component industry. Lecture Notes in Computer Science, 782:259–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Collberg:1994:DHL

[522] Christian S. Collberg. Distributed high-level module binding for flexible encapsulation and fast inter-modular optimization. Lecture Notes in Computer Science, 782:282–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radenski:1994:OSP

[523] Atanas Radenski. Is Oberon as simple as possible? A smaller objectoriented language based on the concept of module type. Lecture Notes in Computer Science, 782:298–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naumann:1994:EO

[524] David A. Naumann. On the essence of oberon. Lecture Notes in Computer Science, 782:313–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lalis:1994:ACO

[525] Spiros Lalis and Beverly A. Sanders. Adding concurrency to the Oberon system. Lecture Notes in Computer Science, 782:328–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goodman:1994:PC

[526] D. J. Goodman. Personal communications. Lecture Notes in Computer Science, 783:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mouly:1994:EG

[527] M. Mouly and M.-B. Pautet. The evolution of GSM. Lecture Notes in Computer Science, 783:13–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merakos:1994:IDT

[528] L. F. Merakos and F. Li. Integrating data traffic in Enhanced-TDMA digital cellular systems. Lecture Notes in Computer Science, 783:21–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baiocchi:1994:GDC

[529] A. Baiocchi, F. Delli Priscoli, F. Grilli, and F. Sestini. The geometric dynamic channel allocation strategy for high traffic FDM/TDMA mobile communications networks. Lecture Notes in Computer Science, 783:33–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hussmann:1994:PED

[530] H. Hussmann. Performance evaluation of the DECT radio resource management. Lecture Notes in Computer Science, 783:45–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Padovani:1994:CCC

[531] R. Padovani. The capacity of CDMA cellular: Reverse link field test results. Lecture Notes in Computer Science, 783:56–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Siveski:1994:AMD

[532] Z. Siveski, Y. Bar-Ness, and D. W. Chen. Adaptive multiuser detector for synchronous code division multiple access applications. Lecture Notes in Computer Science, 783:67–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cideciyan:1994:CCS

[533] R. D. Cideciyan and E. Eleftheriou. Concatenated coding scheme for reliable data transmission in CDMA cellular systems. Lecture Notes in Computer Science, 783:76–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naguib:1994:PCC

[534] A. F. Naguib and A. Paulraj. Performance of CDMA cellular networks with base-station antenna arrays. Lecture Notes in Computer Science, 783:87–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghareeb:1994:JFP

[535] I. Ghareeb and A. Yongacoglu. Joint frequency-phase modulation over Rayleigh fading channels. Lecture Notes in Computer Science, 783:101– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lapidoth:1994:CCF

[536] A. Lapidoth. Convolutional codes and finite interleavers for the block erasure channel. Lecture Notes in Computer Science, 783:113–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayr:1994:FRH

[537] B. Mayr and H. Weinrichter. Fractionalrate hybrid PSK-TCM codes for the Rayleigh fading channel. Lecture Notes in Computer Science, 783:121– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cacopardi:1994:ICD

[538] S. Cacopardi, F. Frescura, and G. Reali. Increased capacity of DS-CDMA mobile systems on frequency-selective slow fading channels. Lecture Notes in Computer Science, 783:131–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drakul:1994:DPD

[539] S. L. Drakul and E. Biglieri. Design and performance of an 8-dimensional modulation scheme. Lecture Notes in Computer Science, 783:141–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patel:1994:ASI

[540] P. R. Patel and J. M. Holtzman. Analysis of successive interference cancellation in M-ary orthogonal DS-CDMA system with single path Rayleigh fading. Lecture Notes in Computer Science, 783:150–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levi:1994:SRC

[541] L. Levi, F. Muratore, and G. Romano. Simulation results for a CDMA interference cancellation technique in a Rayleigh fading channel. Lecture Notes in Computer Science, 783:162– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wichman:1994:MMD

[542] R. Wichman, J. Lilleberg, and Y. Neuvo. Matched median detectors for synchronous direct-sequence multiple-access communications. Lecture Notes in Computer Science, 783: 172–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blanz:1994:CSE

[543] J. Blanz, A. Klein, M. Nabhan, and A. Steil. Cellular spectrum efficiency of a joint detection C/TDMA mobile radio system. Lecture Notes in Computer Science, 783:184–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suzuki:1994:MSD

[544] T. Suzuki and T. Mizuno. Multiplesymbol differential detection scheme for differential amplitude modulation. Lecture Notes in Computer Science, 783: 196–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wautier:1994:PCF

[545] A. Wautier, J.-C. Dany, and C. Mourot. Phase-correcting filter for sub-optimal equalizers. Lecture Notes in Computer Science, 783:208–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valdovinos:1994:PAE

[546] A. Valdovinos and F. J. Casadevall. Performance of adaptive equalization in typical indoor environments including non-linear power amplifiers. Lecture Notes in Computer Science, 783: 219–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nogami:1994:NEP

[547] H. Nogami and G. L. Stuber. New error probability upper bound on maximum likelihood sequence estimation for intersymbol interference channels. Lecture Notes in Computer Science, 783: 230–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schramm:1994:BEP

[548] P. Schramm. Bit error probability of a rapidly-adapting rake receiver as a function of dwell time. Lecture Notes in Computer Science, 783:242–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strom:1994:NFR

[549] E. G. Strom, S. Parkvall, and B. E. Ottersten. Near-far resistant propagation delay estimators for asynchronous direct-sequence code division multiple access systems. Lecture Notes in Computer Science, 783:251–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gfeller:1994:DCP

[550] F. R. Gfeller, P. Bernasconi, W. Hirt, C. Elisii, and B. Weiss. Dynamic cell planning for wireless infrared in-house data transmission. Lecture Notes in Computer Science, 783:261–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hashemi:1994:FRM

[551] H. Hashemi, G. Yun, M. Kavehrad, F. Behbahani, and P. Galko. Frequency response measurements of the wireless indoor channel at infrared optics. Lecture Notes in Computer Science, 783:273–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lomba:1994:PLI

[552] C. R. A. T. Lomba, R. T. Valadas, and A. M. de Oliveira Duarte. Propagation losses and impulse response of the indoor optical channel: A simulation package. Lecture Notes in Computer Science, 783:285–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCullagh:1994:MOW

[553] M. J. McCullagh, D. R. Wisely, P. L. Eardley, and P. P. Smyth. A 50 mbit/s optical wireless LAN link using novel optical and electronic enabling technologies. Lecture Notes in Computer Science, 783:298–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sorbara:1994:PCG

[554] D. Sorbara and M. Visintin. Performance comparison of GMSK and pi/4-DQPSK modulations in a mobile radio environment. Lecture Notes in Computer Science, 783:310–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bennett:1994:EMP

[555] D. W. Bennett, P. B. Kenington, and R. J. Wilkinson. The effect of modulation on the power rating of multichannel linear amplifiers. Lecture Notes in Computer Science, 783:322– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blanke:1994:SGD

[556] G. Blanke. Simulation of GSM data channels in a safety critical railway environment. Lecture Notes in Computer Science, 783:334–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janssen:1994:MBE

[557] G. J. M. Janssen, P. A. Stigter, and R. Prasad. A model for BER evaluation of indoor frequency selective channels using multipath measurement results at 2.4, 4.75 and 11.5 GHz. Lecture Notes in Computer Science, 783: 344–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kattenbach:1994:MBE

[558] R. Kattenbach and H. Fruchting. Measuring bit errors, clock errors and their burst characteristics in mobile communication systems. Lecture Notes in Computer Science, 783:356–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ariyavisitakul:1994:PDF

[559] S. Ariyavisitakul. Performance of dynamic feedback power control in a TDMA radio system. Lecture Notes in Computer Science, 783:365–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saunders:1994:PCR

[560] R. Saunders and L. Lopes. Performance of cellular radio systems with power control in varying shadow fading environments. Lecture Notes in Computer Science, 783:376–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zorzi:1994:PCF

[561] M. Zorzi and L. B. Milstein. Power control on the forward link in cellular CDMA. Lecture Notes in Computer Science, 783:391–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeGaudenzi:1994:PAC

[562] R. De Gaudenzi and F. Giannetti. Performance analysis of a CDMA-based satellite-mobile audio broadcasting system. Lecture Notes in Computer Science, 783:400–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berruto:1994:RPA

[563] E. Berruto, T. Brannlund, J. Gustafsson, D. Maiwald, W. Schott, and P. Whitworth. Radio protocol architecture of the CODIT UMTS system. Lecture Notes in Computer Science, 783:417–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cygan:1994:RIA

[564] D. Cygan, F. David, H. J. Eul, J. Hofmann, N. Metzner, and W. Mohr. RACE-II advanced TDMA mobile access project — an approach for UMTS. Lecture Notes in Computer Science, 783:428–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vandendorpe:1994:MSS

[565] L. Vandendorpe. Multitone spread spectrum communication systems in a multipath rician fading channel. Lecture Notes in Computer Science, 783: 440–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zorzi:1994:SAC

[566] M. Zorzi and R. R. Rao. Slotted ALOHA with capture in a mobile radio environment. Lecture Notes in Computer Science, 783:452–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vvedenskaya:1994:IPM

[567] N. D. Vvedenskaya, J. C. Arnbak, and B. S. Tsybakov. Improved performance of mobile data networks using stack algorithms and receiver capture. Lecture Notes in Computer Science, 783:464–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Babich:1994:CRM

[568] F. Babich. Contention resolution multiple access techniques for handling speech and data traffic on wireless networks. Lecture Notes in Computer Science, 783:476–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ketseoglou:1994:CGS

[569] T. Ketseoglou. Certain generalizations on the slotted collision channel without feedback. Lecture Notes in Computer Science, 783:488–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baumberger:1994:MIS

[570] W. Baumberger and H. Kaufmann. Monolithic integration of a spreadspectrum transmitter and receiver front end for wireless LAN applications. Lecture Notes in Computer Science, 783:501–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nilsson:1994:BSC

[571] P. Nilsson, M. Torkelson, K. Palmkvist, M. Vesterbacka, and L. Wanhammar. A bit-serial CMOS digital IF-Filter for mobile radio using an on-chip clock. Lecture Notes in Computer Science, 783:510–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lingwood:1994:MDC

[572] S. D. Lingwood. A 65-MHz digital chipmatched-filter for DS-Spread spectrum applications. Lecture Notes in Computer Science, 783:522–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1994:AEV

[573] H.-D. Lin. Area-efficient Viterbi decoders for complex rate-k/n convolutional and trellis codes. Lecture Notes in Computer Science, 783:529– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Viterbi:1994:SHE

[574] A. J. Viterbi, A. M. Viterbi, K. S. Gilhousen, and E. Zehavi. Soft handoff extends CDMA cell coverage and increases reverse link capacity. Lecture Notes in Computer Science, 783:541– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Acampora:1994:AMM

[575] A. S. Acampora and M. Naghshineh. An architecture and methodology for mobile-executed cell hand-off in wireless ATM networks. Lecture Notes in Computer Science, 783:552–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kodratoff:1994:IAM

[576] Y. Kodratoff. Industrial applications of ML: Illustrations for the KAML dilemma and the CBR dream. Lecture Notes in Computer Science, 784:3–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neri:1994:KRM

[577] F. Neri and L. Saitta. Knowledge representation in machine learning. Lecture Notes in Computer Science, 784:20–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aha:1994:IIS

[578] D. W. Aha, S. Lapointe, C. X. Ling, and S. Matwin. Inverting implication with small training sets. Lecture Notes in Computer Science, 784:31–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biberman:1994:CSM

[579] Y. Biberman. A context similarity measure. Lecture Notes in Computer Science, 784:49–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borrajo:1994:ILC

[580] D. Borrajo and M. Veloso. Incremental learning of control knowledge for nonlinear problem solving. Lecture Notes in Computer Science, 784:64–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brazdil:1994:CAC

[581] P. Brazdil, J. Gama, and B. Henery. Characterizing the applicability of classification algorithms using meta-level learning. Lecture Notes in Computer Science, 784:83–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emde:1994:ILC

[582] W. Emde. Inductive learning of characteristic concept descriptions from small sets of classified examples. Lecture Notes in Computer Science, 784:103– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Furnkranz:1994:FRR

[583] J. Furnkranz. FOSSIL: A robust relational learner. Lecture Notes in Computer Science, 784:122–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herrmann:1994:MLS

[584] J. Herrmann, R. Ackermann, J. Peters, and D. Reipa. A multistrategy learning system and its integration into an interactive floorplanning tool. Lecture Notes in Computer Science, 784:138–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kohavi:1994:BIO

[585] R. Kohavi. Bottom-up induction of oblivious read-once decision graphs. Lecture Notes in Computer Science, 784:154–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kononenko:1994:EAA

[586] I. Kononenko. Estimating attributes: Analysis and extensions of RELIEF. Lecture Notes in Computer Science, 784:171–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeBlanc:1994:GFA

[587] G. Le Blanc. Generalization and formalization of an algorithm for detecting recursive relations in term sequences. Lecture Notes in Computer Science, 784:183–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malerba:1994:AEC

[588] D. Malerba, G. Semeraro, and F. Esposito. An analytic and empirical comparison of two methods for discovering probabilistic causal relationships. Lecture Notes in Computer Science, 784:198–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nienhuys-Cheng:1994:SPL

[589] S. H. Nienhuys-Cheng and M. Polman. Sample PAC-Learnability in model inference. Lecture Notes in Computer Science, 784:217–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oliver:1994:ADS

[590] J. J. Oliver and D. Hand. Averaging over decision stumps. Lecture Notes in Computer Science, 784:231–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfahringer:1994:CCI

[591] B. Pfahringer. Controlling constructive induction in CIPF: An MDL approach. Lecture Notes in Computer Science, 784:242–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sebag:1994:UCB

[592] M. Sebag. Using constraints to building version spaces. Lecture Notes in Computer Science, 784:257–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stahl:1994:UPI

[593] I. Stahl. On the utility of predicate invention in inductive logic programming. Lecture Notes in Computer Science, 784:272–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stroulia:1994:LPS

[594] E. Stroulia and A. K. Goel. Learning problem-solving concepts by reflecting on problem solving. Lecture Notes in Computer Science, 784:287–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanderLaag:1994:ENC

[595] P. R. J. van der Laag and S. H. Nienhuys-Cheng. Existence and nonexistence of complete refinement operators. Lecture Notes in Computer Science, 784:307–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wettschereck:1994:HNN

[596] D. Wettschereck. A hybrid nearestneighbor and nearest-hyperrectangle algorithm. Lecture Notes in Computer Science, 784:323–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berka:1994:AKA

[597] P. Berka and J. Ivanek. Automated knowledge acquisition for PROSPECTOR-like expert systems. Lecture Notes in Computer Science, 784:339–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brockmann:1994:RML

[598] W. Brockmann. On the role of machine learning in knowledge-based control. Lecture Notes in Computer Science, 784:343–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dzeroski:1994:DDG

[599] S. Dzeroski and I. Petrovksi. Discovering dynamics with genetic programming. Lecture Notes in Computer Science, 784:347–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elomaa:1994:GAF

[600] T. Elomaa and E. Ukkonen. A geometric approach to feature selection. Lecture Notes in Computer Science, 784: 351–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferretti:1994:IUR

[601] C. Ferretti and G. Mauri. Identifying unrecognizable regular languages by queries. Lecture Notes in Computer Science, 784:355–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gunetti:1994:ILL

[602] D. Gunetti and U. Trinchero. Intensional learning of logic programs. Lecture Notes in Computer Science, 784: 359–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirowatari:1994:PIG

[603] E. Hirowatari and S. Arikawa. Partially isomorphic generalization and analogical reasoning. Lecture Notes in Computer Science, 784:363–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jappy:1994:LRT

[604] P. Jappy, M. C. Daniel-Vatonne, O. Gascuel, and C. de la Higuera. Learning from recursive, tree structured examples. Lecture Notes in Computer Science, 784:367–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ketterlin:1994:CFC

[605] A. Ketterlin and J. J. Korczak. Concept formation in complex domains. Lecture Notes in Computer Science, 784:371–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kivinen:1994:ALH

[606] J. Kivinen, H. Mannila, E. Ukkonen, and J. Vilo. An algorithm for learning hierarchical classifiers. Lecture Notes in Computer Science, 784:375–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klopotek:1994:LBN

[607] M. A. Klopotek. Learning belief network structure from data under causal insufficiency. Lecture Notes in Computer Science, 784:379–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knoll:1994:CSP

[608] U. Knoll, G. Nakhaeizadeh, and B. Tausend. Cost-sensitive pruning of decision trees. Lecture Notes in Computer Science, 784:383–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1994:IBL

[609] C. Lee. An instance-based learning method for databases: An information theoretic approach. Lecture Notes in Computer Science, 784:387–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1994:ESG

[610] W. Z. Liu, A. P. White, and M. T. Hallissey. Early screening for gastric cancer using machine learning techniques. Lecture Notes in Computer Science, 784:391–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1994:DSU

[611] J. D. Martin. DP1: Supervised and unsupervised clustering. Lecture Notes in Computer Science, 784:395–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mladenic:1994:UML

[612] D. Mladenic, I. Bratko, R. J. Paul, and M. Grobelnik. Using machine learning techniques to interpret results from discrete event simulation. Lecture Notes in Computer Science, 784:399–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plaza:1994:FIM

[613] E. Plaza and J. L. Arcos. Flexible integration of multiple learning methods into a problem solving architecture. Lecture Notes in Computer Science, 784:403–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sarbo:1994:CS

[614] J. Sarbo and J. Farkas. Concept sublattices. Lecture Notes in Computer Science, 784:407–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulmeister:1994:PLC

[615] B. Schulmeister and F. Wysotzki. The piecewise linear classifier DIPOL92. Lecture Notes in Computer Science, 784:411–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shinohara:1994:CCG

[616] A. Shinohara. Complexity of computing generalized VC-Dimensions. Lecture Notes in Computer Science, 784: 415–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sommer:1994:LRC

[617] E. Sommer. Learning relations without closing the world. Lecture Notes in Computer Science, 784:419–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stahl:1994:PIL

[618] I. Stahl. Properties of inductive logic programming in function-free Horn logic. Lecture Notes in Computer Science, 784:423–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tausend:1994:RBI

[619] B. Tausend. Representing biases for inductive logic programming. Lecture Notes in Computer Science, 784:427– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tausend:1994:BTE

[620] B. Tausend. Biases and their effects in inductive logic programming. Lecture Notes in Computer Science, 784: 431–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vrain:1994:ILN

[621] C. Vrain and L. Martin. Inductive learning of normal clauses. Lecture Notes in Computer Science, 784:435– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goguen:1994:TAS

[622] J. A. Goguen and R. Diaconescu. Towards an algebraic semantics for the object paradigm. Lecture Notes in Computer Science, 785:1–29, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jouannaud:1994:RTS

[623] J.-P. Jouannaud. Rewriting techniques for software engineering. Lecture Notes in Computer Science, 785:30–52, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poigne:1994:IET

[624] A. Poigne. Identity and existence, and types in algebra — A survey of sorts. Lecture Notes in Computer Science, 785:53–78, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1994:OI

[625] E. G. Wagner. Overloading and inheritance. Lecture Notes in Computer Science, 785:79–97, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bettaz:1994:SBK

[626] M. Bettaz and G. Reggio. A SMoLCS based kit for defining high-level algebraic Petri nets. Lecture Notes in Computer Science, 785:98–112, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cerioli:1994:IVA

[627] M. Cerioli and G. Reggio. Institutions for very abstract specifications. Lecture Notes in Computer Science, 785:113– 127, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choppy:1994:ACA

[628] C. Choppy. About the “correctness” and “adequacy” of PLUSS specifications. Lecture Notes in Computer Science, 785:128–143, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clerici:1994:SCS

[629] S. Clerici, R. Jimenez, and F. Orejas. Semantic constructions in the specification language glider. Lecture Notes in Computer Science, 785:144–157, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conrad:1994:CST

[630] S. Conrad. On certification of specifications for TROLL light objects. Lecture Notes in Computer Science, 785:158– 172, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denker:1994:TTL

[631] G. Denker and M. Gogolla. Translating TROLL light concepts to MAUDE. Lecture Notes in Computer Science, 785:173–187, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrig:1994:AHL

[632] H. Ehrig, J. Padberg, and L. Ribeiro. Algebraic high level nets: Petri nets revisited. Lecture Notes in Computer Science, 785:188–206, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grosse-Rhode:1994:CSP

[633] M. Grosse-Rhode and U. Wolter. 2categorical specification of partial algebras. Lecture Notes in Computer Science, 785:207–219, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hennicker:1994:BAF

[634] R. Hennicker and F. Nickl. A behavioural algebraic framework for modular system design with reuse. Lecture Notes in Computer Science, 785:220– 234, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hermida:1994:FAC

[635] C. Hermida. On fibred adjunctions and completeness for fibred categories. Lecture Notes in Computer Science, 785:235–251, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levy:1994:IIN

[636] J. Levy and J. Agusti. Implementing inequality and nondeterministic specifications with bi-rewriting systems. Lecture Notes in Computer Science, 785:252–267, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1994:SBL

[637] J. Liu. A semantic basis of logicindependent transformation. Lecture Notes in Computer Science, 785:268– 279, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mosses:1994:UAA

[638] P. D. Mosses. Unified algebras and abstract syntax. Lecture Notes in Computer Science, 785:280–294, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

ParisiPresicce:1994:SIA

[639] F. Parisi Presicce and A. Pierantonio. Structured inheritance for algebraic class specifications. Lecture Notes in Computer Science, 785:295–309, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rieckhoff:1994:TTA

[640] C. Rieckhoff. Towards a theory for the animation of algebraic specification. Lecture Notes in Computer Science, 785:310–320, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schobbens:1994:SOP

[641] P.-Y. Schobbens. Second-order proof systems for algebraic specification languages. Lecture Notes in Computer Science, 785:321–336, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sernadas:1994:IOB

[642] A. Sernadas, J. F. Costa, and C. Sernadas. An institution of object behaviour. Lecture Notes in Computer Science, 785:337–350, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orbaek:1994:OOA

[643] Peter Orbaek. OASIS: An optimizing action-based compiler generator. Lecture Notes in Computer Science, 786: 1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moura:1994:ATA

[644] Hermano Moura and David A. Watt. Action transformations in the AC-TRESS compiler generator. Lecture Notes in Computer Science, 786:16–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hedin:1994:ODA

[645] Gorel Hedin. An overview of door attribute grammars. Lecture Notes in Computer Science, 786:31–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roussel:1994:CEA

[646] Gilles Roussel, Didier Parigot, and Martin Jourdan. Coupling evaluators for attribute coupled grammars. Lecture Notes in Computer Science, 786: 52–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanus:1994:TGO

[647] Michael Hanus. Towards the global optimization of functional logic programs. Lecture Notes in Computer Science, 786:68–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

George:1994:POB

[648] Lal George, Florent Guillame, and John H. Reppy. A portable and optimizing back end for the SML/NJ compiler. Lecture Notes in Computer Science, 786:83–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hogen:1994:EOC

[649] Guido Hogen and Rita Loogen. Efficient organization of control structures in distributed implementations. Lecture Notes in Computer Science, 786: 98–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ben-Asher:1994:IM

[650] Yosi Ben-Asher, Gudula Runger, Reinhard Wilhelm, and Assaf Schuster. Implementing 2DT on a multiprocessor. Lecture Notes in Computer Science, 786:113–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fauth:1994:GCS

[651] Andreas Fauth, Gunter Hommel, Carsten Muller, and Alois Knoll. Global code selection for directed acyclic graphs. Lecture Notes in Computer Science, 786:128–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Slowik:1994:CNL

[652] Adrian Slowik, Georg Piepenbrock, and Peter Pfahler. Compiling nested loops for limited connectivity VLIWs. Lecture Notes in Computer Science, 786:143–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ertl:1994:DES

[653] M. Anton Ertl and Andreas Krall. Delayed exceptions — speculative execution of trapping instructions. Lecture Notes in Computer Science, 786:158– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janssens:1994:IAI

[654] Gerda Janssens and Wim Simoens. On the implementation of abstract interpretation systems for (constraint) logic programs. Lecture Notes in Computer Science, 786:172–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheatham:1994:SAT

[655] Thomas Cheatham, Haiming Gao, and Dan Stefanescu. A suite of analysis tools based on a general purpose abstract interpreter. Lecture Notes in Computer Science, 786:188–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uhl:1994:FGF

[656] James S. Uhl and R. Nigel Horspool. Flow grammars — A flow analysis methodology. Lecture Notes in Computer Science, 786:203–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lammel:1994:PCP

[657] Ralf Lammel and Gunter Riedewald. Provable correctness of prototype interpreters in LDL. Lecture Notes in Computer Science, 786:218–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poetzsch-Heffter:1994:DEI

[658] Arnd Poetzsch-Heffter. Developing efficient interpreters based on formal language specifications. Lecture Notes in Computer Science, 786:233–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ringstrom:1994:GEC

[659] Johan Ringstrom, Peter Fritzson, and Mikael Pettersson. Generating an efficient compiler for a data parallel language from a denotational specification. Lecture Notes in Computer Science, 786:248–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parr:1994:ASS

[660] Terence J. Parr and Russel W. Quong. Adding semantic and syntactic predicates to LL(k): pred-LL(k). Lecture Notes in Computer Science, 786:263– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alt:1994:CCP

[661] Martin Alt, Uwe Abmann, and Hans van Someren. Cosy compiler phase embedding with the CoSy compiler model. Lecture Notes in Computer Science, 786:278–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franzle:1994:TPC

[662] Martin Franzle and Markus Muller-Olm. Towards provably correct code generation for a hard real-time programming language. Lecture Notes in Computer Science, 786:294–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Justiani:1994:SAD

[663] Justiani and Laurie J. Hendren. Supporting array dependence testing for an optimizing/parallelizing C compiler. Lecture Notes in Computer Science, 786:309–323, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benkner:1994:PAS

[664] Siegfried Benkner, Peter Brezany, and Hans Zima. Processing array statements and procedure interfaces in the PREPARE High Performance Fortran compiler. Lecture Notes in Computer Science, 786:324–338, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pineo:1994:PAS

[665] Patricia Pineo and Mary Lou Soffa. A practical approach to the symbolic debugging of parallelized code. Lecture Notes in Computer Science, 786:339– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duesterwald:1994:RCD

[666] Evelyn Duesterwald, Rajiv Gupta, and Mary Lou Soffa. Reducing the cost of data-flow analysis by congruence partitioning. Lecture Notes in Computer Science, 786:357–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Binkley:1994:ICP

[667] David Binkley. Interprocedural constant propagation using dependence graphs and a data-flow model. Lecture Notes in Computer Science, 786: 374–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reps:1994:SDV

[668] Thomas Reps. Solving demand versions of interprocedural analysis problems. Lecture Notes in Computer Science, 786:389–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mendlson:1994:CTI

[669] Abraham Mendlson, Ruth Shtokhamer, and Shlomit S. Pinter. Compile time instruction cache optimisations. Lecture Notes in Computer Science, 786: 404–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mahadevan:1994:ISR

[670] Uma Mahadevan and Sridhar Ramakrishnan. Instruction scheduling over regions: A framework for scheduling across basic blocks. Lecture Notes in Computer Science, 786:419–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoogerbrugge:1994:TTV

[671] Jan Hoogerbrugge and Henk Corporaal. Transport-triggering vs. operation-triggering. Lecture Notes in Computer Science, 786:435–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Comon:1994:OCT

[672] H. Comon and R. Treinen. Ordering constraints on trees. Lecture Notes in Computer Science, 787:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engelfriet:1994:GGT

[673] J. Engelfriet. Graph grammars and tree transducers. Lecture Notes in Computer Science, 787:15–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alessi:1994:TP

[674] F. Alessi. Type preorders. Lecture Notes in Computer Science, 787:37–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruscoli:1994:CCN

[675] P. Bruscoli, F. Levi, G. Levi, and M. C. Meo. Compilative constructive negation in constraint logic programs. Lecture Notes in Computer Science, 787:52–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cournier:1994:NLA

[676] A. Cournier and M. Habib. A new linear algorithm for modular decomposition. Lecture Notes in Computer Science, 787:68–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deGroote:1994:CTL

[677] P. de Groote. A CPS-Translation of the lambda-mu-calculus. Lecture Notes in Computer Science, 787:85–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drewes:1994:LBG

[678] F. Drewes. A lower bound on the growth of functions computed by tree transductions. Lecture Notes in Computer Science, 787:100–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esparza:1994:DMC

[679] J. Esparza. On the decidability of model checking for several mu-calculi and Petri nets. Lecture Notes in Computer Science, 787:115–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giancarlo:1994:GPT

[680] R. Giancarlo and F. Mignosi. Generalizations of the periodicity theorem of fine and wilf. Lecture Notes in Computer Science, 787:130–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heckmann:1994:PD

[681] R. Heckmann. Probabilistic domains. Lecture Notes in Computer Science, 787:142–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofbauer:1994:SRT

[682] D. Hofbauer, M. Huber, and G. Kucherov. Some results on top-context-free tree languages. Lecture Notes in Computer Science, 787:157–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khasidashvili:1994:HOR

[683] Z. Khasidashvili. On higher order recursive program schemes. Lecture Notes in Computer Science, 787:172– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klarlund:1994:GDT

[684] N. Klarlund and M. I. Schwartzbach. Graphs and decidable transductions based on edge constraints. Lecture Notes in Computer Science, 787:187– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuske:1994:NAC

[685] D. Kuske. Nondeterministic automata with concurrency relations and domains. Lecture Notes in Computer Science, 787:202–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merlini:1994:ACP

[686] D. Merlini, R. Sprugnoli, and M. C. Verri. Algebraic and combinatorial properties of simple, coloured walks. Lecture Notes in Computer Science, 787:218–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Metivier:1994:PAE

[687] Y. Metivier and N. Saheb. Probabilistic analysis of an election algorithm in a tree. Lecture Notes in Computer Science, 787:234–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Noll:1994:FOE

[688] T. Noll. On the first-order equivalence of call-by-name and call-by-value. Lecture Notes in Computer Science, 787:246–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohlebush:1994:MCC

[689] E. Ohlebush. On the modularity of confluence of constructor-sharing term rewriting systems. Lecture Notes in Computer Science, 787:261–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palsberg:1994:GPA

[690] J. Palsberg. Global program analysis in constraint form. Lecture Notes in Computer Science, 787:276–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peled:1994:PSP

[691] D. Peled. On projective and separable properties. Lecture Notes in Computer Science, 787:291–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rambow:1994:RHD

[692] O. Rambow and G. Satta. A rank hierarchy for deterministic tree-walking transducers. Lecture Notes in Computer Science, 787:308–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ratoandromanana:1994:SPL

[693] B. Ratoandromanana and D. Robilliard. Superposition in picture languages. Lecture Notes in Computer Science, 787:322–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sorensen:1994:GBD

[694] M. H. Sorensen. A grammar-based data-flow analysis to stop deforestation. Lecture Notes in Computer Science, 787:335–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abadi:1994:TPOa

[695] M. Abadi and L. Cardelli. A theory of primitive objects: second-order systems. Lecture Notes in Computer Science, 788:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Milner:1994:PNG

[696] R. Milner. Pi-nets: a graphical form of pi-calculus. Lecture Notes in Computer Science, 788:26–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amtoft:1994:LTR

[697] T. Amtoft. Local type reconstruction by means of symbolic fixed point iteration. Lecture Notes in Computer Science, 788:43–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersen:1994:APA

[698] H. R. Andersen and M. Mendler. An asynchronous process algebra with multiple clocks. Lecture Notes in Computer Science, 788:58–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andrews:1994:FII

[699] J. H. Andrews. Foundational issues in implementing constraint logic programming systems. Lecture Notes in Computer Science, 788:74–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berthomieu:1994:PBM

[700] B. Berthomieu and T. Le Sergent. Programming with behaviors in an ML framework: the syntax and semantics of LCS. Lecture Notes in Computer Science, 788:89–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bidoit:1994:CBS

[701] M. Bidoit, R. Hennicker, and M. Wirsing. Characterizing behavioural semantics and abstractor semantics. Lecture Notes in Computer Science, 788:105– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boerio:1994:EPT

[702] L. Boerio. Extending pruning techniques to polymorphic second order lambda-calculus. Lecture Notes in Computer Science, 788:120–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bohm:1994:LDF

[703] C. Bohm, A. Piperno, and S. Guerrini. lambda-definition of function(al)s by normal forms. Lecture Notes in Computer Science, 788:135–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buth:1994:SSD

[704] K.-H. Buth. Simulation of SOS definitions with term rewriting systems. Lecture Notes in Computer Science, 788: 150–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cicerone:1994:SMS

[705] S. Cicerone and F. Parisi Presicce. Strategies in modular system design by interface rewriting. Lecture Notes in Computer Science, 788:165–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corsini:1994:SMC

[706] M.-M. Corsini and A. Rauzy. Symbolic model checking and constraint logic programming: a cross-fertilization. Lecture Notes in Computer Science, 788:180–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiPierro:1994:LDS

[707] A. Di Pierro and C. Palamidessi. A logical denotational semantics for constraint logic programming. Lecture Notes in Computer Science, 788:195– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fradet:1994:CHS

[708] P. Fradet. Compilation of head and strong reduction. Lecture Notes in Computer Science, 788:211–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giegerich:1994:STF

[709] R. Giegerich and S. Kurtz. Suffix trees in the functional programming paradigm. Lecture Notes in Computer Science, 788:225–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hall:1994:TCH

[710] C. Hall, K. Hammond, S. Peyton Jones, and P. Wadler. Type classes in haskell. Lecture Notes in Computer Science, 788:241–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hankin:1994:LTI

[711] C. Hankin and D. Le Metayer. Lazy type inference for the strictness analysis of lists. Lecture Notes in Computer Science, 788:257–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanus:1994:LUS

[712] M. Hanus. Lazy unification with simplification. Lecture Notes in Computer Science, 788:272–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henglein:1994:PBT

[713] F. Henglein and C. Mossin. Polymorphic binding-time analysis. Lecture Notes in Computer Science, 788:287– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jay:1994:STS

[714] C. B. Jay and J. R. B. Cockett. Shapely types and shape polymorphism. Lecture Notes in Computer Science, 788:302–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeuring:1994:BGA

[715] J. Jeuring and D. Swierstra. Bottomup grammar analysis: a functional formulation. Lecture Notes in Computer Science, 788:317–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kahrs:1994:FCP

[716] S. Kahrs. First-class polymorphism for ML. Lecture Notes in Computer Science, 788:333–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kennedy:1994:DT

[717] A. Kennedy. Dimension types. Lecture Notes in Computer Science, 788: 348–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

King:1994:SAS

[718] A. King. A synergistic analysis for sharing and groundness which traces linearity. Lecture Notes in Computer Science, 788:363–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1994:PCS

[719] B. Z. Li. A pi-calculus specification of Prolog. Lecture Notes in Computer Science, 788:379–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1994:LFE

[720] W. Li. A logical framework for evolution of specifications. Lecture Notes in Computer Science, 788:394–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MacQueen:1994:SHO

[721] D. B. MacQueen and M. Tofte. A semantics for higher-order functors. Lecture Notes in Computer Science, 788: 409–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreau:1994:PMA

[722] L. Moreau. The PCKS-machine: an abstract machine for sound evaluation of parallel functional programs with firstclass continuations. Lecture Notes in Computer Science, 788:424–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muck:1994:TCF

[723] A. Muck, T. Streicher, and H. C. R. Lock. A tiny constraint functional logic language and its continuation semantics. Lecture Notes in Computer Science, 788:439–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OHearn:1994:FAT

[724] P. W. O’Hearn and J. G. Riecke. Fully abstract translations and parametric polymorphism. Lecture Notes in Computer Science, 788:454–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prasad:1994:BP

[725] K. V. S. Prasad. Broadcasting with priority. Lecture Notes in Computer Science, 788:469–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sorensen:1994:TUP

[726] M. H. Sorensen, R. Gluck, and N. D. Jones. Towards unifying partial evaluation, deforestation, supercompilation, and GPC. Lecture Notes in Computer Science, 788:485–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walker:1994:APP

[727] D. Walker. Algebraic proofs of properties of objects. Lecture Notes in Computer Science, 788:501–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abramsky:1994:FAP

[728] Samson Abramsky, Pasquale Malacaria, and Radha Jagadeesan. Full abstraction for PCF. Lecture Notes in Computer Science, 789:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dezani-Ciancaglini:1994:FAS

[729] Mariangiola Dezani-Ciancaglini, Ugo de’Liguoro, and Adolfo Piperno. Fully abstract semantics for concurrent lambda-calculus. Lecture Notes in Computer Science, 789:16–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Howe:1994:OAC

[730] Douglas J. Howe and Scott D. Stoller. An operational approach to combining classical set theory and functional programming languages. Lecture Notes in Computer Science, 789:36–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1994:MTE

[731] Mark P. Jones. ML typing, explicit polymorphism and qualified types. Lecture Notes in Computer Science, 789: 56–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Riecke:1994:ETS

[732] Jon G. Riecke and Ramesh Subrahmanyam. Extensions to type systems can preserve operational equivalences. Lecture Notes in Computer Science, 789:76–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kanellakis:1994:CPD

[733] Paris C. Kanellakis and Dina Q. Goldin. Constraint programming and database query languages. Lecture Notes in Computer Science, 789:96–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoa:1994:IRL

[734] Alain Hui Bon Hoa. Intuitionistic resolution for a logic programming language with scoping constructs. Lecture Notes in Computer Science, 789: 121–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertot:1994:PP

[735] Yves Bertot, Gilles Kahn, and Laurent Thery. Proof by pointing. Lecture Notes in Computer Science, 789: 141–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Korver:1994:CCV

[736] Henri Korver and Jan Springintveld. A computer-checked verification of Milner’s scheduler. Lecture Notes in Computer Science, 789:161–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sato:1994:PFL

[737] Masahiko Sato. A purely functional language with encapsulated assignment. Lecture Notes in Computer Science, 789:179–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rao:1994:STH

[738] M. R. K. Krishna Rao. Simple termination of hierarchical combinations of term rewriting systems. Lecture Notes in Computer Science, 789:203– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yan:1994:SAI

[739] Mei Tang Yan and Pierre Jouvelot. Separate abstract interpretation for control-flow analysis. Lecture Notes in Computer Science, 789:224–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cartwright:1994:EDL

[740] Robert Cartwright and Matthias Felleisen. Extensible denotational language specifications. Lecture Notes in Computer Science, 789:244–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsuiki:1994:NCO

[741] Hideki Tsuiki. A normalizing calculus with overloading and subtyping. Lecture Notes in Computer Science, 789:273–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abadi:1994:TPOb

[742] Martin Abadi and Luca Cardelli. A theory of primitive objects — untyped and first-order systems. Lecture Notes in Computer Science, 789:296– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Remy:1994:POM

[743] Didier Remy. Programming objects with ML-ART, an extension to ML with abstract and record. Lecture Notes in Computer Science, 789:321– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kung:1994:TSL

[744] Chen Kung and Martin Odersky. A type system for a lambda calculus with assignments. Lecture Notes in Computer Science, 789:347–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yonezawa:1994:TPC

[745] Akinori Yonezawa. Theory and practice of concurrent object-oriented computing. Lecture Notes in Computer Science, 789:365–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asperti:1994:FRI

[746] Andrea Asperti and Cosimo Laneve. The family relation in interaction systems. Lecture Notes in Computer Science, 789:366–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goltz:1994:SSA

[747] Ursula Goltz, Roberto Gorrieri, and Arend Rensink. On syntactic and semantic action refinement. Lecture Notes in Computer Science, 789:385– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sangiorgi:1994:LTC

[748] Davide Sangiorgi. Locality and trueconcurrency in calculi for mobile processes. Lecture Notes in Computer Science, 789:405–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bosscher:1994:TRP

[749] D. J. B. Bosscher. Term rewriting properties of SOS axiomatisations. Lecture Notes in Computer Science, 789:425– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fokkink:1994:TTF

[750] Willem Jan Fokkink. The tyft/tyxt format reduces to tree rules. Lecture Notes in Computer Science, 789:440– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huttel:1994:UEB

[751] Hans Huttel. Undecidable equivalences for basic parallel processes. Lecture Notes in Computer Science, 789:454– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takahashi:1994:NPT

[752] Masako Takahashi, Y. Akama, and S. Hirokawa. Normal proofs and their grammar. Lecture Notes in Computer Science, 789:465–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbanera:1994:SLC

[753] Franco Barbanera and Stefano Berardi. A symmetric lambda calculus for “classical” program extraction. Lecture Notes in Computer Science, 789:495– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rehof:1994:C

[754] Niels Jakob Rehof and Morten Heine Sorensen. The λ_{Δ}-Calculus. Lecture Notes in Computer Science, 789:516– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ariola:1994:SDU

[755] Zena Ariola, Richard Kennaway, Jan Willem Klop, Ronan Sleep, and Fer-Jan de Vries. Syntactic definitions of undefined: On defining the undefined. Lecture Notes in Computer Science, 789:543–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gardner:1994:DNR

[756] Philippa Gardner. Discovering needed reductions using type theory. Lecture Notes in Computer Science, 789:555– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vardi:1994:NAA

[757] Moshe Y. Vardi. Nontraditional applications of automata theory. Lecture Notes in Computer Science, 789:575– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1994:APO

[758] Kim Ritter Wagner. Abstract preorders. Lecture Notes in Computer Science, 789:598–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Islam:1994:CMR

[759] Amitavo Islam and Wesley Phoa. Categorical models of relational databases I: Fibrational formulation, schema integration. Lecture Notes in Computer Science, 789:618–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kanovich:1994:PNH

[760] Max I. Kanovich. Petri nets, Horn programs, linear logic, and vector games. Lecture Notes in Computer Science, 789:642–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sekiguchi:1994:CTI

[761] Tatsurou Sekiguchi and Akinori Yonezawa. A complete type inference system for subtyped recursive types. Lecture Notes in Computer Science, 789:667–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damm:1994:SUT

[762] Flemming M. Damm. Subtyping with union types, intersection types and recursive types. Lecture Notes in Computer Science, 789:687–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damiani:1994:DIT

[763] Ferruccio Damiani and Paola Giannini. A decidable intersection type system based on relevance. Lecture Notes in Computer Science, 789:707–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manna:1994:TVD

[764] Zohar Manna and Amir Pnueli. Temporal verification diagrams. Lecture Notes in Computer Science, 789:726– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bolignano:1994:STC

[765] Dominique Bolignano and Mourad Debabi. A semantic theory for concurrent ML. Lecture Notes in Computer Science, 789:766–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Honda:1994:RCC

[766] Kohei Honda and Nobuko Yoshida. Replication in concurrent combinators. Lecture Notes in Computer Science, 789:786–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peron:1994:TIN

[767] Adriano Peron and Andrea Maggiolo-Schettini. Transitions as interrupts: A new semantics for timed statecharts. Lecture Notes in Computer Science, 789:806–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonsangue:1994:RMP

[768] Marcello M. Bonsangue and Joost N. Kok. Relating multifunctions and predicate transformers through closure operators. Lecture Notes in Computer Science, 789:822–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fisher:1994:NTO

[769] K. Fisher and J. C. Mitchell. Notes on typed object-oriented programming. Lecture Notes in Computer Science, 789:844–885, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer:1994:OTC

[770] Albert R. Meyer. Observing truly concurrent processes. Lecture Notes in Computer Science, 789:886–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wegner:1994:MPI

[771] P. Wegner. Models and paradigms of interaction. Lecture Notes in Computer Science, 791:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fowler:1994:ACG

[772] R. J. Fowler. Architectural convergence and the granularity of objects in distributed systems. Lecture Notes in Computer Science, 791:33–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jul:1994:SDO

[773] E. Jul. Separation of distribution and objects. Lecture Notes in Computer Science, 791:47–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zeidler:1994:ISO

[774] C. Zeidler and B. Fank. Integrating structural and operational programming to manage distributed systems. Lecture Notes in Computer Science, 791:55–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bourgois:1994:CCC

[775] M. Bourgois, J.-M. Andreoli, and R. Pareschi. Concurrency and communication: Choices in implementing the coordination language LO. Lecture Notes in Computer Science, 791:73–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tokoro:1994:TLF

[776] M. Tokoro and K. Takashio. Toward languages and formal systems for distributed computing. Lecture Notes in Computer Science, 791:93–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wing:1994:DRT

[777] J. M. Wing. Decomposing and recomposing transactional concepts. Lecture Notes in Computer Science, 791:111– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rakotonirainy:1994:DEC

[778] A. Rakotonirainy. DPL to express a concurrency control using transaction and object semantics. Lecture Notes in Computer Science, 791:122–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bijnens:1994:RIS

[779] S. Bijnens, W. Joosen, and P. Verbaeten. A reflective invocation scheme to realise advanced object management. Lecture Notes in Computer Science, 791:139–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aksit:1994:AOI

[780] M. Aksit, K. Wakita, J. Bosch, and L. Bergmans. Abstracting object interactions using composition filters. Lecture Notes in Computer Science, 791: 152–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandt:1994:OOD

[781] S. Brandt and O. L. Madsen. Objectoriented distributed programming in BETA. Lecture Notes in Computer Science, 791:185–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maffeis:1994:FSD

[782] S. Maffeis. A flexible system design to support object-groups and objectoriented distributed programming. Lecture Notes in Computer Science, 791: 213–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garbinato:1994:DPG

[783] B. Garbinato, R. Guerraoui, and K. R. Mazouni. Distributed programming in GARF. Lecture Notes in Computer Science, 791:225–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fazzolare:1994:OOE

[784] M. Fazzolare, B. G. Humm, and R. D. Ranson. Object-oriented extendibility in hermes/ST, a transactional distributed programming environment. Lecture Notes in Computer Science, 791:240–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hayashi:1994:LLC

[785] Susumu Hayashi and Yukihide Takayama. Lifschitz’ logic of calculable numbers and optimizations in program extraction. Lecture Notes in Computer Science, 792:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hagiya:1994:IA

[786] Masami Hagiya and Yozo Toda. On implicit arguments. Lecture Notes in Computer Science, 792:10–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yasugi:1994:FST

[787] Mariko Yasugi and Susumu Hayashi. A functional system with transfinitely defined types. Lecture Notes in Computer Science, 792:31–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakano:1994:NDC

[788] Hiroshi Nakano. The non-deterministic catch and throw mechanism and its subject reduction property. Lecture Notes in Computer Science, 792:61–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sato:1994:CLL

[789] Masahiko Sato and Yukiyoshi Kameyama. Conservativeness of lambda over lambdasigma-calculus. Lecture Notes in Computer Science, 792:73–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nishizaki:1994:MFC

[790] Shin ya Nishizaki. ML with first-class environments and its type inference algorithm. Lecture Notes in Computer Science, 792:95–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takahashi:1994:SPG

[791] Masako Takahashi. A simple proof of the genericity lemma. Lecture Notes in Computer Science, 792:117–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weyhrauch:1994:LFS

[792] Richard W. Weyhrauch and Carolyn Talcott. The logic of FOL systems: Formulated in set theory. Lecture Notes in Computer Science, 792:119– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hasegawa:1994:WOA

[793] Ryu Hasegawa. Well-ordering of algebras and Kruskal’s theorem. Lecture Notes in Computer Science, 792:133– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Igarashi:1994:LAE

[794] Shigeru Igarashi, Tetsuya Mizutani, Takashi Tsuji, and Chiharu Hosono. On locomorphism in analytical equivalence theory. Lecture Notes in Computer Science, 792:173–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tomita:1994:ASH

[795] Kohji Tomita, Takashi Tsuji, and Shigeru Igarashi. Analysis of a software/hardware system by tense arithmetic. Lecture Notes in Computer Science, 792:188–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1994:EPT

[796] Neil D. Jones. The essence of program transformation by partial evaluation and driving. Lecture Notes in Computer Science, 792:206–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mason:1994:PTC

[797] Ian A. Mason and Carolyn Talcott. Program transformation via contextual assertions. Lecture Notes in Computer Science, 792:225–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kobayashi:1994:CTM

[798] Kojiro Kobayashi. On coding theorems with modified length functions. Lecture Notes in Computer Science, 792: 255–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kasai:1994:TFC

[799] Takumi Kasai, Shusaku Sawato, and Shigeki Iwata. Thirty four comparisons are required to sort 13 items. Lecture Notes in Computer Science, 792:260– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blake:1994:CPI

[800] I. F. Blake, S. Gao, and R. Lambert. Constructive problems for irreducible polynomials over finite fields. Lecture Notes in Computer Science, 793:1–23, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Polemi:1994:SAC

[801] D. Polemi and T. Sakkalis. Singular algebraic curves over finite fields. Lecture Notes in Computer Science, 793:24–37, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zarowski:1994:DFS

[802] C. J. Zarowski. A divisionless form of the Schur Berlekamp–Massey algorithm. Lecture Notes in Computer Science, 793:38–44, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sablatash:1994:ECC

[803] M. Sablatash. Error-control coding, modulation and equalization for alldigital advanced television: state of the art and future possibilities. Lecture Notes in Computer Science, 793:45–66, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drolet:1994:FFA

[804] G. Drolet. A finite field arithmetic unit VLSI chip. Lecture Notes in Computer Science, 793:67–77, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hendessi:1994:SAA

[805] F. Hendessi and M. R. Aref. A successful attack against the DES. Lecture Notes in Computer Science, 793:78–90, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1994:LCH

[806] S. Lin, S. Rajpal, and D. J. Rhee. Low-complexity and high-performance multilevel coded modulation for the AWGN and Rayleigh fading channels. Lecture Notes in Computer Science, 793:91–108, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1994:CCF

[807] V. S. Lin and G. J. Pottie. Channel coding for a frequency-hopped wireless transceiver. Lecture Notes in Computer Science, 793:109–124, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boudreau:1994:UVN

[808] D. Boudreau. A unified view of noncoherent detection and differential detection of phase modulated signals. Lecture Notes in Computer Science, 793:125–135, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fattouche:1994:ERC

[809] M. Fattouche and H. Zaghloul. Error reduction of coded DPSK over a fading channel. Lecture Notes in Computer Science, 793:136–152, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferland:1994:PBC

[810] G. Ferland and J.-Y. Chouinard. Performance of BCH codes with DES encryption in a digital mobile channel. Lecture Notes in Computer Science, 793:153–172, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wautier:1994:RCC

[811] A. Wautier, J.-C. Dany, and C. Mourot. Reduced-complexity channels estimation and equalization for TDMA mobile radio systems. Lecture Notes in Computer Science, 793:173–190, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trabelsi:1994:PPS

[812] C. Trabelsi and A. Yongacoglu. Probability of packet success for asynchronous DS/CDMA with block and convolutional codes. Lecture Notes in Computer Science, 793:191–202, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Markman:1994:SDI

[813] I. Markman and J. B. Anderson. Sequential decoding on intersymbol interference channels under the Pe-criterion. Lecture Notes in Computer Science, 793:203–222, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lodge:1994:SCC

[814] J. Lodge, R. Young, and P. Guinand. Separable concatenated codes with iterative map filtering. Lecture Notes in Computer Science, 793:223–240, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roy:1994:ASC

[815] S. Roy and P. Fortier. Adapting the strongly-connected trellis concept for use with trellis-coded modulation. Lecture Notes in Computer Science, 793:241–252, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1994:VRP

[816] F. Chan and D. Haccoun. Variablerate punctured trellis-coded modulation and applications. Lecture Notes in Computer Science, 793:253–269, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sorokine:1994:TBD

[817] V. Sorokine, F. R. Kschischang, and V. Durand. Trellis-based decoding of binary linear block codes. Lecture Notes in Computer Science, 793:270– 286, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Le-Ngoc:1994:PDU

[818] T. Le-Ngoc, M. Jia, and A. Benyamin-Seeyar. Permutation decoding using primitive elements as multipliers. Lecture Notes in Computer Science, 793:287–295, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghazi-Moghaddam:1994:OCH

[819] F. Ghazi-Moghaddam, I. Lambadaris, and J. F. Hayes. Overflow constraint in hybrid nodes with movable boundary scheme. Lecture Notes in Computer Science, 793:296–309, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zeytinoglu:1994:SAW

[820] M. Zeytinoglu and Y. C. Hsu. Statistical analysis of wavelet transform coded multi-channel signals for transmission over ATM packet networks. Lecture Notes in Computer Science, 793:310– 324, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamoun:1994:SAT

[821] F. Kamoun and M. M. Ali. Statistical analysis of the traffic generated by the superposition of N independent interrupted Poisson processes. Lecture Notes in Computer Science, 793:325–342, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ross:1994:TIA

[822] J. A. F. Ross and D. P. Taylor. Throughput increase of ALOHA-based systems with multiuser sequence detection. Lecture Notes in Computer Science, 793:343–358, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1994:ART

[823] X. Huang and H. Leung. Autoregressive transfer function identification using chaos. Lecture Notes in Computer Science, 793:359–367, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inkol:1994:NAS

[824] R. J. Inkol and R. H. Saper. A new algorithm for signal classification using phase information. Lecture Notes in Computer Science, 793:368–384, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khandani:1994:EAF

[825] A. K. Khandani, P. Kabal, and E. Dubois. Efficient algorithms for fixed-rate entropy-coded vector quantization. Lecture Notes in Computer Science, 793:385–394, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trivedi:1994:TTR

[826] K. S. Trivedi, B. R. Haverkort, A. Rindos, and V. Mainkar. Techniques and tools for reliability and performance evaluation: Problems and perspectives. Lecture Notes in Computer Science, 794:1–24, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reed:1994:EAP

[827] D. A. Reed. Experimental analysis of parallel systems: Techniques and open problems. Lecture Notes in Computer Science, 794:25–51, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carmona:1994:PEB

[828] A. Carmona, L. Domingo, R. Macau, and R. Puigjaner. Performance experiences of the Barcelona Olympic Games computer system. Lecture Notes in Computer Science, 794:52–75, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bemmerl:1994:PTI

[829] T. Bemmerl and B. Ries. Performance tools on Intel scalable high performance computing systems. Lecture Notes in Computer Science, 794:76–88, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franken:1994:MAM

[830] L. J. N. Franken, R. H. Pijpers, and B. R. M. Haverkort. Modelling aspects of model-based dynamic QoS management by the performability manager. Lecture Notes in Computer Science, 794:89–110, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braband:1994:WTD

[831] J. Braband. Waiting time distributions for processor sharing queues with state-dependent arrival and service rates. Lecture Notes in Computer Science, 794:111–122, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonkers:1994:QMP

[832] H. Jonkers. Queueing models of parallel applications: The glamis methodology. Lecture Notes in Computer Science, 794:123–138, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malony:1994:ASA

[833] A. D. Malony, V. Mertsiotakis, and A. Quick. Automatic scalability analysis of parallel programs based on modelling techniques. Lecture Notes in Computer Science, 794:139–158, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

AjmoneMarsan:1994:SAC

[834] M. Ajmone Marsan, R. Lo Cigno, M. Munafo, and A. Tonietti. Simulation of ATM computer networks with CLASS. Lecture Notes in Computer Science, 794:159–179, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1994:NTM

[835] V. Gupta and E. Schenfeld. NetSim — A tool for modeling the performance of circuit switched multicomputer networks. Lecture Notes in Computer Science, 794:180–192, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garcia-Haro:1994:AEP

[836] J. Garcia-Haro, R. Marin-Sillue, and J. L. Melus-Moreno. ATMSWSIM: An efficient, portable and expandable ATM SWitch SIMulator tool. Lecture Notes in Computer Science, 794:193– 212, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Serrano:1994:MPE

[837] M. J. Serrano, W. Yamamoto, R. C. Wood, and M. Nemirovsky. A model for performance estimation in a multistreamed superscalar processor. Lecture Notes in Computer Science, 794: 213–230, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferscha:1994:AEP

[838] A. Ferscha and G. Chiola. Accelerating the evaluation of parallel program performance models using distributed simulation. Lecture Notes in Computer Science, 794:231–252, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1994:ALD

[839] T. D. Wagner and B. M. Carlson. An algorithm for off-line detection of phases in execution profiles. Lecture Notes in Computer Science, 794:253– 265, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caselli:1994:ESM

[840] S. Caselli, G. Conte, F. Bonardi, and M. Fontanesi. Experiences on SIMD massively parallel GSPN analysis. Lecture Notes in Computer Science, 794:266–283, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wabnig:1994:PPP

[841] H. Wabnig and G. Haring. PAPS — the parallel program performance prediction toolset. Lecture Notes in Computer Science, 794:284–304, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gunther:1994:QTA

[842] N. J. Gunther. qcomp: A tool for assessing online transaction processing scalability. Lecture Notes in Computer Science, 794:305–320, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bause:1994:QTQ

[843] F. Bause and P. Kemper. QPN-Tool for qualitative and quantitative analysis of queueing Petri nets. Lecture Notes in Computer Science, 794:321–334, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Donatelli:1994:SAE

[844] S. Donatelli, G. Franceschinis, N. Mazzocca, and S. Russo. Software architecture of the EPOCA integrated environment. Lecture Notes in Computer Science, 794:335–352, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gilmore:1994:PWT

[845] S. Gilmore and J. Hillston. The PEPA workbench: A tool to support a process algebra-based approach to performance modelling. Lecture Notes in Computer Science, 794:353–368, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grimsrud:1994:AMR

[846] K. Grimsrud, J. Archibald, R. Frost, and B. Nelson. On the accuracy of memory reference models. Lecture Notes in Computer Science, 794:369– 388, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fahringer:1994:AEN

[847] T. Fahringer. Automatically estimating network contention of parallel programs. Lecture Notes in Computer Science, 794:389–405, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tron:1994:MCC

[848] C. Tron and B. Plateau. Modelling of communication contention in multiprocessors. Lecture Notes in Computer Science, 794:406–424, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanderAalst:1994:UIT

[849] W. M. P. Van der Aalst. Using interval timed coloured Petri nets to calculate performance bounds. Lecture Notes in Computer Science, 794:425–444, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merlo:1994:APT

[850] A. P. Merlo and P. H. Worley. Analyzing PICL trace data with MEDEA. Lecture Notes in Computer Science, 794:445–463, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fox:1994:IPO

[851] G. C. Fox and K. Mills. Information processing and opportunities for HPCN use in industry. Lecture Notes in Computer Science, 796:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Williams:1994:ISD

[852] D. O. Williams. Information superhighways — does Europe need them, can Europe build them? Lecture Notes in Computer Science, 796:15–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hillmann:1994:SMF

[853] M. Hillmann and J. Weiher. Simulation of metal forming processes with respect to MPP-Systems. Lecture Notes in Computer Science, 796:21–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herken:1994:HIQ

[854] R. Herken. High image quality rendering on scalable parallel systems. Lecture Notes in Computer Science, 796: 31–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duncan:1994:PPH

[855] S. H. Duncan, P. L. Gordon, E. J. Zaluska, and S. I. Edwards. Parallel processing in high integrity aircraft engine control. Lecture Notes in Computer Science, 796:32–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garcia:1994:ESU

[856] G. D. Garcia, F. Suarez, M. Garcia, E. Lasso, R. Guzman, T. Carden, D. Watson, C. McGregor, F. Obeso, and M. Tarrio. An embedded solution using high-performance computing for cost effective on-line real-time monitoring of industrial processes. Lecture Notes in Computer Science, 796:40–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Colbrook:1994:EEE

[857] A. Colbrook, M. Lemke, H. Mierendorff, K. Stuben, C.-A. Thole, and O. Thomas. EUROPORT — Esprit European porting projects. Lecture Notes in Computer Science, 796:46–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lonsdale:1994:CRP

[858] G. Lonsdale, J. Clinckemaillie, S. Vlachoutsis, and J. Dubois. Communication requirements in parallel crashworthiness simulation. Lecture Notes in Computer Science, 796:55–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boston:1994:PFC

[859] I. Boston, M. Surridge, and C. Upstill. PAFEC-FE — A commercial parallel finite element package. Lecture Notes in Computer Science, 796:62–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bennani:1994:SLS

[860] M. Bennani, T. Braconnier, and J. C. Dunyach. Solving large-scale nonnormal eigenproblems in the aeronautical industry using parallel BLAS. Lecture Notes in Computer Science, 796:72–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krafczyk:1994:SWC

[861] M. Krafczyk and E. Rank. Supercomputing on a workstation cluster: A parallelized lattice-gas solver for transient Navier–Stokes-Flow. Lecture Notes in Computer Science, 796:78–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grandinetti:1994:PTL

[862] L. Grandinetti, F. Guerriero, and R. Musmanno. Parallel techniques for large-scale nonlinear network optimization. Lecture Notes in Computer Science, 796:84–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watermann:1994:ECP

[863] A. Watermann and J. Altes. Experiences concerning the parallelization of the finite element code SMART. Lecture Notes in Computer Science, 796: 92–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daehn:1994:EEC

[864] W. Daehn. Electrosmog and electromagnetic CAD. Lecture Notes in Computer Science, 796:94–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klix:1994:TDS

[865] W. Klix, R. Dittmann, and R. Stenzel. Three-dimensional simulation of semiconductor devices. Lecture Notes in Computer Science, 796:99–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krauss:1994:DAT

[866] P. A. Krauss. A distributed automatic test pattern generation system. Lecture Notes in Computer Science, 796:105–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luksch:1994:DLS

[867] P. Luksch. Distributed logic simulation — A test environment for the evaluation of different approaches. Lecture Notes in Computer Science, 796:111– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hunger:1994:MHG

[868] A. Hunger and F. Muller. Modelling hierarchy as guideline for parallel simulation. Lecture Notes in Computer Science, 796:117–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daniels:1994:LIF

[869] H. Daniels, A. Peters, O. Bergen, and C. Forkel. Large implicit finite element simulations of time-dependent incompressible flows on scalable parallel systems. Lecture Notes in Computer Science, 796:119–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thevenin:1994:PSR

[870] D. Thevenin, F. Behrendt, U. Maas, and J. Warnatz. Parallel simulation of reacting flows using detailed chemistry. Lecture Notes in Computer Science, 796:125–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haberhauer:1994:UCF

[871] S. Haberhauer. Unstructured computational fluid dynamics on massivelyparallel SIMD-Systems. Lecture Notes in Computer Science, 796:131–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deNicola:1994:APM

[872] C. de Nicola, G. De Pietro, and M. Giuliani. An advanced parallel multiblock code for the solution of 2D flowfields. Lecture Notes in Computer Science, 796:139–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oosterlee:1994:PMR

[873] C. W. Oosterlee, H. Ritzdorf, A. Schuller, and B. Steckel. Parallel multigrid results for Euler equations and grid partitioning into a large number of blocks. Lecture Notes in Computer Science, 796:145–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giraud:1994:PDI

[874] L. Giraud and G. M. Manzini. Parallel distributed implementations of 2D explicit Euler solvers. Lecture Notes in Computer Science, 796:151–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Issman:1994:PME

[875] E. Issman, G. Degrez, and J. de Keyser. A parallel multiblock Euler/Navier– Stokes solver on a cluster of workstations using PVM. Lecture Notes in Computer Science, 796:157–162, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanBeek:1994:PMB

[876] G. van Beek, J. P. Geschiere, and A. R. Sukul. Parallelization of a 3D multi-block Navier–Stokes flow solver on a distributed memory MIMD machine. Lecture Notes in Computer Science, 796:163–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

dAst:1994:PTD

[877] I. d’Ast. Porting of a three-dimensional Navier–Stokes solver to a range of parallel computers and a network of workstations. Lecture Notes in Computer Science, 796:171–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boryczko:1994:LGA

[878] K. Boryczko, M. Bubak, J. Kitowski, J. Moscinski, and R. Slota. Lattice gas automata and molecular dynamics on a network of computers. Lecture Notes in Computer Science, 796:177–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dietrich:1994:PSO

[879] S. Dietrich and I. Boyd. Parallel scalar optimized direct simulation Monte Carlo method for rarefied flows on the SP-1. Lecture Notes in Computer Science, 796:181–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andre:1994:MQM

[880] J. M. Andre, B. Champagne, J. Delhalle, J. G. Fripiat, and D. H. Mosley. Modern quantum mechanical techniques and computations on the electronic structure of polymers. Lecture Notes in Computer Science, 796:183– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bywater:1994:OOS

[881] R. Bywater, W. Joosen, S. Bijnens, P. Verbaeten, T. Larsen, and J. Perram. Object-oriented simulation software for drug design. Lecture Notes in Computer Science, 796:189–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trelles-Salazar:1994:MSS

[882] O. Trelles-Salazar, E. L. Zapata, and J.-M. Carazo. Mapping strategies of sequential sequence comparison algorithms on LAN-Based message passing architectures. Lecture Notes in Computer Science, 796:197–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lischka:1994:PCQ

[883] H. Lischka, H. Dachsel, R. Shepard, and R. J. Harrison. Parallel computing in quantum chemistry — message passing and beyond for a general ab initio program system. Lecture Notes in Computer Science, 796:203–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bleckmann:1994:SDQ

[884] P. Bleckmann and F. H. Walter. Solving dynamic and quantum chemical problems with the help of concurrent processors. Lecture Notes in Computer Science, 796:210–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diercksen:1994:ISO

[885] G. H. F. Diercksen. Intelligent software: The OpenMol program. Lecture Notes in Computer Science, 796:219– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OToole:1994:PMD

[886] E. O’Toole, M. Surridge, and C. Upstill. Parallel molecular dynamics simulation of commercial surfactants. Lecture Notes in Computer Science, 796: 223–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulte:1994:CSM

[887] J. Schulte and G. Seifert. Computer simulations of molecular collision processes. Lecture Notes in Computer Science, 796:226–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braun:1994:PPG

[888] H. Braun, M. Assfalg, K. Weymann, and T. Harvey. Parallel processing for generating large molecular databases. Lecture Notes in Computer Science, 796:232–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clark:1994:APC

[889] D. A. Clark, Ch. Rawlings, and J. Shirazi. Applications of parallel constraint logic programming to molecular biology and genetics. Lecture Notes in Computer Science, 796:236–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tangen:1994:SBE

[890] U. Tangen and H. Weberpals. Simulating biological evolution on a parallel computer. Lecture Notes in Computer Science, 796:238–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kremer:1994:SPR

[891] K. Kremer. Supercomputing in polymere research. Lecture Notes in Computer Science, 796:244–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keunings:1994:PFE

[892] R. Keunings, R. Aggarwal, P. Henriksen, D. Vanderstraeten, and O. Zone. Parallel finite element algorithms applied to polymer flow. Lecture Notes in Computer Science, 796:254–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielsen:1994:PMD

[893] O. H. Nielsen. Performance of a molecular-dynamics algorithm on Connection Machines CM-200 and CM-5. Lecture Notes in Computer Science, 796:261–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robler:1994:PMD

[894] U. K. Robler and H. Teichler. Parallel molecular dynamics simulations of liquid and amorphous metals. Lecture Notes in Computer Science, 796:267– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hohl:1994:CMS

[895] D. Hohl. Computational materials science from first principles. Lecture Notes in Computer Science, 796:274– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerndt:1994:APC

[896] M. Gerndt. Automatic parallelization of a crystal growth simulation program for distributed-memory systems. Lecture Notes in Computer Science, 796:281–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nijmeijer:1994:PMD

[897] M. J. P. Nijmeijer. A parallel molecular-dynamics simulation of crystal growth at a cluster of workstations. Lecture Notes in Computer Science, 796:287–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voogd:1994:SAB

[898] J. M. Voogd, P. M. A. Sloot, and R. v. Dantzig. Simulated annealing for N-body systems. Lecture Notes in Computer Science, 796:293–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barros:1994:MPV

[899] S. R. M. Barros, D. Dent, L. Isaksen, G. Robinson, and F. G. Wollenweber. The message passing version of ECMWF’s weather forecast model. Lecture Notes in Computer Science, 796:299–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gartel:1994:PCW

[900] U. Gartel, W. Joppich, and A. Schuller. Parallel computing for weather prediction. Lecture Notes in Computer Science, 796:305–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerlach:1994:PMM

[901] J. Gerlach and B. Kahler. Parallelization of a mesoscale model of atmospheric circulation — an objectoriented approach. Lecture Notes in Computer Science, 796:311–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tremolet:1994:PSA

[902] Y. Tremolet, F. X. Le Dimet, and D. Trystram. Parallelization of scientific applications: Data assimilation in meteorology. Lecture Notes in Computer Science, 796:317–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kersken:1994:PLS

[903] H.-P. Kersken, B. Fritzsch, O. Schenk, W. Hiller, J. Behrens, and E. Kraube. Parallelization of large scale ocean models by data decomposition. Lecture Notes in Computer Science, 796: 323–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Song:1994:STF

[904] Z. W. Song, K. P. P. Pathirana, D. Roose, and J. Berlamont. Simulation of tidal flows for Southern North Sea and sediment transports for Western Scheldt Estuary on parallel computers. Lecture Notes in Computer Science, 796:329–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kervevan:1994:SPA

[905] Ch. Kervevan, R. Fabriol, A. Guillen, and J.-P. Sauty. A sequential-parallel approach for coupling chemistry and transport in groundwater quality modelling. Lecture Notes in Computer Science, 796:335–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Addison:1994:SOR

[906] C. Addison, T. Christensen, J. Larsen, T. Oliver, and A. Sunderland. Supporting an oil reservoir simulator in a distributed memory environment. Lecture Notes in Computer Science, 796: 340–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yau:1994:PAP

[907] H. W. Yau, K. A. Cliffe, J. E. Sinclair, and P. J. Sumner. The parallelization of the AEA probabilistic safety assessment program, MAS-COT. Lecture Notes in Computer Science, 796:346–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Altevogt:1994:MCS

[908] P. Altevogt and F. Gutbrod. Monte Carlo simulations of lattice gauge theories on multiprocessor systems. Lecture Notes in Computer Science, 796: 352–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steinmetz:1994:FGC

[909] M. Steinmetz. The formation of galaxies: A challenge for supercomputers — A simple task for GRAPE? Lecture Notes in Computer Science, 796:358– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poedts:1994:PMC

[910] S. Poedts, P. M. Meijer, J. P. Goedbloed, H. A. van der Vorst, and A. Jakoby. Parallel magnetohydrodynamics on the CM-5. Lecture Notes in Computer Science, 796:365–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Apostolakis:1994:FRP

[911] J. Apostolakis, C. E. Bruschini, P. Calafiura, F. Gagliardi, M. Metcalf, A. Norton, B. Panzer-Steindel, L. M. Bertolotto, and K. J. Peach. First results from the parallelization of CERN’s NA48 simulation program. Lecture Notes in Computer Science, 796:371–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cirlin:1994:SPR

[912] G. E. Cirlin, G. M. Guryanov, N. P. Korneeva, and Yu. B. Samsonenko. Software package for reflection high energy electron diffraction beam intensity measurement and analysis system. Lecture Notes in Computer Science, 796:377–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petkov:1994:CCI

[913] N. Petkov. Conceiving computationally intensive approaches to vision. Lecture Notes in Computer Science, 796:380– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hagmann:1994:TBV

[914] S. Hagmann, H. Kihl, D. Kuhn, and J.-P. Urban. A Transputer-Based visually-guided robot system using neuro-control. Lecture Notes in Computer Science, 796:390–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vuurpijl:1994:SPP

[915] L. Vuurpijl, T. Schouten, and J. Vytopil. A scalable performance prediction method for parallel neural network simulations. Lecture Notes in Computer Science, 796:396–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vuurpijl:1994:PPR

[916] L. Vuurpijl, T. Schouten, and J. Vytopil. PREENS, a parallel research execution environment for neural systems. Lecture Notes in Computer Science, 796:402–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castelow:1994:PIC

[917] D. A. Castelow and N. D. Gent. Parallel implementation of control tasks for an automated stereo camera mount. Lecture Notes in Computer Science, 796:408–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanzalek:1994:SNN

[918] Z. Hanzalek. Simulating neural networks on telmat T-node. Lecture Notes in Computer Science, 796:416– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lankhorst:1994:GAG

[919] M. M. Lankhorst. A genetic algorithm for grammatical inference. Lecture Notes in Computer Science, 796: 418–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lukowicz:1994:MPI

[920] P. Lukowicz, J. Schiffers, and R. Cober. A massively parallel implementation of the full search vector quantization algorithm. Lecture Notes in Computer Science, 796:420–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cramb:1994:CDA

[921] I. Cramb and C. Upstill. Commercial database applications on parallel systems. Lecture Notes in Computer Science, 796:422–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kerridge:1994:EHP

[922] J. Kerridge, I. Jelly, and Ch. Bates. Evaluation of high-performance parallel database machines. Lecture Notes in Computer Science, 796:424–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kroll:1994:FTR

[923] F. R. Kroll. A free-text retrieval index-server on a massively parallel machine. Lecture Notes in Computer Science, 796:430–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walter:1994:LSD

[924] D. Walter and J. Kerridge. Large scale data management and massively parallel architectures in automatic fingerprint recognition. Lecture Notes in Computer Science, 796:435–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fruscione:1994:PPP

[925] M. Fruscione, L. Marenzi, S. Punzi, and P. Stoffella. PARCS: Parallel picture archiving and communication system. Lecture Notes in Computer Science, 796:441–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1994:PPS

[926] Q. Yang, C. Yu, and C. Liu. The performance of a parallel scheme using the divide-and-conquer method. Lecture Notes in Computer Science, 796: 446–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoffmann:1994:HPC

[927] G.-R. Hoffmann. High-performance computing and networking for numerical weather prediction. Lecture Notes in Computer Science, 797:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmidt:1994:HPC

[928] F. Schmidt. High-performance computing and networking as base element of the Environmental Information System UIS of Baden-Wurttemberg. Lecture Notes in Computer Science, 797:5–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ullmann:1994:HSN

[929] K. Ullmann. High speed networking for the research community in Europe. Lecture Notes in Computer Science, 797:13–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McKeown:1994:NII

[930] M. McKeown. The National Information Infrastructure Testbed (NIIT). Lecture Notes in Computer Science, 797:18–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borst:1994:PKP

[931] P. M. Borst, H.-T. Goetz, P. S. Sapaty, and W. Zorn. Parallel knowledge processing in open networks. Lecture Notes in Computer Science, 797:24–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klauser:1994:DFC

[932] A. Klauser and R. Posch. Distributed file caching in parallel architectures utilizing high speed networks. Lecture Notes in Computer Science, 797:30–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Westphal:1994:AFD

[933] H. Westphal and D. Popovic. Application fundamentals of distributed, high-performance computing and networking systems. Lecture Notes in Computer Science, 797:36–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cap:1994:MPW

[934] C. H. Cap. Massive parallelism with workstation clusters — challange or nonsense? Lecture Notes in Computer Science, 797:42–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peltier:1994:NPC

[935] J. P. Peltier and P. Christ. Networking perspectives for collaborative supercomputing in European aerospace research and industry. Lecture Notes in Computer Science, 797:53–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rill:1994:FAW

[936] S. Rill and R. Grosso. Future aerospace working scenarios using high speed networks and supercomputers applied to flow simulation for complete aircraft. Lecture Notes in Computer Science, 797:60–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lang:1994:SEC

[937] U. Lang. A software environment for cooperative simulation and visualization in the aerospace field. Lecture Notes in Computer Science, 797:70–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Causse:1994:IVT

[938] S. Causse, F. Juaneda, and M. Grave. Implementation of visualization tools in a distributed environment. Lecture Notes in Computer Science, 797:77–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haas:1994:NIP

[939] P. Haas and P. Christ. Networking issues in PAGEIN: The “N” of “HPCN”. Lecture Notes in Computer Science, 797:86–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roux:1994:WIL

[940] F. X. Roux, J. Ryan, B. Troff, T. H. Le, and K. Dang-Tran. Wavelet interactive in-line analysis with pegase. Lecture Notes in Computer Science, 797:94–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scheidegger:1994:EHC

[941] A. Scheidegger. European HPC competitiveness centers for industry. Lecture Notes in Computer Science, 797: 103–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iffert:1994:OHP

[942] R. Iffert and U. Harms. An open highperformance computing infrastructure in industry. Lecture Notes in Computer Science, 797:109–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonniger:1994:CCC

[943] T. Bonniger, R. Esser, and D. Krekel. A catalog of classifying characteristics for massively parallel computers. Lecture Notes in Computer Science, 797: 114–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koski:1994:BPC

[944] K. Koski. Building a parallel computing environment from heterogeneous resources. Lecture Notes in Computer Science, 797:120–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramme:1994:DCC

[945] F. Ramme, T. Romke, and K. Kremer. A distributed computing center software for the efficient use of parallel computer systems. Lecture Notes in Computer Science, 797:129–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fox:1994:ISO

[946] G. C. Fox, E. Bogucz, D. A. Jones, K. Mills, M. Podgorny, and K. A. Hawick. InfoMall: A scalable organisation for the development of highperformance computing and communications — software and systems. Lecture Notes in Computer Science, 797: 137–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uszok:1994:CMS

[947] A. Uszok, A. Krol, and K. Zielinski. Cluster management software for open object-oriented systems. Lecture Notes in Computer Science, 797:140– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boryczko:1994:TRP

[948] K. Boryczko, M. Bubak, M. Gajecki, J. Kitowski, J. Moscinski, and M. Pogoda. Transmission rates and performance of a network of computers. Lecture Notes in Computer Science, 797:142–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Numrich:1994:MCR

[949] R. W. Numrich, P. L. Springer, and J. C. Peterson. Measurement of communication rates on the Cray T3D interprocessor network. Lecture Notes in Computer Science, 797:150–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baker:1994:EMC

[950] M. Baker and R. Cloke. A evaluation of the Meiko CS-2 using the GENE-SIS benchmark suite. Lecture Notes in Computer Science, 797:158–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agarwal:1994:VHP

[951] R. C. Agarwal, F. G. Gustavson, and M. Zubair. A very high-performance algorithm for NAS EP benchmark. Lecture Notes in Computer Science, 797: 164–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lazou:1994:EPO

[952] C. Lazou and G. R. Nudd. Evaluation of parallel operating systems. Lecture Notes in Computer Science, 797:170– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hotovy:1994:JES

[953] S. Hotovy and T. Sterling. JNNIE: Evaluation of scalable parallel processing architectures. Lecture Notes in Computer Science, 797:172–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanderVorst:1994:RDH

[954] H. A. van der Vorst. Recent developments in hybrid CG methods. Lecture Notes in Computer Science, 797:174– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crone:1994:PCG

[955] L. G. C. Crone. The preconditioned conjugate gradient method on distributed memory systems. Lecture Notes in Computer Science, 797:184– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deSturler:1994:CCR

[956] E. de Sturler and H. A. van der Vorst. Communication cost reduction for Krylov methods on parallel computers. Lecture Notes in Computer Science, 797:190–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Booten:1994:PAM

[957] J. G. L. Booten, P. M. Meijer, H. J. J. te Riele, and H. A. van der Vorst. Parallel Arnoldi method for the construction of a Krylov subspace basis: an application in magnetohydrodynamics. Lecture Notes in Computer Science, 797:196–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1994:SPE

[958] G. Cheng, G. C. Fox, and K. A. Hawick. A scalable paradigm for effectively-dense matrix formulated applications. Lecture Notes in Computer Science, 797:202–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeLuca:1994:ADD

[959] L. De Luca and P. Fiorino. Automatic differentiation on distributed memory MIMD systems. Lecture Notes in Computer Science, 797:211–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Czezowski:1994:OMH

[960] A. Czezowski and P. Strazdins. Optimisations for the memory hierarchy of a singular value decomposition algorithm implemented on the MIMD architecture. Lecture Notes in Computer Science, 797:215–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwardmann:1994:PDL

[961] Ulrich Schwardmann. Parallelization and data locality of a multigrid solver on the KSR1. Lecture Notes in Computer Science, 797:217–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kiper:1994:PER

[962] A. Kiper. On the parallel evaluations of recurrences. Lecture Notes in Computer Science, 797:219–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geuder:1994:IPE

[963] U. Geuder, M. Hardtner, B. Worner, and R. Zink. Implementation of a parallel Euler solver with GRIDS. Lecture Notes in Computer Science, 797: 221–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deKeyser:1994:DMS

[964] J. de Keyser and D. Roose. Distributed mapping of SPMD programs with a generalized Kernighan-Lin heuristic. Lecture Notes in Computer Science, 797:227–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johan:1994:MDC

[965] Z. Johan, K. K. Mathur, S. L. Johnsson, and T. J. R. Hughes. Mesh decomposition and communication procedures for finite element applications on the Connection Machine CM-5 system. Lecture Notes in Computer Science, 797:233–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cross:1994:CSP

[966] M. Cross, C. S. Ierotheou, S. P. Johnson, and P. F. Leggett. CAP-Tools — semiautomatic parallelization of mesh-based computational mechanics codes. Lecture Notes in Computer Science, 797:241–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hempel:1994:MSM

[967] R. Hempel. The MPI standard for message passing. Lecture Notes in Computer Science, 797:247–252, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sloot:1994:CIO

[968] P. M. A. Sloot, A. G. Hoekstra, and L. O. Hertzberger. A comparison of the Iserver-Occam, Parix, Express, and PVM programming environments on a Parsytec GCel. Lecture Notes in Computer Science, 797:253–259, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Resch:1994:MPS

[969] M. Resch, A. Geiger, and J. Zikeli. Message-passing-systems on workstation clusters and parallel computers — the impact of software- and networkarchitectures on applications. Lecture Notes in Computer Science, 797:260– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saarinen:1994:EES

[970] S. Saarinen. EASYPVM — an EnhAnced subroutine LibrarY for PVM. Lecture Notes in Computer Science, 797:267–272, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dikken:1994:DDL

[971] L. Dikken, F. van der Linden, J. Vesseur, and P. M. A. Sloot. DynamicPVM — dynamic load balancing on parallel systems. Lecture Notes in Computer Science, 797:273–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strumpen:1994:LSM

[972] V. Strumpen. A large-scale metacomputer approach for distributed parallel computing. Lecture Notes in Computer Science, 797:278–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stellner:1994:DAM

[973] G. Stellner, A. Bode, S. Lamberts, and Th. Ludwig. Developing applications for multicomputer systems on workstation clusters. Lecture Notes in Computer Science, 797:286–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maffeis:1994:SSD

[974] S. Maffeis. System support for distributed computing. Lecture Notes in Computer Science, 797:293–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malyshkin:1994:AED

[975] V. E. Malyshkin. Assembly environment for development of application parallel program. Lecture Notes in Computer Science, 797:302–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hackenberg:1994:MIS

[976] R. G. Hackenberg. MaX — investigating shared virtual memory. Lecture Notes in Computer Science, 797:308– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfenning:1994:EMN

[977] J.-Th. Pfenning. Experiences with the Mether-NFS virtual shared memory system. Lecture Notes in Computer Science, 797:316–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blanco:1994:APP

[978] V. Blanco, F. F. Rivera, and O. Plata. Adapting PICL and ParaGraph tools to Transputer-Based multicomputers. Lecture Notes in Computer Science, 797:324–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arbenz:1994:STL

[979] P. Arbenz, H. P. Luthi, Ch. Sprenger, and S. Vogel. SCIDDLE: A tool for large scale cooperative computing. Lecture Notes in Computer Science, 797: 326–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schafers:1994:TGP

[980] L. Schafers, C. Scheidler, and O. Kramer-Fuhrmann. TRAPPER: A graphical programming environment for parallel systems. Lecture Notes in Computer Science, 797:328–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tritscher:1994:LLP

[981] S. Tritscher, R. Zajcew, and M. Barnett. Load leveling on the paragon multicomputer. Lecture Notes in Computer Science, 797:330–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fonlupt:1994:DPV

[982] C. Fonlupt, P. Marquet, and J.-L. Dekeyser. A data-parallel view of the load balancing — experimental results on MasPar MP-1. Lecture Notes in Computer Science, 797:338–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Philippe:1994:DSM

[983] L. Philippe and G.-R. Perrin. Distributed scheduling for multicomputers. Lecture Notes in Computer Science, 797:344–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hansen:1994:TOP

[984] O. Hansen. A tool for optimizing programs on massively parallel computer architectures. Lecture Notes in Computer Science, 797:350–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwarz:1994:CSP

[985] J. Schwarz and F. A. Rabhi. A case study in the parallel implementation of functional languages using skeletons. Lecture Notes in Computer Science, 797:357–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feng:1994:DTA

[986] T. Feng and Y. Kim. A dynamic task allocation scheme in Benes-type nonblocking networks. Lecture Notes in Computer Science, 797:363–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinti:1994:MDS

[987] A. Pinti and J.-C. Grossetie. A methodology for dimensioning SPMD distributed systems applied to industrial applications. Lecture Notes in Computer Science, 797:365–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naim:1994:DLS

[988] O. Naim and A. J. G. Hey. Doloop-surface: An abstract performance data visualization. Lecture Notes in Computer Science, 797:367–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DePietro:1994:GDM

[989] G. De Pietro, A. Esposito, and G. Reitano. A graphic distributed monitoring system for Unix-Based architectures. Lecture Notes in Computer Science, 797:373–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pouzet:1994:PAD

[990] P. Pouzet, J. Paris, and V. Jorrand. Parallel application design: The simulation approach with HASTE. Lecture Notes in Computer Science, 797: 379–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bemmerl:1994:LDD

[991] T. Bemmerl and R. Wismuller. Online distributed debugging on scaleable multicomputer architectures. Lecture Notes in Computer Science, 797:394– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vounckx:1994:FPD

[992] J. Vounckx, G. Deconinck, R. Lauwereins, G. Viehover, R. Wagner, H. Madeira, J. G. Silva, F. Balbach, J. Altmann, and B. Bieker. The FTMPS-Project: Design and implementation of fault-tolerance techniques for massively parallel systems. Lecture Notes in Computer Science, 797:401– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chapman:1994:HPF

[993] B. Chapman, H. Zima, and P. Mehrotra. High-Performance Fortran languages: Advanced applications and their implementation. Lecture Notes in Computer Science, 797:407–416, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandes:1994:EHP

[994] T. Brandes. Evaluation of High-Performance Fortran on some real applications. Lecture Notes in Computer Science, 797:417–422, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coelho:1994:EHC

[995] F. Coelho. Experiments with HPF compilation for a network of workstations. Lecture Notes in Computer Science, 797:423–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanWaveren:1994:HPF

[996] G. M. van Waveren. High-Performance Fortran. Lecture Notes in Computer Science, 797:429–433, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanxleden:1994:VBD

[997] R. v. Hanxleden, K. Kennedy, and J. Saltz. Value-based distributions in Fortran D. Lecture Notes in Computer Science, 797:434–440, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanson:1994:BLO

[998] R. Hanson and T. Leite. Basic linear operators in Fortran 90. Lecture Notes in Computer Science, 797:441– 444, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sodan:1994:PPS

[999] A. Sodan and H. Bi. parLisp — parallel symbolic processing with Lisp on the distributed memory machine MANNA. Lecture Notes in Computer Science, 797:445–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dion:1994:PCW

[1000] M. Dion, Y. Robert, and J.-L. Philippe.

Parallelizing compilers: What can be achieved? Lecture Notes in Computer Science, 797:447–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clint:1994:FDP

[1001] M. Clint, S. Fitzpatrick, T. J. Harmer, P. L. Kilpatrick, and J. M. Boyle. A family of data-parallel derivations. Lecture Notes in Computer Science, 797: 457–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crooks:1994:ADD

[1002] P. Crooks and R. H. Perrott. Automatic data distribution. Lecture Notes in Computer Science, 797:463– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perrin:1994:SSA

[1003] G.-R. Perrin. Some synthesis aspects for data parallel programming. Lecture Notes in Computer Science, 797: 469–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuster:1994:IFF

[1004] U. Kuster and M. Zurn. Influence of Fortran 90 features on performance on Cray vector computer systems. Lecture Notes in Computer Science, 797:475– 483, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dzwinel:1994:MPW

[1005] W. Dzwinel. Mapping of patterns — the way of insight into multidimensional data structures. Lecture Notes in Computer Science, 797:484– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schimmler:1994:ISA

[1006] M. Schimmler, H.-W. Lang, and R. Maab. The instruction systolic array — implementation of a low-cost parallel architecture as add-on board for personal computers. Lecture Notes in Computer Science, 797:487–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alherbish:1994:BBF

[1007] J. S. Alherbish. The BBN Butterfly family: An overview. Lecture Notes in Computer Science, 797:489–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mamatov:1994:APR

[1008] Y. A. Mamatov and V. P. Yemelin. Architectural principles of the RPM. Lecture Notes in Computer Science, 797: 491–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saini:1994:API

[1009] S. Saini and H. D. Simon. Applications performance on Intel Paragon XP/S-15. Lecture Notes in Computer Science, 797:493–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Casari:1994:SRE

[1010] C. Casari, C. Catelli, M. Guanziroli, M. Mazzeo, G. Meola, S. Punzi, A. L. Scheinine, and P. Stofella. Status report on ESPRIT project P7519 palace: Parallelization of GEANT. Lecture Notes in Computer Science, 797:499– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amato:1994:PEP

[1011] M. Amato, A. Matrone, and P. Schiano. A practical experience in parallelizing a large CFD code: The EN-SOLV flow solver. Lecture Notes in Computer Science, 797:508–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abreu:1994:DAI

[1012] S. Abreu and L. M. Pereira. Design for AKL with intelligent pruning. Lecture Notes in Computer Science, 798:3–10, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alferes:1994:CWA

[1013] J. J. Alferes and L. M. Pereira. Contradiction: When avoidance equals removal, part I. Lecture Notes in Computer Science, 798:11–23, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amrhein:1994:BHT

[1014] B. Amrhein. Birkhoff’s HSP-Theorem for cumulative logic programs. Lecture Notes in Computer Science, 798:24–36, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boley:1994:FDE

[1015] H. Boley. Finite domains and exclusions as first-class citizens. Lecture Notes in Computer Science, 798:37–61, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciampolini:1994:SER

[1016] A. Ciampolini, E. Lamma, P. Mello, and C. Stefanelli. Some experimental results in implementing a logic language with multi-head clauses on a transputer architecture. Lecture Notes in Computer Science, 798:62–76, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dovier:1994:LFS

[1017] A. Dovier. A language with finite sets embedded in the CLP-Scheme. Lecture Notes in Computer Science, 798:77–93, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eriksson:1994:FPI

[1018] L.-H. Eriksson. Finitary partial inductive definitions as a general logic. Lecture Notes in Computer Science, 798: 94–119, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Falkman:1994:PMG

[1019] G. Falkman and O. Torgersson. Programming methodologies in GCLA. Lecture Notes in Computer Science, 798:120–151, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hinkelmann:1994:CCE

[1020] K. Hinkelmann and H. Hintze. Computing cost estimates for proof strategies. Lecture Notes in Computer Science, 798:152–170, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hodas:1994:LPM

[1021] J. S. Hodas. Logic programming with multiple context management schemes. Lecture Notes in Computer Science, 798:171–182, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keronen:1994:NPL

[1022] S. Keronen. Non-procedural logic programming. Lecture Notes in Computer Science, 798:183–195, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kreuger:1994:ADC

[1023] P. Kreuger. Axioms in definitional calculi. Lecture Notes in Computer Science, 798:196–205, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miller:1994:PML

[1024] D. Miller. A proposal for modules in Lambda-Prolog. Lecture Notes in Computer Science, 798:206–221, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Momigliano:1994:RSS

[1025] A. Momigliano and M. Ornaghi. Regular search spaces as a foundation of logic programming. Lecture Notes in Computer Science, 798:222–254, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pearce:1994:ASN

[1026] D. Pearce. Answer sets and nonmonotonic S4. Lecture Notes in Computer Science, 798:255–267, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pereira:1994:CWA

[1027] L. M. Pereira and J. J. Alferes. Contradiction: When avoidance equals removal, part II. Lecture Notes in Computer Science, 798:268–281, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinto:1994:CFL

[1028] L. Pinto. Cut formulae and logic programming. Lecture Notes in Computer Science, 798:282–300, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reddy:1994:HOA

[1029] U. S. Reddy. Higher-order aspects of logic programming. Lecture Notes in Computer Science, 798:301–321, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ribeiro:1994:RIC

[1030] C. Ribeiro and A. Porto. Representation and inference with consistent temporal propositions. Lecture Notes in Computer Science, 798:322–332, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schroeder-Heister:1994:DRC

[1031] P. Schroeder-Heister. Definitional reflection and the completion. Lecture Notes in Computer Science, 798:333– 347, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zlatuska:1994:LLS

[1032] J. Zlatuska. Linear logic semantics for concurrent prolog. Lecture Notes in Computer Science, 798:348–360, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Singh:1994:MST

[1033] M. P. Singh. Multiagent systems: A theoretical framework for intentions, know-how, and communications. Lecture Notes in Computer Science, 799: xxiii + 168, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malladi:1994:EFT

[1034] R. Malladi, J. A. Sethian, and B. C. Vemuri. Evolutionary fronts for topologyindependent shape modeling and recovery. Lecture Notes in Computer Science, 800:3–13, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dickinson:1994:AOR

[1035] S. J. Dickinson, H. I. Christensen, J. K. Tsotsos, and J. G. Olofsson. Active object recognition integrating attention and viewpoint control. Lecture Notes in Computer Science, 800:3–14, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giblin:1994:EFS

[1036] P. J. Giblin and R. S. Weiss. Epipolar fields on surfaces. Lecture Notes in Computer Science, 800:14–23, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vinther:1994:AOR

[1037] S. Vinther and R. Cipolla. Active 3D object recognition using 3D affine invariants. Lecture Notes in Computer Science, 800:15–24, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weinshall:1994:SLV

[1038] D. Weinshall, M. Werman, and N. Tishby. Stability and likelihood of views of three dimensional objects. Lecture Notes in Computer Science, 800: 24–35, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taylor:1994:GAC

[1039] M. J. Taylor and A. Blake. Grasping the apparent contour. Lecture Notes in Computer Science, 800:25–34, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rehg:1994:VTH

[1040] J. M. Rehg and T. Kanade. Visual tracking of high DOF articulated structures: An application to human hand tracking. Lecture Notes in Computer Science, 800:35–46, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Williams:1994:TRS

[1041] L. R. Williams. Topological reconstruction of a smooth manifold-solid from its occluding contour. Lecture Notes in Computer Science, 800:36–50, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crowley:1994:ICR

[1042] J. L. Crowley, J. M. Bedrune, M. Bekker, and M. Schneider. Integration and control of reactive visual processes. Lecture Notes in Computer Science, 800:47–60, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Otte:1994:OFE

[1043] M. Otte and H.-H. Nagel. Optical flow estimation: advances and comparisons. Lecture Notes in Computer Science, 800:51–60, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tistarelli:1994:MCO

[1044] M. Tistarelli. Multiple constraints for optical flow. Lecture Notes in Computer Science, 800:61–70, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soatto:1994:MEE

[1045] S. Soatto, R. Frezza, and P. Perona. Motion estimation on the essential manifold. Lecture Notes in Computer Science, 800:61–72, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papadopoulo:1994:MFC

[1046] T. Papadopoulo and O. Faugeras. Motion field of curves: applications. Lecture Notes in Computer Science, 800: 71–82, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shapiro:1994:MPM

[1047] L. S. Shapiro, A. Zisserman, and M. Brady. Motion from point matches using affine epipolar geometry. Lecture Notes in Computer Science, 800:73–84, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carlsson:1994:SIS

[1048] S. Carlsson. Sufficient image structure for 3-D motion and shape estimation. Lecture Notes in Computer Science, 800:83–94, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beardsley:1994:NUA

[1049] P. A. Beardsley, A. Zisserman, and D. W. Murray. Navigation using affine structure from motion. Lecture Notes in Computer Science, 800:85–96, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xu:1994:CBS

[1050] C. Xu and S. A. Velastin. A comparison between the standard Hough transform and the Mahalanobis distance Hough transform. Lecture Notes in Computer Science, 800:95–100, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poelman:1994:PFM

[1051] C. J. Poelman and T. Kanade. A paraperspective factorization method for shape and motion recovery. Lecture Notes in Computer Science, 800: 97–110, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michaelis:1994:JCM

[1052] M. Michaelis and G. Sommer. Junction classification by multiple orientation detection. Lecture Notes in Computer Science, 800:101–108, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vasselle:1994:FCC

[1053] B. Vasselle, G. Giraudon, and M. Berthod. Following corners on curves and surfaces in the scale space. Lecture Notes in Computer Science, 800:109–114, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sundareswaran:1994:ACS

[1054] V. Sundareswaran, P. Bouthemy, and F. Chaumette. Active camera selforientation using dynamic image parameters. Lecture Notes in Computer Science, 800:111–116, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kube:1994:SSP

[1055] P. Kube and P. Perona. Scale-space properties of quadratic edge detectors. Lecture Notes in Computer Science, 800:115–124, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maver:1994:POS

[1056] J. Maver, A. Leonardis, and F. Solina. Planning the optimal set of views using the max-min principle. Lecture Notes in Computer Science, 800:117–122, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grosso:1994:PAE

[1057] E. Grosso. On perceptual advantages of eye-head active control. Lecture Notes in Computer Science, 800:123–130, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amini:1994:SFF

[1058] A. A. Amini. A scalar function formulation for optical flow. Lecture Notes in Computer Science, 800:125–131, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malandain:1994:IRD

[1059] G. Malandain, S. Fernandez-Vidal, and J. M. Rocchisani. Improving registration of 3-D medical images using a mechanical based method. Lecture Notes in Computer Science, 800:131–136, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tunley:1994:FOO

[1060] H. Tunley and D. Young. First order optic flow from log-polar sampled images. Lecture Notes in Computer Science, 800:132–137, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christmas:1994:NIC

[1061] B. Christmas, J. Kittler, and M. Petrou.

Non-iterative contextual correspondence matching. Lecture Notes in Computer Science, 800:137–142, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Black:1994:RNL

[1062] M. J. Black. Recursive non-linear estimation of discontinuous flow fields. Lecture Notes in Computer Science, 800:138–145, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kita:1994:RMR

[1063] Y. Kita. A registration method for rigid objects without point matching. Lecture Notes in Computer Science, 800:143–150, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giachetti:1994:UOF

[1064] A. Giachetti, M. Campani, and V. Torre. The use of optical flow for the autonomous navigation. Lecture Notes in Computer Science, 800:146–151, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zabih:1994:NPL

[1065] R. Zabih and J. Woodfill. Nonparametric local transforms for computing visual correspondence. Lecture Notes in Computer Science, 800:151– 158, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Germain:1994:IME

[1066] F. Germain and T. Skordas. An image motion estimation technique based on a combined statistical test and spatiotemporal generalised likelihood ratio approach. Lecture Notes in Computer Science, 800:152–160, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manmatha:1994:MAT

[1067] R. Manmatha. Measuring the affine transform using Gaussian filters. Lecture Notes in Computer Science, 800: 159–164, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sinclair:1994:IMS

[1068] D. Sinclair and B. Boufama. Independent motion segmentation and collision prediction for road vehicles. Lecture Notes in Computer Science, 800:161– 166, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sato:1994:EAT

[1069] J. Sato and R. Cipolla. Extracting the affine transformation from texture moments. Lecture Notes in Computer Science, 800:165–172, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luthon:1994:MBM

[1070] F. Luthon, G. V. Popescu, and A. Caplier. An MRF based motion detection algorithm implemented on analog resistive network. Lecture Notes in Computer Science, 800:167–174, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fitzgibbon:1994:LFD

[1071] A. W. Fitzgibbon and R. B. Fisher.

Lack-of-fit detection using the rundistribution test. Lecture Notes in Computer Science, 800:173–178, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geiger:1994:OAM

[1072] D. Geiger and K. J. Diamantaras. Occlusion ambiguities in motion. Lecture Notes in Computer Science, 800:175– 180, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Intille:1994:DSI

[1073] S. S. Intille and A. F. Bobick.

Disparity-space images and large occlusion stereo. Lecture Notes in Computer Science, 800:179–186, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borri:1994:RTM

[1074] A. Borri, G. Bucci, and P. Nesi. A robust tracking of 3D motion. Lecture Notes in Computer Science, 800:181– 188, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gourdon:1994:RCS

[1075] A. Gourdon and N. Ayache. Registration of a curve on a surface using differential properties. Lecture Notes in Computer Science, 800:187–192, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koller:1994:RMC

[1076] D. Koller, J. Weber, and J. Malik. Robust multiple car tracking with occlusion reasoning. Lecture Notes in Computer Science, 800:189–198, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vaillant:1994:GAA

[1077] R. Vaillant and L. Gueguen. Genetic algorithms applied to binocular stereovision. Lecture Notes in Computer Science, 800:193–200, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boufama:1994:SMA

[1078] B. Boufama, D. Weinshall, and M. Werman. Shape from motion algorithms: a comparative analysis of scaled orthography and perspective. Lecture Notes in Computer Science, 800:199–204, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herlin:1994:SEI

[1079] I. L. Herlin, D. Bereziat, G. Giraudon, and C. Nguyen. Segmentation of echocardiographic images with Markov random fields. Lecture Notes in Computer Science, 800:201–206, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lawn:1994:REE

[1080] J. M. Lawn and R. Cipolla. Robust egomotion estimation from affine motion parallax. Lecture Notes in Computer Science, 800:205–210, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Derras:1994:URS

[1081] M. Derras, C. Debain, M. Berducat, and P. Bonton. Unsupervised regions segmentation: Real time control of an upkeep machine of natural spaces. Lecture Notes in Computer Science, 800:207–212, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sull:1994:IAF

[1082] S. Sull and N. Ahuja. Integrated 3D analysis of flight image sequences. Lecture Notes in Computer Science, 800: 211–216, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Younes:1994:SIR

[1083] L. Younes. Synchronous image restoration. Lecture Notes in Computer Science, 800:213–217, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McLauchlan:1994:RAS

[1084] P. F. McLauchlan, I. D. Reid, and D. W. Murray. Recursive affine structure and motion from image sequences. Lecture Notes in Computer Science, 800:217–224, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weidner:1994:PIP

[1085] U. Weidner. Parameterfree informationpreserving surface restoration. Lecture Notes in Computer Science, 800:218– 226, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shen:1994:SMI

[1086] X. Shen and D. Hogg. Shape models from image sequences. Lecture Notes in Computer Science, 800:225–230, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Langer:1994:SVI

[1087] M. S. Langer and S. W. Zucker. Spatially varying illumination: A computational model of converging and diverging sources. Lecture Notes in Computer Science, 800:227–234, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nastar:1994:VMN

[1088] C. Nastar. Vibration modes for nonrigid motion analysis in 3D images. Lecture Notes in Computer Science, 800:231–238, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohta:1994:RIS

[1089] Y. Ohta and Y. Hayashi. Recovery of illuminant and surface colors from images based on the CIE daylight. Lecture Notes in Computer Science, 800:235–246, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lindenbaum:1994:AVD

[1090] M. Lindenbaum and S. Ben-David. Applying VC-dimension analysis to object recognition. Lecture Notes in Computer Science, 800:239–250, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolff:1994:DSU

[1091] L. B. Wolff and E. Angelopoulou. 3-D stereo using photometric ratios. Lecture Notes in Computer Science, 800:247– 258, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Havaldar:1994:EGR

[1092] P. Havaldar, G. Medioni, and F. Stein.

Extraction of groups for recognition. Lecture Notes in Computer Science, 800:251–261, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dupuis:1994:SSP

[1093] P. Dupuis and J. Oliensis. Shape from shading: provably convergent algorithms and uniqueness results. Lecture Notes in Computer Science, 800:259– 268, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hel-Or:1994:MBP

[1094] Y. Hel-Or and M. Werman. Model based pose estimation of articulated and constrained objects. Lecture Notes in Computer Science, 800:262–273, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oren:1994:SBL

[1095] M. Oren and S. K. Nayar. Seeing beyond Lambert’s law. Lecture Notes in Computer Science, 800:269–280, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ballard:1994:SBO

[1096] D. H. Ballard and R. P. N. Rao. Seeing behind occlusions. Lecture Notes in Computer Science, 800:274–285, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fua:1994:UDM

[1097] P. Fua and Y. G. Leclerc. Using 3-dimensional meshes to combine image-based and geometry-based constraints. Lecture Notes in Computer Science, 800:281–294, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moses:1994:FRP

[1098] Y. Moses, Y. Adini, and S. Ullman. Face recognition: the problem of compensating for changes in illumination direction. Lecture Notes in Computer Science, 800:286–298, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Proesmans:1994:DOF

[1099] M. Proesmans, L. J. VanGool, E. Pauwels, and A. Oosterlinck. Determination of optical flow and its discontinuities using non-linear diffusion. Lecture Notes in Computer Science, 800:295–304, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baumberg:1994:LFM

[1100] A. Baumberg and D. Hogg. Learning flexible models from image sequences. Lecture Notes in Computer Science, 800:299–308, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nagel:1994:MBD

[1101] H.-H. Nagel, G. Socher, H. Kollnig, and M. Otte. Motion boundary detection in image sequences by local stochastic tests. Lecture Notes in Computer Science, 800:305–315, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leonardis:1994:DRS

[1102] A. Leonardis, F. Solina, and A. Macerl. A direct recovery of superquadric models in range images using recoverand-select paradigm. Lecture Notes in Computer Science, 800:309–318, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ayer:1994:SMO

[1103] S. Ayer, P. Schroeter, and J. Biguen.

Segmentation of moving objects by robust motion parameter estimation over multiple frames. Lecture Notes in Computer Science, 800:316–327, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zerroug:1994:SRS

[1104] M. Zerroug and R. Nevatia. Segmentation and recovery of SHGCs from a real intensity image. Lecture Notes in Computer Science, 800:319–330, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Torr:1994:SMC

[1105] P. H. S. Torr and D. W. Murray. Stochastic motion clustering. Lecture Notes in Computer Science, 800:328– 337, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davis:1994:RHG

[1106] J. Davis and M. Shah. Recognizing hand gestures. Lecture Notes in Computer Science, 800:331–340, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kollnig:1994:AMV

[1107] H. Kollnig, H.-H. Nagel, and M. Otte. Association of motion verbs with vehicle movements extracted from dense optical flow fields. Lecture Notes in Computer Science, 800:338–350, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Worrall:1994:PRA

[1108] A. D. Worrall, G. D. Sullivan, and K. D. Baker. Pose refinement of active models using forces in 3D. Lecture Notes in Computer Science, 800:341– 352, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaelviaeinen:1994:CPN

[1109] H. Kaelviaeinen, P. Hirvonen, L. Xu, and E. Oja. Comparisons of probabilistic and non-probabilistic Hough transforms. Lecture Notes in Computer Science, 800:351–360, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malik:1994:RSC

[1110] J. Malik and R. Rosenholtz. Recovering surface curvature and orientation from texture distortion: a least squares algorithm and sensitivity analysis. Lecture Notes in Computer Science, 800:353–364, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1994:MRF

[1111] S. Z. Li. Markov random field models in computer vision. Lecture Notes in Computer Science, 800:361–370, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaarding:1994:DEL

[1112] J. Gaarding and T. Lindeberg. Direct estimation of local surface shape in a fixating binocular vision system. Lecture Notes in Computer Science, 800:365–376, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henricsson:1994:RKP

[1113] O. Henricsson and F. Heitger. The role of key-points in finding contours. Lecture Notes in Computer Science, 800:371–382, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robert:1994:DOC

[1114] L. Robert and M. Hebert. Deriving orientation cues from stereo images. Lecture Notes in Computer Science, 800:377–388, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foerstner:1994:FLL

[1115] W. Foerstner. A framework for low level feature extraction. Lecture Notes in Computer Science, 800:383–396, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lindeberg:1994:SAS

[1116] T. Lindeberg and J. Gaarding. Shapeadapted smoothing in estimation of 3-D depth cues from affine distortions of local 2-D brightness structure. Lecture Notes in Computer Science, 800:389–402, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feldmar:1994:RAR

[1117] J. Feldmar and N. Ayache. Rigid and affine registration of smooth surfaces using differential properties. Lecture Notes in Computer Science, 800:397– 406, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zabrodsky:1994:USR

[1118] H. Zabrodsky and D. Weinshall. Utilizing symmetry in the reconstruction of three-dimensional shape from noisy images. Lecture Notes in Computer Science, 800:403–410, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shashua:1994:QRS

[1119] A. Shashua and S. Toelg. The quadric reference surface: Applications in registering views of complex 3D objects. Lecture Notes in Computer Science, 800:407–416, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heyden:1994:CCL

[1120] A. Heyden. Consistency and correction of line-drawing, obtained by projections of piecewise planar objects. Lecture Notes in Computer Science, 800:411–420, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhao:1994:RRB

[1121] C. S. Zhao and R. Mohr. Relative 3D regularized B-spline surface reconstruction through image sequences. Lecture Notes in Computer Science, 800:417– 426, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petitjean:1994:EGA

[1122] S. Petitjean. On the enumerative geometry of aspect graphs. Lecture Notes in Computer Science, 800:421–426, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fiddelaers:1994:GDC

[1123] P. Fiddelaers, E. J. Pauwels, and L. J.

Van Gool. Geometry-driven curve evolution. Lecture Notes in Computer Science, 800:427–432, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delingette:1994:ISP

[1124] H. Delingette. Intrinsic stabilizers of planar curves. Lecture Notes in Computer Science, 800:427–438, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartley:1994:QMM

[1125] R. J. Hartley, A. Noble, J. Grande, and J. Liu. Quantitative measurement of manufactured diamond shape. Lecture Notes in Computer Science, 800:433– 440, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aastroem:1994:APN

[1126] K. Aastroem. Affine and projective normalization of planar curves and regions. Lecture Notes in Computer Science, 800:439–448, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koh:1994:HSR

[1127] E. Koh, D. Metaxas, and N. Badler. Hierarchical shape representation using locally adaptive finite elements. Lecture Notes in Computer Science, 800:441– 448, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daucher:1994:CCS

[1128] N. Daucher, M. Dhome, and J. T. Lapreste. Camera calibration from spheres images. Lecture Notes in Computer Science, 800:449–454, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sapiro:1994:ALP

[1129] G. Sapiro and A. Tannenbaum. Area and length preserving geometric invariant scale-spaces. Lecture Notes in Computer Science, 800:449–458, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horaud:1994:SCS

[1130] R. Horaud, F. Dornaika, B. Boufama, and R. Mohr. Self calibration of a stereo head mounted onto a robot arm. Lecture Notes in Computer Science, 800:455–462, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quan:1994:IPU

[1131] L. Quan. Invariants of 6 points from 3 uncalibrated images. Lecture Notes in Computer Science, 800:459–470, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ponce:1994:AMU

[1132] J. Ponce, D. H. Marimont, and T. A.

Cass. Analytical methods for uncalibrated stereo and motion reconstruction. Lecture Notes in Computer Science, 800:463–470, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartley:1994:SCM

[1133] R. J. Hartley. Self-calibration from multiple views with a rotating camera. Lecture Notes in Computer Science, 800:471–478, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sparr:1994:CFK

[1134] G. Sparr. A common framework for kinetic depth, reconstruction and motion for deformable objects. Lecture Notes in Computer Science, 800:471– 481, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shashua:1994:TVR

[1135] A. Shashua. Trilinearity in visual recognition by alignment. Lecture Notes in Computer Science, 800:479– 484, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faugeras:1994:WCT

[1136] O. Faugeras and L. Robert. What can two images tell us about a third one? Lecture Notes in Computer Science, 800:485–494, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piccioli:1994:RMR

[1137] G. Piccioli, E. De Micheli, and M. Campani. A robust method for road sign detection and recognition. Lecture Notes in Computer Science, 800:495–500, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tan:1994:PDR

[1138] T. N. Tan, G. D. Sullivan, and K. D. Baker. Pose determination and recognition of vehicles in traffic scenes. Lecture Notes in Computer Science, 800:501– 506, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fisher:1994:PCT

[1139] R. B. Fisher. Performance comparison of ten variations on the interpretationtree matching algorithm. Lecture Notes in Computer Science, 800:507–512, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsuno:1994:RHF

[1140] K. Matsuno, C. W. Lee, and S. Tsuji.

Recognition of human facial expressions without feature extraction. Lecture Notes in Computer Science, 800: 513–520, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Derou:1994:PNN

[1141] D. Derou and L. Herault. Pulsed neural networks and perceptive grouping. Lecture Notes in Computer Science, 800:521–526, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanDiest:1994:PIP

[1142] M. Van Diest, L. J. Van Gool, T. Moons, and E. J. Pauwels. Projective invariants for planar contour recognition. Lecture Notes in Computer Science, 800:527–534, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cooper:1994:DWF

[1143] P. R. Cooper, L. A. Birnbaum, D. Halabe, and M. Brand. Divided we fall: Resolving occlusions using causal reasoning. Lecture Notes in Computer Science, 800:535–542, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1994:CCH

[1144] M. Li. Camera calibration of a headeye system for active vision. Lecture Notes in Computer Science, 800:543– 554, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartley:1994:LPC

[1145] R. J. Hartley and R. Gupta. Linear pushbroom cameras. Lecture Notes in Computer Science, 800:555–566, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deriche:1994:RRE

[1146] R. Deriche, Z. Zhang, Q.-T. Luong, and O. Faugeras. Robust recovery of the epipolar geometry for an uncalibrated stereo rig. Lecture Notes in Computer Science, 800:567–576, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luong:1994:SAF

[1147] Q.-T. Luong and O. Faugeras. A stability analysis of the fundamental matrix. Lecture Notes in Computer Science, 800:577–588, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luong:1994:CRG

[1148] Q.-T. Luong and T. Vieville. Canonic representations for the geometries of multiple projective views. Lecture Notes in Computer Science, 800:589– 599, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aczel:1994:FUP

[1149] P. Aczel. Final universes of processes. Lecture Notes in Computer Science, 802:1–28, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Curien:1994:SS

[1150] P.-L. Curien. On the symmetry of sequentiality. Lecture Notes in Computer Science, 802:29–71, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pitts:1994:CAM

[1151] A. M. Pitts. Computational adequacy via “mixed” inductive definitions. Lecture Notes in Computer Science, 802:72–82, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rutten:1994:SCI

[1152] J. Rutten. A structural co-induction theorem. Lecture Notes in Computer Science, 802:83–102, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanBreugel:1994:TMD

[1153] F. Van Breugel. Three metric domains of processes for bisimulation. Lecture Notes in Computer Science, 802:103– 121, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBakker:1994:TMH

[1154] J. W. De Bakker and F. Van Breugel.

Topological models for higher order control flow. Lecture Notes in Computer Science, 802:122–142, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sangiorgi:1994:IFP

[1155] D. Sangiorgi. An investigation into functions as processes. Lecture Notes in Computer Science, 802:143–159, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsen:1994:TAB

[1156] K. G. Larsen and W. Yi. Time abstracted bisimulation: Implicit specifications and decidability. Lecture Notes in Computer Science, 802:160–176, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneider:1994:TRC

[1157] S. Schneider. Timewise refinement for communicating processes. Lecture Notes in Computer Science, 802:177– 214, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1994:ART

[1158] L. Chen. Axiomatising real-time processes. Lecture Notes in Computer Science, 802:215–229, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scholefield:1994:PSR

[1159] D. Scholefield, H. Zedan, and H. Jifeng. A predicative semantics for the refinement of real-time systems. Lecture Notes in Computer Science, 802:230– 249, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Best:1994:CPS

[1160] E. Best and H.-G. Linde-Goeers. Compositional process semantics of Petri boxes. Lecture Notes in Computer Science, 802:250–270, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinna:1994:SER

[1161] G. M. Pinna and A. Poigne. On the specification of elementary reactive behaviour. Lecture Notes in Computer Science, 802:271–292, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeffrey:1994:CAM

[1162] A. Jeffrey. A chemical abstract machine for graph reduction. Lecture Notes in Computer Science, 802:293–303, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mulry:1994:LTK

[1163] P. S. Mulry. Lifting theorems for Kleisli categories. Lecture Notes in Computer Science, 802:304–319, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brookes:1994:SFI

[1164] S. Brookes and S. Geva. Sequential functions on indexed domains and full abstraction for a sub-language of PCF. Lecture Notes in Computer Science, 802:320–332, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bucciarelli:1994:AAS

[1165] A. Bucciarelli. Another approach to sequentiality: Kleene’s unimonotone functions. Lecture Notes in Computer Science, 802:333–358, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stoughton:1994:MLR

[1166] A. Stoughton. Mechanizing logical relations. Lecture Notes in Computer Science, 802:359–377, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bloom:1994:SQV

[1167] S. L. Bloom and Z. Esik. Some quasivarieties of iteration theories. Lecture Notes in Computer Science, 802:378– 409, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heckmann:1994:PPD

[1168] R. Heckmann. Probabilistic power domains, information systems, and locales. Lecture Notes in Computer Science, 802:410–437, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huth:1994:LDL

[1169] M. Huth. Linear domains and linear maps. Lecture Notes in Computer Science, 802:438–453, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1994:UQP

[1170] G.-Q. Zhang. Universal quasi-prime algebraic domains. Lecture Notes in Computer Science, 802:454–473, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blute:1994:HME

[1171] R. F. Blute, P. Panangaden, and R. A. G. Seely. Holomorphic models of exponential types in linear logic. Lecture Notes in Computer Science, 802:474–512, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wadler:1994:SLL

[1172] P. Wadler. A syntax for linear logic. Lecture Notes in Computer Science, 802:513–529, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rabinovich:1994:CAT

[1173] A. Rabinovich. A complete axiomatisation for trace congruence of finite state behaviors. Lecture Notes in Computer Science, 802:530–543, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kopperman:1994:ATC

[1174] R. D. Kopperman and R. C. Flagg.

The asymmetric topology of computer science. Lecture Notes in Computer Science, 802:544–553, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calbrix:1994:UPW

[1175] H. Calbrix, M. Nivat, and A. Podelski. Ultimately periodic words of rational w-languages. Lecture Notes in Computer Science, 802:554–566, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fiech:1994:CDF

[1176] A. Fiech. Category of delta-functors.

Lecture Notes in Computer Science, 802:567–586, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Banerjee:1994:CIL

[1177] A. Banerjee and D. A. Schmidt. A categorical interpretation of Landin’s correspondence principle. Lecture Notes in Computer Science, 802:587–602, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruce:1994:OST

[1178] K. B. Bruce, J. Crabtree, and G. Kanapathy. An operational semantics for TOOPLE: A statically-typed objectoriented programming language. Lecture Notes in Computer Science, 802: 603–626, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Danvy:1994:TBD

[1179] O. Danvy and J. Hatcliff. On the transformation between direct and continuation semantics. Lecture Notes in Computer Science, 802:627–648, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benveniste:1994:DFS

[1180] A. Benveniste, P. Caspi, P. Le Guernic, and N. Halbwachs. Data-flow synchronous languages. Lecture Notes in Computer Science, 803:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergstra:1994:PAB

[1181] J. A. Bergstra, A. Ponse, and J. J. Van Wamel. Process algebra with backtracking. Lecture Notes in Computer Science, 803:46–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boudol:1994:SCA

[1182] G. Boudol. Some chemical abstract machines. Lecture Notes in Computer Science, 803:92–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clarke:1994:VTF

[1183] E. Clarke, O. Grumberg, and D. Long.

Verification tools for finite-state concurrent systems. Lecture Notes in Computer Science, 803:124–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engberg:1994:LLP

[1184] U. Engberg and G. Winskel. Linear logic on Petri nets. Lecture Notes in Computer Science, 803:176–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jensen:1994:ITA

[1185] K. Jensen. An introduction to the theoretical aspects of coloured Petri nets. Lecture Notes in Computer Science, 803:230–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kesten:1994:TVS

[1186] Y. Kesten, Z. Manna, and A. Pnueli.

Temporal verification of simulation and refinement. Lecture Notes in Computer Science, 803:273–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lamport:1994:VSC

[1187] L. Lamport. Verification and specification of concurrent programs. Lecture Notes in Computer Science, 803: 347–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lynch:1994:STP

[1188] N. Lynch. Simulation techniques for proving properties of real-time systems. Lecture Notes in Computer Science, 803:375–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielsen:1994:RBM

[1189] M. Nielsen, V. Sassone, and G. Winskel. Relationships between models of concurrency. Lecture Notes in Computer Science, 803:425–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parrow:1994:ID

[1190] J. Parrow. Interaction diagrams. Lecture Notes in Computer Science, 803: 477–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parrow:1994:ATN

[1191] J. Parrow and D. Sangiorgi. Algebraic theories for name-passing calculi. Lecture Notes in Computer Science, 803:509–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rutten:1994:IAF

[1192] J. Rutten and D. Turi. Initial algebra and final coalgebra semantics for concurrency. Lecture Notes in Computer Science, 803:530–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:1994:LSI

[1193] W. Thomas and H. Lescow. Logical specifications of infinite computations. Lecture Notes in Computer Science, 803:583–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zwiers:1994:POB

[1194] J. Zwiers and W. Janssen. Partial order based design of concurrent systems. Lecture Notes in Computer Science, 803:622–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hernandez:1994:QRS

[1195] D. Hernandez. Qualitative representation of spatial knowledge. Lecture Notes in Computer Science, 804:ix + 202, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pippenger:1994:JN

[1196] N. Pippenger. Juggling networks. Lecture Notes in Computer Science, 805: 1–12, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dixon:1994:OPV

[1197] B. Dixon and R. E. Tarjan. Optimal parallel verification of minimum spanning trees in logarithmic time. Lecture Notes in Computer Science, 805:13–22, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greenlaw:1994:PCA

[1198] R. Greenlaw and J. Machta. The parallel complexity of algorithms for pattern formation models. Lecture Notes in Computer Science, 805:23–34, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waack:1994:PCI

[1199] S. Waack. On the parallel complexity of iterated multiplication in rings of algebraic integers. Lecture Notes in Computer Science, 805:35–44, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liang:1994:NAR

[1200] Y. D. Liang, R. Greenlaw, and G. Manacher. NC02 algorithms regarding Hamiltonian paths and circuits in interval graphs. Lecture Notes in Computer Science, 805:45–58, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dixon:1994:CPQ

[1201] B. Dixon. Concurrency in an O(log log N) priority queue. Lecture Notes in Computer Science, 805:59–72, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bettayeb:1994:EMO

[1202] S. Bettayeb, Z. Miller, T. Peng, and H. Sudborough. Embedding k-D meshes into optimum hypercubes with dilation 2k − 1. Lecture Notes in Computer Science, 805:73–80, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roettger:1994:EDG

[1203] M. Roettger, U.-P. Schroeder, and W. Unger. Embedding 3-dimensional grids into optimal hypercubes. Lecture Notes in Computer Science, 805:81–94, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dehne:1994:DCR

[1204] F. Dehne and R. D. Lins. Distributed cyclic reference counting. Lecture Notes in Computer Science, 805: 95–100, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldy:1994:ERC

[1205] P. Baldy, H. Dicky, R. Medina, and M. Morvan. Efficient reconstruction of the causal relationship in distributed systems. Lecture Notes in Computer Science, 805:101–114, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dehne:1994:SPC

[1206] F. Dehne. Scalable parallel computational geometry. Lecture Notes in Computer Science, 805:115–120, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krizanc:1994:SSA

[1207] D. Krizanc and L. Narayanan. Sorting and selection on arrays with diagonal connections. Lecture Notes in Computer Science, 805:121–136, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ubeda:1994:WOT

[1208] S. Ubeda. Work-optimal thinning algorithm on SIMD machines. Lecture Notes in Computer Science, 805:137– 152, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cung:1994:EIP

[1209] V.-D. Cung and B. Le Cun. An efficient implementation of parallel A. Lecture Notes in Computer Science, 805:153– 168, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bokka:1994:CTC

[1210] V. Bokka, H. Gurla, S. Olariu, and J. L. Schwing. Constant-time convexity problems on reconfigurable meshes. Lecture Notes in Computer Science, 805:169–180, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galtier:1994:DAB

[1211] J. Galtier. Deepness analysis: Bringing optimal fronts to triangular finite element method. Lecture Notes in Computer Science, 805:181–196, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hily:1994:CBN

[1212] A. Hily and D. Solleau. Communications in bus networks. Lecture Notes in Computer Science, 805:197–206, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diks:1994:FTL

[1213] K. Diks and A. Pelc. Fault-tolerant linear broadcasting. Lecture Notes in Computer Science, 805:207–218, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jansen:1994:MBT

[1214] K. Jansen and H. Mueller. The minimum broadcast time problem. Lecture Notes in Computer Science, 805:219– 234, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hromkovic:1994:CSD

[1215] J. Hromkovic, R. Klasing, W. Unger, and H. Wagener. The complexity of systolic dissemination of information in interconnection networks. Lecture Notes in Computer Science, 805:235– 250, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krumme:1994:RGS

[1216] D. W. Krumme. Representations of gossip schemes. Lecture Notes in Computer Science, 805:251–266, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kulasinghe:1994:MTH

[1217] P. Kulasinghe and S. Bettayeb. On the multiply-twisted hypercube. Lecture Notes in Computer Science, 805:267– 278, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Altenkirch:1994:PSN

[1218] T. Altenkirch. Proving strong normalization of CC by modifying realizability semantics. Lecture Notes in Computer Science, 806:3–18, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanBenthemJutting:1994:CAP

[1219] L. S. Van Benthem Jutting, J. McKinna, and R. Pollack. Checking algorithms for pure type systems. Lecture Notes in Computer Science, 806:19–61, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coquand:1994:IOT

[1220] T. Coquand. Infinite objects in type theory. Lecture Notes in Computer Science, 806:62–78, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geuvers:1994:CBL

[1221] H. Geuvers. Conservativity between logics and typed lamda calculi. Lecture Notes in Computer Science, 806: 79–107, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hayashi:1994:LRT

[1222] S. Hayashi. Logic of refinement types. Lecture Notes in Computer Science, 806:108–126, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Helmink:1994:PCD

[1223] L. Helmink, A. Sellink, and F. Vaandrager. Proof-checking a data link protocol. Lecture Notes in Computer Science, 806:127–165, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofmann:1994:EEM

[1224] M. Hofmann. Elimination of extensionality in Martin-Loef type theory. Lecture Notes in Computer Science, 806:166–190, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leclerc:1994:PSC

[1225] F. Leclerc and C. Paulin-Mohring. Programming with streams in Coq — A case study: The sieve of eratosthenes. Lecture Notes in Computer Science, 806:191–212, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Magnusson:1994:APE

[1226] L. Magnusson and B. Nordstroem. The ALF proof editor and its proof engine. Lecture Notes in Computer Science, 806:213–237, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maharaj:1994:ESS

[1227] S. Maharaj. Encoding Z-style schemas in type theory. Lecture Notes in Computer Science, 806:238–262, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miculan:1994:EPS

[1228] M. Miculan. The expressive power of structural operational semantics with explicit assumptions. Lecture Notes in Computer Science, 806:263–290, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parent:1994:DCP

[1229] C. Parent. Developing certified programs in the system Coq — the program tactic. Lecture Notes in Computer Science, 806:291–312, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pollack:1994:CUA

[1230] R. Pollack. Closure under alphaconversion. Lecture Notes in Computer Science, 806:313–332, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raffalli:1994:MD

[1231] C. Raffalli. Machine deduction. Lecture Notes in Computer Science, 806:333– 351, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ranta:1994:TTI

[1232] A. Ranta. Type theory and the informal language of mathematics. Lecture Notes in Computer Science, 806:352– 365, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolfram:1994:SAC

[1233] D. A. Wolfram. Semantics for abstract clauses. Lecture Notes in Computer Science, 806:366–383, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benson:1994:SEA

[1234] G. Benson. A space efficient algorithm for finding the best non-overlapping alignment score. Lecture Notes in Computer Science, 807:1–14, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodlaender:1994:PCS

[1235] H. Bodlaender, R. G. Downey, M. R. Fellows, and H. T. Wareham. The parameterized complexity of sequence alignment and consensus. Lecture Notes in Computer Science, 807:15–30, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chao:1994:CAS

[1236] K.-M. Chao. Computing all suboptimal alignments in linear space. Lecture Notes in Computer Science, 807:31–42, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bafna:1994:AAM

[1237] V. Bafna, E. L. Lawler, and P. A.

Peuzner. Approximation algorithms for multiple sequence alignment. Lecture Notes in Computer Science, 807:43–53, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1994:CDM

[1238] X. Huang. A context dependent method for comparing sequences. Lecture Notes in Computer Science, 807: 54–63, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cobbs:1994:FIA

[1239] A. L. Cobbs. Fast identification of approximately matching substrings. Lecture Notes in Computer Science, 807: 64–74, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jiang:1994:ATA

[1240] T. Jiang, L. Wang, and K. Zhang.

Alignment of trees — an alternative to tree edit. Lecture Notes in Computer Science, 807:75–86, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1994:PRA

[1241] X. Huang, P. A. Pevzner, and W. Miller. Parametric recomputing in alignment graphs. Lecture Notes in Computer Science, 807:87–101, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luczak:1994:LDC

[1242] T. Luczak and W. Szpankowski. A lossy data compression based on string matching: Preliminary analysis and suboptimal algorithms. Lecture Notes in Computer Science, 807:102–112, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manber:1994:TCS

[1243] U. Manber. A text compression scheme that allows fast searching directly in the compressed file. Lecture Notes in Computer Science, 807:113–124, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karpinski:1994:AIO

[1244] M. Karpinski and W. Rytter. An alphabet-independent optimal parallel search for three dimensional pattern. Lecture Notes in Computer Science, 807:125–135, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lestree:1994:URU

[1245] L. Lestree. Unit route upper bound for string-matching on hypercube. Lecture Notes in Computer Science, 807:136– 145, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

RaoKosaraju:1994:CSS

[1246] S. Rao Kosaraju. Computation of squares in a string. Lecture Notes in Computer Science, 807:146–150, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohri:1994:MST

[1247] M. Mohri. Minimization of sequential transducers. Lecture Notes in Computer Science, 807:151–163, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alexander:1994:SCS

[1248] K. S. Alexander. Shortest common superstrings for strings of random letters. Lecture Notes in Computer Science, 807:164–172, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Irving:1994:MCS

[1249] R. W. Irving and C. B. Fraser. Maximal common subsequences and minimal common supersequences. Lecture Notes in Computer Science, 807:173– 183, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breslauer:1994:DMU

[1250] D. Breslauer. Dictionary-matching on unbounded alphabets: Uniformlength dictionaries. Lecture Notes in Computer Science, 807:184–197, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeza-Yates:1994:PMU

[1251] R. Baeza-Yates, W. Cunto, U. Manber, and S. Wu. Proximity matching using fixed-queries trees. Lecture Notes in Computer Science, 807:198–212, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kilpelainen:1994:QPT

[1252] P. Kilpelainen and H. Mannila. Query primitives for tree-structured data. Lecture Notes in Computer Science, 807:213–225, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Idury:1994:MMP

[1253] R. M. Idury and A. A. Schaeffer. Multiple matching of parameterized patterns. Lecture Notes in Computer Science, 807:226–239, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akutsu:1994:ASM

[1254] T. Akutsu. Approximate string matching with don’t care characters. Lecture Notes in Computer Science, 807:240– 249, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tang:1994:MMN

[1255] S. Tang, K. Zhang, and X. Wu. Matching with matrix norm minimization. Lecture Notes in Computer Science, 807:250–258, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chang:1994:ASM

[1256] W. I. Chang and T. G. Marr. Approximate string matching and local similarity. Lecture Notes in Computer Science, 807:259–273, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ito:1994:PTA

[1257] M. Ito, K. Shimizu, M. Nakanishi, and A. Hashimoto. Polynomial-time algorithms for computing characteristic strings. Lecture Notes in Computer Science, 807:274–288, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakakibara:1994:RMR

[1258] Y. Sakakibara, M. Brown, R. Hughey, and I. S. Mian. Recent methods for RNA modeling using stochastic context-free grammars. Lecture Notes in Computer Science, 807:289–306, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kececioglu:1994:EBO

[1259] J. Kececioglu and D. Sankoff. Efficient bounds for oriented chromosome inversion distance. Lecture Notes in Computer Science, 807:307–325, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaerdenfors:1994:RER

[1260] P. Gaerdenfors. The role of expectations in reasoning. Lecture Notes in Computer Science, 808:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hajek:1994:LAR

[1261] P. Hajek. On logics of approximate reasoning. Lecture Notes in Computer Science, 808:17–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liau:1994:GSC

[1262] Churn Jung Liau and B. I.-P. Lin. Gentzen sequent calculus for possibilistic reasoning. Lecture Notes in Computer Science, 808:31–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flach:1994:MIR

[1263] P. A. Flach. A model of inductive reasoning. Lecture Notes in Computer Science, 808:41–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

CorreadaSilva:1994:ARU

[1264] F. S. Correa da Silva, D. S. Robertson, and J. Hesketh. Automated reasoning with uncertainties. Lecture Notes in Computer Science, 808:57–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lawry:1994:AAS

[1265] J. Lawry and G. M. Wilmers. An axiomatic approach to systems of prior distributions in inexact reasoning. Lecture Notes in Computer Science, 808: 81–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pereira:1994:CRS

[1266] L. M. Pereira, J. J. Alferes, and J. N. Aparicio. Contradiction removal semantics with explicit negation. Lecture Notes in Computer Science, 808:91–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pereira:1994:LPN

[1267] L. M. Pereira, J. N. Aparicio, and J. J.

Alferes. Logic programming for nonmonotonic reasoning. Lecture Notes in Computer Science, 808:107–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shoham:1994:AOP

[1268] Y. Shoham. Agent oriented programming. Lecture Notes in Computer Science, 808:123–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakata:1994:ATL

[1269] K. Nakata. An application of temporal logic for representation and reasoning about design. Lecture Notes in Computer Science, 808:131–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Georgatos:1994:KTP

[1270] K. Georgatos. Knowledge theoretic properties of topological spaces. Lecture Notes in Computer Science, 808: 147–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauszer:1994:RLM

[1271] C. M. Rauszer. Rough logic for multi-agent systems. Lecture Notes in Computer Science, 808:161–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cholvy:1994:LAM

[1272] L. Cholvy. A logical approach to multisources reasoning. Lecture Notes in Computer Science, 808:183–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Devlin:1994:STS

[1273] K. Devlin. Situation theory and social structure. Lecture Notes in Computer Science, 808:197–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Massey:1994:SKB

[1274] J. Massey. SAFER K-64: A byteoriented block-ciphering algorithm. Lecture Notes in Computer Science, 809:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daemen:1994:NAB

[1275] J. Daemen, R. Govaerts, and J. Vandewalle. A new approach to block cipher design. Lecture Notes in Computer Science, 809:18–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaliski:1994:FBC

[1276] B. Kaliski and M. Robshaw. Fast block cipher proposal. Lecture Notes in Computer Science, 809:33–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bloecher:1994:FFS

[1277] U. Bloecher and M. Dichtl. Fish: A fast software stream cipher. Lecture Notes in Computer Science, 809:41–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krawczyk:1994:SGS

[1278] H. Krawczyk. The shrinking generator: Some practical considerations. Lecture Notes in Computer Science, 809:45–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1994:MRM

[1279] R. Anderson. A modern rotor machine. Lecture Notes in Computer Science, 809:47–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chambers:1994:TSC

[1280] B. Chambers. Two stream ciphers. Lecture Notes in Computer Science, 809: 51–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rogaway:1994:SOE

[1281] P. Rogaway and D. Coppersmith.

A software-optimised encryption algorithm. Lecture Notes in Computer Science, 809:56–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lomas:1994:ENT

[1282] M. Lomas. Encrypting network traffic. Lecture Notes in Computer Science, 809:64–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Preneel:1994:DPD

[1283] B. Preneel. Design principles for dedicated hash functions. Lecture Notes in Computer Science, 809:71–82, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roe:1994:PSC

[1284] M. Roe. Performance of symmetric ciphers and one-way hash functions. Lecture Notes in Computer Science, 809:83–89, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golic:1994:SSR

[1285] J. Golic. On the security of shift register based keystream generators. Lecture Notes in Computer Science, 809: 90–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ding:1994:DCD

[1286] C. Ding. The differential cryptanalysis and design of natural stream ciphers. Lecture Notes in Computer Science, 809:101–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biham:1994:MO

[1287] E. Biham. On modes of operation. Lecture Notes in Computer Science, 809: 116–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gollmann:1994:CCC

[1288] D. Gollmann. Cryptanalysis of clock controlled shift registers. Lecture Notes in Computer Science, 809:121– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wheeler:1994:BDE

[1289] D. Wheeler. A bulk data encryption algorithm. Lecture Notes in Computer Science, 809:127–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tao:1994:FAO

[1290] R. Tao. On finite automation onekey cryptosystems. Lecture Notes in Computer Science, 809:135–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schnorr:1994:PFH

[1291] C. Schnorr and S. Vaudenay. Parallel FFT-Hashing. Lecture Notes in Computer Science, 809:149–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lai:1994:ADB

[1292] X. Lai and L. Knudsen. Attacks on double block length hash functions. Lecture Notes in Computer Science, 809:157–165, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1994:QMS

[1293] A. Chan, R. Games, and J. Rushanan.

On quadratic M-sequences. Lecture Notes in Computer Science, 809:166– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klapper:1994:ASR

[1294] A. Klapper and M. Goresky. 2-Adic shift registers. Lecture Notes in Computer Science, 809:174–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nyberg:1994:NBM

[1295] K. Nyberg. New bent mappings suitable for fast implementation. Lecture Notes in Computer Science, 809:179– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MacLaren:1994:CPN

[1296] N. MacLaren. Cryptographic pseudorandom numbers in simulation. Lecture Notes in Computer Science, 809: 185–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneier:1994:DNV

[1297] B. Schneier. Description of a new variable-length key, 64-bit block cipher (Blowfish). Lecture Notes in Computer Science, 809:191–204, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.counterpane.com/bfsverlag.
html.

DiPorto:1994:VBC

[1298] A. Di Porto and W. Wolfowicz. VINO: A block cipher including variable permutations. Lecture Notes in Computer Science, 809:205–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knudsen:1994:PSF

[1299] L. Knudsen. Practically secure Feistel ciphers. Lecture Notes in Computer Science, 809:211–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allgayer:1994:CER

[1300] J. Allgayer and E. Franconi. Collective entities and relations in concept languages. Lecture Notes in Computer Science, 810:13–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baader:1994:CET

[1301] F. Baader and B. Hollunder. Computing extensions of terminological default theories. Lecture Notes in Computer Science, 810:30–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bettini:1994:FIB

[1302] C. Bettini. A formalization of intervalbased temporal subsumption in first order logic. Lecture Notes in Computer Science, 810:53–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boutilier:1994:NSA

[1303] C. Boutilier. Normative, subjunctive and autoepistemic defaults. Lecture Notes in Computer Science, 810:74–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Console:1994:ARA

[1304] L. Console and D. T. Dupre. Abductive reasoning with abstraction axioms. Lecture Notes in Computer Science, 810:98–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Donini:1994:QRD

[1305] F. M. Donini, M. Lenzerini, D. Nardi, and W. Nutt. Queries, rules and definitions as epistemic sentences in concept languages. Lecture Notes in Computer Science, 810:113–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gottlob:1994:PBT

[1306] G. Gottlob. The power of beliefs or translating default logic into standard autoepistemic logic. Lecture Notes in Computer Science, 810:133–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greiner:1994:LOA

[1307] R. Greiner and D. Schuurmans. Learning an optimally accurate representation system. Lecture Notes in Computer Science, 810:145–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kakas:1994:DRN

[1308] A. C. Kakas. Default reasoning via negation as failure. Lecture Notes in Computer Science, 810:160–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalinski:1994:WAR

[1309] J. Kalinski. Weak autoepistemic reasoning and well-founded semantics. Lecture Notes in Computer Science, 810:179–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kautz:1994:FCF

[1310] H. Kautz and B. Selman. Forming concepts for fast inference. Lecture Notes in Computer Science, 810:200– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1994:CST

[1311] Y. Q. Lin. A common-sense theory of time. Lecture Notes in Computer Science, 810:216–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Myers:1994:RAR

[1312] K. L. Myers and K. Konolige. Reasoning with analogical representations. Lecture Notes in Computer Science, 810:229–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nejdl:1994:AAP

[1313] W. Nejdl and M. Banagl. Asking about possibilities — revision and update semantics for subjunctive queries. Lecture Notes in Computer Science, 810: 250–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niemelae:1994:ISC

[1314] I. Niemelae and J. Rintanen. On the impact of stratification on the complexity of nonmonotonic reasoning. Lecture Notes in Computer Science, 810: 275–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shoham:1994:LMA

[1315] Y. Shoham and S. B. Cousins. Logics of mental attitudes in AI. Lecture Notes in Computer Science, 810:296– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weydert:1994:HCM

[1316] E. Weydert. Hyperrational conditionals — monotonic reasoning about nested default conditionals. Lecture Notes in Computer Science, 810:310– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Witteveen:1994:REL

[1317] C. Witteveen and C. Jonker. Revision by expansion in logic programs. Lecture Notes in Computer Science, 810: 333–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wasserman:1994:ISM

[1318] A. I. Wasserman. Information systems modernization. Lecture Notes in Computer Science, 811:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conradi:1994:PSC

[1319] R. Conradi, M. Hagaseth, and C. Liu. Planning support for cooperating transactions in EPOS. Lecture Notes in Computer Science, 811:2–13, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marttiin:1994:TFP

[1320] P. Marttiin. Towards flexible process support with a CASE shell. Lecture Notes in Computer Science, 811:14–27, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lonchamp:1994:CPC

[1321] J. Lonchamp. A collaborative processcentered environment kernel. Lecture Notes in Computer Science, 811:28–41, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1994:BWA

[1322] G. Lu and M. Ader. Building workflow applications on top of WooRKS. Lecture Notes in Computer Science, 811: 42–52, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gulla:1994:MCW

[1323] J. A. Gulla and O. I. Lindland.

Modelling cooperative work for workflow management. Lecture Notes in Computer Science, 811:53–65, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tarumi:1994:IGD

[1324] H. Tarumi, A. Tabuchi, and K. Yoshifu.

IKUMI: A groupware development support system with visual environment. Lecture Notes in Computer Science, 811:66–79, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanSwede:1994:CDR

[1325] V. Van Swede and H. Van Vliet. Consistent development: Results of a first empirical study on the relation between project scenario and success. Lecture Notes in Computer Science, 811:80–93, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anton:1994:GDS

[1326] A. I. Anton, W. M. McCracken, and C. Potts. Goal decomposition and scenario analysis in business process reengineering. Lecture Notes in Computer Science, 811:94–104, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofmann:1994:RQC

[1327] H. F. Hofmann and R. Holbein. Reaching out for quality: Considering security requirements in the design of information systems. Lecture Notes in Computer Science, 811:105–118, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

WeiSong:1994:CSO

[1328] W. Wei Song. Capturing semantics by object relativity. Lecture Notes in Computer Science, 811:119–134, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jackson:1994:ASD

[1329] R. B. Jackson, D. W. Embley, and S. N. Woodfield. Automated support for the development of formal object-oriented requirements specifications. Lecture Notes in Computer Science, 811:135–148, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cauvet:1994:AFO

[1330] C. Cauvet and F. Semmak. Abstraction forms in object-oriented conceptual modelling: Localization, aggregation and generalization extensions. Lecture Notes in Computer Science, 811:149–171, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nellborn:1994:CIS

[1331] C. Nellborn and P. Holm. Capturing information systems requirements through enterprise and speech act modelling. Lecture Notes in Computer Science, 811:172–185, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teisseire:1994:DME

[1332] M. Teisseire, P. Poncelet, and R. Cicchetti. Dynamic modelling with events. Lecture Notes in Computer Science, 811:186–199, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johannesson:1994:RCI

[1333] P. Johannesson. Representation and communication in information systems — A speech act based approach. Lecture Notes in Computer Science, 811: 200–213, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurbel:1994:ACG

[1334] K. Kurbel. From analysis to code generation: Experiences from an information engineering project using I-CASE technology. Lecture Notes in Computer Science, 811:214–227, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sheu:1994:KBP

[1335] P. C.-Y. Sheu and S. Yoo. A knowledgebased program transformation system. Lecture Notes in Computer Science, 811:228–241, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonjour:1994:CBS

[1336] M. Bonjour and G. Falquet. Concept bases: A support to information systems integration. Lecture Notes in Computer Science, 811:242–255, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seo:1994:FDP

[1337] D. Seo and P. Loucopoulos. Formalisation of data and process model reuse using hierarchic data types. Lecture Notes in Computer Science, 811:256– 268, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jarzabek:1994:MMV

[1338] S. Jarzabek and Chew Lim Tan. Modelling multiple views of common features in software reengineering for reuse. Lecture Notes in Computer Science, 811:269–282, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freitag:1994:HBT

[1339] B. Freitag. A hypertext-based tool for large scale software reuse. Lecture Notes in Computer Science, 811:283– 296, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thieme:1994:ASI

[1340] C. Thieme and A. Siebes. An approach to schema integration based on transformations and behaviour. Lecture Notes in Computer Science, 811:297– 310, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sancho:1994:DTD

[1341] M. R. Sancho and A. Olive. Deriving transaction from deductive conceptual models of information systems. Lecture Notes in Computer Science, 811:311–324, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keszenheimer:1994:UBA

[1342] L. M. Keszenheimer. Utilizing behavioral abstractions to facilitate maintenance during class evolution. Lecture Notes in Computer Science, 811:325– 338, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prakash:1994:PVM

[1343] N. Prakash. A process view of methodologies. Lecture Notes in Computer Science, 811:339–352, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saeki:1994:SSS

[1344] M. Saeki and K. Wenyin. Specifying software specification and design methods. Lecture Notes in Computer Science, 811:353–366, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreno:1994:GAS

[1345] M. Moreno, C. Rolland, and C. Souveyet. A generic approach to support a way-of-working definition. Lecture Notes in Computer Science, 811:367– 379, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aberer:1994:DUO

[1346] K. Aberer, W. Klas, and A. L. Furtado. Designing a user-oriented query modification facility in object-oriented database systems. Lecture Notes in Computer Science, 811:380–393, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rosengren:1994:UVQ

[1347] P. Rosengren. Using visual ER query systems in real world applications. Lecture Notes in Computer Science, 811: 394–405, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Odberg:1994:CCF

[1348] E. Odberg. Category classes: Flexible classification and evolution in objectoriented databases. Lecture Notes in Computer Science, 811:406–419, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersen:1994:GCT

[1349] N. Andersen and N. D. Jones. Generalizing Cook’s transformation to imperative stack programs. Lecture Notes in Computer Science, 812:1–18, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berstel:1994:RFT

[1350] J. Berstel. A rewriting of Fife’s theorem about overlap-free words. Lecture Notes in Computer Science, 812:19–29, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brauer:1994:RJP

[1351] W. Brauer and U. Brauer. Reconsidering the jeep problem — or how to transport a birthday present to salosauna —. Lecture Notes in Computer Science, 812:30–33, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brueggemann-Klein:1994:LPS

[1352] A. Brueggemann-Klein, P. Fischer, and T. Ottmann. Learning picture sets from examples. Lecture Notes in Computer Science, 812:34–43, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calude:1994:RIN

[1353] C. Calude and H. Juergensen. Randomness as an invariant for number representations. Lecture Notes in Computer Science, 812:44–66, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Csuhaj-Varju:1994:CGS

[1354] E. Csuhaj-Varju. Cooperating grammars’ systems: Power and parameters. Lecture Notes in Computer Science, 812:67–84, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Culik:1994:PPG

[1355] K. Culik and J. Kari. Parallel pattern generation with one-way communications. Lecture Notes in Computer Science, 812:85–96, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrenfeucht:1994:DLS

[1356] A. Ehrenfeucht and G. Rozenberg. Dynamic labeled 2-Structures with variable domains. Lecture Notes in Computer Science, 812:97–123, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engelfriet:1994:DNP

[1357] J. Engelfriet. Deciding the NTS property of context-free grammars. Lecture Notes in Computer Science, 812:124– 130, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gecseg:1994:HRP

[1358] F. Gecseg. Homomorphic representations by products of tree automata. Lecture Notes in Computer Science, 812:131–139, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harju:1994:IT

[1359] T. Harju, H. J. Hoogeboom, and H. C. M. Kleijn. Identities and transductions. Lecture Notes in Computer Science, 812:140–144, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harju:1994:DIL

[1360] T. Harju and G. Rozenberg. Decomposition of infinite labeled 2-Structures. Lecture Notes in Computer Science, 812:145–158, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Honkala:1994:IPL

[1361] J. Honkala. An iteration property of Lindenmayerian power series. Lecture Notes in Computer Science, 812:159– 168, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hromkovic:1994:CDC

[1362] J. Hromkovic, J. Karhumaeki, and A. Lepistoe. Comparing descriptional and computational complexity of infinite words. Lecture Notes in Computer Science, 812:169–182, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ibarra:1994:SOP

[1363] O. H. Ibarra and T. Jiang. On some open problems concerning the complexity of cellular arrays. Lecture Notes in Computer Science, 812:183–196, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kari:1994:PCI

[1364] L. Kari. Power of controlled insertion and deletion. Lecture Notes in Computer Science, 812:197–212, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kelemenova:1994:CEG

[1365] A. Kelemenova and J. Kelemen. From colonies to eco(grammar)systems. Lecture Notes in Computer Science, 812: 213–231, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuich:1994:MEP

[1366] W. Kuich. On the multiplicity equivalence problem for context-free grammars. Lecture Notes in Computer Science, 812:232–250, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Makanin:1994:GSW

[1367] G. S. Makanin and H. Abdulrab. On general solution of word equations. Lecture Notes in Computer Science, 812:251–263, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mateescu:1994:LPS

[1368] A. Mateescu. On (left) partial shuffle. Lecture Notes in Computer Science, 812:264–278, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maurer:1994:LTA

[1369] H. Maurer. Learning theoretical aspects is important but (sometimes) dangerous. Lecture Notes in Computer Science, 812:279–288, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielsen:1994:BGL

[1370] M. Nielsen and C. Clausen. Bisimulation, games, and logic. Lecture Notes in Computer Science, 812:289–306, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niemi:1994:CPV

[1371] V. Niemi and A. Renvall. Cryptographic protocols and voting. Lecture Notes in Computer Science, 812:307– 316, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nurmi:1994:CPA

[1372] H. Nurmi. Cryptographic protocols for auctions and bargaining. Lecture Notes in Computer Science, 812:317–324, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paun:1994:SCC

[1373] G. Paun, J. Dassow, and S. Skalla. On the size of components of cooperating grammar systems. Lecture Notes in Computer Science, 812:325–343, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paz:1994:EAP

[1374] A. Paz. An elementary algorithmic problem from an advanced standpoint. Lecture Notes in Computer Science, 812:344–357, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruohonen:1994:EDO

[1375] K. Ruohonen. Event detection for ODEs and nonrecursive hierarchies. Lecture Notes in Computer Science, 812:358–371, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salomaa:1994:RPM

[1376] K. Salomaa, D. Wood, and S. Yu. Rediscovering pushdown machines. Lecture Notes in Computer Science, 812: 372–385, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simon:1994:SMA

[1377] I. Simon. String matching algorithms and automata. Lecture Notes in Computer Science, 812:386–395, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steinby:1994:CRL

[1378] M. Steinby. Classifying regular languages by their syntactic algebras. Lecture Notes in Computer Science, 812: 396–409, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Turakainen:1994:PME

[1379] P. Turakainen. On polynomial matrix equations X0T = p(X) and X = p(X) where all parameters are nonnegative. Lecture Notes in Computer Science, 812:410–421, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Welzl:1994:GEP

[1380] E. Welzl. Gram’s equation — A probabilistic proof. Lecture Notes in Computer Science, 812:422–424, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ablayev:1994:LBP

[1381] F. Ablayev. Lower bounds for probabilistic space complexity: Communicationautomata approach. Lecture Notes in Computer Science, 813:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersen:1994:MCI

[1382] H. R. Andersen. On model checking infinite-state systems. Lecture Notes in Computer Science, 813:8–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Archangelsky:1994:CPH

[1383] D. A. Archangelsky, M. I. Dekhtyar, E. Kruglov, and I. K. Musikaev. Concurrency problem for Horn fragment of Girard’s linear logic. Lecture Notes in Computer Science, 813:18–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Artemov:1994:RDS

[1384] S. Artemov and V. Krupski. Referential data structures and labeled modal logic. Lecture Notes in Computer Science, 813:23–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asperti:1994:YAC

[1385] A. Asperti and G. Dore. Yet another correctness criterion for multiplicative linear logic with MIX. Lecture Notes in Computer Science, 813:34–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berezin:1994:AEM

[1386] S. A. Berezin and N. V. Shilov. An approach to effective model-checking of real-time finite-state machines in mucalculus. Lecture Notes in Computer Science, 813:47–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1994:AC

[1387] C. Brown and A. Jeffrey. Allegories of circuits. Lecture Notes in Computer Science, 813:56–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1994:CPM

[1388] C.-C. Chen and I.-P. Lin. The complexity of propositional modal theories and the complexity of consistency of propositional modal theories. Lecture Notes in Computer Science, 813:69–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dekhtyar:1994:MLL

[1389] A. M. Dekhtyar. Multiplicative linear logic for resource transformation nets. Lecture Notes in Computer Science, 813:81–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Downey:1994:PCS

[1390] R. G. Downey, M. R. Fellows, B. M. Kapron, and M. T. Hallett. The parameterized complexity of some problems in logic and linguistics. Lecture Notes in Computer Science, 813:89–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galmiche:1994:FPS

[1391] D. Galmiche and G. Perrier. Foundations of proof search strategies design in linear logic. Lecture Notes in Computer Science, 813:101–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ge:1994:EPC

[1392] X. Ge and A. Nerode. On extreme points of convex compact Turing located set. Lecture Notes in Computer Science, 813:114–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghilezan:1994:ATL

[1393] S. Ghilezan. Application of typed lambda calculi in the untyped lambda calculus. Lecture Notes in Computer Science, 813:129–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goncharov:1994:CPE

[1394] S. S. Goncharov and S. A. Badaev.

Classes with pairwise equivalent enumerations. Lecture Notes in Computer Science, 813:140–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeGroote:1994:SNN

[1395] P. De Groote. Strong normalization in a non-deterministic typed lambdacalculus. Lecture Notes in Computer Science, 813:142–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heinemann:1994:ECM

[1396] B. Heinemann. On expressive completeness of modal logic. Lecture Notes in Computer Science, 813:153– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoofman:1994:CMN

[1397] R. Hoofman. Comparing models of the non-extensional typed lambdacalculus. Lecture Notes in Computer Science, 813:164–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacobs:1994:CA

[1398] B. Jacobs. Coalgebras and approximation. Lecture Notes in Computer Science, 813:173–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kanovich:1994:CCM

[1399] M. I. Kanovich. Computational and concurrency models of linear logic. Lecture Notes in Computer Science, 813: 184–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khasidashvili:1994:LPR

[1400] Z. Khasidashvili. The longest perpetual reductions in orthogonal expression reduction systems. Lecture Notes in Computer Science, 813:191–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khoussainov:1994:NRG

[1401] B. Khoussainov and A. Nerode. The notion of rank and games. Lecture Notes in Computer Science, 813:204– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuchuganov:1994:PLW

[1402] M. Kuchuganov. A predicate logic of well-founded actions. Lecture Notes in Computer Science, 813:218–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leivant:1994:PRF

[1403] D. Leivant. Predicative recurrence in finite types. Lecture Notes in Computer Science, 813:227–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Makowsky:1994:AVA

[1404] J. A. Makowsky and Y. B. Pnueli. Arity vs. alternation in second order logic. Lecture Notes in Computer Science, 813:240–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nickau:1994:HSF

[1405] H. Nickau. Hereditarily sequential functionals. Lecture Notes in Computer Science, 813:253–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nitsche:1994:PLT

[1406] U. Nitsche. Propositional linear temporal logic and language homomorphisms. Lecture Notes in Computer Science, 813:265–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perez:1994:APC

[1407] R. P. Perez and C. Even. An abstract property of confluence applied to the study of the lazy partial lambda calculus. Lecture Notes in Computer Science, 813:278–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pliuskeviciene:1994:SDA

[1408] A. Pliuskeviciene and R. Pliuskevicius. On specialization of derivations in axiomatic equality theories. Lecture Notes in Computer Science, 813:291– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rybakov:1994:PAI

[1409] V. V. Rybakov. Preserving of admissible inference rules in modal logic. Lecture Notes in Computer Science, 813: 304–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Severi:1994:PTS

[1410] P. Severi and E. Poll. Pure type systems with definitions. Lecture Notes in Computer Science, 813:316–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sidon:1994:CIP

[1411] T. Sidon. Craig interpolation property in modal logics with provability interpretation. Lecture Notes in Computer Science, 813:329–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Traylor:1994:RNV

[1412] B. Traylor and M. Gelfond. Representing null values in logic programming. Lecture Notes in Computer Science, 813:341–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanBakel:1994:CC

[1413] S. Van Bakel, L. Liquori, S. R. Della Rocca, and P. Urzyczyn. Comparing cubes. Lecture Notes in Computer Science, 813:353–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanderHoek:1994:LC

[1414] W. Van der Hoek, B. Van Linder, and J.-J. C. Meyer. A logic of capabilities. Lecture Notes in Computer Science, 813:366–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanOostrom:1994:WOI

[1415] V. Van Oostrom and F. Van Raamsdonk. Weak orthogonality implies confluence: The higher-order case. Lecture Notes in Computer Science, 813: 379–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Slaney:1994:CFM

[1416] J. Slaney. The crisis in finite mathematics: Automated reasoning as cause and cure. Lecture Notes in Computer Science, 814:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walsh:1994:DC

[1417] T. Walsh. A divergence critic. Lecture Notes in Computer Science, 814:14–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hutter:1994:SIO

[1418] D. Hutter. Synthesis of induction orderings for existence proofs. Lecture Notes in Computer Science, 814:29–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Protzen:1994:LGI

[1419] M. Protzen. Lazy generation of induction hypotheses. Lecture Notes in Computer Science, 814:42–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plaisted:1994:SET

[1420] D. A. Plaisted. The search efficiency of theorem proving strategies. Lecture Notes in Computer Science, 814:57–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bourely:1994:MBM

[1421] C. Bourely, R. Caferra, and N. Peltier. A method for building models automatically: Experiments with an extension of OTTER. Lecture Notes in Computer Science, 814:72–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baumgartner:1994:MEC

[1422] P. Baumgartner and U. Furbach. Model elimination without contrapositives. Lecture Notes in Computer Science, 814:87–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bronsard:1994:IUT

[1423] F. Bronsard, U. S. Reddy, and R. W. Hasker. Induction using term orderings. Lecture Notes in Computer Science, 814:102–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chazarain:1994:MIP

[1424] Jacques Chazarain and Emmanuel Kounalis. Mechanizable inductive proofs for a class of AE formulas. Lecture Notes in Computer Science, 814: 118–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lysne:1994:CBN

[1425] O. Lysne. On the connection between narrowing and proof by consistency. Lecture Notes in Computer Science, 814:133–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paulson:1994:FAI

[1426] L. C. Paulson. A fixedpoint approach to implementing (co)inductive definitions. Lecture Notes in Computer Science, 814:148–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wirth:1994:NIV

[1427] C.-P. Wirth and B. Gramlich. On notions of inductive validity for first-order equational clauses. Lecture Notes in Computer Science, 814:162–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baker:1994:NAE

[1428] S. Baker. A new application for explanation-based generalisation within automated deduction. Lecture Notes in Computer Science, 814:177– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chu:1994:SGF

[1429] H. Chu and D. A. Plaisted. Semantically guided first-order theorem proving using hyper-linking. Lecture Notes in Computer Science, 814:192– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeWaal:1994:ALP

[1430] D. A. De Waal and J. P. Gallagher.

The applicability of logic program analysis and transformation to theorem proving. Lecture Notes in Computer Science, 814:207–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruening:1994:DNP

[1431] S. Bruening. Detecting non-provable goals. Lecture Notes in Computer Science, 814:222–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyer:1994:MPC

[1432] R. S. Boyer, N. G. De Bruijn, G. Huet, and A. Trybulec. A mechanically proof-checked encyclopedia of mathematics: Should we build one? can we? Lecture Notes in Computer Science, 814:237–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sutcliffe:1994:TPL

[1433] G. Sutcliffe, C. Suttner, and T. Yemenis. The TPTP problem library. Lecture Notes in Computer Science, 814: 252–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Domenjoud:1994:CTN

[1434] E. Domenjoud, F. Klay, and C. Ringeissen. Combination techniques for nondisjoint equational theories. Lecture Notes in Computer Science, 814:267– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salzer:1994:PGU

[1435] G. Salzer. Primal grammars and unification modulo a binary clause. Lecture Notes in Computer Science, 814: 282–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iwanuma:1994:CQN

[1436] K. Iwanuma. Conservative query normalization on parallel circumscription. Lecture Notes in Computer Science, 814:296–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fribourg:1994:BED

[1437] L. Fribourg and M. Veloso Peixoto.

Bottom-up evaluation of datalog programs with arithmetic constraints. Lecture Notes in Computer Science, 814: 311–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Royer:1994:IQA

[1438] V. Royer and J. Joachim Quantz.

On intuitionistic query answering in description bases. Lecture Notes in Computer Science, 814:326–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stickel:1994:DCA

[1439] M. Stickel, R. Waldinger, M. Lowry, and T. Pressburger. Deductive composition of astronomical software from subroutine libraries. Lecture Notes in Computer Science, 814:341–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farmer:1994:PSP

[1440] W. M. Farmer, J. D. Guttman, M. E.

Nadel, and F. Javier Thayer. Proof script pragmatics in IMPS. Lecture Notes in Computer Science, 814:356– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kerber:1994:MSK

[1441] M. Kerber and M. Kohlhase. A mechanization of strong Kleene logic for partial functions. Lecture Notes in Computer Science, 814:371–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1994:AFG

[1442] D. Wang. Algebraic factoring and geometry theorem proving. Lecture Notes in Computer Science, 814:386– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

FreitagMcPhee:1994:MPG

[1443] N. Freitag McPhee, S.-C. Chou, and X.-S. Gao. Mechanically proving geometry theorems using a combination of Wu’s method and Collins’ method. Lecture Notes in Computer Science, 814:401–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hines:1994:SI

[1444] L. M. Hines. Str+ve and integers. Lecture Notes in Computer Science, 814: 416–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Platek:1994:WP

[1445] R. Platek. What is a proof? Lecture Notes in Computer Science, 814: 431–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1994:TGI

[1446] U. Martin. Termination, geometry and invariants. Lecture Notes in Computer Science, 814:432–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bachmair:1994:OCT

[1447] L. Bachmair and H. Ganzinger. Ordered chaining for total orderings. Lecture Notes in Computer Science, 814: 435–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Middeldorp:1994:STR

[1448] A. Middeldorp and H. Zantema. Simple termination revisited. Lecture Notes in Computer Science, 814:451–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basin:1994:TOR

[1449] D. A. Basin and T. Walsh. Termination orderings for rippling. Lecture Notes in Computer Science, 814:466–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sturgill:1994:NAP

[1450] D. B. Sturgill and A. M. Segre. A novel asynchronous parallelism scheme for first-order logic. Lecture Notes in Computer Science, 814:484–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goubault:1994:PBC

[1451] J. Goubault. Proving with BDDs and control of information. Lecture Notes in Computer Science, 814:499–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graf:1994:EPI

[1452] P. Graf. Extended path-indexing. Lecture Notes in Computer Science, 814: 514–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Constable:1994:ERA

[1453] R. L. Constable. Exporting and reflecting abstract metamathematics. Lecture Notes in Computer Science, 814:529– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vigneron:1994:ACD

[1454] L. Vigneron. Associative-commutative deduction with constraints. Lecture Notes in Computer Science, 814:530– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nieuwenhuis:1994:ASC

[1455] R. Nieuwenhuis and A. Rubio. ACsuperposition with constraints: No ACunifiers needed. Lecture Notes in Computer Science, 814:545–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hermann:1994:CCP

[1456] M. Hermann and P. G. Kolaitis. The complexity of counting problems in equational matching. Lecture Notes in Computer Science, 814:560–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1994:RPT

[1457] P. Anderson. Representing proof transformations for program optimization. Lecture Notes in Computer Science, 814:575–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jackson:1994:EAA

[1458] P. Jackson. Exploring abstract algebra in constructive type theory. Lecture Notes in Computer Science, 814: 590–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Felty:1994:TTP

[1459] A. Felty and D. Howe. Tactic theorem proving with refinement-tree proofs and metavariables. Lecture Notes in Computer Science, 814:605–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johann:1994:UEL

[1460] P. Johann and M. Kohlhase. Unification in an extensional lambda calculus with ordered function sorts and constant overloading. Lecture Notes in Computer Science, 814:620–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prehofer:1994:DHO

[1461] C. Prehofer. Decidable higher-order unification problems. Lecture Notes in Computer Science, 814:635–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mueller:1994:TPM

[1462] O. Mueller and F. Weber. Theory and practice of minimal modular higherorder E- unification. Lecture Notes in Computer Science, 814:650–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Socher-Ambrosius:1994:RVG

[1463] R. Socher-Ambrosius. A refined version of general E-unification. Lecture Notes in Computer Science, 814:665– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beckert:1994:CBM

[1464] B. Beckert. A completion-based method for mixed universal and rigid E- unification. Lecture Notes in Computer Science, 814:678–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buendgen:1994:PPP

[1465] R. Buendgen. On pot, pans and pudding or how to discover generalised critical pairs. Lecture Notes in Computer Science, 814:693–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klingenbeck:1994:STO

[1466] S. Klingenbeck and R. Haehnle. Semantic tableaux with ordering restrictions. Lecture Notes in Computer Science, 814:708–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Massacci:1994:SAT

[1467] F. Massacci. Strongly analytic tableaux for normal modal logics. Lecture Notes in Computer Science, 814:723– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1994:RPA

[1468] X. Huang. Reconstructing proofs at the assertion level. Lecture Notes in Computer Science, 814:738–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1994:PGF

[1469] J. Zhang. Problems on the generation of finite models. Lecture Notes in Computer Science, 814:753–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clarke:1994:CSC

[1470] E. Clarke and X. Zhao. Combining symbolic computation and theorem proving: Some problems of ramanujan. Lecture Notes in Computer Science, 814:758–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Slaney:1994:SSC

[1471] J. Slaney, E. Lusk, and W. W. McCune.

SCOTT: Semantically constrained OT-TER. Lecture Notes in Computer Science, 814:764–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baumgartner:1994:PPT

[1472] P. Baumgartner and U. Furbach. PRO-TEIN: A PROver with a theory extension INterface. Lecture Notes in Computer Science, 814:769–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schumann:1994:DBP

[1473] J. M. P. Schumann. DELTA — A bottom-up preprocessor for top-down theorem provers. Lecture Notes in Computer Science, 814:774–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goller:1994:SVR

[1474] C. Goller, R. Letz, K. Mayr, and J. M. P. Schumann. SETHEO V3.2: Recent developments. Lecture Notes in Computer Science, 814:778–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bibel:1994:K

[1475] W. Bibel, S. Bruening, U. Egly, and T. Rath. KoMeT. Lecture Notes in Computer Science, 814:783–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1994:MPD

[1476] X. Huang, M. Kerber, M. Kohlhase, and E. Melis. -MKRP: A proof development environment. Lecture Notes in Computer Science, 814:788–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beckert:1994:LLT

[1477] B. Beckert and J. Posegga. leanTAP: Lean tableau-based theorem proving. Lecture Notes in Computer Science, 814:793–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Slaney:1994:FFD

[1478] J. Slaney. FINDER: Finite domain enumerator. Lecture Notes in Computer Science, 814:798–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Portoraro:1994:SAA

[1479] F. D. Portoraro. Symlog: Automated advice in fitch-style proof construction. Lecture Notes in Computer Science, 814:802–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1994:KTA

[1480] X. Huang, M. Kerber, M. Kohlhase, and E. Melis. KEIM: A toolkit for automated deduction. Lecture Notes in Computer Science, 814:807–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfenning:1994:EML

[1481] F. Pfenning. Elf: A meta-language for deductive systems. Lecture Notes in Computer Science, 814:811–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohtani:1994:EIT

[1482] T. Ohtani, H. Sawamura, and T. Minami. EUODHILOS-II on top of GNU epoch. Lecture Notes in Computer Science, 814:816–820, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eriksson:1994:PID

[1483] L.-H. Eriksson. Pi: An interactive derivation editor for the calculus of partial inductive definitions. Lecture Notes in Computer Science, 814:821– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Richards:1994:MGP

[1484] B. L. Richards, I. Kraan, A. Smaill, and G. A. Wiggins. Mollusc: A general proof-development shell for sequent-based logics. Lecture Notes in Computer Science, 814:826–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1994:KAI

[1485] T. C. Wang and A. Goldberg. KITP-93: An automated inference system for program analysis. Lecture Notes in Computer Science, 814:831–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouhoula:1994:SSS

[1486] A. Bouhoula. SPIKE: A system for sufficient completeness and parameterized inductive proofs. Lecture Notes in Computer Science, 814:836–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonacina:1994:DTP

[1487] M. P. Bonacina and W. W. Mc-Cune. Distributed theorem proving by peers. Lecture Notes in Computer Science, 814:841–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clarke:1994:AVF

[1488] E. M. Clarke. Automatic verification of finite-state concurrent systems. Lecture Notes in Computer Science, 815:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiCesare:1994:APN

[1489] F. DiCesare, P. T. Kulp, M. Gile, and G. List. The application of Petri nets to the modeling, analysis and control of intelligent urban traffic networks. Lecture Notes in Computer Science, 815:2–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meseguer:1994:MCP

[1490] J. Meseguer, U. Montanari, and V. Sassone. On the model of computation of place/transition Petri nets. Lecture Notes in Computer Science, 815:16–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anglano:1994:BWA

[1491] C. Anglano and L. Portinale. B-W analysis: a backward reachability analysis for diagnostic problem solving suitable to parallel implementation. Lecture Notes in Computer Science, 815:39–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bennacer:1994:PVR

[1492] N. Bennacer, G. Florin, C. Fraize, and S. Natkin. Probabilistic validation of a remote procedure call protocol. Lecture Notes in Computer Science, 815:59–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berthelot:1994:OGI

[1493] G. Berthelot and H. Boucheneb. Occurrence graphs for interval timed coloured nets. Lecture Notes in Computer Science, 815:79–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Briz:1994:IWP

[1494] J. L. Briz and J. M. Colom. Implementation of weighted place/transition nets based on linear enabling functions. Lecture Notes in Computer Science, 815:99–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buchholz:1994:HHL

[1495] P. Buchholz. Hierarchical high level Petri nets for complex system analysis. Lecture Notes in Computer Science, 815:119–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chiola:1994:MSC

[1496] G. Chiola, G. Franceschinis, and R. Gaeta. Modelling symmetric computer architectures by SWNs. Lecture Notes in Computer Science, 815:139– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christensen:1994:CPN

[1497] S. Christensen and N. D. Hansen. Coloured Petri nets extended with channels for synchronous communication. Lecture Notes in Computer Science, 815:159–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciardo:1994:PNM

[1498] G. Ciardo. Petri nets with markingdependent arc cardinality: properties and analysis. Lecture Notes in Computer Science, 815:179–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Couvreur:1994:NSI

[1499] J. M. Couvreur and E. Paviot-Adet.

New structural invariants for Petri nets analysis. Lecture Notes in Computer Science, 815:199–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diaz:1994:TSP

[1500] M. Diaz and P. Senac. Time stream Petri nets: a model for timed multimedia information. Lecture Notes in Computer Science, 815:219–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dietz:1994:TRP

[1501] C. Dietz and G. Schreiber. A term representation of P/T systems. Lecture Notes in Computer Science, 815: 239–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Donatelli:1994:SGS

[1502] S. Donatelli. Superposed generalized stochastic Petri nets: definition and efficient solution. Lecture Notes in Computer Science, 815:258–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Genrich:1994:VRB

[1503] H. J. Genrich, H.-M. Hanisch, and K. Woellhaf. Verification of recipebased control procedures by means of predicate/transition nets. Lecture Notes in Computer Science, 815:278– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keller:1994:TCS

[1504] R. K. Keller, M. Ozkan, and X. Shen. Towards comprehensive support for the dynamic analysis of Petri net based models. Lecture Notes in Computer Science, 815:298–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koutny:1994:POS

[1505] M. Koutny. Partial order semantics of box expressions. Lecture Notes in Computer Science, 815:318–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lakos:1994:GSA

[1506] C. Lakos and S. Christensen. A general systematic approach to arc extensions for coloured Petri nets. Lecture Notes in Computer Science, 815:338– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lautenbach:1994:LBP

[1507] K. Lautenbach and H. Ridder. Liveness in bounded Petri nets which are covered by t-invariants. Lecture Notes in Computer Science, 815:358–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mortensen:1994:MWF

[1508] K. H. Mortensen and V. Pinci. Modelling the work flow of a nuclear waste management program. Lecture Notes in Computer Science, 815:376–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielsen:1994:DSA

[1509] D. S. Nielsen and L. Kleinrock. Data structures and algorithms for extended state space and structural level reduction of the GSPN model. Lecture Notes in Computer Science, 815:396– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pastor:1994:PNA

[1510] E. Pastor, O. Roig, J. Cortadella, and R. M. Badia. Petri net analysis using Boolean manipulation. Lecture Notes in Computer Science, 815:416– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruckdeschel:1994:MPB

[1511] W. Ruckdeschel and R. Onken. Modelling of pilot behaviour using Petri nets. Lecture Notes in Computer Science, 815:436–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scheschonk:1994:SAD

[1512] G. Scheschonk and M. Timpe. Simulation and analysis of a document storage system. Lecture Notes in Computer Science, 815:454–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sibertin-Blanc:1994:CN

[1513] C. Sibertin-Blanc. Cooperative nets.

Lecture Notes in Computer Science, 815:471–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teruel:1994:WFE

[1514] E. Teruel and M. Silva. Wellformedness of equal conflict systems. Lecture Notes in Computer Science, 815:491–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tiusanen:1994:SSS

[1515] M. Tiusanen. Symbolic, symmetry, and stubborn set searches. Lecture Notes in Computer Science, 815:511–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valmari:1994:CAP

[1516] A. Valmari. Compositional analysis with place-bordered subnets. Lecture Notes in Computer Science, 815:531– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Varpaaniemi:1994:CSS

[1517] K. Varpaaniemi. On combining the stubborn set method with the sleep set method. Lecture Notes in Computer Science, 815:548–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yakovlev:1994:CMH

[1518] A. Yakovlev, M. Kishinevsky, A. Kondratyev, and L. Lavagno. OR causality: modelling and hardware implementation. Lecture Notes in Computer Science, 815:568–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asperti:1994:IS

[1519] A. Asperti and C. Laneve. Interaction systems. Lecture Notes in Computer Science, 816:1–19, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanBakel:1994:SNT

[1520] S. Van Bakel and M. Fernandez. Strong normalization of typeable rewrite systems. Lecture Notes in Computer Science, 816:20–39, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellegarde:1994:TSC

[1521] F. Bellegarde. A transformation system combining partial evaluation with term rewriting. Lecture Notes in Computer Science, 816:40–55, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berghammer:1994:PRS

[1522] R. Berghammer, T. F. Gritzner, and G. Schmidt. Prototyping relational specifications using higher-order objects. Lecture Notes in Computer Science, 816:56–75, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanDeursen:1994:OTH

[1523] A. Van Deursen and T. B. Dinesh.

Origin tracking for higher-order term rewriting systems. Lecture Notes in Computer Science, 816:76–95, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farmer:1994:TIS

[1524] W. H. Farmer. Theory interpretation in simple type theory. Lecture Notes in Computer Science, 816:96–123, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grosu:1994:SS

[1525] R. Grosu and F. Regensburger. The semantics of spectrum. Lecture Notes in Computer Science, 816:124–145, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hearn:1994:ATL

[1526] B. M. Hearn and K. Meinke. ATLAS: A typed language for algebraic specification. Lecture Notes in Computer Science, 816:146–168, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kahrs:1994:CCR

[1527] S. Kahrs. Compilation of combinatory reduction systems. Lecture Notes in Computer Science, 816:169–188, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meinke:1994:SVH

[1528] K. Meinke and L. J. Steggles. Specification and verification in higher-order algebra: A case study of convolution. Lecture Notes in Computer Science, 816:189–222, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moeller:1994:OCM

[1529] B. Moeller. Ordered and continuous models of higher-order specifications. Lecture Notes in Computer Science, 816:223–255, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanOostrom:1994:CCR

[1530] V. Van Oostrom and F. Van Raamsdonk. Comparing combinatory reduction systems and higher-order rewrite systems. Lecture Notes in Computer Science, 816:276–304, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VandePol:1994:TPH

[1531] J. Van de Pol. Termination proofs for higher-order rewrite systems. Lecture Notes in Computer Science, 816:305– 325, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schobbens:1994:EIM

[1532] P.-Y. Schobbens. Extensions of initial models and their second-order proof systems. Lecture Notes in Computer Science, 816:326–344, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferreira:1994:IPR

[1533] A. G. Ferreira and M. D. Grammatikakis. Improved probabilistic routing on generalized hypercubes. Lecture Notes in Computer Science, 817:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1994:EDC

[1534] J. Liu and W.-J. Hsu. Efficient data communication in incomplete hypercubes. Lecture Notes in Computer Science, 817:13–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oehring:1994:ECF

[1535] S. R. Oehring and S. K. Das. Efficient communication in the folded Petersen interconnection networks. Lecture Notes in Computer Science, 817: 25–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lincoln:1994:CRS

[1536] P. Lincoln, N. Marti-Oliet, J. Meseguer, and L. Ricciulli. Compiling rewriting onto SIMD and MIMD/ SIMD machines. Lecture Notes in Computer Science, 817:37–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pande:1994:CTV

[1537] S. Pande and K. Psarris. A compilation technique for varying communication cost NUMA architectures. Lecture Notes in Computer Science, 817:49–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OBoyle:1994:DPA

[1538] M. O’Boyle. A data partitioning algorithm for distributed memory compilation. Lecture Notes in Computer Science, 817:61–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

FiallosAguilar:1994:THP

[1539] M. Fiallos Aguilar and J. Duprat. Towards a high precision massively parallel computer. Lecture Notes in Computer Science, 817:73–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vergis:1994:TPR

[1540] A. Vergis and V. Verykios. On the testability of purely recursive digital filters. Lecture Notes in Computer Science, 817:85–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hopkins:1994:VSC

[1541] R. P. Hopkins. Voting on synchronous communications. Lecture Notes in Computer Science, 817:97–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benslimane:1994:PVP

[1542] A. Benslimane. Protocol validation: A parallel technique to reduce the reachability tree. Lecture Notes in Computer Science, 817:109–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bampis:1994:OPE

[1543] E. Bampis, J.-C. Koenig, and D. Trystram. Optimal parallel execution of complete binary trees and grids into most popular interconnection networks. Lecture Notes in Computer Science, 817:122–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Olariu:1994:TON

[1544] S. Olariu and I. Stojmenovic. Timeoptimal nearest-neighbor computations on enhanced meshes. Lecture Notes in Computer Science, 817:134–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leppaenen:1994:SPM

[1545] V. Leppaenen and M. Penttonen. Simulation of PRAM models on meshes. Lecture Notes in Computer Science, 817:146–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atallah:1994:MPH

[1546] M. J. Atallah and A. Fabri. On the multisearching problem for hypercubes. Lecture Notes in Computer Science, 817:159–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

PratherPineo:1994:CLT

[1547] P. Prather Pineo. Code liberation — A tool for refitting code to a parallel environment. Lecture Notes in Computer Science, 817:167–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prestwich:1994:HLV

[1548] S. D. Prestwich and A. J. Kusalik. High-level visualization of parallel logic program execution. Lecture Notes in Computer Science, 817:180–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hackstadt:1994:NGP

[1549] S. T. Hackstadt and A. D. Malony. Next-generation parallel performance visualization: A prototyping environment for visualization development. Lecture Notes in Computer Science, 817:192–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sharnowski:1994:FAM

[1550] J. L. Sharnowski and B. H. C. Cheng. A formal approach to modeling expected behavior in parallel program visualizations. Lecture Notes in Computer Science, 817:202–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dai:1994:ISD

[1551] H. K. Dai. An improvement in the size-depth tradeoff for strictly nonblocking generalized-concentration networks. Lecture Notes in Computer Science, 817:214–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishikawa:1994:BSH

[1552] T. Ishikawa. Bypass strategy in hypercube multiprocessors by adding bus connections. Lecture Notes in Computer Science, 817:226–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghose:1994:OHM

[1553] K. Ghose. OPTIMUL: A hybrid multiprocessor for distributed and shared memory multiprocessing with WDM optical fiber interconnections. Lecture Notes in Computer Science, 817:238– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loyens:1994:ISL

[1554] L. D. J. C. Loyens and J. R. Moonen. ILIAS: A sequential language for parallel matrix computations. Lecture Notes in Computer Science, 817:250– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

GautierdeLahaut:1994:SCP

[1555] D. Gautier de Lahaut and C. Germain. Static communications in parallel scientific programs. Lecture Notes in Computer Science, 817:262–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauber:1994:LBE

[1556] T. Rauber and G. Ruenger. Load balancing for extrapolation methods on distributed memory multiprocessors. Lecture Notes in Computer Science, 817:277–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ertel:1994:DS

[1557] W. Ertel. On the definition of speedup. Lecture Notes in Computer Science, 817:289–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wabnig:1994:PNP

[1558] H. Wabnig and G. Haring. Petri net performance models of parallel systems — methodology and case study. Lecture Notes in Computer Science, 817: 301–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ziantz:1994:RTO

[1559] L. H. Ziantz, C. C. Oezturan, and B. K. Szymanski. Run-time optimization of sparse matrix-vector multiplication on SIMD machines. Lecture Notes in Computer Science, 817:313–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wakatani:1994:NAA

[1560] A. Wakatani and M. Wolfe. A new approach to array redistribution: Strip mining redistribution. Lecture Notes in Computer Science, 817:323–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carreras:1994:AMC

[1561] C. Carreras, C. A. Lopez, and M. Hermenegildo. Analytic model of a cache only memory architecture. Lecture Notes in Computer Science, 817: 336–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agarwal:1994:MAM

[1562] A. Agarwal, J. V. Guttag, C. N. Hadjicostis, and M. C. Papaefthymiou. Memory assignment for multiprocessor caches through grey coloring. Lecture Notes in Computer Science, 817:351– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nilsson:1994:AUB

[1563] H. Nilsson and P. Stenstroem. An adaptive update-based cache coherence protocol for reduction of miss rate and traffic. Lecture Notes in Computer Science, 817:363–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bennett:1994:EIC

[1564] A. J. Bennett and P. H. J. Kelly.

Eliminating invalidation in coherentcache parallel graph reduction. Lecture Notes in Computer Science, 817: 375–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1994:EPR

[1565] F. Zhang and E. H. D’Hollander. Enhancing parallelism by removing cyclic data dependencies. Lecture Notes in Computer Science, 817:387–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takeuchi:1994:IBL

[1566] K. Takeuchi, K. Honda, and M. Kubo. An interaction-based language and its typing system. Lecture Notes in Computer Science, 817:398–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papadopoulos:1994:PD

[1567] C. V. Papadopoulos. On the parallelism of data. Lecture Notes in Computer Science, 817:414–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

ThudichumVasconcelos:1994:PPC

[1568] V. Thudichum Vasconcelos. Predicative polymorphism in π-calculus. Lecture Notes in Computer Science, 817: 425–437, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shen:1994:IED

[1569] K. Shen. Improving the execution of the dependent And-Parallel Prolog DDAS. Lecture Notes in Computer Science, 817:438–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kacsuk:1994:PPD

[1570] P. Kacsuk. OR-Parallel Prolog on distributed memory systems. Lecture Notes in Computer Science, 817:453– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuchen:1994:EDM

[1571] H. Kuchen, R. Plasmeijer, and H. Stoltze. Efficient distributed memory implementation of a data parallel functional language. Lecture Notes in Computer Science, 817:464–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abramson:1994:MPP

[1572] D. Abramson, M. Cope, and R. McKenzie. Modelling photochemical pollution using parallel and distributed computing platforms. Lecture Notes in Computer Science, 817:478–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bull:1994:PAI

[1573] J. M. Bull and T. L. Freeman. Parallel algorithms and interval selection strategies for globally adaptive quadrature. Lecture Notes in Computer Science, 817:490–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brudaru:1994:SDE

[1574] O. Brudaru and G. M. Megson. Systolic designs for evaluating linear combinations of Chebyshev polynomials. Lecture Notes in Computer Science, 817: 502–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dageville:1994:IKA

[1575] B. Dageville, P. Casadessus, and P. Borla-Salamet. The impact of the KSR1 allcache architecture on the behavior of the DBS3 parallel DBMS. Lecture Notes in Computer Science, 817:514–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liotopoulos:1994:SGH

[1576] F. K. Liotopoulos. Sparse generalized HyperGrids for performance scalability. Lecture Notes in Computer Science, 817:526–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puaut:1994:DGC

[1577] I. Puaut. A distributed garbage collector for active objects. Lecture Notes in Computer Science, 817:539–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferreira:1994:BBP

[1578] A. G. Ferreira, A. G. Vel Lejbman, and S. W. Song. Bus-based parallel computers: A viable way for massive parallelism. Lecture Notes in Computer Science, 817:553–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barrado:1994:IG

[1579] C. Barrado, J. Labarta, and P. Borensztejn. Implementation of GTS. Lecture Notes in Computer Science, 817: 565–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1994:HAM

[1580] C. Kim and J.-L. Gaudiot. A hierarchical activation management technique for fine-grain multithreaded execution. Lecture Notes in Computer Science, 817:577–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mehdiratta:1994:STG

[1581] N. Mehdiratta and K. Ghose. Scheduling task graphs onto distributed memory multiprocessors under realistic constraints. Lecture Notes in Computer Science, 817:589–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papadopoulos:1994:DAO

[1582] C. V. Papadopoulos. A dynamic algorithm for online scheduling of parallel processes. Lecture Notes in Computer Science, 817:601–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boniol:1994:CSC

[1583] F. Boniol. COREA: A synchronous calculus of parallel communicating reactive automata. Lecture Notes in Computer Science, 817:611–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janicki:1994:DHN

[1584] R. Janicki and M. Koutny. Deriving histories of nets with priority relation. Lecture Notes in Computer Science, 817:623–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radestock:1994:WDY

[1585] M. Radestock and S. Eisenbach. What do you get from a Π-calculus semantics? Lecture Notes in Computer Science, 817:635–647, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Viry:1994:REM

[1586] P. Viry. Rewriting: An effective model of concurrency. Lecture Notes in Computer Science, 817:648–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fowler:1994:MOA

[1587] R. J. Fowler and L. Kontothanassis. Mercury: Object-affinity scheduling and continuation passing on multiprocessors. Lecture Notes in Computer Science, 817:661–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smai:1994:DAC

[1588] A.-H. Smai and L.-E. Thorelli. Dynamic allocation of communication bandwidth in multicomputers. Lecture Notes in Computer Science, 817:677– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1994:LBP

[1589] A. K. Gupta and A. G. Photiou. Loadbalanced priority queues on distributed memory machines. Lecture Notes in Computer Science, 817:689–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roch:1994:PCA

[1590] J. L. Roch and G. Villard. Parallel computations with algebraic numbers — A case study: Jordan normal form of matrices. Lecture Notes in Computer Science, 817:701–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferragina:1994:BDA

[1591] P. Ferragina and F. Luccio. Batch dynamic algorithms for two graph problems. Lecture Notes in Computer Science, 817:713–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schollmeyer:1994:GMM

[1592] M. Schollmeyer and B. McMillin. A general method for maximizing the error-detecting ability of distributed algorithms. Lecture Notes in Computer Science, 817:725–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mourlas:1994:MMA

[1593] C. Mourlas and C. Halatsis. MAGE:

A massively agent execution system for logic programming. Lecture Notes in Computer Science, 817:737–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caplain:1994:CSC

[1594] G. Caplain, R. Lalement, and T. Salset.

Checking the serial correctness of control-parallel programs. Lecture Notes in Computer Science, 817:741– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stellner:1994:NPP

[1595] G. Stellner, A. Bode, S. Lamberts, and T. Ludwig. NXLib — A parallel programming environment for workstation clusters. Lecture Notes in Computer Science, 817:745–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hondroudakis:1994:AHF

[1596] A. Hondroudakis and R. Procter. Applying human factors techniques to the design of a tool for parallel program performance analysis and tuning. Lecture Notes in Computer Science, 817:749–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desbiens:1994:ABP

[1597] J. Desbiens, M. Toulouse, M. Lavoie, and S. Pouzyreff. An actor-based programming system for heterogeneous processing. Lecture Notes in Computer Science, 817:753–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pena:1994:IFC

[1598] R. Peña and L. A. Galán. Introducing fairness in CSP systems specified with state variables. Lecture Notes in Computer Science, 817:757–760, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kusakabe:1994:DBM

[1599] S. Kusakabe, E. Takahashi, R.-I. Taniguchi, and M. Amamiya. A dataflow-based massively parallel programming language and its implementation. Lecture Notes in Computer Science, 817:761–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Listl:1994:USC

[1600] A. Listl. Using subpages for coherency control in parallel database systems. Lecture Notes in Computer Science, 817:765–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verdier:1994:NMD

[1601] C. Verdier, A. Lafage, E. Boutillon, and A. Demeure. A new multi-dimensional interconnection network for SIMD architectures. Lecture Notes in Computer Science, 817:769–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neschen:1994:SBS

[1602] M. Neschen and M. Gumm. A scalable bit-sequential SIMD architecture for pattern recognition. Lecture Notes in Computer Science, 817:773–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1994:IAP

[1603] S. P. S. Lam. An iterative array processor architecture for matrix computation. Lecture Notes in Computer Science, 817:777–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kawano:1994:DIF

[1604] T. Kawano, S. Kusakabe, R.-I.

Taniguchi, and M. Amamiya. Datarol-II: A fine-grain massively parallel architecture. Lecture Notes in Computer Science, 817:781–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalantery:1994:BVN

[1605] N. Kalantery, S. C. Winter, and D. R. Wilson. From BSP to a virtual von Neumann machine. Lecture Notes in Computer Science, 817:785–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okada:1994:DPR

[1606] H. Okada, K. Matsumoto, and M. Kitsuregawa. Data parallel relational database processing on the Connection Machine CM-2 with Data Vault. Lecture Notes in Computer Science, 817: 789–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hurson:1994:SSL

[1607] A. R. Hurson, J. T. Lim, B. Shirazi, and K. Kavi. Staggered Scheme: A loop allocation policy. Lecture Notes in Computer Science, 817:793–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borut:1994:UPS

[1608] R. Borut and J. Silc. Using parallel simulated annealing in the mapping problem. Lecture Notes in Computer Science, 817:797–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johansson:1994:NPM

[1609] T. Johansson and E. Bengtsson. A new parallel MIND connected component labeling algorithm. Lecture Notes in Computer Science, 817:801–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Markatos:1994:URC

[1610] E. Markatos and C. E. Chronaki. Using reference counters in update-based coherent memory. Lecture Notes in Computer Science, 817:805–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schepers:1994:CRA

[1611] H. Schepers. Compositional reasoning about real-time distributed systems with limited resources. Lecture Notes in Computer Science, 817:809–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lioupis:1994:DLB

[1612] D. Lioupis and M. Stefanidakis. Dynamic load balancing on a virtuallyshared memory parallel computer system. Lecture Notes in Computer Science, 817:813–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schumann:1994:ALB

[1613] J. Schumann and M. Jobmann.

Analysing the load balancing scheme of a parallel system on multiprocessors — A modeling approach. Lecture Notes in Computer Science, 817:819–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nemawarkar:1994:PIN

[1614] S. S. Nemawarkar, R. Govindarajan, G. R. Gao, and V. K. Agarwal. Performance of interconnection network in multithreaded architectures. Lecture Notes in Computer Science, 817:823– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopes:1994:SAP

[1615] L. M. B. Lopes and F. M. A. Silva. Scheduling algorithms performance with the pSystem parallel programming environment. Lecture Notes in Computer Science, 817:827– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seo:1994:NRS

[1616] S.-W. Seo and T.-Y. Feng. A new routing scheme for concatenating two Omega networks. Lecture Notes in Computer Science, 817:831–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alur:1994:DCT

[1617] R. Alur, L. Fix, and T. A. Henzinger. A determinizable class of timed automata. Lecture Notes in Computer Science, 818:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gorrieri:1994:RTS

[1618] R. Gorrieri and G. Siliprandi. Realtime system verification using P/T nets. Lecture Notes in Computer Science, 818:14–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1994:CSP

[1619] W. K. C. Lam and R. K. Brayton. Criteria for the simple path property in timed automata. Lecture Notes in Computer Science, 818:27–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McMillan:1994:HRD

[1620] K. L. McMillan. Hierarchical representations of discrete functions, with application to model checking. Lecture Notes in Computer Science, 818:41–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boigelot:1994:SVP

[1621] B. Boigelot and P. Wolper. Symbolic verification with periodic sets. Lecture Notes in Computer Science, 818:55–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burch:1994:AVP

[1622] J. R. Burch and D. L. Dill. Automatic verification of pipelined microprocessor control. Lecture Notes in Computer Science, 818:68–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Olivero:1994:UAV

[1623] A. Olivero, J. Sifakis, and S. Yovine.

Using abstractions for the verification of linear hybrid systems. Lecture Notes in Computer Science, 818:81–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puri:1994:DHS

[1624] A. Puri and P. Varaiya. Decidability of hybrid systems with rectangular differential inclusions. Lecture Notes in Computer Science, 818:95–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McManis:1994:SAD

[1625] J. McManis and P. Varaiya. Suspension automata: A decidable class of hybrid automata. Lecture Notes in Computer Science, 818:105–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouajjani:1994:VCF

[1626] A. Bouajjani, R. Echahed, and R. Robbana. Verification of context-free timed systems using linear hybrid observers. Lecture Notes in Computer Science, 818:118–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mihail:1994:RWM

[1627] M. Mihail and C. H. Papadimitriou. On the random walk method for protocol testing. Lecture Notes in Computer Science, 818:132–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernholtz:1994:ATA

[1628] O. Bernholtz, M. Y. Vardi, and P. Wolper. An automata-theoretic approach to branching-time model checking. Lecture Notes in Computer Science, 818:142–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anuchitanukul:1994:RSR

[1629] A. Anuchitanukul and Z. Manna. Realizability and synthesis of reactive modules. Lecture Notes in Computer Science, 818:156–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hungar:1994:MCM

[1630] H. Hungar. Model checking of macro processes. Lecture Notes in Computer Science, 818:169–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beer:1994:MSP

[1631] I. Beer, S. Ben-David, D. Geist, and R. Gewirtzman. Methodology and system for practical formal verification of reactive hardware. Lecture Notes in Computer Science, 818:182–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naik:1994:MVR

[1632] V. G. Naik and A. P. Sistla. Modeling and verification of a real life protocol using symbolic-model checking. Lecture Notes in Computer Science, 818:194–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graf:1994:VDC

[1633] S. Graf. Verification of a distributed cache memory by using abstractions. Lecture Notes in Computer Science, 818:207–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manna:1994:BMC

[1634] Z. Manna. Beyond model checking. Lecture Notes in Computer Science, 818:220–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurshan:1994:MWC

[1635] R. P. Kurshan. Models whose checks don’t explode. Lecture Notes in Computer Science, 818:222–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balarin:1994:ACN

[1636] F. Balarin and A. L. Sangiovanni-Vincentelli. On the automatic computation of network invariants. Lecture Notes in Computer Science, 818:234– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cyrluk:1994:GTL

[1637] D. Cyrluk and P. Narendran. Ground temporal logic: A logic for hardware verification. Lecture Notes in Computer Science, 818:247–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schubert:1994:HMR

[1638] E. T. Schubert. A hybrid model for reasoning about composed hardware systems. Lecture Notes in Computer Science, 818:260–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hazelhurst:1994:CST

[1639] S. Hazelhurst and C.-J. H. Seger. Composing symbolic trajectory evaluation results. Lecture Notes in Computer Science, 818:273–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhu:1994:CHI

[1640] Z. Zhu and C.-J. H. Seger. The completeness of a hardware inference system. Lecture Notes in Computer Science, 818:286–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geist:1994:EMC

[1641] D. Geist and I. Beer. Efficient model checking by automated ordering of transition relation partitions. Lecture Notes in Computer Science, 818:299– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Singhal:1994:VPS

[1642] V. Singhal and C. Pixley. The verification problem for safe replaceability. Lecture Notes in Computer Science, 818:311–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aziz:1994:FDE

[1643] A. Aziz, T. R. Shiple, V. Singhal, and A. L. Sangiovanni-Vincentelli. Formula-dependent equivalence for compositional CTL model checking. Lecture Notes in Computer Science, 818:324–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Long:1994:IAE

[1644] D. E. Long, A. Browne, E. M. Clarke, and S. Jha. An improved algorithm for the evaluation of fixpoint expressions. Lecture Notes in Computer Science, 818:338–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sokolsky:1994:IMC

[1645] O. V. Sokolsky and S. A. Smolka. Incremental model checking in the modal mu-calculus. Lecture Notes in Computer Science, 818:351–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cousin:1994:PIS

[1646] B. Cousin and J. Helary. Performance improvement of state space exploration by regular and differential hashing functions. Lecture Notes in Computer Science, 818:364–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peled:1994:CPO

[1647] D. Peled. Combining partial order reductions with on-the-fly modelchecking. Lecture Notes in Computer Science, 818:377–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hojati:1994:ILC

[1648] R. Hojati, R. Mueller-Thuns, and R. K. Brayton. Improving language containment using fairness graphs. Lecture Notes in Computer Science, 818:391– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1994:PAR

[1649] I. Lee and S. Rajasekaran. A parallel algorithm for relational coarsest partition problems and its implementation. Lecture Notes in Computer Science, 818:404–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clarke:1994:ALL

[1650] E. Clarke, O. Grumberg, and K. Hamaguchi. Another look at LTL model checking. Lecture Notes in Computer Science, 818:415–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Victor:1994:MWT

[1651] B. Victor and F. Moller. The mobility workbench: A tool for the picalculus. Lecture Notes in Computer Science, 818:428–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeSimone:1994:CSE

[1652] R. De Simone and A. Ressouche. Compositional semantics of Esterel and verification by compositional reductions. Lecture Notes in Computer Science, 818:441–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dams:1994:MCU

[1653] D. Dams, R. Gerth, G. Doehmen, and R. Herrmann. Model checking using adaptive state and data abstraction. Lecture Notes in Computer Science, 818:455–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rokicki:1994:AVT

[1654] T. G. Rokicki and C. J. Myers.

Automatic verification of timed circuits. Lecture Notes in Computer Science, 818:468–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yap:1994:DSN

[1655] M. Yap. Developing the Singapore National Information Infrastructure. Lecture Notes in Computer Science, 819: 1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Godfrey:1994:ACD

[1656] P. Godfrey, J. Minker, and L. Novik. An architecture for a cooperative database system. Lecture Notes in Computer Science, 819:3–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Samet:1994:SD

[1657] H. Samet. Spatial databases. Lecture Notes in Computer Science, 819:25–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buchmann:1994:DMR

[1658] A. Buchmann. Database management for real-time applications. Lecture Notes in Computer Science, 819:26–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zloof:1994:EUI

[1659] M. Zloof and R. Krishnamurthy. The evolution of user interface tools for database applications. Lecture Notes in Computer Science, 819:27–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sripada:1994:TDT

[1660] S. M. Sripada, B. L. Rosser, J. M. Bedford, and R. A. Kowalski. Temporal database technology for air traffic flow management. Lecture Notes in Computer Science, 819:28–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johansson:1994:UEM

[1661] O. Johansson. Using an extended ER-Model based data dictionary to automatically generate product modeling systems. Lecture Notes in Computer Science, 819:42–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehlmann:1994:NDA

[1662] B. K. Ehlmann and G. A. Riccardi.

A notation for describing aggregate relationships in an object-oriented data model. Lecture Notes in Computer Science, 819:62–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ketabchi:1994:MAA

[1663] M. Ketabchi, N. Sehgal, S. Dasananda, and R. Mikkilineni. Modeling activities of application domains. Lecture Notes in Computer Science, 819:78–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1994:EIR

[1664] H. Lu, B.-C. Ooi, and Kian Lee Tan.

Efficient image retrieval by color contents. Lecture Notes in Computer Science, 819:95–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1994:SMD

[1665] J. K. Wu, B. M. Mehtre, Y. J. Gao, and P. C. Lam. STAR — A multimedia database system for trademark registration. Lecture Notes in Computer Science, 819:109–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Consens:1994:VQD

[1666] M. P. Consens, M. Z. Hasan, and A. O. Mendelzon. Visualizing and querying distributed event traces with Hy0+. Lecture Notes in Computer Science, 819:123–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arikawa:1994:DMC

[1667] M. Arikawa, H. Kawakita, and Y. Kambayashi. Dynamic maps as composite views of varied geographic database servers. Lecture Notes in Computer Science, 819:142–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kramer:1994:LDE

[1668] R. Kramer and H. Spandl. Limits database for an environmental information system — A fuzzy set-based querying approach. Lecture Notes in Computer Science, 819:158–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baralis:1994:IDS

[1669] E. Baralis, S. Ceri, S. Paraboschi, and G. Monteleone. An intelligent database system application: The design of EMS. Lecture Notes in Computer Science, 819:172–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chandra:1994:UNG

[1670] R. Chandra and A. Segev. Using next generation databases to develop financial applications. Lecture Notes in Computer Science, 819:190–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cochinwala:1994:MSF

[1671] M. Cochinwala, J. Bradley, R. Tanamy, and R. Subramanian. A multidatabase solution for a financial application. Lecture Notes in Computer Science, 819: 204–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orsborn:1994:ANG

[1672] K. Orsborn. Applying next generation object-oriented DBMS to finite element analysis. Lecture Notes in Computer Science, 819:215–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeSmedt:1994:PWA

[1673] P. De Smedt, J. Annevelink, T. Pham, and P. Strong. A physician’s workstation as an application of objectoriented database technology in healthcare. Lecture Notes in Computer Science, 819:234–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Persin:1994:FDR

[1674] M. Persin, J. Zobel, and R. Sacks-Davis. Fast document ranking for large scale information retrieval. Lecture Notes in Computer Science, 819:253– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blake:1994:TRD

[1675] G. E. Blake, M. P. Consens, P. Kilpelaeinen, and P.-A. Larson. Text/ relational database management systems: Harmonizing SQL and SGML. Lecture Notes in Computer Science, 819:267–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thimm:1994:DBM

[1676] H. Thimm and T. C. Rakow. A DBMSbased multimedia archiving teleservice incorporating mail. Lecture Notes in Computer Science, 819:281–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carlis:1994:ZDB

[1677] J. Carlis, J. Riedl, A. Georgopoulos, and G. Wilcox. A zoomable DBMS for brain structure, function and behavior. Lecture Notes in Computer Science, 819:299–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kemp:1994:CCD

[1678] G. J. L. Kemp, Z. Jiao, P. M. D. Gray, and J. E. Fothergill. Combining computation with database access in biomolecular computing. Lecture Notes in Computer Science, 819:317– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakanishi:1994:OOD

[1679] M. Nakanishi, A. Hashimoto, and M. Ito. An object-oriented database of protein structure data. Lecture Notes in Computer Science, 819:336– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Appelrath:1994:ADT

[1680] H.-J. Appelrath, H. Behrends, H. Jasper, and V. Kamp. Active database technology supports cancer clustering. Lecture Notes in Computer Science, 819:351–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matthes:1994:GDI

[1681] F. Matthes, A. Rudloff, J. W. Schmidt, and K. Subieta. A gateway from DBPL to Ingres. Lecture Notes in Computer Science, 819:365–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radeke:1994:FSR

[1682] E. Radeke and M. H. Scholl. Federation and stepwise reduction of database systems. Lecture Notes in Computer Science, 819:381–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aebi:1994:MTD

[1683] D. Aebi and R. Largo. Methods and tools for data value re-engineering. Lecture Notes in Computer Science, 819: 400–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Durand:1994:OVB

[1684] J. Durand, M. Ganti, and R. Salinas. Object view broker: A mediation service and architecture to provide object-oriented views of heterogeneous databases. Lecture Notes in Computer Science, 819:415–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rastogi:1994:EPE

[1685] R. Rastogi and M.-A. Neimat. Enhancing pre-existing data managers with atomicity and durability. Lecture Notes in Computer Science, 819:429– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paradinas:1994:PPD

[1686] P. Paradinas and J.-J. Vandewalle. A personal and portabel database server: The CQL card. Lecture Notes in Computer Science, 819:444–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gallersdoerfer:1994:IND

[1687] R. Gallersdoerfer, M. Jarke, and K. Klabunde. Intelligent networks as a data intensive application (IN-DIA). Lecture Notes in Computer Science, 819:458–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pudlak:1994:UUB

[1688] P. Pudlak. Unexpected upper bounds on the complexity of some communication games. Lecture Notes in Computer Science, 820:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernau:1994:VUL

[1689] H. Fernau and L. Staiger. Valuations and unambiguity of languages, with applications to fractal geometry. Lecture Notes in Computer Science, 820:11–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Siegelmann:1994:CPP

[1690] H. T. Siegelmann. On the computational power of probabilistic and faulty neural networks. Lecture Notes in Computer Science, 820:23–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cerans:1994:DPI

[1691] K. Cerans. Deciding properties of integral relational automata. Lecture Notes in Computer Science, 820:35–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aumann:1994:CRT

[1692] Y. Aumann, J. Bar-Ilan, and U. Feige.

On the cost of recomputing: Tight bounds on pebbling with faults. Lecture Notes in Computer Science, 820:47–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asarin:1994:SRB

[1693] E. Asarin and O. Maler. On some relations between dynamical systems and transition systems. Lecture Notes in Computer Science, 820:59–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Globerman:1994:CRM

[1694] N. Globerman and D. Harel. Complexity results for multi-pebble automata and their logics. Lecture Notes in Computer Science, 820:73–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kanellakis:1994:ACM

[1695] P. C. Kanellakis, G. G. Hillebrand, and H. G. Mairson. An analysis of the Core-ML language: Expressive power and type reconstruction. Lecture Notes in Computer Science, 820:83–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gyssens:1994:EES

[1696] M. Gyssens, J. Van den Bussche, and D. Van Gucht. Expressiveness of efficient semi-deterministic choice constructs. Lecture Notes in Computer Science, 820:106–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graedel:1994:TRC

[1697] E. Graedel and Y. Gurevich. Tailoring recursion for complexity. Lecture Notes in Computer Science, 820:118– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klarlund:1994:DAA

[1698] N. Klarlund, M. Mukund, and M. Sohoni. Determinizing asynchronous automata. Lecture Notes in Computer Science, 820:130–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muscholl:1994:CBA

[1699] A. Muscholl. On the complementation of buechi asynchronous cellular automata. Lecture Notes in Computer Science, 820:142–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corradini:1994:DLC

[1700] F. Corradini and R. De Nicola. Distribution and locality of concurrent systems. Lecture Notes in Computer Science, 820:154–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gawlick:1994:LTU

[1701] R. Gawlick, R. Segala, J. Soegaard-Andersen, and N. Lynch. Liveness in timed and untimed systems. Lecture Notes in Computer Science, 820:166– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanchez-Couso:1994:ACA

[1702] J. R. Sanchez-Couso and M. I.

Fernandez-Camacho. Average-case analysis of pattern-matching in trees under the BST probability model. Lecture Notes in Computer Science, 820: 178–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jiang:1994:ASC

[1703] T. Jiang and M. Li. On the approximation of shortest common supersequences and longest common subsequences. Lecture Notes in Computer Science, 820:191–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hariharan:1994:OPA

[1704] R. Hariharan and S. Muthukrishnan.

Optimal parallel algorithms for prefix matching. Lecture Notes in Computer Science, 820:203–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amir:1994:OTD

[1705] A. Amir, G. Benson, and M. Farach. Optimal two-dimensional compressed matching. Lecture Notes in Computer Science, 820:215–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Italiano:1994:MST

[1706] G. F. Italiano and R. Ramaswami. Maintaining spanning trees of small diameter. Lecture Notes in Computer Science, 820:227–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gil:1994:SFP

[1707] J. Gil and Y. Matias. Simple fast parallel hashing. Lecture Notes in Computer Science, 820:239–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Przytycka:1994:OAT

[1708] T. M. Przytycka and L. L. Larmore. The optimal alphabetic tree problem revisited. Lecture Notes in Computer Science, 820:251–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buhrman:1994:CER

[1709] H. Buhrman and L. Torenvliet. On the cutting edge of relativization: The resource bounded injury method. Lecture Notes in Computer Science, 820: 263–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Birget:1994:PCC

[1710] J.-C. Birget, S. Margolis, J. Meakin, and P. Weil. PSPACE-completeness of certain algorithmic problems on the subgroups of free groups. Lecture Notes in Computer Science, 820:274– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hennessy:1994:HOP

[1711] M. Hennessy. Higher-order processes and their models. Lecture Notes in Computer Science, 820:286–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vergauwen:1994:ELC

[1712] B. Vergauwen and J. Lewi. Efficient local correctness checking for single and alternating Boolean equation systems. Lecture Notes in Computer Science, 820:304–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdulla:1994:UVP

[1713] P. A. Abdulla and B. Jonsson. Undecidable verification problems for programs with unreliable channels. Lecture Notes in Computer Science, 820: 316–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fix:1994:RAP

[1714] L. Fix and F. B. Schneider. Reasoning about programs by exploiting the environment. Lecture Notes in Computer Science, 820:328–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braeuner:1994:MIA

[1715] T. Braeuner. A model of intuitionistic affine logic from stable domain theory. Lecture Notes in Computer Science, 820:340–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plotkin:1994:BBL

[1716] G. Plotkin and G. Winskel. Bistructures, bidomains and linear logic. Lecture Notes in Computer Science, 820: 352–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aziz:1994:EFK

[1717] A. Aziz, V. Singhal, F. Balarin, and R. K. Brayton. Equivalences for fair Kripke structures. Lecture Notes in Computer Science, 820:364–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanBreugel:1994:GFC

[1718] F. Van Breugel. Generalizing finiteness conditions of labelled transition systems. Lecture Notes in Computer Science, 820:376–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Droste:1994:KTR

[1719] M. Droste. A Kleene theorem for recognizable languages over concurrency monoids. Lecture Notes in Computer Science, 820:388–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seidl:1994:LSE

[1720] H. Seidl. Least solutions of equations over N. Lecture Notes in Computer Science, 820:400–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hotz:1994:FUA

[1721] G. Hotz and G. Pitsch. Fast uniform analysis of coupled-context-free languages. Lecture Notes in Computer Science, 820:412–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pin:1994:PCG

[1722] J.-E. Pin. Polynomial closure of group languages and open sets of the hall topology. Lecture Notes in Computer Science, 820:424–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caron:1994:PCS

[1723] A.-C. Caron, H. Comon, J.-L. Coquide, and M. Dauchet. Pumping, cleaning and symbolic constraints solving. Lecture Notes in Computer Science, 820:436–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hintermeier:1994:DTC

[1724] C. Hintermeier, C. Kirchner, and H. Kirchner. Dynamically-typed computations for order-sorted equational presentations. Lecture Notes in Computer Science, 820:450–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiCosmo:1994:CFO

[1725] R. Di Cosmo and D. Kesner. Combining first order algebraic rewriting systems, recursion and extensional lambda calculi. Lecture Notes in Computer Science, 820:462–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Upfal:1994:TIN

[1726] E. Upfal. On the theory of interconnection networks for parallel computers. Lecture Notes in Computer Science, 820:473–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garg:1994:MCD

[1727] N. Garg, V. V. Vazirani, and M. Yannakakis. Multiway cuts in directed and node weighted graphs. Lecture Notes in Computer Science, 820:487–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feige:1994:FRL

[1728] U. Feige. A fast randomized LOGSPACE algorithm for graph connectivity. Lecture Notes in Computer Science, 820:499–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nikoletseas:1994:SVD

[1729] S. Nikoletseas, K. Palem, P. Spirakis, and M. Yung. Short vertex disjoint paths and multiconnectivity in random graphs: Reliable network computing. Lecture Notes in Computer Science, 820:508–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lagergren:1994:SI

[1730] J. Lagergren. The size of an intertwine. Lecture Notes in Computer Science, 820:520–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuster:1994:FEC

[1731] R. Yuster and U. Zwick. Finding even cycles even faster. Lecture Notes in Computer Science, 820:532–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoefting:1994:PTA

[1732] F. Hoefting and E. Wanke. Polynomial time analysis of toroidal periodic graphs. Lecture Notes in Computer Science, 820:544–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonizzoni:1994:TLB

[1733] P. Bonizzoni. A tight lower bound for primitivity in k-structures. Lecture Notes in Computer Science, 820:556– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blundo:1994:RDP

[1734] C. Blundo, A. De Santis, and U. Vaccaro. Randomness in distribution protocols. Lecture Notes in Computer Science, 820:568–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freivalds:1994:LSB

[1735] R. Freivalds and M. Karpinski. Lower space bounds for randomized computation. Lecture Notes in Computer Science, 820:580–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jakoby:1994:ACC

[1736] A. Jakoby, R. Reischuk, C. Schindelhauer, and S. Weis. The average case complexity of the parallel prefix problem. Lecture Notes in Computer Science, 820:593–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golin:1994:PCE

[1737] M. J. Golin and N. Young. Prefix codes: Equiprobable words, unequal letter costs. Lecture Notes in Computer Science, 820:605–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plaxton:1994:SLL

[1738] C. G. Plaxton and T. Suel. A superlogarithmic lower bound for hypercubic sorting networks. Lecture Notes in Computer Science, 820:618–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mei:1994:ESR

[1739] A. Mei and Y. Igarashi. Efficient strategies for robot navigation in unknown environment. Lecture Notes in Computer Science, 820:630–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steels:1994:BO

[1740] Luc Steels. Beyond objects. Lecture Notes in Computer Science, 821: 1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/0821/08210001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0821/08210001.
pdf.

Huersch:1994:SSA

[1741] W. L. Huersch. Should superclasses be abstract? Lecture Notes in Computer Science, 821:12–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hursch:1994:SSA

[1742] Walter L. Hürsch. Should superclasses be abstract? Lecture Notes in Computer Science, 821:12– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0821/08210012.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0821/08210012.
pdf.

Wieringa:1994:RDS

[1743] Roel Wieringa, Wiebren de Jonge, and Paul Spruit. Roles and dynamic subclasses: A modal logic approach. Lecture Notes in Computer Science, 821: 32–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0821/08210032.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0821/08210032.
pdf.

Lamping:1994:MA

[1744] John Lamping and Martin Abadi.

Methods as assertions. Lecture Notes in Computer Science, 821:60– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0821/08210060.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0821/08210060.
pdf.

Lopes:1994:APF

[1745] Cristina Videira Lopes and Karl J. Lieberherr. Abstracting processto-function relations in concurrent object-oriented applications. Lecture Notes in Computer Science, 821:81– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0821/08210081.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0821/08210081.
pdf.

Vasconcelos:1994:TCO

[1746] Vasco T. Vasconcelos. Typed concurrent objects. Lecture Notes in Computer Science, 821:100–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0821/08210100.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0821/08210100.
pdf.

Guerraoui:1994:AOC

[1747] Rachid Guerraoui. Atomic object composition. Lecture Notes in Computer Science, 821:118–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0821/08210118.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0821/08210118.
pdf.

Beck:1994:PGA

[1748] Kent Beck and Ralph Johnson. Patterns generate architectures. Lecture Notes in Computer Science, 821:139– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0821/08210139.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0821/08210139.
pdf.

Pree:1994:MPM

[1749] Wolfgang Pree. Meta patterns — A means for capturing the essentials of reusable object-oriented design. Lecture Notes in Computer Science, 821: 150–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0821/08210150.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0821/08210150.
pdf.

DePauw:1994:MOO

[1750] W. De Pauw, D. Kimelman, and J. Vlissides. Modeling object-oriented program execution. Lecture Notes in Computer Science, 821:163–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pauw:1994:MOO

[1751] Wim De Pauw, Doug Kimelman, and John Vlissides. Modeling objectoriented program execution. Lecture Notes in Computer Science, 821:163– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0821/08210163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0821/08210163.
pdf.

Streitz:1994:POW

[1752] Norbert A. Streitz. Putting objects to work: Hypermedia as the subject matter and the medium for computersupported cooperative work. Lecture Notes in Computer Science, 821:183– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0821/08210183.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0821/08210183.
pdf.

Omicini:1994:OOC

[1753] Andrea Omicini and Antonio Natali. Object-oriented computations in logic programming. Lecture Notes in Computer Science, 821:194–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0821/08210194.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0821/08210194.
pdf.

Bertino:1994:DOD

[1754] Elisa Bertino, Giovanna Guerrini, and Danilo Montesi. Deductive object databases. Lecture Notes in Computer Science, 821:213–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0821/08210213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0821/08210213.
pdf.

Alagic:1994:DOO

[1755] S. Alagic, R. Sunderraman, and R. Bagai. Declarative object-oriented programming: Inheritance, subtyping and prototyping. Lecture Notes in Computer Science, 821:236–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0821/08210236.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0821/08210236.
pdf.

Lopez:1994:COI

[1756] Gus Lopez, Bjørn Freeman-Benson, and Alan Borning. Constraints and object identity. Lecture Notes in Computer Science, 821:260–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0821/08210260.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0821/08210260.
pdf.

Hagimont:1994:PGO

[1757] Daniel Hagimont. Protection in the guide object-oriented distributed system. Lecture Notes in Computer Science, 821:280–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/0821/08210280.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        0821/08210280.pdf.

Okamura:1994:OLC

[1758] Hideaki Okamura and Yutaka Ishikawa. Object location control using metalevel programming. Lecture Notes in Computer Science, 821:299–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0821/08210299.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0821/08210299.
                                        pdf.

Attardi:1994:COA

[1759] Giuseppe Attardi and Tito Flagella. Customising object allocation. Lecture Notes in Computer Science, 821:320– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0821/08210320.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0821/08210320.
                                        pdf.

Moreira:1994:COO

[1760] Ana M. D. Moreira and Robert G. Clark. Combining object-oriented analysis and formal description techniques. Lecture Notes in Computer Science, 821:344–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0821/08210344.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0821/08210344.
pdf.

Briggs:1994:SLO

[1761] Ted L. Briggs and John Werth. A specification language for object- oriented analysis and design. Lecture Notes in Computer Science, 821:365– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0821/08210365.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0821/08210365.
pdf.

Aksit:1994:RTS

[1762] Mehmet Aksit, Jan Bosch, William van der Sterren, and Lodewijk Bergmans. Real-time specification inheritance anomalies and real-time filters. Lecture Notes in Computer Science, 821:386–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0821/08210386.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0821/08210386.pdf.

Chen:1994:EDL

[1763] Weimin Chen, Volker Turau, and Wolfgang Klas. Efficient dynamic look-up strategy for multi-methods. Lecture Notes in Computer Science, 821:408– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0821/08210408.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0821/08210408.
                                        pdf.

Vitek:1994:TMP

[1764] Jan Vitek and R. Nigel Horspool. Taming message passing: Efficient method look-up for dynamically typed languages. Lecture Notes in Computer Science, 821:432–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0821/08210432.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0821/08210432.
                                        pdf.

Nayeri:1994:GDD

[1765] Farshad Nayeri, Ben Hurwitz, and Frank Manola. Generalizing dispatching in a distributed object system. Lecture Notes in Computer Science, 821: 450–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0821/08210450.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0821/08210450.
                                        pdf.

Bruegge:1994:GEM

[1766] Bernd Bruegge and Erik Riedel. A geographic environmental modeling system: Towards an objectoriented framework. Lecture Notes in Computer Science, 821:474–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0821/08210474.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0821/08210474.
pdf.

Freeman:1994:ADV

[1767] S. M. G. Freeman and M. S. Manasse. Adding digital video to an objectoriented user interface toolkit. Lecture Notes in Computer Science, 821:493– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0821/08210493.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0821/08210493.
pdf.

Peltonen:1994:PCA

[1768] Hannu Peltonen, Tomi Männistö, Kari Alho, and Reijo Sulonen. Product configurations — an application for prototype object approach. Lecture Notes in Computer Science, 821:513– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0821/08210513.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0821/08210513.
pdf.

Anonymous:1994:AIc

[1769] Anonymous. Author index. Lecture Notes in Computer Science, 821:535– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/0821/0821ind.pdf.

Felty:1994:GRT

[1770] A. Felty and D. Howe. Generalization and reuse of tactic proofs. Lecture Notes in Computer Science, 822:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giunchiglia:1994:PTL

[1771] F. Giunchiglia and P. Traverso. Program tactics and logic tactics. Lecture Notes in Computer Science, 822:16–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeGroote:1994:RBL

[1772] P. De Groote. On the relation between the lambda-calculus and the syntactic theory of sequential control. Lecture Notes in Computer Science, 822:31–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tupailo:1994:PTA

[1773] S. Tupailo. On a proof-theoretical analysis of 011-AC, 011-DC and cap delta011-CA. Lecture Notes in Computer Science, 822:44–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monroy:1994:PPC

[1774] R. Monroy, A. Bundy, and A. Ireland. Proof plans for the correction of false conjectures. Lecture Notes in Computer Science, 822:54–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Egly:1994:VA

[1775] U. Egly. On the value of antiprenexing. Lecture Notes in Computer Science, 822:69–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vandecasteele:1994:IFD

[1776] H. Vandecasteele and D. De Schreye.

Implementing a finite-domain CLP-Language on top of Prolog: A transformational approach. Lecture Notes in Computer Science, 822:84–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duggan:1994:LC

[1777] D. Duggan. Logical closures. Lecture Notes in Computer Science, 822:114– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goubault:1994:HOR

[1778] J. Goubault. Higher-order rigid Eunification. Lecture Notes in Computer Science, 822:129–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1994:PEL

[1779] P. Anderson. Program extraction in a logical framework setting. Lecture Notes in Computer Science, 822:144– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Despeyroux:1994:HOA

[1780] J. Despeyroux and A. Hirschowitz. Higher-order abstract syntax with induction in coq. Lecture Notes in Computer Science, 822:159–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruening:1994:TEC

[1781] S. Bruening. Towards efficient calculi for resource-oriented deductive planning. Lecture Notes in Computer Science, 822:174–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quaresma:1994:LPF

[1782] P. Quaresma and J. G. Lopes. A logic programming framework for the abductive inference of intentions in cooperative dialogues. Lecture Notes in Computer Science, 822:189–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Darlington:1994:CLP

[1783] J. Darlington and Y. Guo. Constraint logic programming in the sequent calculus. Lecture Notes in Computer Science, 822:200–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Avenhaus:1994:CRS

[1784] J. Avenhaus and C. Loria-Saenz. On conditional rewrite systems with extra variables and deterministic logic programs. Lecture Notes in Computer Science, 822:215–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papp:1994:BRW

[1785] C. Papp. A bottom-up reconstruction of the well-founded semantics for disjunctive logic programs. Lecture Notes in Computer Science, 822:230– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seipel:1994:ECE

[1786] D. Seipel. An efficient computation of the extended generalized closed world assumption by supportfor-negation sets. Lecture Notes in Computer Science, 822:245–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1994:MSR

[1787] D. A. Smith and T. J. Hickey. Multi-SLD resolution. Lecture Notes in Computer Science, 822:260–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beckert:1994:AL

[1788] B. Beckert, R. Haehnle, A. Ramesh, and N. V. Murray. On anti-links. Lecture Notes in Computer Science, 822: 275–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1994:GDD

[1789] L. Lu. A generic declarative diagnoser for normal logic programs. Lecture Notes in Computer Science, 822:290– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Codish:1994:GDV

[1790] M. Codish, M. Garcia de la Banda, M. Bruynooghe, and M. Hermenegildo. Goal dependent vs. goal independent analysis of logic programs. Lecture Notes in Computer Science, 822:305– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mathieu:1994:KAP

[1791] P. Mathieu and J. P. Delahaye. A kind of achievement by parts method. Lecture Notes in Computer Science, 822: 320–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duan:1994:PTL

[1792] Z. Duan, M. Koutny, and C. Holt. Projection in temporal logic programming. Lecture Notes in Computer Science, 822:333–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pollacia:1994:OFM

[1793] L. F. Pollacia and L. M. L. Delcambre. The object flow model: A formal framework for describing the dynamic construction, destruction and interaction of complex objects. Lecture Notes in Computer Science, 823:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dijkstra:1994:COV

[1794] J. Dijkstra. On complex objects and versioning in complex environments. Lecture Notes in Computer Science, 823:13–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morsi:1994:ASP

[1795] M. M. A. Morsi and S. B. Navathe. Application and system prototyping via an extensible object-oriented environment. Lecture Notes in Computer Science, 823:24–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peters:1994:RUB

[1796] R. J. Peters and M. Tamer Oezsu. Reflection in a uniform behavioral object model. Lecture Notes in Computer Science, 823:34–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Behm:1994:RCD

[1797] J. B. Behm and T. J. Teorey. Relative constraints in ER data models. Lecture Notes in Computer Science, 823:46–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Winter:1994:DID

[1798] R. Winter. Design and implementation of derived entities. Lecture Notes in Computer Science, 823:60–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauh:1994:SCS

[1799] O. Rauh and E. Stickel. Searching for compositions in ER schemes. Lecture Notes in Computer Science, 823:74–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zamperoni:1994:EQC

[1800] A. Zamperoni and P. Loehr-Richter.

Enhancing the quality of conceptual database specifications through validation. Lecture Notes in Computer Science, 823:85–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wijsen:1994:FDG

[1801] J. Wijsen, J. Vandenbulcke, and H. Olivie. Functional dependencies generalized for temporal databases that include object-identity. Lecture Notes in Computer Science, 823:99–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goralwalla:1994:TEU

[1802] I. A. Goralwalla and M. Tamer Oezsu. Temporal extensions to a uniform behavioral object model. Lecture Notes in Computer Science, 823:110–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rose:1994:TTO

[1803] E. Rose and A. Segev. TOOSQL — A temporal object-oriented query language. Lecture Notes in Computer Science, 823:122–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roddick:1994:TSV

[1804] J. F. Roddick, N. G. Craske, and T. J. Richards. A taxonomy for schema versioning based on the relational and entity relationship models. Lecture Notes in Computer Science, 823:137– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yen:1994:NCQ

[1805] S.-J. Yen and A. L. P. Chen. Neighborhood/conceptual query answering with imprecise/incomplete data. Lecture Notes in Computer Science, 823: 149–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pernul:1994:ERM

[1806] G. Pernul, W. Winiwarter, and A. Min Tjoa. The entity-relationship model for multilevel security. Lecture Notes in Computer Science, 823:166–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garzotto:1994:HEA

[1807] F. Garzotto, L. Mainetti, and P. Paolini. HDM2: Extending the ER approach to hypermedia application design. Lecture Notes in Computer Science, 823:178–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Metais:1994:DSD

[1808] E. Metais, J.-N. Meunier, and G. Levreau. Database schema design: A perspective from natural language techniques to validation and view integration. Lecture Notes in Computer Science, 823:190–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MinTjoa:1994:TRS

[1809] A. Min Tjoa and L. Berger. Transformation of requirement specifications expressed in natural language into an EER model. Lecture Notes in Computer Science, 823:206–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Storey:1994:CRF

[1810] V. C. Storey, R. C. Goldstein, R. H. L. Chiang, and D. Dey. A commonsense reasoning facility based on the entityrelationship model. Lecture Notes in Computer Science, 823:218–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Falkenberg:1994:DDE

[1811] E. D. Falkenberg. DETERM: Deterministic event-tuned entityrelationship modeling. Lecture Notes in Computer Science, 823:230–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laender:1994:SCM

[1812] A. H. F. Laender and D. J. Flynn. A semantic comparison of the modeling capabilities of the ER and NIAM models. Lecture Notes in Computer Science, 823:242–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VazVelho:1994:ERM

[1813] A. Vaz Velho and R. Carapuca. From entity-relationship models to role-attribute models. Lecture Notes in Computer Science, 823:257–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johannesson:1994:UCG

[1814] P. Johannesson. Using conceptual graph theory to support schema integration. Lecture Notes in Computer Science, 823:283–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koh:1994:IHO

[1815] J.-L. Koh and A. L. P. Chen. Integration of heterogeneous object schemas. Lecture Notes in Computer Science, 823:297–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steele:1994:RMM

[1816] P. M. Steele and A. B. Zaslavsky.

The role of meta models in federating system modeling techniques. Lecture Notes in Computer Science, 823:315– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santucci:1994:MSI

[1817] G. Santucci, C. Batini, and G. Di Battista. Multilevel schema integration. Lecture Notes in Computer Science, 823:327–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castano:1994:ROOa

[1818] S. Castano, V. De Antonellis, and P. San Pietro. Reuse of object-oriented requirement specifications. Lecture Notes in Computer Science, 823:339– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chiang:1994:PER

[1819] R. H. L. Chiang, T. M. Barron, and V. C. Storey. Performance evaluation of reverse engineering relational databases into extended entityrelationship models. Lecture Notes in Computer Science, 823:352–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hainaut:1994:TBD

[1820] J.-L. Hainaut, C. Tonneau, M. Joris, and M. Chandelon. Transformationbased database reverse engineering. Lecture Notes in Computer Science, 823:364–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gogolla:1994:IAO

[1821] M. Gogolla, R. Herzig, S. Conrad, and G. Denker. Integrating the ER approach in an OO environment. Lecture Notes in Computer Science, 823: 376–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Norrie:1994:EER

[1822] M. C. Norrie. An extended entityrelationship approach to data management in object-oriented systems. Lecture Notes in Computer Science, 823: 390–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Narasimhan:1994:MRM

[1823] B. Narasimhan, S. B. Navathe, and S. Jayaraman. On mapping ER and relational models into OO schemas. Lecture Notes in Computer Science, 823: 402–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hurley:1994:RMM

[1824] W. D. Hurley. A repository meta model for interactive systems. Lecture Notes in Computer Science, 823:414– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rosengren:1994:BIR

[1825] P. Rosengren, U. Wingstedt, M. Bern, and P. Kool. ER-Based information retrieval in a mixed database environment. Lecture Notes in Computer Science, 823:426–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akoka:1994:FAC

[1826] J. Akoka and I. Comyn-Wattiau. A framework for automatic clustering of semantic models. Lecture Notes in Computer Science, 823:438–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jaeschke:1994:EMC

[1827] P. Jaeschke, A. Oberweis, and W. Stucky. Extending ER model clustering by relationship clustering. Lecture Notes in Computer Science, 823: 451–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeCastro:1994:SIM

[1828] C. De Castro, F. Grandi, and M. R.

Scalas. Semantic interoperability of multitemporal relational databases. Lecture Notes in Computer Science, 823:463–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1994:MTA

[1829] S.-K. Kim and S. Chakravarthy. Modeling time: Adequacy of three distinct time concepts for temporal databases. Lecture Notes in Computer Science, 823:475–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chudziak:1994:TUL

[1830] J. A. Chudziak, H. Rybinsky, and J. Vorbach. Towards a unifying logic formalism for semantic data models. Lecture Notes in Computer Science, 823:492–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fong:1994:KBA

[1831] J. Fong and M. Ho. Knowledgebased approach for abstracting hierarchical and network schema semantics. Lecture Notes in Computer Science, 823:508–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castilho:1994:SSA

[1832] J. M. Castilho. A state-space approach for database redesign. Lecture Notes in Computer Science, 823:520–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agarwal:1994:CDO

[1833] P. K. Agarwal, M. J. Katz, and M. Sharir. Computing depth orders and related problems. Lecture Notes in Computer Science, 824:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agarwal:1994:SMM

[1834] P. K. Agarwal and S. Sen. Selection in monotone matrices and computing k0t0h nearest neighbors. Lecture Notes in Computer Science, 824:13–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albers:1994:NLA

[1835] S. Albers and H. Koga. New online algorithms for the page replication problem. Lecture Notes in Computer Science, 824:25–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ausiello:1994:SRL

[1836] G. Ausiello, E. Feuerstein, S. Leonardi, and L. Stougie. Serving requests with on-line routing. Lecture Notes in Computer Science, 824:37–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becker:1994:NAC

[1837] P. Becker. A new algorithm for the construction of optimal B-trees. Lecture Notes in Computer Science, 824:49–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBerg:1994:NRB

[1838] M. De Berg, M. De Groot, and M. Overmars. New results on binary space partitions in the plane. Lecture Notes in Computer Science, 824:61–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berman:1994:NOP

[1839] P. Berman and A. Lingas. A nearly optimal parallel algorithm for the Voronoi diagram of a convex polygon. Lecture Notes in Computer Science, 824:73–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biedl:1994:TPG

[1840] T. Biedl, G. Kant, and M. Kaufmann. On triangulating planar graphs under the four-connectivity constraint. Lecture Notes in Computer Science, 824: 83–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Czumaj:1994:PSA

[1841] A. Czumaj, L. Gasieniec, M. Piotrow, and W. Rytter. Parallel and sequential approximation of shortest superstrings. Lecture Notes in Computer Science, 824:95–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Czyzowicz:1994:STP

[1842] J. Czyzowicz, H. Everett, and J.-M. Robert. Separating translates in the plane: Combinatorial bounds and an algorithm. Lecture Notes in Computer Science, 824:107–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Das:1994:FAW

[1843] G. Das, P. J. Heffernan, and G. Narasimhan. Finding all weaklyvisible chords of a polygon in linear time. Lecture Notes in Computer Science, 824:119–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dietz:1994:TLB

[1844] P. F. Dietz, J. I. Seiferas, and J. Zhang. A tight lower bound for on-line monotonic list labeling. Lecture Notes in Computer Science, 824:131–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Felsner:1994:TGG

[1845] S. Felsner, R. Mueller, and L. Wernisch.

Trapezoid graphs and generalizations, geometry and algorithms. Lecture Notes in Computer Science, 824:143– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez-Baca:1994:OPS

[1846] D. Fernandez-Baca and G. Slutzki. Optimal parametric search on graphs of bounded tree-width. Lecture Notes in Computer Science, 824:155–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fredman:1994:LBD

[1847] M. L. Fredman. Lower bounds for dynamic algorithms. Lecture Notes in Computer Science, 824:167–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1994:SPA

[1848] A. Gupta and N. Nishimura. Sequential and parallel algorithms for embedding problems on classes of partial k-trees. Lecture Notes in Computer Science, 824:172–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1994:ISP

[1849] P. Gupta, R. Janardan, and M. Smid.

On intersection searching problems involving curved objects. Lecture Notes in Computer Science, 824:183– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halldorsson:1994:IAI

[1850] M. M. Halldorsson and J. Radhakrishnan. Improved approximations of independent sets in bounded-degree graphs. Lecture Notes in Computer Science, 824:195–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Higham:1994:AOE

[1851] L. Higham and T. Przytycka. Asymptotically optimal election on weighted rings. Lecture Notes in Computer Science, 824:207–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hromkovic:1994:OAB

[1852] J. Hromkovic, R. Klasing, W. Unger, and H. Wagener. Optimal algorithms for broadcast and gossip in the edgedisjoint path modes. Lecture Notes in Computer Science, 824:219–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haastad:1994:RRH

[1853] J. Haastad. Recent results in hardness of approximation. Lecture Notes in Computer Science, 824:231–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juurlink:1994:PHM

[1854] B. H. H. Juurlink and H. A. G. Wijshoff. The parallel hierarchical memory model. Lecture Notes in Computer Science, 824:240–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mulmuley:1994:RGA

[1855] K. Mulmuley. Randomized geometric algorithms. Lecture Notes in Computer Science, 824:252–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palios:1994:CMN

[1856] L. Palios. Connecting the maximum number of grid nodes to the boundary with non-intersecting line segments. Lecture Notes in Computer Science, 824:255–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papatriantafilou:1994:SSW

[1857] M. Papatriantafilou and P. Tsigas. On self-stabilizing wait-free clock synchronization. Lecture Notes in Computer Science, 824:267–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peinado:1994:HGR

[1858] M. Peinado. Hard graphs for randomized subgraph exclusion algorithms. Lecture Notes in Computer Science, 824:278–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Phillips:1994:TSN

[1859] C. Phillips, C. Stein, and J. Wein.

Task scheduling in networks. Lecture Notes in Computer Science, 824:290– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schenk:1994:PDL

[1860] E. Schenk. Parallel dynamic lowest common ancestors. Lecture Notes in Computer Science, 824:302–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuierer:1994:ACK

[1861] S. Schuierer. An O(log log n) algorithm to compute the kernel of a polygon. Lecture Notes in Computer Science, 824:314–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuierer:1994:CLD

[1862] S. Schuierer. Computing the L1diameter and center of a simple rectilinear polygon in parallel. Lecture Notes in Computer Science, 824:326– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sibeyn:1994:ELL

[1863] J. F. Sibeyn and T. Harris. Exploiting locality in LT-RAM computations. Lecture Notes in Computer Science, 824:338–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thorup:1994:EPS

[1864] M. Thorup. Efficient preprocessing of simple binary pattern forests. Lecture Notes in Computer Science, 824:350– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1994:PAE

[1865] X. Zhou, S.-I. Nakano, and T. Nishizeki. A parallel algorithm for edge-coloring partial k-trees. Lecture Notes in Computer Science, 824:359– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dragan:1994:DCD

[1866] F. F. Dragan. Dominating cliques in distance-hereditary graphs. Lecture Notes in Computer Science, 824:370– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faugeras:1994:CMF

[1867] O. Faugeras. Cartan’s moving frame method and its application to the geometry and evolution of curves in the euclidean, affine and projective planes. Lecture Notes in Computer Science, 825:11–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barrett:1994:RTD

[1868] E. Barrett, G. Gheen, and P. Payton.

Representation of three-dimensional object structure as cross-ratios of determinants of stereo image points. Lecture Notes in Computer Science, 825:47–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zisserman:1994:CAE

[1869] A. Zisserman and S. J. Maybank. A case against epipolar geometry. Lecture Notes in Computer Science, 825:69–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mundy:1994:RSI

[1870] J. L. Mundy and A. Zisserman. Repeated structures: Image correspondence constraints and 3D structure recovery. Lecture Notes in Computer Science, 825:89–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gros:1994:HUC

[1871] P. Gros. How to use the cross ratio to compute projective invariants from two images. Lecture Notes in Computer Science, 825:107–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shashua:1994:GAA

[1872] A. Shashua. On geometric and algebraic aspects of 3D affine and projective structures from perspective 2D views. Lecture Notes in Computer Science, 825:127–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carlsson:1994:DAE

[1873] S. Carlsson. The double algebra: An effective tool for computing invariants in computer vision. Lecture Notes in Computer Science, 825:145–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanGool:1994:MPV

[1874] L. Van Gool, T. Moons, M. Van Diest, and E. Pauwels. Matching perspective views of parallel plane structures. Lecture Notes in Computer Science, 825:165–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weiss:1994:IRS

[1875] I. Weiss. Invariants for recovering shape from shading. Lecture Notes in Computer Science, 825:185–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Astroem:1994:FDP

[1876] K. Astroem. Fundamental difficulties with projective normalization of planar curves. Lecture Notes in Computer Science, 825:199–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verri:1994:ISF

[1877] A. Verri and C. Uras. Invariant size functions. Lecture Notes in Computer Science, 825:215–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartley:1994:ERU

[1878] R. I. Hartley. Euclidean reconstruction from uncalibrated views. Lecture Notes in Computer Science, 825:237– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohr:1994:APR

[1879] R. Mohr, B. Boufama, and P. Brand.

Accurate projective reconstruction. Lecture Notes in Computer Science, 825:257–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papadopoulo:1994:AMF

[1880] T. Papadopoulo and O. Faugeras. Applications of motion field of curves. Lecture Notes in Computer Science, 825:

277–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moons:1994:ARP

[1881] T. Moons, L. Van Gool, M. Van Diest, and E. Pauwels. Affine reconstruction from perspective image pairs obtained by a translating camera. Lecture Notes in Computer Science, 825: 297–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zerroug:1994:UIQ

[1882] M. Zerroug and R. Nevatia. Using invariance and quasi-invariance for the segmentation and recovery of curved objects. Lecture Notes in Computer Science, 825:317–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Forsyth:1994:ROI

[1883] D. Forsyth and C. Rothwell. Representations of 3D objects that incorporate surface markings. Lecture Notes in Computer Science, 825:341–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weinshall:1994:MBI

[1884] D. Weinshall. Model-based invariant functions and their use for recognition. Lecture Notes in Computer Science, 825:359–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mao:1994:IMF

[1885] J. Mao, A. K. Jain, and P. J. Flynn. Integration of multiple feature groups and multiple views into a 3D object recognition system. Lecture Notes in Computer Science, 825:381–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rothwell:1994:HOD

[1886] C. A. Rothwell. Hierarchical object description using invariants. Lecture Notes in Computer Science, 825:397– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacobs:1994:GID

[1887] D. W. Jacobs. Generalizing invariants for 3-D to 2-D matching. Lecture Notes in Computer Science, 825: 415–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basri:1994:RCM

[1888] R. Basri. Recognition by combinations of model views: Alignment and invariance. Lecture Notes in Computer Science, 825:435–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maybank:1994:CBC

[1889] S. J. Maybank. Classification based on the cross ratio. Lecture Notes in Computer Science, 825:453–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meer:1994:CCF

[1890] P. Meer, S. Ramakrishna, and R. Lenz.

Correspondence of coplanar features through P02-Invariant representations. Lecture Notes in Computer Science, 825:473–??, 1994. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Keren:1994:IAC

[1891] D. Keren, J. Subrahmonia, and D. B. Cooper. Integrating algebraic curves and surfaces, algebraic invariants and Bayesian methods for 2D and 3D object recognition. Lecture Notes in Computer Science, 825:493–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barker:1994:MOS

[1892] R. Barker. Managing open systems now that the “glasshouse” has gone. Lecture Notes in Computer Science, 826:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gray:1994:KRT

[1893] P. M. D. Gray. Knowledge reuse through networks of large KBS. Lecture Notes in Computer Science, 826:13–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwiderski:1994:ETB

[1894] S. Schwiderski and G. Saake. Expressing temporal behaviour with extended ECA rules. Lecture Notes in Computer Science, 826:23–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soukeras:1994:TDE

[1895] S. Soukeras and P. J. H. King. Temporal databases: an event-oriented approach. Lecture Notes in Computer Science, 826:38–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1994:OCQ

[1896] D. K. C. Chan and P. W. Trinder. Object comprehensions: a query notation for object-oriented databases. Lecture Notes in Computer Science, 826:55–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alagic:1994:ETL

[1897] S. Alagic and R. Sunderraman. Expressivity of typed logic paradigms for object-oriented databases. Lecture Notes in Computer Science, 826:73–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bagai:1994:ACW

[1898] R. Bagai and R. Sunderraman. Algebraic computation of the weak wellfounded model for general deductive databases. Lecture Notes in Computer Science, 826:90–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jelly:1994:BPS

[1899] I. Jelly, J. Kerridge, and C. Bates.

Benchmarking parallel SQL database machines. Lecture Notes in Computer Science, 826:105–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burger:1994:BTT

[1900] A. Burger and P. Thanisch. Branching transactions: a transaction model for parallel databases. Lecture Notes in Computer Science, 826:121–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCarroll:1994:SSI

[1901] N. McCarroll and J. Kerridge. A strategy for semantic integrity enforcement in a parallel database machine. Lecture Notes in Computer Science, 826: 137–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paton:1994:IOO

[1902] N. W. Paton, G. Al-Qaimari, and K. Doan. On interface objects in object-oriented databases. Lecture Notes in Computer Science, 826:153– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kemp:1994:EAF

[1903] G. J. L. Kemp, J. J. Iriarte, and P. M. D. Gray. Efficient access to FDM objects stored in a relational database. Lecture Notes in Computer Science, 826:170–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barclay:1994:CLQ

[1904] P. J. Barclay and J. B. Kennedy. A conceptual language for querying object-oriented data. Lecture Notes in Computer Science, 826:187–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murphy:1994:JSP

[1905] J. P. Murphy and J. B. Grimson. The Jupiter system: a prototype for multi-database inter-operability. Lecture Notes in Computer Science, 826: 205–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeffery:1994:MHD

[1906] K. G. Jeffery, L. Hutchinson, J. Kalmus, and M. Wilson. A model for heterogeneous distributed database systems. Lecture Notes in Computer Science, 826:221–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soerensen:1994:CTS

[1907] M. U. Soerensen, O. E. Hansen, and H. H. Loevengreen. Combining temporal specification techniques. Lecture Notes in Computer Science, 827:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katz:1994:GEP

[1908] S. Katz. Global equivalence proofs for ISTL. Lecture Notes in Computer Science, 827:17–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeten:1994:RTP

[1909] J. C. M. Baeten, J. A. Bergstra, and R. N. Bol. A real time process logic. Lecture Notes in Computer Science, 827:30–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fiadeiro:1994:STS

[1910] J. L. Fiadeiro and T. Maibaum. Sometimes “tomorrow” is “sometime” — action refinement in a temporal logic of objects. Lecture Notes in Computer Science, 827:48–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonner:1994:ATL

[1911] A. J. Bonner and M. Kifer. Applications of transaction logic to knowledge representation. Lecture Notes in Computer Science, 827:67–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doherty:1994:CFFa

[1912] P. Doherty and W. Lukaszewicz. Circumscribing features and fluents. Lecture Notes in Computer Science, 827: 82–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Badaloni:1994:DTG

[1913] S. Badaloni and M. Berati. Dealing with time granularity in a temporal planning system. Lecture Notes in Computer Science, 827:101–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reynolds:1994:AUI

[1914] M. Reynolds. Axiomatizing U and S over integer time. Lecture Notes in Computer Science, 827:117–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goranko:1994:TLR

[1915] V. Goranko. Temporal logic with reference pointers. Lecture Notes in Computer Science, 827:133–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Venema:1994:CTF

[1916] Y. Venema. Completeness through flatness in two-dimensional temporal logic. Lecture Notes in Computer Science, 827:149–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vergauwen:1994:ECN

[1917] B. Vergauwen, J. Lewi, I. Avau, and A. Pote. Efficient computation of nested fix-points, with applications to model checking. Lecture Notes in Computer Science, 827:165–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grumberg:1994:HLC

[1918] O. Grumberg and R. P. Kurshan. How linear can branching-time be? Lecture Notes in Computer Science, 827: 180–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kutty:1994:FOF

[1919] G. Kutty, L. E. Moser, P. M. Melliar-Smith, and L. K. Dillon. First-order future interval logic. Lecture Notes in Computer Science, 827:195–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernholtz:1994:BOG

[1920] O. Bernholtz and O. Grumberg. Buy one, get one free !!! Lecture Notes in Computer Science, 827:210–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blackburn:1994:BFT

[1921] P. Blackburn, C. Gardent, and M. De Rijke. Back and forth through time and events. Lecture Notes in Computer Science, 827:225–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hwang:1994:ITA

[1922] C. H. Hwang and L. K. Schubert. Interpreting tense, aspect and time adverbials: A compositional, unified approach. Lecture Notes in Computer Science, 827:238–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiMaio:1994:SHP

[1923] M. C. Di Maio and A. Zanardo. Synchronized histories in Prior-Thomason representation of branching time. Lecture Notes in Computer Science, 827: 265–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boehlen:1994:CTD

[1924] M. Boehlen and R. Marti. On the completeness of temporal database query languages. Lecture Notes in Computer Science, 827:283–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanBelleghem:1994:AEC

[1925] K. Van Belleghem, M. Denecker, and D. De Schreye. The abductive event calculus as a general framework for temporal databases. Lecture Notes in Computer Science, 827:301–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wooldridge:1994:DPT

[1926] M. Wooldridge and M. Fisher. A decision procedure for a temporal belief logic. Lecture Notes in Computer Science, 827:317–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xu:1994:DDS

[1927] M. Xu. Decidability of deliberative stit theories with multiple agents. Lecture Notes in Computer Science, 827:332– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ribeiro:1994:ATR

[1928] C. Ribeiro and A. Porto. Abduction in temporal reasoning. Lecture Notes in Computer Science, 827:349–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Felder:1994:TLA

[1929] M. Felder and A. Morzenti. A temporal logic approach to implementation and refinement in timed Petri nets. Lecture Notes in Computer Science, 827:365–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mokkedem:1994:SCT

[1930] A. Mokkedem and D. Mery. A stuttering closed temporal logic for modular reasoning about concurrent programs. Lecture Notes in Computer Science, 827:382–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwiatkowska:1994:HPO

[1931] M. Kwiatkowska, D. Peled, and W. Penczek. A hierarchy of partial order temporal properties. Lecture Notes in Computer Science, 827:398– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dixon:1994:GBA

[1932] C. Dixon, M. Fisher, and H. Barringer. A graph-based approach to resolution in temporal logic. Lecture Notes in Computer Science, 827:415–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McGuire:1994:ABD

[1933] H. McGuire, Z. Manna, and R. Waldinger. Annotation-based deduction in temporal logic. Lecture Notes in Computer Science, 827:430–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orgun:1994:OTM

[1934] M. A. Orgun and W. Ma. An overview of temporal and modal logic programming. Lecture Notes in Computer Science, 827:445–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fisher:1994:SCM

[1935] M. Fisher. A survey of concurrent METATEM — the language and its applications. Lecture Notes in Computer Science, 827:480–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chomicki:1994:TQL

[1936] J. Chomicki. Temporal query languages. Lecture Notes in Computer Science, 827:506–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haehnle:1994:ITL

[1937] R. Haehnle and O. Ibens. Improving temporal logic tableaux using integer constraints. Lecture Notes in Computer Science, 827:535–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Melliar-Smith:1994:SAD

[1938] P. M. Melliar-Smith, L. E. Moser, Y. S.

Ramakrishna, and G. Kutty. A system for automated deduction in graphical interval logic. Lecture Notes in Computer Science, 827:540–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertossi:1994:SRS

[1939] L. E. Bertossi and J. C. Ferretti.

SCDBR: A reasoner for specification in the situation calculus of database updates. Lecture Notes in Computer Science, 827:543–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paulson:1994:IGT

[1940] L. C. Paulson. Isabelle: A generic theorem prover. Lecture Notes in Computer Science, 828:xvii + 321, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Artamonov:1994:AAK

[1941] V. A. Artamonov, A. A. Klyachko, V. M. Sidelnikov, and V. V. Yashchenko. Algebraic aspects of key generation systems. Lecture Notes in Computer Science, 829:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alabbadi:1994:SDS

[1942] M. Alabbadi and S. B. Wicker. Susceptibility of digital signature schemes based on error-correcting codes to universal forgery. Lecture Notes in Computer Science, 829:6–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gabidulin:1994:HAS

[1943] E. M. Gabidulin and O. Kjelsen. How to avoid the Sidel’nikov-Shestakov attack. Lecture Notes in Computer Science, 829:25–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blakley:1994:LAA

[1944] G. R. Blakley and G. A. Kabatianskii. Linear algebra approach to secret sharing schemes. Lecture Notes in Computer Science, 829:33–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Helleseth:1994:GGB

[1945] T. Helleseth, T. Kloeve, and O. Yirehus. Generalizations of the Griesmer bound. Lecture Notes in Computer Science, 829:41–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gabidulin:1994:CCT

[1946] E. M. Gabidulin and V. V. Zanin. Codes that correct two-dimensional burst errors. Lecture Notes in Computer Science, 829:53–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyarinov:1994:SCD

[1947] I. M. Boyarinov. Self-checking decoding algorithm for Reed–Solomon codes. Lecture Notes in Computer Science, 829:63–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zyablov:1994:PUM

[1948] V. V. Zyablov and A. E. Ashikhmin.

Partial unit memory codes on the base of subcodes of Hadamard codes. Lecture Notes in Computer Science, 829: 69–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zyablov:1994:PPU

[1949] V. Zyablov and V. Sidorenko. On periodic (partial) unit memory codes with maximum free distance. Lecture Notes in Computer Science, 829:74–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dettmar:1994:CCC

[1950] U. Dettmar, U. K. Sorger, and V. V.

Zyablov. Concatenated codes with convolutional inner and outer codes. Lecture Notes in Computer Science, 829: 80–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taubin:1994:RSD

[1951] F. A. Taubin. Reduced-state decoding for trellis coded modulation on nonlineaer intersymbol interference channels. Lecture Notes in Computer Science, 829:88–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asnis:1994:TCD

[1952] I. L. Asnis, S. V. Fedorenko, E. A.

Krouk, and E. T. Mironchikov. Tables of coverings for decoding by Ssets. Lecture Notes in Computer Science, 829:97–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sidorenko:1994:POD

[1953] V. Sidorenko. Periodicity of onedimensional tilings. Lecture Notes in Computer Science, 829:103–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bocharova:1994:FIR

[1954] I. E. Bocharova and B. D. Kudryashov. Fast infinite response filtering for speech processing. Lecture Notes in Computer Science, 829:109–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bocharova:1994:TCL

[1955] I. E. Bocharova, V. D. Kolesnik, V. Yu Krachkovsky, and B. D. Kudryashov. On trellis codes for linear predictive speech compression. Lecture Notes in Computer Science, 829:115–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drogoul:1994:MAS

[1956] A. Drogoul and J. Ferber. Multi-agent simulation as a tool for modeling societies: Application to social differentiation in ant colonies. Lecture Notes in Computer Science, 830:3–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kearney:1994:EMA

[1957] P. J. Kearney. Experiments in multiagent system dynamics. Lecture Notes in Computer Science, 830:24–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cecconi:1994:SAE

[1958] F. Cecconi, D. Denaro, D. Parisi, and U. Piazzalunga. Social aggregations in evolving neural networks. Lecture Notes in Computer Science, 830:41–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bates:1994:AAE

[1959] J. Bates, A. B. Loyall, and W. S.

Reilly. An architecture for action, emotion, and social behavior. Lecture Notes in Computer Science, 830:55–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lomborg:1994:GTV

[1960] B. Lomborg. Game theory versus multiple agents: The iterated prisoner’s dilemma. Lecture Notes in Computer Science, 830:69–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marsh:1994:TDA

[1961] S. Marsh. Trust in distributed artificial intelligence. Lecture Notes in Computer Science, 830:94–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zlotkin:1994:NII

[1962] G. Zlotkin and J. S. Rosenschein.

Negotiation with incomplete information about worth: Strict versus tolerant mechanisms. Lecture Notes in Computer Science, 830:115–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuwabara:1994:EAD

[1963] K. Kuwabara and T. Ishida. Equilibratory approach to distributed resource allocation: Toward coordinated balancing. Lecture Notes in Computer Science, 830:133–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

CastroCaldas:1994:SIO

[1964] J. Castro Caldas and H. Coelho. Strategic interaction in oligopolistic markets — experimenting with real and artificial agents. Lecture Notes in Computer Science, 830:147–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stirling:1994:MAC

[1965] W. Stirling. Multi-agent coordinated decision-making using epistemic utility theory. Lecture Notes in Computer Science, 830:164–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castelfranchi:1994:MSB

[1966] C. Castelfranchi and E. Werner. The MAAMAW spirit and this book. Lecture Notes in Computer Science, 830: vii–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habib:1994:BVE

[1967] M. Habib and L. Nourine. Bit-vector encoding for partially ordered sets. Lecture Notes in Computer Science, 831: 1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bogart:1994:IOW

[1968] K. P. Bogart. Intervals and orders: What comes after interval orders? Lecture Notes in Computer Science, 831:13–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spinrad:1994:DA

[1969] J. Spinrad. Dimension and algorithms. Lecture Notes in Computer Science, 831:33–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hashemi:1994:UDF

[1970] S. M. Hashemi and I. Rival. Upward drawings to fit surfaces. Lecture Notes in Computer Science, 831:53–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simon:1994:CTO

[1971] K. Simon and P. Trunz. A cleanup on transitive orientation. Lecture Notes in Computer Science, 831:59–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cattel:1994:CGV

[1972] K. Cattel and M. J. Dinneen. A characterization of graphs with vertex cover up to five. Lecture Notes in Computer Science, 831:86–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gustedt:1994:THP

[1973] J. Gustedt and A. Steger. Testing hereditary properties efficiently on average. Lecture Notes in Computer Science, 831:100–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franciosa:1994:OSF

[1974] P. G. Franciosa and M. Talamo. Orders, k-sets and fast halfplane search on paged memory. Lecture Notes in Computer Science, 831:117–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1994:TGT

[1975] Y.-L. Lin. Triangle graphs and their coloring. Lecture Notes in Computer Science, 831:128–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Capelle:1994:ROU

[1976] C. Capelle. Representation of an order as union of interval orders. Lecture Notes in Computer Science, 831: 143–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mitas:1994:MRS

[1977] J. Mitas. Minimal representation of semiorders with intervals of same length. Lecture Notes in Computer Science, 831:162–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Talamo:1994:FLB

[1978] M. Talamo and P. Vocca. Fast lattice browsing on sparse representations. Lecture Notes in Computer Science, 831:186–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aiken:1994:CSC

[1979] A. Aiken, D. Kozen, M. Vardi, and E. Wimmers. The complexity of set constraints. Lecture Notes in Computer Science, 832:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambler:1994:DMC

[1980] S. Ambler, M. Kiviatkowska, and N. Measor. On duality for the modalcalculus. Lecture Notes in Computer Science, 832:18–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baaz:1994:SPT

[1981] M. Baaz and R. Zach. Short proofs of tautologies using the schema of equivalence. Lecture Notes in Computer Science, 832:33–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergstra:1994:PAC

[1982] J. A. Bergstra, I. Bethke, and A. Ponse.

Process algebra with combinators. Lecture Notes in Computer Science, 832: 36–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blass:1994:GSN

[1983] A. Blass. Is game semantics necessary? Lecture Notes in Computer Science, 832:66–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Compton:1994:MTC

[1984] K. J. Compton and R. Uceda-Sosa. The model theory of concurrency. Lecture Notes in Computer Science, 832:78–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coquand:1994:SRM

[1985] C. Coquand. From semantics to rules: a machine assisted analysis. Lecture Notes in Computer Science, 832:91–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goerdt:1994:RAS

[1986] A. Goerdt and U. Kamps. On the reasons for average superlinear speedup in parallel backtrack search. Lecture Notes in Computer Science, 832:106– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gomolko:1994:NIP

[1987] A. Gomolko. Negation as inconsistency in PROLOG via intuitionistic logic. Lecture Notes in Computer Science, 832:128–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graedel:1994:AMP

[1988] E. Graedel and A. Malmstroem. Approximable minimisation problems and optimal solutions on random inputs. Lecture Notes in Computer Science, 832:139–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grohe:1994:BAH

[1989] M. Grohe. Bounded-arity hierarchies in fixed-point logics. Lecture Notes in Computer Science, 832:150–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirshfeld:1994:PNE

[1990] Y. Hirshfeld. Petri nets and the equivalence problem. Lecture Notes in Computer Science, 832:165–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kozen:1994:LAS

[1991] D. Kozen. Logical aspects of set constraints. Lecture Notes in Computer Science, 832:175–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Makowsky:1994:OQ

[1992] J. A. Makowsky and Y. B. Pnueli. Oracles and quantifiers. Lecture Notes in Computer Science, 832:189–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marcinkowski:1994:HCI

[1993] J. Marcinkowski. A Horn clause that implies an undecidable set of Horn clauses. Lecture Notes in Computer Science, 832:223–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Milner:1994:HOA

[1994] R. Milner. Higher-order action calculi. Lecture Notes in Computer Science, 832:238–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ong:1994:GSN

[1995] C.-H. L. Ong and E. Ritter. A generic strong normalisation argument: Application to the calculus of constructions. Lecture Notes in Computer Science, 832:261–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raffalli:1994:DTI

[1996] C. Raffalli. Data types, infinity and equality in system AF2. Lecture Notes in Computer Science, 832:280– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ritter:1994:NTL

[1997] E. Ritter. Normalisation for typed lambda calculi with explicit substitution. Lecture Notes in Computer Science, 832:295–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sieg:1994:PTP

[1998] W. Sieg and S. S. Wainer. Program transformation and proof transformation. Lecture Notes in Computer Science, 832:305–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stewart:1994:IGQ

[1999] I. A. Stewart. Incorporating generalised quantifiers and the least fixed point operator. Lecture Notes in Computer Science, 832:318–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gurevich:1994:ESC

[2000] Y. Gurevich and J. K. Huggins. Errata to “the semantics of the C programming language”. Lecture Notes in Computer Science, 832:334–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1994:BIF

[2001] D. Dubois and H. Prade. Basic issues on fuzzy rules and their application to fuzzy control. Lecture Notes in Computer Science, 833:3–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mukaidono:1994:AST

[2002] M. Mukaidono. Algebraic structures of truth values in fuzzy logic. Lecture Notes in Computer Science, 833:15–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldwin:1994:TMA

[2003] J. F. Baldwin. A theory of mass assignments for artificial intelligence. Lecture Notes in Computer Science, 833:22–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pedrycz:1994:FDS

[2004] W. Pedrycz. Fuzzy dynamic systems. Lecture Notes in Computer Science, 833:37–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1994:HUC

[2005] D. Dubois, J. Lang, and H. Prade.

Handling uncertainty, context, vague predicates, and partial inconsistency in possibilistic logic. Lecture Notes in Computer Science, 833:45–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graham:1994:AFM

[2006] B. Graham and R. Newell. An adaptive fuzzy model-based controller. Lecture Notes in Computer Science, 833:56–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franklin:1994:FRN

[2007] J. Franklin. Fuzzy representations in neural nets. Lecture Notes in Computer Science, 833:69–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Narazaki:1994:MIQ

[2008] H. Narazaki and A. L. Ralescu. A method to implement qualitative knowledge in multi-layered neural networks. Lecture Notes in Computer Science, 833:82–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wakami:1994:FCN

[2009] N. Wakami. Fuzzy control and neural networks: Applications for consumer products. Lecture Notes in Computer Science, 833:99–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yokogawa:1994:CBR

[2010] T. Yokogawa, T. Sakurai, A. Nukuzuma, and T. Takagi. Case-based reasoning for action planning by representing situations at the abstract layers. Lecture Notes in Computer Science, 833:109–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Venkatesh:1994:UFL

[2011] S. Venkatesh and D. Kieronska. A unified formalism for landmark-based representation of maps and navigation plans. Lecture Notes in Computer Science, 833:123–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yen:1994:UFL

[2012] J. Yen and N. Pfluger. Using fuzzy logic in a mobile robot path controller. Lecture Notes in Computer Science, 833:133–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirota:1994:FFL

[2013] K. Hirota. Fundamentals of fuzzy logical circuits. Lecture Notes in Computer Science, 833:145–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1994:PES

[2014] W.-T. Wu. Polynomial equationssolving and its applications. Lecture Notes in Computer Science, 834:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graham:1994:QRC

[2015] R. L. Graham. Quasi-random combinatorial structures. Lecture Notes in Computer Science, 834:10–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klein:1994:HAV

[2016] R. Klein and A. Lingas. Hamiltonian abstract Voronoi diagrams in linear time. Lecture Notes in Computer Science, 834:11–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bose:1994:ESD

[2017] P. Bose, M. E. Houle, and G. Toussaint. Every set of disjoint line segments admits a binary tree. Lecture Notes in Computer Science, 834:20–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iwano:1994:TCP

[2018] K. Iwano, P. Raghavan, and H. Tamaki. The traveling cameraman problem, with applications to automatic optical inspection. Lecture Notes in Computer Science, 834:29–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akutsu:1994:DCP

[2019] T. Akutsu. On determining the congruity of point sets in higher dimensions. Lecture Notes in Computer Science, 834:38–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Book:1994:RHS

[2020] R. V. Book and O. Watanabe. On random hard sets for NP. Lecture Notes in Computer Science, 834:47–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hemaspaandra:1994:CSU

[2021] L. A. Hemaspaandra, A. V. Naik, M. Ogihara, and A. L. Selman. Computing solutions uniquely collapses the polynomial hierarchy. Lecture Notes in Computer Science, 834:56–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santha:1994:VDP

[2022] M. Santha and S. Tan. Verifying the determinant in parallel. Lecture Notes in Computer Science, 834:65–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1994:HAG

[2023] C.-L. Lin. Hardness of approximating graph transformation problem. Lecture Notes in Computer Science, 834:74–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Panconesi:1994:RWF

[2024] A. Panconesi, M. Papatriantafilou, P. Tsigas, and P. Vitanyi. Randomized wait-free naming. Lecture Notes in Computer Science, 834:83–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rajasekaran:1994:UTN

[2025] S. Rajasekaran, W. Chen, and S. Yooseph. Unifying themes for network selection. Lecture Notes in Computer Science, 834:92–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Molcho:1994:RAA

[2026] M. Molcho and S. Zaks. Robust asynchronous algorithms in networks with a fault detection ring. Lecture Notes in Computer Science, 834:101–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jennings:1994:DAF

[2027] E. Jennings and L. Motyckova. Distributed algorithms for finding small cycle covers in arbitrary networks. Lecture Notes in Computer Science, 834: 110–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1994:NNM

[2028] X.-S. Zhang and H.-C. Zhu. A neural network model for quadratic programming with simple upper and lower bounds and its application to linear programming. Lecture Notes in Computer Science, 834:119–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhu:1994:ACT

[2029] D. Zhu, S. Ma, and H. Qiu. Analysis of the convergency of topology preserving neural networks on learning. Lecture Notes in Computer Science, 834:128–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arvind:1994:HIP

[2030] V. Arvind, J. Koebler, and R. Schuler. On helping and interactive proof systems. Lecture Notes in Computer Science, 834:137–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gu:1994:ATC

[2031] J. Gu and Q.-P. Gu. Average time complexity of the SAT 1.2 algorithm. Lecture Notes in Computer Science, 834: 146–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1994:ACP

[2032] J. Chen. Average cost to produce partial orders. Lecture Notes in Computer Science, 834:155–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1994:TNL

[2033] D. Li and S. Ma. Two notes on lowdensity subset sum algorithms. Lecture Notes in Computer Science, 834: 164–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cai:1994:RDT

[2034] J.-Y. Cai and M. D. Hirsch. Rotation distance, triangulations of planar surfaces and hyperbolic geometry. Lecture Notes in Computer Science, 834:172–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1994:FRS

[2035] Y. Chen and T. R. Smith. Finitely representable spatial objects and efficient computation. Lecture Notes in Computer Science, 834:181–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nagamochi:1994:CAS

[2036] H. Nagamochi, K. Nishimura, and T. Ibaraki. Computing all small cuts in undirected networks. Lecture Notes in Computer Science, 834:190–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1994:ECC

[2037] X. Zhou and T. Nishizeki. Edgecoloring and f-coloring for various classes of graphs. Lecture Notes in Computer Science, 834:199–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gopalakrishnan:1994:EDP

[2038] C. P. Gopalakrishnan and C. P. Rangan. Edge-disjoint paths in permutation graphs. Lecture Notes in Computer Science, 834:208–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taoka:1994:MAE

[2039] S. Taoka and T. Watanabe. Minimum augmentation to k-edge-connect specified vertices of a graph. Lecture Notes in Computer Science, 834:217– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Srivastav:1994:ACH

[2040] A. Srivastav and P. Stangier. Algorithmic Chernoff–Hoeffding inequalities in integer programming. Lecture Notes in Computer Science, 834:226–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takaoka:1994:APM

[2041] T. Takaoka. Approximate pattern matching with samples. Lecture Notes in Computer Science, 834:234– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1994:EOA

[2042] X. Lin and M. E. Orlowska. An efficient optimal algorithm for minimizing the overall communication cost in replicated data management. Lecture Notes in Computer Science, 834:243– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gu:1994:MSN

[2043] J. Gu. Multispace search: A new optimization approach. Lecture Notes in Computer Science, 834:252–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arikati:1994:RDS

[2044] S. Arikati and A. Maheshwari. Realizing degree sequences in parallel. Lecture Notes in Computer Science, 834: 261–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kavvadias:1994:ESP

[2045] D. Kavvadias, G. E. Pantziou, P. G.

Spirakis, and C. D. Zaroliagis. Efficient sequential and parallel algorithms for the negative cycle problem. Lecture Notes in Computer Science, 834: 270–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1994:FEO

[2046] D. Z. Chen and X. Hu. Fast and efficient operations on parallel priority queues. Lecture Notes in Computer Science, 834:279–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carlsson:1994:HB

[2047] S. Carlsson, J. Chen, and C. Mattsson. Heaps with bits. Lecture Notes in Computer Science, 834:288–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VonzurGathen:1994:CPC

[2048] J. Von zur Gathen and I. Shparlinski. Components and projections of curves over finite fields. Lecture Notes in Computer Science, 834:297–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Middendorf:1994:AFV

[2049] M. Middendorf. On the approximation of finding various minimal, maximal, and consistent sequences. Lecture Notes in Computer Science, 834: 306–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakanishi:1994:LTA

[2050] M. Nakanishi, M. Hasidume, M. Ito, and A. Hashimoto. A linear-time algorithm for computing characteristic strings. Lecture Notes in Computer Science, 834:315–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Makino:1994:MLI

[2051] K. Makino and T. Ibaraki. The maximum latency and identification of positive Boolean functions. Lecture Notes in Computer Science, 834:324– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaufmann:1994:FDH

[2052] M. Kaufmann, H. Lauer, and H. Schroeder. Fast deterministic hot-potato routing on processor arrays. Lecture Notes in Computer Science, 834:333–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gu:1994:PCF

[2053] Q.-P. Gu and S. Peng. k-pairwise cluster fault tolerant routing in hypercubes. Lecture Notes in Computer Science, 834:342–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaufmann:1994:SMT

[2054] M. Kaufmann, S. Gao, and K. Thulasiraman. On Steiner minimal trees in grid graphs and its application to VLSI routing. Lecture Notes in Computer Science, 834:351–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shen:1994:ORP

[2055] X. Shen, Q. Hu, H. Dai, and X. Wang.

Optimal routing of permutations on rings. Lecture Notes in Computer Science, 834:360–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambos-Spies:1994:RBR

[2056] K. Ambos-Spies, S. A. Terwijn, and X. Zheng. Resource bounded randomness and weakly complete problems. Lecture Notes in Computer Science, 834:369–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krishnan:1994:DOA

[2057] S. C. Krishnan, A. Puri, and R. K.

Brayton. Deterministic omega automata vis-a-vis deterministic Buchi automata. Lecture Notes in Computer Science, 834:378–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mihara:1994:QCN

[2058] T. Mihara and T. Nishino. Quantum computation and NP-Complete problems. Lecture Notes in Computer Science, 834:387–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fortnow:1994:SOW

[2059] L. Fortnow and J. Rogers. Separability and one-way functions. Lecture Notes in Computer Science, 834:396– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akutsu:1994:ALC

[2060] T. Akutsu and M. M. Halldorsson. On the approximation of largest common subtrees and largest common point sets. Lecture Notes in Computer Science, 834:405–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jiang:1994:ASS

[2061] T. Jiang and L. Wang. An approximation scheme for some Steiner tree problems in the plane. Lecture Notes in Computer Science, 834:414–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1994:CDA

[2062] B.-T. Yang, Y.-F. Xu, and Z.-Y. You. A chain decomposition algorithm for the proof of a property on minimum weight triangulations. Lecture Notes in Computer Science, 834:423–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ding:1994:GAQ

[2063] J. Ding. The generating algorithm for quasi-full Steiner trees. Lecture Notes in Computer Science, 834:428– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

ChiKwongHui:1994:ADC

[2064] L. Chi Kwong Hui and C. U. Martel. Analysing deletions in competitive self-adjusting linear list algorithms. Lecture Notes in Computer Science, 834:433–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abuaiadh:1994:FHO

[2065] D. Abuaiadh and J. H. Kingston. Are Fibonacci heaps optimal? Lecture Notes in Computer Science, 834:442– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kao:1994:SEG

[2066] M.-Y. Kao and S.-H. Teng. Simple and efficient graph compression schemes for dense and complement graphs. Lecture Notes in Computer Science, 834:451–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Subramanian:1994:IAC

[2067] C. R. Subramanian. Improved algorithms for coloring random graphs. Lecture Notes in Computer Science, 834:460–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ibarra:1994:PCS

[2068] O. H. Ibarra and N. Q. Tran. On the parallel complexity of solving recurrence equations. Lecture Notes in Computer Science, 834:469–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iwamoto:1994:EGC

[2069] C. Iwamoto and K. Iwama. Extended graph connectivity and its gradually increasing parallel complexity. Lecture Notes in Computer Science, 834:478– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Metaxas:1994:PDB

[2070] P. T. Metaxas, G. E. Pantziou, and A. Symvonis. Parallel h-v drawings of binary trees. Lecture Notes in Computer Science, 834:487–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hsieh:1994:EPS

[2071] S.-Y. Hsieh and C.-W. Ho. An efficient parallel strategy for recognizing series-parallel graphs. Lecture Notes in Computer Science, 834:496–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tokuyama:1994:OQS

[2072] T. Tokuyama. Orthogonal queries in segments and triangles. Lecture Notes in Computer Science, 834:505– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhu:1994:FCG

[2073] B. Zhu. Further computational geometry in secondary memory. Lecture Notes in Computer Science, 834:514– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tan:1994:SSR

[2074] X. Tan and T. Hirata. Shortest safari routes in simple polygon. Lecture Notes in Computer Science, 834:523– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chin:1994:GTP

[2075] F. Y. Chin and Cao An Wang. On greedy tetrahedralization of points in 3D. Lecture Notes in Computer Science, 834:532–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujiyoshi:1994:TPB

[2076] K. Fujiyoshi, Y. Kajitani, and H. Niitsu. The totally-perfect bipartite graph and its construction. Lecture Notes in Computer Science, 834:541– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsai:1994:BCC

[2077] K.-H. Tsai and D. T. Lee. k-best cuts for circular-arc graphs. Lecture Notes in Computer Science, 834:550– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1994:RCA

[2078] L. Chen. Revisiting circular arc graphs. Lecture Notes in Computer Science, 834:559–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rangan:1994:WII

[2079] C. P. Rangan and M.-S. Chang. Weighted irredundance of interval graphs. Lecture Notes in Computer Science, 834:567–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tani:1994:ROO

[2080] S. Tani and H. Imai. A reordering operation for an ordered binary decision diagram and an extended framework for combinatorics of graphs. Lecture Notes in Computer Science, 834:575– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hosaka:1994:SOB

[2081] K. Hosaka, Y. Takenaga, and S. Yajima. On the size of ordered binary decision diagrams representing threshold functions. Lecture Notes in Computer Science, 834:584–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1994:NLS

[2082] H.-Z. Li and G.-Y. Li. Nonuniform lowness and strong nonuniform lowness. Lecture Notes in Computer Science, 834:593–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biehl:1994:DES

[2083] I. Biehl. Definition and existence of super complexity cores. Lecture Notes in Computer Science, 834:600–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Young:1994:EAA

[2084] G. H. Young and C.-L. Chan. Efficient algorithms for assigning chainlike tasks on a chain-like network computer. Lecture Notes in Computer Science, 834:607–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lau:1994:MSS

[2085] Hoong Chuin Lau. Manpower scheduling with shift change constraints. Lecture Notes in Computer Science, 834: 616–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xu:1994:FAT

[2086] Y. Xu and H. N. Gabow. Fast algorithms for transversal matroid intersection problems. Lecture Notes in Computer Science, 834:625–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deng:1994:CHA

[2087] X. Deng. A convex hull algorithm on coarse-grained multiprocessors. Lecture Notes in Computer Science, 834:634– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1994:MMN

[2088] D.-L. Lee and K. E. Batcher. A multiway merging network. Lecture Notes in Computer Science, 834:643–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asano:1994:RCI

[2089] T. Asano. Reporting and counting intersections of lines within a polygon. Lecture Notes in Computer Science, 834:652–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Havas:1994:NPS

[2090] G. Havas and Jin Xian Lian. A new problem in string searching. Lecture Notes in Computer Science, 834:660– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sun:1994:ASC

[2091] S. Sun, D. H. C. Du, and G. Xue. Achieving the shortest clock period by inserting the minimum amount of delay. Lecture Notes in Computer Science, 834:669–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bao:1994:RAF

[2092] F. Bao and Y. Igarashi. A randomized algorithm to finite automata public key cryptosystem. Lecture Notes in Computer Science, 834:678–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hayes:1994:APV

[2093] P. Hayes. Aristotelian and platonic views of knowledge representation. Lecture Notes in Computer Science, 835: 1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Way:1994:CGP

[2094] E. C. Way. Conceptual graphs — past, present, and future. Lecture Notes in Computer Science, 835:11–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Willems:1994:PSC

[2095] M. Willems. Pragmatic semantics by conceptual graphs. Lecture Notes in Computer Science, 835:31–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moulin:1994:TSD

[2096] B. Moulin and S. Dumas. The temporal structure of a discourse and verb tense determination. Lecture Notes in Computer Science, 835:45–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Myaeng:1994:LPT

[2097] S. H. Myaeng, C. Khoo, and M. Li. Linguistic processing of text for a largescale conceptual information retrieval system. Lecture Notes in Computer Science, 835:69–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rassinoux:1994:MAM

[2098] A.-M. Rassinoux, R. H. Baud, and J.-R. Scherrer. A multilingual analyser of medical texts. Lecture Notes in Computer Science, 835:84–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sykes:1994:EEI

[2099] J. T. Sykes, V. Konstantinou, and P. L. R. Morse. Extracting explicit and implicit knowledge from natural language texts. Lecture Notes in Computer Science, 835:97–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mann:1994:RGS

[2100] G. A. Mann. A rational goal-seeking agent using conceptual graphs. Lecture Notes in Computer Science, 835: 113–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tepfenhart:1994:AKP

[2101] W. M. Tepfenhart. Attitudes: Keys to problem identification. Lecture Notes in Computer Science, 835:127– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wermelinger:1994:BCS

[2102] M. Wermelinger and J. G. Lopes. Basic conceptual structures theory. Lecture Notes in Computer Science, 835: 144–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mineau:1994:VMF

[2103] G. W. Mineau. Views, mappings and functions: Essential definitions to the conceptual graph theory. Lecture Notes in Computer Science, 835:160– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esch:1994:CCA

[2104] J. Esch. Contexts and concepts, abstraction duals. Lecture Notes in Computer Science, 835:175–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esch:1994:CCC

[2105] J. Esch. Contexts, canons and coreferent types. Lecture Notes in Computer Science, 835:185–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dick:1994:UCR

[2106] J. P. Dick. Using contexts to represent text. Lecture Notes in Computer Science, 835:196–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghosh:1994:ISC

[2107] B. C. Ghosh and V. Wuwongse. Inference systems for conceptual graph programs. Lecture Notes in Computer Science, 835:214–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levinson:1994:UUD

[2108] R. Levinson. USD: A universal data structure. Lecture Notes in Computer Science, 835:230–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kabbaj:1994:CCO

[2109] A. Kabbaj, C. Frasson, M. Kaltenbach, and J.-Y. Djamen. A conceptual and contextual object-oriented logic programming: The PROLOG++ language. Lecture Notes in Computer Science, 835:251–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cyre:1994:KVC

[2110] W. R. Cyre, S. Balanchandar, and A. Thakar. Knowledge visualization from conceptual structures. Lecture Notes in Computer Science, 835:275– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ellis:1994:EIO

[2111] G. Ellis and F. Lehmann. Exploiting the induced order on typelabeled graphs for fast knowledge retrieval. Lecture Notes in Computer Science, 835:293–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carbonneill:1994:SIR

[2112] B. Carbonneill and O. Haemmerle. Standardizing and interfacing relational databases using conceptual graphs. Lecture Notes in Computer Science, 835:311–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Girard:1994:GI

[2113] J.-Y Girard. Geometry of interaction. Lecture Notes in Computer Science, 836:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uselton:1994:CSS

[2114] A. C. Uselton and S. A. Smolka. A compositional semantics for statecharts using labeled transition systems. Lecture Notes in Computer Science, 836:2–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kiehn:1994:DNI

[2115] A. Kiehn and M. Hennessy. On the decidability of non-interleaving process equivalences. Lecture Notes in Computer Science, 836:18–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mauw:1994:RBS

[2116] S. Mauw and H. Mulder. Regularity of BPA-systems is decidable. Lecture Notes in Computer Science, 836:34–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirshfeld:1994:FAD

[2117] Y. Hirshfeld and F. Moller. A fast algorithm for deciding bisimilarity of normed context-free processes. Lecture Notes in Computer Science, 836:48–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez:1994:TMB

[2118] J. C. Fernandez, J. Sifakis, and R. De Simone. Tutorial: Model-based verification methods and tools. Lecture Notes in Computer Science, 836:64–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Girkar:1994:NRA

[2119] M. Girkar and R. Moll. New results on the analysis of concurrent systems with an indefinite number of processes. Lecture Notes in Computer Science, 836:65–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouajjani:1994:VNT

[2120] A. Bouajjani, R. Echahed, and R. Robbana. Verification of nonregular temporal properties for context-free processes. Lecture Notes in Computer Science, 836:81–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burkart:1994:PPP

[2121] O. Burkart and B. Steffen. Pushdown processes: Parallel composition and model checking. Lecture Notes in Computer Science, 836:98–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hungar:1994:LMC

[2122] H. Hungar. Local model checking for parallel compositions of context-free processes. Lecture Notes in Computer Science, 836:114–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Panangaden:1994:ILL

[2123] P. Panangaden. Invited lecture: The logical structure of concurrent constraint programming languages. Lecture Notes in Computer Science, 836: 129–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiGianantonio:1994:CND

[2124] P. Di Gianantonio, F. Honsell, S. Liani, and G. D. Plotkin. Countable nondeterminism and uncountable limits. Lecture Notes in Computer Science, 836:130–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cockett:1994:SC

[2125] J. R. B. Cockett and D. A. Spooner.

SProc categorically. Lecture Notes in Computer Science, 836:146–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courcoubetis:1994:TTG

[2126] C. Courcoubetis. Tutorial: From timed graphs to hybrid automata. Lecture Notes in Computer Science, 836:160– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dill:1994:HMS

[2127] D. Dill. Hierarchical models of synchronous circuits. Lecture Notes in Computer Science, 836:161–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alur:1994:OPC

[2128] R. Alur, C. Courcoubetis, and T. A.

Henzinger. The observational power of clocks. Lecture Notes in Computer Science, 836:162–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gunawardena:1994:DAT

[2129] J. Gunawardena. A dynamic approach to timed behaviour. Lecture Notes in Computer Science, 836:178–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Winkowski:1994:APT

[2130] J. Winkowski. Algebras of processes of timed Petri nets. Lecture Notes in Computer Science, 836:194–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koutny:1994:OSP

[2131] M. Koutny, J. Esparza, and E. Best. Operational semantics for the Petri box calculus. Lecture Notes in Computer Science, 836:210–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rensink:1994:WSC

[2132] A. Rensink and H. Wehrheim. Weak sequential composition in process algebras. Lecture Notes in Computer Science, 836:226–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kanellakis:1994:EPV

[2133] P. C. Kanellakis, D. Michailidis, and A. A. Shvartsman. Efficient parallelism vs. reliable distribution: A trade-off for concurrent computations. Lecture Notes in Computer Science, 836:242– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xu:1994:UAC

[2134] Q. Xu, A. Cau, and P. Gollette. On unifying assumption-commitment style proof rules for concurrency. Lecture Notes in Computer Science, 836:267– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skakkeboek:1994:LFD

[2135] J. U. Skakkeboek. Liveness and fairness in duration calculus. Lecture Notes in Computer Science, 836:283–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boreale:1994:SSC

[2136] M. Boreale and R. De Nicola. A symbolic semantics for the π-calculus. Lecture Notes in Computer Science, 836: 299–314, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walker:1994:BC

[2137] D. Walker. On bisimulation in the πcalculus. Lecture Notes in Computer Science, 836:315–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1994:CBC

[2138] X. Liu. Characterizing bisimulation congruence in the π-calculus. Lecture Notes in Computer Science, 836:331– 350, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klarlund:1994:LVI

[2139] N. Klarlund. The limit view of infinite computations. Lecture Notes in Computer Science, 836:351–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Back:1994:TRA

[2140] R. J. R. Back and J. Von Wright. Trace refinement of action systems. Lecture Notes in Computer Science, 836:367– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielsen:1994:BMC

[2141] M. Nielsen and C. Clausen. Bisimulation for models in concurrency. Lecture Notes in Computer Science, 836: 385–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bezem:1994:IPA

[2142] M. Bezem and J. F. Groote. Invariants in process algebra with data. Lecture Notes in Computer Science, 836: 401–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cleaveland:1994:TBA

[2143] R. Cleaveland and J. Riely. Testingbased abstractions for value-passing systems. Lecture Notes in Computer Science, 836:417–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verhoef:1994:CTS

[2144] C. Verhoef. A congruence theorem for structured operational semantics with predicates and negative premises. Lecture Notes in Computer Science, 836: 433–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aceto:1994:DCI

[2145] L. Aceto. Deriving complete inference systems for a class of GSOS languages generating regular behaviours. Lecture Notes in Computer Science, 836:449– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeten:1994:PAP

[2146] J. C. M. Baeten and J. A. Bergstra.

Process algebra with partial choice. Lecture Notes in Computer Science, 836:465–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Segala:1994:PSP

[2147] R. Segala and N. Lynch. Probabilistic simulations for probabilistic processes. Lecture Notes in Computer Science, 836:481–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuen:1994:FAC

[2148] S. Yuen, R. Cleaveland, Z. Dayar, and S. A. Smolka. Fully abstract characterizations of testing preorders for probabilistic processes. Lecture Notes in Computer Science, 836:497–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1994:CBP

[2149] S.-H. Wu, S. A. Smolka, and E. W.

Stark. Composition and behaviors of probabilistic I/O automata. Lecture Notes in Computer Science, 836:513– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolodner:1994:UCC

[2150] J. L. Kolodner. Understanding creativity: A case-based approach. Lecture Notes in Computer Science, 837:3–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keane:1994:AAC

[2151] M. T. Keane. Analogical asides on case-based reasoning. Lecture Notes in Computer Science, 837:21–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Veloso:1994:PAA

[2152] M. M. Veloso. Prodigy/analogy: Analogical reasoning in general problem solving. Lecture Notes in Computer Science, 837:33–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Armengol:1994:KLM

[2153] E. Armengol and E. Plaza. A knowledge level model of case-based reasoning. Lecture Notes in Computer Science, 837:53–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakhaeizadeh:1994:LPT

[2154] G. Nakhaeizadeh. Learning prediction of time series — A theoretical and empirical comparison of CBR with some other approaches. Lecture Notes in Computer Science, 837:65–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wess:1994:CBS

[2155] S. Wess and C. Globig. Case-based and symbolic classification. Lecture Notes in Computer Science, 837:77–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bento:1994:SMR

[2156] C. Bento and E. Costa. A similarity metric for retrieval of cases imperfectly explained. Lecture Notes in Computer Science, 837:92–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bunke:1994:SMS

[2157] H. Bunke and B. T. Messmer. Similarity measures for structured representations. Lecture Notes in Computer Science, 837:106–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sebag:1994:RBS

[2158] M. Sebag and M. Schoenauer. A rule-based similarity measure. Lecture Notes in Computer Science, 837:119– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1994:UMM

[2159] M. G. Brown. An underlying memory model to support case retrieval. Lecture Notes in Computer Science, 837: 132–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Myllymaeki:1994:MPC

[2160] P. Myllymaeki and H. Tirri. Massively parallel case-based reasoning with probabilistic similarity metrics. Lecture Notes in Computer Science, 837:144–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rougegrez:1994:SEB

[2161] S. Rougegrez. Similarity evaluation between observed behaviours for the prediction of processes. Lecture Notes in Computer Science, 837:155–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wess:1994:UTI

[2162] S. Wess, K.-D. Althoff, and G. Derwand. Using k-d trees to improve the retrieval step in case-based reasoning. Lecture Notes in Computer Science, 837:167–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergmann:1994:EBS

[2163] R. Bergmann, G. Pews, and W. Wilke. Explanation-based similarity: A unifying approach for integrating domain knowledge into case-based reasoning for diagnosis and planning tasks. Lecture Notes in Computer Science, 837: 182–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boerner:1994:SSG

[2164] K. Boerner. Structural similarity as guidance in case-based design. Lecture Notes in Computer Science, 837:197– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smyth:1994:RAC

[2165] B. Smyth and M. T. Keane. Retrieving adaptable cases: The role of adaptation knowledge in case retrieval. Lecture Notes in Computer Science, 837:209–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chatterjee:1994:ATI

[2166] N. Chatterjee and J. A. Campbell. Adaptation through interpolation for time-critical case-based reasoning. Lecture Notes in Computer Science, 837: 221–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cunningham:1994:KER

[2167] P. Cunningham, D. Finn, and S. Slattery. Knowledge engineering requirements in derivational analogy. Lecture Notes in Computer Science, 837:234– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OHara:1994:IRI

[2168] S. O’Hara and B. Indurkhya. Incorporating (re)-interpretation in case-based reasoning. Lecture Notes in Computer Science, 837:246–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uehara:1994:PPB

[2169] K. Uehara, M. Tanizawa, and S. Maekawa. PBL: Prototype-based learning algorithms. Lecture Notes in Computer Science, 837:261–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aamodt:1994:EDC

[2170] A. Aamodt. Explanation-driven casebased reasoning. Lecture Notes in Computer Science, 837:274–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arcos:1994:RAI

[2171] J. L. Arcos and E. Plaza. A reflective architecture for integrated memorybased learning and reasoning. Lecture Notes in Computer Science, 837:289– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Macchion:1994:HKB

[2172] D. J. Macchion and Dinh Phuoc Vo. A hybrid knowledge-based system for technical diagnosis learning and assistance. Lecture Notes in Computer Science, 837:301–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakatani:1994:TRC

[2173] Y. Nakatani and D. Israel. Tuning rules by cases. Lecture Notes in Computer Science, 837:313–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Portinale:1994:UCB

[2174] L. Portinale, P. Torasso, C. Ortalda, and A. Giardino. Using case-based reasoning to focus model-based diagnostic problem solving. Lecture Notes in Computer Science, 837:325–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ashley:1994:LRR

[2175] K. D. Ashley and V. Aleven. A logical representation for relevance criteria. Lecture Notes in Computer Science, 837:338–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schild:1994:MEP

[2176] U. J. Schild and Y. Kerner. Multiple explanation patterns. Lecture Notes in Computer Science, 837:353–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bradburn:1994:ACB

[2177] C. Bradburn and J. Zeleznikow. The application of case-based reasoning to the tasks of health care planning. Lecture Notes in Computer Science, 837: 365–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hurley:1994:PSM

[2178] N. Hurley. A priori selection of mesh densities for adaptive finite element analysis, using a case-based reasoning approach. Lecture Notes in Computer Science, 837:379–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamp:1994:ISS

[2179] G. Kamp. Integrating semantic structure and technical documentation in case-based service support systems. Lecture Notes in Computer Science, 837:392–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smail:1994:CBI

[2180] M. Smail. Case-based information retrieval. Lecture Notes in Computer Science, 837:404–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stadler:1994:CBR

[2181] M. Stadler. Case-based reasoning for network management. Lecture Notes in Computer Science, 837:414–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wendel:1994:CBR

[2182] O. Wendel. Case-based reasoning in a simulation environment for biological neural networks. Lecture Notes in Computer Science, 837:424–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hua:1994:ICB

[2183] K. Hua, I. Smith, and B. Faltings. Integrated case-based building design. Lecture Notes in Computer Science, 837: 436–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakakoji:1994:CDM

[2184] K. Nakakoji. Case-deliverer: Making cases relevant to the task at hand. Lecture Notes in Computer Science, 837:446–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Napoli:1994:FSC

[2185] A. Napoli and J. Lieber. A first study on case-based planning in organic synthesis. Lecture Notes in Computer Science, 837:458–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gottlob:1994:CML

[2186] G. Gottlob. From Carnap’s modal logic to autoepistemic logic. Lecture Notes in Computer Science, 838:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herre:1994:CPN

[2187] H. Herre. Compactness properties of nonmonotonic inference operations.

Lecture Notes in Computer Science, 838:19–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moinard:1994:APP

[2188] Y. Moinard and R. Rolland. Around a powerful property of circumscriptions. Lecture Notes in Computer Science, 838:34–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dimopoulos:1994:CVJ

[2189] Y. Dimopoulos. The computational value of joint consistency. Lecture Notes in Computer Science, 838:50–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aravindan:1994:BDA

[2190] C. Aravindan and P. M. Dung. Belief dynamics, abduction, and databases. Lecture Notes in Computer Science, 838:66–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Williams:1994:LTB

[2191] M.-A. Williams. On the logic of theory base change. Lecture Notes in Computer Science, 838:86–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alferes:1994:BPL

[2192] J. J. Alferes and L. M. Pereira. Belief, provability, and logic programs. Lecture Notes in Computer Science, 838: 106–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marek:1994:RSM

[2193] V. W. Marek and M. Truszczynski. Revision specifications by means of programs. Lecture Notes in Computer Science, 838:122–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Witteveen:1994:RNM

[2194] C. Witteveen, W. Van der Hoek, and H. De Nivelle. Revision of nonmonotonic theories: Some postulates and an application to logic programming. Lecture Notes in Computer Science, 838:137–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petermann:1994:CCC

[2195] U. Petermann. A complete connection calculus with rigid E-unification. Lecture Notes in Computer Science, 838: 152–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scherl:1994:ECR

[2196] R. Scherl. Equality and constrained resolution. Lecture Notes in Computer Science, 838:167–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demri:1994:ESA

[2197] S. Demri. Efficient strategies for automated reasoning in modal logics. Lecture Notes in Computer Science, 838: 182–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aguilera:1994:TDS

[2198] G. Aguilera, I. P. De Guzman, and M. Ojeda. TAS-D0+0+: Syntactic trees transformations for automated theorem proving. Lecture Notes in Computer Science, 838:198–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeNivelle:1994:UOR

[2199] H. De Nivelle. A unification of ordering refinements of resolution in classical logic. Lecture Notes in Computer Science, 838:217–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orlowska:1994:TLD

[2200] E. Orlowska. Two logical dimensions. Lecture Notes in Computer Science, 838:231–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rintanen:1994:PAL

[2201] J. Rintanen. Prioritized autoepistemic logic. Lecture Notes in Computer Science, 838:232–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brewka:1994:APS

[2202] G. Brewka. Adding priorities and specificity to default logic. Lecture Notes in Computer Science, 838:247–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chatalic:1994:VHT

[2203] P. Chatalic. Viewing hypothesis theories as constrained graded theories. Lecture Notes in Computer Science, 838:261–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engelfriet:1994:TTR

[2204] J. Engelfriet and J. Treur. Temporal theories of reasoning. Lecture Notes in Computer Science, 838:279–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Georgatos:1994:RAK

[2205] K. Georgatos. Reasoning about knowledge on computation trees. Lecture Notes in Computer Science, 838:300– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grosse:1994:PSE

[2206] G. Grosse. Propositional state event logic. Lecture Notes in Computer Science, 838:316–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeGiacomo:1994:DLI

[2207] G. De Giacomo and M. Lenzerini. Description logics with inverse roles, functional restrictions, and N-ary relations. Lecture Notes in Computer Science, 838:332–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertossi:1994:CGO

[2208] L. E. Bertossi and R. Reiter. On the concept of generic object: A nonmonotonic reasoning approach and examples. Lecture Notes in Computer Science, 838:347–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Przymusinski:1994:ALM

[2209] T. C. Przymusinski. Autoepistemic logic of minimal beliefs. Lecture Notes in Computer Science, 838:364– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nonnengart:1994:HUM

[2210] A. Nonnengart. How to use modalities and sorts in Prolog. Lecture Notes in Computer Science, 838:365–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacquet:1994:TRH

[2211] J.-M. Jacquet and L. Monteiro. Towards resource handling in logic programming: the PPL framework and its semantics. Lecture Notes in Computer Science, 838:379–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Costantini:1994:EHC

[2212] S. Costantini, P. Dell’Acqua, and G. A.

Lanzarone. Extending Horn clause theories by reflection principles. Lecture Notes in Computer Science, 838:400– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsui:1994:FEC

[2213] Mitsuru Matsui. The first experimental cryptanalysis of the Data Encryption Standard. Lecture Notes in Computer Science, 839:1–11, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/0839/08390001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0839/08390001.
pdf.

Ohta:1994:LCF

[2214] Kazuo Ohta and Kazumaro Aoki. Linear cryptanalysis of the Fast Data Encipherment Algorithm. Lecture Notes in Computer Science, 839:12– 16, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0839/08390012.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0839/08390012.
pdf.

Langford:1994:DLC

[2215] Susan K. Langford and Martin E. Hellman. Differential-linear cryptanalysis. Lecture Notes in Computer Science, 839:17–25, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0839/08390017.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0839/08390017.
pdf.

Kaliski:1994:LCU

[2216] Burton S. Kaliski Jr. and Matthew J. B. Robshaw. Linear cryptanalysis using multiple approximations. Lecture Notes in Computer Science, 839: 26–39, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0839/08390026.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/0839/08390026.
pdf.

Tillich:1994:HS

[2217] Jean-Pierre Tillich and Gilles Zémor.

Hashing with sl_{2}. Lecture Notes in Computer Science, 839:40–49, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0839/08390040.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0839/08390040.
pdf.

Chao:1994:DEC

[2218] Jinhui Chao, Kazuo Tanada, and Shigeo Tsujii. Design of elliptic curves with controllable lower boundary of extension degree for reduction attacks. Lecture Notes in Computer Science, 839:50–55, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0839/08390050.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0839/08390050.
pdf.

Biehl:1994:CPB

[2219] Ingrid Biehl, Johannes A. Buchmann, and Christoph Thiel. Cryptographic protocols based on discrete logarithms in real-quadratic orders. Lecture Notes in Computer Science, 839:56– 60, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0839/08390056.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0839/08390056.
                                        pdf.

Okamoto:1994:DCS

[2220] Tatsuaki Okamoto. Designated confirmer signatures and public-key encryption are equivalent. Lecture Notes in Computer Science, 839:61– 74, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0839/08390061.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0839/08390061.
                                        pdf.

Bleichenbacher:1994:DAG

[2221] Daniel Bleichenbacher and Ueli M. Maurer. Directed acyclic graphs, oneway functions and digital signatures. Lecture Notes in Computer Science, 839:75–82, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/0839/08390075.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/0839/08390075.
                                        pdf.

Delos:1994:IBS

[2222] Olivier Delos and Jean-Jacques Quisquater. An identity-based signature scheme with bounded life-span. Lecture Notes in Computer Science, 839:83– 94, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0839/08390083.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/0839/08390083.
pdf.

Lim:1994:MFE

[2223] Chae Hoon Lim and Pil Joong Lee. More flexible exponentiation with precomputation. Lecture Notes in Computer Science, 839:95–107, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0839/08390095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0839/08390095.
pdf.

Horvath:1994:PPM

[2224] Tamás Horváth, Spyros S. Magliveras, and Tran van Trung. A parallel permutation multiplier for a PGM crypto-chip. Lecture Notes in Computer Science, 839:108–113, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0839/08390108.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0839/08390108.
pdf.

Davis:1994:CRA

[2225] Don Davis, Ross Ihaka, and Philip Fenstermacher. Cryptographic randomness from air turbulence in disk drives. Lecture Notes in Computer Science, 839:114–120, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0839/08390114.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/0839/08390114.
                                        pdf.

Gehrmann:1994:CGN

[2226] Christian Gehrmann. Cryptanalysis of the Gemmell and Naor multiround authentication protocol. Lecture Notes in Computer Science, 839:121– 128, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0839/08390121.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0839/08390121.
                                        pdf.

Krawczyk:1994:LBH

[2227] Hugo Krawczyk. LFSR-based hashing and authentication. Lecture Notes in Computer Science, 839:129–139, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0839/08390129.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0839/08390129.
                                        pdf.

Kurosawa:1994:NBA

[2228] Kaoru Kurosawa. New bound on authentication code with arbitration. Lecture Notes in Computer Science, 839: 140–149, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/0839/08390140.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/0839/08390140.
                                        pdf.

Blundo:1994:MSS

[2229] Carlo Blundo, Alfredo De Santis, Giovanni Di Crescenzo, Antonio Giorgio Gaggia, and Ugo Vaccaro. Multisecret sharing schemes. Lecture Notes in Computer Science, 839:150–163, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0839/08390150.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0839/08390150.
pdf.

Stern:1994:DIS

[2230] Jacques Stern. Designing identification schemes with keys of short size. Lecture Notes in Computer Science, 839:164– 173, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0839/08390164.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0839/08390164.
pdf.

Cramer:1994:PPK

[2231] Ronald Cramer, Ivan Bjerre Damgård, and Berry Schoenmakers. Proof of partial knowledge and simplified design of witness hiding protocols. Lecture Notes in Computer Science, 839:174– 187, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0839/08390174.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0839/08390174.
pdf.

Itoh:1994:LDS

[2232] Toshiya Itoh, Yuji Ohta, and Hiroki Shizuya. Language dependent secure bit commitment. Lecture Notes in Computer Science, 839:188–201, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0839/08390188.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0839/08390188.
                                        pdf.

Girault:1994:LCH

[2233] Marc Girault and Jacques Stern. On the length of cryptographic hash-values used in identification schemes. Lecture Notes in Computer Science, 839:202– 215, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0839/08390202.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0839/08390202.
                                        pdf.

Bellare:1994:ICC

[2234] Mihir Bellare, Oded Goldreich, and Shafi Goldwasser. Incremental cryptography: The case of hashing and signing. Lecture Notes in Computer Science, 839:216–233, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0839/08390216.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0839/08390216.
                                        pdf.

Dwork:1994:EEU

[2235] Cynthia Dwork and Moni Naor. An efficient existentially unforgeable signature scheme and its applications. Lecture Notes in Computer Science, 839: 234–246, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0839/08390234.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0839/08390234.
pdf.

Bierbrauer:1994:BRF

[2236] Jürgen Bierbrauer, K. Gopalakrishnan, and Douglas R. Stinson. Bounds for resilient functions and orthogonal arrays. Lecture Notes in Computer Science, 839:247–256, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0839/08390247.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0839/08390247.
pdf.

Chor:1994:TT

[2237] Benny Chor, Amos Fiat, and Moni Naor. Tracing traitors. Lecture Notes in Computer Science, 839:257– 270, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0839/08390257.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0839/08390257.
pdf.

Maurer:1994:TEB

[2238] Ueli M. Maurer. Towards the equivalence of breaking the Diffie– Hellman protocol and computing discrete logarithms. Lecture Notes in Computer Science, 839:271–281, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0839/08390271.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0839/08390271.
                                        pdf.

Mihailescu:1994:FGP

[2239] Preda Mihailescu. Fast generation of provable primes using search in arithmetic progressions. Lecture Notes in Computer Science, 839:282–293, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0839/08390282.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0839/08390282.
                                        pdf.

Coppersmith:1994:ACS

[2240] Don Coppersmith. Attack on the cryptographic scheme NIKS–TAS. Lecture Notes in Computer Science, 839:294– 307, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0839/08390294.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0839/08390294.
                                        pdf.

Burmester:1994:ROD

[2241] Mike V. D. Burmester. On the risk of opening distributed keys. Lecture Notes in Computer Science, 839:308– 317, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0839/08390308.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0839/08390308.
pdf.

Beth:1994:CCB

[2242] Thomas Beth, D. E. Lazic, and A. Mathias. Cryptanalysis of cryptosystems based on remote chaos replication. Lecture Notes in Computer Science, 839:318–331, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0839/08390318.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0839/08390318.pdf.

Massey:1994:FTA

[2243] James L. Massey and Shirlei Serconek. A Fourier transform approach to the linear complexity of nonlinearly filtered sequences. Lecture Notes in Computer Science, 839:332–340, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0839/08390332.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0839/08390332.
pdf.

Bellare:1994:SCB

[2244] Mihir Bellare, Joe Kilian, and Phillip Rogaway. The security of cipher block chaining. Lecture Notes in Computer Science, 839:341–358, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0839/08390341.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0839/08390341.
pdf.

Gilbert:1994:CPA

[2245] Henri Gilbert and Pascal Chauvaud. A chosen plaintext attack of the 16round Khufu cryptosystem. Lecture Notes in Computer Science, 839:359– 368, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0839/08390359.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0839/08390359.
pdf.

Tsunoo:1994:COA

[2246] Yukiyasu Tsunoo, Eiji Okamoto, and Tomohiko Uyematsu. Ciphertext only attack for one-way function of the MAP using one ciphertext. Lecture Notes in Computer Science, 839:369– 382, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0839/08390369.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0839/08390369.
pdf.

Seberry:1994:PDS

[2247] Jennifer Seberry, Xian-Mo Zhang, and Yuliang Zheng. Pitfalls in designing substitution boxes. Lecture Notes in Computer Science, 839:383–396, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0839/08390383.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0839/08390383.
pdf.

Kushilevitz:1994:RRT

[2248] Eyal Kushilevitz and Adi Rosén. A randomness-rounds tradeoff in private computation. Lecture Notes in Computer Science, 839:397–410, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0839/08390397.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0839/08390397.
pdf.

Sako:1994:SVU

[2249] Kazue Sako and Joe Kilian. Secure voting using partially compatible homomorphisms. Lecture Notes in Computer Science, 839:411–424, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0839/08390411.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0839/08390411.
pdf.

Canetti:1994:MSP

[2250] Ran Canetti and Amir Herzberg. Maintaining security in the presence of transient faults. Lecture Notes in Computer Science, 839:425–438, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0839/08390425.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0839/08390425.
                                        pdf.

Anonymous:1994:AId

[2251] Anonymous. Author index. Lecture Notes in Computer Science, 839:439– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/0839/0839indx.pdf.

Reinelt:1994:I

[2252] Gerhard Reinelt. 1 introduction. Lecture Notes in Computer Science, 840: 1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0840/08400001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0840/08400001.
                                        pdf.

Reinelt:1994:BC

[2253] Gerhard Reinelt. 2 basic concepts. Lecture Notes in Computer Science, 840: 4–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/0840/08400004.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0840/08400004.
pdf.

Reinelt:1994:RPA

[2254] Gerhard Reinelt. 3 related problems and applications. Lecture Notes in Computer Science, 840:31– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0840/08400031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0840/08400031.
pdf.

Reinelt:1994:GC

[2255] Gerhard Reinelt. 4 geometric concepts.

Lecture Notes in Computer Science, 840:42–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0840/08400042.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0840/08400042.
pdf.

Reinelt:1994:CS

[2256] Gerhard Reinelt. 5 candidate sets. Lecture Notes in Computer Science, 840: 64–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0840/08400064.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0840/08400064.
pdf.

Reinelt:1994:CH

[2257] Gerhard Reinelt. 6 construction heuristics. Lecture Notes in Computer Science, 840:73–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0840/08400073.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0840/08400073.
                                        pdf.

Reinelt:1994:IS

[2258] Gerhard Reinelt. 7 improving solutions. Lecture Notes in Computer Science, 840:100–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/0840/08400100.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        0840/08400100.pdf.

Reinelt:1994:FHL

[2259] Gerhard Reinelt. 8 fast heuristics for large geometric problems. Lecture Notes in Computer Science, 840:133– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0840/08400133.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0840/08400133.
                                        pdf.

Reinelt:1994:FHA

[2260] Gerhard Reinelt. 9 further heuristic approaches. Lecture Notes in Computer Science, 840:153–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0840/08400153.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0840/08400153.
pdf.

Reinelt:1994:LB

[2261] Gerhard Reinelt. 10 lower bounds. Lecture Notes in Computer Science, 840: 161–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0840/08400161.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0840/08400161.
pdf.

Reinelt:1994:CST

[2262] Gerhard Reinelt. 11 A case study:

TSPs in PCB production. Lecture Notes in Computer Science, 840:187– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0840/08400187.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0840/08400187.
pdf.

Reinelt:1994:PTS

[2263] Gerhard Reinelt. 12 practical TSP solving. Lecture Notes in Computer Science, 840:200–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0840/08400200.
htm;   http://link.springer-ny.

                                        com/link/service/series/0558/papers/
                                        0840/08400200.pdf.

Reinelt:1994:AT

[2264] Gerhard Reinelt. Appendix: TSPLIB. Lecture Notes in Computer Science, 840:211–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/0840/08400211.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/0840/08400211.
                                        pdf.

Anonymous:1994:R

[2265] Anonymous. References. Lecture Notes in Computer Science, 840:214– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/0840/0840refs.pdf.

Anonymous:1994:I

[2266] Anonymous. Index. Lecture Notes in Computer Science, 840:222–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/0840/0840indx.pdf.

Reinelt:1994:TSC

[2267] G. Reinelt. The traveling salesman: Computational solutions for TSP applications. Lecture Notes in Computer Science, 840:viii + 223, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheatham:1994:MLC

[2268] T. Cheatham. Models, languages, and compiler technology for high performance computers. Lecture Notes in Computer Science, 841:3–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cosnard:1994:NRC

[2269] M. Cosnard and M. Matamala. On NC-Real complexity classes for additive circuits and their relations with NC. Lecture Notes in Computer Science, 841:27–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duris:1994:CSD

[2270] P. Duris and J. D. P. Rolim. Ecomplete sets do not have optimal polynomial time approximations. Lecture Notes in Computer Science, 841:38–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bloom:1994:SPF

[2271] S. L. Bloom and Z. Esik. Solving polynomial fixed point equations. Lecture Notes in Computer Science, 841:52–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Makowsky:1994:CCC

[2272] J. A. Makowsky. Capturing complexity classes with lindstroem quantifiers. Lecture Notes in Computer Science, 841:68–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monien:1994:CTI

[2273] B. Monien, R. Diekmann, and R. Lueling. Communication throughput of interconnection networks. Lecture Notes in Computer Science, 841:72–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lipton:1994:ICC

[2274] J. Lipton and M. J. O’Donnell. Intuitive counterexamples for constructive fallacies. Lecture Notes in Computer Science, 841:87–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orejas:1994:AMC

[2275] F. Orejas, E. Pino, and H. Ehrig. Algebraic methods in the compositional analysis of logic programs. Lecture Notes in Computer Science, 841:112– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paterson:1994:LCS

[2276] M. Paterson and V. Dancik. Longest common subsequences. Lecture Notes in Computer Science, 841:127–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pevzner:1994:TDS

[2277] P. A. Pevzner and R. J. Lipshutz. Towards DNA sequencing chips. Lecture Notes in Computer Science, 841:143– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sgall:1994:LSP

[2278] J. Sgall. On-line scheduling of parallel jobs. Lecture Notes in Computer Science, 841:159–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Winskel:1994:SBM

[2279] G. Winskel. Stable bistructure models of PCF. Lecture Notes in Computer Science, 841:177–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albers:1994:CAL

[2280] S. Albers. A competitive analysis of the list update problem with lookahead. Lecture Notes in Computer Science, 841:201–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alessi:1994:MMC

[2281] F. Alessi, M. Dezani-Ciancaglini, and U. De’Liguoro. May and must convergency in concurrent lambda-calculus. Lecture Notes in Computer Science, 841:211–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambos-Spies:1994:GME

[2282] K. Ambos-Spies, H.-C. Neis, and S. A.

Terwijn. Genericity and measure for exponential time. Lecture Notes in Computer Science, 841:221–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeza-Yates:1994:ABD

[2283] R. A. Baeza-Yates. Analysis of bounded disorder. Lecture Notes in Computer Science, 841:233–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertoni:1994:LAS

[2284] A. Bertoni, C. Mereghetti, and G. Pighizzini. On languages accepted with simultaneous complexity bounds and their ranking problem. Lecture Notes in Computer Science, 841:245–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBoer:1994:PSA

[2285] F. S. De Boer and M. Van Hulst. A proof system for asynchronously communicating deterministic processes. Lecture Notes in Computer Science, 841:256–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouabdallah:1994:ECB

[2286] A. Bouabdallah, M. C. Heydemann, J. Opatrny, and D. Sotteau. Embedding complete binary trees into star networks. Lecture Notes in Computer Science, 841:266–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damm:1994:ICB

[2287] C. Damm and M. Holzer. Inductive counting below LOGSPACE. Lecture Notes in Computer Science, 841:276– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dassow:1994:DPE

[2288] J. Dassow. Decision problems for edge grammars. Lecture Notes in Computer Science, 841:286–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dehne:1994:BSP

[2289] F. Dehne and R. Klein. “the big sweep”: On the power of the wavefront approach to Voronoi diagrams. Lecture Notes in Computer Science, 841: 296–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denejko:1994:RMF

[2290] P. Denejko, K. Diks, A. Pelc, and M. Piotrow. Reliable minimum finding comparator networks. Lecture Notes in Computer Science, 841:306–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dessmark:1994:PCM

[2291] A. Dessmark, A. Lingas, and O. Garrido. On parallel complexity of maximum f-matching and the degree sequence problem. Lecture Notes in Computer Science, 841:316–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dietzfelbinger:1994:CTL

[2292] M. Dietzfelbinger, J. Hromkovic, and G. Schnitger. A comparison of two lower bound methods for communication complexity. Lecture Notes in Computer Science, 841:326–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

EyonoObono:1994:ESW

[2293] S. Eyono Obono, P. Goralcik, and M. Maksimenko. Efficient solving of the word equations in one variable. Lecture Notes in Computer Science, 841:336–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferrari:1994:CES

[2294] G. L. Ferrari, U. Montanari, and P. Quaglia. A π-calculus with explicit substitutions: the late semantics. Lecture Notes in Computer Science, 841:342–351, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Forti:1994:PH

[2295] M. Forti, F. Honsell, and M. Lenisa. Processes and hyperuniverses. Lecture Notes in Computer Science, 841:352– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frayman:1994:CCF

[2296] F. Frayman, V. Kanevsky, and W. Kirchherr. The combinatorial complexity of a finite string. Lecture Notes in Computer Science, 841:364– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gold:1994:FAS

[2297] R. Gold. A fully abstract semantics for termination in Petri nets. Lecture Notes in Computer Science, 841:373– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

HalfeldFerrariAlves:1994:PAR

[2298] M. Halfeld Ferrari Alves, D. Laurent, and N. Spyratos. Passive and active rules in deductive databases. Lecture Notes in Computer Science, 841:383– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hegedus:1994:CRC

[2299] T. Hegedus. Combinatorial results on the complexity of teaching and learning. Lecture Notes in Computer Science, 841:393–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hermo:1994:DRE

[2300] M. Hermo. Degrees and reducibilities of easy tally sets. Lecture Notes in Computer Science, 841:403–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoogeboom:1994:MDT

[2301] H. J. Hoogeboom and P. Ten Pas.

MSO definable text languages. Lecture Notes in Computer Science, 841: 413–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hromkovic:1994:TLB

[2302] J. Hromkovic, J. Kari, L. Kari, and D. Pardubska. Two lower bounds on distributive generation of languages. Lecture Notes in Computer Science, 841:423–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inverardi:1994:RR

[2303] P. Inverardi and M. Venturini Zilli.

Rational rewriting. Lecture Notes in Computer Science, 841:433–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karhumaeki:1994:SIS

[2304] J. Karhumaeki and W. Plandowski. On the size of independent systems of equations in semigroups. Lecture Notes in Computer Science, 841:443– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karpinski:1994:STP

[2305] M. Karpinski and W. Rytter. On a sublinear time parallel construction of optimal binary search trees. Lecture Notes in Computer Science, 841:453– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kavvadias:1994:HED

[2306] D. Kavvadias, G. E. Pantziou, P. G. Spirakis, and C. D. Zaroliagis. Hammock-on-ears decomposition: A technique for the efficient parallel solution of shortest paths and other problems. Lecture Notes in Computer Science, 841:462–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klasing:1994:RBG

[2307] R. Klasing. The relationship between gossiping in vertex-disjoint paths mode and bisection width. Lecture Notes in Computer Science, 841:473–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurka:1994:CFC

[2308] P. Kůrka. A comparison of finite and cellular automata. Lecture Notes in Computer Science, 841:484–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lange:1994:EA

[2309] K.-J. Lange and K. Reinhardt. Empty alternation. Lecture Notes in Computer Science, 841:494–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeSaec:1994:MPO

[2310] B. Le Saec and I. Litovsky. On the minimization problem for omegaautomata. Lecture Notes in Computer Science, 841:504–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meinel:1994:CCO

[2311] C. Meinel and A. Slobodova. On the complexity of constructing optimal ordered binary decision diagrams. Lecture Notes in Computer Science, 841: 515–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Narayanan:1994:AAM

[2312] H. Narayanan, S. Roy, and S. Patkar.

Approximation algorithms for min-k-overlap problems using the principal lattice of partitions approach. Lecture Notes in Computer Science, 841:525– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeNicola:1994:CTN

[2313] R. De Nicola and A. Labella. A completeness theorem for nondeterministic Kleene algebras. Lecture Notes in Computer Science, 841:536–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papatriantafilou:1994:HRC

[2314] M. Papatriantafilou and P. Tsigas. How a rainbow colouring function can simulate wait-free handshaking. Lecture Notes in Computer Science, 841:546– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plump:1994:CPT

[2315] D. Plump. Critical pairs in term graph rewriting. Lecture Notes in Computer Science, 841:556–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Priami:1994:RWC

[2316] C. Priami and D. Yankelevich. Readwrite causality. Lecture Notes in Computer Science, 841:567–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Richomme:1994:STM

[2317] G. Richomme. Some trace monoids where both the star problem and the finite power property problem are decidable. Lecture Notes in Computer Science, 841:577–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salomaa:1994:CES

[2318] K. Salomaa, D. Wood, and S. Yu. Complexity of EOL structural equivalence. Lecture Notes in Computer Science, 841:587–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sibeyn:1994:SQP

[2319] J. F. Sibeyn, B. S. Chlebus, and M. Kaufmann. Shorter queues for permutation routing on meshes. Lecture Notes in Computer Science, 841:597– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sieber:1994:FAS

[2320] K. Sieber. Full abstraction for the second order subset of an ALGOL-like language. Lecture Notes in Computer Science, 841:608–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber:1994:FHP

[2321] A. Weber and T. Head. The finest homophonic partition and related code concepts. Lecture Notes in Computer Science, 841:618–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kloks:1994:TCA

[2322] T. Kloks. Treewidth: Computations and approximations. Lecture Notes in Computer Science, 842:ix + 209, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Szepietowski:1994:TMS

[2323] A. Szepietowski. Turing machines with sublogarithmic space. Lecture Notes in Computer Science, 843:viii + 114, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palamidessi:1994:CCPa

[2324] C. Palamidessi. Concurrent constraint programming. Lecture Notes in Computer Science, 844:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miller:1994:SUMa

[2325] D. Miller. Specifications using multiple-conclusion logic programs. Lecture Notes in Computer Science, 844:3–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paige:1994:VPTa

[2326] R. Paige. Viewing a program transformation system at work. Lecture Notes in Computer Science, 844:5–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chikayama:1994:PEI

[2327] T. Chikayama, T. Fujise, and D. Sekita. A portable and efficient implementation of KL1. Lecture Notes in Computer Science, 844:25–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1994:WMS

[2328] D. A. Smith. Why Multi-SLD beats SLD (even on a uniprocessor). Lecture Notes in Computer Science, 844:40–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zaffanella:1994:ASC

[2329] E. Zaffanella, R. Giacobazzi, and G. Levi. Abstracting synchronization in concurrent constraint programming. Lecture Notes in Computer Science, 844:57–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tarau:1994:NTC

[2330] P. Tarau and U. Neumerkel. A novel term compression scheme and data representation in the BinWAM. Lecture Notes in Computer Science, 844:73–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bevemyr:1994:SEC

[2331] J. Bevemyr and T. Lindgren. A simple and efficient copying garbage collector for Prolog. Lecture Notes in Computer Science, 844:88–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulte:1994:DRR

[2332] W. Schulte. Deriving residual reference count garbage collectors. Lecture Notes in Computer Science, 844:102– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pettersson:1994:RNL

[2333] M. Pettersson. RML — A new language and implementation for natural semantics. Lecture Notes in Computer Science, 844:117–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamel:1994:TPC

[2334] L. H. Hamel and J. A. Goguen. Towards a provably correct compiler for OBJ3. Lecture Notes in Computer Science, 844:132–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voronkov:1994:ITC

[2335] A. Voronkov. An implementation technique for a class of bottom-up procedures. Lecture Notes in Computer Science, 844:147–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glueck:1994:PDD

[2336] R. Glueck and M. H. Soerensen. Partial deduction and driving are equivalent. Lecture Notes in Computer Science, 844:165–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreau:1994:PCD

[2337] L. Moreau and C. Queinnec. Partial continuations as the difference of continuations. A duumvirate of control operators. Lecture Notes in Computer Science, 844:182–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Birkedal:1994:HPG

[2338] L. Birkedal and M. Welinder. Handwriting program generator generators. Lecture Notes in Computer Science, 844:198–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bugliesi:1994:LEO

[2339] M. Bugliesi and H. M. Jamil. A logic for encapsulation in object oriented languages. Lecture Notes in Computer Science, 844:215–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fruehwirth:1994:ACL

[2340] T. Fruehwirth. Annotated constraint logic programming applied to temporal reasoning. Lecture Notes in Computer Science, 844:230–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Codognet:1994:CBC

[2341] P. Codognet and D. Diaz. clp(B):

Combining simplicity and efficiency in Boolean constraint solving. Lecture Notes in Computer Science, 844:244– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fages:1994:HPC

[2342] F. Fages, J. Fowler, and T. Sola. Handling preferences in constraint logic programming with relational optimization. Lecture Notes in Computer Science, 844:261–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breitinger:1994:ISJ

[2343] S. Breitinger and H. C. R. Lock. Improving search for job-shop scheduling with CLP (FD). Lecture Notes in Computer Science, 844:277–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brogi:1994:IPC

[2344] A. Brogi, A. Chiarelli, P. Mancarella, and V. Mazzotta. Implementations of program composition operations. Lecture Notes in Computer Science, 844:292–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Codognet:1994:SVH

[2345] C. Codognet, P. Codognet, V. Loia, and M. Quaggetto. Sleepers: a versatile high-level control mechanism. Lecture Notes in Computer Science, 844: 308–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Metzemakers:1994:IAP

[2346] T. Metzemakers, A. Miniussi, D. Sherman, and R. Strandh. Improving arithmetic performance using fine-grain unfolding. Lecture Notes in Computer Science, 844:324–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gergatsoulis:1994:UFT

[2347] M. Gergatsoulis and M. Katzouraki. Unfold/fold transformations for definite clause programs. Lecture Notes in Computer Science, 844:340–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maraist:1994:GRT

[2348] J. Maraist and F. S. K. Silbermann. A graph reduction technique with sharing across narrowings for functional-logic languages. Lecture Notes in Computer Science, 844:355–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanus:1994:CLN

[2349] M. Hanus. Combining lazy narrowing and simplifications. Lecture Notes in Computer Science, 844:370–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arenas-Sanchez:1994:CLN

[2350] P. Arenas-Sanchez, A. Gil-Luezas, and F. J. Lopez-Fraguas. Combining lazy narrowing with disequality constraints. Lecture Notes in Computer Science, 844:385–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janssens:1994:ANV

[2351] G. Janssens, M. Bruynooghe, and V. Englebert. Abstracting numerical values in CLP(H,N). Lecture Notes in Computer Science, 844:400–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Field:1994:DDT

[2352] J. Field and F. Tip. Dynamic dependence in term rewriting systems and its application to program slicing. Lecture Notes in Computer Science, 844:415–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boulanger:1994:ASU

[2353] D. Boulanger, M. Bruynooghe, and M. Denecker. Abstracting S-semantics using A model-theoretic approach. Lecture Notes in Computer Science, 844: 432–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Serrano:1994:UHO

[2354] M. Serrano. Using higher order control flow analysis when compiling functional languages. Lecture Notes in Computer Science, 844:447–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kluzniak:1994:RRR

[2355] F. Kluzniak and M. Milkowska. Readable, runnable requirements specifications in SPILL-2. Lecture Notes in Computer Science, 844:449–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leal:1994:SVI

[2356] J. P. Leal. A system to visualize and interact with Prolog programs. Lecture Notes in Computer Science, 844: 451–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hundehege:1994:SBS

[2357] J.-B. Hundehege, W.-M. Lippe, and U. Xanten. Semantic-based static program transformations for memory space and run time optimization in hybrid languages. Lecture Notes in Computer Science, 844:453–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grivas:1994:TCF

[2358] G. Grivas. Towards a constraint functional logic compiler for symbolic computation systems. Lecture Notes in Computer Science, 844:455–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rodrigues:1994:PPM

[2359] O. Rodrigues and M. Benevides. PRO-MAL: Programming in modal action logic. Lecture Notes in Computer Science, 844:457–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bsaies:1994:ISP

[2360] K. Bsaies. Implementing the synthesis of properties in unfold/fold transformations. Lecture Notes in Computer Science, 844:459–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saenz:1994:SMS

[2361] F. Saenz, W. Hans, J. J. Ruz, and S. Winkler. Shared memory system for babel: A VHDL specification. Lecture Notes in Computer Science, 844:461–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belkhir:1994:TIF

[2362] A. Belkhir and N. Nemouche. Towards integrating functional and logic styles using relation. Lecture Notes in Computer Science, 844:463–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barklund:1994:TFD

[2363] J. Barklund. Tabulation of functions in definite clause programs. Lecture Notes in Computer Science, 844:465– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mantsivoda:1994:FSN

[2364] A. Mantsivoda. Flang system: A new version. Lecture Notes in Computer Science, 844:467–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Snyder:1994:ADC

[2365] W. Snyder. Automated deduction with constraints. Lecture Notes in Computer Science, 845:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lynch:1994:LS

[2366] C. Lynch. Local simplification. Lecture Notes in Computer Science, 845:3–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barth:1994:SCS

[2367] P. Barth. Simplifying clausal satisfiability problems. Lecture Notes in Computer Science, 845:19–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uribe:1994:OBD

[2368] T. E. Uribe and M. E. Stickel. Ordered binary decision diagrams and the Davis–Putnam procedure. Lecture Notes in Computer Science, 845:34–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smolka:1994:FHO

[2369] G. Smolka. A foundation for higherorder concurrent constraint programming. Lecture Notes in Computer Science, 845:50–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielson:1994:CPB

[2370] F. Nielson and A. R. Nielson. Constraints for polymorphic behaviours of concurrent ML. Lecture Notes in Computer Science, 845:73–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niehren:1994:CRC

[2371] J. Niehren and G. Smolka. A confluent relational calculus for higher-order programming with constraints. Lecture Notes in Computer Science, 845:89–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quian:1994:MAU

[2372] Z. Quian and K. Wang. Modular AC-Unification of higher-order patterns. Lecture Notes in Computer Science, 845:105–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lugiez:1994:HOD

[2373] D. Lugiez. Higher-order disunification:

Some decidable cases. Lecture Notes in Computer Science, 845:121–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boudet:1994:SAU

[2374] A. Boudet and E. Contejean. “syntactic” AC-Unification. Lecture Notes in Computer Science, 845:136–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simonis:1994:ACL

[2375] H. Simonis. Applications of constraint logic programming. Lecture Notes in Computer Science, 845:152–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michaylov:1994:RRI

[2376] S. Michaylov. Repeated redundant inequalities in constraint logic programming. Lecture Notes in Computer Science, 845:154–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsutsumi:1994:EAL

[2377] F. Tsutsumi. An efficient algorithm of logic programming with constraint hierarchy. Lecture Notes in Computer Science, 845:170–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beckmann:1994:ACL

[2378] R. Beckmann, U. Bieker, and I. Markhof. Application of constraint logic programming for VLSI CAD tools. Lecture Notes in Computer Science, 845:183–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Codognet:1994:PCC

[2379] P. Codognet and G. Nardiello. Path consistency in clp(FD). Lecture Notes in Computer Science, 845:201– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dauchet:1994:SCT

[2380] M. Dauchet. Symbolic constraints and tree automata. Lecture Notes in Computer Science, 845:217–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hense:1994:RCP

[2381] A. V. Hense and G. Smolka. A record calculus with principal types. Lecture Notes in Computer Science, 845:219– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haberstrau:1994:EEC

[2382] M. Haberstrau. ECOLOG: an environment for COnstraint LOGics. Lecture Notes in Computer Science, 845:237– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prehofer:1994:MTR

[2383] C. Prehofer. On modularity in term rewriting and narrowing. Lecture Notes in Computer Science, 845:253– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Avenhaus:1994:HOC

[2384] J. Avenhaus and C. Loria-Saenz.

Higher-order conditional rewriting and narrowing. Lecture Notes in Computer Science, 845:269–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bachmair:1994:BAC

[2385] L. Bachmair and H. Ganzinger. Buchberger’s algorithm: a constraint-based completion procedure. Lecture Notes in Computer Science, 845:285–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kozen:1994:SCL

[2386] D. Kozen. Set constraints and logic programming. Lecture Notes in Computer Science, 845:302–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charatonik:1994:SCS

[2387] W. Charatonik. Set constraints in some equational theories. Lecture Notes in Computer Science, 845:304–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Backofen:1994:HWG

[2388] R. Backofen and R. Treinen. How to win a game with features. Lecture Notes in Computer Science, 845:320– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gilleron:1994:SND

[2389] R. Gilleron, S. Tison, and M. Tommassi. Some new decidability results on positive and negative set constraints. Lecture Notes in Computer Science, 845:336–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johann:1994:SSO

[2390] P. Johann and R. Socher-Ambrosius. Solving simplification ordering constraints. Lecture Notes in Computer Science, 845:352–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McAuley:1994:OSS

[2391] D. R. McAuley. Operating system support for the desk area network. Lecture Notes in Computer Science, 846:13–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1994:ART

[2392] M. B. Jones. Adaptive real-time resource management supporting modular composition of digital multimedia services. Lecture Notes in Computer Science, 846:21–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfeifer:1994:MMN

[2393] T. Pfeifer. Micronet machines — new architectural approaches for multimedia end-systems. Lecture Notes in Computer Science, 846:29–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nieh:1994:SUS

[2394] J. Nieh, J. G. Hanko, J. D. Northcutt, and G. A. Wall. SVR4 UNIX scheduler unacceptable for multimedia applications. Lecture Notes in Computer Science, 846:41–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coulson:1994:ECM

[2395] G. Coulson, G. S. Blair, P. Robin, and D. Shepherd. Extending the Chorus micro-kernel to support continuous media applications. Lecture Notes in Computer Science, 846:54–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oikawa:1994:ULR

[2396] S. Oikawa and H. Tokuda. Userlevel real-time threads: An approach towards high performance multimedia threads. Lecture Notes in Computer Science, 846:66–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grinham:1994:IDP

[2397] J. Grinham and M. Spratt. IEEE 802.12 demand priority and multimedia. Lecture Notes in Computer Science, 846:77–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saran:1994:SDA

[2398] H. Saran, S. Keshav, C. R. Kalmanek, and S. P. Morgan. A scheduling discipline and admission control policy for Xunet 2. Lecture Notes in Computer Science, 846:89–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hyman:1994:MVV

[2399] J. M. Hyman, A. A. Lazar, and G. Pacifici. Modelling VC, VP and VN bandwidth assignment strategies in broadband networks. Lecture Notes in Computer Science, 846:102–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tokuda:1994:DQC

[2400] H. Tokuda and T. Kitayama. Dynamic QOS control based on real-time threads. Lecture Notes in Computer Science, 846:114–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Campbell:1994:MET

[2401] A. Campbell, G. Coulson, and D. Hutchison. A multimedia enhanced transport service in a quality of service architecture. Lecture Notes in Computer Science, 846:124–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fry:1994:FIN

[2402] M. Fry, A. Richards, and A. Seneviratne. Framework for implementing the next generation of communication protocols. Lecture Notes in Computer Science, 846:138–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stone:1994:QMD

[2403] D. L. Stone and K. Jeffay. Queue monitoring: A delay jitter management policy. Lecture Notes in Computer Science, 846:149–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keller:1994:PBD

[2404] R. Keller, W. Effelsberg, and B. Lamparter. Performance bottlenecks in digital movie systems. Lecture Notes in Computer Science, 846:161–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramanathan:1994:FIP

[2405] S. Ramanathan, P. V. Rangan, and H. M. Vin. Frame-induced packet discarding: An efficient strategy for video networking. Lecture Notes in Computer Science, 846:173–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hagsand:1994:ALS

[2406] O. Hagsand and S. Pink. ATM as a link in an ST-2 Internet. Lecture Notes in Computer Science, 846:185–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delgrossi:1994:RPI

[2407] L. Delgrossi, R. G. Herrtwich, C. Vogt, and L. C. Wolf. Reservation protocols for internetworks: A comparison of ST-II and RSVP. Lecture Notes in Computer Science, 846:195–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Little:1994:PAM

[2408] T. D. C. Little and D. Venkatesh. Probabilistic assignment of movies to storage devices in a video-on-demand system. Lecture Notes in Computer Science, 846:204–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramakrishnan:1994:OSS

[2409] K. K. Ramakrishnan, L. Vaitzblit, C. Gray, and U. Vahalia. Operating system support for a video-on-demand file service. Lecture Notes in Computer Science, 846:216–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keeton:1994:EVL

[2410] K. Keeton and R. H. Katz. The evaluation of video layout strategies on a high-bandwidth file server. Lecture Notes in Computer Science, 846:228– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoffman:1994:NSD

[2411] D. Hoffman, M. Speer, and G. Fernando. Network support for dynamically scaled multimedia data streams. Lecture Notes in Computer Science, 846:240–??, 1994. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1994:UFM

[2412] Z. Wang and J. Crowcroft. A unified framework for multicast forwarding. Lecture Notes in Computer Science, 846:252–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pasquale:1994:FPD

[2413] J. C. Pasquale, G. C. Polyzos, E. W.

Anderson, and V. P. Kompella. Filter propagation in dissemination trees: Trading off bandwidth and processing in continuous media networks. Lecture Notes in Computer Science, 846:259– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berenji:1994:FRL

[2414] H. R. Berenji. Fuzzy reinforcement learning and dynamic programming. Lecture Notes in Computer Science, 847:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldwin:1994:FSF

[2415] J. F. Baldwin. Fuzzy sets, fuzzy clustering and fuzzy rules in AI. Lecture Notes in Computer Science, 847:10–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saffiotti:1994:RER

[2416] A. Saffiotti, E. H. Ruspini, and K. Konolige. Robust execution of robot plans using fuzzy logic. Lecture Notes in Computer Science, 847:24–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouchon-Meunier:1994:FLI

[2417] B. Bouchon-Meunier, M. Ramdani, and L. Valverde. Fuzzy logic, inductive and analogical reasoning. Lecture Notes in Computer Science, 847:38–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1994:PTB

[2418] D. Dubois and H. Prade. Possibility theory, belief revision and nonmonotonic logic. Lecture Notes in Computer Science, 847:51–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldwin:1994:QDF

[2419] J. F. Baldwin, M. R. Coyne, and T. P. Martin. Querying a database with fuzzy attribute values by iterative updating of the selection criteria. Lecture Notes in Computer Science, 847:62–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Woods:1994:LFM

[2420] K. Woods, D. Cook, L. Hall, and L. Stark. Learning fuzzy membership functions in a function-based object recognition system. Lecture Notes in Computer Science, 847:77–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foong:1994:FSA

[2421] S. B. Foong and P. Z. Wang. A factor space approach to concept representation. Lecture Notes in Computer Science, 847:97–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kasabov:1994:CFP

[2422] N. K. Kasabov. Connectionist fuzzy production systems. Lecture Notes in Computer Science, 847:114–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krommer:1994:NIA

[2423] A. R. Krommer and C. W. Ueberhuber. Numerical integration on advanced computer systems. Lecture Notes in Computer Science, 848:xiii + 341, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hermann:1994:FMT

[2424] M. Hermann and W. Hoffmann. Fault modeling and test generation for FP-GAs. Lecture Notes in Computer Science, 849:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeOliveiraDuarte:1994:TMA

[2425] R. De Oliveira Duarte and M. Nicolaidis. A test methodology applied to cellular logic programmable gate arrays. Lecture Notes in Computer Science, 849:11–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Servit:1994:ILS

[2426] M. Z. Servit and Z. Muzikar. Integrated layout synthesis for FPGAs. Lecture Notes in Computer Science, 849:23–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robert:1994:ILB

[2427] M. Robert, L. Torres, F. Moraes, and D. Auvergne. Influence of logic block layout architecture on FPGA performance. Lecture Notes in Computer Science, 849:34–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haritaoglu:1994:GRH

[2428] I. Haritaoglu and C. Aykanat. A global routing heuristic for FPGAs based on mean field annealing. Lecture Notes in Computer Science, 849:45–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roy:1994:PDD

[2429] K. Roy and S. Prasad. Power dissipation driven FPGA place and route under delay constraints. Lecture Notes in Computer Science, 849:57–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farrahi:1994:FTM

[2430] A. H. Farrahi and M. Sarrafzadeh. FPGA technology mapping for power minimization. Lecture Notes in Computer Science, 849:66–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brand:1994:SSC

[2431] H.-J. Brand, D. Mueller, and W. Rosenstiel. Specification and synthesis of complex arithmetic operators for FP-GAs. Lecture Notes in Computer Science, 849:78–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miyazaki:1994:STS

[2432] T. Miyazaki, H. Nakada, A. Tsutsui, and K. Yamada. A speed-up technique for synchronous circuits realized as LUT-based FPGAs. Lecture Notes in Computer Science, 849:89–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naseer:1994:ETM

[2433] A. R. Naseer, M. Balakrishnan, and A. Kumar. An efficient technique for mapping RTL structures onto FP-GAs. Lecture Notes in Computer Science, 849:99–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamann:1994:TDM

[2434] V. Hamann. A testbench design method suitable for FPGA-Based prototyping of reactive systems. Lecture Notes in Computer Science, 849:111– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldberg:1994:UCC

[2435] E. Goldberg and L. Krasilnikova. Using consensusless covers for fast operating on Boolean functions. Lecture Notes in Computer Science, 849:114–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bartos:1994:FVT

[2436] T. Bartos and N. Fristacky. Formal verification of timing rules in design specifications. Lecture Notes in Computer Science, 849:117–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hlawiczka:1994:OSS

[2437] A. Hlawiczka and J. Binda. Optimized synthesis of self-testable finite state machines (FSM) using BIST-PST structures in altera structures. Lecture Notes in Computer Science, 849:120– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lichtermann:1994:HSR

[2438] J. Lichtermann and G. Neustaedter. A high-speed rotation processor. Lecture Notes in Computer Science, 849:123– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gramata:1994:MMD

[2439] P. Gramata, P. Trebaticky, and E. Gramatova. The MD5 messagedigest algorithm in the XILINX FPGA. Lecture Notes in Computer Science, 849:126–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fagin:1994:RPF

[2440] B. Fagin and P. Chintrakulchai. A reprogrammable processor for fractal image compression. Lecture Notes in Computer Science, 849:129–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jebelean:1994:IGS

[2441] T. Jebelean. Implementing GCD systolic arrays on FPGA. Lecture Notes in Computer Science, 849:132–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hughes:1994:FCT

[2442] R. B. Hughes and G. Musgrave. Formal CAD techniques for safety-critical FPGA design and deployment in embedded subsystems. Lecture Notes in Computer Science, 849:135–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saluvere:1994:DSS

[2443] T. Saluvere, D. Kerek, and H. Tenhunen. Direct sequence spread spectrum digital radio DSP prototyping using Xilinx FPGAs. Lecture Notes in Computer Science, 849:138–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguyen:1994:FBR

[2444] R. Nguyen and P. Nguyen. FPGA based reconfigurable architecture for a compact vision system. Lecture Notes in Computer Science, 849:141– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartenstein:1994:NFA

[2445] R. W. Hartenstein, R. Kress, and H. Reinig. A new FPGA architecture for word-oriented datapaths. Lecture Notes in Computer Science, 849:144– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Athanas:1994:IPC

[2446] P. M. Athanas and A. L. Abbott. Image processing on a custom computing platform. Lecture Notes in Computer Science, 849:156–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iseli:1994:SRP

[2447] C. Iseli and E. Sanchez. A superscalar and reconfigurable processor. Lecture Notes in Computer Science, 849:168– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salapura:1994:FFI

[2448] V. Salapura, M. Gschwind, and O. Maischberger. A fast FPGA implementation of a general purpose neuron. Lecture Notes in Computer Science, 849:175–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ast:1994:DPL

[2449] A. Ast, J. Becker, R. W. Hartenstein, and R. Kress. Data-procedural languages for FPL-Based machines. Lecture Notes in Computer Science, 849: 183–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daumas:1994:ILA

[2450] M. Daumas, J.-M. Muller, and J. Vuillemin. Implementing on line arithmetic on PAM. Lecture Notes in Computer Science, 849:196–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1994:SEW

[2451] X.-Y. Chen, X.-P. Ling, and H. Amano. Software environment for WASMII: a data driven machine with a virtual hardware. Lecture Notes in Computer Science, 849:208–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Newman:1994:CBH

[2452] M. Newman, W. Luk, and I. Page.

Constraint-based hierarchical placement of parallel programs. Lecture Notes in Computer Science, 849:220– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Njoelstad:1994:ZZL

[2453] T. Njoelstad, J. Pihl, and J. Hofstad.

ZAREPTA: A zero lead-time, all reconfigurable system for emulation, prototyping and testing of ASICs. Lecture Notes in Computer Science, 849:230– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wieler:1994:SSD

[2454] R. W. Wieler, Z. Zhang, and R. D.

McLeod. Simulating static and dynamic faults in BIST structures with a FPGA based emulator. Lecture Notes in Computer Science, 849:240– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benner:1994:FBP

[2455] T. Benner, R. Ernst, I. Koenenkamp, and U. Holtmann. FPGA based prototyping for verification and evaluation in hardware-software cosynthesis. Lecture Notes in Computer Science, 849:251–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dollas:1994:FBL

[2456] A. Dollas, B. Ward, and J. D. S. Babcock. FPGA based low cost generic reusable module for the rapid prototyping of subsystems. Lecture Notes in Computer Science, 849:259–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fawcett:1994:FDT

[2457] B. K. Fawcett and S. H. Kelem. FPGA development tools: Keeping pace with design complexity. Lecture Notes in Computer Science, 849:271–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kelem:1994:MBL

[2458] S. H. Kelem. Meaningful benchmarks for logic optimization of table-lookup FPGAs. Lecture Notes in Computer Science, 849:274–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1994:EUF

[2459] D. Lam. Educational use of field programmable gate arrays. Lecture Notes in Computer Science, 849:277– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fawcett:1994:HRF

[2460] B. K. Fawcett, N. Sawyer, and T. Williams. HardWire: A risk-free FPGA-to-ASIC migration path. Lecture Notes in Computer Science, 849: 280–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toon:1994:RHP

[2461] N. Toon. Reconfigurable hardware from programmable logic devices. Lecture Notes in Computer Science, 849:

283–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katona:1994:SLX

[2462] A. Katona and P. Szolgay. On some limits of XILINX based control logic implementations. Lecture Notes in Computer Science, 849:286–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cadek:1994:EUX

[2463] G. R. Cadek, P. C. Thorwartl, and G. P. Westphal. Experiences of using XBLOX for implementing a digital filter algorithm. Lecture Notes in Computer Science, 849:289–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toon:1994:CIP

[2464] N. Toon. Continuous interconnect provides solution to density/performance trade-off in programmable logic. Lecture Notes in Computer Science, 849: 292–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agrawal:1994:HDC

[2465] O. P. Agrawal. A high density complex PLD family optimized for flexibility, predictability and 100% routability. Lecture Notes in Computer Science, 849:295–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lysaght:1994:DEF

[2466] P. Lysaght, D. McConnell, and H. Dick.

Design experience with fine-grained FPGAs. Lecture Notes in Computer Science, 849:298–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leaver:1994:FRS

[2467] A. Leaver. FPGA routing structures from real circuits. Lecture Notes in Computer Science, 849:303–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klindworth:1994:TSS

[2468] A. Klindworth. A tool-set for simulating Altera-PLDs using VHDL. Lecture Notes in Computer Science, 849:306– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hallas:1994:CTD

[2469] J. A. Hallas, E. P. Mariatos, M. K. Birbas, and A. N. Birbas. A CAD tool for the development of an extra-fast fuzzy logic controller based on FPGAs and memory modules. Lecture Notes in Computer Science, 849:309–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cowen:1994:PCM

[2470] C. P. Cowen and S. Monaghan. Performance characteristics of the Monte-Carlo clustering processor (MCCP)

— a field programmable logic based custom computing machine. Lecture Notes in Computer Science, 849:312– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

BoegelVom:1994:DEE

[2471] G. Boegel Vom, P. Nauber, and J. Winkler. A design environment with emulation of prototypes for hardware/software systems using XILINX FPGA. Lecture Notes in Computer Science, 849:315–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Isoaho:1994:DDF

[2472] J. Isoaho, A. Jantsch, and H. Tenhunen. DSP development with fullspeed prototyping based on HW/SW codesign techniques. Lecture Notes in Computer Science, 849:318–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Danecek:1994:AGP

[2473] J. Danecek, A. Pluhacek, and M. Z.

Servit. The architecture of a generalpurpose processor cell. Lecture Notes in Computer Science, 849:321–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gschwind:1994:DSB

[2474] M. Gschwind and C. Mautner. The design of a stack-based microprocessor. Lecture Notes in Computer Science, 849:326–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akil:1994:IPE

[2475] M. Akil and M. Alves de Barros. Implementation and performance evaluation of an image pre-processing chain on FPGA. Lecture Notes in Computer Science, 849:332–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalosha:1994:STP

[2476] E. P. Kalosha, V. N. Yarmolik, and M. G. Karpovsky. Signature testability of PLA. Lecture Notes in Computer Science, 849:335–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MatBin:1994:FPP

[2477] I. Mat Bin and J. M. Noras. A FPL prototyping package with a C++ interface for the PC bus. Lecture Notes in Computer Science, 849:338–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rodriguez-Andina:1994:DSS

[2478] J. J. Rodriguez-Andina, J. Alvarez, and E. Mandado. Design of safety systems using field programmable gate arrays. Lecture Notes in Computer Science, 849:341–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Debize:1994:JDC

[2479] J. C. Debize and R. J. Glaise. A job dispatcher-collector made of FP-GAs for a centralized voice server. Lecture Notes in Computer Science, 849: 344–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Depreitere:1994:ODF

[2480] J. Depreitere, H. Neefs, H. Van Marck, and J. Van Campenhout. An optoelectronic 3-D field programmable gate array. Lecture Notes in Computer Science, 849:352–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roy:1994:CAR

[2481] K. Roy and S. Nag. On channel architecture and routability for FPGAs under faulty conditions. Lecture Notes in Computer Science, 849:361–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Isshiki:1994:HPD

[2482] T. Isshiki and W. W.-M. Dai. Highperformance datapath implementation on field-programmable multi-chip module (FPMCM). Lecture Notes in Computer Science, 849:373–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gehring:1994:LDD

[2483] S. Gehring, S. Ludwig, and N. Wirth.

A laboratory for a digital design course using FPGAs. Lecture Notes in Computer Science, 849:385–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer-Base:1994:CRD

[2484] U. Meyer-Base, A. Meyer-Baese, and W. Hilberg. Coordinate rotation digital computer (CORDIC) synthesis for FPGA. Lecture Notes in Computer Science, 849:397–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kempa:1994:MMN

[2485] G. J. Kempa and P. Rieger. MARC:

A Macintosh NUBUS-Expansion board based reconfigurable test system for validating communication systems. Lecture Notes in Computer Science, 849:409–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lysaght:1994:ANN

[2486] P. Lysaght, J. Stockwood, J. Law, and D. Girma. Artificial neural network implementation on a fine-grained FPGA. Lecture Notes in Computer Science, 849:421–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palamidessi:1994:CCPb

[2487] C. Palamidessi. Concurrent constraint programming. Lecture Notes in Computer Science, 850:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miller:1994:SUMb

[2488] D. Miller. Specifications using multiple-conclusion logic programs. Lecture Notes in Computer Science, 850:3–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paige:1994:VPTb

[2489] R. Paige. Viewing a program transformation system at work. Lecture Notes in Computer Science, 850:5–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Codish:1994:PIA

[2490] M. Codish and G. Mashevitzky. Proving implications by algebraic approximation. Lecture Notes in Computer Science, 850:6–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouhoula:1994:SCP

[2491] A. Bouhoula. Sufficient completeness and parameterized proofs by induction. Lecture Notes in Computer Science, 850:23–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bidoit:1994:PBT

[2492] M. Bidoit and R. Hennicker. Proving behavioural theorems with standard first-order logic. Lecture Notes in Computer Science, 850:41–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Werner:1994:HRL

[2493] A. Werner, A. Bockmayr, and S. Krischer. How to realize LSE narrowing. Lecture Notes in Computer Science, 850:59–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alpuente:1994:CAE

[2494] M. Alpuente, M. Falaschi, and G. Vidal. Compositional analysis for equational Horn programs. Lecture Notes in Computer Science, 850:77–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balbiani:1994:ESP

[2495] P. Balbiani. Equation solving in projective planes and planar ternary rings. Lecture Notes in Computer Science, 850:95–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bueno:1994:EAL

[2496] F. Bueno, M. Hermenegildo, U. Montanari, and F. Rossi. From eventual to atomic and locally atomic CC programs: A concurrent semantics. Lecture Notes in Computer Science, 850:

114–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Volpe:1994:CLP

[2497] P. Volpe. Concurrent logic programming as uniform linear proofs. Lecture Notes in Computer Science, 850:133– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teusink:1994:TVC

[2498] F. Teusink. Three-valued completion for abductive logic programs. Lecture Notes in Computer Science, 850:150– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoy:1994:SRS

[2499] S. Antoy and A. Middeldorp. A sequential reduction strategy. Lecture Notes in Computer Science, 850:168– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gramlich:1994:MTC

[2500] B. Gramlich. On modularity of termination and confluence properties of conditional rewrite systems. Lecture Notes in Computer Science, 850:186– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferreira:1994:SAT

[2501] M. C. F. Ferreira and H. Zantema. Syntactical analysis of total termination. Lecture Notes in Computer Science, 850:204–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marchiori:1994:LPT

[2502] M. Marchiori. Logic programs as term rewriting systems. Lecture Notes in Computer Science, 850:223–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1994:HOM

[2503] N. D. Jones and M. Rosendahl. Higherorder minimal function graphs. Lecture Notes in Computer Science, 850: 242–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kesner:1994:RAL

[2504] D. Kesner. Reasoning about layered, wildcard and product patterns. Lecture Notes in Computer Science, 850: 253–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bossi:1994:PUT

[2505] A. Bossi and N. Cocco. Prserving universal termination through unfold/fold. Lecture Notes in Computer Science, 850:269–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marchiori:1994:LVA

[2506] E. Marchiori. A logic for variable aliasing in logic programs. Lecture Notes in Computer Science, 850:287–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dickerson:1994:SIS

[2507] K. R. Dickerson and J. P. Chester. Standards for integrated services and networks. Lecture Notes in Computer Science, 851:5–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haallstrand:1994:IRS

[2508] J. Haallstrand and D. Martin. Industrial requirements on a service creation environment. Lecture Notes in Computer Science, 851:17–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keane:1994:DSS

[2509] D. Keane, C. Treanor, C. Licciardi, and T. Mota. Defining and structuring service requirements on the IN in order to evaluate IN architecture initiatives. Lecture Notes in Computer Science, 851:27–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MacanAirchinnigh:1994:SEV

[2510] M. Mac an Airchinnigh and H.-J. Kugler. Service engineering versus software engineering — A foundation study. Lecture Notes in Computer Science, 851:39–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hine:1994:IAP

[2511] N. A. Hine, W. Beattie, I. A. S. Gordon, and J. L. Arnott. Integration of adaptations for people with special needs. Lecture Notes in Computer Science, 851:51–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Androutsopoulos:1994:SPI

[2512] D. Androutsopoulos, P. Kaijser, S. Katsikas, and K. Presttun. Surveillance and protection in IBC management: The applicability of two RACE security projects — SecureNet II and SESAME. Lecture Notes in Computer Science, 851:61–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Condon:1994:MLS

[2513] C. Condon and S. Keuneke. Metaphors and layers of signification: The consequences for advanced user service interfaces. Lecture Notes in Computer Science, 851:75–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abramovici:1994:GUI

[2514] M. Abramovici and N. Klussmann.

Graphical user interface style guide for mobile communication services. Lecture Notes in Computer Science, 851: 89–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boecker:1994:AVS

[2515] M. Boecker, K. Hopf, and D. Runde.

Advanced videocommunications with stereoscopy and individual perspectives. Lecture Notes in Computer Science, 851:99–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haake:1994:HAD

[2516] J. M. Haake, J. Hannemann, and M. Thuering. Hypermedia applications for distance education and training. Lecture Notes in Computer Science, 851:109–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jakobs:1994:FCN

[2517] K. Jakobs. Future communication networks — what (public) service providers should provide. Lecture Notes in Computer Science, 851:123– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braux:1994:MSP

[2518] H. Braux, J. Dreteler, M. Grasdijk, and J.-L. Le Bail. Modelling services in the portfolio from a service provisioning perspective. Lecture Notes in Computer Science, 851:133–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bisseling:1994:EPC

[2519] H. Bisseling, J. Den Hartog, and B. Schrijver. Enhanced IN for personal communications. Lecture Notes in Computer Science, 851:145–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1994:CTT

[2520] H. Li. A component theory for telecommunication architectures. Lecture Notes in Computer Science, 851: 157–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haallstrand:1994:BDS

[2521] J. Haallstrand, M. M. Marques, and L. Ponten. Building dedicated service creation environments for reuse based production. Lecture Notes in Computer Science, 851:169–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carvalho:1994:MON

[2522] A. Carvalho, F. De Carvalho, and G. Pavlou. Management of optical networks: SNMP agents enabling TMN. Lecture Notes in Computer Science, 851:181–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bjerring:1994:RID

[2523] L. H. Bjerring and M. Tschichholz.

Requirements of inter-domain management and their implications for TMN architecture and implementation. Lecture Notes in Computer Science, 851: 193–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Donnelly:1994:SRT

[2524] W. Donnelly and S. O’Connell. Security requirements of the TMN Xinterface within end-to-end service management of virtual private networks. Lecture Notes in Computer Science, 851:207–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carr:1994:HLA

[2525] A. Carr, G. Pavlou, and T. Tin.

High-level access APIs in the OS-IMIS TMN platform: Harnessing and hiding. Lecture Notes in Computer Science, 851:219–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clarke:1994:URA

[2526] A. M. Clarke and S. M. Pomfrett. User requirements for advanced communication services. Lecture Notes in Computer Science, 851:235–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niebert:1994:PCW

[2527] N. Niebert and E. Geulen. Personal communications — what is beyond radio? Lecture Notes in Computer Science, 851:247–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davison:1994:SPM

[2528] R. Davison and P. O’Brien. Service provisioning in a multi-provider environment. Lecture Notes in Computer Science, 851:259–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DelaFuente:1994:ATC

[2529] L. A. De la Fuente, J. Pavon, and N. Singer. Application of TINA-C architecture to management services. Lecture Notes in Computer Science, 851:273–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beires:1994:EAT

[2530] N. Beires, T. Magedanz, and M. Kockelmans. An evolutionary approach for TMN management of IN services. Lecture Notes in Computer Science, 851: 285–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeCaluwe:1994:UTA

[2531] I. De Caluwe, P. Leever, and J. Wester.

The use of TMN as an architectural framework for value added services management. Lecture Notes in Computer Science, 851:295–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bjerring:1994:EES

[2532] L. H. Bjerring and J. M. Schneider.

End-to-end service management with multiple providers. Lecture Notes in Computer Science, 851:305–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonsson:1994:PCN

[2533] U. Jonsson and S. Kleier. Personal communication — network aspects and implementations. Lecture Notes in Computer Science, 851:321–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mudhar:1994:SCE

[2534] P. Mudhar. A service creation environment for a future intelligent network. Lecture Notes in Computer Science, 851:333–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prevedourou:1994:PSW

[2535] D. Prevedourou, G. D. Stamoulis, I. Toennby, and T. An. Providing services in a world of IBC resources: An architectural approach. Lecture Notes in Computer Science, 851:343– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berndt:1994:SSC

[2536] H. Berndt, P. Graubmann, and M. Wakano. Service specification concepts in TINA-C. Lecture Notes in Computer Science, 851:355–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carvalho:1994:ADT

[2537] P. Carvalho, M. Mesquita, L. Da Silva, and F. Roque. Application of distributed techniques to the modelling of IN architectures and support of service execution. Lecture Notes in Computer Science, 851:367–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lacoste:1994:DTP

[2538] G. Lacoste. Distributed transaction processing in the IBC. Lecture Notes in Computer Science, 851:377–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouma:1994:FPF

[2539] W. Bouma, W. Levelt, A. Melisse, and K. Middelburg. Formalisation of properties for feature interaction detection: Experience in a real-life situation. Lecture Notes in Computer Science, 851:393–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielsen:1994:DTM

[2540] P. S. Nielsen and B. Loenvig. Development of telecommunications management systems using OO methods and CASE tool support. Lecture Notes in Computer Science, 851:407–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Newcombe:1994:DSS

[2541] A. Newcombe, D. D. Botvitch, F. Lodge, and T. Curran. A decision support system for assurance of quality of service in intelligent network service provisioning. Lecture Notes in Computer Science, 851:419–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keller:1994:FDO

[2542] J. Keller and O. Dubuisson. Formal description of OSI management information structure as a prerequisite for formal specifications of TMN interfaces. Lecture Notes in Computer Science, 851:433–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lioupas:1994:SPT

[2543] Z. Lioupas, Y. Manolessos, and M. Theologou. Structuring principles for total quality management in IBCN. Lecture Notes in Computer Science, 851:445–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doyle:1994:SRD

[2544] D. C. Doyle and D. S. Cochrane.

Service/resource discovery and allocation. Lecture Notes in Computer Science, 851:455–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strick:1994:SPE

[2545] L. Strick and J. Meinkoehn. Specifying Pan-European management systems. Lecture Notes in Computer Science, 851:467–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dahle:1994:CMU

[2546] E. Dahle and P. L. Giganti. Computational modelling — UPT split charging management case. Lecture Notes in Computer Science, 851:479–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oelnes:1994:SUM

[2547] J. Oelnes, P. Spilling, and A. Berquist. Security in use and management of VPN — a banking scenario. Lecture Notes in Computer Science, 851:493– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baum-Waidner:1994:SSM

[2548] B. Baum-Waidner, H. Bunz, and C. Capellaro. SAMSON, security management in a health care scenario. Lecture Notes in Computer Science, 851: 503–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maillot:1994:SSS

[2549] D. Maillot, J. Oelnes, and P. Spilling. IN service security and service management security and their relationships — using UPT as a case study. Lecture Notes in Computer Science, 851: 513–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruno:1994:SDV

[2550] G. Bruno, F. Lucidi, J. Insulander, and U. Larsson. A service-driven vision of integrated broadband communications: the OSA approach. Lecture Notes in Computer Science, 851:529–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Redlich:1994:CFS

[2551] J.-P. Redlich and D. Balfanz. Concepts for a flexible service architecture. Lecture Notes in Computer Science, 851:539–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sullivan:1994:CPO

[2552] D. Sullivan and P. McLaughlin. A comparison of the PRISM and OMNI-Point methodologies for the specification of management systems. Lecture Notes in Computer Science, 851:553–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MacanAirchinnigh:1994:FMS

[2553] M. M. Mac an Airchinnigh, G. O’Regan, and D. Belsnes. Formal methods and service specification. Lecture Notes in Computer Science, 851:563–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hiltunen:1994:MAF

[2554] M. A. Hiltunen and R. D. Schlichting. A model for adaptive fault-tolerant systems. Lecture Notes in Computer Science, 852:3–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fabre:1994:DSR

[2555] J.-C. Fabre, Y. Deswartes, and B. Randell. Designing secure and reliable applications using fragmentationredundancy-scattering: An objectoriented approach. Lecture Notes in Computer Science, 852:21–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Silva:1994:FTM

[2556] J. G. Silva, L. M. Silva, H. Madeira, and J. Bernardino. A fault-tolerant mechanism for simple controllers. Lecture Notes in Computer Science, 852: 39–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernardeschi:1994:FRF

[2557] C. Bernardeschi, A. Fantechi, and L. Simoncini. Formal reasoning on fault coverage of fault tolerant techniques. Lecture Notes in Computer Science, 852:77–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chiaradonna:1994:PME

[2558] S. Chiaradonna, A. Bondavalli, and L. Strigini. On performability modeling and evaluation of software fault tolerance structures. Lecture Notes in Computer Science, 852:97–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antonelli:1994:ODF

[2559] C. Antonelli and V. Grassi. Optimal design of fault-tolerant soft-real-time systems with imprecise computations. Lecture Notes in Computer Science, 852:115–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bobbio:1994:CRS

[2560] A. Bobbio and M. Telek. Computational restrictions for SPN with generally distributed transition times. Lecture Notes in Computer Science, 852: 131–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ubar:1994:TGD

[2561] R. Ubar. Test generation for digital systems based on alternative graphs. Lecture Notes in Computer Science, 852: 151–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Renovell:1994:CRM

[2562] M. Renovell, P. Huc, and Y. Bertrand.

The configuration ratio: A model for simulating CMOS intra-gate bridge with variable logic thresholds. Lecture Notes in Computer Science, 852:165– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krasniewski:1994:CDF

[2563] A. Krasniewski and L. B. Wronski.

Coverage of delay faults: When 13% and 99% mean the same. Lecture Notes in Computer Science, 852:178– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Madeira:1994:RGP

[2564] H. Madeira, M. Rela, F. Moreira, and J. G. Silva. RIFLE: A general purpose pin-level fault injector. Lecture Notes in Computer Science, 852:199– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johansson:1994:SEU

[2565] R. Johansson. On single event upset error manifestation. Lecture Notes in Computer Science, 852:217–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhu:1994:IFE

[2566] H. Zhu, P. A. V. Hall, J. H. R. May, and T. Cockram. Injecting faults into environment simulators for testing safety critical software. Lecture Notes in Computer Science, 852:235–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thevenod-Fosse:1994:SST

[2567] P. Thevenod-Fosse, C. Mazuet, and Y. Crouzet. On statistical structural testing of synchronous data flow programs. Lecture Notes in Computer Science, 852:250–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Masseboeuf:1994:HTA

[2568] G. Masseboeuf, J. Pulou, and J. L. Rainard. Hierarchical test analysis of VLSI circuits for random BIST. Lecture Notes in Computer Science, 852: 271–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boehlau:1994:ZAC

[2569] P. Boehlau. Zero aliasing compression based on groups of weakly independent outputs in circuits with high complexity for two fault models. Lecture Notes in Computer Science, 852:

289–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lovric:1994:SDD

[2570] T. Lovric. Systematic and design diversity — software techniques for hardware fault detection. Lecture Notes in Computer Science, 852:309–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerber:1994:DPH

[2571] S. Gerber and M. Goessel. Detection of permanent hardware faults of a floating point adder by pseudoduplication. Lecture Notes in Computer Science, 852:327–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watanabe:1994:MML

[2572] A. Watanabe and K. Sakamura. MLDD (multi-layered design diversity) architecture for achieving high design fault tolerance capabilities. Lecture Notes in Computer Science, 852:336–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bieker:1994:RCM

[2573] B. Bieker, E. Maehle, G. Deconinck, and J. Vounckx. Reconfiguration and checkpointing in massively parallel systems. Lecture Notes in Computer Science, 852:353–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Altmann:1994:AHS

[2574] J. Altmann, F. Balbach, and A. Hein.

An approach for hierarchical system level diagnosis of massively parallel computers combined with a simulationbased method for dependability analysis. Lecture Notes in Computer Science, 852:371–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Majzik:1994:HCM

[2575] I. Majzik, A. Pataricza, M. Dal Cin, and W. Hohl. Hierarchical checking of multiprocessors using watchdog processors. Lecture Notes in Computer Science, 852:386–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laprie:1994:DCF

[2576] J.-C. Laprie. Dependability: The challenge for the future of computing and communication technologies. Lecture Notes in Computer Science, 852:407– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Randell:1994:SLS

[2577] B. Randell. Some lessons from the SW2000 workshop. Lecture Notes in Computer Science, 852:414–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmitter:1994:DCI

[2578] E. Schmitter. Dependable computing and its industrial use. Lecture Notes in Computer Science, 852:417–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1994:ERP

[2579] Y.-Y. Chen, C.-H. Cheng, and Y.-C. Chou. An effective reconfiguration process for fault-tolerant VLSI/WSI array processors. Lecture Notes in Computer Science, 852:421–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tahir:1994:CED

[2580] J. M. Tahir, S. S. Dlay, R. N. Gorgui-Naguib, and O. R. Hinton. Concurrent error detection in fast FNT networks. Lecture Notes in Computer Science, 852:439–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LaForge:1994:FRQ

[2581] L. E. LaForge. Feasible regions quantify the configuration power of arrays with multiple fault types. Lecture Notes in Computer Science, 852:453– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaaniche:1994:SRA

[2582] M. Kaaniche, K. Kanoun, M. Cukier, and M. Bastos Martini. Software reliability analysis of three successive generations of a switching system. Lecture Notes in Computer Science, 852: 473–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muller:1994:PCC

[2583] G. Muller, M. Hue, and N. Peyrouze. Performance of consistent checkpointing in a modular operating system: Results of the FTM experiment. Lecture Notes in Computer Science, 852:491–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Park:1994:RBN

[2584] J.-H. Park, H.-K. Lee, and J.-H. Cho. Ring-banyan network: A fault tolerant multistage interconnection network and its fault diagnosis. Lecture Notes in Computer Science, 852:511–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Avresky:1994:RFH

[2585] D. R. Avresky and K. M. Al-Tawil. Reconfiguration of faulty hypercubes. Lecture Notes in Computer Science, 852:529–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1994:FTB

[2586] C.-S. Yang and S.-Y. Wu. Faulttolerance on Boolean n-cube architectures. Lecture Notes in Computer Science, 852:546–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leu:1994:RSF

[2587] M. Leu. Relative signatures for fault tolerance and their implementation. Lecture Notes in Computer Science, 852:563–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Folliot:1994:GFT

[2588] B. Folliot and P. Sens. GATOSTAR: A fault tolerant load sharing facility for parallel applications. Lecture Notes in Computer Science, 852:581–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanderStok:1994:HMP

[2589] P. D. V. Van der Stok, M. M. M.

P. J. Claessen, and D. Alstein. A hierarchical membership protocol for synchronous distributed systems. Lecture Notes in Computer Science, 852:599– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karamcheti:1994:DFR

[2590] V. Karamcheti and A. A. Chien. Do faster routers imply faster communication? Lecture Notes in Computer Science, 853:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becker:1994:FAD

[2591] M. E. Becker and T. F. Knight. Fast arbitration in dilated routers. Lecture Notes in Computer Science, 853:16–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguyen:1994:PAM

[2592] Thu Duc Nguyen and L. Snyder. Performance analysis of a minimal adaptive router. Lecture Notes in Computer Science, 853:31–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duato:1994:PEA

[2593] J. Duato and P. Lopez. Performance evaluation of adaptive routing algorithms for k-ary n-cubes. Lecture Notes in Computer Science, 853:45–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chinn:1994:PAR

[2594] D. D. Chinn. The performance of adaptive routers on worst case permutations. Lecture Notes in Computer Science, 853:60–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Konstantinidou:1994:EQS

[2595] S. Konstantinidou. On the effect of queues sizes and channel scheduling policies in the segment router. Lecture Notes in Computer Science, 853:72–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wills:1994:OCO

[2596] D. Scott Wills, W. Stephen Lacy, and J. Cruz-Rivera. The offset cube: An optoelectronic interconnection network. Lecture Notes in Computer Science, 853:86–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguyen:1994:TDN

[2597] J. Nguyen, J. Pezaris, G. Pratt, and S. Ward. Three-dimensional network topologies. Lecture Notes in Computer Science, 853:101–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rexford:1994:SMC

[2598] J. Rexford and K. G. Shin. Support for multiple classes of traffic in multicomputer routers. Lecture Notes in Computer Science, 853:116–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Panda:1994:MMP

[2599] D. K. Panda, S. Singal, and P. Prabhakaran. Multidestination message passing mechanism conforming to base wormhole routing scheme. Lecture Notes in Computer Science, 853:131– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chiang:1994:MAE

[2600] C.-M. Chiang and L. M. Ni. Multiaddress encoding for multicast. Lecture Notes in Computer Science, 853: 146–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abali:1994:RAI

[2601] B. Abali and C. Aykanat. Routing algorithms for IBM SP1. Lecture Notes in Computer Science, 853:161–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heller:1994:CFR

[2602] S. Heller. Congestion-free routing on the CM-5 data router. Lecture Notes in Computer Science, 853:176–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nesson:1994:RRC

[2603] T. Nesson and L. Johnsson. ROMM routing: A class of efficient minimal routing algorithms. Lecture Notes in Computer Science, 853:185–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chong:1994:PMH

[2604] F. T. Chong, E. A. Brewer, F. T. Leighton, and T. F. Knight. Packaging and multiplexing of hierarchical scalable expanders. Lecture Notes in Computer Science, 853:200–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eslick:1994:GIT

[2605] I. Eslick, A. DeHon, and T. Knight. Guaranteeing idempotence for tightlycoupled, fault-tolerant networks. Lecture Notes in Computer Science, 853: 215–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bolding:1994:DRF

[2606] K. Bolding and W. Yost. Design of a router for fault-tolerant networks. Lecture Notes in Computer Science, 853: 226–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dally:1994:RRR

[2607] W. J. Dally, L. R. Dennison, D. Harris, and K. Kan. The reliable router: A reliable and high-performance communication substrate for parallel computers. Lecture Notes in Computer Science, 853:241–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubnicki:1994:NIS

[2608] C. Dubnicki, K. Li, and M. Mesarina. Network interface support for user-level buffer management. Lecture Notes in Computer Science, 853:256–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McKenzie:1994:CIM

[2609] N. R. McKenzie, K. Bolding, C. Ebeling, and L. Snyder. CRANIUM: An interface for message passing on adaptive packet routing networks. Lecture Notes in Computer Science, 853:266– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scott:1994:ORC

[2610] S. Scott and G. Thorson. Optimized routing in the Cray T3D. Lecture Notes in Computer Science, 853:281– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davis:1994:RDA

[2611] A. Davis, R. Hodgson, I. Robinson, and L. Cherkasova. R2: A damped adaptive router design. Lecture Notes in Computer Science, 853:295–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boughton:1994:ARC

[2612] G. A. Boughton. Arctic routing chip. Lecture Notes in Computer Science, 853:310–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hwang:1994:SPM

[2613] K. Hwang. Scalability and programmability of massively parallel processors. Lecture Notes in Computer Science, 854:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foster:1994:LCR

[2614] I. Foster and C. Kesselman. Language constructs and runtime systems for compositional parallel programming. Lecture Notes in Computer Science, 854:5–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1994:LOP

[2615] M. S. Lam. Locality optimizations for parallel machines. Lecture Notes in Computer Science, 854:17–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohr:1994:TPP

[2616] B. Mohr, D. Brown, and A. Malony. TAU: A portable parallel program analysis environment for pC++. Lecture Notes in Computer Science, 854: 29–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wabnig:1994:CPB

[2617] H. Wabnig, G. Haring, D. Kranzlmueller, and J. Volkert. Communication pattern based performance prediction on the nCUBE 2 multiprocessor system. Lecture Notes in Computer Science, 854:41–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferscha:1994:POD

[2618] A. Ferscha and J. Johnson. Performance oriented development of SPMD programs based on task structure specifications. Lecture Notes in Computer Science, 854:53–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grabner:1994:MDB

[2619] S. Grabner and D. Kranzlmueller.

Monitoring for detecting bugs and blocking communication. Lecture Notes in Computer Science, 854:66–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levrouw:1994:MLS

[2620] L. J. Levrouw and K. M. R. Audenaert. Minimizing the log size for execution replay of shared-memory programs. Lecture Notes in Computer Science, 854:76–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kastner:1994:GMS

[2621] W. Kastner, S. Stoeckler, and R. Zeitlberger. GOLD MINE: Specification language for event-based monitoring. Lecture Notes in Computer Science, 854:88–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouge:1994:EWP

[2622] L. Bouge, Y. Le Guyadec, G. Uiard, and B. Virot. On the expressivity of a weakest precondition calculus for a simple data-parallel programming language. Lecture Notes in Computer Science, 854:100–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Violard:1994:RP

[2623] E. Violard and G.-R. Perrin. Reduction in PEI. Lecture Notes in Computer Science, 854:112–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arvind:1994:TSA

[2624] D. K. Arvind and B. R. J. McConnell. Towards the synthesis of asynchronous concurrent programs. Lecture Notes in Computer Science, 854:124–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geerling:1994:FDS

[2625] A. M. Geerling. Formal derivation of SIMD parallelism from non-linear recursive specifications. Lecture Notes in Computer Science, 854:136–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fitzpatrick:1994:DEP

[2626] S. Fitzpatrick, T. J. Harmer, and J. M. Boyle. Deriving efficient parallel implementations of algorithms operating on general sparse matrices using automatic program transformation. Lecture Notes in Computer Science, 854:148– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kindermann:1994:FPA

[2627] S. Kindermann. Flexible program and architecture specification for massively parallel systems. Lecture Notes in Computer Science, 854:160–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sedukhin:1994:SAS

[2628] S. G. Sedukhin and I. S. Sedukhin. Systematic approach and software tool for systolic design. Lecture Notes in Computer Science, 854:172–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lau:1994:AAD

[2629] L. Lau, M. Rezny, J. Belward, and K. Burrage. ADVISE — agricultural developmental visualisation interactive software environment. Lecture Notes in Computer Science, 854:184–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Detert:1994:TTS

[2630] U. Detert and M. Gerndt. TOP02 tool suite for the development and testing of parallel applications. Lecture Notes in Computer Science, 854:196–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amor:1994:VRR

[2631] M. Amor, M. J. Martin, D. Blanco, and O. G. Plata. Vectorization of the radix r self-sorting FFT. Lecture Notes in Computer Science, 854:208–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clint:1994:EGS

[2632] M. Clint, J. S. Weston, and J. B.

Flannagan. Efficient Gram–Schmidt orthogonalisation on an array processor. Lecture Notes in Computer Science, 854:218–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stojcev:1994:PAI

[2633] M. K. Stojcev, E. I. Milovanovic, M. D. Mihajlovic, and I. Z. Milovanovic. Parallel algorithm for inverting tridiagonal matrix on linear processor array. Lecture Notes in Computer Science, 854:229–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jebelean:1994:SAL

[2634] T. Jebelean. Systolic algorithms for long integer GCD computation. Lecture Notes in Computer Science, 854: 241–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiMartino:1994:PNS

[2635] B. Di Martino and G. Iannello. Parallelization of non-simultaneous iterative methods for systems of linear equations. Lecture Notes in Computer Science, 854:253–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauber:1994:HIP

[2636] T. Rauber and G. Ruenger. Hypercube implementation and performance analysis for extrapolation models. Lecture Notes in Computer Science, 854:265–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zimmermann:1994:AMI

[2637] W. Zimmermann and W. Loewe. An approach to machine-independent parallel programming. Lecture Notes in Computer Science, 854:277–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prestwich:1994:LPP

[2638] S. Prestwich. On logic program parallelisation strategies. Lecture Notes in Computer Science, 854:289–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akl:1994:PAS

[2639] S. G. Akl and L. F. Lindon. Paradigms admitting superunitary behaviour in parallel computation. Lecture Notes in Computer Science, 854:301–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Limongelli:1994:ESL

[2640] C. Limongelli and R. Pirastu. Exact solution of linear systems over rational numbers by parallel p-adic arithmetic. Lecture Notes in Computer Science, 854:313–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hong:1994:PCM

[2641] H. Hong and H. W. Loidl. Parallel computation of modular multivariate polynomial resultants on a shared memory machine. Lecture Notes in Computer Science, 854:325–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tanaka:1994:PMG

[2642] Y. Tanaka, S. Matsui, A. Maeda, and M. Nakanishi. Partial marking GC. Lecture Notes in Computer Science, 854:337–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nikolovoulos:1994:PRL

[2643] S. D. Nikolovoulos. Parallel recognition and location algorithms for chordal graphs using distance matrices. Lecture Notes in Computer Science, 854: 349–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dion:1994:CTP

[2644] M. Dion, M. Gengler, and S. Ubeda. Comparing two probabilistic models of the computational complexity of the branch and bound algorithm. Lecture Notes in Computer Science, 854:359– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1994:PHC

[2645] J. Chen. Parallel heap construction using multiple selection. Lecture Notes in Computer Science, 854:371–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gorton:1994:EHQ

[2646] I. Gorton, T. S. Chan, and I. Jelly. Engineering high quality parallel software using PARSE. Lecture Notes in Computer Science, 854:381–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spalt:1994:MPV

[2647] A. Spalt, E. Spiegl, and T. Meikl. Massively parallel volume rendering. Lecture Notes in Computer Science, 854: 393–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stuerzlinger:1994:PVC

[2648] W. Stuerzlinger and C. Wild. Parallel visibility computations for parallel radiosity. Lecture Notes in Computer Science, 854:405–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valero:1994:MAS

[2649] M. Valero, M. Peiron, and E. Ayguade.

Memory access synchronization in vector multiprocessors. Lecture Notes in Computer Science, 854:414–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yousif:1994:SCD

[2650] M. S. Yousif and C. R. Das. A switch cache design for MIN-Based shared-memory multiprocessors. Lecture Notes in Computer Science, 854: 426–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kattner:1994:MCC

[2651] R. Kattner, M. Eger, and C. Mueller-Schloer. Modeling cache coherence overhead with geometric objects. Lecture Notes in Computer Science, 854: 438–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luepke:1994:AAS

[2652] S. Luepke. Accelerated access to shared distributed arrays on distributed memory systems by access objects. Lecture Notes in Computer Science, 854:449– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berrendorf:1994:CSV

[2653] R. Berrendorf, M. Gerndt, Z. Lahjomri, and T. Priol. A comparison of shared virtual memory and message passing programming techniques based on a finite element application. Lecture Notes in Computer Science, 854:461– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dohmen:1994:CBM

[2654] R. Dohmen. Comparison between a message passing and a global shared memory parallel computer by means of a concrete Monte-Carlo algorithm. Lecture Notes in Computer Science, 854: 473–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Milikowski:1994:NHP

[2655] R. Milikowski and W. G. Vree. Nonhomogeneous parallel memory operations in a VLIW machine. Lecture Notes in Computer Science, 854:485– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buelck:1994:EIC

[2656] T. Buelck, A. Held, W. Kluge, and S. Pantke. Experience with the implementation of a concurrent graph reduction system on an nCube/2 platform. Lecture Notes in Computer Science, 854:497–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lincoln:1994:RRM

[2657] P. Lincoln, J. Meseguer, and L. Ricciulli. The rewrite rule machine node architecture and its performance. Lecture Notes in Computer Science, 854: 509–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hammond:1994:ASS

[2658] K. Hammond, J. Mattson, and S. L.

Peyton Jones. Automatic spark strategies and granularity for a parallel functional language reducer. Lecture Notes in Computer Science, 854:521– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schreiner:1994:VTP

[2659] W. Schreiner. Virtual tasks for the PACLIB kernel. Lecture Notes in Computer Science, 854:533–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeGall:1994:MMP

[2660] J.-Y. Le Gall and F. Simon. MUSE: a message passing concurrent computer for on-board space systems. Lecture Notes in Computer Science, 854:545– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hogg:1994:STC

[2661] R. S. Hogg, D. W. Lloyd, and W. I.

Hughes. Self-timed communication strategies for massively parallel systolic architectures. Lecture Notes in Computer Science, 854:557–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feng:1994:DLT

[2662] M. D. Feng, Y. Q. Gao, and C. K.

Yuen. Distributed Linda tuplespace algorithms and implementations. Lecture Notes in Computer Science, 854:581– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bailey:1994:SCF

[2663] P. Bailey, M. Newey, D. Sitsky, and R. Stanton. Supporting coarse and fine grain parallelism in an extension of ML. Lecture Notes in Computer Science, 854:593–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calvin:1994:TMC

[2664] C. Calvin, L. Colombet, F. Desprez, and B. Jargot. Towards mixed computation/communication in parallel scientific libraries. Lecture Notes in Computer Science, 854:605–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bagley:1994:CFU

[2665] R. A. Bagley, P. A. Wilsey, and N. B. Abu-Ghazaleh. Composing functional unit blocks for efficient interpretation of MIMD code sequences on SIMD processors. Lecture Notes in Computer Science, 854:616–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Llosa:1994:USO

[2666] J. Llosa, M. Valero, J. A. B. Fortes, and E. Ayguade. Using sacks to organize registers in VLIW machines. Lecture Notes in Computer Science, 854: 628–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Govindarajan:1994:FRC

[2667] R. Govindarajan, E. R. Altman, and G. R. Gao. A framework for resource-constrained rate-optimal software pipelining. Lecture Notes in Computer Science, 854:640–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berthou:1994:DMP

[2668] J.-Y. Berthou. Design of a metaparallelizer for large scientific applications. Lecture Notes in Computer Science, 854:652–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chapman:1994:SAM

[2669] B. Chapman, P. Mehrotra, J. V.

Rosendale, and H. Zima. A software architecture for multidisciplinary applications: Integrating task and data parallelism. Lecture Notes in Computer Science, 854:664–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Griebl:1994:SST

[2670] M. Griebl and C. Lengauer. On scanning space-time mapped while loops. Lecture Notes in Computer Science, 854:677–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Werth:1994:OCU

[2671] M. R. Werth and P. Feautrier. Optimizing communications by using compile time analysis. Lecture Notes in Computer Science, 854:689–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andonov:1994:OTS

[2672] R. Andonov and S. Rajopadhye. Optimal tile sizing. Lecture Notes in Computer Science, 854:701–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouchitte:1994:EAE

[2673] V. Bouchitte, P. Boulet, A. Darte, and Y. Robert. Evaluating array expressions on massively parallel machines with communication/ computation overlap. Lecture Notes in Computer Science, 854:713–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johnson:1994:UIS

[2674] S. P. Johnson, C. S. Ierotheou, M. Cross, and P. F. Leggett. User interaction and symbolic extensions to dependence analysis. Lecture Notes in Computer Science, 854:725–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maslov:1994:SPC

[2675] V. Maslov and W. Pugh. Simplifying polynomial constraints over integers to make dependence analysis more precise. Lecture Notes in Computer Science, 854:737–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oehring:1994:AOC

[2676] S. R. Oehring, S. K. Das, and D. H. Hohndel. Asymptotically optimal communication algorithms on faulty folded Petersen networks. Lecture Notes in Computer Science, 854:749–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trobec:1994:ODE

[2677] R. Trobec and I. Jerebic. Optimization of diagnostic examination. Lecture Notes in Computer Science, 854: 761–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1994:LMN

[2678] G. Chen and F. C. M. Lau. Laying out midimew networks with constant dilation. Lecture Notes in Computer Science, 854:773–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fraigniaud:1994:OIR

[2679] P. Fraigniaud and C. Gavoille. Optimal interval routing. Lecture Notes in Computer Science, 854:785–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferreira:1994:BBI

[2680] A. Ferreira, A. Goldman vel Lejbman, and S. W. Song. Broadcasting in bus interconnection networks. Lecture Notes in Computer Science, 854:797– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deserable:1994:BAT

[2681] D. Deserable. Broadcasting in the arrowhead torus. Lecture Notes in Computer Science, 854:808–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haritaoglu:1994:EMH

[2682] I. Haritaoglu and C. Aykanat. An efficient mapping heuristic for meshconnected parallel architectures based on mean field annealing. Lecture Notes in Computer Science, 854:820– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwehm:1994:MSG

[2683] M. Schwehm and T. Walter. Mapping and scheduling by genetic algorithms. Lecture Notes in Computer Science, 854:832–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer:1994:SOP

[2684] J. W. Meyer. Self-organizing processes. Lecture Notes in Computer Science, 854:842–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeSergent:1994:BLU

[2685] T. Le Sergent and B. Berthomieu. Balancing load under large and fast load changes in distributed computing systems. Lecture Notes in Computer Science, 854:854–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elleuch:1994:DLB

[2686] A. Elleuch, R. Kanawati, T. Muntean, and E.-G. Talbi. Dynamic load balancing mechanisms for a parallel operating system kernel. Lecture Notes in Computer Science, 854:866–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roch:1994:NLP

[2687] J. L. Roch, A. Vermeerbergen, and G. Villard. A new load-prediction scheme based on algorithmic cost functions. Lecture Notes in Computer Science, 854:878–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldberg:1994:OAL

[2688] A. V. Goldberg. Optimization algorithms for large networks. Lecture Notes in Computer Science, 855:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gonnet:1994:WBA

[2689] G. H. Gonnet. Wanna buy an algorithm? cheap! or: Algorithms for text searching which could have commercial value. Lecture Notes in Computer Science, 855:10–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garg:1994:PDA

[2690] A. Garg and R. Tamassia. Planar drawings and angular resolution: Algorithms and bounds. Lecture Notes in Computer Science, 855:12–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biedl:1994:BHO

[2691] T. Biedl and G. Kant. A better heuristic for orthogonal graph drawings. Lecture Notes in Computer Science, 855:24–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arkin:1994:HTF

[2692] E. M. Arkin, M. Held, J. S. B. Mitchell, and S. S. Skiena. Hamiltonian triangulations for fast rendering. Lecture Notes in Computer Science, 855:36–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arya:1994:ECB

[2693] S. Arya and M. Smid. Efficient construction of a bounded degree spanner with low weight. Lecture Notes in Computer Science, 855:48–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berman:1994:AAR

[2694] P. Berman, U. Foessmeier, M. Karpinski, and M. Kaufmann. Approaching the 5/4-Approximation for rectilinear Steiner trees. Lecture Notes in Computer Science, 855:60–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brodnik:1994:MCT

[2695] A. Brodnik and J. I. Munro. Membership in constant time and minimum space. Lecture Notes in Computer Science, 855:72–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersson:1994:FST

[2696] A. Andersson and S. Nilsson. Faster searching in tries and quadtrees — an analysis of level compression. Lecture Notes in Computer Science, 855:82–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poblete:1994:AHS

[2697] P. V. Poblete, A. Viola, and J. I. Munro. The analysis of a hashing scheme by the diagonal Poisson transform. Lecture Notes in Computer Science, 855:94–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carlsson:1994:SLB

[2698] S. Carlsson and J. Chen. Some lower bounds for comparison-based algorithms. Lecture Notes in Computer Science, 855:106–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1994:EAE

[2699] X. Zhou and T. Nishizeki. An efficient algorithm for edge-ranking trees. Lecture Notes in Computer Science, 855: 118–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weihe:1994:EDP

[2700] K. Weihe. Edge-disjoint (s, t)-paths in undirected planar graphs in linear time. Lecture Notes in Computer Science, 855:130–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stoer:1994:SMC

[2701] M. Stoer and F. Wagner. A simple min cut algorithm. Lecture Notes in Computer Science, 855:141–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cho:1994:AAM

[2702] J. D. Cho, S. Raje, and M. Sarrafzadeh.

Approximation algorithm on multi-way maxcut partitioning. Lecture Notes in Computer Science, 855:148–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chepoi:1994:LTA

[2703] V. Chepoi and F. Dragan. A lineartime algorithm for finding a central vertex of a chordal graph. Lecture Notes in Computer Science, 855:159–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Israeli:1994:TCU

[2704] A. Israeli and A. Shirazi. The time complexity of updating snapshot memories. Lecture Notes in Computer Science, 855:171–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parlati:1994:NES

[2705] G. Parlati and M. Yung. Nonexploratory self-stabilization for constantspace symmetry-breaking. Lecture Notes in Computer Science, 855:183– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lund:1994:LDD

[2706] C. Lund, N. Reingold, J. Westbrook, and D. Yan. On-line distributed data management. Lecture Notes in Computer Science, 855:202–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asano:1994:USD

[2707] T. Asano, N. Katoh, and T. Tokuyama. A unified scheme for detecting fundamental curves in binary edge images. Lecture Notes in Computer Science, 855:215–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burnikel:1994:HCV

[2708] C. Burnikel, K. Mehlhorn, and S. Schirra. How to compute the Voronoi diagram of line segments: Theoretical and practical results. Lecture Notes in Computer Science, 855:227–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Overmars:1994:RSP

[2709] M. H. Overmars and A. F. Van der Stappen. Range searching and point location among fat objects. Lecture Notes in Computer Science, 855:240– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boissonnat:1994:CTB

[2710] J.-D. Boissonnat, J. Czyzowicz, O. Devillers, and J.-M. Robert. Convex tours of bounded curvature. Lecture Notes in Computer Science, 855: 254–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chiang:1994:OSP

[2711] Y.-J. Chiang and R. Tamassia. Optimal shortest path and minimum-link path queries in the presence of obstacles. Lecture Notes in Computer Science, 855:266–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1994:FAC

[2712] P. Gupta, R. Janardan, and M. Smid.

Fast algorithms for collision and proximity problems involving moving geometric objects. Lecture Notes in Computer Science, 855:278–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schiermeyer:1994:RFO

[2713] I. Schiermeyer. Reverse-fit: A 2-Optimal algorithm for packing rectangles. Lecture Notes in Computer Science, 855:290–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1994:OAP

[2714] B. Chen, A. Van Vliet, and G. J. Woeginger. An optimal algorithm for preemptive on-line scheduling. Lecture Notes in Computer Science, 855:300– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Srivastav:1994:TAR

[2715] A. Srivastav and P. Stangier. Tight approximations for resource constrained scheduling problems. Lecture Notes in Computer Science, 855:307–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wedelin:1994:APA

[2716] D. Wedelin. An algorithm for 0-1 programming with application to airline crew scheduling. Lecture Notes in Computer Science, 855:319–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Das:1994:WEP

[2717] S. K. Das and P. Ferragina. An o(n) work EREW parallel algorithm for updating MST. Lecture Notes in Computer Science, 855:331–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franciosa:1994:SDF

[2718] P. G. Franciosa, G. Gambosi, and U. Nanni. On the structure of DFS-Forests on directed graphs and the dynamic maintenance of DFS on DAG’s. Lecture Notes in Computer Science, 855:343–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alon:1994:FCG

[2719] N. Alon, R. Yuster, and U. Zwick.

Finding and counting given length cycles. Lecture Notes in Computer Science, 855:354–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ben-Aroya:1994:GHP

[2720] I. Ben-Aroya and A. Schuster. Greedy hot-potato routing on the mesh. Lecture Notes in Computer Science, 855: 365–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sibeyn:1994:DMS

[2721] J. F. Sibeyn. Desnakification of mesh sorting algorithms. Lecture Notes in Computer Science, 855:377–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pietracaprina:1994:TBD

[2722] A. Pietracaprina and G. Pucci. Tight bounds on deterministic PRAM emulations with constant redundancy. Lecture Notes in Computer Science, 855: 391–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chlebus:1994:PCR

[2723] B. S. Chlebus, A. Gambin, and P. Indyk. PRAM computations resilient to memory faults. Lecture Notes in Computer Science, 855:401–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bilardi:1994:ALB

[2724] G. Bilardi and P. Bay. An area lower bound for a class of fat-trees. Lecture Notes in Computer Science, 855:413– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hunt:1994:UAA

[2725] H. B. Hunt, M. V. Marathe, V. Radhakrishnan, and S. S. Ravi. A unified approach to approximation schemes for NP- and PSPACE-hard problems for geometric graphs. Lecture Notes in Computer Science, 855:424–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greenlaw:1994:PCE

[2726] R. Greenlaw and J. Machta. The parallel complexity of Eden growth, solid-on-solid growth and ballistic deposition. Lecture Notes in Computer Science, 855:436–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schoenhage:1994:NAR

[2727] A. Schoenhage and E. Vetter. A new approach to resultant computations and other algorithms with exact division. Lecture Notes in Computer Science, 855:448–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plandowski:1994:TEM

[2728] W. Plandowski. Testing equivalence of morphisms on context-free languages. Lecture Notes in Computer Science, 855:460–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gasieniec:1994:WTO

[2729] L. Gasieniec and K. Park. Worktime optimal parallel prefix matching. Lecture Notes in Computer Science, 855:471–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breslauer:1994:ECS

[2730] D. Breslauer, L. Colussi, and L. Toniolo. On the exact complexity of the string prefix-matching problem. Lecture Notes in Computer Science, 855: 483–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferragina:1994:ITE

[2731] P. Ferragina. Incremental text editing:

A new data structure. Lecture Notes in Computer Science, 855:495–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsichritzis:1994:RCI

[2732] D. Tsichritzis. Research as a catalyst for industrial innovation. Lecture Notes in Computer Science, 856:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dechamboux:1994:POO

[2733] P. Dechamboux and C. Roncancio. Peplom0d: an object oriented database programming language extended with deductive capabilities. Lecture Notes in Computer Science, 856:2–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hohenstein:1994:OOD

[2734] U. Hohenstein, R. Lauffer, and P. Weikert. Object-oriented database systems: How much SQL do they understand? Lecture Notes in Computer Science, 856:15–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ketabchi:1994:OCA

[2735] M. A. Ketabchi, S. Dasananda, R. Mikkilineni, and X. Li. An objectcentered application development system. Lecture Notes in Computer Science, 856:27–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Novak:1994:FFS

[2736] M. Novak, G. Gardarin, and P. Valduriez. Flora: A functional-style language for object and relational algebra. Lecture Notes in Computer Science, 856:37–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1994:OTE

[2737] Y. Chen and T. Haerder. On the optimal top-down evaluation of recursive queries. Lecture Notes in Computer Science, 856:47–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kumar:1994:SSC

[2738] A. Kumar. A study of spatial clustering techniques. Lecture Notes in Computer Science, 856:57–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lorentzos:1994:EMD

[2739] N. Lorentzos and Y. Manolopoulos. Efficient management of 2-d interval relations. Lecture Notes in Computer Science, 856:72–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burger:1994:MSD

[2740] F. Burger, G. Quirchmayr, S. Reich, and A. M. Tjoa. Managing structured documents in distributed publishing environments. Lecture Notes in Computer Science, 856:83–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hawryszkiewycz:1994:GSM

[2741] I. Hawryszkiewycz. A generalized semantic model for CSCW systems. Lecture Notes in Computer Science, 856: 93–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoppe:1994:CTI

[2742] H. U. Hoppe and J. Zhao. C-TORI:

An interface for cooperative database retrieval. Lecture Notes in Computer Science, 856:103–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rupietta:1994:OMC

[2743] W. Rupietta. Organization models for cooperative office applications. Lecture Notes in Computer Science, 856:114– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castellanos:1994:SER

[2744] M. Castellanos, F. Saltor, and M. Garcia-Solaco. Semantically enriching relational databases into an object oriented semantic model. Lecture Notes in Computer Science, 856:125– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1994:CCE

[2745] X. Li and Z. Tari. Consistency checking of evolving methods. Lecture Notes in Computer Science, 856:135–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakamoto:1994:DIH

[2746] N. Sakamoto and K. Ushijima. Designing and integrating human genome databases with object-oriented technology. Lecture Notes in Computer Science, 856:145–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spanoudakis:1994:EFS

[2747] G. Spanoudakis and P. Constantopoulos. On evidential feature salience. Lecture Notes in Computer Science, 856: 153–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ng:1994:RLT

[2748] V. Ng and T. Kameda. The R-link tree:

A recoverable index structure for spatial data. Lecture Notes in Computer Science, 856:163–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papadias:1994:RDR

[2749] D. Papadias, Y. Theodoridis, and T. Sellis. The retrieval of direction relations using R-trees. Lecture Notes in Computer Science, 856:173–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsatalos:1994:UFI

[2750] O. G. Tsatalos and Y. E. Ioannidis.

A unified framework for indexing in database systems. Lecture Notes in Computer Science, 856:183–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Macedo:1994:ISI

[2751] J. Macedo and R. R. Usano. An intelligent system to identify the appropriate organization for a factory. Lecture Notes in Computer Science, 856: 193–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tatsiopoulos:1994:OOD

[2752] I. P. Tatsiopoulos. Object-oriented design of manufacturing database helps simplify production management in small firms. Lecture Notes in Computer Science, 856:204–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xerokostas:1994:SPS

[2753] D. A. Xerokostas and K. G. Aravossis. A solution procedure for solving the non-rectangular cutting stock problem of the clothing industry using expert knowledge within an intelligent CAD environment. Lecture Notes in Computer Science, 856:214–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Edelweiss:1994:OOA

[2754] N. Edelweiss, J. P. M. De Oliveira, and B. Pernici. An object-oriented approach to a temporal query language. Lecture Notes in Computer Science, 856:225–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kokkotos:1994:FDT

[2755] S. Kokkotos and C. D. Spyropoulos. A framework for developing temporal databases. Lecture Notes in Computer Science, 856:236–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1994:ESS

[2756] C.-K. Li and S. Wang. Efficient storage structures for temporal object-oriented databases. Lecture Notes in Computer Science, 856:246–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soukeras:1994:TTF

[2757] S. Soukeras and P. J. H. King. TFDL: A temporal functional language for the management of historical databases. Lecture Notes in Computer Science, 856:259–??, 1994. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Theodoulidis:1994:OTD

[2758] B. Theodoulidis, A. Ait-Braham, and G. Karvelis. The ORES temporal DBMS and the ERT-SQL query language. Lecture Notes in Computer Science, 856:270–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonfatti:1994:OOC

[2759] F. Bonfatti, P. D. Monari, and P. Paganelli. Object-oriented constraint analysis in complex applications. Lecture Notes in Computer Science, 856: 280–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krotopoulou:1994:CDA

[2760] A. Krotopoulou, P. Spirakis, D. Terpou, and A. Tsakalidis. A conceptual database approach for modelling 3D objects of irregular geometry. Lecture Notes in Computer Science, 856: 290–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teisseire:1994:ALE

[2761] M. Teisseire and R. Cicchetti. An algebraic language for event-driven modelling. Lecture Notes in Computer Science, 856:300–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toval:1994:POO

[2762] A. Toval, I. Ramos, and O. Pastor. Prototyping object-oriented specifications in an algebraic environment. Lecture Notes in Computer Science, 856:310– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeKorvin:1994:IPU

[2763] A. De Korvin, G. Quirchmayr, and S. Hashemi. Identifying precedents under uncertainty. Lecture Notes in Computer Science, 856:320–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merkl:1994:CCA

[2764] D. Merkl, E. Schweighofer, and W. Winiwarter. CONCAT — connotation analysis of thesauri based on the interpretation of context meaning. Lecture Notes in Computer Science, 856: 329–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spinellis:1994:DIL

[2765] D. Spinellis. The design and implementation of a legal text database. Lecture Notes in Computer Science, 856:339– 348, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanderWees:1994:MJL

[2766] J. G. L. Van der Wees. Maintenance of juricas legal computer advice systems. Lecture Notes in Computer Science, 856:349–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mylopoulos:1994:ESI

[2767] J. Mylopoulos. Expert systems for industrial process control: A study in technology transfer. Lecture Notes in Computer Science, 856:359–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Defude:1994:PAD

[2768] B. Defude and H. Martin. From a passive to an active database supporting exceptions. Lecture Notes in Computer Science, 856:360–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kawano:1994:HDK

[2769] H. Kawano, S. Nishio, J. Han, and T. Hasegawa. How does knowledge discovery cooperate with active database techniques in controlling dynamic environment? Lecture Notes in Computer Science, 856:370–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naqui:1994:EAK

[2770] W. Naqui and M. T. Ibrahim. EECA:

An active knowledge model. Lecture Notes in Computer Science, 856:380– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palopoli:1994:APR

[2771] L. Palopoli and R. Torlone. Active and passive rules for complex object databases. Lecture Notes in Computer Science, 856:390–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gancarski:1994:MEV

[2772] S. Gancarski and G. Jomier. Managing entity versions within their contexts: A formal approach. Lecture Notes in Computer Science, 856:400–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hahn:1994:TEC

[2773] U. Hahn and M. Klenner. Tracking the evolution of concepts in dynamic worlds. Lecture Notes in Computer Science, 856:410–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rolland:1994:TEA

[2774] C. Rolland and N. Prakash. Tracing the evolution of artifacts. Lecture Notes in Computer Science, 856:420–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

TerHofstede:1994:SID

[2775] A. H. M. Ter Hofstede, H. A. Proper, and T. P. Van der Weide. Supporting information disclosure in an evolving environment. Lecture Notes in Computer Science, 856:433–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdelmoty:1994:GDH

[2776] A. I. Abdelmoty, N. W. Paton, M. H. Williams, and A. A. A. Fernandes. Geographic data handling in a deductive object-oriented database. Lecture Notes in Computer Science, 856:445– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kettal:1994:ESK

[2777] E.-Y. Kettal. Expert spatial knowledge: An expert geographical information system in diffusion of water. Lecture Notes in Computer Science, 856: 455–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Medeiros:1994:UVG

[2778] C. B. Medeiros and G. Jomier. Using versions in GIS. Lecture Notes in Computer Science, 856:465–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Metakides:1994:SRI

[2779] G. Metakides. The strategic role of information technology towards the information society. Lecture Notes in Computer Science, 856:475–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gardarin:1994:FOO

[2780] G. Gardarin. Federating objectoriented and relational databases. Lecture Notes in Computer Science, 856: 476–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaplan:1994:BGB

[2781] R. M. Kaplan. Bridging the gap between knowledge engineering and efficient implementation in an intelligent tutoring system. Lecture Notes in Computer Science, 856:477–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Landes:1994:MSK

[2782] D. Landes and R. Studer. Mechanisms for structuring knowledge-based systems. Lecture Notes in Computer Science, 856:488–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nikolaidu:1994:DSI

[2783] M. Nikolaidu, D. Lelis, D. Mouzakis, and P. Georgiadis. Distributed system intelligent design. Lecture Notes in Computer Science, 856:498–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaviotis:1994:MHB

[2784] I. Gaviotis and D. Christodoulakis. A model for hypermedia based production systems. Lecture Notes in Computer Science, 856:509–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hara:1994:HDS

[2785] Y. Hara and R. A. Botafogo. Hypermedia databases: A specification and formal language. Lecture Notes in Computer Science, 856:521–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hatzimanikatis:1994:IOH

[2786] A. Hatzimanikatis, A. Papadopoulos, and D. Christodoulakis. Implementing open hypermedia services on top of the Postgres extended relational database management system. Lecture Notes in Computer Science, 856:531–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsalgatidov:1994:SSB

[2787] A. Tsalgatidov, Z. Palaskas, C. Halatsis, and M. Hatzopoulos. A support system for building hypermedia training applications. Lecture Notes in Computer Science, 856:540–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bayer:1994:OMD

[2788] R. Bayer, P. Vogel, and S. Wiesener.

OMNIS/Myriad document retrieval and its database requirements. Lecture Notes in Computer Science, 856: 550–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Layaida:1994:CIR

[2789] R. Layaida, M. Boughanem, and A. Caron. Constructing an information retrieval system with neural networks. Lecture Notes in Computer Science, 856:561–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yoo:1994:HFN

[2790] J. S. Yoo, Y.-J. Lee, J. W. Chang, and M. H. Kim. The HS file: A new dynamic signature file method for efficient information retrieval. Lecture Notes in Computer Science, 856:571–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Billet-Coat:1994:MAA

[2791] S. Billet-Coat and D. Herin-Aime. A multi-agent architecture for an evolving expert system module. Lecture Notes in Computer Science, 856:581–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1994:IIC

[2792] S. Y. Lee and T. W. Ling. Improving integrity constraint checking for stratified deductive databases. Lecture Notes in Computer Science, 856:591– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prichard:1994:RRT

[2793] J. J. Prichard, L. C. DiPippo, J. Peckham, and V. F. Wolfe. RTSORAC: A real-time object-oriented database model. Lecture Notes in Computer Science, 856:601–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ulusoy:1994:ADT

[2794] O. Ulusoy. Applying database technology to real-time systems. Lecture Notes in Computer Science, 856:611– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aggarwal:1994:DED

[2795] S. Aggarwal, I. Kim, and W. Meng. Database exploration with dynamic abstractions. Lecture Notes in Computer Science, 856:621–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bosc:1994:SQN

[2796] P. Bosc, L. Lietard, and O. Pivert. Soft querying, a new feature for database management systems. Lecture Notes in Computer Science, 856:631–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gunter:1994:SFD

[2797] E. Gunter and L. Libkin. OR-SML: A functional database programming language for disjunctive information and its applications. Lecture Notes in Computer Science, 856:641–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1994:SQO

[2798] Y.-W. Lee and S. I. Yoo. Semantic query optimization in OODB systems. Lecture Notes in Computer Science, 856:651–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goble:1994:MTS

[2799] C. Goble, P. Crowther, and D. Solomon. A medical terminology server. Lecture Notes in Computer Science, 856:661–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Houstis:1994:BBI

[2800] C. Houstis, T. S. Papatheodorou, V. Verykios, and A. Floratos. BIND: a biomedical interoperable database system. Lecture Notes in Computer Science, 856:671–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pangalos:1994:DSM

[2801] G. Pangalos, A. Pomporisis, L. Bozios, and M. Khair. Development of secure medical database systems. Lecture Notes in Computer Science, 856: 680–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanDyne:1994:UIM

[2802] M. Van Dyne, C. Tsatsoulis, and J. Thorp. Using inductive machine learning, expert systems and case based reasoning to predict preterm delivery in pregnant women. Lecture Notes in Computer Science, 856:690–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drescher:1994:IFR

[2803] P. Drescher, M. Holetha, R. Kruschinski, and G. Laufkoetter. Integrating frames, rules and uncertainty in a database-coupled knowledgerepresentation system. Lecture Notes in Computer Science, 856:703–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamel:1994:PRB

[2804] M. N. Kamel. A prototype rule-based front end system for integrity enforcement in relational databases: An application to the Naval Aircraft Flight Records database. Lecture Notes in Computer Science, 856:713–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lakshmanan:1994:MUD

[2805] V. S. Lakshmanan and F. Sadri. Modeling uncertainty in deductive databases. Lecture Notes in Computer Science, 856:724–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papakonstantinou:1994:DDB

[2806] G. Papakonstantinou, C. Voliotis, and N. M. Sgouros. Dependency-directed binding of variables for constraint logic programming. Lecture Notes in Computer Science, 856:734–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Afsarmanesh:1994:FDI

[2807] H. Afsarmanesh, M. Wiedijk, and L. O. Hertzberger. Flexible and dynamic integration of multiple information bases. Lecture Notes in Computer Science, 856:744–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boeszoermenyi:1994:PPD

[2808] L. Boeszoermenyi, J. Eder, and C. Weich. PPOST: A parallel database in main memory. Lecture Notes in Computer Science, 856:754–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hameurlain:1994:EIO

[2809] A. Hameurlain and F. Morvan. Exploiting inter-operation parallelism for SQL query optimization. Lecture Notes in Computer Science, 856:759– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berndtsson:1994:ROO

[2810] M. Berndtsson. Reactive objectoriented databases and CIM. Lecture Notes in Computer Science, 856:769– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eberl:1994:ACI

[2811] M. Eberl. Advanced computer integrated technology for furniture manufacturing. Lecture Notes in Computer Science, 856:779–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marik:1994:DSC

[2812] V. Marik, J. Lazansky, J. Koutnik, and T. Hazdra. A distributed system for CIM. Lecture Notes in Computer Science, 856:785–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Starzacher:1994:ICS

[2813] P. Starzacher and W. Woess. Integratedcompany-software (INCOSOFT) — a special solution for small- and mediumsized companies. Lecture Notes in Computer Science, 856:795–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Babaoglu:1994:RFM

[2814] O. Babaoglu, A. Bartoli, and G. Dini. Replicated file management in largescale distributed systems. Lecture Notes in Computer Science, 857:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keller:1994:RPP

[2815] J. Keller, W. J. Paul, and D. Scheerer. Realization of PRAMs: Processor design. Lecture Notes in Computer Science, 857:17–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lenstra:1994:F

[2816] A. K. Lenstra. Factoring. Lecture Notes in Computer Science, 857:28–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schiper:1994:PPV

[2817] A. Schiper and A. Sandoz. Primary partition “virtually-synchronous communication” harder than consensus. Lecture Notes in Computer Science, 857:39–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yung:1994:SDC

[2818] M. Yung. Secure distributed computing: Theory and practice. Lecture Notes in Computer Science, 857:53–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaudhuri:1994:FIR

[2819] S. Chaudhuri and M. R. Tuttle. Fast increment registers. Lecture Notes in Computer Science, 857:74–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choy:1994:CC

[2820] M. Choy and A. K. Singh. Coping with contention. Lecture Notes in Computer Science, 857:89–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cori:1994:EBC

[2821] R. Cori and S. Moran. Exotic behaviour of consensus numbers. Lecture Notes in Computer Science, 857:101– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haldar:1994:SOC

[2822] S. Haldar and P. S. Subramanian.

Space-optimum conflict-free construction of 1-Writer 1-Reader multivalued atomic variable. Lecture Notes in Computer Science, 857:116–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inoue:1994:LTS

[2823] M. Inoue, W. Chen, T. Masuzawa, and N. Tokura. Linear-time snapshot using multi-writer multi-reader registers. Lecture Notes in Computer Science, 857:130–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moir:1994:FLL

[2824] M. Moir and J. H. Anderson. Fast, long-lived renaming. Lecture Notes in Computer Science, 857:141–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rachman:1994:AWF

[2825] O. Rachman. Anomalies in the waitfree hierarchy. Lecture Notes in Computer Science, 857:156–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alimonti:1994:FED

[2826] P. Alimonti, P. Flocchini, and N. Santoro. Finding the extreme of a distributed multiset. Lecture Notes in Computer Science, 857:164–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Attiya:1994:LHR

[2827] H. Attiya and R. Rappoport. The level of handshake required for establishing a connection. Lecture Notes in Computer Science, 857:179–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Attiya:1994:LLR

[2828] H. Attiya, H. Shachnai, and T. Tamir. Local labeling and resource allocation using preprocessing. Lecture Notes in Computer Science, 857:194–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cidon:1994:SAR

[2829] I. Cidon, O. Gerstel, and S. Zaks. A scalable approach to routing in ATM networks. Lecture Notes in Computer Science, 857:209–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shabtay:1994:LCN

[2830] L. Shabtay and A. Segall. Low complexity network synchronization. Lecture Notes in Computer Science, 857: 223–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Swaminathan:1994:IDA

[2831] B. Swaminathan and K. J. Goldman. An incremental distributed algorithm for computing biconnected components. Lecture Notes in Computer Science, 857:238–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garay:1994:RMA

[2832] J. A. Garay. Reaching (and maintaining) agreement in the presence of mobile faults. Lecture Notes in Computer Science, 857:253–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoepman:1994:UDS

[2833] J.-H. Hoepman. Uniform deterministic self-stabilizing ring-orientation on odd-length rings. Lecture Notes in Computer Science, 857:265–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lo:1994:UFD

[2834] W.-K. Lo and V. Hadzilacos. Using failure detectors to solve consensus in asynchronous shared-memory systems. Lecture Notes in Computer Science, 857:280–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsay:1994:AOF

[2835] Y.-K. Tsay and R. L. Bagrodia. An algorithm with optimal failure locality for the dining philosophers problem. Lecture Notes in Computer Science, 857:296–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allenberg-Navony:1994:ARC

[2836] N. Allenberg-Navony, A. Itai, and S. Moran. Average and randomized complexity of distributed problems. Lecture Notes in Computer Science, 857:311–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Awerbuch:1994:SSL

[2837] B. Awerbuch, B. Patt-Shamir, G. Varghese, and S. Dolev. Self-stabilization by local checking and global reset. Lecture Notes in Computer Science, 857: 326–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basten:1994:BTD

[2838] T. Basten. Breakpoints and time in distributed computations. Lecture Notes in Computer Science, 857:340– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruget:1994:CMC

[2839] F. Ruget. Cheaper matrix clocks. Lecture Notes in Computer Science, 857: 355–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Exertier:1994:IER

[2840] F. Exertier and S. Haj Houssain. Issues in extending a relational system with object-oriented features. Lecture Notes in Computer Science, 858:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herzig:1994:SLQ

[2841] R. Herzig and M. Gogolla. An SQL-Like query calculus for object-oriented database systems. Lecture Notes in Computer Science, 858:20–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1994:SEO

[2842] J.-L. Chen, D. McLeod, and D. O’Leary.

Schema evolution for object-based accounting database systems. Lecture Notes in Computer Science, 858:40–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carzaniga:1994:DII

[2843] A. Carzaniga, G. P. Picco, and G. Vigna. Designing and implementing inter-client communication in the O2 object-oriented database management system. Lecture Notes in Computer Science, 858:53–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreira:1994:ROO

[2844] A. M. D. Moreira and R. G. Clark. Rigorous object-oriented analysis. Lecture Notes in Computer Science, 858:65–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rouxel:1994:OOM

[2845] C. Rouxel, J.-P. Velu, M. Texier, and D. Leroey. Object-oriented methodologies for large scale projects: Does it work? Lecture Notes in Computer Science, 858:79–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pomberger:1994:QQA

[2846] G. Pomberger and W. Pree. Quantitative and qualitative aspects of objectoriented software development. Lecture Notes in Computer Science, 858:96–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonfatti:1994:TGP

[2847] F. Bonfatti and P. D. Monari. Towards a general purpose approach to objectoriented analysis. Lecture Notes in Computer Science, 858:108–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liddle:1994:SMO

[2848] S. W. Liddle, D. W. Embley, and S. N. Woodfield. A seamless model for object-oriented systems development. Lecture Notes in Computer Science, 858:123–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ebert:1994:SBV

[2849] J. Ebert and G. Engels. Structural and behavioral views on OMT-Classes. Lecture Notes in Computer Science, 858:142–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mathiassen:1994:CTA

[2850] L. Mathiassen, A. Munk-Madsen, P. A.

Nielsen, and J. Stage. Combining two approaches to object-oriented analysis. Lecture Notes in Computer Science, 858:158–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montesi:1994:RTI

[2851] D. Montesi and R. Torlone. A rewriting technique for implementing active object systems. Lecture Notes in Computer Science, 858:171–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kappel:1994:BCM

[2852] G. Kappel, S. Rausch-Schott, and W. Retschitzegger. Beyond coupling modes: Implementing active concepts on top of a commercial OODBMS. Lecture Notes in Computer Science, 858: 189–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

HuffmanHayes:1994:TOO

[2853] J. Huffman Hayes. Testing of object-oriented programming systems (OOPS): A fault-based approach. Lecture Notes in Computer Science, 858: 205–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shrivastava:1994:ARA

[2854] C. Shrivastava, D. L. Carver, and R. Shrivastava. An ambiguity resolution algorithm. Lecture Notes in Computer Science, 858:221–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kochhar:1994:PHL

[2855] S. Kochhar. On providing a high-level C interface for an object-oriented C++ system. Lecture Notes in Computer Science, 858:232–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paltrinieri:1994:IOC

[2856] M. Paltrinieri. Integrating objects with constraint-programming languages. Lecture Notes in Computer Science, 858:248–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boccignone:1994:OOR

[2857] G. Boccignone, A. Chianese, M. De Santo, and A. Picariello. Objectoriented representation of shape information. Lecture Notes in Computer Science, 858:266–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papa:1994:DOO

[2858] M. P. Papa, G. Ragucci, G. Corrente, and M. Ferrise. The development of an object-oriented multimedia information system. Lecture Notes in Computer Science, 858:278–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damiani:1994:IIS

[2859] M. L. Damiani. An intelligent information system for heterogeneous data exploration. Lecture Notes in Computer Science, 858:298–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graham:1994:EOC

[2860] P. Graham and K. Barker. Effective optimistic concurrency control in multiversion object bases. Lecture Notes in Computer Science, 858:313–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parisi-Presicce:1994:ROO

[2861] F. Parisi-Presicce and A. Pierantonio.

Reusing object-oriented design: An algebraic approach. Lecture Notes in Computer Science, 858:329–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castano:1994:ROOb

[2862] S. Castano and V. De Antonellis.

Reuse in object-oriented information systems development. Lecture Notes in Computer Science, 858:346–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corbellini:1994:OOA

[2863] P. Corbellini, P. Della Vigna, F. Mercalli, and M. Pugliese. An objectoriented approach to the integration of online services into office automation environments. Lecture Notes in Computer Science, 858:359–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mich:1994:LAD

[2864] L. Mich and R. Garigliano. A linguistic approach to the development of object-oriented systems using the NL system LOLITA. Lecture Notes in Computer Science, 858:371–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agerholm:1994:LEH

[2865] S. Agerholm. LCF examples in HOL. Lecture Notes in Computer Science, 859:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersen:1994:GTP

[2866] F. Andersen, K. D. Petersen, and J. S. Pettersson. A graphical tool for proving UNITY progress. Lecture Notes in Computer Science, 859:17–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Angelo:1994:RAC

[2867] C. M. Angelo, L. Claesen, and H. De Man. Reasoning about a class of linear systems of equations in HOL. Lecture Notes in Computer Science, 859:33–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodeveix:1994:THT

[2868] J.-P. Bodeveix, M. Filali, and P. Roche. Towards a HOL theory of memory. Lecture Notes in Computer Science, 859: 49–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brackin:1994:PTS

[2869] S. H. Brackin. Providing tractable security analyses in HOL. Lecture Notes in Computer Science, 859:65–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBruijn:1994:HLF

[2870] N. G. De Bruijn. Highlighting the lambda-free fragment of automath. Lecture Notes in Computer Science, 859:81–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Busch:1994:FOA

[2871] H. Busch. First-order automation for higher-order-logic theorem proving. Lecture Notes in Computer Science, 859:97–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Camilleri:1994:SAP

[2872] J. Camilleri and V. Zammit. Symbolic animations as a proof tool. Lecture Notes in Computer Science, 859:113– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chapman:1994:DL

[2873] N. Chapman, S. Finn, and M. P. Fourman. Datatypes in L2. Lecture Notes in Computer Science, 859:128–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chou:1994:FTU

[2874] C.-T. Chou. A formal theory of undirected graphs in higher-order logic. Lecture Notes in Computer Science, 859:144–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chou:1994:MVD

[2875] C.-T. Chou. Mechanical verification of distributed algorithms in higherorder logic. Lecture Notes in Computer Science, 859:158–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Curzon:1994:TDC

[2876] P. Curzon. Tracking design changes with formal verification. Lecture Notes in Computer Science, 859:177– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Forster:1994:WSS

[2877] T. Forster. Weak systems of set theory related to HOL. Lecture Notes in Computer Science, 859:193–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fura:1994:ISC

[2878] D. A. Fura and A. K. Somani. Intervalsemantic component models and the efficient verification of transaction-level circuit behavior. Lecture Notes in Computer Science, 859:205–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graham:1994:INH

[2879] B. T. Graham. An interpretation of NODEN in HOL. Lecture Notes in Computer Science, 859:221–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanna:1994:RAR

[2880] K. Hanna. Reasoning about real circuits. Lecture Notes in Computer Science, 859:235–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harrison:1994:BDD

[2881] J. Harrison. Binary decision diagrams as a HOL derived rule. Lecture Notes in Computer Science, 859:254–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Homeier:1994:TTT

[2882] P. V. Homeier and D. F. Martin. Trustworthy tools for trustworthy programs: A verified verification condition generator. Lecture Notes in Computer Science, 859:269–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joyce:1994:MRS

[2883] J. Joyce, N. Day, and M. Donat. S:

A machine readable specification notation based on higher order logic. Lecture Notes in Computer Science, 859: 285–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsson:1994:EAF

[2884] M. Larsson. An engineering approach to formal digital system design. Lecture Notes in Computer Science, 859: 300–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1994:GDU

[2885] J.-Y. Lu and S.-K. Chin. Generating designs using an algorithmic register transfer language with formal semantics. Lecture Notes in Computer Science, 859:316–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laangbacka:1994:HFT

[2886] T. Laangbacka. A HOL formalisation of the temporal logic of actions. Lecture Notes in Computer Science, 859:332–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maharaj:1994:SMM

[2887] S. Maharaj and E. Gunter. Studying the ML module system in HOL. Lecture Notes in Computer Science, 859: 346–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prasetya:1994:TMS

[2888] I. S. W. B. Prasetya. Towards a mechanically supported and compositional calculus to design distributed algorithms. Lecture Notes in Computer Science, 859:362–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reetz:1994:SDE

[2889] R. Reetz and T. Kropf. Simplifying deep embedding: A formalised code generator. Lecture Notes in Computer Science, 859:378–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneider:1994:AVF

[2890] K. Schneider, R. Kumar, and T. Kropf.

Automating verification by functional abstraction at the system level. Lecture Notes in Computer Science, 859: 391–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Slind:1994:PPM

[2891] K. Slind. A parameterized proof manager. Lecture Notes in Computer Science, 859:407–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tahar:1994:IIV

[2892] S. Tahar and R. Kumar. Implementational issues for verifying RISC-Pipeline conflicts in HOL. Lecture Notes in Computer Science, 859:424– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Windley:1994:SIS

[2893] P. J. Windley. Specifying instructionset architectures in HOL: A primer. Lecture Notes in Computer Science, 859:440–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VonWright:1994:RHO

[2894] J. Von Wright. Representing higher order logic proofs in HOL. Lecture Notes in Computer Science, 859:456– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laughton:1994:CIT

[2895] M. F. Laughton. Communications and information technology for persons with disabilities — the Canadian National strategy as an example. Lecture Notes in Computer Science, 860:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kieninger:1994:IIS

[2896] T. Kieninger, N. Kuhn, K. Seidenschwann, and W. Weiss. An intelligent information system for blind people — AI technology and philosophical aspects. Lecture Notes in Computer Science, 860:9–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karshmer:1994:AGU

[2897] A. J. Karshmer, B. Ogden, P. Brawner, and K. Kaugars. Adapting graphical user interfaces for use by visually handicapped computer users: Current results and continuing research. Lecture Notes in Computer Science, 860:16–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber:1994:TBP

[2898] G. Weber, H. Petrie, D. Kochanek, and S. Morley. Training blind people in the use of graphical user interfaces. Lecture Notes in Computer Science, 860:25–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bothe:1994:AVS

[2899] H. H. Bothe and E. A. Wieden. Artificial visual speech, synchronized with a speech synthesis system. Lecture Notes in Computer Science, 860:32–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1994:NRS

[2900] J. Lee. Notational representation of sign language: A structural description of hand configuration. Lecture Notes in Computer Science, 860:38–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oakey:1994:FAR

[2901] J. E. Oakey. Further advances in realtime voice to text with steno interpreters. Lecture Notes in Computer Science, 860:46–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buehler:1994:UCI

[2902] C. Buehler, H. Heck, and R. Wallbruch. A uniform control interface for various electronic aids. Lecture Notes in Computer Science, 860:51–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tronconi:1994:ISC

[2903] A. Tronconi and M. Billi. An integrated system for communication and equipments control using radio link. Lecture Notes in Computer Science, 860:59–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flachberger:1994:AFE

[2904] C. Flachberger, P. Panek, and W. L.

Zagler. AUTONOMY — A flexible and easy-to-use assistive system to support the independence of handicapped and elderly persons. Lecture Notes in Computer Science, 860:65–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thatcher:1994:SRP

[2905] J. W. Thatcher. Screen reader/2 — programmed access to the GUI. Lecture Notes in Computer Science, 860:76–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kochanek:1994:DOM

[2906] D. Kochanek. Designing an OffScreen model for a GUI. Lecture Notes in Computer Science, 860:89–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graziani:1994:SRW

[2907] P. Graziani and B. Breschi. Screen reader for Windows based on speech output. Lecture Notes in Computer Science, 860:96–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmitt:1994:NWL

[2908] W. Schmitt and W. Zang. The new wireless LinguControl. Lecture Notes in Computer Science, 860:101–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kerstner:1994:FMC

[2909] W. Kerstner, G. Pigel, and M. Tscheligi. The FeelMouse: Making computer screens feelable. Lecture Notes in Computer Science, 860:106–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bowes:1994:UBV

[2910] D. R. Bowes. Unexpected benefits of voice type computing. Lecture Notes in Computer Science, 860:115–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eftring:1994:RCM

[2911] H. Eftring. Robot control methods using the RAID workstation. Lecture Notes in Computer Science, 860:120– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neverya:1994:UNR

[2912] H. Neverya and G. Bolmsjoe. The ultrasonic navigating robot, WALKY. Lecture Notes in Computer Science, 860:129–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levine:1994:NES

[2913] S. P. Levine, D. A. Bell, and Y. Koren.

NavChair: An example of a sharedcontrol system for assistive technologies. Lecture Notes in Computer Science, 860:136–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crispien:1994:USA

[2914] K. Crispien, W. Wuerz, and G. Weber. Using spatial audio for the enhanced presentation of synthesized speech within screen-readers for blind computer users. Lecture Notes in Computer Science, 860:144–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vigouroux:1994:MCN

[2915] N. Vigouroux and B. Oriola. Multimodal concept for a new generation of screen reader. Lecture Notes in Computer Science, 860:154–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kennel:1994:AEU

[2916] A. R. Kennel. Auditory extension of user interfaces. Lecture Notes in Computer Science, 860:162–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cooper:1994:ADF

[2917] R. J. Cooper and J. C. Senge. An attempt to define fully-accessible workstation levels of accessibility. Lecture Notes in Computer Science, 860:164– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalcher:1994:GBC

[2918] J. Kalcher, D. Flotzinger, S. Goelly, and C. Neuper. Graz brain-computer interface (BCI) II. Lecture Notes in Computer Science, 860:170–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cherry:1994:HCI

[2919] A. D. Cherry, M. S. Hawley, M. Freeman, and P. A. Cudd. Humancomputer interfacing for the severely physically disabled. Lecture Notes in Computer Science, 860:177–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Busby:1994:DAT

[2920] G. Busby. Day and assessment training technology centres. Lecture Notes in Computer Science, 860:185–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leneway:1994:NAA

[2921] R. J. Leneway. The North America Association of Rehabilitation Programs in Computer Technology seeks to network with other training programs. Lecture Notes in Computer Science, 860:189–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1994:DRD

[2922] Q. Li and M. Guo. Disability and rehabilitation database in Chinese language. Lecture Notes in Computer Science, 860:192–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lane:1994:TSG

[2923] J. P. Lane. Toward a single global market for assistive technology. Lecture Notes in Computer Science, 860: 195–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ayres:1994:USW

[2924] N. Ayres and T. Wesley. Using structure within electronic documents to make editors more accessible. Lecture Notes in Computer Science, 860:198– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lokowandt:1994:DPT

[2925] G. Lokowandt and W. Schweikhardt.

Distinguishing pattern-types in printed documents. Lecture Notes in Computer Science, 860:206–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bauwens:1994:SDK

[2926] B. Bauwens, J. Engelen, F. Evenepoel, and C. Tobin. Structuring documents: The key to increasing access to information for the print disabled. Lecture Notes in Computer Science, 860:214–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klaus:1994:SCV

[2927] J. Klaus. Study center for visually impaired persons — supportive system for blind and partially sighted students at the University of Karlsruhe/ Germany. Lecture Notes in Computer Science, 860:222–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moravcikova:1994:SCV

[2928] L. Moravcikova. Support center for visually impaired students. Lecture Notes in Computer Science, 860:229– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1994:EEC

[2929] R. Wagner, B. Stoeger, and K. Miesenberger. Educational endeavour “computer science for the blind” — state of the art and experiences in supporting visually handicapped students. Lecture Notes in Computer Science, 860:236– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stoeger:1994:MIB

[2930] B. Stoeger and K. Miesenberger. Modellversuch “Informatik für Blinde”.

Lecture Notes in Computer Science, 860:244–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barnes:1994:CIN

[2931] K. M. S. Barnes, R. W. Jotham, and C. I. G. Sherman. Cottage industry at NewLink. Lecture Notes in Computer Science, 860:246–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zampiceni:1994:THP

[2932] M. Zampiceni. Telework for handicapped people: An experience. Lecture Notes in Computer Science, 860: 249–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramseur:1994:ARA

[2933] J. H. Ramseur, L. S. Gardner, and D. S. Painter. AbleProfessionals: A recruiting and accommodation service for Atlanta employers. Lecture Notes in Computer Science, 860:252–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Resele:1994:NCD

[2934] P. Resele and W. Resele. “News-Reader” — A comfortable-digital newspaper and bookreading system. Lecture Notes in Computer Science, 860: 255–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hansson:1994:DTB

[2935] K. Hansson, L. Soennebo, and J. Lindholm. Digital talking books — A report from a practical, ongoing project. Lecture Notes in Computer Science, 860: 262–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneider:1994:EKA

[2936] A. Schneider and C. Pescia. The electronic kiosk — accessing newspapers with electronic media. Lecture Notes in Computer Science, 860:269–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nagel:1994:AAT

[2937] H.-H. Nagel. Al approaches towards sensor-based driver support in road vehicles. Lecture Notes in Computer Science, 861:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bornscheuer:1994:RCA

[2938] S.-E. Bornscheuer and M. Thielscher. Representing concurrent actions and solving conflicts. Lecture Notes in Computer Science, 861:16–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goos:1994:PSC

[2939] K. Goos. Preselection strategies for case based classification. Lecture Notes in Computer Science, 861:28–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stopp:1994:USR

[2940] E. Stopp, K.-P. Gapp, G. Herzog, and T. Laengle. Utilizing spatial relations for natural language access to an autonomous mobile robot. Lecture Notes in Computer Science, 861:39–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baader:1994:CRC

[2941] F. Baader, M. Buchheit, and B. Hollunder. Cardinality restrictions on concepts. Lecture Notes in Computer Science, 861:51–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pomplun:1994:ANN

[2942] M. Pomplun, B. Velichkovsky, and H. Ritter. An artificial neural network for high precision eye movement tracking. Lecture Notes in Computer Science, 861:63–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuznetsov:1994:KKL

[2943] V. Kuznetsov. A Kripke-Kleene logic over general logic programs. Lecture Notes in Computer Science, 861:70–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dix:1994:SSV

[2944] J. Dix and M. Mueller. The stable semantics and its variants: A comparison of recent approaches. Lecture Notes in Computer Science, 861:82–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gloeckner:1994:TCS

[2945] C. Gloeckner, A. Grieszl, M. Mueller, and M. Ronthaler. TabVer: A case study in table verbalization. Lecture Notes in Computer Science, 861:94–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jameson:1994:CND

[2946] A. Jameson, B. Kipper, A. Ndiaye, and R. Schaefer. Cooperating to be noncooperative: The dialog system PRACMA. Lecture Notes in Computer Science, 861:106–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfahringer:1994:RCI

[2947] B. Pfahringer. Robust constructive induction. Lecture Notes in Computer Science, 861:118–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuhr:1994:ESN

[2948] T. Fuhr, F. Kummert, and G. Sagerer. Enriching a semantic network language by integrating qualitative reasoning techniques. Lecture Notes in Computer Science, 861:130–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haarslev:1994:CST

[2949] V. Haarslev, R. Moeller, and C. Schroeder. Combining spatial and terminological reasoning. Lecture Notes in Computer Science, 861:142–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaaf:1994:DGC

[2950] J. W. Schaaf. Detecting gestalts in CAD-Plans to be used as indices for case-retrieval in architecture. Lecture Notes in Computer Science, 861:154– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schikuta:1994:NST

[2951] E. Schikuta. The NeuDB-System: Towards the integration of neural networks and database systems. Lecture Notes in Computer Science, 861:166– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quantz:1994:WDD

[2952] J. Quantz and S. Suska. Weighted defaults in description logics: Formal properties and proof theory. Lecture Notes in Computer Science, 861:178– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becker:1994:EQC

[2953] A. Becker and G. Lakemeyer. Epistemic queries in classic. Lecture Notes in Computer Science, 861:190– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanLinder:1994:CRA

[2954] B. Van Linder, W. Van der Hoek, and J.-J. C. Meyer. Communicating rational agents. Lecture Notes in Computer Science, 861:202–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer-Fujara:1994:KLM

[2955] J. Meyer-Fujara, B. Heller, S. Schlegelmilch, and I. Wachsmuth. Knowledgelevel modularization of a complex knowledge base. Lecture Notes in Computer Science, 861:214–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lukaszewicz:1994:PVT

[2956] W. Lukaszewicz and E. Madalinska-Bugaj. Program verification techniques as a tool for reasoning about action and change. Lecture Notes in Computer Science, 861:226–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DelCerro:1994:CLU

[2957] L. F. Del Cerro and A. Herzig. A conditional logic for updating in the possible models approach. Lecture Notes in Computer Science, 861:237–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schurz:1994:PJD

[2958] G. Schurz. Probabilistic justification of default reasoning. Lecture Notes in Computer Science, 861:248–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zaverucha:1994:PCD

[2959] G. Zaverucha. A prioritized contextual default logic: Curing anomalous extensions with a simple abnormality default theory. Lecture Notes in Computer Science, 861:260–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delgrande:1994:ISC

[2960] J. P. Delgrande and T. H. Schaub. Incorporating specificity into circumscriptive theories. Lecture Notes in Computer Science, 861:272–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rott:1994:CCE

[2961] H. Rott. Coherent choice and epistemic entrenchment. Lecture Notes in Computer Science, 861:284–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1994:NTL

[2962] Z. Lin and W. Li. A note on tableaux of logic of paradox. Lecture Notes in Computer Science, 861:296–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asher:1994:WNC

[2963] N. Asher and J. Lang. When nonmonotonicity comes from distances. Lecture Notes in Computer Science, 861:308– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becher:1994:RUC

[2964] G. Becher and U. Petermann. Rigid unification by completion and rigid paramodulation. Lecture Notes in Computer Science, 861:319–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kohlhase:1994:USL

[2965] M. Kohlhase. Unification in a sorted lambda-calculus with term declarations and function sorts. Lecture Notes in Computer Science, 861:331–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denzinger:1994:GOE

[2966] J. Denzinger and M. Fuchs. Goal oriented equational theorem proving using team work. Lecture Notes in Computer Science, 861:343–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gent:1994:HRS

[2967] I. P. Gent and T. Walsh. The hardest random SAT problems. Lecture Notes in Computer Science, 861:355– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Madden:1994:FMA

[2968] P. Madden. Formal methods for automated program improvement. Lecture Notes in Computer Science, 861: 367–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1994:AMN

[2969] X. Huang, M. Kerber, M. Kohlhase, and J. Richts. Adapting methods to novel tasks in proof planning. Lecture Notes in Computer Science, 861:379– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amtrup:1994:UCT

[2970] J. W. Amtrup. Using charts for transfer in MT. Lecture Notes in Computer Science, 861:391–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baaz:1994:NFC

[2971] M. Baaz and K. Hoerwein. A new frame for common-sense reasoning — towards local inconsistencies. Lecture Notes in Computer Science, 861:392–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cordier:1994:PTU

[2972] M.-O. Cordier and P. Siegel. Prioritized transitions for updates. Lecture Notes in Computer Science, 861:393– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eckerle:1994:OBS

[2973] J. Eckerle. An optimal bidirectional search algorithm. Lecture Notes in Computer Science, 861:394–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Filipic:1994:LDP

[2974] B. Filipic, I. Zun, and M. Perpar.

Learning to discriminate phases in gasliquid flow. Lecture Notes in Computer Science, 861:395–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giesl:1994:SSC

[2975] J. Giesl and I. Neumann. Strategies for semantical contractions. Lecture Notes in Computer Science, 861:396– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Penders:1994:CSI

[2976] J. S. J. H. Penders. Conflicts in the spatial interaction of autonomous agents. Lecture Notes in Computer Science, 861:397–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rapp:1994:URS

[2977] S. Rapp, M. Jessen, and G. Dogil. Using rough sets theory to predict German word stress. Lecture Notes in Computer Science, 861:398–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaedler:1994:GTA

[2978] K. Schaedler. Graph-theoretical algorithms and knowledge-based design. Lecture Notes in Computer Science, 861:399–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ternovskaia:1994:ISC

[2979] E. Ternovskaia. Interval situation calculus. Lecture Notes in Computer Science, 861:400–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Itai:1994:LMP

[2980] A. Itai. Learning morphology — practice makes good. Lecture Notes in Computer Science, 862:5–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takada:1994:HLF

[2981] Y. Takada. A hierarchy of language families learnable by regular language learners. Lecture Notes in Computer Science, 862:16–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dupont:1994:WSS

[2982] P. Dupont, L. Miclet, and E. Vidal. What is the search space of the regular inference? Lecture Notes in Computer Science, 862:25–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sempere:1994:CEL

[2983] J. M. Sempere and P. Garcia. A characterization of even linear languages and its application to the learning problem. Lecture Notes in Computer Science, 862:38–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciortuz:1994:OOI

[2984] L.-V. Ciortuz. Object-oriented inferences in a logical framework for feature grammars. Lecture Notes in Computer Science, 862:45–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jardino:1994:ADS

[2985] M. Jardino and G. Adda. Automatic determination of a stochastic bi-gram class language model. Lecture Notes in Computer Science, 862:57–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeMarcken:1994:ALP

[2986] C. De Marcken. The acquisition of a lexicon from paired phoneme sequences and semantic representations. Lecture Notes in Computer Science, 862:66–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DellaPietra:1994:IEL

[2987] S. Della Pietra, V. Della Pietra, J. Gillet, and J. Lafferty. Inference and estimation of a long-range trigram model. Lecture Notes in Computer Science, 862:78–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castellanos:1994:AOM

[2988] A. Castellanos, I. Galiano, and E. Vidal. Application of OSTIA to machine translation tasks. Lecture Notes in Computer Science, 862:93–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stolcke:1994:IPG

[2989] A. Stolcke and S. Omohundro. Inducing probabilistic grammars by Bayesian model merging. Lecture Notes in Computer Science, 862:106–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Casacuberta:1994:SES

[2990] F. Casacuberta. Statistical estimation of stochastic context-free grammars using the inside-outside algorithm and a transformation on grammars. Lecture Notes in Computer Science, 862:119–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanchez:1994:SIL

[2991] J. A. Sanchez and J. M. Benedi. Statistical inductive learning of regular formal languages. Lecture Notes in Computer Science, 862:130–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carrasco:1994:LSR

[2992] R. C. Carrasco and J. Oncina. Learning stochastic regular grammars by means of a state merging method. Lecture Notes in Computer Science, 862:139– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ahonen:1994:FGS

[2993] H. Ahonen, H. Mannila, and E. Nikunen. Forming grammars for structured documents: an application of grammatical inference. Lecture Notes in Computer Science, 862:153–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lucas:1994:CSS

[2994] S. Lucas, E. Vidal, A. Amiri, and S. Hanlon. A comparison of syntactic and statistical techniques for offline OCR. Lecture Notes in Computer Science, 862:168–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruiz:1994:ARK

[2995] J. Ruiz and P. Garcia. The algorithms RT and k-TTI: a first comparison. Lecture Notes in Computer Science, 862:180–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roques:1994:DGR

[2996] M. Roques. Dynamic grammatical representations in guided propagation networks. Lecture Notes in Computer Science, 862:189–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alquezar:1994:HCS

[2997] R. Alquezar and A. Sanfeliu. A hybrid connectionist-symbolic approach to regular grammatical inference based on neural learning and hierarchical clustering. Lecture Notes in Computer Science, 862:203–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giordano:1994:ICF

[2998] J.-Y. Giordano. Inference of contextfree grammars by enumeration: Structural containment as an ordering bias. Lecture Notes in Computer Science, 862:212–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wyard:1994:RIC

[2999] P. Wyard. Representational issues for context free grammar induction using genetic algorithms. Lecture Notes in Computer Science, 862:222–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dupont:1994:RGI

[3000] P. Dupont. Regular grammatical inference from positive and negative samples by genetic search: the GIG method. Lecture Notes in Computer Science, 862:236–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wright:1994:TAI

[3001] J. H. Wright, G. J. F. Jones, and H. Lloyd-Thomas. Training and application of integrated grammar/bigram language models. Lecture Notes in Computer Science, 862:246–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Osborne:1994:LUB

[3002] M. Osborne and D. Bridge. Learning unification-based grammars using the spoken English corpus. Lecture Notes in Computer Science, 862:260– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dennis:1994:SOP

[3003] M. D. Dennis, A. M. Wallington, and G. R. Sampson. Stochastic optimization of a probabilistic language model. Lecture Notes in Computer Science, 862:271–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Young:1994:CAG

[3004] S. J. Young and H.-H. Shih. Computer assisted grammar construction. Lecture Notes in Computer Science, 862: 282–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fix:1994:HVE

[3005] L. Fix and F. B. Schneider. Hybrid verification by exploiting the environment. Lecture Notes in Computer Science, 863:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hooman:1994:CRT

[3006] J. Hooman. Correctness of real time systems by construction. Lecture Notes in Computer Science, 863:19–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lamport:1994:SVF

[3007] L. Lamport and S. Merz. Specifying and verifying fault-tolerant systems. Lecture Notes in Computer Science, 863:41–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pnueli:1994:DHS

[3008] A. Pnueli. Development of hybrid systems. Lecture Notes in Computer Science, 863:77–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1994:LDI

[3009] C. Zhou. Linear duration invariants. Lecture Notes in Computer Science, 863:86–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arora:1994:ERT

[3010] A. Arora. Efficient reconfiguration of trees: A case study in methodical design of nonmasking fault-tolerant programs. Lecture Notes in Computer Science, 863:110–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VonderBeeck:1994:CSV

[3011] M. Von der Beeck. A comparison of statecharts variants. Lecture Notes in Computer Science, 863:128–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benveniste:1994:CSS

[3012] A. Benveniste, B. C. Levy, E. Fabre, and P. Le Guernic. A calculus of stochastic systems for the specification, simulation, and hidden state estimation of hybrid stochastic/nonstochastic systems. Lecture Notes in Computer Science, 863:149–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bosscher:1994:VAC

[3013] D. Bosscher, I. Polak, and F. Vaandrager. Verification of an audio control protocol. Lecture Notes in Computer Science, 863:170–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouajjani:1994:VIP

[3014] A. Bouajjani, R. Echahed, and R. Robbana. Verifying invariance properties of timed systems with duration variables. Lecture Notes in Computer Science, 863:193–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boucheneb:1994:PLT

[3015] H. Boucheneb and G. Berthelot. Predicting logical and temporal properties of real-time systems using synchronized elementary nets. Lecture Notes in Computer Science, 863:211–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bradley:1994:DIC

[3016] S. Bradley, W. Henderson, D. Kendall, and A. Robson. Designing and implementing correct real-time systems. Lecture Notes in Computer Science, 863: 228–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broy:1994:SRF

[3017] M. Broy and K. Stoelen. Specification and refinement of finite dataflow networks — a relational approach. Lecture Notes in Computer Science, 863: 247–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gafni:1994:AOS

[3018] V. Gafni, A. Yehudai, and Y. A. Feldman. Activation-oriented specification of real-time systems. Lecture Notes in Computer Science, 863:268–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

He:1994:PCS

[3019] J. He, C. A. R. Hoare, M. Fraenzle, and M. Mueller-Olm. Provably correct systems. Lecture Notes in Computer Science, 863:288–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

He:1994:SAP

[3020] J. He and J. Zheng. Simulation approach to provably correct hardware compilation. Lecture Notes in Computer Science, 863:336–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henzinger:1994:VMD

[3021] T. A. Henzinger and P. W. Kopke. Verification methods for the divergent runs of clock systems. Lecture Notes in Computer Science, 863:351–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janowski:1994:FTB

[3022] T. Janowski. Fault-tolerant bisimulation and process transformations. Lecture Notes in Computer Science, 863: 373–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janssen:1994:LRT

[3023] W. Janssen, M. Poel, Q. Xu, and J. Zwiers. Layering of real-time distributed processes. Lecture Notes in Computer Science, 863:393–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonsson:1994:TRN

[3024] B. Jonsson, C. Ho-Stuart, and W. Yi.

Testing and refinement for nondeterministic and probabilistic processes. Lecture Notes in Computer Science, 863:418–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapur:1994:PSP

[3025] A. Kapur, T. A. Henzinger, Z. Manna, and A. Pnueli. Proving safety properties of hybrid systems. Lecture Notes in Computer Science, 863:431–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kearney:1994:LRT

[3026] P. Kearney and M. Utting. A layered real-time specification of a RISC processor. Lecture Notes in Computer Science, 863:455–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kishore:1994:RTF

[3027] J. K. Kishore, R. S. Manjunatha, V. K.

Agrawal, and N. K. Malik. A real time fault tolerant microprocessor based onboard computer system for INSAT-2 spacecraft. Lecture Notes in Computer Science, 863:476–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lakhneche:1994:RAD

[3028] Y. Lakhneche and J. Hooman. Reasoning about durations in metric temporal logic. Lecture Notes in Computer Science, 863:488–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeLann:1994:SCR

[3029] G. Le Lann. Scheduling in critical realtime systems: a manifesto. Lecture Notes in Computer Science, 863:511– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1994:SDF

[3030] Z. Liu and M. Joseph. Stepwise development of fault-tolerant reactive systems. Lecture Notes in Computer Science, 863:529–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maffeis:1994:DIS

[3031] O. Maffeis and P. Le Guernic. Distributed implementation of SIGNAL: Scheduling and graph clustering. Lecture Notes in Computer Science, 863: 547–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mori:1994:DIC

[3032] R. Mori and N. Yonezaki. Derivation of the input conditional formula from a reactive system specification in temporal logic. Lecture Notes in Computer Science, 863:567–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nadjm-Tehrani:1994:PMC

[3033] S. Nadjm-Tehrani and J.-E. Stroemberg. From physical modelling to compositional models of hybrid systems. Lecture Notes in Computer Science, 863:583–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schenke:1994:STR

[3034] M. Schenke. Specification and transformation of reactive systems with time restrictions and concurrency. Lecture Notes in Computer Science, 863:605– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shyamasundar:1994:LRS

[3035] R. K. Shyamasundar and S. Ramesh.

Languages for reactive specifications: Synchrony vs asynchrony. Lecture Notes in Computer Science, 863:621– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sipma:1994:SVC

[3036] H. B. Sipma and Z. Manna. Specification and verification of controlled systems. Lecture Notes in Computer Science, 863:641–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skakkebaek:1994:TDC

[3037] Jens Ulrik Skakkebaek and N. Shankar.

Towards a duration calculus proof assistant in PVS. Lecture Notes in Computer Science, 863:660–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yi:1994:ARR

[3038] W. Yi. Algebraic reasoning for realtime probabilistic processes with uncertain information. Lecture Notes in Computer Science, 863:680–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilke:1994:STS

[3039] T. Wilke. Specifying timed state sequences in powerful decidable logics and timed automata. Lecture Notes in Computer Science, 863:694–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yu:1994:CHS

[3040] H. Yu, P. K. Pandya, and Y. Sun. A calculus for hybrid sampled data systems. Lecture Notes in Computer Science, 863:716–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yu:1994:FDH

[3041] X. Yu, J. Wang, C. Zhou, and P. K. Pandya. Formal design of hybrid systems. Lecture Notes in Computer Science, 863:738–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zheng:1994:FPD

[3042] Y. Zheng and C. Zhou. A formal proof of the deadline driven scheduler. Lecture Notes in Computer Science, 863:756–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Getzinger:1994:CBA

[3043] T. W. Getzinger. The costs and benefits of abstract interpretation-driven Prolog optimization. Lecture Notes in Computer Science, 864:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanus:1994:MAF

[3044] M. Hanus and F. Zartmann. Mode analysis of functional logic programs. Lecture Notes in Computer Science, 864:26–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aiken:1994:DTC

[3045] A. Aiken and T. K. Lakshman. Directional type checking of logic programs. Lecture Notes in Computer Science, 864:43–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boulanger:1994:SCA

[3046] D. Boulanger and M. Bruynooghe. A systematic construction of abstract domains. Lecture Notes in Computer Science, 864:61–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agesen:1994:CBT

[3047] O. Agesen. Constraint-based type inference and parametric polymorphism. Lecture Notes in Computer Science, 864:78–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mueller:1994:EFA

[3048] F. Mueller and D. B. Whalley. Efficient on-the-fly analysis of program behavior and static cache simulation. Lecture Notes in Computer Science, 864: 101–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harcourt:1994:PTS

[3049] E. Harcourt, J. Mauney, and T. Cook. From processor timing specifications to static instruction scheduling. Lecture Notes in Computer Science, 864:116– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Banerjee:1994:SST

[3050] A. Banerjee and D. A. Schmidt. Stackability in the simply-typed call-by-value lambda calculus. Lecture Notes in Computer Science, 864:131–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goubault:1994:GBC

[3051] J. Goubault. Generalized boxings, congruences and partial inlining. Lecture Notes in Computer Science, 864:147– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steckler:1994:ST

[3052] P. Steckler and M. Wand. Selective thunkification. Lecture Notes in Computer Science, 864:162–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halbwachs:1994:ASP

[3053] N. Halbwachs. About synchronous programming and abstract interpretation. Lecture Notes in Computer Science, 864:179–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garza-Salazar:1994:UCA

[3054] D. Garza-Salazar and W. Boehm.

Uniqueness and completeness analysis of array comprehensions. Lecture Notes in Computer Science, 864:193– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lisper:1994:EAD

[3055] B. Lisper and J.-F. Collard. Extent analysis of data fields. Lecture Notes in Computer Science, 864:208–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halbwachs:1994:VLH

[3056] N. Halbwachs, Y.-E. Proy, and P. Raymond. Verification of linear hybrid systems by means of convex approximations. Lecture Notes in Computer Science, 864:223–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yi:1994:CTD

[3057] K. Yi. Compile-time detection of uncaught exceptions in standard ML programs. Lecture Notes in Computer Science, 864:238–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Debray:1994:ECC

[3058] S. K. Debray, P. Lopez Garcia, M. Hermenegildo, and N.-W. Lin. Estimating the computational cost of logic programs. Lecture Notes in Computer Science, 864:255–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Armstrong:1994:BFD

[3059] T. Armstrong, K. Marriott, P. Schachte, and H. Soendergaard. Boolean functions for dependency analysis: Algebraic properties and efficient representation. Lecture Notes in Computer Science, 864:266–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Codish:1994:DPT

[3060] M. Codish and B. Demoen. Deriving polymorphic type dependencies for logic programs using multiple incarnations of prop. Lecture Notes in Computer Science, 864:281–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gras:1994:ENS

[3061] D. C. Gras and M. Hermenegildo. Extracting non-strict independent andparallelism using sharing and freeness information. Lecture Notes in Computer Science, 864:297–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vergauwen:1994:EFC

[3062] B. Vergauwen, J. Wauman, and J. Lewi. Efficient FixPoint computation. Lecture Notes in Computer Science, 864:314–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joergensen:1994:FFF

[3063] N. Joergensen. Finding fixpoints in finite function spaces using neededness analysis and chaotic iteration. Lecture Notes in Computer Science, 864:329– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jensen:1994:ESA

[3064] K. D. Jensen, P. Hjoeresen, and M. Rosendahl. Efficient strictness analysis of haskell. Lecture Notes in Computer Science, 864:346–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mauborgne:1994:AIU

[3065] L. Mauborgne. Abstract interpretation using TDGs. Lecture Notes in Computer Science, 864:363–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hankin:1994:TBF

[3066] C. Hankin and D. Le Metayer. A typebased framework for program analysis. Lecture Notes in Computer Science, 864:380–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henglein:1994:IFP

[3067] F. Henglein. Iterative fixed point computation for type-based strictness analysis. Lecture Notes in Computer Science, 864:395–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Solberg:1994:STA

[3068] K. L. Solberg, H. R. Nielson, and F. Nielson. Strictness and totality analysis. Lecture Notes in Computer Science, 864:408–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Consel:1994:FSA

[3069] C. Consel. Fast strictness analysis via symbolic fixpoint iteration. Lecture Notes in Computer Science, 864:423– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glueck:1994:GTD

[3070] R. Glueck and J. Joergensen. Generating transformers for deforestation and supercompilation. Lecture Notes in Computer Science, 864:432–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khoshnevisan:1994:MEC

[3071] H. Khoshnevisan and M. Afshar. Mechanical elimination of commutative redundancy. Lecture Notes in Computer Science, 864:449–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radcliffe:1994:FMA

[3072] N. J. Radcliffe and P. D. Surry. Formal memetic algorithms. Lecture Notes in Computer Science, 865:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shapiro:1994:SMF

[3073] J. Shapiro, A. Pruegel-Bennett, and M. Rattray. A statistical mechanical formulation of the dynamics of genetic algorithms. Lecture Notes in Computer Science, 865:17–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yates:1994:ESS

[3074] D. F. Yates and A. Fairley. Evolutionary stability in simple classifier systems. Lecture Notes in Computer Science, 865:28–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Field:1994:NTG

[3075] P. Field. Nonbinary transforms for genetic algorithm problems. Lecture Notes in Computer Science, 865:38–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paton:1994:EEC

[3076] R. Paton. Enhancing evolutionary computation using analogues of biological mechanisms. Lecture Notes in Computer Science, 865:51–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miller:1994:EMC

[3077] G. F. Miller. Exploiting mate choice in evolutionary computation: Sexual selection as a process of search, optimisation and diversification. Lecture Notes in Computer Science, 865:65–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hancock:1994:ECS

[3078] P. J. B. Hancock. An empirical comparison of selection methods in evolutionary algorithms. Lecture Notes in Computer Science, 865:80–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bull:1994:ESG

[3079] L. Bull and T. C. Fogarty. An evolution strategy and genetic algorithm hybrid: An initial implementation and first results. Lecture Notes in Computer Science, 865:95–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coyne:1994:GAD

[3080] J. Coyne and R. Paton. Genetic algorithms and directed adaptation. Lecture Notes in Computer Science, 865: 103–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reeves:1994:GAN

[3081] C. R. Reeves. Genetic algorithms and neighbourhood search. Lecture Notes in Computer Science, 865:115– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapsalis:1994:UPP

[3082] A. Kapsalis, G. D. Smith, and V. J. Rayward-Smith. A unified paradigm for parallel genetic algorithms. Lecture Notes in Computer Science, 865: 131–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Husbands:1994:DCG

[3083] P. Husbands. Distributed coevolutionary genetic algorithms for multicriteria and multi-constraint optimisation. Lecture Notes in Computer Science, 865:150–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fairley:1994:IOR

[3084] A. Fairley and D. F. Yates. Inductive operators and rule repair in a hybrid genetic learning system: Some initial results. Lecture Notes in Computer Science, 865:166–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patel:1994:ALR

[3085] M. J. Patel and M. Dorigo. Adaptive learning of a robot arm. Lecture Notes in Computer Science, 865:180– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fogarty:1994:CEC

[3086] T. C. Fogarty. Co-evolving co-operative populations of rules in learning control systems. Lecture Notes in Computer Science, 865:195–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carse:1994:LAB

[3087] B. Carse. Learning anticipatory behaviour using a delayed action classifier system. Lecture Notes in Computer Science, 865:210–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Melhuish:1994:ARM

[3088] C. Melhuish and T. C. Fogarty. Applying a restricted mating policy to determine state space niches using immediate and delayed reinforcement. Lecture Notes in Computer Science, 865: 224–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pipe:1994:CBT

[3089] A. G. Pipe and B. Carse. A comparison between two architectures for searching and learning in maze problems. Lecture Notes in Computer Science, 865:238–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corne:1994:FPE

[3090] D. Corne, P. Ross, and H.-L. Fang. Fast practical evolutionary timetabling. Lecture Notes in Computer Science, 865:250–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paechter:1994:OPT

[3091] B. Paechter. Optimising a presentation timetable using evolutionary algorithms. Lecture Notes in Computer Science, 865:264–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cartwright:1994:GAF

[3092] H. M. Cartwright and A. L. Tuson. Genetic algorithms and flowshop scheduling: Towards the development of a real-time process control system. Lecture Notes in Computer Science, 865:277–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

White:1994:GAD

[3093] M. S. White and S. J. Flockton. Genetic algorithms for digital signal processing. Lecture Notes in Computer Science, 865:291–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beasley:1994:CRU

[3094] D. Beasley, D. R. Bull, and R. R.

Martin. Complexity reduction using expansive coding. Lecture Notes in Computer Science, 865:304–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Howard:1994:AGP

[3095] E. Howard and N. Oakley. The application of genetic programming to the investigation of short, noisy, chaotic data series. Lecture Notes in Computer Science, 865:320–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Whitley:1994:LEB

[3096] D. Whitley, V. S. Gordon, and K. Mathias. Lamarckian evolution, the baldwin effect and function optimization. Lecture Notes in Computer Science, 866:6–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mattfeld:1994:CPP

[3097] D. C. Mattfeld, H. Kopfer, and C. Bierwirth. Control of parallel population dynamics by social-like behavior of GA-Individuals. Lecture Notes in Computer Science, 866:16–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chiva:1994:SGP

[3098] E. Chiva and P. Tarroux. Studying genotype-phenotype interactions: A model of the evolution of the cell regulation network. Lecture Notes in Computer Science, 866:26–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yoshida:1994:DGA

[3099] Y. Yoshida and N. Adachi. A diploid genetic algorithm for preserving population diversity — pseudo-meiosis GA. Lecture Notes in Computer Science, 866:36–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paredis:1994:CEC

[3100] J. Paredis. Co-evolutionary constraint satisfaction. Lecture Notes in Computer Science, 866:46–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beyer:1994:TTE

[3101] H.-G. Beyer. Towards a theory of evolution strategies: Results for (1+, λ)- strategies on (nearly) arbitrary fitness functions. Lecture Notes in Computer Science, 866:58–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dzubera:1994:ACA

[3102] J. Dzubera and D. Whitley. Advanced correlation analysis of operators for the traveling salesman problem. Lecture Notes in Computer Science, 866:68–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eiben:1994:GAM

[3103] A. E. Eiben, P.-E. Raue, and Z. Ruttkay. Genetic algorithms with multiparent recombination. Lecture Notes in Computer Science, 866:78–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asoh:1994:MCT

[3104] H. Asoh and H. Muehlenbein. On the mean convergence time of evolutionary algorithms without selection and mutation. Lecture Notes in Computer Science, 866:88–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asoh:1994:EHD

[3105] H. Asoh and H. Muehlenbein. Estimating the heritability by decomposing the genetic variance. Lecture Notes in Computer Science, 866:98–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bui:1994:AHS

[3106] T. N. Bui and B.-R. Moon. Analyzing hyperplane synthesis in genetic algorithms using clustered schemata. Lecture Notes in Computer Science, 866: 108–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thierens:1994:CMG

[3107] D. Thierens and D. Goldberg. Convergence models of genetic algorithm selection schemes. Lecture Notes in Computer Science, 866:119–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakano:1994:OPS

[3108] R. Nakano, Y. Davidor, and T. Yamada. Optimal population size under constant computation cost. Lecture Notes in Computer Science, 866: 130–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rudolph:1994:EAI

[3109] G. Rudolph. An evolutionary algorithm for integer programming. Lecture Notes in Computer Science, 866: 139–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horn:1994:LPP

[3110] J. Horn, D. E. Goldberg, and K. Deb. Long path problems. Lecture Notes in Computer Science, 866:149–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hammel:1994:ESN

[3111] U. Hammel and T. Baeck. Evolution strategies on noisy functions: How to improve convergence properties. Lecture Notes in Computer Science, 866:159–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pal:1994:SSS

[3112] K. F. Pal. Selection schemes with spatial isolation for genetic optimization. Lecture Notes in Computer Science, 866:170–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tang:1994:MER

[3113] A. Y.-C. Tang and K.-S. Leung. A modified edge recombination operator for the travelling salesman problem. Lecture Notes in Computer Science, 866:180–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ostermeier:1994:SSA

[3114] A. Ostermeier, A. Gawelczyk, and N. Hansen. Step-size adaptation based on non-local use of selection information. Lecture Notes in Computer Science, 866:189–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schlierkamp-Voosen:1994:SAC

[3115] D. Schlierkamp-Voosen and H. Muehlenbein. Strategy adaptation by competing subpopulations. Lecture Notes in Computer Science, 866:199–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sebag:1994:CCT

[3116] M. Sebag and M. Schoenauer. Controlling crossover through inductive learning. Lecture Notes in Computer Science, 866:209–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lorenz:1994:DOP

[3117] D. H. Lorenz. Derivative operators for preference predicate evolution. Lecture Notes in Computer Science, 866:219– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

White:1994:ACU

[3118] T. White and F. Oppacher. Adaptive crossover using automata. Lecture Notes in Computer Science, 866:229– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakanashi:1994:CDG

[3119] H. Sakanashi, K. Suzuki, and Y. Kakazu. Controlling dynamics of GA through filtered evaluation function. Lecture Notes in Computer Science, 866:239–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Potter:1994:CCA

[3120] M. A. Potter and K. A. De Jong. A cooperative coevolutionary approach to function optimization. Lecture Notes in Computer Science, 866:249–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carse:1994:FCS

[3121] B. Carse and T. C. Fogarty. A fuzzy classifier system using the pittsburgh approach. Lecture Notes in Computer Science, 866:260–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giani:1994:QLE

[3122] A. Giani, F. Baiardi, and A. Starita. Q-learning in evolutionary rule-based systems. Lecture Notes in Computer Science, 866:270–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartmann:1994:CLC

[3123] U. Hartmann. On the complexity of learning in classifier systems. Lecture Notes in Computer Science, 866:280– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wineberg:1994:RSP

[3124] M. Wineberg and F. Oppacher. A representation scheme to perform program induction in a canonical genetic algorithm. Lecture Notes in Computer Science, 866:292–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iba:1994:GPL

[3125] H. Iba, H. De Garis, and T. Sato.

Genetic programming with local hillclimbing. Lecture Notes in Computer Science, 866:302–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gathercole:1994:DTS

[3126] C. Gathercole and P. Ross. Dynamic training subset selection for supervised learning in genetic programming. Lecture Notes in Computer Science, 866:312–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Banzhaf:1994:GPM

[3127] W. Banzhaf. Genotype-phenotype-mapping and neutral variation — A case study in genetic programming. Lecture Notes in Computer Science, 866:322–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacob:1994:GSP

[3128] C. Jacob. Genetic L-system programming. Lecture Notes in Computer Science, 866:334–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Das:1994:GAD

[3129] R. Das, M. Mitchell, and J. P. Crutchfield. A genetic algorithm discovers particle-based computation in cellular automata. Lecture Notes in Computer Science, 866:344–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeOliveira:1994:SEB

[3130] P. P. B. De Oliveira. Simulation of exaptive behaviour. Lecture Notes in Computer Science, 866:354–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inayoshi:1994:ASP

[3131] H. Inayoshi. Artificial spacing patterns in a network of interacting celloids. Lecture Notes in Computer Science, 866:365–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muruzabal:1994:DPL

[3132] J. Muruzabal and A. Muthoz. Diffuse pattern learning with fuzzy ARTMAP and PASS. Lecture Notes in Computer Science, 866:376–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heistermann:1994:DLA

[3133] J. Heistermann. Different learning algorithms for neural networks — A comparative study. Lecture Notes in Computer Science, 866:386–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OReilly:1994:PSH

[3134] U.-M. O’Reilly and F. Oppacher. Program search with a hierarchical variable length representation: Genetic programming, simulated annealing and hill climbing. Lecture Notes in Computer Science, 866:397–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laursen:1994:PIP

[3135] P. S. Laursen. Problem-independent parallel simulated annealing using selection and migration. Lecture Notes in Computer Science, 866:408–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeck:1994:POE

[3136] T. Baeck. Parallel optimization of evolutionary algorithms. Lecture Notes in Computer Science, 866:418–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1994:PSA

[3137] H. Chen and N. S. Flann. Parallel simulated annealing and genetic algorithms: A space of hybrid methods. Lecture Notes in Computer Science, 866:428–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braun:1994:EMH

[3138] H. Braun and P. Zagorski. ENZO-M — A hybrid approach for optimizing neural networks by evolution and learning. Lecture Notes in Computer Science, 866:440–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ronald:1994:GLE

[3139] E. Ronald and M. Schoenauer. Genetic lander: An experiment in accurate neuro-genetic control. Lecture Notes in Computer Science, 866:452–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1994:EOR

[3140] B.-T. Zhang. Effects of Occam’s razor in evolving sigma-pi neural nets. Lecture Notes in Computer Science, 866: 462–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Born:1994:DNN

[3141] J. Born, I. Santibathez-Koref, and H.- M. Voigt. Designing neural networks by adaptively building blocks in cascades. Lecture Notes in Computer Science, 866:472–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pipe:1994:HAH

[3142] A. G. Pipe, T. C. Fogarty, and A. Winfield. Hybrid adaptive heuristic critic architectures for learning in mazes with continuous search spaces. Lecture Notes in Computer Science, 866:482– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Utecht:1994:MOS

[3143] U. Utecht and K. Trint. Mutation operators for structure evolution of neural networks. Lecture Notes in Computer Science, 866:492–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seredynski:1994:LCD

[3144] F. Seredynski. Loosely coupled distributed genetic algorithms. Lecture Notes in Computer Science, 866:514– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Higuchi:1994:AEH

[3145] T. Higuchi, H. Iba, and B. Manderick. Applying evolvable hardware to autonomous agents. Lecture Notes in Computer Science, 866:524–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marin:1994:GAL

[3146] F. J. Marin, O. Trelles-Salazar, and F. Sandoval. Genetic algorithms on LAN-Message passing architectures using application to the routing problem. Lecture Notes in Computer Science, 866:534–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hill:1994:GAB

[3147] A. M. Hill and S.-M. Kang. Genetic algorithm based design optimization of CMOS VLSI circuits. Lecture Notes in Computer Science, 866:546–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ross:1994:IET

[3148] P. Ross, D. Corne, and H.-L. Fang. Improving evolutionary timetabling with delta evaluation and directed mutation. Lecture Notes in Computer Science, 866:556–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanWezel:1994:GIR

[3149] M. C. Van Wezel, J. N. Kok, J. Van den Berg, and W. Van Kampen. Genetic improvement of railway timetables. Lecture Notes in Computer Science, 866:566–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cobb:1994:UGA

[3150] H. G. Cobb and P. Bock. Using a genetic algorithm to search for the representational bias of a collective reinforcement learner. Lecture Notes in Computer Science, 866:576–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lienig:1994:EAR

[3151] J. Lienig and H. Brandt. An evolutionary algorithm for the routing of multi-chip modules. Lecture Notes in Computer Science, 866:588–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pullen:1994:SDU

[3152] S. P. Pullen and B. W. Parkinson. System design under uncertainty: Evolutionary optimization of the gravity probe-B spacecraft. Lecture Notes in Computer Science, 866:598–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karcz-Duleba:1994:SSD

[3153] I. Karcz-Duleba. Soft selection in D-optimal designs. Lecture Notes in Computer Science, 866:608–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inayoshi:1994:WGB

[3154] H. Inayoshi and B. Manderick. The weighted graph bi-partitioning problem: A look at GA performance. Lecture Notes in Computer Science, 866: 617–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Surry:1994:RLP

[3155] P. D. Surry and N. J. Radcliffe. RPL2:

A language and parallel framework for evolutionary computing. Lecture Notes in Computer Science, 866:628– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schreiber:1994:CCC

[3156] G. Schreiber, B. Wielinga, H. Akkermans, and W. Van de Velde. CML: The CommonKADS conceptual modelling language. Lecture Notes in Computer Science, 867:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruiz:1994:EFM

[3157] F. Ruiz, F. Van Harmelen, M. Aben, and J. Van de Plassche. Evaluating a formal modelling language. Lecture Notes in Computer Science, 867:26–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fensel:1994:CTA

[3158] D. Fensel and K. Poeck. A comparison of two approaches to model-based knowledge acquisition. Lecture Notes in Computer Science, 867:46–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dieng:1994:ABK

[3159] R. Dieng, O. Corby, and S. Labidi. Agent-based knowledge acquisition. Lecture Notes in Computer Science, 867:63–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmalhofer:1994:BKL

[3160] F. Schmalhofer, J. Stuart Aitken, and L. E. Bourne. Beyond the knowledge level: Descriptions of rational behavior for sharing and reuse. Lecture Notes in Computer Science, 867:83–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Compton:1994:LPP

[3161] P. Compton, P. Preston, B. Kang, and T. Yip. Local patching produces compact knowledge bases. Lecture Notes in Computer Science, 867:104–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breuker:1994:CPS

[3162] J. Breuker. Components of problem solving and types of problems. Lecture Notes in Computer Science, 867: 118–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benjamins:1994:RPS

[3163] R. Benjamins. On a role of problem solving methods in knowledge acquisition — experiments with diagnostic strategies. Lecture Notes in Computer Science, 867:137–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cathamero:1994:MPR

[3164] D. Cathamero. Modelling plan recognition for decision support. Lecture Notes in Computer Science, 867:158– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanHeijst:1994:COB

[3165] G. Van Heijst and G. Schreiber. CUE:

Ontology-based knowledge acquisition. Lecture Notes in Computer Science, 867:178–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pirlein:1994:KIE

[3166] T. Pirlein and R. Studer. KARO: An integrated environment for reusing ontologies. Lecture Notes in Computer Science, 867:200–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geldof:1994:VMG

[3167] S. Geldof and A. Slodzian. From verification to modelling guidelines. Lecture Notes in Computer Science, 867: 226–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andrienko:1994:AAC

[3168] G. L. Andrienko and N. V. Andrienko.

AFORIZM approach: Creating situations to facilitate expertise transfer. Lecture Notes in Computer Science, 867:244–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aussenac-Gilles:1994:HCD

[3169] N. Aussenac-Gilles. How to combine data abstraction and model refinement: a methodological contribution in MACAO. Lecture Notes in Computer Science, 867:262–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brazier:1994:UCK

[3170] F. M. T. Brazier and J. Treur. User centered knowledge-based system design: a formal modelling approach. Lecture Notes in Computer Science, 867: 283–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tourtier:1994:VES

[3171] P.-A. Tourtier and S. Boyera. Validating at early stages with a causal simulation tool. Lecture Notes in Computer Science, 867:303–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yost:1994:SFA

[3172] G. R. Yost, G. Klinker, M. Linster, and D. Marques. The SBF framework, 1989–1994: From applications to workplaces. Lecture Notes in Computer Science, 867:318–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Major:1994:ARM

[3173] N. Major, J. Cupit, and N. Shadbolt. Applying the REKAP methodology to situation assessment. Lecture Notes in Computer Science, 867:342–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arcos:1994:ILK

[3174] J. L. Arcos and E. Plaza. Integration of learning into a knowledge modelling framework. Lecture Notes in Computer Science, 867:355–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rouveirol:1994:KLM

[3175] C. Rouveirol and P. Albert. Knowledge level model of a configurable learning system. Lecture Notes in Computer Science, 867:374–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stroulia:1994:RSA

[3176] E. Stroulia and A. K. Goel. Reflective, self-adaptive problem solvers. Lecture Notes in Computer Science, 867:394– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Georganas:1994:SSF

[3177] N. D. Georganas. Self-similar (“fractal”) traffic driving an ATM networks. Lecture Notes in Computer Science, 868:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nahrstedt:1994:SKM

[3178] K. Nahrstedt and J. Smith. A service kernel for multimedia endstations. Lecture Notes in Computer Science, 868: 8–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reinhardt:1994:ARN

[3179] W. Reinhardt. Advance reservation of network resources for multimedia applications. Lecture Notes in Computer Science, 868:23–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaeppner:1994:MSD

[3180] T. Kaeppner and L. C. Wolf. Media scaling in distributed multimedia object services. Lecture Notes in Computer Science, 868:34–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Achilleoudis:1994:ICT

[3181] N. Achilleoudis, J. Lamour, and F. Daffara. An interactive cable television network for multimedia applications. Lecture Notes in Computer Science, 868:44–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertin:1994:EIP

[3182] C. Bertin. Eurescom IMS1 projects (integrated multimedia services at about 1 mbit/s). Lecture Notes in Computer Science, 868:53–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sung:1994:HIR

[3183] J.-J. Sung, M.-Y. Huh, H.-J. Kim, and J.-H. Hahm. Hypermedia information retrieval system using MHEG coded representation in a networked environment. Lecture Notes in Computer Science, 868:67–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Droz:1994:TAL

[3184] P. Droz and J.-Y. Le Boudec. Transparent ATM LAN interconnection over ISDN. Lecture Notes in Computer Science, 868:78–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wittig:1994:CUM

[3185] H. Wittig, L. C. Wolf, and C. Vogt.

CPU utilization of multimedia processes: HEIPOET — the Heidelberg Predictor of Execution Times Measurement Tool. Lecture Notes in Computer Science, 868:92–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barth:1994:ERM

[3186] I. Barth. Extending the rate-monotonic scheduling algorithm to get shorter delays. Lecture Notes in Computer Science, 868:104–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1994:MAA

[3187] W. Liu. A multimedia application adaptation layer (MAAL) protocol. Lecture Notes in Computer Science, 868:115–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carle:1994:ALG

[3188] G. Carle. Adaptation layer and group communication server for reliable multipoint services in ATM networks. Lecture Notes in Computer Science, 868: 124–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boecking:1994:TTI

[3189] S. Boecking and P. Vindeby. TIP: A transport and internetworking package for ATM. Lecture Notes in Computer Science, 868:139–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pronios:1994:MHS

[3190] N. B. Pronios and T. Bozios. Multimedia and hypermedia synchronisation: A unified framework. Lecture Notes in Computer Science, 868:153–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koehler:1994:MPS

[3191] D. Koehler and H. Mueller. Multimedia playout synchronisation using buffer level control. Lecture Notes in Computer Science, 868:167–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lamparter:1994:ECC

[3192] B. Lamparter and W. Effelsberg. eXtended color cell compression — A runtime-efficient compression scheme for software video. Lecture Notes in Computer Science, 868:181–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yeadon:1994:QAF

[3193] N. Yeadon, F. Garcia, A. Campbell, and D. Hutchison. QoS adaptation and flow filtering in ATM networks. Lecture Notes in Computer Science, 868:191–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frings:1994:ORM

[3194] J. Frings and T. Bauschert. Optimal resource management in ATM transport networks supporting multimedia traffic. Lecture Notes in Computer Science, 868:203–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bahr:1994:ISF

[3195] K. Bahr, E. Hinsch, and G. Schulze. Incorporating security functions in multimedia conferencing applications in the context of the MICE project. Lecture Notes in Computer Science, 868:216– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bunz:1994:SMA

[3196] H. Bunz, A. Bertsch, M. Jurecic, and B. Baum-Waidner. Secure multimedia applications and teleservices — security requirements and prototype for health care. Lecture Notes in Computer Science, 868:224–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Altenhofen:1994:BMT

[3197] M. Altenhofen, J. Schaper, and S. Thomas. The BERKOM multimedia teleservices. Lecture Notes in Computer Science, 868:237–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rozek:1994:CMCa

[3198] A. Rozek and P. Christ. The CIO multimedia communication platform. Lecture Notes in Computer Science, 868: 251–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mauthe:1994:RSG

[3199] A. Mauthe, D. Hutchison, G. Coulson, and S. Namuye. From requirements to services: Group communication support for distributed multimedia systems. Lecture Notes in Computer Science, 868:266–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blakowski:1994:HIF

[3200] G. Blakowski and W. Steinbeck. Harmonization of an infrastructure for flexible distance learning in Europe with CTA. Lecture Notes in Computer Science, 868:280–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Velez:1994:DMT

[3201] J. M. Velez, M. R. Gomes, B. Tritsch, and C. Hornung. Dedicated — modular training system (DELTA project D2014). Lecture Notes in Computer Science, 868:300–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pusztaszeri:1994:MTT

[3202] Y.-H. Pusztaszeri, M. Alou, E. W. Biersack, and P. Dubois. Multimedia teletutoring over a Trans-European ATM network. Lecture Notes in Computer Science, 868:315–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmidt:1994:FST

[3203] T. Schmidt, J. Schweitzer, and M. Weber. A framework for synchronous tele-cooperation. Lecture Notes in Computer Science, 868:328–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leydekkers:1994:MCS

[3204] P. Leydekkers and V. Gay. Multimedia conferencing services in an open distributed environment. Lecture Notes in Computer Science, 868:339–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeMiguel:1994:IED

[3205] T. P. De Miguel, S. Pavon, J. Salvachua, and J. Quemada Vives. IS-ABEL — experimental distributed cooperative work application over broadband networks. Lecture Notes in Computer Science, 868:353–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strack:1994:DIC

[3206] R. Strack, R. Cordes, and D. C. Sutcliffe. Demonstrating image communication within open distributed environments. Lecture Notes in Computer Science, 868:363–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dempsey:1994:DIH

[3207] B. J. Dempsey, M. T. Lucas, and A. C. Weaver. Design and implementation of a high quality video distribution system using XTP reliable multicast. Lecture Notes in Computer Science, 868:376–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strayer:1994:OOI

[3208] W. T. Strayer, S. Gray, and R. E. Cline. An object-oriented implementation of the Xpress transfer protocol. Lecture Notes in Computer Science, 868:387–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rakow:1994:DMA

[3209] T. C. Rakow, P. Dettling, F. Moser, and B. Paul. Development of a multimedia archiving teleservice using the DFR standard. Lecture Notes in Computer Science, 868:401–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaeckenhoff:1994:MMO

[3210] R. Kaeckenhoff, D. Merten, and K. Meyer-Wegener. MOSS as a multimedia-object server. Lecture Notes in Computer Science, 868:413– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gervais:1994:UPT

[3211] M.-P. Gervais. The universal personal telecommunication service in a multioperator environment. Lecture Notes in Computer Science, 868:426–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eckardt:1994:PCI

[3212] T. Eckardt and T. Magedanz. On the personal communication impacts on multimedia teleservices. Lecture Notes in Computer Science, 868:435– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Campbell:1994:DES

[3213] W. J. Campbell, N. M. Short, P. Coronado, and R. F. Cromp. Distributed earth science validation centers for mission to planet Earth. Lecture Notes in Computer Science, 869:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chandrasekaran:1994:CUR

[3214] B. Chandrasekaran. Causal understanding in reasoning about the world. Lecture Notes in Computer Science, 869:13–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fox:1994:HMI

[3215] E. A. Fox. How to make intelligent digital libraries. Lecture Notes in Computer Science, 869:27–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meersman:1994:SMR

[3216] R. A. Meersman. Some methodology and representation problems for the semantics of prosaic application domains. Lecture Notes in Computer Science, 869:39–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jamil:1994:RCE

[3217] H. M. Jamil and F. Sadri. Recognizing credible experts in inaccurate databases. Lecture Notes in Computer Science, 869:46–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kundu:1994:FLL

[3218] S. Kundu and J. Chen. Fuzzy logic or Lukasiewicz logic: a clarification. Lecture Notes in Computer Science, 869: 56–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1994:LSA

[3219] T. Y. Lin, Q. Liu, and Y. Y. Yao. Logic systems for approximate reasoning: via rough sets and topology. Lecture Notes in Computer Science, 869:65–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1994:SFF

[3220] J. J. Lu, N. V. Murray, and E. Rosenthal. Signed formulas and fuzzy operator logics. Lecture Notes in Computer Science, 869:75–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Polkowski:1994:RM

[3221] L. Polkowski and A. Skowron. Rough mereology. Lecture Notes in Computer Science, 869:85–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shi:1994:NRU

[3222] S. L. Shi, M. E. Hull, and D. A. Bell. A new rule for updating evidence. Lecture Notes in Computer Science, 869:95–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1994:GMA

[3223] Z. W. Wang and S. K. M. Wong. A global measure of ambiguity for classification. Lecture Notes in Computer Science, 869:105–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zilberstein:1994:MLC

[3224] S. Zilberstein. Meta-level control of approximate reasoning: a decision theoretic approach. Lecture Notes in Computer Science, 869:114–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michalewicz:1994:GAK

[3225] Z. Michalewicz and J. Arabas. Genetic algorithms for the 0/1 knapsack problem. Lecture Notes in Computer Science, 869:134–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andreasen:1994:WFR

[3226] T. Andreasen and O. Pivert. On the weakening of fuzzy relational queries. Lecture Notes in Computer Science, 869:144–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chang:1994:TQR

[3227] Y. Chang, L. Raschid, and B. Dorr.

Transforming queries from a relational schema to an object schema: a prototype based on F-logic. Lecture Notes in Computer Science, 869:154–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheung:1994:KDD

[3228] D. W.-I. Cheung, A. W.-C. Fu, and J. Han. Knowledge discovery in databases: a rule-based attributeoriented approach. Lecture Notes in Computer Science, 869:164–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cross:1994:CBR

[3229] S. E. Cross, D. F. Roberts, A. M. Mulvehill, and J. A. Sears. Case-based reasoning applied to a force generation decision aid. Lecture Notes in Computer Science, 869:174–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fouque:1994:CBR

[3230] G. Fouque, W. W. Chu, and H. Yau. A case-based reasoning approach for associative query answering. Lecture Notes in Computer Science, 869:183– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greco:1994:EER

[3231] S. Greco and C. Zaniolo. Efficient execution of recursive queries through controlled binding propagation. Lecture Notes in Computer Science, 869:193– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeStrugeon:1994:TDM

[3232] E. Le Strugeon, R. Mandiau, and G. Libert. Towards a dynamic multiagent organization. Lecture Notes in Computer Science, 869:203–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lever:1994:PPA

[3233] J. Lever and B. Richards. parc-Plan: a planning architecture with parallel actions, resources and constraints. Lecture Notes in Computer Science, 869:213–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quafafou:1994:GII

[3234] M. Quafafou. GAITS II: an intelligent system for computer-aided education. Lecture Notes in Computer Science, 869:223–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhong:1994:GDS

[3235] N. Zhong and S. Ohsuga. The GLS discovery system: its goal, architecture and current results. Lecture Notes in Computer Science, 869:233–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Analyti:1994:DSC

[3236] A. Analyti and S. Pramanik. Declarative semantics for contradictory modular logic programs. Lecture Notes in Computer Science, 869:245–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brusoni:1994:LGP

[3237] V. Brusoni, L. Console, B. Pernici, and P. Terenziani. LaTeR: a general purpose manager of temporal information. Lecture Notes in Computer Science, 869:255–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gan:1994:USC

[3238] H. Gan. Understanding a story with causal relationships. Lecture Notes in Computer Science, 869:265–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Terenziani:1994:DQQ

[3239] P. Terenziani. Dealing with qualitative and quantitative temporal information concerning periodic events. Lecture Notes in Computer Science, 869: 275–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xiang:1994:DMA

[3240] Y. Xiang. Distributed multi-agent probabilistic reasoning with Bayesian networks. Lecture Notes in Computer Science, 869:285–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calmet:1994:BBB

[3241] J. Calmet and I. A. Tjandra. Building bridges between knowledge representation and algebraic specification. Lecture Notes in Computer Science, 869: 295–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hertzberg:1994:TAF

[3242] J. Hertzberg and S. Thiebaux. Turning an action formalism into a planner — essentials of a case study. Lecture Notes in Computer Science, 869:305– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Komorowski:1994:TRD

[3243] J. Komorowski and S. Trcek. Towards refinement of definite logic programs. Lecture Notes in Computer Science, 869:315–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lowry:1994:AAP

[3244] M. Lowry, A. Philpot, T. Pressburger, and I. Underwood. AMPHION: automatic programming for scientific subroutine libraries. Lecture Notes in Computer Science, 869:326–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ade:1994:RIT

[3245] H. Ade, B. Malfait, and L. De Raedt.

RUTH: an ILP theory revision system. Lecture Notes in Computer Science, 869:336–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bazan:1994:DRT

[3246] J. G. Bazan, A. Skowron, and P. Synak.

Dynamic reducts as a tool for extracting laws from decisions tables. Lecture Notes in Computer Science, 869:346– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Botta:1994:LFO

[3247] M. Botta. Learning first order theories. Lecture Notes in Computer Science, 869:356–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eick:1994:EEB

[3248] C. F. Eick and E. Toto. Evaluation and enhancement of Bayesian rule-sets in a genetic algorithm learning environment for classification tasks. Lecture Notes in Computer Science, 869:366–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esposito:1994:TPW

[3249] F. Esposito, D. Malerba, G. Semeraro, and C. Brunk. Traps and pitfalls when learning logical definitions from relations. Lecture Notes in Computer Science, 869:376–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hu:1994:DRS

[3250] X. Hu, N. Shan, N. Cercone, and W. Ziarko. DBROUGH: a rough set based knowledge discovery system. Lecture Notes in Computer Science, 869:386–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopez-Suarez:1994:RRB

[3251] A. Lopez-Suarez and M. Kamel. Restructuring rule bases to improve performance. Lecture Notes in Computer Science, 869:396–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCluskey:1994:LHO

[3252] T. L. McCluskey and J. M. Porteous. Learning heuristics for ordering plan goals through static operator analysis. Lecture Notes in Computer Science, 869:406–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michalski:1994:LPO

[3253] R. S. Michalski and I. F. Imam. Learning problem-oriented decision structures from decision rules: the AQDT-2 system. Lecture Notes in Computer Science, 869:416–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nissan:1994:TFA

[3254] E. Nissan, H. Siegelmann, A. Galperin, and S. Kimhi. Towards full automation of the discovery of heuristics in a nuclear engineering project: integration with a neural information language. Lecture Notes in Computer Science, 869:427–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Troxel:1994:CHR

[3255] M. Troxel, K. Swarm, R. Zembowicz, and J. M. Zytkow. Concept hierarchies: a restricted form of knowledge derived from regularities. Lecture Notes in Computer Science, 869:437–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1994:DDA

[3256] J. Zhang and H.-H. Lu. A datadriven approach to feature construction. Lecture Notes in Computer Science, 869:448–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asher:1994:RAA

[3257] N. Asher. Reasoning about action and time with epistemic conditionals. Lecture Notes in Computer Science, 869:458–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buvac:1994:SPC

[3258] S. Buvac, V. Buvac, and I. A. Mason. The semantics of propositional contexts. Lecture Notes in Computer Science, 869:468–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1994:GLO

[3259] J. Chen. The generalized logic of only knowing (GOL) that covers the notion of epistemic specifications. Lecture Notes in Computer Science, 869: 478–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cholvy:1994:RAS

[3260] L. Cholvy, R. Demolombe, and A. Jones. Reasoning about the safety of information: from logical formalization to operational definition. Lecture Notes in Computer Science, 869:488– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dimopoulos:1994:CMN

[3261] Y. Dimopoulos. Classical methods in nonmonotonic reasoning. Lecture Notes in Computer Science, 869:500– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dix:1994:PER

[3262] J. Dix and M. Mueller. Partial evaluation and relevance for approximations of the stable semantics. Lecture Notes in Computer Science, 869:511– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doherty:1994:CFFb

[3263] P. Doherty and W. Lukaszewicz. Circumscribing features and fluents: a fluent logic for reasoning about action and change. Lecture Notes in Computer Science, 869:521–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghosh:1994:PBN

[3264] S. Ghosh. Paraconsistency and beyond: a new approach to inconsistency handling. Lecture Notes in Computer Science, 869:531–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goubault:1994:BAD

[3265] J. Goubault and J. Posegga. BDDs and automated deduction. Lecture Notes in Computer Science, 869:541–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hsia:1994:PBP

[3266] Y.-T. Hsia. A possibility-based propositional logic of conditionals. Lecture Notes in Computer Science, 869:551– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1994:IPL

[3267] Y.-N. Huang, V. Dahl, and J. Han. Incremental processing of logic database relations. Lecture Notes in Computer Science, 869:561–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jiang:1994:RBA

[3268] Y. J. Jiang and Y. Aramkulchai. On the relationship between assumptionbased framework and autoepistemic logic. Lecture Notes in Computer Science, 869:571–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaub:1994:CQP

[3269] T. Schaub. Computing queries from prioritized default theories. Lecture Notes in Computer Science, 869:584– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sim:1994:BB

[3270] K. M. Sim. Beliefs and bilattices. Lecture Notes in Computer Science, 869: 594–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stachniak:1994:FTD

[3271] Z. Stachniak. Fast termination of the deductive process in resolution proof systems for non-classical logics. Lecture Notes in Computer Science, 869: 604–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greenfield:1994:DPP

[3272] J. S. Greenfield. Distributed programming paradigms with cryptography applications. Lecture Notes in Computer Science, 870:xi + 182, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hibbard:1994:VAD

[3273] W. L. Hibbard, C. R. Dyer, and B. E. Paul. The VIS-AD data model: Integrating metadata and polymorphic display with a scientific programming language. Lecture Notes in Computer Science, 871:37–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kao:1994:ESM

[3274] D. T. Kao, R. D. Bergeron, and T. M.

Sparr. An extended schema model for scientific data. Lecture Notes in Computer Science, 871:69–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ryan:1994:DIV

[3275] K. L. Ryan. Data integration for visualization systems. Lecture Notes in Computer Science, 871:83–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walther:1994:ILS

[3276] S. Walther. Inherent logical structure of computational data: Its role in storage and retrieval strategies to support user queries. Lecture Notes in Computer Science, 871:94–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kochevar:1994:DMD

[3277] P. Kochevar. Database management for data visualization. Lecture Notes in Computer Science, 871:109–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1994:DEI

[3278] J. P. Lee. Data exploration interactions and the ExBase system. Lecture Notes in Computer Science, 871:118– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vasudevan:1994:DRS

[3279] V. Vasudevan. Database requirements for supporting end-user visualizations. Lecture Notes in Computer Science, 871:138–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wierse:1994:SAD

[3280] A. Wierse, U. Lang, and R. Ruehle. A system architecture for data-oriented visualization. Lecture Notes in Computer Science, 871:148–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Woyna:1994:HID

[3281] M. A. Woyna, D. G. Zawada, K. Lee Simunich, and J. H. Christiansen. A hyperspectral image data exploration workbench for environmental science applications. Lecture Notes in Computer Science, 871:160–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyle:1994:DUI

[3282] J. Boyle, J. E. Fothergill, and P. M. D. Gray. Design of a 3D user interface to a database. Lecture Notes in Computer Science, 871:173–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eick:1994:VRD

[3283] S. G. Eick, E. E. Sumner, and G. J. Wills. Visualizing reference databases. Lecture Notes in Computer Science, 871:186–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hemmje:1994:BUI

[3284] M. Hemmje. A 3D based user interface for information retrieval systems. Lecture Notes in Computer Science, 871:194–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keim:1994:UVS

[3285] D. A. Keim and H.-P. Kriegel. Using visualization to support data mining of large existing databases. Lecture Notes in Computer Science, 871: 210–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barzdins:1994:TEIc

[3286] J. Barzdins. Towards efficient inductive synthesis from input/output examples. Lecture Notes in Computer Science, 872:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bibel:1994:DPG

[3287] W. Bibel and M. Thielscher. Deductive plan generation. Lecture Notes in Computer Science, 872:2–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dershowitz:1994:SPI

[3288] N. Dershowitz. From specifications to programs: Induction in the service of synthesis. Lecture Notes in Computer Science, 872:6–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zeugmann:1994:ACA

[3289] T. Zeugmann. Average case analysis of pattern language learning algorithms. Lecture Notes in Computer Science, 872:8–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambainis:1994:ECT

[3290] A. Ambainis and J. Smotrovs. Enumerable classes of total recursive functions: Complexity of inductive inference. Lecture Notes in Computer Science, 872:10–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Apsitis:1994:DSI

[3291] K. Apsitis. Derived sets and inductive inference. Lecture Notes in Computer Science, 872:26–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arnold:1994:TPG

[3292] O. Arnold and K. P. Jantke. Therapy plan generation as program synthesis. Lecture Notes in Computer Science, 872:40–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bai:1994:CLC

[3293] S. Bai. A calculus for logical clustering. Lecture Notes in Computer Science, 872:56–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baliga:1994:LHO

[3294] G. Baliga and J. Case. Learning with higher order additional information. Lecture Notes in Computer Science, 872:64–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brazma:1994:ELR

[3295] A. Brazma and K. Cerans. Efficient learning of regular expressions from good examples. Lecture Notes in Computer Science, 872:76–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freivalds:1994:INM

[3296] R. Freivalds, O. Botuscharov, and R. Wiehagen. Identifying nearly minimal goedel numbers from additional information. Lecture Notes in Computer Science, 872:91–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freivalds:1994:CLF

[3297] R. Freivalds, D. Gobleja, M. Karpinski, and C. H. Smith. Co-learnability and FIN-Identifiability of enumerable classes of total recursive functions. Lecture Notes in Computer Science, 872: 100–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Globig:1994:CBR

[3298] C. Globig and S. Lange. On case-based representability and learnability of languages. Lecture Notes in Computer Science, 872:106–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirata:1994:RGA

[3299] K. Hirata. Rule-generating abduction for recursive Prolog. Lecture Notes in Computer Science, 872:121–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iwatani:1994:FAB

[3300] T. Iwatani, S. Tano, A. Inoue, and W. Okamoto. Fuzzy analogy based reasoning and classification of fuzzy analogies. Lecture Notes in Computer Science, 872:137–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koga:1994:EBR

[3301] Y. Koga, E. Hirowatari, and S. Arikawa. Explanation-based reuse of Prolog programs. Lecture Notes in Computer Science, 872:149–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mofizur:1994:CIR

[3302] C. R. Mofizur and M. Numao. Constructive induction for recursive programs. Lecture Notes in Computer Science, 872:161–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abe:1994:TRT

[3303] N. Abe. Towards realistic theories of learning. Lecture Notes in Computer Science, 872:187–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Richter:1994:UAI

[3304] M. M. Richter. A unified approach to inductive logic and case-based reasoning. Lecture Notes in Computer Science, 872:210–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1994:TDT

[3305] C. H. Smith. Three decades of team learning. Lecture Notes in Computer Science, 872:211–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Auer:1994:LLM

[3306] P. Auer and N. Cesa-Bianchi. On-line learning with malicious noise and the closure algorithm. Lecture Notes in Computer Science, 872:229–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ben-David:1994:LRF

[3307] S. Ben-David and E. Dichterman.

Learnability with restricted focus of attention guarantees noise-tolerance. Lecture Notes in Computer Science, 872:248–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brazma:1994:EAL

[3308] A. Brazma. Efficient algorithm for learning simple regular expressions from noisy examples. Lecture Notes in Computer Science, 872:260–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1994:NLD

[3309] Z. Chen. A note on learning DNF formulas using equivalence and incomplete membership queries. Lecture Notes in Computer Science, 872:272– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferretti:1994:IRL

[3310] C. Ferretti and G. Mauri. Identifying regular languages over partiallycommutative monoids. Lecture Notes in Computer Science, 872:282–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gasarch:1994:CUI

[3311] W. I. Gasarch, M. G. Pleszkoch, and M. Velauthapillai. Classification using information. Lecture Notes in Computer Science, 872:290–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishino:1994:LET

[3312] A. Ishino and A. Yamamoto. Learning from examples with typed equational programming. Lecture Notes in Computer Science, 872:301–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishizaka:1994:FTP

[3313] H. Ishizaka, H. Arimura, and T. Shinohara. Finding tree patterns consistent with positive and negative examples using queries. Lecture Notes in Computer Science, 872:317–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jain:1994:PSP

[3314] S. Jain. Program synthesis in the presence of infinite number of inaccuracies. Lecture Notes in Computer Science, 872:333–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jain:1994:MSL

[3315] S. Jain and A. Sharma. On monotonic strategies for learning r.e. languages. Lecture Notes in Computer Science, 872:349–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapur:1994:LLU

[3316] S. Kapur. Language learning under various types of constraint combinations. Lecture Notes in Computer Science, 872:365–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kimura:1994:SAR

[3317] S. Kimura, A. Togashi, and N. Shiratori. Synthesis algorithm for recursive processes by μ-calculus. Lecture Notes in Computer Science, 872:379– 394, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kinber:1994:MVE

[3318] E. Kinber. Monotonicity versus efficiency for learning languages from texts. Lecture Notes in Computer Science, 872:395–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kobayashi:1994:LCL

[3319] S. Kobayashi and T. Yokomori. Learning concatenations of locally testable languages from positive data. Lecture Notes in Computer Science, 872:407– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lange:1994:LLG

[3320] S. Lange, J. Nessel, and R. Wiehagen. Language learning from good examples. Lecture Notes in Computer Science, 872:423–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lange:1994:MDP

[3321] S. Lange and P. Watson. Machine discovery in the presence of incomplete or ambiguous data. Lecture Notes in Computer Science, 872:438–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lange:1994:SDR

[3322] S. Lange and T. Zeugmann. Set-driven and rearrangement-independent learning of recursive languages. Lecture Notes in Computer Science, 872:453– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsumoto:1994:RPA

[3323] S. Matsumoto and A. Shinohara. Refutably probably approximately correct learning. Lecture Notes in Computer Science, 872:469–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mukouchi:1994:IIA

[3324] Y. Mukouchi. Inductive inference of an approximate concept from positive data. Lecture Notes in Computer Science, 872:484–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakamura:1994:EDF

[3325] A. Nakamura, N. Abe, and J. Takeuchi. Efficient distribution-free population learning of simple concepts. Lecture Notes in Computer Science, 872:500– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okubo:1994:CPM

[3326] Y. Okubo and M. Haraguchi. Constructing predicate mappings for goaldependent abstraction. Lecture Notes in Computer Science, 872:516–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakakibara:1994:LLC

[3327] Y. Sakakibara, K. P. Jantke, and S. Lange. Learning languages by collecting cases and tuning parameters. Lecture Notes in Computer Science, 872:532–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takimoto:1994:MIG

[3328] E. Takimoto, I. Tajika, and A. Maruoka. Mutual information gaining algorithm and its relation to PAC-learning algorithm. Lecture Notes in Computer Science, 872:547–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tanida:1994:IIM

[3329] N. Tanida and T. Yokomori. Inductive inference of monogenic pure contextfree languages. Lecture Notes in Computer Science, 872:560–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Back:1994:ASM

[3330] R. J. R. Back and K. Sere. From action systems to modular systems. Lecture Notes in Computer Science, 873:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dehbonei:1994:FMR

[3331] B. Dehbonei and F. Mejia. Formal methods in the railways signalling industry. Lecture Notes in Computer Science, 873:26–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fitzgerald:1994:FIS

[3332] J. S. Fitzgerald, T. M. Brookes, M. A. Green, and P. G. Larsen. Formal and informal specifications of a secure system component: First results in a comparative study. Lecture Notes in Computer Science, 873:35–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

King:1994:FBR

[3333] T. King. Formalising British Rail’s signalling rules. Lecture Notes in Computer Science, 873:45–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guttman:1994:TAF

[3334] J. D. Guttman and D. M. Johnson. Three applications of formal methods at MITRE. Lecture Notes in Computer Science, 873:55–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cherkaoui:1994:SAS

[3335] O. Cherkaoui, N. Rico, and A. Bernardi. Specification and analysis of a security management system. Lecture Notes in Computer Science, 873:66–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1994:VTL

[3336] U. Martin and M. Thomas. Verification techniques for LOTOS. Lecture Notes in Computer Science, 873:83–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ritchie:1994:EUA

[3337] B. Ritchie, J. Bicarregui, and H. Haughton. Experiences in using the abstract machine notation in a GKS case study. Lecture Notes in Computer Science, 873:93–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bowen:1994:SMM

[3338] J. P. Bowen and M. G. Hinchey. Seven more myths of formal methods: Dispelling industrial prejudices. Lecture Notes in Computer Science, 873:105– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clement:1994:CAD

[3339] T. Clement. Comparing approaches to data reification. Lecture Notes in Computer Science, 873:118–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maung:1994:TFP

[3340] I. Maung, J. R. Howse, and R. J.

Mitchell. Towards a formalization of programming-by-difference. Lecture Notes in Computer Science, 873:134– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paech:1994:NCR

[3341] B. Paech and B. Rumpe. A new concept of refinement used for behaviour modelling with automata. Lecture Notes in Computer Science, 873:154– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elvang-Goransson:1994:EVR

[3342] M. Elvang-Goransson and R. E. Fields. An extended VDM refinement relation. Lecture Notes in Computer Science, 873:175–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lindsay:1994:TVV

[3343] P. A. Lindsay. On transferring VDM verification techniques to Z. Lecture Notes in Computer Science, 873:190– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ledru:1994:PBD

[3344] Y. Ledru. Proof-based development of specifications with KIDS/VDM. Lecture Notes in Computer Science, 873: 214–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

GormLarsen:1994:EUE

[3345] P. Gorm Larsen. Evaluation of underdetermined explicit definitions. Lecture Notes in Computer Science, 873:233– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ross:1994:PEB

[3346] K. J. Ross and P. A. Lindsay. A precise examination of the behaviour of process models. Lecture Notes in Computer Science, 873:251–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duke:1994:TP

[3347] D. J. Duke and M. D. Harrison. A theory of presentations. Lecture Notes in Computer Science, 873:271–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keane:1994:ACF

[3348] J. A. Keane, J. Sa, and B. C. Warboys. Applying a concurrent formal framework to process modelling. Lecture Notes in Computer Science, 873: 291–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cordeiro:1994:MER

[3349] V. A. O. Cordeiro, A. Sampaio, and S. L. Meira. From MooZ to Eiffel — A rigorous approach to system development. Lecture Notes in Computer Science, 873:306–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mens:1994:OFA

[3350] T. Mens, K. Mens, and P. Steyaert.

OPUS: a formal approach to objectorientation. Lecture Notes in Computer Science, 873:326–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Storey:1994:SPV

[3351] A. C. Storey and H. P. Haughton. A strategy for the production of verifiable code using the B method. Lecture Notes in Computer Science, 873: 346–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Evans:1994:SVC

[3352] A. S. Evans. Specifying and verifying concurrent systems using Z. Lecture Notes in Computer Science, 873: 366–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Collinson:1994:CLF

[3353] R. Collinson. A critical look at functional specifications. Lecture Notes in Computer Science, 873:381–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naftalin:1994:ISD

[3354] M. Naftalin. Informal strategies in design by refinement. Lecture Notes in Computer Science, 873:401–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anlauff:1994:ESS

[3355] M. Anlauff, S. Jaehnichen, and M. Simons. An experimental support system for formal mathematical reasoning. Lecture Notes in Computer Science, 873:421–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber:1994:LMD

[3356] M. Weber. Literate mathematical development of a revision management system. Lecture Notes in Computer Science, 873:441–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mosses:1994:ASM

[3357] P. D. Mosses and M. Musicante. An action semantics for ML concurrency primitives. Lecture Notes in Computer Science, 873:461–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mukherjee:1994:SNV

[3358] P. Mukherjee. A semantics for NewSpeak in VDM-SL. Lecture Notes in Computer Science, 873:480– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OHalloran:1994:ESZ

[3359] C. O’Halloran. Evaluation semantics in Z. Lecture Notes in Computer Science, 873:502–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jackson:1994:AMC

[3360] D. Jackson. Abstract model checking of infinite specifications. Lecture Notes in Computer Science, 873:519–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seidel:1994:SRP

[3361] K. Seidel. Specification and refinement of the PI-Bus. Lecture Notes in Computer Science, 873:532–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Erasmy:1994:SRC

[3362] F. Erasmy and E. Sekerinski. Stepwise refinement of control software — A case study using RAISE. Lecture Notes in Computer Science, 873:547–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gortz:1994:SSP

[3363] J. Gortz. Specifying safety and progress properties with RSL. Lecture Notes in Computer Science, 873:567–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hansen:1994:VRI

[3364] K. M. Hansen. Validation of a railway interlocking model. Lecture Notes in Computer Science, 873:582–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simpson:1994:FSA

[3365] A. Simpson. A formal specification of an automatic train protection system. Lecture Notes in Computer Science, 873:602–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fidge:1994:ART

[3366] C. Fidge. Adding real time to formal program development. Lecture Notes in Computer Science, 873:618–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cuellar:1994:CDI

[3367] J. Cuellar, I. Wildgruber, and D. Barnard. Combining the design of industrial systems with effective verification techniques. Lecture Notes in Computer Science, 873:639–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1994:RRP

[3368] F. Wang and A. K. Mok. RTL and refutation by positive cycles. Lecture Notes in Computer Science, 873:659– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fencott:1994:FSW

[3369] P. C. Fencott, A. J. Galloway, M. A.

Lockyer, and S. J. O’Brien. Formalising the semantics of ward/mellor SA/ RT essential models using a process algebra. Lecture Notes in Computer Science, 873:681–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBarros:1994:DRD

[3370] R. S. M. De Barros. Deriving relational database programs from formal specifications. Lecture Notes in Computer Science, 873:703–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeavons:1994:SOC

[3371] P. Jeavons, D. Cohen, and M. Cooper.

A substitution operation for constraints. Lecture Notes in Computer Science, 874:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sabin:1994:CCW

[3372] D. Sabin and E. C. Freuder. Contradicting conventional wisdom in constraint satisfaction. Lecture Notes in Computer Science, 874:10–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jiang:1994:NBM

[3373] Y. Jiang, T. Richards, and B. Richards. Nogood backmarking with min-conflict repair in constraint satisfaction and optimization. Lecture Notes in Computer Science, 874:21–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haroud:1994:GCC

[3374] D. Haroud and B. Faltings. Global consistency for continuous constraints. Lecture Notes in Computer Science, 874:40–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hosobe:1994:LSC

[3375] H. Hosobe, K. Miyashita, S. Takahashi, and S. Matsuoka. Locally simultaneous constraint satisfaction. Lecture Notes in Computer Science, 874:51–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sannella:1994:ADH

[3376] M. Sannella. Analyzing and debugging hierarchies of multi-way local propagation constraints. Lecture Notes in Computer Science, 874:63–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thennarangam:1994:IDC

[3377] S. Thennarangam and G. Singh. Inferring 3-dimensional constraints with DEVI. Lecture Notes in Computer Science, 874:78–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jaffar:1994:BFD

[3378] J. Jaffar, M. J. Maher, P. J. Stuckey, and R. H. C. Yap. Beyond finite domains. Lecture Notes in Computer Science, 874:86–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pesant:1994:QCR

[3379] G. Pesant and M. Boyer. QUAD-CLP(R): Adding the power of quadratic constraints. Lecture Notes in Computer Science, 874:95–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rajasekar:1994:ACL

[3380] A. Rajasekar. Applications in constraint logic programming with strings. Lecture Notes in Computer Science, 874:109–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hermenegildo:1994:SMI

[3381] M. Hermenegildo. Some methodological issues in the design of CIAO, a generic, parallel concurrent constraint logic programming system. Lecture Notes in Computer Science, 874:123– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulte:1994:ESC

[3382] C. Schulte, G. Smolka, and J. Wuertz.

Encapsulated search and constraint programming in oz. Lecture Notes in Computer Science, 874:134–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montanari:1994:TCS

[3383] U. Montanari, F. Rossi, F. Bueno, and M. Garcia de la Banda. Towards a concurrent semantics based analysis of CC and CLP. Lecture Notes in Computer Science, 874:151–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montanari:1994:CPB

[3384] U. Montanari, F. Rossi, and V. Saraswat. CC programs with both inand non-determinism: A concurrent semantics. Lecture Notes in Computer Science, 874:162–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1994:ECT

[3385] A. Gupta, Y. Sagiv, J. D. Ullman, and J. Widom. Efficient and complete tests for database integrity constraint checking. Lecture Notes in Computer Science, 874:173–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Afrati:1994:LVP

[3386] F. Afrati, S. S. Cosmadakis, S. Grumbach, and G. M. Kuper. Linear vs. polynomial constraints in database query languages. Lecture Notes in Computer Science, 874:181–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ross:1994:FAC

[3387] K. A. Ross, D. Srivastava, P. J. Stuckey, and S. Sudarshan. Foundations of aggregation constraints. Lecture Notes in Computer Science, 874: 193–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baudinet:1994:CGD

[3388] M. Baudinet, J. Chomicki, and P. Wolper. Constraint-generating dependencies. Lecture Notes in Computer Science, 874:205–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Srivastava:1994:CO

[3389] D. Srivastava, R. Ramakrishnan, and P. Z. Revesz. Constraint objects. Lecture Notes in Computer Science, 874: 218–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1994:SVC

[3390] Y. Zhang and A. K. Mackworth. Specification and verification of constraintbased dynamic systems. Lecture Notes in Computer Science, 874:229– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ginsberg:1994:GDB

[3391] M. L. Ginsberg and D. A. McAllester.

GSAT and dynamic backtracking. Lecture Notes in Computer Science, 874: 243–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koubarakis:1994:FIC

[3392] M. Koubarakis. Foundations of indefinite constraint databases. Lecture Notes in Computer Science, 874:266– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heintze:1994:SCS

[3393] N. Heintze and J. Jaffar. Set constraints and set-based analysis. Lecture Notes in Computer Science, 874: 281–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paltrinieri:1994:DCS

[3394] M. Paltrinieri. On the design of constraint satisfaction problems. Lecture Notes in Computer Science, 874:299– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pugh:1994:ECB

[3395] W. Pugh and D. Wonnacott. Experiences with constraint-based array dependence analysis. Lecture Notes in Computer Science, 874:312–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aiken:1994:SCR

[3396] A. Aiken. Set constraints: Results, applications and future directions. Lecture Notes in Computer Science, 874: 326–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hooker:1994:LBM

[3397] J. N. Hooker. Logic-based methods for optimization. Lecture Notes in Computer Science, 874:336–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurlander:1994:PSC

[3398] D. Kurlander, J.-F. Puget, and J. Heisserman. Panel session: Commercial applications of constraint programming. Lecture Notes in Computer Science, 874:350–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beth:1994:VTO

[3399] T. Beth, M. Borcherding, and B. Klein.

Valuation of trust in open networks. Lecture Notes in Computer Science, 875:3–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zorkadis:1994:SVP

[3400] V. Zorkadis. Security versus performance requirements in data communication systems. Lecture Notes in Computer Science, 875:19–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roscoe:1994:NIT

[3401] A. W. Roscoe, J. C. P. Woodcock, and L. Wulf. Non-interference through determinism. Lecture Notes in Computer Science, 875:33–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Banatre:1994:CTD

[3402] J.-P. Banatre, C. Bryce, and D. Le Metayer. Compile-time detection of information flow in sequential programs. Lecture Notes in Computer Science, 875:55–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OHalloran:1994:STT

[3403] C. O’Halloran and C. T. Sennett. Security through type analysis. Lecture Notes in Computer Science, 875:75–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyd:1994:DSK

[3404] C. Boyd and W. Mao. Designing secure key exchange protocols. Lecture Notes in Computer Science, 875:93–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hauser:1994:RSP

[3405] R. Hauser, P. Janson, R. Molva, and G. Tsudik. Robust and secure password and key change method. Lecture Notes in Computer Science, 875:107– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jiwa:1994:BBA

[3406] A. Jiwa, J. Seberry, and Y. Zheng. Beacon based authentication. Lecture Notes in Computer Science, 875:125– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hardjono:1994:AMS

[3407] T. Hardjono and J. Seberry. Authentication via multi-service tickets in the Kuperee server. Lecture Notes in Computer Science, 875:143–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1994:OS

[3408] L. Chen. Oblivious signatures. Lecture Notes in Computer Science, 875:

161–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maurer:1994:CSC

[3409] U. M. Maurer and P. E. Schmid. A calculus for secure channel establishment in open networks. Lecture Notes in Computer Science, 875:175–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mao:1994:SAP

[3410] W. Mao and C. Boyd. On strengthening authentication protocols to foil cryptanalysis. Lecture Notes in Computer Science, 875:193–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Camenisch:1994:EEP

[3411] J. L. Camenisch, J. M. Piveteau, and M. A. Stadler. An efficient electronic payment system protecting privacy. Lecture Notes in Computer Science, 875:207–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boly:1994:EPC

[3412] J.-P. Boly, A. Bosselaers, R. Cramer, and R. Michelsen. The ESPRIT project CAFE — high security digital payment systems. Lecture Notes in Computer Science, 875:217–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1994:LCS

[3413] R. J. Anderson. Liability and computer security: Nine principles. Lecture Notes in Computer Science, 875: 231–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DAusbourg:1994:ISD

[3414] B. D’Ausbourg. Implementing secure dependencies over a network by designing a distributed security Sub-System. Lecture Notes in Computer Science, 875:249–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Siron:1994:SMA

[3415] P. Siron and B. D’Ausbourg. A secure medium access control protocol: Security vs performances. Lecture Notes in Computer Science, 875:267–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calas:1994:DFS

[3416] C. Calas. Distributed file system over a multilevel secure architecture — problems and solutions. Lecture Notes in Computer Science, 875:281–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sandhu:1994:EPU

[3417] R. S. Sandhu and S. Ganta. On the expressive power of the unary transformation model. Lecture Notes in Computer Science, 875:301–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dacier:1994:PGE

[3418] M. Dacier and Y. Deswarte. Privilege graph: an extension to the typed access matrix model. Lecture Notes in Computer Science, 875:319–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robinson:1994:CMO

[3419] C. L. Robinson and S. R. Wiseman. A consideration of the modes of operation for secure systems. Lecture Notes in Computer Science, 875:335–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciampichetti:1994:MSG

[3420] A. Ciampichetti, E. Bertino, and L. Mancini. Mark-and-sweep garbage collection in multilevel secure objectoriented database systems. Lecture Notes in Computer Science, 875:359– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boulahia-Cuppens:1994:DMO

[3421] N. Boulahia-Cuppens, F. Cuppens, A. Gabillon, and K. Yazdanian. Decomposition of multilevel objects in an object-oriented database. Lecture Notes in Computer Science, 875:375– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:1994:SOB

[3422] R. K. Thomas and R. S. Sandhu. Supporting object-based high-assurance write-up in multilevel databases for the replicated architecture. Lecture Notes in Computer Science, 875:403– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Motro:1994:ARD

[3423] A. Motro, D. G. Marks, and S. Jajodia.

Aggregation in relational databases:

Controlled disclosure of sensitive information. Lecture Notes in Computer Science, 875:431–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cuppens:1994:IFC

[3424] F. Cuppens and G. Trouessin. Information flow controls vs. interference controls: An integrated approach. Lecture Notes in Computer Science, 875: 447–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uzilevsky:1994:EUI

[3425] G. Uzilevsky. Ergosemiotics of user interface research and design: Foundations, objectives, potential. Lecture Notes in Computer Science, 876:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertelsen:1994:FLD

[3426] O. Bertelsen. Fitts’ Law as a design artefact: A paradigm case of theory in software design. Lecture Notes in Computer Science, 876:11–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rueckert:1994:ESU

[3427] C. Rueckert and S. Klein. Empirical study on the use of a knowledgebased system for selecting standard engineering components. Lecture Notes in Computer Science, 876:19–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauterberg:1994:PES

[3428] M. Rauterberg and E. Styger. Positive effects of sound feedback during the operation of a plant simulator. Lecture Notes in Computer Science, 876:35–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer:1994:RUA

[3429] B. Meyer. Retail user assistant: Evaluation of a user-adapted performance support system. Lecture Notes in Computer Science, 876:45–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tudor:1994:GIS

[3430] L. G. Tudor. Growing an icon set:

User acceptance of abstract and concrete icon styles. Lecture Notes in Computer Science, 876:56–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Modugno:1994:EGF

[3431] F. Modugno and B. A. Myers. Exploring graphical feedback in a demonstrational visual shell. Lecture Notes in Computer Science, 876:64–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conway:1994:AAD

[3432] D. Conway. ARGS: Autogenerating documented command line interfaces. Lecture Notes in Computer Science, 876:77–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hix:1994:IES

[3433] D. Hix and H. R. Hartson. IDEAL: An environment to support usability engineering. Lecture Notes in Computer Science, 876:95–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schreiber:1994:SGU

[3434] S. Schreiber. Specification and generation of user interfaces with the BOSSsystem. Lecture Notes in Computer Science, 876:107–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Donskoy:1994:AOO

[3435] M. V. Donskoy. An architecture for object oriented user interfaces. Lecture Notes in Computer Science, 876:121– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kosbie:1994:EPD

[3436] D. S. Kosbie and B. A. Myers. Extending programming by demonstration with hierarchical event histories. Lecture Notes in Computer Science, 876: 128–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levin:1994:HDU

[3437] M. S. Levin. Hierarchical design of user interfaces. Lecture Notes in Computer Science, 876:140–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kotenko:1994:CRC

[3438] I. V. Kotenko. Conflict resolution in computer-supported cooperative design. Lecture Notes in Computer Science, 876:152–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hurts:1994:CEM

[3439] K. Hurts and P. De Greef. Cognitive ergonomics of multi-agent systems: Observations, principles and research issues. Lecture Notes in Computer Science, 876:164–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellamy:1994:SCL

[3440] R. K. E. Bellamy, E. B. W. Cooper, and R. D. Borovoy. Supporting collaborative learning through the use of electronic conversational props. Lecture Notes in Computer Science, 876: 181–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nitsche-Ruhland:1994:KBA

[3441] D. Nitsche-Ruhland. A knowledgebased authoring system for hypermediabased learning environments. Lecture Notes in Computer Science, 876:192– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brusilovsky:1994:EVE

[3442] P. Brusilovsky. Explanatory visualization in an educational programming environment: Connecting examples with general knowledge. Lecture Notes in Computer Science, 876:202–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chelnokov:1994:CHS

[3443] V. M. Chelnokov and V. L. Zephyrova. Cultivating and harvesting semantic coherence in SMIsC hypertext networks. Lecture Notes in Computer Science, 876:213–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leventhal:1994:HBK

[3444] L. Leventhal, K. Instone, and B. Teasley. Hypertext-based kiosk systems: Seven challenges and an empirical study. Lecture Notes in Computer Science, 876:226–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alford:1994:DFR

[3445] W. R. Alford, A. Granville, and C. Pomerance. On the difficulty of finding reliable witnesses. Lecture Notes in Computer Science, 877:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bosma:1994:DCR

[3446] W. Bosma and P. Stevenhagen. Density computations for real quadratic 2-class groups. Lecture Notes in Computer Science, 877:17–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golliver:1994:LST

[3447] R. A. Golliver, A. K. Lenstra, and K. S. McCurley. Lattice sieving and trial division. Lecture Notes in Computer Science, 877:18–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adleman:1994:SADb

[3448] L. M. Adleman, J. DeMarrais, and M.-D. Huang. A subexponential algorithm for discrete logarithms over the rational subgroup of the Jacobians of large genus hyperelliptic curves over finite fields. Lecture Notes in Computer Science, 877:28–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dodson:1994:CRG

[3449] B. A. Dodson and M. J. Haines. Computing rates of growth of division fields on CM abelian varieties. Lecture Notes in Computer Science, 877:41–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paulus:1994:ACF

[3450] S. Paulus. Algorithms for CM-fields.

Lecture Notes in Computer Science, 877:42–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Couveignes:1994:SAI

[3451] J.-M. Couveignes and F. Morain.

Schoof’s algorithm and isogeny cycles. Lecture Notes in Computer Science, 877:43–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stephens:1994:IPR

[3452] N. Stephens. Integer points on rational elliptic curves. Lecture Notes in Computer Science, 877:59–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lehmann:1994:CNP

[3453] F. Lehmann, M. Maurer, V. Mueller, and V. Shoup. Counting the number of points on elliptic curves over finite fields of characteristic greater than three. Lecture Notes in Computer Science, 877:60–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lipton:1994:SLC

[3454] R. J. Lipton. Straight-line complexity and integer factorization. Lecture Notes in Computer Science, 877:71–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kozen:1994:DAF

[3455] D. Kozen, S. Landau, and R. Zippel. Decomposition of algebraic functions. Lecture Notes in Computer Science, 877:80–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rubinfeld:1994:NMI

[3456] R. Rubinfeld and R. Zippel. A new modular interpolation algorithm for factoring multivariate polynomials. Lecture Notes in Computer Science, 877:93–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adleman:1994:FHS

[3457] L. M. Adleman. The function held sieve. Lecture Notes in Computer Science, 877:108–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elkies:1994:HPC

[3458] N. D. Elkies. Heegner point computations. Lecture Notes in Computer Science, 877:122–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cremona:1994:CDM

[3459] J. E. Cremona. Computing the degree of a modular parametrization. Lecture Notes in Computer Science, 877:134– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McConnell:1994:GRC

[3460] M. McConnell. Galois representations from the cohomology of SL(3,Z). Lecture Notes in Computer Science, 877: 143–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daude:1994:AGA

[3461] H. Daude, P. Flajolet, and B. Vallee.

An analysis of the Gaussian algorithm for lattice reduction. Lecture Notes in Computer Science, 877:144–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaib:1994:FVG

[3462] M. Kaib. A fast variant of the Gaussian reduction algorithm. Lecture Notes in Computer Science, 877:159–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buchmann:1994:RLB

[3463] J. Buchmann. Reducing lattice bases by means of approximations. Lecture Notes in Computer Science, 877:160– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shallit:1994:ALS

[3464] J. Shallit and J. Sorenson. Analysis of a left-shift binary GCD algorithm. Lecture Notes in Computer Science, 877:169–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Majewski:1994:CGC

[3465] B. S. Majewski and G. Havas. The complexity of greatest common divisor computations. Lecture Notes in Computer Science, 877:184–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanderPoorten:1994:EFU

[3466] A. J. Van der Poorten. Explicit formulas for units in certain quadratic number fields. Lecture Notes in Computer Science, 877:194–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Evdokimov:1994:FPF

[3467] S. Evdokimov. Factorization of polynomials over finite fields in subexponential time under GRH. Lecture Notes in Computer Science, 877:209–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gao:1994:OON

[3468] S. Gao and S. A. Vanstone. On orders of optimal normal basis generators. Lecture Notes in Computer Science, 877:220–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Volcheck:1994:CJP

[3469] E. J. Volcheck. Computing in the Jacobian of a plane algebraic curve. Lecture Notes in Computer Science, 877: 221–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thiel:1994:UAG

[3470] C. Thiel. Under the assumption of the generalized Riemann Hypothesis verifying the class number belongs to NPCo-NP. Lecture Notes in Computer Science, 877:234–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hajir:1994:CCN

[3471] F. Hajir. Calculating the class number of certain Hilbert class fields. Lecture Notes in Computer Science, 877: 248–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adleman:1994:ECC

[3472] L. M. Adleman, M.-D. Huang, and K. Kompella. Efficient checking of computations in number theory. Lecture Notes in Computer Science, 877:249– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lay:1994:CEC

[3473] G.-J. Lay and H. G. Zimmer. Constructing elliptic curves with given group order over large finite fields. Lecture Notes in Computer Science, 877: 250–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deleglise:1994:CXM

[3474] M. Deleglise and J. Rivat. Computing (x), M(x) and psi(x). Lecture Notes in Computer Science, 877:264–??, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shparlinski:1994:SAF

[3475] I. E. Shparlinski. On some applications of finitely generated semigroups. Lecture Notes in Computer Science, 877:265–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pritchard:1994:IIP

[3476] P. Pritchard. Improved incremental prime number sieves. Lecture Notes in Computer Science, 877:280–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shor:1994:PTA

[3477] P. W. Shor. Polynomial time algorithms for discrete logarithms and factoring on a quantum computer. Lecture Notes in Computer Science, 877: 289–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ji:1994:DMC

[3478] G. Ji and H. Lu. On dispersion and Markov constants. Lecture Notes in Computer Science, 877:290–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adleman:1994:OPN

[3479] L. M. Adleman and K. S. McCurley. Open problems in number theoretic complexity, II. Lecture Notes in Computer Science, 877:291–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aalund:1994:CFD

[3480] A. Aalund, P. Loetstedt, and R. Ryden. Computational fluid dynamics on workstation clusters in industrial environments. Lecture Notes in Computer Science, 879:1–10, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersen:1994:PIA

[3481] B. S. Andersen, P. Kaae, C. Keable, and W. Owczarz. PVM implementations of advection-chemistry modules of air pollution models. Lecture Notes in Computer Science, 879:11–16, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antonuccio-Delogu:1994:PTB

[3482] V. Antonuccio-Delogu and U. Becciani.

A parallel tree N-body code for heterogeneous clusters. Lecture Notes in Computer Science, 879:17–32, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Badia:1994:PCE

[3483] J. M. Badia and A. M. Vidal. Parallel computation of the eigenstructure of Toeplitz-plus-Hankel matrices on multicomputers. Lecture Notes in Computer Science, 879:33–40, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balle:1994:SVD

[3484] S. M. Balle and P. M. Pedersen. Singular value decomposition on the Connection Machine CM-5/CM-5E. Lecture Notes in Computer Science, 879:41–50, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barker:1994:CES

[3485] V. A. Barker and C. Yingqun. Computing eigenvalues of sparse matrices on the Connection Machine. Lecture Notes in Computer Science, 879:51–58, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bendtsen:1994:PEE

[3486] C. Bendtsen. Preliminary experiences with extrapolation methods for parallel solution of differential algebraic equations. Lecture Notes in Computer Science, 879:59–66, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berkhin:1994:TAF

[3487] P. Berkhin and J. Brown. A transform approach to fast matrix multiplication. Lecture Notes in Computer Science, 879:67–79, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Billing:1994:QMT

[3488] G. D. Billing. Quantum mechanical treatment of reactive scattering problems. Lecture Notes in Computer Science, 879:80–88, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bode:1994:VCD

[3489] A. Bode, N. Eicker, T. Lippert, and K. Schilling. Visualizing the critical dynamics of monopole current clusters — global/local programming on the Connection Machine CM-5. Lecture Notes in Computer Science, 879:89–99, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boryczko:1994:LBP

[3490] K. Boryczko, J. Kitowski, and J. Moscinski. Load-balancing procedure for distributed short-range molecular dynamics. Lecture Notes in Computer Science, 879:100–109, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1994:CMP

[3491] J. Brown, P. C. Hansen, J. Wasniewski, and Z. Zlatev. Comparison of massively parallel SIMD computers using air pollution models. Lecture Notes in Computer Science, 879:110–126, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bubak:1994:PDD

[3492] M. Bubak, J. Moscinski, M. Pogoda, and W. Zdechlikiewicz. Parallel distributed 2-D short-range molecular dynamics on networked workstations. Lecture Notes in Computer Science, 879:127–135, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bubak:1994:IPL

[3493] M. Bubak, J. Moscinski, and R. Slota.

Implementation of parallel lattice gas program on workstations under PVM. Lecture Notes in Computer Science, 879:136–146, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choi:1994:CNS

[3494] J. Choi, J. J. Dongarra, R. Pozo, and D. W. Walker. Constructing numerical software libraries for highperformance computing environments.

Lecture Notes in Computer Science, 879:147–168, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dayde:1994:CE

[3495] M. J. Dayde and I. S. Duff. The CER-FACS experience. Lecture Notes in Computer Science, 879:169–176, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dohmen:1994:TAP

[3496] R. Dohmen. Towards an automatic portation of programs from nCUBE2 to KSR1. Lecture Notes in Computer Science, 879:177–189, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elinov:1994:SDP

[3497] V. D. Elinov, O. G. Monakhov, and I. I. Ushatinsky. Seismic data processing: Parallel and conservative computations on the multitransputer system. Lecture Notes in Computer Science, 879:190–193, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fet:1994:HPC

[3498] Y. I. Fet and A. P. Vazhenin. Heterogeneous processing: A combined approach. Lecture Notes in Computer Science, 879:194–206, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Filippone:1994:ENS

[3499] S. Filippone and M. L. Sales. Experiences in numerical software on IBM distributed memory architectures. Lecture Notes in Computer Science, 879:207– 218, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freeman:1994:SMM

[3500] T. L. Freeman and J. M. Bull. Shared memory and message passing implementations of parallel algorithms for numerical integration. Lecture Notes in Computer Science, 879:219–228, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gajecki:1994:NAT

[3501] M. Gajecki and J. Moscinski. A new algorithm for the traveling salesman problem on networked workstations. Lecture Notes in Computer Science, 879:229–235, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geist:1994:CCW

[3502] G. A. Geist. Cluster computing: The wave of the future? Lecture Notes in Computer Science, 879:236–246, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grosz:1994:NFE

[3503] L. Grosz, C. Roll, and W. Schoenauer.

Nonlinear finite element problems on parallel computers. Lecture Notes in Computer Science, 879:247–261, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hakula:1994:FEM

[3504] H. Hakula, J. Malinen, P. Kallberg, and P. Valve. The finite element method applied to the exterior Helmholtz problem on the IBM SP-1. Lecture Notes in Computer Science, 879:262–269, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hammarling:1994:CPL

[3505] S. Hammarling. The challenge of portable libraries for high performance machines. Lecture Notes in Computer Science, 879:270–286, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hansen:1994:PIR

[3506] P. C. Hansen. Parallel issues of regularization problems. Lecture Notes in Computer Science, 879:287–295, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hansen:1994:TEP

[3507] P. C. Hansen, T. Ostromsky, and Z. Zlatev. Two enhancements in a partitioned sparse code. Lecture Notes in Computer Science, 879:296–303, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harrod:1994:NAC

[3508] W. J. Harrod and M. Sidani. Numerical algorithms for the Cray T3D. Lecture Notes in Computer Science, 879:304– 311, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hunding:1994:NDB

[3509] A. Hunding. Nonlinear dynamics in biochemical control systems. Lecture Notes in Computer Science, 879:312– 317, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johnsson:1994:SSL

[3510] S. L. Johnsson and K. K. Mathur. Scientific software libraries for scalable architectures. Lecture Notes in Computer Science, 879:318–332, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kienhoefer:1994:ESP

[3511] J. G. Kienhoefer. The EFECT system for parallel programming. Lecture Notes in Computer Science, 879:333– 343, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsen:1994:OOD

[3512] T. Larsen, W. Joosen, and J. Perram. The object-oriented development of a parallel application in polymer dynamics. Lecture Notes in Computer Science, 879:344–348, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leszczynski:1994:SDA

[3513] H. Leszczynski and J. Pieniazek. On some difference analogues of PDEs with a delay. Lecture Notes in Computer Science, 879:349–357, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Madsen:1994:PIM

[3514] K. Madsen and O. Toft. Parallel interval methods for perturbed linear systems. Lecture Notes in Computer Science, 879:358–361, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McBryan:1994:LFH

[3515] O. A. McBryan. Limiting factors in high performance computing. Lecture Notes in Computer Science, 879:362– 373, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michielse:1994:PCE

[3516] P. Michielse. Programming the convex exemplar series SPP system. Lecture Notes in Computer Science, 879:374– 382, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Museth:1994:FQM

[3517] K. Museth, G. D. Billing, and J. Linderberg. The full quantum mechanical three body problem of H0+2. Lecture Notes in Computer Science, 879:383– 391, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielsen:1994:MDE

[3518] O. H. Nielsen. Molecular dynamics: experiences with Connection Machines CM-200 and CM-5. Lecture Notes in Computer Science, 879:392–396, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nordling:1994:SOD

[3519] P. Nordling and P. Fritzson. Solving ordinary differential equations on parallel computers — applied to dynamic rolling bearings simulation. Lecture Notes in Computer Science, 879:397– 415, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paprzycki:1994:SLR

[3520] M. Paprzycki and P. Stpiczynski. Solving linear recurrence systems on a Cray Y-MP. Lecture Notes in Computer Science, 879:416–424, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pedersen:1994:STE

[3521] P. M. Pedersen and S. M. Balle. Selected techniques for efficient parallel implementation of dense linear algebra algorithms on the Connection Machine CM-5/CM-5E. Lecture Notes in Computer Science, 879:425–433, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skaalin:1994:PHM

[3522] R. Skaalin, D. Bjoerge, N. Gustafsson, and I. Lie. Parallelization of HIRLAM on MIMD platforms. Lecture Notes in Computer Science, 879:434–457, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soerensen:1994:SFD

[3523] J. N. Soerensen, M. O. L. Hansen, and E. Jensen. Simulation of fluid dynamical flow problems. Lecture Notes in Computer Science, 879:458–468, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Szularz:1994:PAP

[3524] M. Szularz, J. Weston, K. Murphy, and M. Clint. Parallel algorithms for the partial eigensolution of large sparse matrices on novel architecture computers. Lecture Notes in Computer Science, 879:469–482, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tengwall:1994:SAP

[3525] C. G. Tengwall. SP2 architecture and performance. Lecture Notes in Computer Science, 879:483–492, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomsen:1994:RTS

[3526] P. G. Thomsen. Real time simulation in a cluster computing environment. Lecture Notes in Computer Science, 879:493–497, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toxvaerd:1994:PCM

[3527] S. Toxvaerd. Parallel computations in molecular dynamics. Lecture Notes in Computer Science, 879:498–504, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vazhenin:1994:EHA

[3528] A. P. Vazhenin. Efficient high-accuracy computations in massively parallel systems. Lecture Notes in Computer Science, 879:505–519, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yalamov:1994:SPB

[3529] P. Y. Yalamov. Stability of parallel bidiagonal solvers. Lecture Notes in Computer Science, 879:520–533, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zemla:1994:WTC

[3530] A. Zemla. Wavelet transforms computing on PVM. Lecture Notes in Computer Science, 879:534–546, 1994.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zlatev:1994:PPS

[3531] Z. Zlatev and J. Wasniewski. PARAS-PAR: Parallel solvers for sparse linear algebraic systems. Lecture Notes in Computer Science, 879:547–556, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zrimsek:1994:FMF

[3532] A. Zrimsek and T. Smukavec. Fair metacomputer — FMC. Lecture Notes in Computer Science, 879:557–566, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kozen:1994:ERS

[3533] D. Kozen. Efficient resolution of singularities of plane curves. Lecture Notes in Computer Science, 880:1–11, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Couveignes:1994:ICG

[3534] J.-M. Couveignes, J. F. Diaz-Frias, M. De Rougemont, and M. Santha. On the interactive complexity of graph reliability. Lecture Notes in Computer Science, 880:12–23, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dietzfelbinger:1994:MUL

[3535] M. Dietzfelbinger and M. Huehne.

Matching upper and lower bounds for simulations of several tapes on one multidimensional tape. Lecture Notes in Computer Science, 880:24–35, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caussinus:1994:CCQ

[3536] H. Caussinus and F. Lemieux. The complexity of computing over quasigroups. Lecture Notes in Computer Science, 880:36–47, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mahajan:1994:NCC

[3537] M. Mahajan and V. Vinay. Noncommutative computation, depth reduction and skew circuits. Lecture Notes in Computer Science, 880:48–59, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coquand:1994:IDT

[3538] T. Coquand and P. Dybjer. Inductive definitions and type theory an introduction (preliminary version). Lecture Notes in Computer Science, 880:60–76, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broy:1994:IVF

[3539] M. Broy, U. Hinkel, T. Nipkow, and C. Prehofer. Interpreter verification for a functional language. Lecture Notes in Computer Science, 880:77–88, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lakshmanan:1994:EFL

[3540] V. S. Lakshmanan. An epistemic foundation for logic programming with uncertainty. Lecture Notes in Computer Science, 880:89–100, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsuiki:1994:TCM

[3541] H. Tsuiki. On typed calculi with a merge operator. Lecture Notes in Computer Science, 880:101–112, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frigioni:1994:IAS

[3542] D. Frigioni, A. Marchetti-Spaccamela, and U. Nanni. Incremental algorithms for the single-source shortest path problem. Lecture Notes in Computer Science, 880:113–124, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

RaoArikati:1994:ARD

[3543] S. Rao Arikati and A. Maheshwari. An O(n) algorithm for realizing degree sequences. Lecture Notes in Computer Science, 880:125–136, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Subramanian:1994:CSR

[3544] C. R. Subramanian and C. E. Veni Madhavan. Coloring semi-random graphs in polynomial expected time. Lecture Notes in Computer Science, 880:137–148, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:1994:FSS

[3545] W. Thomas. Finite-state strategies in regular infinite games. Lecture Notes in Computer Science, 880:149–158, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nandy:1994:LLE

[3546] S. C. Nandy, A. Sinha, and B. B. Bhattacharya. Location of the largest empty rectangle among arbitrary obstacles. Lecture Notes in Computer Science, 880:159–170, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dahlhaus:1994:EPL

[3547] E. Dahlhaus. Efficient parallel and linear time sequential split decomposition. Lecture Notes in Computer Science, 880:171–180, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biswas:1994:ACV

[3548] S. Biswas, D. C. Prasad, and S. P. Pal. Algorithms for convex visibility problems. Lecture Notes in Computer Science, 880:181–192, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sen:1994:LBP

[3549] S. Sen. Lower bounds for parallel algebraic decision trees, complexity of convex hulls and related problems. Lecture Notes in Computer Science, 880:193–204, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amadio:1994:LF

[3550] R. M. Amadio and S. Prasad. Localities and failures. Lecture Notes in Computer Science, 880:205–216, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Natarajan:1994:PAP

[3551] V. Natarajan, I. Christoff, L. Christoff, and R. Cleaveland. Priority and abstraction in process algebra. Lecture Notes in Computer Science, 880:217– 230, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

NarayanKumar:1994:CPO

[3552] K. Narayan Kumar and P. K. Pandya.

On the computational power of operators in ICSP with fairness. Lecture Notes in Computer Science, 880:231– 242, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yi:1994:DTL

[3553] W. Yi and B. Jonsson. Decidability of timed language-inclusion for networks of real-time communicating sequential processes. Lecture Notes in Computer Science, 880:243–255, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fortnow:1994:MFT

[3554] L. Fortnow. My favorite ten complexity theorems of the past decade. Lecture Notes in Computer Science, 880:256– 275, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaji:1994:SUP

[3555] Y. Kaji, T. Fujiwara, and T. Kasami.

Solving a unification problem under constrained substitutions using tree automata. Lecture Notes in Computer Science, 880:276–287, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramesh:1994:ADE

[3556] R. Ramesh, I. V. Ramakrishnan, and R. C. Sekar. Automata-driven efficient subterm unification. Lecture Notes in Computer Science, 880:288–299, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raghavan:1994:RAA

[3557] P. Raghavan. Randomized approximation algorithms in combinatorial optimization. Lecture Notes in Computer Science, 880:300–317, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arora:1994:LBG

[3558] V. Arora, V. Santosh, H. Saran, and V. V. Vazirani. A limited-backtrack greedy schema for approximation algorithms. Lecture Notes in Computer Science, 880:318–329, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crescenzi:1994:ASP

[3559] P. Crescenzi and L. Trevisan. On approximation scheme preserving reducibility and its applications. Lecture Notes in Computer Science, 880:330– 341, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hunt:1994:ASU

[3560] H. B. Hunt, M. V. Marathe, V. Radhakrishnan, and S. S. Ravi. Approximation schemes using L-reductions. Lecture Notes in Computer Science, 880:342–353, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Subramanian:1994:ES

[3561] A. Subramanian. An explanation of splaying. Lecture Notes in Computer Science, 880:354–365, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desel:1994:PNR

[3562] J. Desel and M.-D. Radola. Proving non-reachability by modulo-place-invariants. Lecture Notes in Computer Science, 880:366–377, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knapp:1994:SCU

[3563] E. Knapp. Soundness and completeness of UNITY logic. Lecture Notes in Computer Science, 880:378–389, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Savicky:1994:EAT

[3564] P. Savicky and I. Wegener. Efficient algorithms for the transformation between different types of binary decision diagrams. Lecture Notes in Computer Science, 880:390–401, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Siegel:1994:ELS

[3565] M. Siegel and F. A. Stomp. Extending the limits of sequentially phased reasoning. Lecture Notes in Computer Science, 880:402–413, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Estivill-Castro:1994:FFE

[3566] V. Estivill-Castro and D. Wood. Foundations for faster external sorting. Lecture Notes in Computer Science, 880: 414–425, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hooker:1994:BRS

[3567] J. N. Hooker and V. Vinay. Branching rules for satisfiability. Lecture Notes in Computer Science, 880:426–437, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapur:1994:ULA

[3568] D. Kapur and M. Subramaniam. Using linear arithmetic procedure for generating induction schemes. Lecture Notes in Computer Science, 880:438–449, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kervella:1994:TCM

[3569] B. Kervella, V. Gay, and E. Horlait. Towards a complete multimedia mail: Use of MHEG in standard messaging systems. Lecture Notes in Computer Science, 882:1–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thimm:1994:MBT

[3570] H. Thimm, K. Roehr, and T. C. Rakow. A mail-based teleservice architecture for archiving and retrieving dynamically composable multimedia documents. Lecture Notes in Computer Science, 882:14–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blum:1994:GSS

[3571] C. Blum and L. Neumann. The global store server — A multimedia teleservice component. Lecture Notes in Computer Science, 882:35–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leopold:1994:BNS

[3572] H. Leopold, K. Frimpong-Ansah, and N. Singer. From broadband network services to a distributed multimedia support-environment. Lecture Notes in Computer Science, 882:47–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shenker:1994:MSE

[3573] S. Shenker, A. Weinrib, and E. Schooler.

Managing shared ephemeral teleconferencing state: Policy and mechanism. Lecture Notes in Computer Science, 882:69–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loge:1994:CCS

[3574] C. Loge, V. Gay, and E. Horlait. Computational components for synchronous cooperation on multimedia information. Lecture Notes in Computer Science, 882:89–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lazar:1994:BAM

[3575] A. A. Lazar, S. K. Bhonsle, and K. S.

Lim. A binding architecture for multimedia networks. Lecture Notes in Computer Science, 882:103–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fedaoui:1994:IEE

[3576] L. Fedaoui, A. Seneviratne, and E. Horlait. Implementation of an end-to-end quality of service management scheme. Lecture Notes in Computer Science, 882:124–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bowman:1994:FDT

[3577] H. Bowman, L. Blair, G. S. Blair, and A. G. Chetwynd. A formal description technique supporting expression of quality of service and media synchronisation. Lecture Notes in Computer Science, 882:145–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yen:1994:SMM

[3578] W. Yen and I. F. Akyildiz. On the synchronization mechanisms for multimedia integrated services networks. Lecture Notes in Computer Science, 882: 168–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Szyperski:1994:ESM

[3579] C. Szyperski and G. Ventre. Efficient support for multiparty communication. Lecture Notes in Computer Science, 882:185–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mathy:1994:QNM

[3580] L. Mathy and O. Bonaventure. QoS negotiation for multicast communications. Lecture Notes in Computer Science, 882:199–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carle:1994:SHP

[3581] G. Carle, J. Schiller, and C. Schmidt. Support for high-performance multipoint multimedia services. Lecture Notes in Computer Science, 882:219– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schatzmayr:1994:PSD

[3582] R. Schatzmayr and R. Popescu-Zeletin.

Providing support for data transfer in a new networking environment. Lecture Notes in Computer Science, 882: 241–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakatani:1994:CAV

[3583] T. Sakatani. Congestion avoidance for video over IP networks. Lecture Notes in Computer Science, 882:256– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wittig:1994:NLS

[3584] H. Wittig, J. Winckler, and J. Sandvoss. Network layer scaling: Congestion control in multimedia communication with heterogeneous networks and receivers. Lecture Notes in Computer Science, 882:274–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamdi:1994:RRV

[3585] M. Hamdi, P. Rolin, Y. Duboc, and M. Ferry. Resource requirements for VBR MPEG traffic in interactive applications. Lecture Notes in Computer Science, 882:294–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andrade:1994:TMA

[3586] T. Andrade and A. P. Alves. Transmission of MPEG2 applications over ATM networks. Lecture Notes in Computer Science, 882:310–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clark:1994:PMC

[3587] W. J. Clark. Protocols for multimedia conferencing — an introduction to the ITU-T T.120 series. Lecture Notes in Computer Science, 882:322–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dermler:1994:PMT

[3588] G. Dermler, T. Gutekunst, E. Ostrowski, and N. Pires. A platform for multimedia telecooperation bridging endsystem heterogeneity. Lecture Notes in Computer Science, 882:334– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pronios:1994:SMH

[3589] N. B. Pronios and T. Bozios. A scheme for multimedia and hypermedia synchronisation. Lecture Notes in Computer Science, 882:340–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Riexinger:1994:IEA

[3590] D. Riexinger and K. Werner. Integration of existing applications into a conference system. Lecture Notes in Computer Science, 882:346–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rozek:1994:CMCb

[3591] A. Rozek and P. Christ. The CIO multimedia communication platform. Lecture Notes in Computer Science, 882: 356–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blanco:1994:BMC

[3592] M. A. Blanco, R. Montero, F. Almerico, and G. Venuti. Broadband multimedia and collaboration tools. idea project. Lecture Notes in Computer Science, 882:368–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hunt:1994:FVM

[3593] Warren A. Hunt. FM8501, a verified microprocessor, volume 795 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 0-387-57960-5 (New York), 3-540-57960-5 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 333 pp. LCCN QA73.8.F6 H86 1994.

Ishida:1994:PDM

[3594] Toru Ishida. Parallel, distributed, and multiagent production systems, volume 878 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CO-DEN LNCSD9. ISBN 3-540-58698-9 (Berlin), 0-387-58698-9 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). xvi + 166 pp. LCCN QA7.58 .I84 1994. DM46.00.